A Complete Bibliography of Publications in Computer Networks (Amsterdam, Netherlands: 2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

30 September 2025 Version 1.56

Title word cross-reference

0 [2569]. 1 [1690]. 1 + 1 [237, 1096]. 2 [1558, 1592, 1971]. 3 [600, 937, 1147, 1282, 1340, 1663, 1979, 2205, 2642]. 5 [3, 14, 15, 19, 32, 78, 108, 114, 131, 140, 149, 160, 170, 174, 186, 211–213, 219, 231]. ^{2} [1629, 2578]. ^{3} [1323, 2223]. _{16} [2019]. _{2} [2060]. _{K} [340]. EON [491]. H [2883]. K [723, 1068, 1640, 2064, 2728]. κ [1920]. m [1858]. M_{K} [340]. H^{2} [469]. μ [1877]. n [2994]. p [2214]. P^{2} [599]. Q [1789, 1804, 2187, 2619, 2622]. z [2883].

* [760]. *-flow [760].

-ABC [2214]. -anonymity [1068, 2064]. -based [412, 880, 1903, 2372]. -Coverage [1920]. -D [1282, 1340, 1690]. -day [2569]. -dimensional [1592]. -gram [2994]. -learning [1804]. -learning-based [2187, 2619]. -learning-enabled [1789]. -reconstruction [1903]. -SHARP [599]. -UAV [2578]. -V2X [1903, 2372]. -view [2728].

/1/M [340]. /Hyper [340].

1 [2975]. 100 [653]. 107082 [357]. 107186 [271]. 107699 [527]. 109133 [1571]. 109529 [1764]. 109750 [1759]. 11 [49, 480, 952]. 13 [2362]. 14 [88, 523, 1000]. 15 [26, 449, 929]. 169 [357]. 16th [3012]. 173 [271]. 185 [527]. 19 [185, 217, 386, 636, 1092, 1369, 1496].

2 [3066]. 2.0 [287, 2039, 2089]. 20 [221, 688, 1137]. 2019 [24]. 2020 [26, 49, 71, 88, 106, 126, 147, 167, 185, 201, 221, 244, 270, 296, 324, 361, 397, 422]. 2021 [449, 480, 501, 523, 547, 570, 592, 614, 636, 665, 688, 710, 741, 777, 813, 848, 879, 912]. 2022 [929, 952, 972, 1000, 1024, 1044, 1058, 1092, 1110, 1137, 1169, 1204, 1221, 1262, 1305, 1343, 1380]. 2023 [1416, 1441, 1465, 1482, 1520, 1550, 1565, 1588, 1614, 1635, 1650, 1674, 1694, 1726, 1766, 1808, 1843, 1905]. 2024 [1926, 1944, 1967, 2002, 2042, 2078, 2105, 2140, 2157, 2194, 2221, 2241, 2301, 2347, 2380, 2414, 2512, 2566]. 2025 [2567, 2585, 2660, 2703, 2742, 2779, 2812, 2849, 2880, 2912, 2927, 2948, 2966, 3033, 3121, 3190]. 214 [1571]. 22 [126, 167, 570, 614, 1044]. 24 [324, 422, 813, 912, 1262, 1380]. 256 [768]. 26 [71, 501, 972]. 26th [1195]. 2IoT [1478].

3 [1712, 3167]. 3.0 [2858, 2962]. 360-degree [598, 2317]. 3C [2149]. 3D [3001]. 3GPP [64, 76, 2116, 2574]. 3GPP-compliant [2574]. 3MRP [1151]. 3P [333]. 3R [1872].

4 [244, 270, 710, 741, 1169, 1204, 2024]. 4.0 [1896]. 40/100 [1720]. 40802.11 [132]. 4G [482, 675, 797, 884, 970, 1575, 1866, 2933]. 4G/5G [482, 675, 884, 970, 2933].

5 [201, 665, 698, 1071, 1110]. 5.0 [1078, 2017]. 5G [175, 255, 282, 297, 326, 334, 339, 357, 391, 392, 410, 415, 460, 482, 490, 525, 530, 564, 580, 581, 628, 652, 667, 692, 707, 744, 754, 775, 803, 808, 828, 898, 944, 983, 1030, 1034, 1048, 1065, 1082, 1089, 1101, 1156, 1198, 1214, 1215, 1219, 1245, 1291, 1331, 1370, 1438, 1444, 1450, 1548, 1583, 1590, 1750, 1755, 1816, 1846, 1857, 1869, 1889, 1896, 1961, 1965, 1985, 2005, 2081, 2116, 2184, 2186, 2242, 2250, 2259, 2279, 2349, 2462, 2562, 2600, 2671, 2678, 2707, 2739, 2830, 2844, 2854, 2871, 2895, 3001, 3038, 3040, 3098, 3155, 3176]. 5G [234, 251, 276, 283, 310, 318, 367, 377, 379, 424, 453, 477, 481, 529, 562, 568, 668, 675, 729, 785, 794, 827, 833, 859, 863, 884, 893, 907, 908, 945, 970, 1028, 1130, 1259, 1297, 1298, 1312, 1334, 1404, 1425, 1445, 1446, 1449, 1582, 1680, 1709, 1772, 1818, 1850, 1864, 1866, 1939, 1954, 1980, 2051, 2086, 2087, 2126, 2170, 2173, 2216, 2218, 2230, 2233, 2276, 2285, 2290, 2365, 2389, 2428, 2483, 2544, 2549, 2592, 2613, 2657, 2690, 2696, 2715, 2724, 2756, 2771, 2908, 2933, 2953, 3025, 3029, 3093, 3151, 3186]. 5G-air-simulator [234]. 5G-AKA [794, 1334, 3040]. 5G-and-Beyond [334, 1582]. 5G-based [1312, 2216]. 5G-Edge [2276]. 5G-Enabled [175, 453, 707, 1156]. 5G-Flow [808]. 5G-GRAKA [3038]. 5G-MEC [1954]. 5G-NB-IoT [827]. 5G-RAN [1219]. 5G-satellite [1198]. 5G-VIOS [1980]. 5G/6G [1101, 1298, 1425, 1444, 2230, 2349, 2365, 2657, 2908]. 5G/6G-Enabled [2349]. 5G/B5G [367, 1869, 2389]. 5GMEC [1312]. 5GMEC-DP [1312]. 5GUK [251].

6 [334, 1425]. 60 [474, 716]. 64-antenna [334]. 6G [399, 575, 1101, 1147, 1177, 1187, 1286, 1298, 1408, 1425, 1444, 1544, 1571, 1591, 1600, 1607, 1705, 1882, 1923, 2027, 2035, 2102, 2112, 2213, 2230, 2261, 2365, 2370, 2384, 2418, 2426, 2457, 2467, 2506, 2526, 2547, 2580, 2653, 2657, 2902, 2908, 2934, 2960, 2989, 3014, 3028, 3050, 3052, 3086, 3095, 3104, 3115, 3117, 3166, 3176]. 6G-connected [2526]. 6G-enabled [2261, 2349, 3115]. 6G-IoT [2102]. 6G-Tactile [1923]. 6Graph [943]. 6LoWPAN [2251, 2534]. 6LoWPAN-ND [2251]. ‘6LoWPAN-RPL [2534]. 6Search [1797]. 6TiSCH [305, 2436].

7 [106, 547, 1024, 1940, 3147].

8 [147, 592, 1058, 3147]. 802.11 [1083]. 802.11ac [327, 513, 763]. 802.11ad [1636]. 802.11ah [199, 1053, 2003]. 802.11ax [376]. 802.11be [2695]. 802.11p [364]. 802.15.4a [902]. 802.15.4e [1653]. 802.15.4e-based [1653]. 802.1Qbv [2765]. 802.1X [2089]. 8219 [2637].

9 [296, 361, 397, 777, 848, 879, 1221, 1305, 1343].

A-DECS [650]. AAE [1810]. AAE-DSVDD [1810]. AAPC [240]. AAR [3023]. AAR-Log [3023]. ABAC [1140]. ABC [2214]. ABE [612, 2443]. ability [2868]. abnormality [896]. above [653]. abstraction [429, 1603]. abundance [774]. ABUV [2630]. accelerated [401, 1082, 1344, 2074]. Accelerating [1406, 1449, 1908, 1999, 2355, 2553, 2852, 2972, 3036, 3061, 3170]. acceleration [1186, 1501, 2308, 2767, 2920]. Accelerator [240, 1717, 1845]. Accelerator-aware [240]. acceptance [699, 2170]. Access [34, 56, 64, 145, 172, 173, 205, 216, 236, 267, 297, 302, 316, 331, 339, 365, 385, 388, 411, 415, 416, 455, 481, 524, 542, 558, 604, 632, 698, 733, 744, 770, 820, 927, 933, 958, 959, 975, 1010, 1016, 1032, 1034, 1082, 1130, 1133, 1142, 1152, 1198, 1259, 1267, 1333, 1347, 1353, 1361, 1447, 1479, 1511, 1512, 1545, 1582, 1591, 1593, 1601, 1625, 1658, 1667, 1672, 1746, 1749, 1802, 1855, 1951, 1955, 1968, 2003, 2035, 2051, 2054, 2116, 2144, 2148, 2149, 2153, 2250, 2278, 2322, 2353, 2439, 2443, 2472, 2494, 2592, 2596, 2617, 2634, 2686, 2688, 2745, 2835, 2856, 2930, 3011, 3021, 3043]. access [1975, 2348]. accessible [2019]. according [982]. account [2450]. account-based [2450]. accountability [1671]. accountable [612, 1622]. Accounting [631, 1611, 1927, 2067]. accuracy [72, 214, 536, 1354, 2217, 2452]. Accurate [247, 255, 683, 804, 904, 1273, 1606, 2090, 2188, 2900, 3007, 3110]. Achievable [1690, 1996]. Achieve [772, 1446, 2611]. Achieving [513, 745, 2064, 2189, 2388, 2826]. ACHO [303]. ACK [112]. ACL [446]. ACM [543]. ACoPE [4]. Acoustic [120, 159, 336, 711, 759, 1134, 1272, 1992, 2015, 2034, 2072, 2148, 2167, 2320, 2332, 2368, 2456, 2619]. Acoustic-based [336]. acquisition [210, 1467]. across [86, 489, 1285, 2316, 2498]. action [1265, 1799, 2019, 2052, 2549, 2863, 3096]. action-selection [2863]. activations [1006]. Active [866, 1521, 1791, 2517, 2601, 2621, 2646, 2687, 2758, 2793, 2820, 2843, 3110, 3113]. activeness [1561]. activities [354]. Activity [1369, 2090, 2901, 2988]. Activity-based [2901]. activity-level [1369]. Actor [1831, 2361, 2549, 2794, 3164]. Actor-Critic [2361, 2549, 3164]. actors [1831]. Ad [17, 124, 141, 459, 542, 617, 671, 1094, 1151, 1503, 1566, 1576, 1608, 1632, 1684, 1781, 1836, 1989, 2084, 2235, 2312, 2494, 2545, 2693, 2717, 2769, 2814, 3116, 3118, 3135]. Ad-Hoc [17, 671, 1503, 1632, 1684, 1836, 2084, 2312, 2545, 2769, 2814, 3116]. ADA [1578]. ADA-PC [1578]. Adapt [1767]. adaptability [2247]. adaptable [342, 949]. adaptation [92, 865, 1321, 1507, 1628, 1815, 2006, 2094, 2134, 3000, 3167]. adaption [1789]. Adaptive [4, 77, 171, 187, 190, 252, 275, 342, 384, 552, 568, 582, 650, 796, 871, 880, 921, 977, 1009, 1051, 1079, 1080, 1114, 1131, 1236, 1315, 1317, 1327, 1331, 1388, 1436, 1445, 1547, 1576, 1677, 1686, 1699, 1710, 1745, 1767, 1915, 1955, 1961, 2071, 2151, 2167, 2181, 2252, 2272, 2279, 2332, 2333, 2368, 2399, 2421, 2429, 2498, 2506, 2518, 2525, 2556, 2570, 2627, 2630, 2647, 2695, 2717, 2768, 2801, 2803, 2815, 2838, 2899, 2942, 2959, 2963, 2977, 2993, 3030, 3138, 3152, 3159, 3179, 3186]. Adaptively [841, 1683]. AdaptPUD [2900]. added [1155]. addition [258]. additional [1971]. additive [2125]. Address [780, 943, 949, 998, 1112, 1328, 1453, 1993, 2482, 2718, 2758]. addresses [647]. Addressing [949, 1417, 2690, 2744]. AdFeed [583]. adjacency [826]. adjacent [3149]. adjustable [97]. adjusted [1916]. adjusting [1768]. adjustment [685, 1466, 1608, 2026]. Admission [543, 683, 2549]. Adoption [1864, 2056]. ADPF [2766]. ads [452, 1127]. ADS-B [452]. ADSim [2163]. ADT [1392]. advance [1941]. Advanced [566, 1429, 1492, 1567, 2122, 2281, 2702]. advancements [367, 1938]. Advances [948, 958, 1108, 1142, 1314, 1991, 2481, 3012]. Advancing [2375, 2945]. advantages [2419]. Adversarial [332, 585, 682, 872, 1170, 1249, 1317, 2062, 2404, 2460, 2468, 2540, 2618, 2668, 2865, 2894, 3023, 3143, 3163]. adversaries [2487]. adversary [2800]. adverse [3108]. advertising [583]. advisor [846]. AE [2772]. Aerial [46, 110, 170, 322, 496, 512, 640, 714, 973, 1097, 1108, 1408, 1503, 1532, 2107, 2141, 2444, 2545, 2610, 2706, 2771, 2928, 2987]. aerial-vehicular [2141]. Aerial/Ground [1108]. AERO [2647]. AERPAW [666]. AES [806]. AF [2843, 3110]. AF-Detector [3110]. AF-MCDC [2843]. AFDX [2154]. affinity [1038]. Against [239, 309, 311, 382, 585, 608, 925, 986, 1060, 1422, 1429, 1437, 1574, 1669, 1753, 1764, 1879, 2022, 2064, 2125, 2282, 2289, 2342, 2426, 2445, 2460, 2463, 2497, 2625, 2668, 2749, 2810, 2828, 2900, 2945, 2998, 3009, 3058]. Age [1554, 2218, 2358, 2745, 2955]. age-sensitive [1554]. Agent [222, 318, 695, 778, 863, 1061, 1292, 1414, 1435, 1442, 1670, 1872, 2020, 2037, 2159, 2187, 2201, 2247, 2262, 2356, 2387, 2614, 2632, 2727, 2789, 2826, 2836, 2919, 3015, 3049, 3096]. agents [274, 2071]. Aggregate [452, 557, 3048]. aggregated [245, 3021]. Aggregation [192, 218, 241, 262, 453, 833, 841, 1443, 1878, 2160, 2202, 2257, 2420, 2464, 2494, 2736, 2798, 2896, 3058, 3087]. aggressive [2814]. aggressiveness [1011]. agnostic [1965, 2167]. agree [873]. agreement [400, 1222, 1503, 2827, 3038, 3057, 3068, 3069]. agricultural [512, 1039]. agriculture [56, 130]. AHDom [2447]. Ahead [391, 458]. AHP [207, 1789]. AI [282, 395, 481, 488, 494, 703, 944, 1089, 1101, 1248, 1291, 1390, 1392, 1410, 1475, 1597, 1850, 1965, 2043, 2180, 2230, 2292, 2370, 2498, 2514, 2518, 2526, 2718, 2731, 2844, 2897, 2902, 2969, 2997, 3047, 3118, 3151]. AI-assisted [2292]. AI-based [282, 481, 703, 2043, 2969, 3151]. AI-Driven [494, 1291, 1392, 1965, 2370, 2498, 2718, 2844, 2902]. AI-enabled [395, 1597]. AI/ML [1089, 1410, 1850, 2180, 2230, 2526]. AI/ML-aided [2230]. AI/ML-based [2526]. AIB [2997]. AIB-Metaverse [2997]. aided [12, 722, 837, 966, 1036, 1125, 1243, 1247, 1265, 1337, 1485, 1573, 1609, 1733, 1778, 2035, 2230, 2433, 2605, 2650, 2690, 2725, 2783, 2862, 2868, 3030, 3159]. AIGC [2719]. aimed [107]. Air [149, 234, 263, 638, 884, 924, 1066, 1329, 1609, 1979, 2076, 2340, 2467, 2486, 2658, 3085, 3096, 3127, 3136]. air-ground [2486, 3085]. air-to-ground [1609]. aircraft [2164]. airtime [22]. AKA [326, 794, 898, 1334, 1590, 3040]. Alarm [619, 1402]. Alert [2251, 2258]. Alexa [997]. Algorithm [6, 17, 28, 34, 37, 75, 97, 120, 135, 143, 150, 153, 203, 204, 226, 239, 256, 280, 305, 338, 384, 453, 457, 491, 552, 582, 585, 621, 645, 668, 680, 692, 714, 725, 743, 791, 821, 825, 1020, 1046, 1079, 1158, 1172, 1326, 1375, 1408, 1466, 1475, 1539, 1552, 1578, 1592, 1616, 1620, 1653, 1727, 1806, 1845, 1870, 1876, 1888, 1911, 1977, 2000, 2020, 2026, 2032, 2063, 2107, 2135, 2150, 2160, 2190, 2219, 2225, 2267, 2269, 2293, 2295, 2333, 2360, 2406, 2438, 2472, 2501, 2509, 2559, 2625, 2654, 2655, 2665, 2667, 2699, 2737, 2751, 2774, 2885, 2904, 2937, 3001, 3002, 3006, 3015, 3017, 3070, 3140, 3172]. algorithm [1805, 2080, 2802, 2934]. Algorithmically [2447]. Algorithms [151, 177, 497, 679, 778, 805, 906, 1039, 1179, 1191, 1543, 1702, 1832, 1928, 1952, 2111, 3164]. align [1452, 3041]. Aligned [3123]. alignment [2865]. all-I [1347]. all-in-one [2895]. alleviation [1448]. allocate [1831]. Allocating [1579]. Allocation [14, 22, 43, 111, 128, 219, 230, 233, 265, 302, 346, 348, 373, 383, 416, 460, 472, 491, 525, 530, 559, 573, 574, 579, 603, 628, 658, 663, 690, 730, 900, 950, 979, 1034, 1160, 1161, 1215, 1229, 1237, 1238, 1268, 1274, 1308, 1321, 1324, 1340, 1355, 1396, 1432, 1435, 1450, 1475, 1518, 1573, 1582, 1653, 1663, 1667, 1670, 1672, 1686, 1709, 1770, 1833, 1869, 1906, 1935, 2004, 2040, 2094, 2106, 2155, 2187, 2219, 2294, 2296, 2313, 2386, 2439, 2448, 2456, 2466, 2510, 2522, 2529, 2531, 2576, 2577, 2673, 2685, 2806, 2830, 2878, 2907, 2919, 2921, 2950, 2974, 3021, 3029, 3056, 3067, 3081, 3127, 3166]. allocation [34, 153, 192, 226, 440, 488, 588, 729, 731, 733, 785, 945, 981, 1036, 1206, 1285, 1292, 1338, 1447, 1472, 1485, 1666, 1687, 1728, 1739, 1907, 1937, 2247, 2255, 2345, 2459, 2589, 2602, 2610, 2629, 2632, 2655, 2726, 2756, 2835, 2850, 2873, 2932, 3020, 3172]. allocations [2149]. allreduce [2986]. ALOHA [818, 1748, 1780]. ALOHA-NUI [1780]. Alone [329, 834]. Alpha [2975]. Altair [2595]. alternate [1955]. Alternating [3187]. alternative [210, 428, 877, 2953]. alternatives [631]. altitude [109, 1861, 2467, 2673]. Amazon [997]. Ambient [554, 1235, 1335, 1938, 2129, 2309]. amid [2055]. AMLFN [1576]. AMLFN-AD [1576]. among [69, 253, 1035]. amplifier [963]. AMQP [1523]. analog [2867]. Analyses [1575, 2035]. Analysis [21, 44, 63, 89, 94, 102, 121, 194, 224, 284, 298, 300, 304, 316, 364, 368, 370, 424, 430, 443, 456, 462, 504, 555, 671, 684, 719, 732, 763, 774, 786, 805, 865, 923, 964, 968, 1014, 1033, 1054, 1071, 1083, 1158, 1181, 1293, 1322, 1325, 1355, 1402, 1407, 1480, 1505, 1543, 1589, 1691, 1708, 1736, 1737, 1740, 1748, 1754, 1777, 1788, 1802, 1814, 1830, 1866, 1902, 1941, 1971, 1985, 2028, 2045, 2056, 2112, 2117, 2169, 2178, 2180, 2197, 2208, 2227, 2264, 2266, 2275, 2400, 2427, 2521, 2542, 2546, 2830, 2867, 2973, 2975, 2994, 2996, 3042, 3091, 3125, 3151]. analysis [267, 287, 505, 792, 993, 1461, 1991, 2369, 2663, 2674, 2687, 2764, 2875, 3043, 3103, 3136]. analysis-based [1158]. Analytical [132, 207, 255, 314, 1003, 1692, 2278, 2691].

Analytics [56, 208, 788, 983, 1327, 1767, 1999, 2094, 2972, 3052]. analyze [2081]. analyzer [2959]. Analyzing [909, 1005, 1276, 1476, 1773, 2603]. Anchor [1569, 1937]. Anchor-based [1569]. anchored [2995]. Android [42, 52, 430, 1279, 2956]. angle [1411]. anisotropic [289]. ANN [151]. annotation [2720]. announcements [531]. anomalies [531]. Anomaly [29, 74, 311, 434, 456, 572, 616, 860, 934, 993, 1067, 1175, 1256, 1365, 1557, 1762, 1929, 2029, 2163, 2236, 2316, 2343, 2360, 2527, 2591, 2593, 2594, 2729, 2753, 2770, 2886, 2909, 2958, 2960, 3023, 3025, 3050, 3080, 3162]. anonymity [326, 590, 1068, 1439, 1461, 2064, 2401]. anonymity-preserving [1439]. Anonymization [1821, 1877]. Anonymous [400, 452, 838, 1468, 1508, 1671, 1883, 2165, 2259, 2367, 2431, 2497, 2617, 2831, 3057]. answering [1827]. Ant [478, 743, 1784, 1876, 2234, 2509]. antagonistic [855]. antenna [139, 334, 811, 1229, 1770, 1957, 3097]. antennas [1397, 1684]. Anti [135, 439, 503, 1173, 1486, 1747, 1903, 2225, 2372, 2648, 2766, 3020]. anti-collision [135, 2648]. anti-cybercrime [1747]. Anti-inference [2766]. Anti-jamming [439, 503, 1173, 3020]. anti-long-range [2225]. anti-LPB [1903]. anti-LS [2372]. Anti-quantum [1486]. anticipated [2126]. AntidoteFL [2998]. Anycast [2711]. AoI [2129, 2507, 2709]. AOPDDQN [2829]. AP [1940]. APCC [3113]. APGNN [1402]. API [2749]. APIs [42]. App [290, 870, 903, 1256, 1273, 1596]. applicability [281]. applicable [806]. Application [58, 388, 420, 548, 658, 746, 877, 940, 1029, 1068, 1270, 1479, 1498, 2297, 2307, 2386, 2425, 2578, 2646, 2976]. Application-Aware [2578]. Applications [130, 175, 285, 313, 325, 336, 375, 518, 557, 577, 606, 651, 726, 838, 889, 922, 1015, 1026, 1074, 1113, 1135, 1148, 1223, 1235, 1288, 1373, 1494, 1496, 1513, 1524, 1543, 1580, 1585, 1625, 1636, 1671, 1687, 1692, 1771, 1825, 1862, 1930, 1938, 1949, 2035, 2061, 2086, 2091, 2206, 2309, 2324, 2327, 2451, 2455, 2470, 2473, 2498, 2526, 2545, 2620, 2638, 2639, 2701, 2722, 2762, 2801, 2816, 2818, 2855, 3082, 3147, 3169]. applied [828, 2934]. Applying [1740, 2212]. apprenticeship [2459]. Approach [4, 98, 134, 142, 144, 191, 223, 265, 291, 312, 327, 347, 371, 383, 409, 413, 428, 455, 495, 551, 561, 571, 595, 612, 633, 692, 697, 738, 778, 810, 844, 853, 897, 991, 1025, 1032, 1117, 1126, 1251, 1275, 1339, 1347, 1360, 1382, 1426, 1435, 1470, 1516, 1530, 1553, 1561, 1605, 1677, 1796, 1797, 1804, 1820, 1877, 1885, 1891, 1973, 2012, 2015, 2071, 2072, 2081, 2091, 2120, 2136, 2144, 2164, 2169, 2170, 2172, 2227, 2232, 2294, 2353, 2387, 2468, 2496, 2624, 2644, 2723, 2815, 2884, 2902, 2958, 2967, 2989, 3011, 3034, 3072, 3094, 3096, 3118, 3145, 3160, 3171, 3182]. approach [3, 102, 247, 308, 317, 365, 431, 468, 506, 617, 729, 755, 830, 907, 943, 1019, 1106, 1141, 1171, 1206, 1321, 1335, 1354, 1414, 1494, 1514, 1525, 1558, 1644, 1661, 1734, 1739, 1756, 1793, 1831, 1960, 1989, 2074, 2162, 2215, 2237, 2314, 2335, 2385, 2417, 2425, 2465, 2549, 2598, 2679, 2718, 2729, 2814, 2820, 2829, 2900, 2932, 2964, 3020, 3039, 3084, 3162]. Approaches [100, 387, 482, 749, 823, 828, 904, 1071, 1282, 1372, 1716, 1934, 2012, 2017, 2028, 2044, 2495, 2695, 2734, 2769, 2793]. Approver [819]. Approximate [738, 937, 1705, 1950, 2876, 2934]. approximation [300, 924, 1889]. apps [1369, 2045]. APQA [2617]. April [106, 126, 547, 570, 1024, 1044, 1520, 1550, 2042, 2078, 2742, 2779]. APT [2800, 3076]. APT-ATT [3076]. APTs [2453]. AQM [196, 2082]. AQRS [1486]. Aquila [1686]. AR-GAIL [1317]. ARA [2816]. Arbiter [672]. Arbiter-based [672]. Architecting [859]. architectural [510]. Architecture [8, 78, 117, 160, 194, 318, 351, 357, 486, 517, 539, 624, 634, 721, 792, 808, 907, 1069, 1100, 1132, 1165, 1266, 1298, 1371, 1423, 1652, 1708, 1727, 1825, 1864, 1879, 1922, 2052, 2184, 2223, 2233, 2238, 2275, 2338, 2370, 2454, 2627, 2676, 2801, 2825, 2854, 3014, 3016, 3100, 3160, 3176]. architecture-based [2052]. architecture-level [351]. Architectures [32, 61, 279, 410, 606, 658, 909, 1015, 1059, 1160, 1375, 2277, 2418, 2545, 2546, 2739, 3059]. ArchW3 [2801]. Area [153, 229, 242, 289, 625, 1105, 1290, 1508, 1600, 1639, 1651, 1909, 1977, 2202, 2377, 2446, 2638, 2784, 2938, 3114, 3119]. area-scalable [229]. areas [321, 1880]. Arena [334]. ARES [1315]. ARIMA [657, 2044]. ARIMA-BP [657]. ARM [806]. ARM-based [806]. armed [302, 1640, 2015, 2144, 2245]. Arnold [2622]. arrangement [418]. array [2320]. arrival [214]. Art [391, 399, 458, 537, 615, 646, 1007, 1488, 1531, 2027, 2429, 3012]. artifacts [1005]. Artificial [380, 399, 405, 434, 548, 696, 1007, 1260, 1378, 1513, 1583, 2027, 2044, 2293, 3117]. AS-level [1461]. AS-T [2223]. ASAP [2496]. ASIRIA [1098]. aspect [2263]. Aspects [1585, 3147]. ASRPH [746]. assess [1316]. Assessing [148, 284, 601, 1783, 2229]. Assessment [154, 210, 271, 565, 675, 1026, 1147, 1524, 1598, 1602, 1605, 1822, 1852, 2096, 2103, 2122, 2136, 2390, 2739, 2853, 2892]. asset [811, 856]. assets [913, 1322, 1969]. Assigning [712]. Assignment [133, 136, 177, 265, 484, 713, 753, 1002, 1056, 1123, 1323, 1477, 1491, 1685, 1711, 1744, 1974, 2223, 2234, 2629, 2888, 2901, 2917, 2944, 2979, 3027]. assist [2128]. assistance [626, 3039]. assistant [2712]. Assisted [59, 76, 195, 371, 404, 416, 460, 560, 609, 730, 817, 863, 979, 1039, 1040, 1073, 1099, 1211, 1260, 1349, 1370, 1413, 1414, 1503, 1526, 1544, 1631, 1651, 1654, 1663, 1760, 1788, 1790, 1796, 1809, 1856, 1873, 1888, 1957, 2037, 2091, 2094, 2129, 2130, 2132, 2142, 2181, 2201, 2205, 2257, 2292, 2294, 2314, 2317, 2394, 2407, 2408, 2443, 2471, 2489, 2510, 2552, 2563, 2578, 2629, 2667, 2673, 2681, 2709, 2722, 2727, 2740, 2829, 2845, 2921, 2943, 2967, 2979, 3008, 3020, 3034, 3049, 3065, 3077, 3089, 3111, 3144, 3166]. Assisting [238, 1965]. association [199, 233, 365, 628, 744, 755, 1053, 1125, 1271, 1670, 2128, 2489, 2840, 3041]. associations [563, 1724, 2400]. assumptions [148]. assurance [1735, 2895]. ASVTuw [2332]. asymmetric [658, 2326]. AsyncDefender [3000]. Asynchronous [824, 825, 1152, 1325, 1578, 2154, 2189, 2935, 3000, 3016]. AT-GCN [1838]. Atlas [165]. ATM [3126]. ATSSS [2005]. ATT [3076]. ATT&CK [2122]. ATT&CK-based [2122]. Attack [80, 151, 189, 400, 499, 516, 533, 541, 585, 779, 782, 806, 809, 964, 1097, 1165, 1180, 1205, 1257, 1377, 1456, 1499, 1570, 1576, 1598, 1628, 1662, 1804, 1819, 1838, 2083, 2127, 2186, 2225, 2243, 2266, 2357, 2445, 2496, 2593, 2599, 2668, 2671, 2717, 2749, 2787, 2804, 2823, 2836, 2838, 2883, 2894, 2943, 2960, 3037, 3075, 3102, 3160, 3163, 3167, 3177]. Attack-adaptive [2838]. Attack-data [2668]. attack-resilience [400]. attacker [1345]. Attackers [1326]. attacking [2303]. Attacks [11, 44, 112, 148, 155, 164, 304, 311, 312, 497, 503, 561, 596, 672, 751, 807, 822, 834, 857, 884, 984, 1041, 1053, 1060, 1158, 1251, 1437, 1444, 1525, 1688, 1706, 1753, 1754, 1764, 1828, 1947, 1981, 1993, 2122, 2125, 2282, 2289, 2342, 2403, 2426, 2460, 2463, 2502, 2540, 2569, 2579, 2603, 2625, 2633, 2668, 2774, 2800, 2804, 2847, 2872, 2884, 2987, 2998, 3009, 3023, 3051, 3058, 3088, 3098, 3125]. Attention [1175, 1533, 1574, 1770, 1792, 1800, 2011, 2059, 2090, 2204, 2350, 2682, 2823, 2839, 3103, 3150, 3153]. Attention-based [2059]. attention-driven [2823]. attentive [737]. Attribute [702, 885, 1140, 1307, 1683, 1802, 1839, 1883, 2153, 2265, 2447, 2523, 2535, 2634, 2688, 2888, 3141]. Attribute-Based [702, 885, 1683, 1839, 1883, 2153, 2265, 2523, 2535, 2634]. attribute-level [2688]. attributed [2192]. attributes [936, 1452]. attribution [3076]. Auction [73, 236, 297, 398, 549, 1285, 1459, 1532, 2164, 2327, 2353, 2475, 2522, 2726, 2897, 3090, 3092, 3127, 3185]. auction-based [1459, 2327, 2726, 2897, 3127]. audience [2898]. Auditable [2322]. auditing [708, 1252, 1601, 1840, 2114, 2722, 2828, 3063]. augmentation [1602, 2865, 3007]. Augmented [697, 822, 865, 2789, 2952, 3011]. Augmenting [918, 2663]. August [244, 710, 1169, 1694, 2301, 2966]. AURA [628]. AURA-5G [628]. Aurora [60]. authenticated [2827, 3069]. Authentication [100, 122, 242, 245, 258, 326, 328, 400, 407, 452, 475, 519, 526, 599, 649, 728, 768, 795, 838, 840, 845, 885, 898, 1013, 1018, 1027, 1115, 1185, 1200, 1222, 1283, 1287, 1300, 1341, 1350, 1387, 1439, 1478, 1503, 1508, 1517, 1544, 1548, 1551, 1590, 1641, 1671, 1834, 1914, 1919, 1923, 1927, 1969, 1975, 2013, 2072, 2089, 2102, 2143, 2165, 2175, 2176, 2185, 2195, 2261, 2354, 2367, 2397, 2431, 2494, 2505, 2617, 2666, 2689, 2716, 2721, 2776, 2810, 2822, 2847, 2877, 2924, 2999, 3030, 3038, 3057, 3065, 3068, 3079, 3131, 3135, 3138]. authorities [3083]. authority [434, 702, 1288, 1671, 1683, 1791, 2265, 2443, 2523]. Authorization [1593, 1927, 2451]. Auto [327, 333, 353, 805, 1180, 1539, 1557, 2038, 2404]. Auto-3P [333]. auto-encoder [1180, 1557, 2404]. auto-provisioning [2038]. auto-scaling [353]. Auto-tune [805, 1539]. Autoencoder [719, 1025, 2770, 2772]. autoencoder-based [719]. Autoencoders [816]. Automated [1310, 1657, 1724, 2066, 2222, 2306, 2555, 2939, 3066]. Automatic [268, 883, 914, 1166, 2097, 2285, 2496, 2720, 3054]. automating [2390]. automation [2057]. automotive [649]. Autonomous [309, 333, 477, 494, 505, 626, 859, 1108, 1371, 1395, 1444, 1539, 2197, 2325, 2369, 2440, 2526, 2561, 2776, 2789, 2929, 3128]. AUTOSAR [1313]. AUTOSAR-based [1313]. Autoscaling [1289]. AUV [2783].

AUV-aided [2783]. availability [463, 1046, 2670]. available [60]. avatar [2341]. avatars [2822]. average [826, 1502, 2166, 2272, 3129]. averaging [3004]. avionic [2154]. Avoidance [131, 485, 759, 1114, 1577, 2209, 2489, 2648, 2808, 2856, 2957]. avoided [2028]. Aware [9, 16, 65, 69, 136, 137, 206, 246, 250, 279, 319, 339, 346, 349, 405, 413, 439, 444, 551, 583, 625, 655, 678, 695, 701, 724, 739, 781, 796, 837, 865, 938, 974, 975, 1051, 1056, 1064, 1126, 1129, 1164, 1166, 1184, 1199, 1227, 1258, 1306, 1315, 1324, 1388, 1507, 1516, 1521, 1527, 1540, 1600, 1610, 1654, 1663, 1713, 1792, 1814, 1833, 1867, 1891, 1901, 1935, 1945, 1946, 1973, 1975, 2028, 2040, 2053, 2065, 2128, 2147, 2163, 2172, 2248, 2267, 2323, 2335, 2364, 2578, 2614, 2640, 2650, 2714, 2732, 2749, 2752, 2792, 2830, 2841, 2898, 2917, 2925, 2928, 2949, 2974, 2996, 3046, 3104]. aware [38, 76, 91, 111, 124, 133, 168, 177, 186, 188, 193, 210, 240, 283, 363, 385, 403, 476, 504, 553, 645, 668, 676, 809, 821, 851, 859, 989, 1032, 1038, 1088, 1151, 1153, 1238, 1246, 1286, 1318, 1434, 1463, 1494, 1698, 1732, 1757, 1794, 1812, 1921, 1949, 1952, 1968, 1970, 2063, 2173, 2234, 2237, 2247, 2276, 2348, 2417, 2508, 2550, 2556, 2592, 2619, 2631, 2679, 2715, 2759, 2834, 2840, 2861, 2937, 2954, 2955, 2992, 3019, 3183]. Awareness [882, 914, 1047, 1102, 1836, 1884, 2199, 2308, 2495, 2615]. AWS [443]. Azure [443].

B [452]. B2H [1078]. B5G [367, 541, 955, 973, 1243, 1271, 1705, 1869, 2389, 2577, 2580, 2707, 3014, 3151]. B5G-driven [955]. B5G/5G [2707]. B5G/6G [2580]. BaaS [371]. back [1466]. backbone [1211, 2454]. backbone-assisted [1211]. backend [2019]. Background [2275]. backhaul [416, 798, 963, 988, 1030, 1034, 1438, 1447, 1627, 2184]. backhauling [562, 1297]. backoff [524, 2003]. Backpressure [3035]. backscatter [195, 554, 1178, 1235, 1485, 1676, 2008, 2129, 2309, 2672]. backscatter-aided [1485]. backscatter-assisted [195, 2129]. backstage [1574]. backstepping [196]. backtracking [491]. backup [484, 535, 1935, 2480, 3183]. backward [1902]. badger [2774]. bag [1949]. bag-of-tasks [1949]. Balanar [1798]. Balanced [563, 950, 1102, 1701, 2293]. Balancing [77, 102, 222, 261, 286, 339, 447, 543, 601, 793, 987, 1011, 1035, 1084, 1183, 1210, 1330, 1522, 1572, 1620, 1798, 1891, 1964, 2152, 2160, 2489, 2539, 2551, 2616, 2676, 2699, 2731, 2970, 2991]. ballot [2831]. Band [65, 308, 388, 450, 493, 698, 1095, 1228, 1233, 1310, 1333, 1471, 1830, 1835, 1988, 2033, 2082, 2292, 2611, 2944, 2954, 2983, 3114]. bandit [1640, 2015, 2144, 2245]. bandits [302]. bands [2000, 2049]. Bandwidth [4, 92, 226, 305, 390, 451, 622, 773, 792, 871, 970, 1011, 1248, 1327, 1471, 1527, 1653, 1687, 1699, 1948, 2070, 2106, 2164, 2573, 2713]. Bandwidth-aware [1527]. bandwidth-constrained [1948]. Bandwidth-efficient [1248, 1327]. bandwidth-optimal [622]. bargaining [230]. barrier [1592]. Bartolomeu [86]. Base [109, 110, 231, 632, 818, 1691, 1772, 1838, 1951, 2074, 2724, 2771, 3001, 3107]. baseband [2126, 2592]. baseband-function [2126]. Based [19, 113, 205, 223, 264, 279, 282, 289, 312, 326, 334, 336, 340, 345, 347, 380, 451, 455, 478, 528, 579, 590, 611, 699, 711, 724, 726, 728, 734, 747, 751, 802, 840, 946, 976, 983, 1018, 1026, 1048, 1072, 1080, 1123, 1157, 1205, 1287, 1288, 1310, 1313, 1339, 1374, 1382, 1420, 1460, 1466, 1475, 1517, 1552, 1598, 1632, 1652, 1662, 1704, 1719, 1782, 1785, 1825, 1863, 1910, 1984, 2022, 2047, 2072, 2081, 2098, 2143, 2150, 2341, 2370, 2427, 2442, 2468, 2483, 2488, 2490, 2495, 2529, 2620, 2624, 2633, 2634, 2730, 2752, 2843, 2886, 2957, 3002, 3112, 3128, 3170]. based [64, 179, 181, 204, 207, 233, 237, 333, 411, 459, 538, 544, 561, 582, 585, 604, 627, 672, 685, 703, 730, 758, 764, 768, 806, 931, 965, 990, 1104, 1133, 1160, 1166, 1173, 1180, 1185, 1190, 1201, 1258, 1308, 1309, 1331, 1361, 1405, 1455, 1474, 1483, 1504, 1516, 1518, 1541, 1543, 1607, 1623, 1679, 1683, 1731, 1839, 1869, 1889, 1903, 1929, 1933, 1961, 2053, 2085, 2106, 2138, 2232, 2236, 2246, 2266, 2267, 2313, 2337, 2351, 2372, 2384, 2398, 2454, 2503, 2523, 2526, 2628, 2644, 2662, 2707, 2709, 2749, 2755, 2780, 2830, 2836, 2884, 2897, 2907, 2919, 2973, 2980, 3062, 3067, 3076, 3094, 3135, 3181, 3185]. based [6, 23, 55, 60, 89, 171, 229, 230, 254, 294, 311, 313, 316, 375, 409, 446, 450, 543, 602, 610, 612, 713, 715, 739, 759, 864, 906, 915, 1065, 1071, 1083, 1094, 1105, 1132, 1164, 1174, 1197, 1215, 1278, 1283, 1345, 1462, 1470, 1505, 1532, 1540, 1569, 1573, 1631, 1636, 1641, 1646, 1664, 1697, 1770, 1792, 1804, 1817, 1822, 1872, 1912, 1914, 1916, 1923, 1958, 1976, 1977, 1985, 2003, 2045, 2067, 2107, 2169, 2187, 2195, 2202, 2259, 2350, 2393, 2434, 2497, 2540, 2558, 2597, 2656, 2671, 2705, 2717, 2777, 2806, 2819, 2859, 2877, 2893, 2969, 2977, 3024, 3047, 3081, 3102, 3109, 3118, 3158–3160]. based [13, 93, 196, 245, 259, 290, 299, 402, 417, 437, 469, 498, 508, 549, 564, 589, 642, 663, 705, 782, 826, 845, 926, 979, 992, 1009, 1021, 1059, 1079, 1121, 1178, 1186, 1257, 1274, 1279, 1294, 1303, 1351, 1403, 1443, 1530, 1572, 1581, 1584, 1663, 1671, 1711, 1718, 1769, 1797, 1826, 1827, 1837, 1838, 1840, 1849, 1859, 1870, 1874, 1883, 1917, 1953, 1974, 2004, 2039, 2043, 2052, 2061, 2063, 2076, 2127, 2182, 2190, 2226, 2235, 2310, 2331, 2353, 2366, 2371, 2399, 2443, 2451, 2469, 2479, 2533, 2536, 2581, 2590, 2627, 2642, 2665, 2667, 2720, 2725, 2837, 2904, 2916, 2938, 2952, 2999, 3001, 3016, 3030, 3038, 3054, 3057, 3064, 3065, 3072, 3074, 3127, 3133, 3137, 3140, 3145, 3148, 3178]. based [18, 46, 180, 321, 350, 371, 457, 466, 471, 488, 519, 595, 714, 719, 721, 765, 785, 790, 794, 809, 817, 821, 838, 861, 903, 908, 916, 937, 973, 995, 1014, 1032, 1047, 1062, 1103, 1130, 1144, 1158, 1159, 1206, 1210, 1217, 1222, 1238, 1254, 1267, 1296, 1302, 1326, 1365, 1375, 1394, 1398, 1423, 1463, 1489, 1511, 1629, 1653, 1656, 1677, 1702, 1707, 1717, 1724, 1742, 1756, 1763, 1781, 1794, 1847, 1864, 1880, 1928, 1934, 1952, 1962, 2199, 2209, 2216, 2253, 2262, 2272, 2306, 2316, 2342, 2400, 2422, 2494, 2525, 2550, 2700, 2715, 2754, 2863, 2960, 3123, 3125, 3130, 3132, 3162]. based [28, 45, 117, 122, 153, 156, 157, 252, 273, 304, 322, 358, 387, 389, 412, 418, 453, 454, 465, 481, 516, 599, 616, 648, 654, 655, 671, 702, 708, 732, 733, 760, 860, 869, 875, 877, 936, 981, 993, 1066, 1074, 1086, 1181, 1216, 1236, 1240–1242, 1263, 1272, 1328, 1336, 1338, 1341, 1362, 1364, 1408, 1412, 1436, 1459, 1471, 1542, 1644, 1682, 1800, 1802, 1806, 1832, 1841, 1848, 1871, 1898, 1987, 2083, 2084, 2152, 2200, 2231, 2255, 2265, 2274, 2357, 2376, 2417, 2440, 2450, 2504, 2534, 2552, 2561, 2604, 2666, 2691, 2718, 2724, 2727, 2737, 2831, 2846, 2917, 2937, 2991, 3007, 3151, 3175]. based [2, 37, 54, 139, 158, 226, 235, 267, 318, 338, 425, 431, 442, 533, 552, 598, 626, 657, 660, 669, 686, 689, 729, 766, 784, 858, 885, 1011, 1013, 1070, 1077, 1149, 1183, 1224, 1290, 1292, 1334, 1457, 1655, 1732, 1746, 1779, 1812, 1836, 1862, 1876, 1892, 1907, 1918, 1955, 1970, 2013, 2020, 2021, 2031, 2096, 2130, 2135, 2151, 2153, 2183, 2204, 2268, 2270, 2289, 2291, 2327, 2339, 2471, 2491, 2530, 2535, 2575, 2613, 2617, 2641, 2645, 2655, 2669, 2684, 2689, 2696, 2710, 2719, 2726, 2750, 2751, 2758, 2788, 2791, 2805, 2810, 2823, 2828, 2839, 2847, 2850, 2873, 2882, 2885, 2891, 2896, 2936, 2981, 3049, 3083, 3085, 3106, 3136, 3141, 3152, 3172, 3179]. based [67, 74, 141, 182, 257, 285, 317, 520, 647, 762, 799, 880, 882, 932, 945, 987, 1038, 1140, 1188, 1230, 1312, 1449, 1525, 1566, 1601, 1689, 1757, 1805, 1824, 1835, 1856, 1868, 1919, 1947, 1956, 2006, 2011, 2059, 2080, 2086, 2095, 2122, 2165, 2202, 2213, 2215, 2314, 2321, 2363, 2411, 2431, 2493, 2520, 2569, 2606, 2619, 2643, 2708, 2713, 2722, 2738, 2764, 2829, 2900, 2901, 3089]. based-Context [655]. Basic [1313]. basics [1483]. BASM [819]. BASTION [2939]. batch [1974, 2195, 3058, 3065]. batch-aggregation [3058]. batch-based [1974]. batched [2093]. Battery [306, 1460, 1970, 2993]. battery-constricted [1970]. Battery-less [306]. Battlefield [1854, 2777]. Bayesian [28, 390, 1231, 2663, 3004, 3089]. BBR [1376, 1446, 2485]. BBR-R [2485]. BBRv2 [1011]. BBU [231, 339, 1210]. BC [1350]. BC-enabled [1350]. BCN [3035]. BCSA [1552]. BD [2310, 3165]. BD-RIS [3165]. BD-TTS [2310]. be [767, 1990]. Beacon [1072]. Beam [783, 1073, 1191, 1294, 1397, 1957, 2480, 2657, 3041]. beam-align [3041]. beam-hopping [2657]. beam-width [1957]. beamformer [1661]. Beamforming [521, 716, 755, 770, 985, 1597, 1657, 2510, 2694, 2867, 3097, 3168]. beams [3097]. BeamSense [2694]. beat [143]. bee [322, 405, 2293]. BEE-DRONES [322]. behavior [1176, 1476, 1837, 1841, 1902, 2222, 2266, 2282, 2322, 2372]. behavioral [1322, 2973]. behaviors [836, 1572, 2973]. behaviour [3102]. behaviour-based [3102]. bell [708]. benchmarking [537, 2637]. Benders [2074]. beneficial [2818]. benefits [123, 194, 1095]. BER [136, 2407]. BERT [2425]. Best [143, 819, 1695, 1829]. best-known [143]. Betastack [3107]. better [91, 1371, 2611]. between [278, 768, 920, 1259, 2244, 3066]. betweenness [3145]. Beyond [174, 334, 377, 575, 692, 744, 775, 785, 807, 808, 893, 908, 944, 945, 1028, 1065, 1089, 1177, 1291, 1331, 1389, 1571, 1582, 1583, 1597, 1709, 1850, 1864, 1882, 1939, 1940, 1946, 1965, 2051, 2116, 2173, 2184, 2186, 2214, 2233, 2242, 2279, 2690, 2739, 2844, 2871, 2939]. BFL [2376]. BFLS [1495]. BFT [953, 1348, 2222, 2802, 2977]. BFT-based [2977]. BFTDiagnosis [2222]. BGP [531, 925, 1004, 1175, 1302, 1399, 1959, 2029, 2591]. BGPattacks [1646]. BGSD [1405]. bi [352, 2236]. bi-level [352]. Bi-LSTM [2236]. Bi-LSTM/GRU-based [2236]. BIAS [2431]. biased [907]. bidirectional [1711, 2223, 2781, 3143]. big [56, 62, 602, 631, 942]. BigFooT [953]. Bilateral [2762]. Binary [984, 1552]. bins [321]. biometric [2431]. biometric-based [2431]. bipartite [572]. bistatic [1676]. Bit [31, 2326]. BiTCN [2061]. Bitcoin [127, 991, 1012, 1363, 2505, 2670]. bitmap [544]. bitmap-based [544]. bitrate [1131, 2630, 3109]. bivariate [1240]. Black [309, 2126, 2307, 2737, 3088]. Black-box [309, 2126, 2307, 3088]. black-winged [2737]. blacklisting [3157]. BLAW [3008]. BLE [1162, 1462, 3008]. Blending [1528]. blind [2497]. block [142, 910, 1012, 2506, 2767]. Blockage [798]. blockages [1657]. Blockchain [264, 285, 414, 602, 606, 612, 707, 715, 721, 749, 768, 802, 819, 826, 856, 861, 863, 864, 873, 909, 913, 916, 932, 946, 953, 965, 995, 1042, 1078, 1105, 1144, 1152, 1172, 1178, 1197, 1288, 1302, 1336, 1345, 1398, 1423, 1460, 1477, 1486, 1495, 1513, 1517, 1585, 1622, 1646, 1697, 1714, 1730, 1740, 1790, 1840, 1849, 1859, 1912, 1934, 1962, 1998, 2013, 2066, 2067, 2088, 2102, 2114, 2119, 2132, 2185, 2191, 2259, 2261, 2295, 2310, 2319, 2371, 2374, 2376, 2394, 2398, 2454, 2536, 2634, 2773, 2798, 2801, 2828, 2831, 2877, 2884, 2889, 2892, 2916, 2962, 2980, 3024, 3030, 3057, 3065, 3083, 3100, 3115, 3135, 3146]. blockchain [179, 454, 605, 858, 875, 945, 955, 1013, 1216, 1250, 1362, 1454, 1503, 1611, 1682, 1774, 1802, 1854, 1900, 2031, 2095, 2133, 2153, 2215, 2223, 2312, 2423, 2450, 2520, 2733, 2750, 2847, 2866, 2885, 2896, 3048, 3055, 3172]. Blockchain-assisted [1790, 2132, 2394, 3065]. Blockchain-based [179, 602, 612, 715, 802, 858, 861, 864, 916, 932, 946, 965, 995, 1013, 1105, 1144, 1178, 1302, 1336, 1345, 1460, 1517, 1697, 1840, 1849, 1859, 1962, 2013, 2067, 2398, 2454, 2828, 2831, 2847, 2896, 2916, 3024, 3057, 3083, 3135, 3172]. blockchain-empowered [1854]. Blockchain-Enabled [414, 819, 873, 945, 955, 2102, 2153, 2866, 2962, 3115]. Blockchain-enhanced [1477]. blockchain-federated [2798].

Blockchain-native [913]. blockchain-oriented [707]. blockchain-powered [749, 3146]. Blockchained [960, 2115, 2521]. blockchains [596, 905, 1104, 2573, 3006]. blocking [1184, 2664]. blocklength [2629, 2913]. blocks [842, 2775, 3087]. Bloom [720, 1656]. Blueprint [580]. Bluetooth [342, 693, 980, 1071, 1093]. BMS [1527]. Board [25, 48, 70, 87, 105, 125, 146, 166, 184, 200, 220, 243, 269, 295, 323, 360, 396, 421, 448, 479, 500, 522, 546, 569, 591, 613, 635, 664, 687, 709, 740, 776, 812, 847, 878, 911, 928, 951, 971, 999, 1023, 1043, 1057, 1075, 1091, 1109, 1136, 1168, 1203, 1220, 1261, 1304, 1342, 1379, 1415, 1440, 1464, 1481, 1519, 1549, 1564, 1587, 1613, 1634, 1649, 1673, 1693, 1725, 1765, 1807, 1842, 1904, 1925, 1943, 1966, 2001, 2041, 2077, 2104, 2139, 2156, 2193, 2220, 2240, 2300, 2346, 2379, 2565]. Board [2413, 2511, 2568, 2586, 2661, 2704, 2743, 2778, 2811, 2848, 2879, 2911, 2926, 2947, 2965, 3032, 3120, 3189]. boards [2024]. Body [121, 242, 1105, 1508, 1580, 2377, 2446, 2938]. bonding [712]. boosted [1702]. Boosting [1574, 2334, 3005]. border [2637]. bot [1595, 2477]. both [135, 597]. botnet [864, 1303, 1421, 1800, 2478, 2787, 2994]. botnets [382]. bottleneck [651, 1841, 1853, 2598, 2599]. bound [1211, 1665]. Boundary [1920]. bounded [1690, 2831, 3100]. bounded-time [3100]. box [309, 2126, 2307, 3088]. BP [139, 657, 1240, 2223]. BP-MGKM [1240]. brain [116, 2532]. branch [731, 2052, 2366, 2839, 3022, 3188]. Breadth [679]. Breadth-first [679]. breaking [2773]. Broadband [227, 288, 427, 1052, 2148, 2816, 2944]. Broadcast [371, 381, 1103, 1683, 2218, 2407]. Broadcasting [248, 1165]. broker [1492, 1976]. Brownian [1377]. browser [425]. browser-based [425]. browsing [206, 788, 1389]. BRP [226]. BS [851, 2771].

BSS [302]. BTG [2973]. BTG-RF [2973]. BTMH [3146]. BTSM [819]. budget [1469, 1546, 2327, 2992]. budget-constrained [2327]. budgeted [1973]. Buffer [12, 567, 837, 2990]. Buffer-aided [12, 837]. buffering [45]. Building [55, 182, 905, 1284, 1576, 2775]. buildings [817]. Bulwark [2018]. bundling [2806]. bundling-based [2806]. Burn [640, 856]. Burn-to-Claim [856]. burst [3045]. bursty [314, 436, 531, 2280]. bus [2716]. Business [1226]. busy [2542]. Butterworth [1958]. BWpredictor [2713]. BwShare [92]. Bytes [669]. Bytes-based [669]. ByteSGAN [872]. Byzantine [825, 1454, 1879, 2333, 2420, 2488, 2625, 2904, 2995, 3000]. Byzantine-robust [2420, 3000].

C [207, 314, 318, 1095, 1210, 1424, 2145, 2184, 2561, 2944]. C-ITS [2145]. C-RAN [314, 318, 1210, 2184]. C-V2X [1424, 2561]. CA [524, 726, 2317]. CA-Live360 [2317]. CAAST [2711]. CAB [534]. Cable [1449]. cables [2007]. Cache [23, 123, 300, 514, 645, 806, 829, 877, 1061, 1727, 1804, 1945, 2454, 2686, 2711, 2846, 2862, 2881, 2983, 3049]. cache-aided [2862]. Cache-assisted [3049]. Cache-based [877]. cache-enabled [1945]. Cache-INT [2983]. CacheNet [954]. caches [256, 351, 895, 1140]. Caching [128, 168, 188, 294, 366, 426, 534, 558, 632, 722, 742, 836, 862, 1131, 1149, 1156, 1159, 1253, 1360, 1553, 1589, 1642, 1700, 1716, 1787, 1844, 1874, 1897, 1945, 1973, 1984, 2020, 2080, 2093, 2135, 2144, 2161, 2226, 2270, 2280, 2288, 2387, 2396, 2438, 2448, 2529, 2610, 2628, 2635, 2700, 2840, 2845, 2862, 2983, 3018]. caching-enabled [2983]. CAD [3080]. CAEAID [2788]. CAI [655]. CAIN [1600]. Calculation [1263]. calculations [2297]. calendaring [823]. Calibration [716]. Call [419]. camera [2728]. camouflaged [758]. Campus [94, 217, 1861, 2075, 2098, 2729, 3113]. can [143, 1513, 2716]. cancellation [1955]. candidate [1789]. capabilities [484, 2051, 2141]. capability [554, 2901]. capable [2158]. capacitated [348, 764]. Capacity [255, 408, 985, 1433, 1676, 1816, 1847, 1989, 2124, 2230, 2305, 2388]. CapRadar [1699]. capsule [766]. CAPTAIN [120]. carbon [2067]. cardinality [1865]. care [242, 519, 746, 1490]. Carlo [1924]. carrying [2079]. cars [2527]. CART [31]. cascaded [2786]. cascading [239, 2533]. case [47, 115, 131, 443, 578, 797, 859, 1104, 1378, 1383, 1717, 2036, 2123, 2147, 2316, 2527, 2546]. cases [234]. cash [248]. Catchers [675]. categorization [1917]. category [2036]. Causal [1132, 3139]. cause [1402, 2580]. CBF [1719]. CBNet [1153]. CC [1200]. CCI [1736]. CDN [975, 1294]. ceiling [334]. Cell [5, 128, 211, 266, 413, 415, 888, 1400, 1408, 1476, 1739, 1844, 2098, 2128, 2161, 2351, 2353, 2840, 2930, 3019, 3086]. cell-edge [2930]. cell-free [2353]. cell-less [3086]. CellOS [301]. cells [562, 1019]. Cellular [9, 29, 64, 102, 134, 188, 265, 301, 358, 359, 367, 416, 424, 440, 462, 482, 515, 579, 773, 828, 851, 1048, 1155, 1189, 1280, 1281, 1292, 1306, 1395, 1500, 1506, 1512, 1663, 1707, 1886, 1937, 1957, 2012, 2345, 2351, 2364, 2459, 2764, 2833, 2964, 3150]. cellular-connected [367, 2459, 2964]. cellular/WiFi [773]. Censorship [2360]. Center [118, 187, 272, 307, 474, 483, 644, 760, 1086, 1217, 1233, 1433, 1723, 1799, 1952, 2026, 2097, 2990]. centers [319, 1507]. central [1260]. Centrality [800, 2022]. Centralized [823, 1537, 1708]. Centric [5, 102, 119, 294, 462, 539, 584, 608, 939, 1149, 1787, 1998, 2153, 2335, 2370, 2384, 2534, 2833, 3086]. Certificate [604, 2722]. Certificate-based [2722]. Certificateless [191, 452, 833, 1103, 1252, 1790, 2530, 2575, 2721, 3006, 3051]. Certifying [1460, 2889]. CETAnalytics [208]. CFmMIMOs [1631]. CFR [2794]. CGAN [2342]. CGAN-based [2342].

Chain [15, 90, 190, 250, 316, 355, 510, 692, 900, 919, 950, 1106, 1150, 1258, 1285, 1314, 1368, 1426, 1457, 1459, 1469, 1998, 2014, 2026, 2406, 2426, 2647, 2691, 2720, 2869]. chain-based [2691]. Chain-of-Thought [2720]. chained [329, 1665]. Chaining [240, 291, 378, 676, 844, 1051, 1114, 1121, 1126, 1161, 1214, 1230, 1417, 1846, 1982, 2631, 2826]. Chains [782, 873, 1200, 1268, 1430, 1484, 1515, 1703, 1795, 2053, 2211, 2304, 2864]. Challenges [32, 58, 81, 193, 276, 367, 368, 399, 492, 646, 704, 788, 802, 941, 958, 1022, 1120, 1124, 1142, 1209, 1235, 1241, 1298, 1314, 1372, 1373, 1417, 1456, 1483, 1488, 1496, 1630, 1715, 1716, 1735, 1802, 1815, 1829, 1882, 1938, 2027, 2309, 2359, 2384, 2389, 2428, 2453, 2477, 2484, 2564, 2651, 2653, 2741, 2757, 2818, 2871, 2946, 3050, 3095, 3147]. challenging [2804]. change [2272]. changing [1640]. Channel [65, 97, 109, 111, 121, 142, 302, 439, 456, 554, 579, 698, 712, 719, 758, 805, 807, 900, 902, 984, 1041, 1049, 1065, 1173, 1225, 1239, 1338, 1539, 1653, 1682, 1690, 1739, 1757, 1835, 1863, 1937, 1951, 2003, 2034, 2187, 2264, 2277, 2519, 2632, 2647, 2663, 2755, 2854, 2856, 2873, 2914, 3051, 3089, 3133, 3177]. channel-aware [1757]. channeled [1427]. channels [281, 407, 607, 649, 712, 871, 1351, 1736, 1835, 2034, 2055, 2602]. chaos [203]. chaotic [2176]. characteristics [1861, 3148]. Characterization [443, 877, 1665, 2207, 2343]. Characterizing [343]. Charge [2046]. charger [103]. Charging [897, 1102, 1212, 2314, 2544, 2673, 2727]. chat [3066]. ChatGPT [3066]. cheating [2828]. Chebyshev [2176]. checking [2882]. checking-based [2882]. chemical [739]. China [85]. Chinese [2827]. chip [937, 1971, 2607, 2797]. chips [2151]. chit [3066]. chit-chat [3066]. choices [115]. Choose [297, 767]. CI GRU [1800]. Cipher [291]. ciphertext [2265, 2710]. ciphertext-policy [2265]. ciphertext-selected [2710]. CIPO [2110]. circles [552]. circuit [760, 1987]. circular [1373]. circulation [913]. Citadel [609]. Cities [209, 2636, 2877]. City [487, 721, 831, 1141, 1189, 1383, 2468, 2529]. city-scale [1189]. CL [3061]. Claim [856]. clarification [661]. class [57, 630, 1181, 1810, 2047, 2120, 2838, 3013]. classes [918, 1117]. classical [2458]. Classification [31, 117, 152, 158, 208, 262, 347, 354, 378, 430, 538, 589, 711, 737, 832, 843, 872, 921, 967, 1117, 1150, 1279, 1282, 1367, 1369, 1385, 1421, 1493, 1504, 1529, 1533, 1599, 1786, 1800, 1810, 1811, 1893, 1899, 1915, 1917, 1948, 2061, 2062, 2118, 2137, 2190, 2286, 2287, 2348, 2350, 2411, 2424, 2425, 2460, 2482, 2492, 2618, 2651, 2674, 2720, 2781, 2892, 2903, 2918, 2994, 3013, 3022, 3064, 3078, 3094]. classifier [182, 957, 1150, 1205, 1303, 2281]. classifiers [2683]. classify [163]. Classifying [1322, 2601]. clean [1360]. clean-slate [1360]. Client [18, 328, 1065, 1125, 1895, 2070, 2237, 2245, 2344, 2589, 2843, 3036, 3067]. client-based [18]. client-server [328]. clients [590, 1756, 2851]. clipping [963, 2768]. clique [2667]. CLLS [2530]. clock [10, 757, 2393]. clogging [475]. clone [964]. cloning [561]. closed [1444]. CLoSER [1844]. Cloud [7, 60, 92, 118, 138, 169, 211, 215, 228, 231, 236, 264, 274, 279, 282, 320, 331, 339, 353, 365, 371, 434, 437, 443, 498, 516, 519, 557, 690, 696, 708, 739, 795, 801, 814, 869, 889, 891, 892, 899, 932, 977, 1014, 1121, 1132, 1138, 1160, 1205, 1252, 1287, 1307, 1339, 1375, 1432, 1467, 1501, 1503, 1507, 1534, 1542, 1594, 1601, 1659, 1672, 1728, 1732, 1742, 1763, 1785, 1840, 1908, 1922, 1994, 2013, 2038, 2108, 2200, 2206, 2244, 2258, 2265, 2271, 2283, 2305, 2330, 2367, 2395, 2443, 2448, 2676, 2722, 2749, 2777, 2787, 2837, 2846, 2855, 2872, 2920, 2953, 2972, 2982, 3063, 3067, 3083, 3182]. Cloud [197, 2074]. cloud-assisted [1503, 2443, 2722]. Cloud-Based [519, 1339]. cloud-edge [1132, 1467, 1594, 1732, 1763, 2108, 2777]. cloud-edge-end [1742]. cloud-edge-fog [2283]. Cloud-Enabled [437]. cloud-Fog [279, 2013]. cloud-native [1994, 2271, 2920]. Cloud-RAN [1728]. Cloud-RANs [138, 2074]. cloud-to-things [795]. cloud-to-user [443]. clouded [2671]. Cloudlets [1035]. CloudPilot [1501]. clouds [335, 942, 1218, 1528, 1713, 2117, 2228, 2327, 2735, 2828]. CLS [2575]. Cluster [2, 102, 459, 498, 508, 714, 966, 1412, 1628, 2206, 2294, 2321]. cluster-aided [966]. Cluster-based [2, 508, 2321]. Clustered [373, 559, 726, 1338, 1384, 1714, 1957, 2092, 2899]. Clustering [17, 37, 207, 254, 273, 298, 365, 502, 868, 993, 1037, 1038, 1129, 1319, 1340, 1476, 1781, 2438, 2560, 2830, 3170]. Clustering-based [993, 3170]. clusters [2731, 2949]. CM [2256]. CN [114]. CNCAA [135]. CNN [55, 57, 881, 1186, 1958, 2061, 2959]. CNN-based [55]. CNN-DAG-Editor [2959]. CNN-LSTM [1958]. CNN-SSDI [881]. CNNParted [1621]. CNNs [1006]. CNP [2179]. co [2060, 2671]. co-resident [2671]. coalitional [1242]. CoAP [771]. Cobalt [1921]. code [1247, 1283, 1591, 2148, 2363]. code-based [1283]. Coded [31, 722, 888, 966, 1012, 1384, 2092, 2396, 3087]. Codes [748, 2034, 2358, 2914]. Coding [99, 171, 192, 511, 713, 1436, 2428]. coefficient [2560]. Coercion [179]. Coercion-free [179]. coevolutionary [2804]. coexistence [698, 1425, 1637, 2000, 2562, 2771]. Coffee [3018]. coflow [2166, 2373]. coflows [441, 1199]. Cognisseum [2487]. Cognition [1633, 1722]. Cognitive [2, 43, 79, 97, 249, 308, 310, 392, 435, 439, 553, 554, 701, 736, 933, 969, 1048, 1173, 1178, 1338, 1511, 1736, 2213, 2487, 2677, 2845, 2873, 2875]. Collaboration [1013, 1132, 1390, 1568, 1594, 1633, 1727, 1742, 1763, 2658, 2836, 2869]. collaboration-driven [1390].

Collaborative [141, 192, 643, 650, 695, 715, 778, 892, 1094, 1159, 1186, 1318, 1525, 1655, 1657, 1873, 1910, 1931, 1962, 2013, 2016, 2023, 2063, 2065, 2080, 2226, 2296, 2312, 2314, 2328, 2437, 2492, 2522, 2523, 2762, 2774, 2777, 2846, 2909, 3000, 3054, 3096, 3105, 3115]. Collection [41, 120, 425, 707, 831, 1036, 1409, 1410, 1412, 1413, 1455, 2162, 2275, 2636, 2709, 2745, 2783, 2785, 2918, 3101]. collective [993]. collector [1586]. collided [135]. Collision [76, 131, 135, 485, 821, 1020, 1386, 1780, 1830, 2440, 2648, 2808, 3161]. collision-aware [821]. Collision-free [1020, 1780]. colluding [1480, 2249]. collusion [1877, 2721]. collusion-resistance [2721]. Colony [405, 478, 1784, 1876, 2234, 2293]. Colony-inspired [1784]. Colosseum [2487]. column [729]. CoMap [1242]. combination [151, 2438, 3099]. combinatorial [1640, 2015, 2245]. Combine [2137]. combined [564, 1269, 2065, 2847]. Combining [1182, 1295, 1328, 1633, 2404, 2729, 2958]. COMEC [2930]. commerce [411]. commercial [600, 686, 803]. commercial-off-the-shelf [600]. commit [2214]. commit-and-prove [2214]. commitments [2750]. commodity [261, 2058, 2177]. Common [347, 873, 1729]. commons [2442]. commons-based [2442]. Communication [38, 52, 111, 139, 140, 150, 245, 265, 350, 356, 377, 388, 405, 440, 536, 577, 590, 619, 620, 629, 653, 663, 686, 758, 768, 769, 791, 824, 902, 937, 992, 1064, 1066, 1087, 1101, 1145, 1247, 1259, 1369, 1378, 1386, 1512, 1605, 1624, 1626, 1627, 1637, 1675, 1681, 1690, 1718, 1750, 1759, 1775, 1779, 1790, 1863, 1895, 1896, 1903, 1911, 1971, 1997, 2008, 2017, 2131, 2142, 2189, 2237, 2238, 2345, 2363, 2368, 2377, 2407, 2458, 2506, 2513, 2529, 2549, 2554, 2587, 2597, 2605, 2607, 2650, 2672, 2705, 2746, 2754, 2764, 2797, 2825, 2940, 2951, 3006, 3038, 3109, 3117, 3138, 3166, 3185]. communication-channel-based [758]. Communication-efficient [824, 1895, 2554, 2705]. Communications [3, 76, 116, 159, 163, 174, 212, 288, 410, 457, 575, 736, 837, 852, 859, 935, 1099, 1195, 1235, 1241, 1243, 1280, 1291, 1527, 1580, 1609, 1733, 1778, 1822, 1867, 1884, 2006, 2012, 2036, 2037, 2039, 2046, 2113, 2181, 2218, 2229, 2231, 2279, 2429, 2470, 2584, 2711, 2712, 2871, 3012, 3056, 3079, 3095, 3115, 3136, 3156]. communities [75]. community [68, 153, 1294, 1743, 2127, 2442, 2679]. community-based [1294]. CoMP [67, 1873, 2421]. CoMP-Based [67]. compact [2320]. Comparative [210, 430, 532, 1033, 1170, 1777, 1802, 1814, 1971, 2044, 2178, 2264, 2600, 2633, 2663, 2830, 3043, 3091]. Comparison [58, 273, 408, 618, 2034, 2462, 2562]. comparisons [1384]. Compatible [898, 1144, 1619]. compensated [10]. Compensation [67, 1382]. competition [1206, 2485]. Competitive [304, 567, 1156, 1562, 2804, 2959]. compilation [130, 2363]. compiler [2019]. completion [343, 1085, 2166, 2334, 2806]. Complex [4, 13, 75, 572, 913, 1598, 1605, 1774, 1827, 1978, 2252, 2558, 2728]. Complex-network [2558]. complex-policy [4]. Complexity [1211, 1646, 1875, 2226]. compliant [939, 2574, 2637]. COMPNW [1571, 1764]. Component [1350, 2533]. Component-based [2533]. Component-wise [1350]. components [234, 846]. composing [2508]. Composite [541, 2823]. composition [16, 21, 1090, 3184]. compound [2598]. Comprehensive [161, 208, 287, 615, 682, 760, 761, 827, 828, 1017, 1054, 1066, 1124, 1146, 1154, 1347, 1372, 1483, 1715, 1771, 1776, 1788, 1866, 1882, 1934, 1939, 1965, 2056, 2100, 2117, 2183, 2184, 2344, 2389, 2477, 2651, 2702, 2784, 2821, 2858, 2892, 2946, 2991, 3044, 3094, 3136]. compressed [51, 2705]. compression [1427, 2190, 2716, 2767, 3134]. Compressive [54, 3089]. compromise [3173]. Computation [223, 257, 387, 389, 558, 733, 742, 830, 945, 956, 981, 1190, 1237, 1381, 1470, 1532, 1655, 1910, 1936, 1954, 2108, 2141, 2149, 2235, 2270, 2314, 2340, 2538, 2604, 2610, 2622, 2635, 2850, 2899, 2971, 3105, 3154, 3181]. computational [787, 959]. computationally [69]. computations [2854]. compute [557, 2356, 2916]. compute-aggregate [557]. Computer [271, 357, 527, 996, 1598, 1648, 1759, 1893, 2813, 2958, 3169, 3186]. computers [1195]. Computing [16, 145, 209, 343, 359, 401, 475, 557, 559, 574, 584, 611, 642, 667, 677, 696, 697, 703, 706, 781, 787, 789, 814, 823, 839, 840, 956, 976, 979, 1003, 1015, 1064, 1084, 1085, 1100, 1135, 1179, 1186, 1190, 1205, 1209, 1227, 1274, 1287, 1289, 1340, 1366, 1370, 1401, 1455, 1470, 1475, 1484, 1496, 1552, 1555, 1559, 1582, 1623, 1678, 1683, 1705, 1710, 1731, 1758, 1816, 1906, 1910, 1930, 1938, 2016, 2107, 2108, 2113, 2144, 2155, 2198, 2201, 2242, 2294, 2296, 2310, 2331, 2356, 2394, 2415, 2438, 2461, 2510, 2518, 2522, 2572, 2590, 2640, 2667, 2805, 2870, 2910, 2917, 2930, 2935, 2953, 3011, 3067, 3095, 3096]. computing [212, 257, 366, 387, 395, 398, 404, 434, 436, 516, 549, 558, 650, 680, 690, 702, 733, 784, 785, 795, 824, 862, 892, 955, 959, 981, 1032, 1040, 1059, 1060, 1066, 1087, 1106, 1118, 1216, 1238, 1248, 1312, 1318, 1337, 1357, 1359, 1381, 1394, 1406, 1423, 1429, 1432, 1472, 1473, 1479, 1525, 1568, 1654, 1687, 1728, 1732, 1760, 1805, 1833, 1856, 1975, 2017, 2021, 2040, 2051, 2073, 2076, 2142, 2200, 2262, 2288, 2308, 2312, 2324, 2340, 2365, 2367, 2466, 2467, 2475, 2492, 2604, 2605, 2626, 2643, 2679, 2684–2686, 2726, 2762, 2799, 2820, 2835, 2840, 2861, 2872, 2937, 2962, 2964, 2985, 3030, 3043, 3061, 3144, 3172, 3183]. Computing [2270]. computing-based [642]. computing-enabled [2626]. Computing-endowed [2324]. CoMSeC [845]. concatenated [1247]. concave [292]. concept [378, 973, 2250]. Concepts [1850]. conceptual [2112]. Concordit [2520]. Concurrent [14, 73, 302, 686, 1591, 2664, 2672]. Condition [65, 2205, 2593]. conditional [2721]. conditions [902]. Conference [1647, 3012]. Conferencing [1745]. Confidence [957, 3005]. confidential [2490]. configurable [1771]. Configuration [327, 865, 1603, 1936, 2048, 2337]. configurations [238]. confirmation [127]. conflict [1374, 3128]. confusable [1158]. Congestion [196, 447, 482, 618, 866, 1029, 1094, 1114, 1174, 1388, 1664, 1698, 1707, 1708, 1823, 1853, 1870, 1928, 1964, 2111, 2128, 2147, 2179, 2182, 2272, 2534, 2537, 2624, 2698, 2714, 2957, 2986, 3113]. Congestion-aware [1698, 2128, 2147, 2714]. congestion-induced [1964]. Conjunctive [1453]. Connect [2277]. Connected [263, 313, 367, 1108, 1200, 1232, 1329, 2459, 2526, 2732, 2956, 2964]. Connecting [1930]. Connection [342, 1656, 2708]. connections [2256]. Connectivity [433, 684, 738, 1039, 1074, 1081, 1147, 1358, 1363, 1386, 1578, 1684, 2197, 2206, 2547, 2563, 3041, 3163]. ConPoolUBF [1656]. conquer [807]. consensus [285, 507, 825, 826, 873, 905, 953, 1042, 1172, 1288, 1454, 1740, 1956, 2096, 2295, 2333, 2733, 2885, 2977, 2995, 3055, 3100]. consequences [189]. conservation [2462]. Conservative [1276]. consideration [2925]. considering [150, 484, 1620, 1894, 1935, 2655]. Consistency [158, 1132, 1430, 1610, 2018, 2171, 2681]. Consistent [659, 2650]. consolidation [169, 231, 1738, 1885]. Consortium [1454, 2031]. Constant [924, 3029]. constellation [1013, 1687, 1911, 2437, 2657]. constellations [1535, 2730]. constrained [10, 124, 320, 398, 824, 889, 1001, 1029, 1116, 1359, 1366, 1386, 1393, 1398, 1527, 1948, 2144, 2323, 2327, 2700, 2763, 2972, 2993, 3005]. constraint [2245, 2375, 2977, 2992]. constraints [261, 441, 567, 1081, 1466, 1795, 2302, 2319]. constricted [1970]. Construction [714, 994, 1608, 2486]. consumer [539, 1775]. consumption [278, 693, 1040, 1258, 1404, 1436, 1537, 1692, 1728, 1862, 2081, 2117, 2283, 2507, 2541, 2552, 2573, 3039, 3156, 3187]. consumption-oriented [3187]. Contact [645]. Containerized [1936, 2117, 2427, 2932]. containers [3174]. containment [1697]. contemporary [467, 527]. Content [91, 225, 248, 294, 331, 498, 514, 742, 836, 865, 887, 895, 976, 1063, 1068, 1117, 1328, 1510, 1645, 1672, 1679, 1700, 1716, 1844, 1874, 1897, 1945, 2161, 2267, 2898]. Content-aware [865]. content-based [1328, 1679, 2267]. Contention [901, 1448, 1719, 2385, 2949]. contention-based [1719]. contention-free [901]. contents [426]. Context [20, 210, 655, 1298]. Context-aware [210]. contexts [1474]. Contextual [1369, 2504, 2893]. Continuity [825, 2354]. Continuous [774, 863, 2247, 2681, 3078]. continuous-time [2247]. continuum [2038, 2676]. contract [1269, 1516, 1646, 1923, 2011, 2232, 2257, 2753]. contract-based [1923, 2232]. ContractBox [1622]. Contractive [2770]. contracts [1104, 2603, 3157]. contrast [2788, 3007]. Contrastive [1961, 2286, 2753, 3099, 3153]. ConTrib [631]. Control [19, 112, 116, 196, 209, 342, 385, 482, 494, 504, 508, 542, 543, 550, 553, 571, 573, 604, 618, 649, 696, 734, 850, 866, 880, 914, 1029, 1094, 1125, 1133, 1174, 1230, 1286, 1309, 1353, 1361, 1388, 1486, 1522, 1593, 1601, 1664, 1701, 1707, 1802, 1823, 1853, 1855, 1858, 1870, 1886, 1924, 1928, 2003, 2022, 2043, 2152, 2179, 2182, 2233, 2252, 2272, 2322, 2348, 2358, 2402, 2443, 2455, 2537, 2549, 2596, 2624, 2634, 2698, 2936, 2938, 2969, 2986, 3016, 3035, 3043, 3097, 3113, 3178]. control-theory-based [1309]. controlled [1130]. controller [222, 325, 476, 623, 642, 763, 764, 797, 975, 1037, 1158, 1701, 1732, 1909, 2010, 2271, 2293, 2784]. controllers [473, 489, 601, 793, 1346, 1927, 2474, 2970]. controlling [1632]. controls [2688]. conventional [461]. convergence [725, 1959, 3170]. converging [1956]. conversations [2905]. conversion [724]. convex [1256]. Convolution [881, 1504, 2350]. Convolutional [117, 235, 247, 487, 669, 934, 1320, 1621, 1770, 2226, 2351, 2772, 2959]. cooperated [2453]. Cooperation [2, 5, 862, 888, 933, 2448, 2590, 3181]. Cooperation-based [2590]. Cooperative [12, 108, 159, 252, 274, 454, 558, 645, 770, 837, 1035, 1061, 1223, 1272, 1414, 1420, 1455, 1472, 1480, 1512, 1733, 1832, 1836, 1858, 1876, 1897, 1955, 2020, 2055, 2129, 2145, 2155, 2196, 2336, 2387, 2393, 2495, 2628, 2632, 2689, 2725, 2826, 2930, 3159]. CoopeRIS [2196]. CoopFL [1406]. coordinate [2814]. Coordinated [5, 1260, 1308, 1447, 1933, 2621, 2695]. Coordinates [2915]. coordinating [2495]. coordination [5, 507, 717, 756, 1680, 1830, 1940, 2907]. cope [1756]. copper [172]. copper-to-fiber [172]. Coral [2319]. Core [1259, 1459, 2032, 2081, 2555, 2895, 3027]. Core-selecting [1459]. COREC [2664]. corpus [3123]. Corr [3007]. correct [767]. Correcting [748, 2914]. correction [1247, 2034, 2736]. Correlated [1990, 2516]. Correlation [99, 920, 1602, 1723, 2497, 2540, 2581, 2830, 3007]. correlation-based [2581, 2830]. correlations [113, 598, 2641]. Corridor [268]. Corrigendum [271, 357, 527, 1571, 1759, 1764, 1893]. Cos [3134]. cosine [1164, 2420]. COSMOS [820, 1749]. Cost [53, 137, 330, 331, 473, 695, 711, 764, 850, 974, 1003, 1064, 1126, 1300, 1562, 1604, 1667, 1672, 1734, 1949, 1968, 1976, 2005, 2021, 2026, 2047, 2161, 2170, 2228, 2237, 2399, 2626, 2631, 2700, 2735, 2953, 2979, 3018, 3144]. cost-aware [695, 1064, 1126, 1949, 1968, 2237]. Cost-effective [2953, 3018]. Cost-efficient [2005, 2228, 2700, 2735]. cost-energy-QoE [137]. cost-latency [1003]. cost-provable [2626]. COTS [1]. Count [1276, 1684]. counter [524]. countermeasure [1180, 2289]. countermeasures [596, 1639, 2378]. counters [1369]. counting [1328]. coupled [2183]. Coupling [169, 1884]. Coupon [1606]. covariance [1377]. cover [679, 1347]. Coverage [186, 462, 810, 1073, 1074, 1081, 1290, 1411, 1548, 1562, 1592, 1600, 1691, 1784, 1866, 1920, 2728, 2757, 2868]. coverage-based [1290]. coverage-centric [462]. Covert [407, 607, 649, 992, 1668, 1682, 2249, 2914]. COVID [217, 386, 1369, 1496]. COVID-19 [217, 386, 1369, 1496]. CP [2443]. CP-ABE [2443]. CPA [3160]. CPN [2572]. CPS [1668]. CPU [1665]. CPU-bound [1665]. CR [770, 1686, 2500]. CR-IoT [2500]. CR-NOMA [1686]. CRAN [2106]. CRC [1247]. CRC-aided [1247]. creation [3066]. Credential [1018, 1883]. credentials [2328, 2924]. credibility [661]. credit [644, 1174, 1962, 2066, 2520]. credit-based [1174, 2520]. criteria [234, 457]. Critic [2361, 2549, 2794, 3164]. critical [358, 1138, 2154, 2181, 2299, 2532, 2794]. CRLM [1832]. CRN [2129]. CRNs [705]. Cross [11, 19, 309, 371, 510, 617, 1069, 1314, 1377, 1477, 1546, 1570, 1680, 1730, 1825, 1855, 1917, 2086, 2277, 2498, 2535, 2599, 2600, 2607, 2644, 2682, 2714, 2748, 2797, 2859, 2869, 2896, 2925, 3003, 3015, 3026, 3083, 3138]. cross-chain [1314, 2869]. Cross-Connect [2277]. Cross-domain [510, 1377, 1570, 1730, 1855, 2535, 2748, 2859, 2896, 3003, 3083]. cross-ISP [1477]. Cross-layer [371, 617, 1069, 1546, 1825, 2086, 2498, 2600, 2644, 2714, 3138]. Cross-layering [2925]. cross-link [1680]. cross-platform [3026]. cross-producting [1917]. Cross-Silo [3015]. Cross-site [11]. cross-task [309]. cross-traffic [2599]. crosschain [852]. crossing [1239]. Crosstalk [346, 1434, 1894, 2028, 2234]. Crosstalk- [2234]. crosstalk-avoided [2028]. crosstalk-aware [1434, 2028]. crosstalk-derived [1894]. Crowd [638, 1102, 1232, 1455, 1744, 1747, 2199, 2317, 2820, 3158]. Crowd-assisted [2317]. crowd-sourced [1232, 1747]. crowded [64]. Crowds [2915]. Crowdsensing [122, 133, 406, 1152, 1443, 1633, 1722, 1859, 2132, 2313, 2369, 2614, 2746, 2806, 2868, 2901, 3090, 3099, 3185]. crowdsourced [248, 1189, 1208]. Crowdsourcing [73, 119, 176, 409, 444, 718, 1216, 1491, 1711, 1974, 2452, 2888, 3070]. CRQF [2174]. cruising [815]. crypto [608]. crypto-ransomware [608]. cryptocurrency [1363, 3073]. cryptographic [2410, 3006]. cryptographically [503]. cryptosystem [328]. CSI [521, 594, 755, 1480, 1733, 2345, 2407, 2589, 2824, 3069, 3154]. CSMA [524, 726, 732, 1244]. CSMA/CA [524, 726]. CTAP [786]. CTC [2978]. CTGAN [3076]. CTI [1747]. CTS [900]. CUBIC [1500]. Cuckoo [1552, 1892]. CUPS [459]. CUPS-based [459]. Current [646, 802, 1580, 1630, 1639, 2453]. CUSTOM [260]. Customized [401]. CUSUM [1301]. CWT [1350]. CWT-DPA [1350]. CYBELE [56]. Cyber [82, 464, 532, 609, 662, 1041, 1050, 1316, 1319, 1495, 1547, 1602, 1638, 1819, 2174, 2208, 2342, 2374, 2453, 2504, 2789, 2800, 2945]. cyber-attacks [2800]. Cyber-Physical [464, 1041, 1316, 1319, 1547, 2208]. cyber-physical-social [2453]. Cyber-security [82, 662]. Cyberattacks [2929]. cybercrime [1747]. cybersecurity [855, 1120, 1857, 2378, 2918]. Cybersecurity-vs-QoS [1857]. cyberspace [3124]. cycle [417, 1062, 1546]. cycles [2113]. cycling [2048].

D [583, 600, 937, 1147, 1282, 1340, 1558, 1663, 1680, 1690, 1971, 1979, 2205, 3188]. D-AdFeed [583]. D-PhishNet [3188]. D-TDD [1680]. D2D [9, 76, 79, 137, 188, 245, 265, 405, 416, 440, 663, 695, 1064, 1087, 1089, 1099, 1101, 1280, 1281, 1419, 1450, 1512, 1548, 1616, 1867, 2249, 2279, 2294, 2345, 2764, 3020, 3056]. D2D-assisted [416, 2294]. D2D-based [2764]. D2D-enabled [9, 188, 1064]. D2DLive [1616]. D3QN [2837]. D3QN-based [2837]. DA-GNN [2011]. DACP [1855]. DADCNF [1453]. DAG [789, 889, 1803, 2327, 2959]. DAG-SFCs [1803]. damages [113]. Danger [835]. DAP [683]. DAP-Sketch [683]. dApps [2969]. dark [1054]. Darknet [2227]. DARVAN [1468]. DASH [99, 1782]. Data [8, 47, 120, 156, 262, 278, 332, 337, 341, 411, 426, 451, 453, 483, 556, 561, 602, 608, 644, 703, 706, 707, 720, 726, 750, 773, 829, 831, 890, 910, 935, 944, 968, 1025, 1055, 1086, 1105, 1115, 1277, 1349, 1355, 1409, 1410, 1431, 1455, 1462, 1499, 1507, 1516, 1522, 1533, 1559, 1606, 1622, 1632, 1643, 1690, 1771, 1773, 1775, 1804, 1878, 1881, 1897, 1941, 1990, 2097, 2114, 2119, 2146, 2281, 2298, 2351, 2366, 2394, 2398, 2419, 2446, 2470, 2513, 2516, 2579, 2595, 2628, 2670, 2676, 2687, 2709, 2736, 2749, 2760, 2806, 2836, 2870, 2889, 2923, 2956, 2978, 3052, 3071, 3075, 3096, 3173]. data [9, 22, 29, 41, 56, 104, 118, 154, 155, 187, 192, 241, 261, 271, 272, 307, 319, 342, 463, 468, 469, 474, 502, 626, 650, 694, 708, 760, 765, 842, 858, 896, 942, 995, 1036, 1070, 1071, 1135, 1141, 1149, 1167, 1196, 1217, 1230, 1233, 1350, 1412, 1413, 1436, 1443, 1446, 1451, 1452, 1529, 1572, 1626, 1656, 1663, 1723, 1738, 1774, 1779, 1796, 1799, 1840, 1851, 1952, 1973, 1998, 2010, 2026, 2029, 2068, 2132, 2153, 2160, 2162, 2199, 2278, 2280, 2339, 2357, 2358, 2360, 2399, 2451, 2491, 2535, 2536, 2636, 2668, 2748, 2783, 2809, 2856, 2869, 2876, 2885, 2946, 2958, 2990, 3030, 3046, 3083, 3101, 3126, 3130, 3141, 3155, 3175]. data [62, 168, 173, 227, 366, 423, 932, 966, 1312, 1467, 1865, 1868, 2215, 2528, 2711, 2716, 3070]. data-aware [1663]. data-blocks [842]. Data-centric [608]. Data-driven [154, 271, 750, 2360, 2687]. data-efficient [910]. data-power [3071]. data-scarce [3130]. data-sharing [1105]. data-smart [703]. databases [1022]. Datacenter [143, 950, 1153, 1199, 1798, 1823, 2182, 2432, 2539, 2598, 2616, 2986, 3035]. datacenters [93, 344, 739, 1174, 1285, 1946, 2111, 2624]. dataplanes [429]. Dataset [183, 235, 640, 751, 832, 920, 996, 998, 1234, 1421, 1560, 1567, 1660, 2075, 2103, 2131, 2392, 2918]. Datasets [321, 394, 532, 718, 1747, 2316, 2557, 2674, 2796, 3171]. Day [312, 2203, 2569, 2883, 2960]. DBSRP [2980]. DBSRP-ML [2980]. DBUP [2423]. DC [1750]. DCCS [2182]. DCDTS [3003]. DCs [2750]. DDoS [497, 499, 541, 816, 857, 1060, 1080, 1097, 1180, 1205, 1250, 1301, 1525, 1715, 1828, 1838, 1947, 2023, 2083, 2243, 2463, 2781, 2943, 3037, 3167]. DDoS-MQTT-IoT [1660]. DDoSBERT [2781]. DDoSBlocker [2718]. DDPS [2642]. DDQN [2426]. DDQN-SFCAG [2426]. DDS [2765]. DE-GNN [2118]. de-randomization [998, 1328]. deadline [1199, 1798, 2319]. deadline-aware [1199]. deadlines [174, 2021]. debris [1880]. December [24, 397, 422, 879, 912, 1343, 1380, 1905, 2512, 2566]. decentralization [826, 2088, 2115, 2634]. decentralize [1738]. Decentralized [302, 507, 631, 715, 823, 885, 901, 938, 1034, 1090, 1144, 1152, 1199, 1468, 1507, 1645, 1646, 1708, 1812, 1836, 1854, 1931, 2382, 2733, 2780, 2798, 2841, 2842, 2981, 3055, 3126, 3170]. deception [2342, 2525]. deceptive [2900]. Decision [554, 1514, 1552, 1593, 1963, 3086]. Decision-Making [1552, 1593, 3086]. DEcisions [429, 2799]. decoder [2854]. decoding [176]. decomposition [35, 860, 1734, 2074, 3002]. deconfounded [2570]. decoupled [1142, 1479, 2435, 2451]. DECS [650]. Dedicated [1202, 2500]. deduplication [2132, 2722]. Deep [61, 116, 163, 290, 291, 311, 370, 385, 477, 495, 564, 571, 605, 627, 697, 718, 744, 765, 866, 870, 880, 896, 926, 944, 957, 1021, 1032, 1037, 1060, 1120, 1194, 1215, 1237, 1249, 1281, 1282, 1309, 1310, 1320, 1321, 1369, 1372, 1389, 1470, 1475, 1505, 1511, 1529, 1557, 1574, 1612, 1623, 1631, 1670, 1677, 1700, 1707, 1723, 1753, 1805, 1829, 1837, 1953, 1976, 2023, 2037, 2052, 2082, 2084, 2087, 2106, 2132, 2150, 2201, 2209, 2231, 2235, 2262, 2270, 2337, 2381, 2387, 2470, 2502, 2561, 2597, 2622, 2640, 2700, 2709, 2746, 2777, 2786, 2850, 2938, 2950, 2957, 2961, 2987, 3011, 3030, 3047, 3084, 3152, 3160, 3182, 3185]. deep [272, 403, 465, 587, 640, 643, 730, 737, 755, 801, 830, 959, 967, 981, 1047, 1061, 1086, 1088, 1176, 1202, 1206, 1265, 1292, 1408, 1515, 1525, 1661, 1732, 1739, 1799, 1856, 1879, 1947, 1995, 2006, 2151, 2159, 2183, 2274, 2288, 2291, 2385, 2459, 2537, 2559, 2605, 2632, 2641, 2761, 2770, 2799, 2864, 2872, 2873, 2942, 2964, 3049, 3053, 3172]. deep-reinforcement-learning [3182]. deep-space [2537]. DeepLoad [436]. deeply [758]. DeepMetricCorr [1723]. Defeating [2824]. defence [2668, 2789]. Defending [1158, 1753]. Defense [282, 609, 1050, 1060, 1250, 1429, 1514, 1638, 1656, 1688, 1715, 2282, 2445, 2454, 2573, 2625, 2774, 2894, 2945, 2998, 3000]. defenses [112, 1053, 1981, 2929]. Defined [80, 84, 95, 164, 218, 222, 325, 356, 423, 442, 470, 476, 508, 516, 520, 543, 593, 609, 623, 624, 662, 673, 690, 745, 790, 849, 981, 1037, 1069, 1100, 1105, 1121, 1165, 1183, 1254, 1540, 1610, 1640, 1688, 1749, 1787, 1805, 1875, 1901, 1908, 1909, 1972, 1994, 2083, 2108, 2206, 2284, 2290, 2293, 2488, 2524, 2730, 2991, 3004, 3114, 3116]. Defining [62, 2697]. Definition [2330]. DeFlow [2616]. Defragmentation [237]. degradation [66]. degree [598, 826, 986, 2317, 3018]. DEHO [1205]. Delay [10, 73, 150, 304, 320, 349, 385, 389, 398, 482, 513, 525, 552, 642, 763, 781, 917, 956, 1011, 1016, 1077, 1078, 1116, 1211, 1306, 1318, 1359, 1366, 1397, 1494, 1500, 1527, 1591, 1698, 1754, 1782, 1818, 1841, 1968, 2053, 2063, 2086, 2219, 2299, 2348, 2406, 2681, 2693, 2802, 2840, 2862, 2975, 2992, 3125, 3129]. Delay-aware [385, 781, 2063, 2348, 2992]. delay-based [1011, 1782, 3125]. Delay-constrained [320, 1116, 1359, 1366, 1527]. Delay-differences [2681]. Delay-energy-aware [2840]. delay-guaranteed [389, 1698]. delay-sensitive [2219]. delay-tolerant [1078]. Delayed [1642, 2208]. delays [2280]. Delegation [833, 1502, 1883, 2374]. Delivering [364]. Delivery [116, 278, 358, 514, 707, 742, 965, 976, 1177, 1510, 1571, 1745, 1874, 1897, 1945, 2198, 2278, 2317, 2740, 3070, 3184]. delta [1471]. delta-based [1471]. Demand [23, 307, 346, 364, 1153, 1535, 1741, 1946, 2005, 2026, 2065, 2714, 2861, 2864, 3017]. Demand-aware [1153, 1946]. demands [969, 1652, 2977]. Democratizing [213]. Dendrimer [1646]. deniable [1914]. Denial [148, 164, 311, 516, 1080, 1251, 1444, 1456, 2599]. Denial-of-Service [164, 2599]. Dense [34, 204, 416, 562, 563, 579, 730, 893, 1332, 1381, 1400, 1739, 2262, 2385, 2430, 2930, 2981]. dense-interference [2981]. DenseNet [2381]. DenseNet-Transformer [2381]. Density [193, 1125, 1748, 2438, 2915, 2938]. Density-aware [193]. density-distance [2915]. departure [214]. dependency [2835]. Dependent [1466, 1478, 1669, 3011]. deployed [224, 1004, 1264, 1293]. Deploying [1116, 1803, 2211]. Deployment [90, 102, 110, 562, 764, 796, 814, 851, 897, 990, 1138, 1227, 1258, 1368, 1403, 1536, 1631, 1663, 1669, 1749, 1760, 1821, 1942, 2053, 2063, 2158, 2205, 2590, 2595, 2626, 2638, 2790, 2862, 2968, 2979, 3001, 3127, 3145, 3183]. deployments [1225, 1746, 1866, 2555, 2739]. depth [1355, 2168]. derived [1894]. description [691, 2442]. Design [36, 40, 84, 115, 232, 234, 258, 300, 315, 367, 417, 439, 505, 521, 580, 622, 677, 678, 701, 722, 747, 763, 792, 919, 954, 975, 985, 1029, 1069, 1083, 1138, 1173, 1254, 1442, 1453, 1461, 1492, 1573, 1636, 1652, 1695, 1724, 1727, 1749, 1801, 1825, 1835, 1847, 1887, 1972, 1986, 2049, 2238, 2284, 2405, 2510, 2587, 2607, 2658, 2691, 2719, 2776, 2789, 2816, 2913, 2964, 3028, 3077]. designated [1252]. Designing [355, 781, 1313, 2457, 2752]. designs [1012]. DESiRED [2082]. DESLR [1757]. destination [3163]. destructive [1275]. Detail [419]. detect [809, 1205, 1646, 2599]. detected [2679]. Detecting [42, 75, 531, 607, 758, 1070, 1196, 1668, 1837, 1947, 2045, 2069, 2203, 2453, 2502, 2800, 2803, 2804, 2872, 3075, 3110, 3124]. Detection [36, 61, 151, 282, 311, 312, 420, 456, 495, 497, 533, 541, 561, 608, 610, 630, 669, 723, 751, 780, 857, 881, 896, 910, 915, 926, 941, 960, 964, 978, 991, 1045, 1080, 1098, 1170, 1180, 1201, 1251, 1256, 1278, 1319, 1339, 1402, 1424, 1453, 1495, 1499, 1685, 1706, 1743, 1748, 1773, 1820, 1824, 1830, 1853, 1857, 1929, 1932, 1992, 2023, 2101, 2138, 2163, 2168, 2186, 2207, 2243, 2266, 2281, 2297, 2329, 2434, 2447, 2477, 2594, 2671, 2701, 2717, 2736, 2753, 2770, 2772, 2784, 2786, 2838, 2843, 2886, 2894, 2902, 2952, 2987, 3037, 3047, 3050, 3073, 3080, 3094, 3115, 3143, 3160, 3180, 3188]. detection [29, 57, 63, 72, 144, 409, 434, 437, 438, 442, 464, 532, 564, 572, 616, 640, 648, 681, 715, 761, 766, 801, 835, 860, 904, 934, 961, 993, 1067, 1175, 1232, 1236, 1239, 1257, 1301, 1303, 1365, 1421, 1530, 1542, 1557, 1576, 1594, 1595, 1737, 1754, 1793, 1829, 1871, 1921, 1934, 2029, 2083, 2134, 2192, 2212, 2254, 2289, 2303, 2306, 2316, 2357, 2360, 2366, 2391, 2404, 2440, 2469, 2478, 2519, 2527, 2540, 2591, 2612, 2665, 2689, 2729, 2751, 2758, 2781, 2788, 2796, 2798, 2839, 2883, 2893, 2909, 2943, 2958, 2960, 2963, 2994, 3004, 3023, 3025, 3060, 3078, 3088, 3102, 3103, 3118, 3119, 3152, 3155, 3162, 3167]. detection [74, 141, 182, 235, 317, 1188, 1269, 1374, 1525, 1570, 1762, 1941, 2011, 2291, 2823, 2900, 3179]. detector [55, 247, 864, 1702, 3110]. detectors [309, 2178]. deterioration [681]. Determining [278]. Deterministic [683, 825, 1171, 1658, 1918, 2582, 2761, 2765, 3003, 3100]. DetNet [1096]. detouring [1569]. development [1410, 2775, 2882, 2951]. deviation [1231]. Device [38, 76, 183, 595, 629, 898, 998, 1222, 1328, 1527, 1847, 1884, 1924, 1942, 2062, 2492, 3034, 3046]. device-free [3034]. Device-to-Device [38, 629, 898, 1884]. Devices [52, 90, 100, 253, 268, 490, 597, 600, 686, 728, 735, 806, 947, 961, 1041, 1119, 1133, 1283, 1692, 1825, 1834, 1970, 2072, 2121, 2253, 2630, 2890, 2972, 3010, 3124]. DeX [2231]. DFS [1626]. DG [2299]. DGA [1421, 1800, 2994]. DGPAS [3149]. DHR [2363]. DHR-based [2363]. diagnose [1390]. Diagnoser [1453]. diagnosis [791, 2183, 2236]. DID [2085]. DID-based [2085]. differences [2681]. different [594, 854, 884, 1669, 2600]. Differential [226, 855, 1505, 1514, 1990, 2111, 2162, 2266, 2374, 2491, 2516, 2616, 2768, 2935]. Differentially [1312, 2327, 2766]. differentiated [2111, 2707]. differentiating [2499]. Differentiation [118, 853, 2515]. differentiations [982]. Diffie [2400, 2906]. DiffServ [918]. DIFFUSE [1090]. diffusion [304, 2315, 2369, 2642, 2719, 2749, 2758, 3081]. diffusion-based [2719]. Digital [438, 612, 716, 775, 1284, 1373, 1490, 1654, 1815, 1989, 2036, 2100, 2330, 2343, 2370, 2583, 2699, 2738, 3046, 3079, 3130, 3132]. Digital-care [1490]. digitally [3030]. Dikaios [2995]. dimension [1080]. dimensional [754, 1255, 1569, 1592, 1806, 1878, 2090, 2549, 2642, 2814, 2883, 3141]. dimensioning [1438]. directed [2533]. direction [298, 1662]. directional [1684, 1979]. directions [577, 1425, 1490, 1580, 1777, 2389, 2428, 2651, 2652, 2741, 2793, 2871]. DIS [2403].

DIS-Guard [2403]. disaggregated [1801, 2592]. disagreement [718]. disaster [293, 319, 432, 1526]. disaster-aware [319]. DiSAuth [2451]. disciplines [115]. disclosure [2214]. Discontinuous [1162]. Discovering [566, 572, 2611]. Discovery [76, 101, 181, 253, 597, 691, 849, 1405, 1797, 1979, 2279, 2708, 3090, 3158]. discrepancy [2834]. discrepancy-aware [2834]. Discrete [2794, 2814]. Discrimination [246, 1301]. Discrimination-aware [246]. dispatch [2544]. dispersal [2189]. dispersed [2962, 3106]. disposable [431]. Disruptive [428, 2370]. Dissecting [1006]. Dissemination [22, 83, 379, 887, 914, 1246, 1632, 1836, 2792, 2978, 3072]. Distance [871, 1377, 2669, 2737, 2915, 2994]. distillation [2435, 2515, 2985]. distinct [511]. Distinguishing [906]. distribute [2030]. Distributed [46, 93, 111, 141, 252, 258, 305, 311, 319, 365, 414, 478, 498, 507, 568, 639, 744, 778, 793, 830, 834, 869, 877, 889, 927, 981, 1010, 1051, 1060, 1084, 1089, 1101, 1126, 1132, 1160, 1186, 1202, 1207, 1251, 1285, 1288, 1338, 1357, 1375, 1405, 1442, 1444, 1456, 1475, 1492, 1559, 1566, 1578, 1620, 1706, 1718, 1768, 1776, 1830, 1877, 1879, 1979, 2016, 2043, 2052, 2060, 2064, 2168, 2187, 2199, 2214, 2228, 2248, 2270, 2275, 2279, 2302, 2324, 2333, 2415, 2468, 2473, 2549, 2553, 2606, 2621, 2646, 2719, 2769, 2814, 2817, 2834, 2907, 2949, 2970, 3041, 3071, 3072, 3149, 3157, 3168, 3180]. distributed [27, 285, 637, 885, 966, 1028, 1080, 1090, 1525, 1626, 1664, 2125, 2596, 2695]. distribution [191, 204, 498, 680, 805, 1019, 1181, 1672, 1865, 1993, 2075, 2399, 2474, 2899, 2971]. distributions [986]. disturbed [2497]. DIVA [2085]. dive [1176, 2933]. Diver [2131]. Diver-robot [2131]. diverse [502, 1692, 2137, 2588]. Diversity [142, 174, 198, 426, 583, 1268, 2032, 2199, 2690]. diversity-aware [583]. Diversity-improved [426]. divide [807]. divide-and-conquer [807]. Divisible [1401]. division [177, 346, 1056, 1324, 1393, 1703, 2099, 2148]. DiWi [3132]. DKD [2435]. DL [880, 1382, 1835, 2943]. DL-based [1382, 1835]. DLP [2609]. DLT [1401]. DM [375]. DM-GKM [375]. DMatrix [804]. DMO [1162]. DMZ [2010]. DNA [1278]. DNN [908, 1406, 1655, 1678, 1814, 2308, 2553, 3054, 3149]. DNN-based [908]. DNS [761, 834, 1070, 1176, 1595, 1877, 2178, 2360, 2451, 2454, 2708, 2881]. DNS-based [2451]. DNSxP [694]. DOCUS [1301]. DOCUS-DDoS [1301]. Does [2060]. DoH [1127, 1737, 3088]. Domain [250, 251, 284, 431, 467, 510, 527, 538, 597, 648, 833, 1126, 1341, 1377, 1423, 1541, 1570, 1593, 1730, 1815, 1855, 1980, 2023, 2134, 2159, 2447, 2519, 2535, 2748, 2859, 2865, 2896, 2994, 3003, 3083, 3150, 3167]. domain-based [1423]. domains [489, 2881]. Dominant [42, 1562]. donations [248]. DONES [1717]. DoS/DDoS [1660]. DoS/DDoS-MQTT-IoT [1660]. Double [61, 549, 868, 1247, 1289, 1292, 2475, 2522, 2725, 2750, 3152]. double-stage [868]. DoubleR [2303]. Douyin [2973]. down [966, 2486]. down-link [2486]. Downlink [373, 698, 732, 763, 1131, 1142, 1330, 1479, 2035, 2130, 2490]. downlink-uplink [1142]. downlink/uplink [373]. downloading [2298]. DP [1312, 2935]. DP-SAFL [2935]. DPA [1350]. DPDK [2250]. DPDK-integrated [2250]. DPoS [826, 2225]. DPoW [3055]. DPP [2609]. DQN [2835, 3074, 3149]. DQN-enhanced [3074]. DQN-GRU [3149]. DR3DH [2906]. DRAFTEE [2977]. Dragonfly [2931]. DrDoS [2289]. drift [1529, 2054, 2844]. drifted [1596]. Driven [15, 154, 211, 271, 494, 495, 677, 750, 944, 955, 1177, 1267, 1291, 1332, 1390, 1392, 1395, 1557, 1571, 1695, 1761, 1884, 1929, 1965, 2095, 2111, 2137, 2253, 2280, 2332, 2337, 2360, 2370, 2454, 2498, 2547, 2643, 2646, 2687, 2705, 2718, 2731, 2798, 2799, 2823, 2844, 2866, 2869, 2902, 3028, 3112]. driver [53, 2664]. Driving [309, 626, 2196, 2989]. DRL [1190, 1888, 2095, 2310, 2331, 2335, 2356, 2655, 2684, 2789, 2846, 3096]. DRL-based [2310, 2331, 2655, 2684]. DRL-driven [2095]. DRLAR [2151]. DRLO [2961]. DROI [6]. Drone [63, 778, 845, 1370, 1506, 1544, 1886, 3082, 3184]. drone-assisted [1370, 1544]. Drone-enabled [845]. dronecells [2696]. drones [322, 604, 965, 1018, 1784, 1919, 1930, 2175, 3068]. Drops [1196, 1375]. drug [116]. drug-delivery [116]. DSMLB [1183]. DSO [1375]. DSSS [503]. DSVDD [1810]. DT [2506]. DT-Block [2506]. DTLShps [1224]. DTN [2537]. DTPBFT [2295]. DTS [2713]. DTS-BWpredictor [2713]. Dual [139, 462, 766, 850, 987, 1031, 1322, 1358, 1444, 1601, 1973, 1975, 2011, 2118, 2182, 2263, 2373, 2533, 2594, 2610, 2622, 2713, 2736, 2839, 2896, 3002, 3006, 3022, 3064, 3099, 3135, 3137, 3188]. dual- [2622]. dual-antenna [139]. dual-aspect [2263]. dual-branch [2839, 3022, 3188]. dual-grained [1322]. dual-granularity [2594]. dual-layer [1444, 2610, 2896, 3137]. dual-link [850]. dual-modal [3064]. Dual-Queues [2373]. dual-radio [1031]. Dual-scale [2713]. dual-slope [462]. dual-view [3099]. dual-weight [987, 2533]. DUdetector [2594]. due [446]. Dueling [3152]. Duplex [12, 142, 551, 820, 1083, 1281, 1830, 2409, 2440, 2542]. Duplicate [780, 1453]. duplication [1750]. DUPS [2579]. duration [645]. duration-aware [645]. during [217, 445, 1369, 1496, 2069]. duty [1546]. DV [1072, 1806, 2669]. DV-Hop [1072, 1806, 2669]. DVB [1835]. DVB-S2X [1835]. DWT [1180]. Dynamic [5–7, 30, 34, 111, 171, 258, 280, 294, 338, 375, 378, 385, 406, 451, 491, 514, 530, 535, 542, 562, 587, 627, 668, 676, 685, 736, 953, 979, 996, 1037, 1046, 1139, 1183, 1184, 1200, 1210, 1214, 1292, 1336, 1409, 1419, 1462, 1469, 1497, 1511, 1534, 1568, 1598, 1666, 1667, 1710, 1732, 1750, 1751, 1781, 1792, 1813, 1816, 1855, 1869, 1876, 1946, 1968, 1987, 1994, 2015, 2106, 2210, 2232, 2247, 2276, 2295, 2335, 2399, 2407, 2423, 2577, 2655, 2667, 2678, 2696, 2715, 2727, 2744, 2756, 2782, 2795, 2843, 2885, 2888, 2899, 2915, 2958, 2959, 2961, 2963, 2980, 3000, 3019, 3027, 3031, 3073, 3083]. dynamic [218, 647, 921, 1248, 1699, 2082, 2322, 2713, 3036, 3063, 3153]. dynamical [2228, 2252, 2539]. Dynamically [918, 1819]. dynamics [304, 855, 1476, 2197, 2208, 2591]. DynamicTuple [921]. DynGATT [1462]. DySO [1498].

e-commerce [411]. E-health [264, 932, 2443]. e-Learning [217]. EagerCC [1823]. EAOS [2603]. EAP [2089, 2924]. EAP-FIDO [2924]. EAP-OAUTH [2089]. EAPT [2618]. Early [319, 437, 681, 874, 1729, 1830, 1932, 2988]. earth [1687, 2324]. eavesdroppers [1480, 1733]. Eavesdropping [2048]. eBPF [469, 1449, 2058]. EC [1504, 2289]. EC-GCN [1504]. EC-NTD [2289]. ECC [1489, 2666]. ECC-based [1489, 2666]. ECCPWS [1489]. ECG [2668]. ECN [767]. Economic [516, 1035, 2544]. economics [68]. economy [123, 1280, 1373]. ecosystem [909, 1054]. ecosystems [1016]. ED [1818]. ED-VNE [1818]. Edge [47, 194, 215, 359, 436, 559, 574, 611, 667, 677, 696, 697, 742, 781, 785, 787, 789, 814, 823, 863, 872, 947, 956, 979, 1015, 1084, 1131, 1156, 1190, 1193, 1250, 1289, 1366, 1368, 1370, 1427, 1431, 1455, 1474, 1475, 1484, 1552, 1579, 1582, 1591, 1622, 1623, 1654, 1713, 1731, 1878, 1910, 1951, 1975, 2016, 2038, 2065, 2076, 2080, 2108, 2113, 2144, 2155, 2172, 2198, 2201, 2224, 2242, 2244, 2288, 2296, 2323, 2324, 2338, 2364, 2395, 2438, 2461, 2492, 2518, 2532, 2564, 2578, 2635, 2755, 2785, 2805, 2835, 2845, 2846, 2861, 2870, 2897, 2898, 2906, 2910, 2917, 2953, 2961, 3011, 3031, 3043, 3067, 3095, 3172]. edge [46, 212, 266, 268, 387, 395, 398, 401, 513, 549, 632, 642, 645, 680, 692, 824, 869, 938, 974, 1032, 1040, 1059–1061, 1071, 1085, 1111, 1118, 1121, 1132, 1135, 1139, 1186, 1274, 1340, 1344, 1345, 1357, 1359, 1394, 1432, 1470, 1473, 1479, 1528, 1553, 1555, 1568, 1593, 1642, 1700, 1710, 1730, 1742, 1758, 1760, 1763, 1803, 1833, 1844, 1973, 2017, 2032, 2040, 2051, 2063, 2094, 2146, 2170, 2200, 2262, 2283, 2310, 2331, 2356, 2365, 2383, 2387, 2433, 2448, 2466, 2473, 2510, 2514, 2590, 2604, 2605, 2640, 2667, 2681, 2685, 2700, 2777, 2790, 2799, 2840, 2868, 2930, 2935, 2971, 2972, 2985, 3018, 3030, 3142, 3144, 3149, 3182]. edge [90, 211, 257, 363, 366, 404, 533, 558, 637, 650, 733, 862, 882, 934, 942, 955, 959, 981, 1106, 1216, 1248, 1312, 1318, 1337, 1381, 1406, 1428, 1467, 1472, 1525, 1594, 1601, 1655, 1732, 1805, 1856, 2020, 2021, 2073, 2086, 2142, 2158, 2270, 2276, 2308, 2576, 2589, 2631, 2643, 2679, 2684, 2686, 2719, 2726, 2909, 2932, 2933, 2964, 2982, 2992, 3061, 3183]. edge-aided [2868]. edge-assisted [2094, 2578, 2681]. edge-based [1071, 1474]. Edge-caching [2845]. edge-cloud [215, 2038, 2200, 2395, 2448, 3182]. edge-cloud-based [1601]. edge-computing [781, 2016]. Edge-device [2492]. edge-edge [650, 2308]. Edge-enabled [2473, 3031]. edge-end [1568]. Edge-enhanced [1878]. Edge-to-Cloud [3067]. edge-to-core [2032]. EdgeBoost [3005]. EdgeML [1344]. edges [90, 977, 1936, 2071]. EDiPSo [849]. editable [2536]. Editor [2959]. Editorial [24, 25, 48, 70, 87, 105, 125, 146, 166, 175, 184, 200, 220, 243, 269, 295, 323, 360, 396, 421, 448, 479, 500, 522, 546, 569, 591, 613, 635, 662, 664, 687, 703, 709, 740, 776, 812, 847, 878, 911, 928, 948, 951, 971, 999, 1023, 1042, 1043, 1057, 1075, 1091, 1108, 1109, 1136, 1168, 1195, 1197, 1203, 1220, 1261, 1304, 1342, 1379, 1415, 1440, 1464, 1481, 1519, 1549, 1564, 1587, 1613, 1634, 1649, 1673, 1693, 1725, 1751, 1765, 1807, 1842, 1904, 1925, 1943, 1966, 2001, 2041, 2077, 2565]. Editorial [2104, 2139, 2156, 2193, 2220, 2240, 2300, 2346, 2379, 2413, 2511, 2568, 2586, 2661, 2704, 2743, 2778, 2811, 2848, 2879, 2911, 2926, 2947, 2965, 3032, 3120, 3189]. EDTP [917]. education [1860]. EDVFL [3126]. EEM [1679]. Effect [203, 963, 1411, 1443, 1920]. effect-based [1443]. Effective [63, 93, 160, 172, 208, 235, 382, 683, 760, 909, 1518, 1662, 1706, 2188, 2303, 2601, 2729, 2814, 2953, 2978, 3018, 3088]. Effectiveness [148, 2260, 2302]. effects [1684]. Efficiency [138, 169, 202, 272, 310, 559, 560, 632, 818, 854, 1308, 1498, 1796, 1869, 1900, 1997, 2025, 2130, 2217, 2352, 2423, 2433, 2584, 2717, 2760, 2933, 3021, 3044]. Efficient [6, 30, 51, 53, 101, 179, 292, 349, 389, 403, 420, 461, 463, 534, 541, 576, 579, 633, 714, 749, 781, 800, 807, 849, 868, 910, 1018, 1021, 1106, 1145, 1185, 1192, 1194, 1205, 1229, 1281, 1287, 1326, 1426, 1427, 1435, 1505, 1507, 1527, 1539, 1541, 1594, 1680, 1681, 1691, 1703, 1714, 1720, 1745, 1853, 1933, 1983, 2005, 2053, 2099, 2110, 2121, 2186, 2190, 2228, 2242, 2289, 2298, 2339, 2349, 2375, 2409, 2421, 2439, 2446, 2474, 2500, 2505, 2506, 2515, 2530, 2575, 2595, 2647, 2662, 2705, 2708, 2736, 2748, 2754, 2763, 2785, 2830, 2854, 2857, 2887, 2888, 2920, 3058, 3075, 3076, 3103, 3134, 3161, 3179]. efficient [17, 38, 63, 92, 182, 231, 236, 252, 311, 404, 564, 589, 637, 660, 705, 736, 739, 743, 759, 772, 805, 824, 839, 871, 876, 888, 892, 937, 949, 957, 989, 1019, 1029, 1086, 1143, 1167, 1172, 1240, 1242, 1243, 1248, 1253, 1271, 1294, 1300, 1302, 1307, 1321, 1327, 1377, 1408, 1413, 1454, 1528, 1610, 1621, 1626, 1629, 1718, 1739, 1757, 1758, 1797, 1799, 1800, 1895, 1906, 1911, 1937, 1942, 1995, 2017, 2068, 2142, 2175, 2195, 2223, 2255, 2345, 2371, 2406, 2417, 2431, 2480, 2508, 2509, 2535, 2554, 2559, 2688, 2700, 2735, 2842, 2885, 2903, 2968, 2995, 2999, 3007, 3038, 3054, 3063, 3083, 3084, 3101, 3138, 3184]. efficiently [2256]. effort [1695]. efforts [284]. egress [576]. EH [2129, 3181]. EH-CRN [2129]. eHDDP [597]. ehealth [958]. EHF [734]. EHR [2523]. EID [1246]. Elastic [27, 77, 136, 138, 177, 237, 346, 504, 545, 584, 724, 752, 753, 871, 950, 1184, 1255, 1323, 1324, 1393, 1396, 1540, 1679, 1703, 1888, 1894, 2028, 2047, 2079, 2099, 2326, 2475, 2609, 2944, 2954, 2974]. Elastically [1908]. election [2494]. electric [2544]. Electromagnetic [181, 2928]. Electrosense [176]. elimination [420]. emails [1021]. eMBB [975, 2562, 2756]. embedded [437, 1069, 1621, 1834, 1899, 2478]. Embedding [6, 280, 409, 472, 699, 1226, 1309, 1391, 1393, 1541, 1703, 1734, 1933, 2118, 2170, 2260, 2335, 2508, 2631, 2751]. emergencies [2408]. Emergency [379, 432, 1157, 1386, 1742, 1910, 2597, 2638, 3109, 3167]. emergency/remote [2638]. emerging [581, 1066, 1347, 2100, 3147]. EMF [515]. emission [2067, 2928]. emission-aware [2928]. emissions [2060]. emphasis [2231]. Empirical [594, 1529, 1845, 1985, 2012, 2633, 2933]. Employing [1409]. empowered [1281, 1648, 1854, 1963, 2027, 2051, 2102, 3046]. Empowering [76, 905, 2733]. EMTD [3179]. emulation [216, 666, 1059, 2248]. emulators [1860]. Enable [716, 1701]. Enabled [9, 15, 19, 56, 59, 62, 110, 157, 160, 175, 188, 209, 282, 373, 379, 395, 399, 414, 437, 453, 487, 539, 559, 707, 819, 845, 871, 873, 919, 945, 947, 955, 989, 1008, 1016, 1019, 1027, 1064, 1088, 1145, 1156, 1246, 1306, 1340, 1350, 1368, 1381, 1450, 1472, 1494, 1570, 1597, 1610, 1627, 1732, 1758, 1789, 1815, 1837, 1871, 1945, 1952, 1998, 2092, 2102, 2141, 2149, 2153, 2187, 2261, 2281, 2349, 2367, 2421, 2449, 2455, 2467, 2473, 2507, 2584, 2626, 2790, 2866, 2878, 2902, 2928, 2962, 2983, 3031, 3087, 3101, 3111, 3115, 3172]. enablers [174, 1850]. Enabling [174, 222, 367, 833, 1078, 1090, 1364, 1801, 1881, 1882, 1995, 2271, 2287, 2441, 2765, 2785, 2858, 2969, 3100]. eNCache [1897]. encapsulation [45]. encoder [1180, 1557, 2404, 2781, 2854]. encoder/decoder [2854]. encoding [2351, 2391, 2971]. encounter [338]. Encrypted [208, 290, 425, 589, 843, 872, 903, 1117, 1148, 1150, 1493, 1504, 1529, 1533, 1596, 1599, 1786, 1811, 1893, 1902, 1921, 2045, 2118, 2120, 2137, 2192, 2286, 2291, 2425, 2493, 2543, 2560, 2618, 2651, 2688, 2891, 3013, 3060, 3064, 3148, 3177, 3179]. Encrypted-OFDM [2543]. Encryption [393, 702, 875, 885, 1283, 1307, 1389, 1683, 1851, 2119, 2125, 2253, 2265, 2307, 2376, 2523, 2535, 3030, 3051]. EncryptoVision [3064]. End [107, 228, 251, 482, 1147, 1537, 1547, 1568, 1655, 1658, 1735, 1742, 1835, 2154, 2165, 2250, 2496, 2499, 2599, 2846, 3112, 3129]. end-host [2599]. end-system [2154]. End-to-edge [1655]. End-to-end [228, 251, 482, 1147, 1537, 1547, 1735, 1835, 2165, 2250, 2496, 2499, 3112, 3129]. endogenous [673]. endowed [2324]. endpoint [2297]. endpoints [1005]. Energy [41, 43, 69, 107, 137, 138, 169, 202, 206, 231, 278, 310, 349, 359, 373, 381, 389, 398, 560, 693, 695, 705, 714, 743, 759, 781, 854, 868, 892, 917, 937, 956, 969, 980, 1018, 1071, 1088, 1093, 1129, 1145, 1167, 1172, 1192, 1229, 1243, 1253, 1281, 1286, 1287, 1308, 1326, 1507, 1527, 1600, 1631, 1681, 1691, 1692, 1720, 1739, 1757, 1758, 1796, 1825, 1867, 1873, 1906, 1913, 1933, 1937, 1977, 1992, 1997, 2053, 2099, 2195, 2283, 2331, 2345, 2349, 2352, 2364, 2409, 2421, 2446, 2462, 2500, 2507, 2509, 2541, 2584, 2685, 2714, 2717, 2754, 2842, 3021, 3039, 3071, 3084, 3101]. energy [6, 17, 38, 142, 252, 275, 342, 404, 445, 632, 637, 654, 660, 818, 839, 1019, 1032, 1038, 1040, 1086, 1238, 1271, 1321, 1413, 1436, 1494, 1511, 1629, 1728, 1799, 1818, 1831, 1952, 1955, 1970, 2025, 2068, 2130, 2142, 2245, 2278, 2333, 2358, 2552, 2559, 2684, 2700, 2724, 2759, 2840, 2933, 3187]. Energy-aware [38, 69, 1038, 1088, 1238, 1600, 1970, 2759]. energy-constrained [2700]. energy-efficiency [632]. Energy-Efficient [6, 231, 252, 349, 389, 404, 637, 660, 759, 781, 868, 892, 1018, 1086, 1145, 1167, 1172, 1192, 1253, 1271, 1281, 1287, 1321, 1326, 1507, 1527, 1757, 1758, 1906, 1933, 1937, 2195, 2345, 2349, 2421, 2509, 2559, 2842, 3084, 3101]. energy-harvesting [41, 373, 695, 969, 2278]. Energy-latency [2685]. energy-optimal [1831]. energy-saving [1977]. energy-sensitive [2331]. energy-spectral [1308]. ENERO [1194]. enforcement [4, 656, 1050, 1166, 1267]. Enforcing [1855]. engine [1363, 1999]. Engineering [461, 571, 576, 1052, 1217, 1644, 1995, 2159, 2735, 2852, 2942]. English [3123]. enhance [598, 2146, 2440, 2806]. Enhanced [28, 35, 80, 477, 597, 650, 794, 806, 992, 1072, 1313, 1477, 1483, 1599, 1624, 1878, 1912, 2039, 2082, 2084, 2224, 2225, 2266, 2385, 2386, 2422, 2443, 2534, 2547, 2612, 2659, 2695, 2721, 2759, 2784, 2902, 2950, 2978, 2994, 3035, 3057, 3069, 3073, 3074, 3089, 3131, 3152]. Enhanced-SETL [2385]. Enhancement [933, 1039, 1201, 1458, 1590, 1730, 2444, 2769, 2819, 2878, 3133]. enhancements [1119]. Enhancing [8, 694, 1244, 1306, 1498, 1981, 1993, 2025, 2089, 2259, 2378, 2403, 2472, 2529, 2634, 2644, 2702, 2718, 2740, 2750, 2804, 2869, 2870, 2998, 3107, 3162]. eNodeB [1801]. eNodeBs [578]. enrichment [735]. Ensemble [13, 74, 182, 957, 1354, 2163, 2243, 2360, 2708, 2774, 2903, 2988]. Ensemble-model-based [13]. ensure [2000]. Ensuring [355, 2832]. enterprise [751, 1322, 1374, 1706]. entities [2889]. entity [1769, 2889, 3060]. Entropy [986, 1845, 2083]. entry [499]. Environment [207, 257, 268, 498, 516, 526, 564, 707, 809, 917, 976, 1027, 1085, 1135, 1138, 1200, 1205, 1238, 1341, 1510, 1579, 1593, 1771, 1774, 1861, 1949, 2087, 2306, 2462, 2819, 3060]. environmental [2043, 2476]. environments [16, 56, 64, 317, 607, 755, 771, 811, 869, 1090, 1100, 1188, 1307, 1335, 1367, 1423, 1475, 1690, 1785, 1855, 1885, 1896, 1914, 1915, 2024, 2145, 2196, 2264, 2417, 2488, 2646, 2671, 2728, 2804, 2820, 2837, 2872, 2943, 3067]. envisioned [1802]. EON [177, 2052]. EONs [1434, 3027]. EPC [2903]. ephemeral [277]. epidemic [1486, 2881]. epidemiological [304]. EPIoT [1912]. equality [3051]. equalization [2654]. Equalizing [1526]. equation [1841]. Equilibrium [383, 1420]. equipped [1077]. equivalence [50]. era [62, 213, 276, 1124, 1389, 1544, 2112, 2944]. ERASMIS [2666]. erasure [966, 3087]. erasure-coded [966, 3087]. ERO [2883]. ERO-Day [2883]. ERRANT [216]. Error [748, 1247, 1382, 2034, 2455, 2914]. error-correcting [2914]. ES-SDPC [2692]. ESN [1657]. ESRAS [1300]. Establishment [2400]. estimate [255, 908, 1845]. estimating [1659, 1791, 1865, 1989]. Estimation [444, 738, 805, 890, 891, 1065, 1345, 1363, 1835, 1863, 2642, 2663, 2669, 3089]. ESTK [3060]. ESTK-JC [3060]. Ethereum [1902, 2482, 2751]. Ethernet [226, 1717, 1720]. Ethernet-based [1717]. ethical [3052]. ETSI [1719]. EUAV [3057]. Europe [337]. evaluable [3126]. evaluated [578]. Evaluating [351, 941, 1660, 2115, 2316, 2548, 2635, 2760]. Evaluation [18, 33, 36, 39, 89, 210, 259, 376, 392, 393, 402, 429, 524, 579, 594, 634, 666, 682, 699, 771, 846, 942, 1050, 1122, 1255, 1296, 1325, 1333, 1336, 1346, 1407, 1589, 1630, 1740, 1821, 1928, 1962, 2024, 2054, 2145, 2284, 2405, 2449, 2521, 2613, 2752, 2789, 2910, 2922, 3109, 3140, 3171]. evaluations [719]. Event [418, 641, 1246, 1679, 2267, 2918]. event-participant [418]. events [778, 1275]. Everything [212, 2691, 3115]. EVFL [2750]. EVFL-DCs [2750]. eviction [256, 1642]. evidence [2012]. evil [1574]. EvoIoT [1367]. Evolution [615, 868, 1748, 1841, 2056, 2524, 2739, 2963, 3147, 3176]. evolutionary [75, 151, 278, 473, 1367, 2107]. Evolving [61, 1817, 2177, 2720]. Exact [300]. Examples [332, 2865]. exceeding [3017]. excessive [1632]. Exchange [251, 287, 639, 728, 2123]. EXCLF [2357]. Executable [3180]. execution [498, 536, 576, 695, 1475, 2331, 2541]. exemplar [1493]. exfiltration [694, 1070]. existing [884, 2833, 3098]. exit [2988]. eXO [2442]. Expediting [1174]. experience [206, 225, 1782, 1886, 3029]. experiment [580]. Experimental [33, 367, 429, 653, 750, 783, 948, 1051, 1122, 1543, 1783, 1861, 2033, 2054, 2096, 2745, 2833]. experimentation [56, 427, 983, 1749, 1801, 2951]. Expiration [2675]. Explainable [1378, 2066, 2343, 2770, 3013, 3047, 3118]. explainers [2683]. explaining [1251]. explicit [2936]. exploit [183]. exploitation [129, 1972, 2390, 2432, 3161]. Exploiting [90, 99, 598, 887, 1104, 1130, 2141, 2819, 2874, 2914]. exploits [2203, 3066]. exploration [23, 1309, 2168, 2702]. Exploring [1077, 1452, 1537, 1684, 1777, 2197, 2260, 2678, 2757, 2893, 2976, 3059]. exponentially [743]. exponentially-ant [743]. Exposing [2603]. exposure [515, 1245, 2515, 2810]. expression [51, 2355]. Extended [1869, 1948, 2740, 2906, 3012]. Extending [281, 1847, 2214]. extensible [867]. extensions [64, 284]. extensive [920]. external [2350]. extract [1275]. extracting [1171]. extraction [782, 883, 1257, 1522, 2366, 2478, 2504]. extremely [281, 1471]. eXtremes [2231].

F [1824]. F-NIDS [1824]. F2DC [1279]. F4Tele [483]. fabric [792]. Face [1529, 2468]. facial [2103]. Facilitating [506]. facilities [1717]. facing [2487]. factor [242, 384, 838, 1478, 1544, 2176, 2494, 3057]. factories [1742, 2513, 2856, 3039, 3137]. factorization [738, 891]. factory [2723]. fading [142, 1736, 1835, 1858, 2055]. Failure [446, 622, 789, 1055, 1254, 1297, 1378, 1457, 1669, 2168, 2292, 2517, 2533, 2786, 2825]. Failure-resilient [789]. failures [239, 445, 623, 681, 1176, 1422, 1879, 1909, 1935, 2734, 3019]. Fair [689, 712, 1049, 1518, 1554, 1555, 1900, 1950, 2000, 2037, 2255, 2336, 2995]. Fairness [2, 124, 485, 631, 916, 1011, 1358, 1466, 1506, 1798, 2908, 2923, 3010]. Fairness-Aware [124]. false [2186]. falsification [44]. family [430, 2994]. FAMTDS [2306]. FANET [799, 2022, 2292, 2528, 2541]. FANETs [799, 1317, 1366]. FARASP [2974]. farming [46, 56, 58, 62, 129, 817]. Fast [50, 156, 199, 247, 255, 659, 784, 804, 890, 1053, 1062, 1254, 1362, 1497, 1723, 1786, 1868, 1893, 1899, 2063, 2217, 2278, 2393, 2469]. FastCredit [1174]. FastDet [2582]. Faster [57, 2582, 2608]. FastTraffic [1786, 1893]. FASUS [199]. fat [1217]. fat-tree [1217]. Fatigue [2258]. Fault [7, 95, 187, 545, 791, 825, 1129, 1234, 1268, 1299, 1319, 1402, 1611, 1785, 1868, 1906, 2183, 2304, 2333, 2473, 2486, 2488, 2501, 3082, 3116]. Fault-tolerant [1268, 1611, 1785, 2486, 3082, 3116]. FCNR [659]. FDMA [1190, 3077]. FDMA-based [1190]. FDR [1480]. FDSS [2885]. feasibility [1689]. feasible [915]. Feature [144, 158, 182, 430, 456, 857, 883, 904, 910, 1201, 1257, 1303, 1365, 1542, 1581, 1889, 1899, 2360, 2478, 2493, 2791, 2834, 2865, 2893, 3133, 3188]. feature-embedded [1899]. Features [566, 589, 926, 1266, 1948, 2061, 2103, 2137, 2269, 2348, 2391, 2796, 2823, 2874, 3045, 3060]. featuring [1755]. February [49, 71, 480, 501, 952, 972, 1441, 1465, 1764, 1944, 1967, 2585, 2660]. FedAAA [1927]. FedAAA-SDN [1927]. Federated [401, 489, 633, 765, 824, 841, 916, 941, 955, 960, 961, 973, 1035, 1067, 1126, 1155, 1182, 1216, 1344, 1366, 1406, 1427, 1474, 1495, 1531, 1537, 1555, 1652, 1756, 1761, 1762, 1772, 1820, 1895, 1927, 2012, 2066, 2070, 2101, 2112, 2115, 2121, 2181, 2190, 2191, 2224, 2232, 2237, 2257, 2282, 2283, 2341, 2344, 2376, 2382, 2435, 2445, 2491, 2502, 2506, 2521, 2550, 2554, 2571, 2700, 2701, 2705, 2710, 2736, 2750, 2755, 2763, 2768, 2798, 2805, 2819, 2834, 2841, 2842, 2851, 2874, 2899, 2910, 2935, 2960, 2985, 2998, 3000, 3009, 3013, 3015, 3016, 3024, 3025, 3036, 3047, 3048, 3058, 3061, 3067, 3104, 3155, 3170, 3173, 3175, 3180, 3184]. federated [1128, 1467, 1824, 1970, 2073, 2125, 2245, 2420, 2579, 2589, 2625, 2766, 2866, 2940]. Federated-Learning [2101]. Federation [1065, 1740]. FedLLMGuard [3025]. FedMP [2625]. FedPA [841]. FedSBS [2101]. FedShelter [2763]. FedSSC [1895]. FedXAI [3047]. fee [1178]. Feedback [1408, 1781, 2621, 2694, 2843]. Feedback-Based [1408, 1781, 2843]. fees [127]. femto [579]. FemtoCaching [1414]. femtocell [67]. Few [779, 1377, 1599, 1628, 3080]. Few-shot [779, 1377, 1599, 1628, 3080]. FFRLI [1868]. FHSS [2659]. Fi [94, 506, 698, 797, 998, 1446, 1617, 1940, 2000, 2274, 2385, 2694, 2824, 3008, 3147]. fiber [172, 963]. fiber-optic [963]. FIBs [2464]. Fidelity [1352, 2248, 2574]. Fidelity-aware [2248]. FIDO [2924]. FIDO2 [2924]. Field [347, 350, 1069, 1186, 1779, 1847, 2663, 2740]. field-based [1186]. Field-of-View [2740]. fieldable [1069]. fields [289]. file [1181, 1840]. filling [2716]. Filter [155, 720, 1006, 1606, 1656, 1892, 1958, 2675]. Filtering [582, 1021, 2063, 2323, 2393, 2527, 2558, 2608, 3042, 3140]. Filtering-based [2063]. Finding [341, 431, 1704, 1892]. Fine [117, 542, 1189, 1318, 1385, 1601, 1652, 1712, 1902, 2118, 2539, 2580, 2767, 2781]. Fine-grained [542, 1189, 1318, 1385, 1601, 1652, 1712, 1902, 2118, 2539, 2580, 2767]. fine-tuned [117]. Fine-tuning [2781]. Fingerprint [1916, 2192, 2325, 3133]. fingerprint-based [1916]. Fingerprinting [394, 779, 834, 874, 1006, 1273, 1372, 1377, 1584, 1595, 1628, 1630, 1662, 1688, 1811, 1993, 2497, 2807, 3034]. Fingerprinting-assisted [3034]. fingerprints [903, 1328]. finite [2913]. FINT [1233, 2292]. fire [437]. firefly [778]. Firewall [1207, 2608, 2729]. First [165, 679, 1244, 1487, 1662, 1886, 1951, 2916]. first-person-view [1886]. fitness [548]. fitting [578]. five [778]. five-step [778]. FIWARE [711]. Fixed [815, 826]. FL [1182, 1652]. FL-based [1652]. flagging [374]. FLAME [640]. flash [1301]. FLchain [1325]. FLCom [3009]. fledged [1972]. Flex [1200]. Flex-CC [1200]. flexgrid [348]. Flexible [16, 303, 466, 633, 876, 923, 949, 968, 1135, 1200, 1215, 1224, 1233, 1438, 1446, 1491, 2165, 2723, 3130]. Flight [404, 1036, 2408, 2885]. Flipping [2445]. Flood [487, 499, 1656]. flooding [1947]. floor [2503]. Flow [40, 42, 143, 155, 187, 192, 218, 380, 403, 446, 461, 508, 509, 589, 603, 685, 760, 808, 891, 1008, 1273, 1351, 1501, 1521, 1696, 1708, 1723, 1773, 1865, 1908, 1948, 2137, 2171, 2177, 2188, 2228, 2287, 2365, 2375, 2485, 2497, 2723, 2891, 2923, 2990, 3007, 3010, 3035, 3134]. Flow-level [187, 461, 1273, 2171]. FlowFight [723]. flowing [2707]. flowlet [2539, 2616]. flows [402, 499, 1096, 1306, 1426, 1640, 1658, 1803, 1845, 1892, 2154, 2203, 2212, 2582, 2675, 2794, 2813, 3110]. fluctuating [2977]. fluctuations [1957]. FLV [2416]. Flying [109, 1503, 1608, 1661, 2084, 2235, 3116]. flying-beamformer [1661]. FlyingLoRa [1413]. FMORT [2541]. FOCCA [2676]. focus [1738, 2146]. Fog [209, 225, 279, 317, 343, 389, 475, 689, 702, 706, 722, 839, 840, 946, 976, 1003, 1016, 1027, 1087, 1100, 1140, 1149, 1179, 1188, 1209, 1227, 1238, 1296, 1401, 1423, 1429, 1494, 1542, 1615, 1683, 1785, 1891, 1906, 1949, 2013, 2143, 2283, 2294, 2304, 2676, 2705, 2872, 2877, 2937, 3044]. fog-aided [722]. Fog-based [317, 389, 1140, 1149, 1296, 2143]. fog-blockchain [2877]. Fog-cloud [1542, 2676]. Fog-driven [2705]. fog-enabled [1016, 1027, 1494]. Fog-oriented [946]. fog/cloud [1785]. following [2400]. footprints [1704]. Forecasting [59, 487, 657, 931, 1189, 1534, 1772, 1816, 2012, 2044, 2050, 2271, 2292, 3062].

Forensic [2956]. forest [809, 2791]. Form [1453]. formal [178, 287, 964]. Formalization [260, 557]. format [273, 1434]. formation [69, 2074]. formatting [1626]. formulation [256]. formulations [417]. Formullar [466]. forward [1185, 1900, 2238, 3131]. forward-secure [1185, 1900]. forwarder [1422, 3026]. Forwarding [23, 50, 156, 472, 762, 829, 1030, 1360, 1719, 1812, 1977, 2814]. fostering [56]. FoT [1719]. Foundation [2976]. four [190]. four-stage [190]. FPGA [466, 1082, 2854]. FPGA-accelerated [1082]. FPGA-based [466]. FPGAs [31]. fractal [1080]. fractional [1996]. fracture [2820]. Fragmentation [136, 177, 346, 1255, 1894, 2234, 2954, 2974]. Fragmentation-aware [136, 346, 2234, 2974]. Frame [481, 1293, 1745, 2049, 2659]. Framed [1352]. frames [1680]. Framework [163, 253, 256, 266, 274, 283, 303, 314, 325, 333, 419, 496, 512, 518, 541, 564, 583, 619, 628, 715, 796, 819, 846, 861, 946, 968, 1009, 1035, 1050, 1059, 1067, 1089, 1101, 1105, 1144, 1175, 1199, 1228, 1316, 1352, 1366, 1405, 1427, 1442, 1460, 1477, 1499, 1504, 1552, 1606, 1621, 1646, 1678, 1679, 1697, 1762, 1763, 1787, 1839, 1849, 1901, 1931, 1933, 2054, 2102, 2174, 2261, 2310, 2342, 2371, 2374, 2375, 2398, 2400, 2421, 2449, 2451, 2550, 2593, 2640, 2662, 2730, 2773, 2818, 2842, 2903, 2928, 2952, 2977, 3028, 3042, 3071, 3073, 3092, 3096, 3112, 3115, 3138, 3157, 3170, 3173, 3178]. framework [27, 418, 510, 588, 643, 647, 867, 1265, 1404, 1449, 1471, 1524, 1534, 1576, 1721, 1854, 2095, 2151, 2175, 2196, 2203, 2222, 2308, 2517, 2646, 2658, 2692, 2698, 2735, 2788, 2794, 2882, 2909, 2968, 3031]. frameworks [129]. fraud [3073]. fraudsters [1574]. free [179, 447, 868, 901, 1020, 1445, 1507, 1514, 1566, 1625, 1780, 1914, 1959, 2197, 2353, 2358, 2393, 2531, 2737, 2871, 3034]. free-ride [2358]. FreeBSD [2362]. FreeBSD-13 [2362]. freed [767]. Freezing [2554]. frequencies [288]. Frequency [890, 1372, 1488, 1704, 1869, 1916, 2490, 2682, 2696, 2865, 2954, 3133, 3150]. Frequency-domain [2865, 3150]. Freshness [2161, 2614]. Friendship [693, 894, 2662, 2859]. FROG [2608]. Fronthaul [2106, 2184, 2381, 2402]. fruit [57]. FSO [483, 1411, 2706, 3136]. FSO-based [3136]. FT [60]. FT-Aurora [60]. FTG [2243]. FTG-Net-E [2243]. FTTx [678]. Full [12, 45, 551, 820, 1083, 1281, 1830, 1972, 2409, 2440, 2542, 2634]. Full-duplex [12, 551, 820, 1281, 1830, 2409, 2542]. full-fledged [1972]. fully [365, 612, 716, 1468, 2306]. fully-digital [716]. Function [15, 190, 250, 378, 412, 510, 525, 658, 668, 676, 692, 699, 854, 907, 989, 1090, 1106, 1160, 1161, 1218, 1219, 1258, 1264, 1268, 1368, 1417, 1422, 1430, 1457, 1459, 1469, 1484, 1515, 1669, 1703, 1795, 1830, 1846, 1942, 1983, 2005, 2014, 2026, 2047, 2053, 2126, 2150, 2211, 2236, 2304, 2406, 2426, 2508, 2569, 2571, 2592, 2593, 2631, 2715, 2733, 2752, 2826, 2864, 3052]. function-based [2047]. functionality [1111]. Functions [21, 262, 303, 330, 691, 931, 1375, 1641, 1665, 2722]. Fundamentals [2484, 2832]. funnel [2188]. funnel-shaped [2188]. fusing [2068, 3060]. Fusion [390, 626, 726, 782, 926, 1128, 1135, 1201, 1296, 1499, 1576, 1581, 1722, 2061, 2360, 2435, 2493, 2519, 2649, 2744, 3007, 3064, 3139, 3188]. fusion-based [2061, 3064]. Future [32, 228, 280, 294, 652, 727, 1124, 1286, 1366, 1395, 1483, 1490, 1531, 1580, 1600, 1705, 1777, 1863, 2100, 2216, 2238, 2384, 2389, 2419, 2428, 2477, 2651, 2652, 2741, 2793, 2807, 2832, 2871, 2902, 3147]. fuzziness [826]. fuzzing [2307]. Fuzzy [497, 520, 714, 839, 1003, 1167, 1975, 2557, 2910]. fuzzy-based [520].

G [3, 14, 15, 19, 32, 78, 108, 114, 131, 140, 149, 160, 170, 174, 186, 207, 211– 213, 219, 231, 1151, 2975]. G-3MRP [1151].

G-air-simulator [149]. G-enabled [160]. GA [1331, 1473]. GA-based [1331]. GAIL [1317]. Gain [2881]. Gait [1617]. gambling [2045]. Game [34, 215, 223, 230, 257, 413, 471, 617, 654, 663, 785, 855, 869, 956, 1097, 1151, 1225, 1455, 1514, 1530, 1531, 1640, 1744, 1859, 1876, 2076, 2107, 2147, 2294, 2495, 2522, 2725, 2755, 2919, 3015, 3159]. Game-based [223, 869]. game-theoretic [215, 785, 2522]. game-theoretical [1151]. games [1118, 3084]. gaming [977, 1659, 2087, 2855]. GAN [1009]. GAN-based [1009]. GANs [682, 2482]. GAT [1405, 2729]. GAT-based [1405]. gated [1557]. GateKeeper [2263]. Gateway [162, 872, 1458, 1527, 1951]. gateway-to-device [1527]. gateways [384, 1431, 1881, 1908, 1924, 2263, 2305]. Gathering [104, 1796, 2068]. GATT [1462]. GATT-based [1462]. Gaussian [235]. Gbps [1720]. GCACS [604]. GCACS-IoD [604]. GCN [1504, 1838, 2729, 2784, 2846, 2958, 3011, 3162]. GCN-based [2846]. GCN-LOF [3162]. GCP [2476]. gen [2361, 2844]. GEN-DRIFT [2844]. General [122, 1471, 1524, 1618]. generalisation [2134]. Generalizable [3102]. generalization [2826]. Generalized [383, 857, 1249, 2827]. generated [2447, 2905]. Generating [2918]. Generation [183, 505, 511, 593, 639, 648, 729, 787, 819, 1009, 1050, 1353, 1490, 1612, 1620, 1819, 1980, 2057, 2216, 2315, 2340, 2351, 2429, 2939, 2952, 3016, 3084]. Generative [682, 872, 1317, 1493, 2062, 2404, 2432, 2844, 3081, 3117, 3143]. Generator [1438, 2548, 2608, 2809]. generic [604, 1316, 1865, 2136]. Genetic [207, 621, 645, 680, 692, 2937, 3001]. genetical [1299]. GenNP [2809]. geo [319, 1132, 1285, 1290, 2228]. geo-distributed [319, 1285, 2228]. geo-indistinguishability [1290]. geo-replication [1132]. geofenced [2123].

Geographic [886, 1391]. geographical [660, 994, 1227]. geographical-aware [1227]. geographically [93, 498, 502]. geography [2060]. geolocalization [2915]. geolocation [1769, 2645, 2984, 3125]. geometric [3034]. Geometry [2764, 2967]. geosynchronous [1928]. GERT [2655]. gesture [996, 2131]. GFDA [1257]. GFDA-WGAN [1257]. GFL [2160]. GFL-LFF [2160]. GHz [334, 474, 653, 698, 716, 783, 1896]. GI [2975]. GI/G/1 [2975]. GKM [375]. GLADS [1533]. GlareShell [2138]. global [638, 1439, 1533, 2064, 2881]. global-local [1533]. GNN [1572, 1769, 2011, 2118, 2794, 3145]. GNN-based [1769]. gNodeB [1896]. goal [575]. goal-oriented [575]. Going [1597]. Good [540]. gossiping [1182]. Gov [1721]. governance [1721]. GPIO [1604]. GPS [321, 1232, 2452]. GPT [2905, 3073]. GPT-enhanced [3073]. GPU [2949]. GR [139]. Grace [1604]. Gradient [1201, 1427, 2540, 2554, 2768, 3036]. Gradient-Wise [2554]. gradual [172]. GRAIN [1150]. grained [542, 1189, 1318, 1322, 1385, 1601, 1652, 1712, 1902, 2118, 2539, 2580, 2767]. GRAKA [3038]. gram [2994]. GramMatch [883]. Grano [425]. Grano-GT [425]. Grant [1445, 1625, 2531, 2871]. Grant-free [1445, 1625, 2531, 2871]. Granular [425, 1150]. granularity [2277, 2594]. Graph [344, 403, 472, 538, 804, 919, 943, 1014, 1175, 1273, 1284, 1311, 1320, 1338, 1402, 1504, 1574, 1598, 1785, 1792, 1819, 1827, 1862, 1929, 2011, 2045, 2059, 2118, 2138, 2192, 2204, 2212, 2217, 2226, 2243, 2365, 2434, 2447, 2469, 2570, 2606, 2624, 2730, 2751, 2800, 2803, 2973, 3020, 3062, 3073, 3082, 3099, 3103, 3130, 3139, 3153, 3162, 3167, 3171, 3179]. Graph-aware [403]. Graph-based [1014, 1785, 1862, 2045, 2434, 3130, 3162]. graph-theoretic [943]. GraphCC [2624]. Graphical [1264]. graphlet [75].

GraphNEI [1769]. GraphPowerNet [1862]. graphs [2560, 2667]. GraphVeri [3178]. grasshopper [203]. GRE [1391]. Greedy [1391, 1420]. Green [372, 518, 817, 827, 1172, 1286, 1511, 1607, 1709, 2377, 2937]. Green-PoW [1172]. greenhouses [47]. Grid [82, 334, 577, 585, 641, 1074, 1777, 2455]. Grids [625, 1625, 2453, 2676]. GRL [2730]. GRL-RR [2730]. grooming [77, 491]. Ground [322, 425, 528, 626, 1013, 1108, 1609, 1861, 2076, 2467, 2486, 2584, 2617, 2658, 3085, 3096, 3127, 3136]. Ground-Integrated [3127]. ground-to-air [1609]. Group [205, 245, 358, 375, 488, 795, 898, 1185, 1240, 1256, 1443, 1718, 2253, 2354, 2591, 2913, 2995, 3038]. Group-based [205, 358, 488, 2253]. grouping [553, 2003, 2004]. GRU [3149]. GRU-based [2236]. GSDRL [2982]. GT [425]. guarantee [92, 663, 1798, 1873, 2406]. Guaranteed [389, 470, 476, 549, 1591, 1698, 2009]. guaranteeing [1982]. guarantees [353, 1667, 2299, 3104]. guard [2049, 2403]. Guest [662, 1042, 1108, 1195]. guide [123]. guided [1394, 2288, 2390, 3149, 3179]. Guifi.net [68]. GWO [2791]. GWO-XGBoost [2791]. GWPF [2554]. Gym [1410, 1442].

H [1366]. H-HOME [1366]. HAC [2467]. HAC-SAGIN [2467]. half [142]. half-duplex [142]. hand [996, 2131]. Handling [95, 630, 909, 1025, 2054, 2319, 2844, 3128]. handoff [701, 970]. Handover [204, 210, 369, 415, 457, 734, 1019, 1198, 1293, 1623, 2039, 2098, 2259, 2354, 3019, 3053]. handovers [1575]. handshake [1224, 2046]. HAP [2025, 2538]. HAP-LEO [2025]. hard [174]. hardening [1046]. hardware [45, 534, 681, 801, 846, 1788, 1845, 2058, 2921]. hardware-assisted [1788]. Harnessing [1496, 2361]. harvesting [41, 43, 142, 275, 373, 695, 956, 969, 1511, 1629, 1631, 1825, 1955, 1997, 2278, 2446, 2500, 3071]. Hash [156, 599, 2636]. HashXor [490]. HawkEye [2599]. hay [341]. HCNs [3056]. HDPoA [1288]. head [22, 2773]. header [2350]. headerless [2659]. headers [2425]. healing [1299, 2254, 2495]. Health [242, 264, 519, 590, 746, 749, 932, 1998, 2443, 2644]. health-care [242, 519]. HealthBlock [858]. Healthcare [279, 400, 611, 858, 863, 932, 995, 1078, 1115, 1197, 1496, 1923, 2191, 2209, 2309, 3146]. HealthRec [1998]. HealthRec-Chain [1998]. heavy [2675, 2975]. heavy-tailed [2975]. HECC [1222]. HELAD [74]. Hellinger [2994]. Hellman [2400, 2906]. helper [2456]. herd [835]. HERO [2883]. Het-MEC [1064]. HetECN [2489]. Heterogeneity [3046]. Heterogeneous [37, 67, 74, 102, 163, 219, 230, 289, 310, 313, 316, 320, 340, 352, 362, 369, 401, 409, 462, 476, 477, 621, 797, 810, 836, 963, 1081, 1085, 1135, 1213, 1293, 1332, 1351, 1400, 1404– 1406, 1414, 1479, 1570, 1746, 1763, 1853, 1913, 1924, 1949, 2045, 2128, 2201, 2262, 2386, 2447, 2469, 2472, 2474, 2491, 2612, 2649, 2662, 2698, 2736, 2764, 2771, 2840, 2851, 2856, 2874, 2930, 2935, 3076, 3099, 3103, 3155]. HeteroTiC [1351]. HetIoT [2205]. hetnets [27, 79, 137, 210, 233, 413, 1271, 1620, 2421, 2696]. Heuristic [680, 2472, 2541]. HGC [3062]. HGC-LSTM [3062]. HIBS [2483]. Hidden [224, 1008, 1045, 2542, 2791]. hide [2045]. Hiding [1875, 3094]. HIDRA [1423]. Hierarchal [207]. Hierarchical [452, 863, 910, 1117, 1761, 1789, 1899, 2062, 2064, 2107, 2243, 2277, 2338, 2474, 2483, 2535, 2621, 2723, 2873, 3031, 3152, 3182]. hierarchy [1003]. High [4, 10, 408, 455, 482, 513, 632, 644, 700, 723, 745, 763, 784, 792, 921, 1031, 1071, 1125, 1143, 1306, 1446, 1619, 1668, 1676, 1865, 1892, 1948, 2034, 2084, 2399, 2424, 2453, 2467, 2574, 2675, 2890, 2938, 2986, 3026, 3056, 3141, 3142, 3186]. High-altitude [2467]. high-bandwidth [4, 792]. High-Capacity [408]. high-density [1125, 2938]. high-DER-penetrated [2453]. high-dimensional [3141]. high-fidelity [2574]. high-mobility [3056]. high-performance [921, 1143, 1619]. High-precision [10, 2986]. high-rate [763]. high-resolution [3186]. high-speed [644, 700, 1668, 1865, 1892, 1948, 2675, 3026]. high-throughput [1446, 2084]. Higher [414, 699, 1900, 1978]. Higher-Level [414]. higher-order [1978]. Highly [60, 1247, 2295, 2363, 2429, 2713]. highway [1719]. hijacking [189, 925]. Histogram [2886]. Histogram-based [2886]. History [646]. hit [862]. HITLinQ [2124]. HL [2225]. HL-DPoS [2225]. Hoc [17, 124, 141, 459, 542, 617, 671, 1094, 1151, 1503, 1566, 1608, 1632, 1684, 1781, 1836, 1989, 2084, 2235, 2312, 2494, 2545, 2693, 2717, 2769, 2814, 3116, 3118, 3135]. HoDiNT [2275]. hole [259, 2254]. Holistic [853]. home [162, 1027, 1222, 1366, 1601, 2013, 2232, 2253]. homomorphic [2246]. Honesty [1288]. Honesty-based [1288]. honey [1638, 2774]. Honey-X [1638]. honeypot [2968]. Hop [10, 65, 1020, 1030, 1034, 1072, 1143, 1578, 1664, 1667, 1684, 1806, 1946, 1951, 2015, 2124, 2437, 2669, 2732, 2808]. hopping [97, 439, 991, 1173, 2490, 2657]. hopping-based [2490]. Horizontal [1289]. hospital [811]. host [1097, 1461, 2281, 2599]. host- [1461]. hotspot [3124]. hotspots [1488, 2832]. HotStuff [2802]. HOXC [2277]. HPC [56]. HSA [933]. HSA-SPC [933]. HSGAN [2062]. HSGAN-IoT [2062]. HSWF [1993]. HTML [3188]. HTTP [411, 1712, 2416, 3167]. HTTP-FLV [2416]. HTTP-level [411]. HTTP/3 [1712, 3167]. HTTPS [1070, 1921, 2178, 2612, 2708]. hub [352]. hub-spoke [352]. hull [1256]. Human [1, 121, 229, 996, 1187, 1245, 1365, 2090, 2330, 2988, 2997, 3086]. human-AI [2997]. human-based [229]. Human-centric [3086]. human-computer [996]. Human-in-the-loop [1187, 2997]. Hunter [2800]. HVVU [2636]. Hybrid [36, 134, 151, 159, 205, 308, 317, 335, 446, 463, 469, 508, 543, 550, 586, 593, 597, 615, 704, 753, 786, 790, 844, 857, 863, 871, 906, 933, 1283, 1557–1559, 1638, 1661, 1799, 1947, 2021, 2023, 2044, 2050, 2098, 2124, 2130, 2158, 2244, 2281, 2502, 2509, 2517, 2525, 2534, 2605, 2693, 2729, 2774, 2784, 2840, 2932, 2937, 2941, 2950, 2958, 2979, 2982, 3004, 3069, 3075, 3096, 3162, 3165, 3166, 3174, 3183]. hybrid-action [3096]. hybrid-powered [2950]. hybrid-task [2840]. Hyper [340]. hypergraph [93, 1258]. hyperparameter [857]. Hypervisor [899].

I/O [899]. I/Q [394]. I2P [684]. IaaS [60]. IAB [729, 1330, 2405]. IBDA [1914]. IBN [3112]. IBTA [1165]. IBTA-IoV [1165]. ICAN [1804, 3160]. ICMP [516]. ICN [511, 1068, 1716, 2464]. ICN-IoT [1716]. ICNIRP [515]. ICS [2342]. ICSDN [1360]. ICT [827]. ICV [1579]. ID [1736, 2829]. ID-AOPDDQN-based [2829]. Identification [1, 113, 277, 330, 881, 883, 903, 1311, 1378, 1596, 1617, 1769, 1810, 1902, 1916, 2061, 2252, 2263, 2269, 2297, 2416, 2493, 2651, 2687, 2786, 2836, 2891, 3148, 3161, 3171, 3177]. identification/classification [2651]. identified [2593]. identifiers [1645]. identifies [940]. identify [1640]. Identifying [44, 284, 1921]. Identity [490, 728, 822, 1287, 1361, 1839, 1914, 2089, 2322, 2483, 3068]. identity-augmented [822]. Identity-Based [728, 1914, 2483]. idle [3008]. IDS [630, 1163, 1871, 2952]. IEEE [132, 327, 364, 524, 594, 902, 1053, 1083, 1195, 1636, 1653, 2003, 2695, 2765]. IEoT [2338]. IFIP [1647]. IFMIF [1717]. IFMIF-DONES [1717]. igh [2883]. igh-dimensional [2883]. II [2080, 2488]. IIoT [585, 589, 629, 1702, 2160, 2436, 2559]. ILETC [1493]. illegal [2045]. ILP [417].

Image [117, 1863, 2432]. Image-based [117]. imagery [640]. images [55, 2351]. imaging [2315]. imbalance [630, 1964]. imbalanced [235, 2601]. IMCFN [117]. imitation [1317, 2643]. Impact [33, 136, 214, 369, 386, 515, 516, 698, 925, 1077, 1084, 1487, 1754, 1909, 1950, 2115, 2117, 2180, 2218, 2277, 2382, 2402, 2410, 2892, 3125]. impairment [724]. impairment-aware [724]. impairments [2103, 2921]. imperfect [521, 1131, 1480, 1736, 2589, 3154]. implantable [2377]. Implementability [1707]. Implementation [10, 36, 402, 408, 505, 701, 720, 1442, 1451, 1492, 1695, 1801, 1847, 1965, 1972, 2049, 2238, 2330, 2362, 2639, 2816]. implementation-agnostic [1965]. implementations [2214]. Implementing [1307, 1313, 1975]. implications [803, 1355, 2115]. importance [259, 1201, 1248, 1868, 1978]. important [2611]. Improve [99, 377, 851, 1130, 1513, 1782, 2133, 2623, 3090]. Improved [28, 75, 155, 259, 426, 630, 680, 799, 1049, 1129, 1286, 1334, 1841, 2047, 2080, 2170, 2198, 2217, 2293, 2297, 2476, 2737, 2904, 3029, 3063]. improvement [795, 1295, 1707]. improvements [1500]. Improving [248, 272, 378, 536, 768, 818, 884, 897, 1140, 1164, 1354, 1719, 1729, 1897, 2124, 2391, 2435, 2452, 2485, 2615, 2865, 2908, 2923, 2931, 2940, 3164]. Impulse [121]. imputation [2676]. IMSI [675]. In-band [493, 1228, 1233, 1471, 1830, 1988, 2082, 2292, 2611, 2983, 3114]. In-depth [1355, 2168]. in-kernel [1963]. in-kernel/user-space [1963]. In-network [1359, 1559, 1879, 2517, 2571, 2626, 2983, 2986, 3087]. in-protocol [2573]. In-Vehicle [2716, 2956]. Incentive [119, 173, 1156, 1671, 1761, 1859, 1970, 2065, 2232, 2341, 2454, 2520, 2746, 2805, 2866, 2896]. incentive-based [1671]. Incentive-driven [1761, 2454]. incentivization [2916]. Incentivizing [454, 2155, 2897]. incident [1411]. incomplete [1199, 1390, 2345]. inconsistencies [2115]. incorporating [1920, 2868, 3006]. Increased [138, 1676]. Increasing [1600, 2305]. Incremental [1493, 1497, 2120, 2286, 2788, 2838, 3013]. INCS [2337]. Independent [124, 362, 503, 517, 2109, 2668, 2904]. Index [1780, 2541]. indicators [1659, 2504, 3122]. indices [3082]. indistinguishability [1290]. Indoor [268, 437, 1124, 1848, 1896, 2004, 2325, 2503, 2807, 3091]. Indoor/Outdoor [1124]. induced [963, 1964]. induction [159, 350]. inductive [3130]. Industrial [56, 372, 407, 486, 517, 518, 584, 588, 595, 605, 633, 818, 819, 876, 1128, 1426, 1636, 1828, 1856, 2138, 2146, 2176, 2338, 2462, 2593, 2622, 2649, 2699, 2723, 2748, 2823, 3016, 3046]. industries [1331]. industry [1896, 2017, 2100, 2242]. inertial [783]. iNet [1442]. infeasibility [2014]. infection [1595]. inference [299, 368, 578, 1186, 1248, 1621, 1655, 1678, 1704, 1706, 1902, 2572, 2621, 2766, 3005, 3054, 3096, 3142]. inference-Sector [578]. Influence [1743, 1773]. Information [20, 135, 155, 208, 304, 528, 539, 641, 651, 655, 782, 914, 939, 958, 1017, 1084, 1094, 1130, 1149, 1199, 1232, 1246, 1278, 1356, 1390, 1458, 1787, 1884, 1965, 2086, 2123, 2124, 2161, 2189, 2218, 2515, 2612, 2764, 2792, 2873, 2955]. information-aware [2955]. Information-Centric [539, 939]. information-theoretic [2124]. informed [2169]. Informer [1779]. Informer-based [1779]. Infotainment [2956]. Infrastructure [85, 203, 457, 505, 784, 859, 877, 1266, 1605, 1636, 2127, 2174, 2330]. infrastructures [283, 1051]. initial [267, 1198]. initialization [680]. initiated [3101]. initiation [189]. Initiatives [802]. injection [1753, 2222]. innovation [213, 1364, 2019, 2989]. innovations [367, 506]. Innovative [2387, 2488, 3127]. input [651, 735]. inputs [783]. inRCL [2580]. insensitive [2802]. insider [503, 2434, 2968]. insightful [2168]. insights [2881, 3043]. inspection [2318]. inspired [280, 881, 978, 1736, 1784]. installation [685]. instance [158, 395, 1106, 1158, 2236]. instances [2495]. Instant [2097]. instantaneous [2657]. institution [1962]. INT [1741, 2983]. INTaaS [1988]. integral [196]. Integrated [79, 170, 288, 481, 528, 679, 909, 1013, 1019, 1034, 1046, 1251, 1447, 1576, 1666, 1670, 2076, 2250, 2463, 2467, 2486, 2507, 2538, 2587, 2617, 2658, 2659, 2709, 2867, 2982, 3106, 3117, 3127, 3136, 3165]. Integrating [1924, 2405, 2541, 2545, 2784]. integration [194, 2066, 2330, 2547, 2548, 2764, 2870, 2941, 3044, 3050]. integrity [355, 842, 916, 1277, 1840, 2316]. integrity-preserving [1277]. intellectual [2514]. Intelligence [380, 399, 434, 548, 609, 696, 697, 811, 978, 1267, 1335, 1368, 1378, 1495, 1513, 1583, 1602, 1747, 1938, 2027, 2044, 2162, 2254, 2374, 2457, 2514, 2517, 2564, 2576, 2997, 3076, 3117]. intelligence-based [380]. Intelligent [98, 318, 345, 380, 437, 453, 460, 470, 496, 520, 550, 731, 863, 866, 942, 1007, 1052, 1289, 1298, 1359, 1412, 1442, 1534, 1544, 1576, 1600, 1631, 1661, 1701, 1702, 1793, 1846, 1934, 1958, 1980, 1986, 1991, 2009, 2035, 2048, 2196, 2262, 2389, 2438, 2439, 2444, 2544, 2606, 2615, 2640, 2878, 3014, 3039, 3093, 3168]. Intent [1121, 1374, 1395, 1540, 2337, 2627, 2720, 3112, 3123, 3128]. Intent-based [1121, 1540, 2627, 3128]. Intent-driven [1395, 2337, 3112]. intention [2643]. intention-driven [2643]. intents [1171]. inter [339, 459, 471, 731, 950, 1013, 1199, 1798, 1980, 2137, 2152, 2185]. inter-BBU [339]. inter-cluster [459]. inter-constellation [1013]. inter-datacenter [950, 1199, 1798]. inter-domain [1980]. inter-flow [2137]. inter-numerology [731]. inter-server [471]. inter-slice [2152]. interaction [1581]. interactions [861, 2997]. interactive [1118]. interactivity [1118]. Interception [2238]. interdomain [86, 756]. Interest [6, 1165, 2644]. interests [418, 511, 2859]. interface [149, 884, 1527, 1667]. interfaces [996]. Interference [33, 420, 731, 732, 747, 1680, 1939, 1955, 2039, 2209, 2706, 2981, 3074]. interference-based [2039]. interfering [132]. Interlayer [936]. interleaved [1814]. intermittent [2113]. Internal [45, 1326]. International [3012]. Internet [160–163, 175, 183, 191, 207, 246, 290, 294, 306, 386, 433, 445, 447, 450, 487, 495, 517, 518, 521, 548, 582, 584, 604, 633, 765, 770, 787, 834, 853, 864, 876, 886, 894, 930, 938, 939, 941, 960, 964, 1005, 1018, 1039, 1128, 1154, 1156, 1161, 1296, 1320, 1326, 1333, 1398, 1418, 1470, 1513, 1538, 1583, 1652, 1697, 1784, 1787, 1820, 1829, 1856, 1890, 1912, 1923, 1930, 1959, 1964, 2061, 2100, 2119, 2138, 2147, 2176, 2209, 2216, 2219, 2246, 2275, 2330, 2366, 2442, 2454, 2525, 2536, 2570, 2633, 2640, 2741, 2744, 2748, 2775, 2777, 2815, 3068, 3125, 3140, 3171]. Internet [20, 157, 249, 258, 276, 412, 425, 453, 465, 660, 885, 943, 1248, 1315, 1436, 1568, 1676, 1721, 1738, 1793, 1802, 1841, 1854, 1919, 1991, 2080, 2175, 2284, 2311, 2328, 2471, 2610, 2691, 2850, 2873, 2896, 2991, 3046, 3070, 3175]. Internet-of-Things [258, 487, 1912, 2147, 2219]. Internet-scale [938]. Internet-wide [943]. interoperability [856, 2352]. interoperable [1474, 1751]. interpretability [737]. Interpretable [2269, 2770]. interpretation [2834]. interrupt [984]. interval [342]. interworking [876]. IntOpt [1228]. Intra [459, 1580, 2137, 2185]. Intra- [459, 2137]. Intra-body [1580]. Intraflow [2581]. Introducing [1941, 2160]. Introduction [1407, 1583, 1647]. Intrusion [61, 141, 144, 182, 235, 282, 317, 438, 442, 464, 495, 532, 610, 630, 669, 715, 751, 766, 910, 915, 941, 978, 1080, 1170, 1188, 1201, 1278, 1530, 1542, 1594, 1685, 1793, 1820, 1824, 1829, 1857, 1871, 1934, 1941, 2101, 2212, 2266, 2281, 2404, 2701, 2736, 2784, 2788, 2796, 2823, 2838, 2894, 2952, 3004, 3047, 3103, 3115, 3118, 3119, 3155, 3162]. intrusions [1660]. invariant [141]. invasive [2894]. Inventory [2551]. inverse [1870, 2459]. Inversion [1950]. investigation [784, 2324, 2956]. INVISIBLE [2039]. invocation [1922]. IoBT [2777]. IoD [604, 707, 1919, 2175]. IoE [2576]. IoE-edge [2576]. IoEPM [2261]. IoHT’ [2534]. IoMT [1167, 2153, 2209, 2827, 3146]. IoT [58, 145, 176, 178, 209, 242, 253, 254, 279, 291, 447, 450, 486, 490, 526, 543, 590, 606, 620, 627, 685, 768, 819, 840, 901, 917, 947, 976, 990, 1007, 1041, 1071, 1124, 1140, 1166, 1192, 1270, 1288, 1366, 1367, 1405, 1431, 1444, 1474, 1475, 1483, 1523, 1527, 1533, 1543, 1569, 1604, 1612, 1615, 1660, 1685, 1692, 1767, 1820, 1825, 1828, 1839, 1852, 1873, 1924, 1931, 1945, 2003, 2062, 2069, 2202, 2263, 2361, 2398, 2500, 2502, 2529, 2601, 2620, 2662, 2709, 2763, 2792, 2808, 2832, 2838, 2842, 2870, 2889, 2902, 2936, 3044, 3047, 3050, 3059, 3157]. IoT [3, 46, 62, 107, 202, 241, 267, 355, 358, 372, 389, 437, 444, 481, 488, 503, 533, 543, 588, 595, 605, 648, 671, 771, 809, 817, 818, 827, 838, 846, 863, 908, 909, 961, 1016, 1020, 1029, 1088, 1123, 1155, 1193, 1224, 1243, 1283, 1303, 1335, 1336, 1338, 1373, 1478, 1494, 1536, 1542, 1585, 1594, 1716, 1721, 1729, 1751, 1796, 1834, 1871, 1878, 1906, 1949, 1977, 2024, 2102, 2121, 2165, 2253, 2261, 2367, 2412, 2449, 2455, 2501, 2622, 2659, 2666, 2672, 2698, 2702, 2764, 2776, 2804, 2819, 2847, 2869, 2876, 2890, 2937, 2943, 3022, 3031, 3043, 3048, 3101, 3102, 3172, 3174]. IoT [317, 643, 882, 1149, 1183, 1188, 1409, 1746, 1805, 1937, 2095, 2215, 2556, 2788, 2823]. IoT-based [46, 279, 437, 590, 838, 976, 990, 1303, 2202, 2620]. IoT-blockchain [1288]. IoT-ecosystem [909]. IoT-enabled [62, 2455]. IoT-FoG [1615, 1906, 3044]. IoT-Gov [1721]. IoT-networks [488]. IoT-oriented [588, 2412]. IoT/WoT [2889]. IOTA [819, 2085]. IoTGeM [3102]. IoTs [721, 927, 1036, 2273, 2331, 2408]. IoV [642, 1165, 1307, 2266, 2721, 2866, 2981, 3185]. IoVs [599]. IP [370, 478, 610, 949, 1234, 1769, 1948, 2350, 2463, 2645, 2984]. IP-network [1234]. IP/MPLS [478]. IPFS [1998, 2767, 2819]. IPO [3014]. IPO-ZTA [3014]. IPro [98]. IPsec [1724, 2307, 2400]. IPv4 [1418]. IPv4/IPv6 [1418]. IPv6 [943, 1418, 1797, 2758]. iRED [2082]. IRIS [1546]. Irregular [818, 1748]. irrigation [47]. IRS [1809, 2142, 2510, 2584, 2829, 3166]. IRS-assisted [1809, 2142, 3166]. IRS-enabled [2584]. IRS-UAV [2829]. IRSA [1625]. ISAC [3159]. ISAC-aided [3159]. ISCC [1195]. Island [1535]. isolated [1039, 2783]. isolation [114, 899, 1326, 2136, 2375, 3093]. ISP [123, 1477]. ISPPFL [2341]. Issue [662, 1583, 1647, 3012]. issues [81, 284, 581, 1022, 1456, 1509, 2484]. ITA [2418]. ITA-NTN [2418]. ITD [3124]. item [341]. iteration [2841]. Iterative [1616, 1734]. itineraries [2747]. IXPs [368].

Jailhouse [2136]. Jain [1798]. jammer [738, 3084]. Jamming [439, 503, 1053, 1173, 1272, 1733, 1957, 2022, 2205, 2489, 3020]. January [26, 449, 929, 1416, 1926, 2567]. JANUS [2096]. JANUS-based [2096]. JAPPI [2297]. JC [3060]. job [517, 587, 2550]. job-shop [587]. jobs [441, 1678]. joining [2069]. joins [3082]. Joint [14, 22, 54, 192, 233, 359, 365, 366, 377, 416, 440, 573, 603, 628, 638, 733, 755, 757, 785, 831, 854, 897, 919, 1034, 1036, 1064, 1065, 1099, 1111, 1118, 1161, 1178, 1229, 1268, 1321, 1340, 1400, 1450, 1485, 1522, 1536, 1573, 1626, 1627, 1663, 1666, 1728, 1746, 1760, 1831, 1857, 1858, 1874, 1887, 1895, 1913, 1918, 1936, 1954, 2004, 2009, 2063, 2149, 2235, 2296, 2386, 2408, 2409, 2448, 2456, 2459, 2466, 2471, 2507, 2510, 2544, 2589, 2602, 2610, 2629, 2636, 2657, 2673, 2746, 2790, 2835, 2840, 2841, 2846, 2898, 2919, 2944, 2951, 3021, 3045, 3056, 3067, 3071, 3081, 3096, 3168]. JointCloud [2417, 3060]. JRC [2951]. JT [2421]. JT-CoMP [2421]. Judging [982]. July [201, 221, 665, 688, 1110, 1137, 1650, 1674, 2241, 2927, 2948]. June [185, 636, 1092, 1614, 1635, 2157, 2194, 2221, 2880, 2912]. JurisNN [982]. Just [2860].

Ka-band [1835]. Kalman [2393]. kangaroo [442]. kangaroo-based [442]. Keeping [47]. Kernel [469, 607, 2192, 3133]. Kernel-level [607]. Kernel/eBPF [469]. kernel/user [1963]. Key [191, 287, 291, 326, 375, 400, 728, 772, 794, 833, 1019, 1146, 1222, 1240, 1334, 1503, 1659, 1813, 2030, 2031, 2153, 2810, 2817, 2827, 3030, 3038, 3051, 3057, 3068, 3069, 3122, 3169]. key-agreement [400]. key-aided [3030]. key-value [2817]. keypads [2109]. keys [2030]. KeySFC [30]. keystroke [2109]. keyword [1851, 2239]. KFTO [689]. kicking [39]. kite [2737]. Knacks [1557]. Knowing [91, 2329]. Knowledge [214, 873, 1827, 1834, 1838, 2259, 2373, 2435, 2504, 2680, 2717, 2985, 3060]. Knowledge-based [2717]. known [143]. Kolmogorov [2622]. KP [2794]. KP-GNN [2794]. Krill [835]. KTWIN [2738]. Kubernetes [2738]. Kubernetes-based [2738]. Kuhn [689].

L [445, 1095, 2944]. L-band [1095]. L-RPL [445]. L2 [2061]. L2-BiTCN-CNN [2061]. L3TM [2945]. LA-Trickle [725]. lab [2816]. Label [718, 1150, 1682, 2411, 2445, 2674, 2980]. labeled [1025, 1560, 2339]. labelled [998]. Lagrange [35]. LAKA [1503]. LAKA-UAV [1503]. LAN [727]. Language [788, 2291, 2905, 2946, 2952, 3025, 3107, 3169]. LANs [34, 2429]. LAPCHS [519]. Laplacian [445]. Large [18, 50, 56, 278, 386, 427, 491, 768, 829, 891, 1001, 1319, 1435, 1451, 1512, 1559, 1881, 2009, 2127, 2248, 2437, 2465, 2502, 2845, 2860, 2905, 2934, 2946, 2952, 3003, 3025, 3107, 3161, 3169]. large-scale [56, 278, 386, 768, 1001, 1319, 1435, 1512, 1559, 1881, 2009, 2248, 2437, 2465, 2502, 2845, 2860, 3003, 3161]. laser [1411]. last [2437]. last-hop [2437]. lasting [1399]. latencies [343]. Latency [31, 85, 283, 353, 359, 386, 443, 466, 549, 567, 659, 769, 938, 953, 984, 1003, 1010, 1032, 1051, 1138, 1230, 1260, 1265, 1446, 1487, 1577, 1750, 1782, 1823, 1960, 1961, 2009, 2150, 2172, 2184, 2293, 2302, 2331, 2438, 2598, 2664, 2684, 2685, 2829, 2855, 2971, 3104, 3111, 3144, 3165]. latency-aware [283, 938, 1051]. latency-critical [1138]. Latency-energy [359, 2684]. Latency-Guaranteed [549]. latency-sensitive [1265]. Latent [2749]. lateral [2800]. Latin [747]. Lattice [1185, 1283, 2265, 2530, 2575, 2617, 2780]. Lattice-based [1185, 2265, 2530, 2575, 2780]. Lawful [2238]. laws [3129]. Layer [19, 99, 112, 371, 450, 598, 603, 617, 667, 781, 880, 1030, 1069, 1121, 1394, 1444, 1546, 1825, 1848, 1898, 2027, 2072, 2086, 2190, 2294, 2391, 2418, 2490, 2498, 2529, 2600, 2610, 2644, 2677, 2714, 2725, 2768, 2773, 2823, 2851, 2871, 2884, 2896, 3137, 3138]. Layer-based [2884]. layer-wise [2768, 2851]. layered [282, 800, 1427, 1757, 1956, 2945, 3151]. layering [596, 2925]. layers [884]. LBSs [2859]. LCA [2957]. LDA [1478]. LDA-2IoT [1478]. LDoS [2357]. LDPC [2428, 2854]. leader [2841]. Leaderless [1348, 1454]. leadership [2372]. leak [1098]. Leakage [155, 728, 1239, 3051]. Leakage-Resilient [728, 3051]. learned [788, 918, 2639]. Learning [163, 274, 291, 333, 345, 401, 430, 433, 436, 446, 556, 571, 611, 639, 749, 751, 790, 839, 841, 844, 857, 870, 880, 905, 926, 927, 941, 957, 960, 1037, 1045, 1065, 1112, 1125, 1194, 1249, 1310, 1315, 1339, 1369, 1442, 1448, 1474, 1475, 1495, 1502, 1507, 1537, 1583, 1648, 1670, 1704, 1731, 1762, 1820, 1829, 1961, 2012, 2070, 2071, 2082, 2087, 2101, 2120, 2121, 2150, 2232, 2258, 2270, 2280, 2341, 2361, 2433, 2445, 2468, 2470, 2502, 2506, 2547, 2550, 2571, 2606, 2624, 2701, 2730, 2746, 2753, 2805, 2826, 2838, 2928, 2957, 2987, 3000, 3013, 3046, 3047, 3097, 3104, 3112, 3170]. learning [61, 311, 327, 354, 495, 532, 538, 551, 618, 627, 633, 672, 695, 697, 715, 744, 759, 832, 906, 915, 1170, 1215, 1226, 1282, 1309, 1316, 1353, 1366, 1372, 1427, 1470, 1505, 1531, 1532, 1555, 1559, 1573, 1612, 1623, 1631, 1664, 1752, 1795, 1804, 1837, 1872, 1882, 1889, 1891, 1931, 1933, 1976, 1988, 2015, 2023, 2050, 2058, 2065, 2066, 2106, 2112, 2138, 2163, 2169, 2181, 2187, 2201, 2281, 2329, 2337, 2435, 2521, 2572, 2597, 2601, 2640, 2671, 2705, 2709, 2741, 2755, 2763, 2777, 2786, 2819, 2841, 2842, 2874, 2897, 2910, 2919, 2935, 2949, 2961, 3011, 3067, 3080, 3094, 3115, 3155, 3156, 3160, 3173, 3180, 3182, 3185]. learning [4, 15, 222, 371, 385, 387, 403, 465, 497, 529, 640, 765, 824, 866, 896, 916, 961, 973, 1010, 1021, 1032, 1033, 1047, 1061, 1062, 1067, 1088, 1120, 1155, 1182, 1206, 1237, 1298, 1317, 1320, 1321, 1344, 1364, 1375, 1389, 1394, 1463, 1494, 1511, 1529, 1595, 1629, 1657, 1677, 1702, 1707, 1709, 1723, 1739, 1753, 1756, 1761, 1772, 1789, 1794, 1797, 1870, 1879, 1953, 2037, 2052, 2170, 2190, 2209, 2231, 2235, 2237, 2262, 2274, 2287, 2344, 2362, 2365, 2381, 2387, 2390, 2459, 2496, 2525, 2570, 2605, 2614, 2700, 2798, 2899, 2937, 2938, 2950, 2960, 2985, 2998, 3009, 3015, 3016, 3030, 3053, 3072, 3084, 3130]. learning [27, 54, 74, 152, 211, 252, 272, 377, 477, 535, 587, 605, 643, 648, 718, 729, 737, 755, 801, 830, 955, 959, 967, 981, 1002, 1086, 1202, 1216, 1231, 1265, 1292, 1406, 1408, 1414, 1467, 1493, 1515, 1525, 1566, 1570, 1599, 1661, 1712, 1732, 1799, 1805, 1832, 1898, 1947, 1970, 2006, 2010, 2020, 2021, 2073, 2084, 2115, 2125, 2130, 2151, 2183, 2212, 2224, 2257, 2282, 2291, 2324, 2385, 2491, 2537, 2554, 2559, 2561, 2579, 2589, 2619, 2622, 2625, 2632, 2641, 2643, 2646, 2696, 2710, 2734, 2750, 2761, 2766, 2768, 2788, 2793, 2834, 2850, 2866, 2873, 2909, 2942, 2964, 2981, 3020, 3036, 3039, 3048, 3049, 3058, 3061, 3085, 3099, 3153, 3172, 3175]. learning [217, 799, 1028, 1128, 1626, 1824, 1856, 1995, 2159, 2245, 2288, 2376, 2382, 2420, 2799, 2872, 2932, 2940, 3024]. Learning-aided [1125]. Learning-Based [345, 371, 387, 611, 627, 648, 790, 1310, 1364, 1394, 1470, 1525, 1532, 1623, 1629, 1631, 1756, 1797, 1805, 1933, 2130, 2138, 2151, 2209, 2231, 2270, 2525, 2624, 2671, 2700, 2730, 2788, 2919, 2957, 2960, 2981, 3030, 3047, 3094, 3112, 3160, 3185]. Learning-driven [15, 211, 495, 2280, 2547, 2646]. learning-enabled [2281]. learning-guided [2390]. learning-powered [2058]. LearningTuple [2424]. learnt [1829]. leasing [236]. least [28]. least-squares [28]. LEDBAT [1122, 1376]. ledger [639]. legs [1936]. length [33, 2272]. lens [1327, 2458, 2984]. lenses [3022]. LEO [150, 1107, 1191, 1535, 1822, 2025, 2384, 2437, 2538, 2677, 2730, 2762, 2782, 2875, 2987, 3108, 3164]. LEO-based [1822, 2384]. Less [91, 306, 1376, 1695, 1900, 2364, 3086]. LESS-ON [2364]. less-than-best-effort [1695]. lessons [788, 1829, 2639]. Let [2045]. Level [187, 351, 352, 411, 414, 461, 585, 607, 793, 870, 926, 1273, 1323, 1369, 1426, 1461, 1478, 1576, 1619, 1685, 1785, 1813, 1911, 1929, 1941, 2023, 2052, 2171, 2299, 2405, 2483, 2688, 3027]. Leveraging [485, 517, 954, 1938, 2146, 2312, 2458, 2482, 2601, 2765]. LFF [2160]. LFGurad [2445]. Libra [2986]. Licensed [297]. licensing [2760]. LiDAR [1424]. Life [207, 1062, 1460]. lifecycle [1839]. lifelong [2909]. Lifetime [180, 1074, 2455]. Light [845, 1054, 1411, 1448, 1619, 1675, 1759, 2039]. light-weight [845, 1448]. LightGBM [2479]. lighting [437]. LightLog [934]. lightning [772, 2060]. Lightpath [1894]. lightpaths [1112]. LightPoW [1398]. Lightweight [36, 101, 242, 250, 286, 355, 490, 519, 647, 728, 840, 842, 843, 934, 946, 973, 1224, 1283, 1300, 1326, 1403, 1461, 1503, 1508, 1544, 1548, 1606, 1786, 1834, 1878, 1893, 1919, 1963, 2110, 2253, 2261, 2312, 2333, 2397, 2398, 2401, 2517, 2596, 2776, 2827, 3040, 3048, 3068, 3173]. like [277]. LiKey [2109]. LiM [207]. LiM-AHP-G-C [207]. Limitations [64, 1777]. limited [90, 728]. limits [515, 1617]. Line [1592, 2478, 2669]. linear [44, 560, 1238, 1401, 1534, 2465]. linguistics [431]. linguistics-based [431]. Link [13, 29, 327, 445, 459, 660, 679, 681, 850, 936, 1321, 1347, 1680, 1812, 1868, 1903, 1960, 1987, 2006, 2124, 2292, 2486, 2734, 2761, 2791, 3019, 3153]. link-Configuration [327]. link-state [1347]. linkable [1849]. Linkage [2505]. links [446, 963, 1618, 1968, 1971, 2079, 2599, 2706, 2713]. Linux [428, 607]. LIO [630]. LIO-IDS [630]. lion [743]. list [153]. listening [3008]. literature [107, 978, 1188, 1395, 2872, 2929]. Live [248, 751, 865, 1063, 1388, 1616, 1782, 1886, 2198, 2317, 2416, 2898, 3018, 3029]. Live360 [2317]. liveness [2018]. livestock [56]. living [2816]. Llama [3066]. LLDP [3075]. LLM [2789]. LLMs [2683]. LLNs [3059]. LM [2800]. LM-Hunter [2800]. Load [77, 102, 222, 261, 286, 339, 439, 447, 498, 543, 563, 601, 680, 793, 908, 987, 1035, 1084, 1183, 1184, 1210, 1401, 1522, 1572, 1620, 1701, 1732, 1891, 1952, 1964, 2152, 2247, 2293, 2364, 2474, 2489, 2539, 2551, 2616, 2676, 2699, 2731, 2970, 2991]. Load-aware [439, 1184, 1732, 2247, 2364]. Load-balanced [563, 1701]. Load-balancing [1035, 1522]. LOADER [429]. loading [2576]. loads [1401, 1785]. Local [123, 305, 429, 638, 756, 1533, 1844, 2266, 2491, 2516, 2560, 2693, 2760, 2841, 3119]. localisation [321, 1992]. locality [886, 954, 2287]. locality-sensitive [2287]. Localization [28, 66, 113, 214, 289, 293, 821, 1079, 1124, 1354, 1543, 1558, 1806, 2320, 2384, 2517, 2580, 2669, 2737, 3034, 3091]. Locally [3105, 3141]. Locating [578, 853]. Location [65, 122, 352, 419, 594, 621, 660, 738, 810, 859, 1068, 1159, 1239, 1340, 1393, 1491, 1990, 2064, 2109, 2162, 2268, 2495, 2516, 2732, 2807, 2887]. location-authentication [122]. Location-aware [65, 2732]. location-based [1159]. Location-independent [2109]. location-related [1068]. LoCoNOMA [2531]. LOF [3162]. log [934, 2371, 2918, 3023]. Logchain [197]. Logging [610]. Logic [497, 734, 1598]. logic-based [1598]. logistic [1841]. logits [2874]. LOGO [3061]. LOGO-CL [3061]. logs [197, 411, 2729]. Long [658, 1399, 1469, 1546, 1682, 1761, 2225, 2245, 2990, 3111]. Long-lasting [1399]. long-range [1546]. long-term [1469, 1761, 2245, 3111]. Longest [1244]. Longitudinal [94]. looking [2238]. looking-forward [2238]. lookup [341, 1908]. loop [1187, 1418, 2997]. loops [1418, 1444]. LoRa [1413, 1431, 1881, 1972, 2503, 3045]. LoRaWAN [384, 485, 920, 1225, 1244, 1435, 1522, 1689, 1783, 2044]. LoRaWAN-based [1689]. loss [462, 658, 2663, 2665, 2808, 3173]. losses [2410]. Losslessly [1352]. Lossy [286, 700, 2665, 3059]. Low [31, 109, 286, 322, 342, 466, 482, 513, 659, 711, 721, 723, 752, 763, 769, 877, 980, 1016, 1071, 1093, 1260, 1300, 1306, 1446, 1471, 1500, 1546, 1604, 1646, 1687, 1750, 1782, 1823, 1861, 1881, 1960, 1968, 1992, 2086, 2150, 2226, 2287, 2399, 2438, 2599, 2664, 2693, 2709, 2809, 2820, 2971, 3059, 3104, 3109, 3110, 3165]. low-altitude [1861]. Low-AoI [2709]. low-bitrate [3109]. Low-complexity [2226]. Low-cost [711, 1300, 1604, 2399]. Low-delay [763, 1016, 1500, 1968, 2086]. low-infrastructure [877]. Low-latency [769, 1260, 1446, 1782, 1960, 2150, 2438, 2971, 3165]. low-overhead [3110]. Low-power [286, 322, 1881, 3059]. Low-rate [2599]. low-resource [2820]. low-threshold [2809]. lowest [1562]. LPB [1903]. LPPAC [2596]. LPulse [2406]. LPWAN [291, 362]. LPWANs [492]. LR [144, 678, 2659]. LR-FHSS [2659]. LR-PON [678]. LRS PKI [1849]. LRU [1984]. LRU-based [1984]. LS [2372]. LSA [44]. LSTM [390, 630, 1160, 1785, 1958, 2204, 2784, 2958, 3062]. LSTM-based [1160, 1785, 2204]. LSTM/GRU [2236]. LTE [38, 76, 154, 224, 271, 413, 563, 797, 884, 1045, 1590, 2000, 2483, 2687]. LTE-A [76]. LTE-U [797]. LTE-WiFi [38]. LTE/5G [1590, 2483]. LTE/WLAN [413]. LWSBFT [1348]. Lyapunov [979, 2237, 2288]. Lyapunov-guided [2288].

M [340]. M3F [1584]. MA [612]. MA-ABE [612]. MAB [3090]. MAC [112, 124, 786, 901, 998, 1083, 1328, 1352, 1676, 2167, 2542, 2693]. MAC-agnostic [2167]. MAC-layer [112]. Machine [15, 163, 174, 333, 345, 354, 377, 387, 388, 430, 433, 446, 497, 529, 532, 535, 556, 611, 618, 672, 715, 739, 751, 844, 857, 915, 1028, 1033, 1045, 1062, 1063, 1065, 1170, 1231, 1298, 1303, 1339, 1448, 1507, 1583, 1595, 1626, 1648, 1664, 1702, 1704, 1709, 1882, 1889, 1931, 2010, 2058, 2169, 2258, 2281, 2287, 2362, 2468, 2547, 2671, 2696, 2734, 2741, 2793, 2871, 2928, 2949, 3038, 3094, 3156]. Machine-Learning [1339, 1595]. Machine-type [174, 2871, 3038]. machines [1935, 2654]. macro [579]. macro-femto [579]. MADDPG [2040]. made [284]. MADRL [2471]. MADRLOM [2108]. MAESP [363]. MAGD [1819]. Magnetic [159, 350, 1282]. Magteon [2945]. Magteon-Turing [2945]. MAID [2792]. Maintaining [623, 631, 3163]. maintenance [2325]. major [364]. Making [554, 612, 1552, 1593, 2019, 3086]. MalCon [1697]. Malicious [277, 758, 906, 926, 960, 1236, 1345, 1737, 1756, 1773, 1837, 2069, 2192, 2222, 2263, 2282, 2291, 2306, 2391, 2482, 2493, 2519, 2674, 2689, 2772, 2803, 2843, 2881, 3060, 3078]. Malware [42, 72, 117, 564, 801, 961, 1005, 1279, 1584, 1697, 1788, 2208, 2269, 2469, 2513, 3179, 3180]. man [3125]. man-in-the-middle [3125]. Management [3, 17, 19, 39, 78, 91, 118, 160, 210, 246, 274, 315, 318, 325, 357, 362, 375, 483, 514, 567, 589, 612, 624, 633, 641, 652, 654, 670, 734, 772, 778, 829, 833, 858, 861, 866, 876, 894, 930, 1062, 1115, 1146, 1191, 1240, 1266, 1355, 1374, 1395, 1521, 1547, 1637, 1755, 1775, 1813, 1839, 1854, 1856, 1867, 1895, 1924, 1939, 2004, 2010, 2153, 2215, 2216, 2285, 2311, 2349, 2427, 2439, 2473, 2561, 2715, 2724, 2793, 2815, 2981, 2990, 3052, 3128, 3146, 3164, 3169, 3186]. manager [60]. Managers [1266]. Managing [667]. MANET [1914]. manifolds [1263]. manipulators [1574]. MANO [489]. manufacturing [587, 588]. Many [188, 347, 1212]. many-field [347]. many-objective [1212]. Many-to-many [188]. map [1151, 1208, 2176, 2305, 2325, 2637, 2922, 3001]. map-aware [1151]. MAP-T [2637]. MAP/PH/1/n [2922]. Mapping [919, 1210, 1242, 2718, 2984]. MapReduce [2166]. Maptperf [2637]. March [88, 523, 1000, 1482, 2002, 2703]. Marconi [1442]. margin [1184]. maritime [565, 2102, 2242, 2546, 2850, 3174]. mark [767]. market [549, 2370]. market-based [2370]. market-oriented [549]. marketplace [974, 1370]. markets [1459]. marking [370, 767]. Markov [316, 370, 392, 1008, 2025, 2691]. Markovian [79, 1362]. MARL [2255]. Mars [1147]. MARVEL [222, 2121]. mask [2839]. Massive [76, 205, 267, 394, 433, 896, 908, 1116, 1229, 1770, 2428, 2531, 2830, 2871]. Massive-Scale [394, 1116]. match [2019]. match-action [2019]. Matching [51, 188, 413, 1242, 1258, 1311, 1679, 2267, 2353, 2355, 3158]. matchings [1946]. Matchmaker [623]. MATE [2159]. MATEC [843]. materialized [942]. Mathematical [304, 392, 696, 2542]. MATILDA [667]. Matrix [891, 2029, 2642, 2854]. matter [2060]. max [1592]. maximal [2032]. Maximization [38, 202, 559, 705, 862, 1032, 1074, 1193, 1274, 1381, 1743, 1997, 2092, 2230, 2433, 2455, 2486, 2604, 3077]. maximize [1085, 1522, 1803, 2993, 3158]. Maximizing [72, 207, 583, 1190, 1796, 1869, 3181]. Maximum [133, 1420, 2667]. May [147, 167, 592, 614, 1058, 1076, 1565, 1588, 1759, 1990, 2105, 2140, 2812, 2849]. MAYA [1452]. MC [1032, 2519]. MC-Det [2519]. MC-NOMA-based [1032]. MCDC [2843]. MCEN [1420]. MCS [1321, 2162, 2334]. MDCHD [801]. MDFD [1499]. MDN [2463]. MDP [1210]. means [2975]. measure [75, 158, 957, 3124]. Measurement [466, 629, 677, 683, 896, 1148, 1418, 1563, 1741, 1847, 2081, 2117, 2188, 2392, 2600]. measurement-based [2081]. Measurement-driven [677]. measurements [134, 227, 738, 1189, 1791, 2075, 2274]. measures [2702]. Measuring [515, 1127, 1659, 1964, 2697]. MEC [15, 131, 145, 194, 313, 539, 830, 1064, 1130, 1206, 1306, 1371, 1485, 1510, 1607, 1677, 1702, 1727, 1954, 2091, 2141, 2149, 2280, 2314, 2472, 2507, 2538, 2552, 2646, 2673, 2756, 2829, 2878, 2917, 2950, 2961, 2979, 3105, 3111, 3144, 3181]. MEC-assisted [2091]. MEC-based [313, 1130]. MEC-enabled [15, 539, 1306]. MEC-IIoT [1702]. MEC-NFV [2756]. Mechanism [80, 119, 199, 225, 285, 297, 406, 470, 543, 568, 685, 693, 766, 809, 826, 913, 914, 1149, 1156, 1286, 1296, 1314, 1324, 1336, 1390, 1494, 1555, 1658, 1823, 1912, 1918, 1970, 2065, 2108, 2200, 2202, 2323, 2463, 2466, 2475, 2491, 2494, 2520, 2613, 2668, 2731, 2746, 2805, 2836, 2866, 2869, 2896, 2916, 3053, 3055, 3137, 3140, 3185]. mechanism-driven [2866]. mechanisms [173, 282, 1285, 1459, 1740, 1857, 2446, 2529, 2726, 2824, 3151]. Media [362, 1353]. mediated [2712]. Medical [355, 875, 965, 1104, 1115, 1872, 2119, 2209, 2620, 2666, 2847]. medium [836, 927, 2348]. Meet [1652, 1979, 2977]. meets [27, 142, 618, 866, 1226, 2159, 2572]. mega [1911, 2730]. mega-constellations [2730]. MEH [1954]. membership [953]. MEMENTO [2120]. memetic [239]. Memory [658, 723, 762, 792, 1063, 2217, 2857]. MemOry-based [762]. Memory-efficient [2857]. Merkle [2999]. Mesh [693, 1100, 1162, 1971, 2388, 2607, 2907, 3129]. Message [83, 89, 835, 1157, 2410, 2716, 3072]. messages [273, 379, 1836, 2483, 2615, 2889]. messaging [1335]. Meta [1599, 2020, 2269, 2472, 2541, 3080, 3179]. Meta-CAD [3080]. meta-features [2269]. Meta-Heuristic [2472, 2541]. meta-learning [1599, 3080]. meta-path [3179]. metadata [2708]. Metaheuristic [61, 699, 714, 753, 1008, 1375, 1832]. metaheuristic-based [699]. metaNet [2269]. Metaversal [2997]. metaverse [2185, 2341, 2473, 2818, 2822, 2846, 2962, 2997]. Method [151, 310, 382, 407, 456, 524, 584, 651, 657, 724, 737, 758, 782, 801, 819, 842, 889, 1047, 1223, 1233, 1303, 1331, 1457, 1505, 1593, 1611, 1633, 1663, 1696, 1729, 1769, 1779, 1786, 1795, 1813, 1817, 1852, 1893, 1915, 1958, 1962, 2011, 2037, 2101, 2106, 2177, 2262, 2263, 2268, 2331, 2363, 2381, 2404, 2426, 2497, 2503, 2541, 2553, 2556, 2599, 2642, 2656, 2708, 2720, 2800, 2873, 2901, 2915, 2924, 2950, 2971, 2978, 3007, 3023, 3085, 3110, 3133, 3148, 3152]. Methodology [677, 774, 1256, 1975, 2274, 2297, 2760, 3123]. Methods [368, 380, 430, 532, 560, 1589, 1814, 1930, 2548]. metric [1255, 1723, 2813]. metrics [565, 625, 2050, 2785]. metropolitan [2123]. metropolitan-wide [2123]. MFFusion [926]. MFO [2306]. MFO-based [2306]. MFSI [2891]. MFTE [2910]. MGGPT [3073]. MGKM [1240]. MGM [1182]. MGM-4-FL [1182]. MI [2834]. MI-VFL [2834]. MICN [511]. micro [2195, 2919]. micro-segmentation [2195]. MicroCM [1922]. microservice [1864, 1922, 1929, 2837]. microservice-based [1864]. microservices [507, 2038, 2427]. microservices-based [2427]. microwave [29, 1378]. middle [3125]. middlebox [656]. middleboxes [484]. MIF [782]. Migration [330, 339, 403, 668, 987, 1047, 1063, 1183, 1457, 1469, 1654, 1983, 2021, 2063, 2228, 2260, 2276, 2640, 2715]. migrations [931, 1219]. military [1544]. Millimeter [14, 786, 798, 963, 985, 1073, 1749, 2008]. millimeter-wave [14, 786, 798, 963, 1749]. MIMO [159, 316, 376, 521, 1229, 1243, 1607, 1770, 2428, 2574, 2587, 2694, 2830, 2913]. MIMO-NOMA [1243, 2830]. Min [1276, 1592]. min-max [1592]. mined [566]. mini [698]. mini-slots [698]. Minimal [969, 1819]. minimisation [995]. minimization [289, 642, 731, 969, 1873, 1976, 2021, 2129, 2331, 2507, 2552, 2829, 2862, 3039]. Minimize [446, 545, 2407]. minimized [188, 346]. Minimizing [286, 331, 850, 956, 1040, 1578, 1672, 1728, 2166, 2609, 2646, 3111, 3144]. Minimum [525, 661, 1181, 1211, 1347, 1667, 1734, 2187, 2293]. Minimum-cost [1734]. Mining [42, 227, 943, 1700, 2029, 2269, 2675]. Misconfiguration [2180]. mismatch [1993]. MISO [1809, 3168]. missing [2528, 3161]. mistrusting [285]. mitigate [1680, 2569]. Mitigating [984, 1909, 2258, 2532, 2808]. Mitigation [11, 80, 516, 608, 747, 816, 925, 1297, 1301, 1326, 1543, 1804, 2357, 2463, 2534, 2542, 2787, 2836, 2968, 3074, 3160]. mix [1821]. mix-zones [1821]. mixed [38, 1874, 2055, 2154, 2963]. mixed-critical [2154]. mixed-timescale [1874]. Mixer [2411]. mixture [235]. MJOA [1655]. MJOA-MU [1655]. ML [1089, 1101, 1410, 1483, 1816, 1850, 1985, 2152, 2180, 2316, 2943, 2980]. ML-aided [2230]. ML-based [1483, 1985, 2152, 2316, 2526]. MLP [2411]. MLP-Mixer [2411]. MMT [455]. MMT-based [455]. mMTC [385, 1010]. mmWave [186, 416, 729, 783, 988, 1526, 1661, 2141, 2205, 2353, 2489, 2908, 2951, 3041]. mmWave/sub [2908]. mmWave/sub-THz [2908]. mmWaves [140]. MobilE [22, 193, 332, 359, 390, 398, 451, 490, 496, 559, 574, 583, 590, 642, 697, 700, 701, 703, 773, 787, 823, 831, 885, 901, 956, 963, 970, 974, 977, 1003, 1015, 1041, 1145, 1190, 1212, 1256, 1273, 1340, 1355, 1361, 1369, 1382, 1419, 1427, 1432, 1443, 1455, 1476, 1484, 1496, 1506, 1510, 1548, 1552, 1555, 1561, 1590, 1596, 1623, 1633, 1710, 1711, 1744, 1758, 1791, 1859, 1881, 1910, 1953, 2045, 2056, 2072, 2106, 2113, 2167, 2169, 2199, 2201, 2238, 2269, 2296, 2313, 2331, 2370, 2433, 2461, 2466, 2494, 2510, 2555, 2574, 2590, 2630, 2644, 2667, 2790, 2792, 2806, 2842, 2870, 2906, 2910, 2917, 3015, 3030, 3062, 3158]. mobile [17, 91, 100, 119, 133, 139, 173, 211, 212, 227, 257, 266, 363, 366, 387, 395, 427, 436, 502, 586, 600, 645, 733, 755, 785, 821, 862, 870, 888, 896, 924, 981, 1040, 1061, 1077, 1106, 1130, 1216, 1312, 1381, 1429, 1437, 1439, 1472, 1473, 1479, 1545, 1592, 1654, 1722, 1732, 1760, 1764, 1862, 2020, 2021, 2142, 2262, 2270, 2288, 2365, 2383, 2483, 2604, 2605, 2679, 2686, 2700, 2726, 2759, 2814, 2820, 2901, 2933, 2964, 3099, 3130, 3144]. mobile-app [870, 1273, 1596]. Mobile-edge [642]. mobile-IoT [2842]. Mobility [78, 180, 229, 233, 279, 357, 362, 363, 369, 492, 539, 595, 887, 939, 1071, 1102, 1383, 1431, 1561, 1654, 1750, 1775, 1792, 1833, 2276, 2279, 2308, 2349, 2351, 2392, 2600, 2640, 2678, 2792, 2841, 2917, 3056, 3104, 3132, 3164]. Mobility-Aware [279, 1654, 1833, 2276, 2640, 2792, 2841, 3104]. mobility-awareness [2308]. mobility-compliant [939]. modal [2550, 2963, 3064]. Mode [315, 893, 1089, 1162, 1485, 1775, 2003, 2440]. Model [13, 39, 79, 132, 154, 183, 229, 260, 271, 340, 369, 462, 464, 488, 556, 582, 639, 732, 840, 841, 857, 868, 915, 926, 1008, 1072, 1132, 1182, 1268, 1336, 1367, 1370, 1430, 1448, 1451, 1507, 1533, 1545, 1557, 1581, 1618, 1636, 1655, 1686, 1730, 1770, 1792, 1800, 1810, 1832, 1848, 1886, 1894, 1903, 1949, 1965, 1982, 1984, 2061, 2107, 2115, 2137, 2209, 2226, 2257, 2357, 2360, 2372, 2469, 2476, 2553, 2558, 2574, 2594, 2658, 2691, 2749, 2777, 2784, 2786, 2787, 2791, 2834, 2839, 2861, 2864, 2877, 2922, 2952, 2957, 2963, 3004, 3009, 3017, 3025, 3037, 3062, 3064, 3075, 3076, 3081, 3130, 3137, 3142, 3181]. model [74, 235, 691, 1437, 1566, 1757, 1764, 1947, 2049, 2122, 2158, 2618, 2708, 2758, 2774, 2882, 3153]. model-based [2749, 3081]. model-free [1507, 1566]. Modeling [21, 149, 344, 370, 681, 773, 910, 1046, 1356, 1572, 1690, 1692, 1857, 1962, 1985, 2025, 2028, 2036, 2145, 2435, 2533, 2562, 2663, 2821, 2922]. Modelling [109, 329, 674, 964, 1264]. Models [89, 210, 226, 345, 509, 536, 565, 896, 1048, 1112, 1384, 1452, 1495, 1889, 2079, 2208, 2264, 2297, 2315, 2362, 2439, 2502, 2514, 2650, 2889, 2905, 2946, 2976, 3043, 3102, 3107, 3155, 3169]. models-aware [2650]. modern [3124]. modes [854]. Modified [726, 1008, 1301, 2737]. modular [55, 1801, 2890, 3119]. modulated [2428]. Modulation [77, 832, 1323, 1434, 1488, 1556, 2234, 3027]. Modulation-Level [1323, 3027]. module [1313, 2090]. molecular [1690]. Monero [992]. Monero-based [992]. monetary [248]. monitor [1228]. Monitoring [40, 98, 202, 223, 322, 417, 469, 593, 638, 679, 711, 800, 863, 924, 1066, 1209, 1270, 1547, 1689, 1922, 2219, 2258, 2261, 2441, 2476, 2976, 3106, 3174]. monitoring-cycle [417]. Monkey [1226]. Monte [1924]. Montgomery [2890]. months [1487]. morphological [1080]. MoTor [2920]. mounted [2510]. movements [2800]. MOVES [762]. moving [477, 494]. MP [644, 3134]. MP-Cos [3134]. MP-CREDIT [644]. MPC [2773]. MPC-in-the-Head [2773]. MPLS [478, 593]. MpScope} [2441]. MPTCP [2551]. MQTT [771, 1492, 1660, 1871]. MQTT-enabled [1871]. MR [1442]. MR-iNet [1442]. MRAI [1959]. MRAI-free [1959]. MSCCS [992]. MSDQ [2373]. MSN [2427]. MSTFCAN [2682]. MSYM [52]. MT [1323]. MTC [205, 2531]. MTCR [2772]. MTCR-AE [2772]. MTD [2968]. MU [376, 1607, 1655, 2694]. MU-MIMO [376, 1607, 2694]. MuHoW [2016]. MUlti [53, 65, 75, 97, 198, 232, 282, 340, 374, 413, 435, 457, 627, 644, 695, 778, 781, 800, 808, 814, 926, 1034, 1046, 1055, 1126, 1131, 1245, 1258, 1270, 1280, 1397, 1424, 1427, 1428, 1435, 1442, 1455, 1499, 1502, 1504, 1518, 1527, 1544, 1568, 1582, 1593, 1618, 1667, 1670, 1671, 1683– 1685, 1704, 1774, 1785, 1801, 1873, 1876, 1878, 1885, 1896, 1949, 1982, 2015, 2023, 2124, 2144, 2149, 2161, 2177, 2201, 2239, 2244, 2296, 2326, 2356, 2394, 2397, 2411, 2418, 2471, 2503, 2519, 2523, 2690, 2732, 2761, 2780, 2790, 2826, 2835, 2836, 2856, 2857, 2888, 2891, 2919, 2934, 2959, 2980, 3011, 3017, 3073, 3096]. multi [3, 5, 10, 35, 72, 196, 222, 225, 228, 250, 278, 302, 343, 352, 353, 369, 456, 489, 651, 680, 738, 782, 811, 832, 833, 869, 903, 907, 962, 1020, 1030, 1032, 1061, 1138, 1175, 1206, 1225, 1240, 1294, 1340, 1347, 1352, 1394, 1426, 1432, 1484, 1539, 1584, 1591, 1609, 1620, 1640, 1664, 1677, 1706, 1711, 1730, 1795, 1796, 1813, 1833, 1872, 1888, 1940, 1946, 1962, 2035, 2037, 2061, 2090, 2127, 2159, 2187, 2206, 2262, 2265, 2306, 2366, 2386, 2387, 2391, 2396, 2441, 2443, 2444, 2492, 2549, 2550, 2569, 2577, 2614, 2621, 2727, 2808, 2865, 2916, 2930, 2945, 2949, 2963, 3015, 3064, 3078, 3082, 3105, 3129, 3134]. multi [57, 66, 108, 118, 211, 251, 257, 318, 441, 473, 554, 558, 621, 686, 702, 731, 733, 830, 959, 1134, 1143, 1150, 1248, 1269, 1292, 1295, 1341, 1414, 1576, 1578, 1599, 1727, 1746, 1809, 1851, 1954, 1960, 1975, 2020, 2021, 2135, 2144, 2171, 2245, 2247, 2314, 2373, 2385, 2476, 2485, 2491, 2493, 2508, 2563, 2598, 2625, 2631, 2632, 2674, 2686, 2714, 2823, 2840, 2913, 2954, 2992, 3007, 3041, 3049, 3152]. Multi-access [558, 733, 959, 1032, 1582, 1591, 1975, 2144, 2149, 2686, 2835, 2930, 3011].

Multi-agent [222, 318, 695, 778, 1061, 1292, 1414, 1435, 1442, 1670, 2020, 2037, 2159, 2187, 2201, 2247, 2262, 2356, 2387, 2614, 2632, 2826, 2836, 2919, 3015, 3049, 3096]. multi-agent-based [2727]. Multi-ant [1876]. multi-antenna [811]. multi-AP [1940]. multi-armed [302, 1640, 2015, 2144, 2245]. Multi-Attack [2836]. multi-attribute [2888]. Multi-authority [702, 1671, 1683, 2265, 2443, 2523]. multi-band [65, 2954]. multi-beam [1294, 1397]. multi-blockchain [1730]. multi-bottleneck [2598]. multi-branch [731, 2366]. multi-cache [2686]. Multi-cell [5, 211, 2161, 2840]. multi-centrality [800]. Multi-channel [456, 554, 2519, 2856]. multi-channeled [1427]. multi-class [57]. multi-classification [2061, 3064]. multi-cloud [869, 1138]. multi-cluster [2206]. multi-commodity [2177]. multi-connectivity [3041]. Multi-criteria [457]. Multi-dimensional [1878, 2090, 2549]. Multi-domain [250, 251, 833, 1126, 1341, 1593, 2023, 2159]. multi-domains [489]. multi-edge [869]. multi-environment [2306]. multi-factor [1544]. Multi-failure [1055]. multi-feature [2493]. multi-floor [2503]. Multi-flow [2485, 2891]. multi-frequency [1704]. multi-gNodeB [1896]. Multi-Graph [3073]. multi-group [1240, 2913]. Multi-hop [10, 65, 1020, 1030, 1034, 1143, 1578, 1664, 1667, 1684, 1946, 2015, 2124, 2732, 2808]. multi-information [782]. multi-input [651]. multi-institution [1962]. Multi-interface [1527, 1667]. multi-jammer [738]. Multi-keyword [1851, 2239]. multi-label [1150, 2411, 2674, 2980]. Multi-layer [781, 1394, 2391, 2418, 2823]. Multi-layered [282, 2945]. Multi-level [926, 1426, 1576, 1685, 1785, 1813].

Multi-link [1960, 2761]. Multi-MEC [1727]. multi-modal [2550, 2963]. multi-node [2127]. Multi-Objective [35, 75, 278, 352, 435, 473, 621, 680, 1046, 1245, 1269, 1796, 1888, 1949, 1954, 2135, 2326, 2471, 2934, 2959, 3017, 3152]. Multi-observable [374]. multi-operator [1225]. multi-optimality [627]. multi-pair [2037]. Multi-Party [2394, 2780]. Multi-path [644, 962, 1134, 1539]. multi-pipeline [2441]. multi-player [1455]. multi-point [343, 2621]. multi-privacy [2491]. multi-pronged [2625]. multi-protocol [1584, 3134]. Multi-provider [1502]. multi-purpose [1270]. multi-quality [343]. multi-radio [97]. Multi-RAT [413, 808, 1620]. Multi-rate [2690]. multi-region [228]. Multi-resource [1484, 1795, 1885]. multi-RIS [1295]. multi-router [196]. multi-scale [456, 1504, 2391, 2865]. multi-scanner [72]. Multi-Scheduling [2373]. Multi-Sensor [340]. multi-server [2386, 2396]. MUlti-Service [118, 907, 1428]. multi-session [1584]. Multi-site [2244]. multi-smartphone [903]. Multi-source [1499, 1982]. multi-split [1801]. Multi-stage [441, 1706, 1711, 1774, 2992]. multi-step [782]. multi-strategy [2476]. multi-swarm [3082]. Multi-Table [2171]. multi-task [832, 1248, 1599, 3078]. multi-tasks [1677]. Multi-tenancy [232]. multi-tenant [66, 353, 1258, 2577, 2857, 2949]. multi-tier [225, 369, 814, 1280, 2577, 3129]. Multi-timescale [800]. multi-to-multi [2916]. multi-topology [1347]. multi-traffic [3]. Multi-UAV [830, 1618, 2397, 2563, 2714, 3105]. multi-UAV-assisted [1873, 2314]. Multi-UAV-enabled [1340, 2790]. Multi-User [198, 257, 686, 1131, 1295, 1352, 1432, 1518, 1609, 1746, 1809, 1833, 1851, 2021, 2035, 2296, 2444]. multi-users [1206]. multi-variable [2385]. multi-vendor [489, 2508, 2569, 2631]. Multi-view [1175, 2411, 2492, 3007]. multibeam [2655]. multibiometric [328]. Multicast [198, 713, 1099, 1200, 1294, 1323, 1573, 1888, 1982, 2052, 2484, 3027]. multicasting [1048, 1143, 2690]. Multicell [1308, 2967]. multichannel [52, 2500, 2693]. multicolor [2004]. multidimensional [1296, 3045]. Multifactor [2910]. multiform [1452]. multigraph [299]. Multihop [1164, 1330, 1920]. multilayer [478, 1743, 3150]. Multimedia [158, 252, 332, 395, 703, 967, 1151, 1177, 1571, 2946]. multimetric [1151]. Multimodal [1369, 1533, 1597, 1633, 1722, 2137]. multiparameter [1396]. Multipath [23, 77, 165, 346, 447, 451, 455, 518, 545, 603, 673, 773, 803, 829, 1031, 1213, 1253, 1388, 1681, 1853, 1968, 2005, 2086, 2099, 2326, 2362]. Multiple [50, 86, 129, 274, 275, 331, 384, 388, 415, 455, 511, 552, 623, 744, 770, 793, 889, 936, 1039, 1113, 1238, 1390, 1430, 1511, 1512, 1817, 1881, 1955, 2035, 2070, 2148, 2183, 2249, 2409, 2557, 2602, 2745, 2813, 2828, 3083, 3087]. Multiple-level [793]. multiplexed [177, 346, 2099]. multiplexing [1056, 1324, 1393, 1703]. multiplier [2890]. Multiscale [2682, 2772, 3150]. multiservice [432, 674]. Multisite [2087]. multisource [626]. multistory [817]. multitask [870, 1533]. Multiuser [933, 1748, 2690]. multivariate [2050, 2360]. Munkres [689]. MUSCAT [2187]. MuSLi [1424]. MUTAA [1337]. Mutation [2556]. Mutual [728, 845, 1341, 1508, 1639, 2397, 3131]. MVC [3007]. MVC-Corr [3007]. Mx [2732]. Mx-TORU [2732].

n [2922]. Nakagami [1858]. Nakagami- [1858]. name [467, 527, 538, 595, 2519]. name-based [595]. Named [8, 155, 168, 426, 720, 829, 1165, 1771, 1775, 1804, 1897, 2628, 2836, 2870]. named-data [1771]. names [431]. nano [181]. nano-networks [181]. Nanonetwork [1880]. Nanonetwork-based [1880]. nanonetworks [292, 2430]. NAPT [2289]. NAR [3178]. NAR-based [3178]. narrow [450, 1333]. Narrowband [161, 2207]. Nash [383]. native [913, 1994, 2206, 2271, 2920]. Natural [788, 2291]. Navigation [268, 496, 2929]. NB [450, 817, 827, 1692, 1937, 2361]. NB-IoT [450, 817, 1692, 1937, 2361]. NBA [595]. NC [171, 1019]. NC-enabled [1019]. ND [2251]. NDN [8, 156, 171, 1812, 2484, 2870, 3026, 3160]. NDN-based [1812]. Near [211, 256, 350, 1116, 2048]. Near-optimal [211, 1116, 2048]. near-optimally [256]. NEAT [3123]. necessary [2251]. needles [341]. neighbor [1979, 2279]. NeighborGeo [2645]. Neighborhood [259, 625, 1226, 1780]. Neighborhood-Understood [1780]. neighbors [2645]. NEMa [2349]. NemesisGuard [984]. nerve [116]. nervous [1580]. NESEPRIN [1192]. NEST [1803, 2499]. nested [42]. Net [2243]. NetBoost [2515]. NetDiffus [2315]. Nets [2613]. NetShield [1879]. NetTiSA [1948]. NetView [307]. Network [19, 32, 171, 260, 291, 300, 310, 325, 330, 483, 493, 517, 528, 659, 668, 669, 671, 699, 711, 717, 748, 751, 764, 840, 860, 872, 881, 899, 931, 933, 968, 976, 993, 1025, 1082, 1086, 1147, 1193, 1226, 1228, 1233, 1278, 1289, 1382, 1384, 1391, 1402, 1502, 1515, 1665, 1688, 1740, 1762, 1791, 1824, 1913, 1920, 1988, 2003, 2011, 2033, 2056, 2059, 2082, 2092, 2103, 2150, 2163, 2170, 2184, 2243, 2244, 2315, 2329, 2427, 2442, 2445, 2479, 2482, 2488, 2527, 2529, 2572, 2574, 2592, 2771, 2772, 2787, 2803, 2821, 2886, 2949, 2959, 3002, 3013, 3106, 3114, 3116, 3130, 3143, 3157]. network [6, 30, 61, 68, 94, 127, 150, 203, 303, 490, 623, 634, 658, 667, 819, 828, 843, 915, 954, 1037, 1078, 1081, 1113, 1133, 1164, 1180, 1219, 1245, 1250, 1281, 1284, 1309, 1345, 1392, 1411, 1427, 1541, 1548, 1591, 1605, 1651, 1685, 1714, 1752, 1755, 1817, 1845, 1860, 1889, 1903, 1929, 1933, 1958, 1961, 1963, 2050, 2057, 2069, 2091, 2108, 2118, 2128, 2154, 2169, 2236, 2248, 2252, 2260, 2273, 2337, 2350, 2372, 2384, 2447, 2460, 2505, 2513, 2515, 2546, 2558, 2577, 2594, 2662, 2670, 2707, 2763, 2785, 2824, 2827, 2838, 2905, 2920, 3017, 3062, 3077, 3094, 3097, 3134, 3156, 3171]. network [14, 43, 69, 144, 148, 180, 196, 202, 228, 230, 238, 255, 290, 307, 311, 344, 354, 390, 409, 466, 511, 543, 564, 568, 597, 619, 663, 713, 739, 786, 905, 937, 938, 973, 994, 1062, 1069, 1097, 1103, 1107, 1170, 1178, 1257, 1260, 1311, 1322, 1347, 1351, 1359, 1365, 1375, 1395, 1470, 1510, 1559, 1572, 1581, 1596, 1619, 1624, 1663, 1707, 1735, 1741, 1769, 1792, 1803, 1827, 1837, 1888, 1915, 1934, 1948, 1952, 1980, 1982, 1983, 1992, 1999, 2004, 2014, 2076, 2262, 2272, 2302, 2353, 2382, 2581, 2612, 2626, 2682, 2715, 2720, 2723, 2783, 2883, 2925, 2945, 2958, 3015, 3021, 3034, 3042, 3087, 3107, 3137, 3155, 3174, 3188]. network [21, 63, 66, 76, 117, 139, 154, 186, 247, 271, 418, 445, 454, 473, 516, 529, 621, 632, 655, 657, 683, 772, 811, 830, 877, 888, 896, 902, 907, 934, 936, 1033, 1067, 1088, 1111, 1114, 1171, 1175, 1198, 1218, 1234, 1242, 1243, 1246, 1263, 1265, 1292, 1344, 1354, 1381, 1471, 1473, 1511, 1545, 1577, 1644, 1763, 1774, 1871, 1879, 1899, 1942, 2026, 2060, 2151, 2203, 2204, 2212, 2255, 2282, 2285, 2318, 2381, 2390, 2412, 2476, 2494, 2517, 2564, 2569, 2580, 2604, 2613, 2617, 2623, 2651, 2691, 2697, 2699, 2731, 2762, 2813, 2853, 2864, 2891, 2924, 2936, 2960, 2963, 2976, 3085, 3093, 3110, 3150, 3151]. network [74, 235, 338, 441, 691, 730, 742, 882, 982, 989, 1001, 1012, 1149, 1160, 1165, 1183, 1267, 1371, 1374, 1393, 1433, 1563, 1570, 1621, 1746, 1921, 1941, 1956, 1991, 2059, 2083, 2158, 2250, 2293, 2411, 2426, 2436, 2457, 2556, 2571, 2611, 2619, 2767, 2809, 2825, 2982, 2983, 2986, 2988, 2991, 3089, 3111, 3143, 3152, 3167]. network-accelerated [1344]. network-assisted [1651]. network-aware [739]. Network-based [139, 877, 915, 1871, 3062, 3106]. network-centric [2384]. Network-coded [888, 1384, 2092]. network-coding-based-multicast-wireless-sensor [713]. network-on-chip [937]. network-on-chips [2151]. network-slicing [517]. network-specific [3107]. network-wide [307, 2337]. networked [958, 1063, 1414, 2071, 2499, 2777]. Networking [8, 80, 155, 160, 168, 294, 426, 466, 476, 624, 625, 682, 720, 829, 944, 1093, 1363, 1374, 1426, 1450, 1610, 1647, 1657, 1754, 1771, 1775, 1804, 1897, 1960, 2019, 2043, 2049, 2210, 2284, 2359, 2393, 2395, 2664, 2707, 2761, 2765, 2836, 2870, 2916, 2925, 2940, 3004, 3012, 3123, 3169]. Networks [47, 75, 78, 136, 143, 205, 209, 265, 271, 282, 301, 315, 348, 357, 391, 435, 455, 478, 481, 487, 527, 539, 541, 579, 617, 682, 724, 725, 730, 734, 752, 787, 849, 880, 891, 901, 960, 1034, 1045, 1073, 1094, 1100, 1280, 1284, 1323, 1355, 1387, 1428, 1442, 1490, 1506, 1508, 1517, 1607, 1618, 1632, 1660, 1667, 1759, 1830, 1871, 1873, 1893, 1906, 1910, 1957, 2016, 2028, 2072, 2271, 2277, 2312, 2349, 2361, 2370, 2419, 2500, 2502, 2609, 2622, 2628, 2635, 2657, 2678, 2732, 2854, 2870, 2898, 2906, 2967, 3012, 3035, 3052, 3059, 3113, 3114, 3139, 3176]. networks [96, 97, 124, 134, 224, 237, 262, 289, 328, 373, 376, 383, 392, 407, 415, 416, 439, 460, 521, 544, 559, 563, 627, 701, 753, 773, 781, 1048, 1081, 1126, 1145, 1166, 1177, 1192, 1270, 1277, 1281, 1306, 1378, 1384, 1444, 1448, 1466, 1500, 1504, 1546, 1571, 1603, 1616, 1648, 1651, 1670, 1684, 1731, 1816, 1853, 1857, 1869, 1909, 1979, 2015, 2047, 2053, 2070, 2092, 2098, 2186, 2286, 2294, 2299, 2326, 2354, 2368, 2377, 2397, 2458, 2490, 2522, 2547, 2555, 2578, 2714, 2769, 2784, 2830, 2838, 2841, 2845, 2857, 2897, 2907, 2919, 2934, 2941, 2974, 2978, 3073, 3086, 3165]. networks [10, 67, 77, 120, 128, 181, 239, 242, 316, 331, 346, 359, 393, 459, 482, 513, 515, 551, 565, 678, 695, 810, 828, 868, 895, 919, 920, 927, 1030, 1046, 1101, 1184, 1215, 1273, 1293, 1308, 1331, 1332, 1424, 1435, 1462, 1497, 1522, 1523, 1526, 1540, 1541, 1561, 1569, 1573, 1574, 1590, 1612, 1627, 1672, 1678, 1752, 1872, 1886, 1891, 1945, 1971, 2054, 2055, 2057, 2062, 2097, 2099, 2100, 2124, 2147, 2177, 2181, 2213, 2249, 2259, 2336, 2351, 2364, 2439, 2446, 2481, 2539, 2576, 2597, 2647, 2675, 2755, 2786, 2808, 2842, 2844, 2860, 2902, 2923, 2971, 3003, 3044, 3062, 3080, 3082, 3100, 3105, 3180]. networks [9, 13, 33, 41, 83, 89, 104, 114, 219, 280, 327, 339, 352, 356, 369, 384, 508, 525, 572, 652, 692, 700, 713, 754, 759, 770, 798, 808, 897, 950, 1020, 1056, 1079, 1116, 1146, 1212, 1214, 1255, 1279, 1299, 1324, 1368, 1396, 1400, 1451, 1476, 1530, 1608, 1639, 1664, 1706, 1710, 1723, 1743, 1750, 1758, 1777, 1813, 1823, 1844, 1874, 1895, 1901, 1953, 1965, 1985, 2040, 2063, 2127, 2148, 2159, 2164, 2182, 2187, 2202, 2388, 2399, 2404, 2484, 2486, 2509, 2533, 2540, 2649, 2653, 2656, 2665, 2693, 2717, 2728, 2782, 2790, 2930, 2938, 2990, 2996, 2999, 3038, 3047, 3057, 3065, 3071, 3118, 3119, 3129, 3131, 3149, 3164]. networks [4, 22, 29, 95, 142, 164, 172, 180, 188, 193, 236, 286, 298, 378, 385, 389, 488, 503, 504, 566, 568, 595, 653, 671, 673, 690, 783, 785, 790, 791, 854, 893, 908, 970, 1014, 1039, 1121, 1129, 1130, 1151, 1155, 1187, 1217, 1222, 1254, 1298, 1320, 1330, 1338, 1395, 1463, 1469, 1511, 1582, 1600, 1653, 1736, 1739, 1781, 1789, 1794, 1815, 1831, 1875, 2032, 2039, 2088, 2167, 2197, 2280, 2306, 2332, 2389, 2405, 2444, 2449, 2474, 2544, 2659, 2700, 2715, 2739, 2757, 2783, 2795, 2799, 2863, 2895, 2931, 2939, 2953, 2955, 3019, 3025, 3028, 3101, 3104, 3122, 3184]. networks [15, 17, 28, 38, 66, 76, 102, 187, 216, 222, 251, 252, 272, 275, 305, 320, 358, 365, 379, 424, 462, 470, 474, 506, 514, 542, 545, 562, 575, 586, 616, 622, 645, 649, 722, 743, 760, 836, 871, 985, 1002, 1010, 1061, 1134, 1139, 1231, 1259, 1286, 1364, 1404, 1407, 1542, 1545, 1578, 1640, 1734, 1768, 1799, 1806, 1832, 1894, 1898, 1911, 1939, 1978, 1987, 1989, 1997, 2010, 2051, 2068, 2084, 2141, 2173, 2239, 2254, 2276, 2285, 2289, 2304, 2324, 2365, 2385, 2457, 2465, 2467, 2472, 2501, 2561, 2616, 2727, 2756, 2759, 2814, 2840, 2867, 2933, 3053, 3144, 3163, 3186]. networks [7, 37, 54, 59, 139, 141, 170, 177, 192, 199, 218, 267, 363, 366, 423, 440, 442, 477, 520, 533, 552, 554, 593, 609, 637, 647, 660, 661, 668, 679, 689, 729, 745, 766, 945, 969, 989, 1038, 1053, 1077, 1211, 1356, 1393, 1413, 1414, 1437, 1446, 1447, 1458, 1467, 1503, 1512, 1514, 1562, 1566, 1594, 1658, 1666, 1668, 1676, 1688, 1689, 1696, 1699, 1703, 1732, 1764, 1836, 1858, 1907, 1918, 1937, 2009, 2020, 2086, 2122, 2171, 2233, 2314, 2321, 2345, 2358, 2426, 2471, 2483, 2508, 2538, 2598, 2602, 2631, 2632, 2658, 2672, 2677, 2684, 2689, 2711, 2719, 2729, 2734, 2788, 2875, 2944, 2986, 3020, 3041, 3095, 3103, 3136, 3153]. networks [108, 427, 496, 1013, 1031, 1105, 1140, 1153, 1173, 1272, 1539, 1542, 1575, 1681, 2126, 2235, 2524, 2545, 2887, 2954, 3127, 3135]. Neural [117, 139, 235, 247, 290, 344, 390, 487, 564, 566, 639, 657, 658, 669, 718, 843, 881, 905, 1180, 1260, 1273, 1279, 1281, 1284, 1289, 1402, 1451, 1542, 1621, 1767, 1871, 1929, 1963, 2011, 2118, 2212, 2243, 2351, 2395, 2411, 2447, 2803, 2959, 3002, 3062, 3097, 3139, 3167, 3171]. Neuro [714]. Neuro-Fuzzy [714]. neutrality [982]. neutrosophic [1789]. Never [834]. Next [419, 505, 593, 639, 648, 787, 819, 1353, 1490, 1612, 1620, 1980, 2057, 2351, 2361, 2429, 2939, 3016]. Next-cell [2351]. Next-gen [2361]. next-generation [593, 787, 1353, 1612, 2057, 2429, 2939]. NextG [1597]. NFMI [350]. NFStream [968]. NFV [32, 59, 131, 190, 194, 250, 282, 320, 353, 378, 403, 472, 537, 658, 754, 854, 919, 1051, 1177, 1266, 1368, 1422, 1459, 1502, 1571, 1610, 2053, 2569, 2588, 2756]. NFV-based [2053, 2569]. NFV-enabled [59, 919, 989, 1368, 1610, 1837]. NFV/SDN [250, 378]. NGSO [3074]. NIC [1999]. NIDS [1824]. Nile [3123]. Nile-English [3123]. NLOS [1411, 1543]. NLOS-FSO [1411]. NLP [2800]. NLP-powered [2800]. NNSplit [1451]. NNSplit-SØREN [1451]. nnWeb [3054]. no [143, 321]. No-GPS [321]. no-scheduling [143]. Node [224, 259, 266, 349, 382, 800, 862, 960, 964, 1072, 1079, 1452, 1541, 1787, 1994, 2127, 2305, 2501, 2542, 2698, 2904, 3085, 3173]. node-independent [2904]. node-removal [382]. Nodes [10, 69, 381, 757, 1045, 1611, 2069, 2323, 2532, 2602, 2611, 2646]. noise [1847, 2207]. NOMA [219, 373, 521, 553, 559, 581, 733, 785, 817, 1032, 1073, 1131, 1211, 1237, 1243, 1295, 1384, 1450, 1473, 1480, 1607, 1609, 1686, 1736, 1874, 1907, 2004, 2092, 2141, 2149, 2213, 2421, 2444, 2472, 2490, 2529, 2605, 2629, 2690, 2830, 2867, 2902, 2913, 3165, 3181]. NOMA-assisted [2629]. NOMA-based [733, 785, 1874, 2213, 2529, 3181]. NOMA-enabled [559, 2149, 2902]. NOMA-FDR [1480]. NOMA-MEC-enabled [2141]. NOMA/OMA [3165]. Non [44, 132, 135, 347, 364, 388, 415, 491, 509, 560, 1035, 1362, 1367, 1512, 1534, 1658, 1670, 1736, 1937, 1985, 2164, 2249, 2336, 2359, 2405, 2495, 2547, 2632, 2656, 2657, 2664, 2725, 2745, 2894, 2923, 3084]. non-anchor [1937]. non-backtracking [491]. non-blocking [2664]. non-collided [135]. non-colluding [2249]. Non-cooperative [1035, 2495, 2632, 2725]. non-deterministic [1658]. non-ID [1736]. non-IEEE [132]. non-invasive [2894]. non-IoT [1367]. non-linear [44, 560, 1534]. non-Markovian [1362]. Non-Orthogonal [388, 415, 1512, 2745]. Non-parametric [509]. non-programmable [2923]. non-SDN [597]. non-standalone [1985]. Non-Terrestrial [1670, 2164, 2336, 2359, 2405, 2547, 2656, 2657]. non-urban [364]. non-wildcard [347]. non-zero-sum [3084]. Nonlinear [763, 1401, 1955, 1997, 3181]. nonnegative [891]. nonorthogonal [770, 1955]. nonuniform [256]. NORD [1541]. Normal [1453]. north [2034]. NoSQL [1022]. notice [1762]. Notification [778, 1708, 3035].

Novel [38, 74, 76, 280, 310, 407, 485, 549, 565, 589, 651, 662, 670, 672, 699, 701, 725, 792, 801, 979, 1069, 1101, 1133, 1159, 1165, 1255, 1296, 1303, 1323, 1397, 1404, 1551, 1561, 1584, 1603, 1640, 1784, 1849, 1857, 1870, 1898, 1960, 2000, 2120, 2177, 2269, 2306, 2349, 2425, 2431, 2550, 2662, 2729, 2752, 2773, 2873, 2924, 3043, 3118, 3124, 3148, 3172]. November [361, 848, 1305, 1808, 1843, 1893, 2414]. nowhere [2045]. NPN [1438]. NR [698, 1750, 1755, 1869, 2218, 2428, 2462, 2480, 2724]. NR-DC [1750]. NR-U [698]. NR-Unlicensed [2480]. NR-V2X [2218]. NSGA [2080, 2488]. NSGA-II-based [2488]. NSGA2 [144]. NSGA2-LR [144]. NTD [2289]. NTN [2418, 3074]. NTNs [2862]. NTRU [2575]. NTRU-CLS [2575]. NTS4PTP [1146]. Nucleo [2024]. NUI [1780]. number [1515, 2548]. numeric [2109]. numerology [698, 731, 1869]. NUV [238]. Nyquist [963].

O [525, 899, 1298, 1410, 2098, 2180, 2352, 2770, 2951]. O-JRC [2951]. O-RAN [525, 1298, 1410, 2180, 2352, 2770]. O-RAN-based [2098]. OAUTH [2089]. Ob [1246]. Ob-EID [1246]. obfuscated [2612, 3143]. obfuscation [1208, 2615, 2824]. object [309, 494, 648, 1135]. Objective [35, 75, 278, 352, 435, 473, 621, 680, 1046, 1212, 1245, 1269, 1796, 1888, 1949, 1954, 2135, 2326, 2471, 2752, 2934, 2959, 3017, 3152]. objectives [298]. objects [2728]. observable [374, 3105]. observation [1687, 2324]. observer [2064]. Obstacle [1246]. obstructing [2055]. occasions [2480]. occupancy [1173, 2097]. OCDMA [560]. OCDMA-PON [560]. octagonal [2607, 2797]. octagonal-cross-by-pass-mesh [2607]. Octagonal-Cross-By-Pass-Torus [2797]. October [296, 324, 777, 813, 1221, 1262, 1766, 2380, 3121, 3190]. ODL [1346]. ODM [1552]. ODM-BCSA [1552]. ODNS [1877]. OEM [2267]. OFDM [553, 963, 1847, 2543, 2754]. OFDM-based [1847, 2754]. OFDM-NOMA [553]. OFDMA [315, 376, 1332, 1518, 2604]. OFDMA-based [2604]. Off [23, 560, 600, 1003, 1422, 1739, 2576, 2647, 2817, 3019]. off-chain [2647]. off-loading [2576]. off-path [23, 2817]. Off-site [1422]. offline [1361, 2014]. offload [1498, 2107]. offloaded [1579]. Offloading [211, 223, 359, 378, 387, 389, 398, 451, 534, 559, 611, 696, 733, 742, 773, 781, 787, 815, 830, 839, 869, 956, 959, 981, 1040, 1064, 1085, 1088, 1123, 1167, 1179, 1187, 1190, 1193, 1237, 1394, 1406, 1432, 1466, 1470, 1475, 1485, 1536, 1552, 1663, 1677, 1710, 1731, 1742, 1758, 1760, 1814, 1906, 1910, 1942, 1975, 2040, 2076, 2108, 2113, 2128, 2135, 2141, 2149, 2152, 2155, 2200, 2235, 2262, 2270, 2283, 2288, 2296, 2356, 2448, 2461, 2466, 2471, 2472, 2518, 2577, 2588, 2622, 2635, 2667, 2684, 2685, 2726, 2732, 2799, 2817, 2840, 2878, 2930, 2938, 2950, 2959, 2971, 2992, 3005, 3011, 3031, 3049, 3054, 3081, 3105, 3154, 3182]. offloading [173, 257, 366, 558, 689, 922, 945, 1318, 1472, 1805, 2314, 2538, 2610, 2643, 2686, 2762, 2835, 2850, 3187]. offloading-efficiency [559]. offloading/sharing [223]. offs [278, 985]. Offspeeding [404]. olive [1977]. OMA [2421, 3165]. OMA/NOMA [2421]. omnidirectional [2283]. on-chip [1971]. on-chip-communication [2607, 2797]. On-demand [23, 307, 346, 1741, 2714]. on-field [1847]. on-NIC [1999]. ON/OFF [3019]. onboard [792]. one [630, 1810, 2452, 2683, 2895]. one-class [1810]. one-shot [2683]. one-vs-one [630]. Online [73, 90, 236, 256, 274, 327, 354, 411, 637, 668, 671, 699, 737, 843, 914, 967, 989, 1033, 1080, 1084, 1237, 1337, 1356, 1435, 1467, 1640, 1711, 1899, 1918, 1933, 1973, 1976, 1988, 2014, 2052, 2091, 2093, 2094, 2135, 2144, 2158, 2171, 2237, 2313, 2335, 2350, 2496, 2572, 2681, 2719, 2825, 2868, 2897, 3031, 3080, 3092]. online-learning [2496]. only [1662]. ONOS [1346]. OPC [2285]. opcode [2603]. Open [149, 301, 321, 391, 525, 534, 750, 820, 923, 983, 1002, 1054, 1142, 1148, 1163, 1209, 1367, 1372, 1456, 1621, 1749, 1802, 1811, 1829, 1915, 1938, 1972, 2019, 2036, 2054, 2056, 2117, 2250, 2329, 2392, 2402, 2484, 2493, 2653, 2739, 2951, 2969]. Open-access [820, 1749]. Open-RAN [1002]. open-set [1811, 2329]. open-source [149, 923, 1163, 1972, 2019, 2117, 2250]. OpenAirInterface [213, 2989]. OpenFlow [45, 1901, 2922]. OpenRAN [1410]. OpenSSL [2548]. operands [2603]. operated [53]. operating [574, 806, 1630]. Operation [763, 1162, 2267]. operation-aware [2267]. Operations [851, 918, 1433, 1880, 2024, 2056, 3169]. operator [173, 907, 1225, 1280, 2919]. operator-biased [907]. operator-owned [173]. Opportunistic [83, 338, 349, 447, 544, 900, 924, 1123, 1140, 1386, 1412, 1512, 1538, 2321, 2679]. Opportunities [193, 399, 577, 788, 1120, 1357, 1456, 1882, 2027, 2389, 2545, 2564, 3050, 3169]. opposite [1420]. opposition [2937]. opposition-based [2937]. oppositional [1164]. Opt [543]. Opt-ACM [543]. optic [963]. Optical [77, 120, 136, 177, 237, 346, 348, 478, 504, 545, 724, 752, 753, 760, 784, 871, 950, 1095, 1184, 1202, 1255, 1323, 1324, 1393, 1396, 1540, 1703, 1888, 1894, 2006, 2028, 2032, 2047, 2070, 2099, 2184, 2277, 2326, 2381, 2542, 2602, 2609, 2706, 2786, 2944, 2954, 2974]. optical-acoustic [120]. Optimal [103, 123, 150, 168, 211, 398, 404, 496, 554, 622, 678, 785, 814, 945, 953, 988, 1020, 1116, 1299, 1432, 1532, 1591, 1592, 1685, 1687, 1755, 1803, 1831, 1859, 1906, 2009, 2032, 2034, 2048, 2144, 2189, 2246, 2588, 2615, 2863, 2993, 3030, 3141]. optimal-latency [953]. optimality [627]. optimally [256]. optimisation [632, 1245, 1980]. Optimistic [2093]. Optimization [14, 83, 203, 239, 278, 289, 310, 344, 351, 359, 405, 435, 478, 550, 628, 638, 677, 739, 743, 764, 769, 810, 815, 819, 835, 844, 897, 902, 919, 959, 986, 1046, 1081, 1099, 1118, 1179, 1194, 1212, 1219, 1228, 1264, 1284, 1299, 1340, 1375, 1396, 1473, 1627, 1663, 1678, 1698, 1761, 1796, 1857, 1873, 1874, 1898, 1913, 1949, 1965, 2040, 2113, 2144, 2149, 2161, 2170, 2209, 2234, 2263, 2273, 2326, 2336, 2361, 2385, 2386, 2395, 2409, 2507, 2509, 2510, 2552, 2577, 2595, 2673, 2732, 2740, 2759, 2790, 2813, 2841, 2862, 2863, 2934, 2955, 2959, 3001, 3017, 3053, 3071, 3096, 3112, 3134, 3144, 3165, 3166]. optimization [377, 823, 867, 892, 1095, 1198, 1337, 1404, 1806, 1809, 1818, 1918, 2087, 2126, 2130, 2135, 2205, 2211, 2247, 2423, 2471, 2538, 2559, 2631, 2684, 2829, 2846, 2852, 3099, 3111, 3172, 3187]. Optimization-based [478, 764]. optimization-Machine [844]. optimize [2541]. Optimized [293, 294, 509, 543, 657, 917, 1003, 1062, 1069, 1166, 1231, 1686, 1746, 1770, 2184, 2529, 2587, 3042, 3182]. Optimizer [2611]. Optimizing [2, 119, 473, 1207, 1213, 1280, 1320, 1506, 2277, 2283, 2294, 2537, 2544, 2584, 2669, 2707, 2711, 2961, 2979, 3021, 3053, 3092]. Optimum [1401, 1558]. option [1385]. orbit [820, 1687]. Orbital [2324]. orchestrate [1423]. Orchestrating [313, 2897]. Orchestration [30, 250, 251, 303, 320, 510, 529, 988, 1051, 1111, 1266, 1484, 1980, 2038, 2146, 2555, 2611, 2719, 2939, 2982, 3014, 3042, 3112]. orchestrator [225]. Order [1978, 2608]. ordering [753, 2995]. OrderSketch [890]. organic [3176]. organizing [717, 828, 1714, 1722]. oriented [14, 160, 424, 549, 575, 588, 707, 946, 1353, 1818, 2412, 2853, 3187]. Orthogonal [388, 415, 1488, 1512, 2745]. ORTP [2352]. OSDL [1202]. OSPF [44]. OSWireless [1875]. OTFS [2428]. other [162]. OUSN [3072]. Outage [67, 413, 1618, 1736]. Outage-aware [413]. Outdoor [266, 796, 1079, 1124, 3001]. outlook [1705, 2807]. output [651, 2252]. outsourced [702]. outsourcing [264, 925, 1532]. over-the-air [263, 1329]. Overcoming [3167]. overhead [1471, 2033, 2287, 3110]. overlapped [842]. overlapping [153]. overlaps [218]. overlay [938, 2071]. overlaying [440]. Overview [276, 492, 577, 666, 1093, 1314, 1589, 1802, 2419, 2561]. OVS [534]. OVS-CAB [534]. owned [173]. owner [2688, 2828]. owner-level [2688].

P [2802]. P-HotStuff [2802]. P2P

[382, 1181, 1363]. P4 [780, 947, 1026, 1119, 1228, 1230, 1250, 1254, 1586, 1643, 2019, 2082, 2292, 2588, 2639, 2923, 2943, 3116]. P4-AQM [2082]. P4-Assisted [2943]. P4-based [1230, 1254]. P4-programmable [1119, 2923]. P4Resilience [1055]. P4Rex [2355]. P4TE [1217]. pack [37]. Packet [33, 104, 157, 347, 351, 370, 483, 533, 616, 669, 870, 921, 1020, 1113, 1114, 1196, 1382, 1397, 1521, 1662, 1750, 1753, 1778, 1892, 1917, 1941, 1963, 2110, 2202, 2299, 2318, 2350, 2424, 2425, 2430, 2608, 2808, 2857, 2903]. packet-based [616]. packet-chaining [1114]. packet-in-packet [104]. Packet-level [870, 1941, 2299]. packets [249, 567]. packing [1352]. PADAL [1992]. paging [2480]. pair [2037]. pairing [1887, 1914]. pairing-free [1914]. pairs [2409, 3141]. pairwise [1578]. Pandemic [217, 386, 973, 1369, 1496]. Papers [1195, 1647, 3012]. Paradigm [1478, 2162, 2340, 2753, 2962]. paradigms [662]. Parallel [535, 1447, 1870, 1911, 2133, 2802, 2854, 2890, 3087, 3103]. parallelism [889]. parallelized [2053]. Parameter [2376, 2553, 2554]. parameterization [3019]. Parameterized [1799, 1987]. Parameters [114, 1382, 1445, 2044, 2541, 2551]. parametric [509]. PARFAIT [1016]. parity [2654]. parking [990, 1783]. Parser [2171]. parsing [2857]. Part [2813]. PartChain [2450]. Partial [257, 841, 1081, 1132, 1190, 1983]. partially [1004, 1025]. participant [122, 418, 2101, 2868]. participant-selection [2101]. Participants [1409]. participatory [2369]. partition [93]. partitioned [2450]. partitioning [237, 347, 1406, 1621, 1911, 2127, 2210, 2326, 2974]. parts [135]. Party [36, 2394, 2780]. Pass [2607, 2797]. Passive [1617, 1630, 1992, 2070, 2320, 2517]. passive-active [2517]. Past [727]. Path [23, 110, 157, 162, 181, 186, 299, 445, 462, 638, 644, 962, 1004, 1034, 1134, 1539, 1784, 1837, 1838, 1876, 1887, 1954, 2032, 2124, 2408, 2459, 2583, 2656, 2662, 2817, 3106, 3179]. paths [86, 496, 1116, 1213]. Patient [1998, 2153, 2534]. Patient-centric [1998, 2153, 2534]. patrol [2964]. pattern [943, 1704, 1722, 2497]. patterns [502, 887, 1700, 2442]. PAutoBotCatcher [864]. PAveMENT [496]. Paving [2242]. PAWR [1410, 1749]. Payload [2716]. payment [127, 1539, 1900, 2060, 2647, 2780, 2973]. payments [2596]. PBAS [3065]. PBCNN [669]. PBFT [1104, 3100]. PBFT-based [1104]. PC [1578]. PCA [2886]. PCNP [2179]. PD [1211, 1668]. PD-CPS [1668]. PD-NOMA [1211]. PDMA [2888]. peak [2438]. peaks [1953]. PEDDA [1706]. peer [90, 738, 2060, 2176]. peer-to-peer [738, 2060, 2176]. Pegasus [3026]. Penalty [724]. Penalty-method [724]. PeNCache [2628]. penetrated [2453]. penetration [1981]. PEP [1928]. PEP-based [1928]. per-flow [1865]. per-sample [2768]. perceived [2687]. perception [1898]. perceptron [3150]. perceptual [2198]. perfect [521, 3131]. Performance [18, 66, 102, 139, 154, 210, 232, 267, 271, 316, 329, 333, 364, 376, 402, 408, 462, 469, 473, 504, 524, 537, 579, 606, 634, 666, 698, 723, 724, 732, 771, 773, 784, 792, 846, 897, 902, 921, 933, 989, 1077, 1130, 1143, 1147, 1286, 1295, 1306, 1333, 1384, 1407, 1408, 1473, 1500, 1513, 1532, 1537, 1605, 1609, 1619, 1665, 1719, 1729, 1740, 1822, 1830, 1866, 1964, 1985, 2024, 2035, 2133, 2205, 2277, 2284, 2382, 2386, 2392, 2444, 2472, 2485, 2521, 2562, 2600, 2623, 2739, 2785, 2809, 2870, 2878, 2903, 2922, 2931, 2940, 3090, 3109, 3136, 3162]. performance-aware [989]. performance-low [723]. performance-optimal [1532]. performances [1223]. period [1658]. Periodic [1720, 2917]. Periscope [2881]. permanent [1100]. Permissioned [2520, 3100]. permissionless [1363, 1611, 2505, 2573]. permutation [1192]. persistence [2218]. Persistent [341, 566, 1275, 1418, 1429, 1567, 1892, 2122]. person [1886]. Personalized [418, 1067, 1516, 1762, 2435, 2747, 2985, 3036]. personnel [1617]. Perspective [85, 215, 299, 372, 387, 596, 967, 1425, 1615, 1981, 2258, 2338, 2358, 2522, 2852, 2940, 3015, 3044]. Perspectives [620, 1124, 2384, 2412]. perturbation [309, 2768]. pervasive [1267, 1938, 2643, 2861]. PETIT [2449]. PETNet [1921]. Petri [2613]. PF [1914]. PF-IBDA [1914]. PFL [2435]. PFL-DKD [2435]. PGCE [1132]. PH [2922]. pharmaceutical [355]. phase [230, 287, 465, 701, 2069, 2829, 3079]. Pheromone [2751]. Pheromone-based [2751]. Phishing [247, 904, 1021, 1339, 2134, 2751, 2839, 2900, 3188]. PhishNet [3188]. PhishNot [1339]. phone [1862]. phones [2956]. PHP [2138]. Physical [412, 464, 503, 732, 913, 1030, 1041, 1316, 1319, 1547, 1935, 2027, 2072, 2208, 2453, 2490, 2677, 2707, 2725, 2871]. Physically [1641, 2650]. Physically-consistent [2650]. Pi [716, 2024]. Pi-Radio [716]. PIE [115]. PIFO [1950]. piggybacking [1836]. piggybacking-based [1836]. pile [640]. PiNcH [63]. ping [1575]. ping-pong [1575]. Pinning [571, 2252]. pipe [2084]. pipeline [113, 2441, 3149]. pipelined [2890]. PISA [1217]. PKI [1144, 1849]. Placement [7, 15, 103, 123, 168, 240, 331, 333, 363, 395, 473, 476, 525, 530, 624, 642, 645, 680, 692, 739, 797, 844, 854, 888, 895, 907, 966, 989, 1037, 1106, 1139, 1214, 1270, 1359, 1375, 1458, 1507, 1713, 1732, 1817, 1833, 1846, 1936, 2091, 2126, 2172, 2293, 2335, 2386, 2406, 2465, 2501, 2571, 2592, 2646, 2686, 2861, 2932, 2961, 2992]. places [1054]. placing [668]. Plaintext [1921]. Plaintext-aware [1921]. plan [576]. Plane [261, 423, 469, 504, 694, 944, 1055, 1196, 1230, 1350, 1451, 1592, 1643, 1701, 1846, 1868, 2146, 2233, 2357, 2595, 3075, 3178]. planes [241, 1572, 1656, 1941, 2923]. planet [827]. Planning [110, 154, 186, 271, 293, 638, 678, 714, 778, 811, 917, 1431, 1433, 1711, 1758, 1784, 2014, 2201, 2408, 2459, 2489, 2563, 2790, 3092, 3105, 3106]. Planning-based [1711]. plants [2282]. Platform [119, 288, 334, 414, 526, 528, 667, 750, 781, 846, 899, 990, 1090, 1141, 1637, 2284, 2635, 2657, 2738, 2895, 2951, 3026]. platform-centric [119]. platforms [107, 228, 409, 477, 494, 606, 620, 948, 983, 1314, 1410, 1532, 1974, 2112, 2185, 2889]. platoon [458, 2561]. platoon-based [2561]. play [1002, 2090]. player [1455]. playout [3029]. PLC [463, 1847]. PLC/RF [463]. plenty [513]. plot [1800]. PLS [1222]. PLS-HECC-based [1222]. Plug [2046, 2090]. Plug-and-play [2090]. pluggable [2362]. plus [2745]. PMIPv6 [284]. Po-Fi [506]. Poa [1956]. PoiEvent [1275]. Point [34, 108, 302, 343, 810, 1352, 2439, 2472, 2621, 2792, 2859, 2972, 3021]. point-of-interests [2859]. point-to-point [1352]. pointer [1470]. points [173, 499, 744, 1270]. POIs [805]. poisoning [2125, 2454, 2579, 2625, 2749, 2763, 2998, 3009, 3058, 3075]. Poisson [300]. polar [2363]. PoLe [905]. policies [1928, 2635]. policing [2527]. Policy [4, 124, 446, 656, 683, 756, 1166, 1374, 1419, 1855, 1874, 1984, 2265, 2495, 3010, 3014]. Policy-based [1166, 2495]. Policy-rich [756]. political [372]. Pollard [1231]. polling [1260]. pollution [1804, 2261]. Polynomial [1240, 2219]. Polynomial-time [2219]. PON [226, 560, 678, 2106]. pong [1575]. pool [991]. pool-hopping [991]. pooling [1656]. poor [281]. popular [426]. Popularity [91, 294, 2628, 2898]. Popularity-aware [91]. populations [18]. PoR [2133]. pornography [555]. port [312, 1668]. port-based [312]. portable [3124, 3180]. portion [347]. portion-based [347]. Position [757, 767, 1157, 1632, 2186, 2205, 2510, 2995]. Position-anchored [2995]. position-based [1157, 1632]. Positioning [1072, 1135, 1691, 1848, 2290, 2325, 2422, 2503]. Post [432, 1283, 1526, 2304, 2390, 2617, 3024, 3040]. post-disaster [432, 1526]. post-exploitation [2390]. Post-fault [2304]. post-quantum [1283, 3024, 3040]. posterior [2642]. postquantum [2521]. Potential [34, 82, 1777, 1924, 1972]. PoVF [2733]. PoW [1172, 1398]. Powder [750]. Power [103, 195, 265, 272, 275, 286, 315, 322, 342, 356, 383, 440, 553, 663, 705, 713, 739, 850, 888, 963, 985, 989, 1095, 1162, 1211, 1229, 1292, 1308, 1321, 1381, 1404, 1450, 1522, 1546, 1686, 1739, 1826, 1858, 1862, 1881, 2081, 2108, 2117, 2282, 2456, 2459, 2522, 2556, 2572, 2602, 2813, 2830, 2933, 2938, 3059, 3071, 3166]. power-efficiency [272]. power-efficient [739, 989]. power-Splitting [275]. powered [12, 202, 445, 554, 705, 749, 1178, 1190, 1485, 1526, 1809, 2058, 2149, 2213, 2331, 2507, 2510, 2604, 2673, 2800, 2950, 2993, 3146]. powerful [2809]. pp [1571]. PPTA [1491]. PQBFL [3024]. PQCLP [1987]. Pr [655]. Pr-CAI [655]. PRACH [908]. Practicable [1098]. Practical [10, 108, 561, 600, 719, 1050, 1668, 1706, 1720, 1724, 2165, 2400, 2624, 2904, 3026]. Practicality [1963]. practice [873]. practices [1829]. pragmatic [1356]. Pre [1543, 1708, 2609, 2618, 3078]. pre-congestion [1708]. pre-mitigation [1543]. Pre-provisioning [2609]. pre-trained [2618]. pre-training [3078]. preamble [1215]. Precise [466, 887, 2179, 3113]. precision [10, 56, 130, 2986, 3142]. Predicted [338]. predicting [419, 502, 1223, 2734]. Prediction [13, 139, 233, 330, 333, 345, 380, 390, 509, 601, 650, 870, 889, 933, 936, 987, 1001, 1008, 1102, 1291, 1420, 1428, 1456, 1457, 1525, 1534, 1572, 1581, 1699, 1779, 1785, 1792, 1840, 1841, 1903, 1953, 1958, 1961, 1987, 2059, 2065, 2204, 2231, 2274, 2351, 2365, 2372, 2381, 2432, 2479, 2553, 2581, 2656, 2682, 2713, 2728, 2791, 2855, 2864, 3002, 3107, 3130, 3139, 3150, 3153, 3184]. Prediction-based [987]. predictions [970, 1049]. Predictive [887, 1094, 1561, 1700, 1781, 1885, 2264, 2287, 2368, 2439, 3104]. predictor [658]. preemption [2049, 2996]. preference [1516, 1945, 3158]. preference-aware [1945]. preferences [2232]. prefetch [156]. prefix [925]. preliminary [666]. premise [873]. PrePass [261, 446]. PrePass-Flow [446]. preprocessing [1349]. prescription [1104]. presence [1657, 1736, 2706]. Present [727]. preservation [96, 477, 1115, 1912, 2563, 2807]. preserve [594]. preserved [1013]. Preserving [212, 406, 512, 599, 864, 932, 955, 962, 973, 1016, 1087, 1224, 1252, 1277, 1439, 1443, 1491, 1517, 1551, 1553, 1744, 1839, 1974, 1998, 2030, 2073, 2125, 2191, 2223, 2224, 2298, 2313, 2339, 2341, 2374, 2468, 2560, 2596, 2705, 2710, 2721, 2726, 2763, 2776, 2780, 2859, 2866, 2888, 3090, 3126, 3135, 3175]. prevent [1646]. prevention [282, 437, 438, 751, 1188, 2783, 2960]. Price [137]. Price-aware [137]. Pricing [406, 785, 869, 945, 1274, 1459, 2147, 2200, 2735, 2782, 2805, 3127]. Pricing-based [945, 1274, 2805]. PriDPM [406]. primal [1973]. primal-dual [1973]. primitive [287, 1788]. principle [954]. Principles [1235, 1298, 1373]. prior [2373]. prioritization [1149, 1712]. prioritized [39]. prioritizing [889]. Priority [267, 655, 1047, 1077, 1324, 1952, 2679, 2937, 3065]. Priority-aware [1324, 2937]. Priority-awareness [1047]. Priority-based [267, 3065]. priority-queue [1077]. Priv [2374]. Priv-Share [2374]. Privacy [96, 212, 250, 276, 406, 467, 477, 490, 512, 527, 590, 594, 599, 606, 620, 864, 914, 916, 932, 955, 962, 973, 980, 1013, 1016, 1022, 1087, 1115, 1159, 1208, 1216, 1224, 1252, 1356, 1389, 1425, 1443, 1491, 1509, 1516, 1517, 1551, 1553, 1730, 1738, 1744, 1814, 1821, 1839, 1891, 1912, 1974, 1990, 1998, 2030, 2065, 2073, 2125, 2160, 2162, 2191, 2223, 2224, 2239, 2259, 2266, 2268, 2298, 2313, 2339, 2341, 2374, 2468, 2491, 2516, 2560, 2596, 2702, 2705, 2710, 2721, 2722, 2726, 2763, 2776, 2780, 2807, 2819, 2859, 2866, 2887, 2888, 2935, 3044, 3090, 3126, 3135, 3175]. Privacy-aware [250, 1814, 1891, 2065]. privacy-preservation [1115]. privacy-preserved [1013]. Privacy-preserving [406, 512, 864, 955, 962, 1016, 1087, 1224, 1252, 1443, 1491, 1517, 1551, 1553, 1744, 1839, 1974, 2030, 2073, 2125, 2191, 2223, 2298, 2313, 2339, 2341, 2374, 2560, 2596, 2705, 2710, 2721, 2726, 2763, 2776, 2780, 2859, 2866, 2888, 3090, 3126, 3135, 3175]. Privacy-utility [1208]. Private [727, 914, 1068, 1152, 1312, 1383, 2031, 2123, 2266, 2285, 2327, 2766, 2768, 3063, 3094, 3141, 3151, 3186]. PRLAP [1919]. PRLAP-IoD [1919]. Proactive [29, 95, 353, 685, 752, 851, 942, 1253, 1433, 1814, 2152, 2514, 2611, 2700]. proactively [1701]. Probabilistic [370, 791, 1553]. Probability [181, 299, 338, 1073, 1173, 1184, 1345, 1526, 1618, 1669, 2169, 3085]. Probability-based [181, 299]. probability-informed [2169]. probe [791, 998, 2611]. Probe-Optimizer [2611]. ProbInfer [299]. probing [1011, 2279]. problem [224, 348, 352, 621, 797, 844, 1434, 2177, 2591, 2773]. Problems [377, 557, 1142, 1357]. procedure [524, 734]. procedures [410, 1093]. Process [102, 207, 487, 681, 1003, 1362, 2025, 2853]. Process-oriented [2853]. processes [392, 2518, 2792, 2796]. Processing [27, 46, 351, 517, 619, 637, 703, 788, 1113, 1141, 1392, 1485, 1559, 1606, 1963, 2058, 2113, 2133, 2291, 2423]. processing/offloading [1485]. Processor [2716]. procurement [2327]. producting [1917]. production [1316, 2734]. Profile [183, 1133, 2029]. profiling [1665, 1862]. Profit [601, 1818]. profit-oriented [1818]. profitable [699]. Programmability [623, 694, 1230, 2146, 3116]. Programmable [241, 391, 429, 622, 1026, 1055, 1119, 1196, 1451, 1498, 1521, 1572, 1643, 1656, 1941, 1942, 2110, 2171, 2299, 2305, 2355, 2357, 2401, 2857, 2920, 2923, 3075]. Programming [280, 338, 429, 560, 944, 2465]. programs [906, 2595]. Progress [1120, 1483]. progressive [1706]. prompt [1827, 2504]. prompt-based [1827]. Prompt-WNQA [1827]. pronged [2625]. Proof [822, 905, 973, 1172, 1288, 2018, 2250, 2259, 3055]. proof-of-concept [973]. Proof-of-Stake [822, 2018]. proof-of-work [1172, 3055]. ProofChain [1144]. Propagation [10, 89, 109, 1012, 1038, 1356, 1362, 1402, 1562, 1783, 1861, 1884, 2122, 2208, 2218, 2293, 2533, 2802, 3171, 3179]. properties [1115, 2882]. property [1252, 2514]. proportional [1049, 1554, 2356]. proportional-fair [1049]. Proposal [2324]. proposed [338]. Prospects [958, 1864]. protect [834, 2162]. protected [237, 1096]. Protecting [2451, 2483]. protection [189, 203, 319, 337, 484, 694, 1098, 1159, 1216, 1312, 1422, 1444, 1669, 2031, 2145, 2211, 2239, 2268, 2326, 2514, 2609, 2722]. PROTECTOR [337]. Protocol [181, 189, 207, 245, 258, 273, 286, 292, 326, 400, 452, 471, 511, 519, 539, 590, 599, 700, 708, 726, 728, 743, 746, 759, 786, 794, 838, 845, 849, 856, 883, 898, 953, 975, 1018, 1019, 1083, 1129, 1151, 1154, 1164, 1173, 1185, 1224, 1288, 1317, 1341, 1348, 1362, 1387, 1397, 1439, 1462, 1463, 1489, 1517, 1546, 1548, 1551, 1569, 1584, 1590, 1600, 1625, 1641, 1644, 1660, 1681, 1781, 1872, 1914, 1919, 2016, 2018, 2030, 2068, 2096, 2132, 2143, 2165, 2189, 2195, 2253, 2259, 2319, 2339, 2349, 2352, 2354, 2397, 2573, 2666, 2680, 2693, 2732, 2759, 2776, 2980, 3024, 3038, 3040, 3057, 3134]. protocol [597, 660, 799, 1029, 1334, 1461, 1601, 1664, 1956, 2111, 2321, 2401, 2557, 2619, 2722, 2766]. protocols [58, 64, 162, 178, 275, 355, 852, 901, 1042, 1084, 1538, 1676, 1822, 1832, 1930, 2123, 2222, 2307, 2348, 2410, 2882, 3178]. prototype [2286]. prototypes [2496, 2739]. provable [452, 2626, 2776]. Provably [708, 1334, 1914, 2143, 2246, 2508]. prove [2214]. Provenance [157, 3103]. Provenance-enabled [157]. provide [914, 1335, 1366]. provider [593, 1364, 1502, 2442]. providers [2089]. Providing [2582]. provision [356, 969]. Provisioning [27, 138, 545, 752, 950, 960, 977, 1118, 1202, 1285, 1331, 1459, 1510, 1894, 1936, 1988, 2038, 2047, 2475, 2609, 2691, 2954, 3142]. proximity [2123]. Proxy [1054, 1306, 2069, 2119, 2523, 2931, 3177]. PRSD [2339]. Pseudonym [2620]. PSI [2339]. PSM [1162]. PSM-DMO [1162]. pSMART [250]. PSO [61, 1473, 2479]. PSO-LightGBM-TM [2479]. PTAP [1551]. PTP [1146]. Pub [1335, 1409, 1492, 1751, 2267]. Pub/sub [1335, 1409, 1492, 1751, 2267]. Public [191, 708, 833, 946, 1252, 1383, 1671, 1774, 3051, 3063]. public/private [3063]. publish [840, 939, 1360, 1651, 1679]. publish-subscribe [840, 939]. publish/subscribe [1679]. publishing [96, 1763]. PUF [412, 672, 845, 1341, 1718, 1919, 2165, 2397, 2449, 2714]. PUF-based [845, 1341, 1919, 2165]. PUF-enabled [2449]. PUF-secured [2714]. pull [304]. pull-based [304]. pulsing [148]. purification [158]. purpose [1270]. purposes [1791]. push [1775]. Pushing [1617]. pyramid [2391, 2823]. Pythia [1113].

Q [394, 730, 755, 759, 799, 867, 1010, 1463, 1507, 1511, 1804, 3152, 3160]. Q-FANET [799]. Q-ICAN [1804, 3160]. Q-learning [755, 759, 799, 1010, 1507, 1511]. Q-learning-based [1463]. Q-Network [730, 3152]. Q-SR [867]. QBDD [2132]. QC [2854]. QC-LDPC [2854]. QMIX [3081]. QoE [27, 137, 154, 271, 553, 1177, 1524, 1571, 1973, 2103]. QoE-aware [553, 1973]. QoE-driven [1177, 1571]. QoS [14, 16, 111, 134, 186, 470, 476, 515, 550, 617, 625, 663, 701, 851, 967, 1056, 1291, 1330, 1353, 1388, 1620, 1667, 1779, 1857, 1873, 1907, 1909, 1924, 2012, 2040, 2335, 2417, 2499, 2687, 2749]. QoS-Aware [16, 186, 851, 1056, 1388, 2040, 2335, 2417, 2749]. QoS-based [1907]. QoS-guaranteed [470, 476]. QoS-Oriented [14, 1353]. QoT [1112, 1184]. QRNG [2030]. QTAR [1463]. Quadtree [471]. Quadtree-based [471]. Quality [169, 202, 206, 225, 343, 444, 455, 497, 504, 660, 924, 1066, 1332, 1371, 1382, 1659, 1782, 1791, 1821, 1886, 2050, 2198, 2313, 2798, 2806, 2868, 2933]. quality-based [2313]. Quality-driven [1332, 2798]. Quantification [2174, 2984, 3004]. Quantifying [1978]. quantile [1112, 2395]. Quantitative [2546, 2975]. quantization [2851]. Quantum [1154, 1283, 1486, 1624, 1890, 1987, 2132, 2359, 2410, 2415, 2419, 2422, 2458, 2465, 2548, 2575, 2617, 2775, 2810, 2821, 2925, 2934, 2941, 3001, 3024, 3040]. Quantum-resistant [2132, 2575, 2810]. Quantum-Safe [2941]. quantum-secured [2925]. quasi [314]. quasi-random [314]. queries [804, 942, 1277]. Query [183, 834, 1176, 1900, 2371, 2560, 2876]. question [1827]. queue [115, 551, 866, 1077, 1521, 2097, 2272, 2664, 2793, 3010]. queue-aware [551]. queueing [1950, 2922]. queues [1785, 2373, 2975]. Queuing [340, 2079, 2299, 2908]. QUIC [874, 1213, 1358, 1523]. Quick [513].

R [57, 2485, 3112]. R-CNN [57]. R-IBN [3112]. Radar [1637, 2951]. Radar-Communication [2951]. radiated [2207]. Radiation [2275]. Radio [3, 43, 97, 109, 154, 176, 216, 236, 249, 271, 310, 318, 322, 331, 334, 339, 365, 383, 392, 394, 435, 439, 553, 554, 652, 701, 716, 736, 851, 933, 969, 1031, 1048, 1082, 1130, 1173, 1178, 1267, 1297, 1338, 1372, 1382, 1511, 1545, 1637, 1672, 1680, 1728, 1755, 1783, 1791, 1916, 2036, 2051, 2054, 2213, 2250, 2561, 2562, 2592, 2712, 2845, 3019, 3101, 3133]. radio- [851]. radio-based [1338]. radios [1069, 1749, 1972, 2290, 2487]. Raft [1956]. Raft-Poa [1956]. RAGN [2803]. Raiju [2390]. railway [578]. rainfall [1382]. RAN [19, 114, 231, 314, 318, 525, 808, 859, 1002, 1210, 1215, 1219, 1264, 1298, 1306, 1410, 1709, 1728, 1801, 2056, 2098, 2180, 2184, 2352, 2392, 2402, 2739, 2770, 2969]. RAN-aware [1306]. Random [89, 205, 314, 316, 809, 1010, 1198, 1333, 1347, 1625, 2548, 2706, 2791]. random-based [89]. randomization [594, 998, 1328]. Randomized [1768]. randomly [1293]. randomness [2458]. Range [738, 1079, 1277, 1546, 1900, 2225, 2229, 2737]. range-based [1079]. Range-free [2737]. ranging [2167]. Rank [1326, 1521, 2752]. rank-aware [1521]. ranked [2239]. ranking [910, 1541]. Ranking-based [1541]. RANs [138, 2074]. RansomCare [608]. Ransomware [608, 1932]. Rapid [1306, 1442, 2352, 3035]. raptor [1591, 1625]. Raptor-IRSA [1625]. RAQ [2655]. RAQ-GERT [2655]. Raspberry [2024]. RAT [413, 808, 1620]. rate [112, 192, 342, 369, 513, 573, 763, 1032, 1071, 1085, 1190, 1308, 1522, 1690, 1773, 1960, 1996, 2010, 2035, 2092, 2170, 2249, 2326, 2334, 2478, 2486, 2599, 2604, 2690, 2806, 2936, 3181]. rate-splitting [2035]. rateless [2971]. Rateless802.11 [281]. rates [699, 3029]. ratio [862]. rational [2031]. raw [1279, 2003]. Ray [2430]. Rayleigh [142]. RB [1869]. RBT [2003]. RCAP [786]. RCAP/CTAP [786]. RCO [2403]. RD [2332]. RD-ASVTuw [2332]. RDMA [2256, 2665]. RDMA-CM [2256]. RDP [2864]. Re [62, 291, 303, 752, 1242, 2119, 2523]. Re-defining [62]. Re-Encryption [2119, 2523]. re-mapping [1242]. re-orchestration [303]. re-provisioning [752]. re-synchronization [291]. reachability [238]. Reaching [507]. reaction [739]. reactive [1814]. readable [1365]. readers [811]. reading [821]. ready [1747]. Real [379, 619, 675, 726, 850, 957, 1141, 1194, 1303, 1349, 1383, 1514, 1547, 1575, 1595, 1699, 1737, 1745, 1993, 2173, 2217, 2231, 2247, 2264, 2274, 2320, 2356, 2541, 2644, 2855, 2856, 2958, 2969, 3010, 3042, 3119, 3160, 3174, 3186]. Real-time [379, 619, 726, 850, 957, 1141, 1194, 1303, 1349, 1514, 1547, 1595, 1699, 1745, 2173, 2217, 2231, 2247, 2320, 2356, 2644, 2855, 2856, 2958, 2969, 3010, 3042, 3119, 3160, 3174, 3186]. Realistic [148, 216, 887, 1801, 1822, 2145, 2905]. Reality [697, 865, 1187, 1779, 2303, 2740]. Realizing [1622]. Reallocation [451]. Realtime [390, 970, 2181]. reasoning [1644]. reasoning-based [1644]. rebalancing [86]. receive [2664]. Receiver [142, 1310, 1695, 1812, 2111, 2332, 2659, 2672, 2731, 3101]. Receiver-based [1812]. Receiver-driven [1695, 2111, 2332, 2731]. receiver-initiated [3101]. receivers [176]. recently [1892]. reception [2430, 3045]. receptive [1186]. rechargeable [897, 1212, 2727]. Recognition [290, 996, 2090, 2109, 2131, 2468, 2988, 3064, 3133]. Recognizing [2973]. Recommendation [582, 863, 1633, 2417, 2662, 2852, 2859, 3145]. recommendation-based [2662]. recommender [2525, 2749]. Recommending [2747]. reconfigurable [670, 954, 1153, 1986, 2019, 2196, 2444]. Reconfiguration [659, 988, 1207, 1214, 1497, 1515, 1816, 1846, 1983, 2795, 2825, 2980, 3142]. reconnaissance [2342]. reconnecting [2662]. reconstruction [197, 782, 1903, 2497]. record [875]. Records [419]. Recovery [7, 622, 1055, 1254, 1669, 1868, 2292, 2366, 2426, 2648, 2659, 2665, 2680]. recruitment [122, 1290, 3090]. rectangles [747]. rectangles-based [747]. recurrence [1800]. Recurrent [639, 1061, 1542, 1557, 2772]. Redact4Trace [2114]. Redactable [2114, 2215, 2398, 2520]. RedCap [2462]. redistribution [2681]. reduce [725, 1420]. reduced [127, 755]. reducer [654]. Reducing [693, 1184, 1397, 2033, 2060, 2598]. reduction [381, 468, 497, 1436, 3019]. reduction-of-quality [497]. redundancy [188, 468, 1268]. redundant [1605]. reference [234, 1079]. reflecting [2035, 3168]. reflective [2048]. regenerative [2025]. Region [228, 1989, 2323, 2644, 2907]. Region-aware [2323]. Region-based [2907]. Region-of-Interest [2644]. RegionFilter [2323]. regions [6, 1996]. Registered [2003]. Registered-backoff-time [2003]. registration [189]. regression [419, 1238]. regression-based [1238]. regret [2417]. regular [51, 2355]. Regularized [152]. regulation [982]. Regulatory [1791]. reinforced [1581]. Reinforcement [222, 272, 274, 385, 403, 551, 571, 627, 695, 697, 749, 765, 839, 866, 927, 1032, 1037, 1047, 1061, 1088, 1194, 1206, 1215, 1226, 1309, 1320, 1321, 1353, 1375, 1394, 1442, 1470, 1475, 1494, 1502, 1515, 1573, 1612, 1623, 1629, 1670, 1677, 1707, 1731, 1739, 1752, 1794, 1795, 1797, 1829, 1837, 1870, 1872, 1891, 1898, 1976, 1995, 2037, 2052, 2071, 2082, 2084, 2087, 2130, 2150, 2159, 2170, 2181, 2201, 2235, 2262, 2270, 2288, 2337, 2361, 2385, 2387, 2390, 2459, 2506, 2561, 2597, 2605, 2606, 2614, 2640, 2700, 2709, 2730, 2746, 2777, 2799, 2826, 2919, 2932, 2938, 2942, 2950, 2957, 2961, 3011, 3084, 3112, 3115, 3160, 3182]. reinforcement [252, 587, 605, 648, 729, 830, 959, 1086, 1265, 1292, 1414, 1566, 1661, 1712, 1732, 1799, 1805, 1832, 1856, 2020, 2151, 2537, 2559, 2622, 2632, 2761, 2850, 2873, 2964, 2981, 3020, 3039, 3049, 3085]. Reinterpreting [2227]. reject [1385]. rekey [1724]. relabelling [1596]. related [492, 1068, 2024, 2956]. Relational [1644]. relationship [1234, 1273]. Relay [12, 142, 266, 454, 837, 1073, 1077, 1145, 1272, 1295, 1384, 1548, 1608, 1686, 1789, 1887, 1977, 1994, 2015, 2051, 2429, 2444, 2637, 2867, 2967]. Relay-Assisted [1073, 2967]. Relay-empowered [2051]. relaying [195, 1512, 2409, 2602, 2677, 2875]. relays [275, 1412, 1419, 1506, 2092]. release [2516]. Reliability [283, 356, 745, 1591, 1676, 1920, 1982, 2875, 3183]. Reliability-aware [3183]. reliability-guaranteed [1591]. Reliable [381, 468, 625, 769, 1072, 1134, 1157, 1219, 1247, 1341, 1468, 1555, 1632, 1750, 1826, 1872, 1896, 2032, 2278, 2321, 2325, 2348, 2363, 2367, 2429, 2626, 2783, 3004, 3101]. relocation [1540]. Remainder [2827]. remote [53, 55, 1027, 1130, 2638, 2722, 2762]. remote-controlled [1130]. removal [382, 2532]. renewable [2282]. renewal [1698]. Rényi [2083]. repair [1776, 3087]. Repeated [654]. repeater [2465]. repetition [818, 1748]. replacement [1642, 2268, 2387, 2569]. Replay [1493, 2034, 2633]. Replicated [429]. replication [60, 1132, 1437, 1764]. replications [931]. representation [1599, 1782, 2519, 2863, 2886, 2893, 3076]. representations [430, 2781]. Reputation [285, 374, 533, 1336, 1956, 1970, 2085, 2869, 2995]. Reputation-based [285, 1956]. reputation-driven [2869]. request [1139, 1713]. requesters [1390]. requests [300, 436, 998, 2093, 2158]. required [545]. Requirements [84, 1490, 2201]. Rerouting [1643, 2794]. Res [1793]. Res-TranBiLSTM [1793]. rescue [1526, 1880]. Research [58, 298, 334, 528, 540, 577, 750, 935, 1124, 1241, 1425, 1483, 1802, 1860, 1977, 2112, 2389, 2503, 2775, 2919, 3070, 3095]. reservation [305, 1114]. resident [2671]. Resilience [381, 400, 503, 706, 1055, 1458, 2174, 2403, 2613, 2697]. Resilient [285, 670, 673, 728, 789, 2578, 2597, 2730, 2825, 3051, 3138]. resistance [1877, 2721, 2763]. resistant [277, 475, 2132, 2575, 2810, 2847, 2906, 3040, 3051]. resisting [672, 3023]. resolution [1327, 1374, 1863, 2094, 2198, 2360, 3186]. resolvers [1148]. Resource [10, 39, 43, 128, 138, 219, 233, 265, 274, 315, 373, 383, 416, 460, 472, 484, 525, 530, 549, 559, 573, 574, 579, 584, 628, 658, 728– 730, 823, 824, 950, 979, 1034, 1143, 1160, 1161, 1206, 1258, 1268, 1274, 1315, 1318, 1340, 1435, 1450, 1475, 1484, 1518, 1537, 1547, 1573, 1582, 1627, 1637, 1670, 1714, 1755, 1770, 1795, 1833, 1885, 1895, 1906, 2016, 2065, 2094, 2149, 2155, 2294, 2296, 2323, 2336, 2361, 2386, 2439, 2448, 2510, 2522, 2529, 2576, 2577, 2595, 2673, 2732, 2763, 2830, 2878, 2907, 2919– 2921, 2959, 2972, 2974, 2982, 3020, 3029, 3044, 3067, 3081, 3127, 3144, 3156]. resource [3, 14, 27, 90, 236, 261, 318, 441, 529, 652, 690, 721, 733, 785, 945, 981, 1002, 1237, 1238, 1286, 1324, 1357, 1447, 1472, 1485, 1528, 1534, 1666, 1687, 1709, 1728, 1907, 1980, 2004, 2040, 2247, 2285, 2327, 2345, 2399, 2408, 2466, 2475, 2531, 2561, 2589, 2610, 2629, 2632, 2655, 2685, 2696, 2715, 2726, 2756, 2820, 2835, 2850, 2864, 2873, 2916, 2932, 2950, 2981, 3005, 3021, 3056, 3172, 3184, 3186]. resource-allocation [1485]. Resource-Aware [1315]. resource-constrained [10, 824, 2323, 2763, 2972]. Resource-efficient [1143, 1528, 2595, 2920]. Resource-limited [90, 728]. resources [545, 1432, 1579, 1935, 2211, 2255, 2707, 2712, 2925, 3017]. resources-aware [2925]. responding [2158]. response [121, 973, 2408]. restoration [478, 752, 2014, 2304]. restricted [349, 2278]. restrictive [2931]. results [1589, 2033]. retention [2898]. retention-aware [2898]. Rethinking [2383, 2694, 3130]. Retina [923]. Retraction [1762]. retraining [2826]. retransmissions [2410]. Retrieval [1068, 1140, 2339, 2557, 2952]. retrievals [1700]. Reunion [2731]. Reuse [1106, 1918, 1940, 2695, 2696]. Reuse-based [1918]. Reusing [195]. Reveal [1713]. Revelation [599]. Revenue [215, 1746]. reverse [1644, 2372, 2518, 3090, 3185]. reverse-leadership [2372]. Review [107, 178, 180, 284, 472, 795, 894, 941, 978, 1124, 1188, 1209, 1266, 1395, 1483, 1488, 1639, 1934, 1991, 2112, 2116, 2273, 2377, 2436, 2477, 2701, 2796, 2821, 2872, 2929]. revision [287]. revisited [1630]. Revisiting [115, 922, 1114, 1434]. revocable [702]. revocation [1883, 2265, 2620, 3083]. reward [1502, 3152]. reward-enhanced [3152]. rewriting [984]. RF [463, 554, 1178, 1372, 1597, 1903, 2184, 2372, 2973, 2993, 3021]. RF-powered [554, 1178, 2993]. RFC [2637]. RFFI [1916]. RFID [1, 355, 561, 811, 821, 1287, 1300, 1641, 2367, 2648, 3161]. RFID-based [821]. RFID-enabled [2367]. RFTRUST [809]. RICERCANDO [227]. rich [756]. ride [2358]. right [612]. ring [1486, 1849, 2810]. RIS [1099, 1295, 1483, 1573, 1609, 1631, 2130, 2407, 2433, 2563, 2605, 2629, 2650, 2725, 2833, 2921, 3039, 3077, 3089, 3111, 3165]. RIS-aided [1573, 1609, 2650, 2725]. RIS-assisted [1099, 1631, 2130, 2407, 2563, 3077, 3089]. RIS-enhanced [1483]. risk [1598, 1852, 2122, 2297, 2752]. risk-based [2752]. RL [1767]. RL-Adapt [1767]. RLFE [2952]. RLFE-IDS [2952]. RLS [1629]. RNN [2072]. RNN-based [2072].

RNTI [1355]. Road [131, 391, 458, 648, 1001, 1157, 1177, 1571, 2145, 2536]. road-network [1001]. roads [364]. roadside [798]. roaming [1439]. robot [268, 505, 2131]. robotic [57, 1923, 3039]. robotics [194]. robots [1130]. Robust [1, 21, 41, 63, 171, 219, 311, 406, 463, 622, 649, 663, 705, 712, 810, 868, 940, 953, 1218, 1271, 1309, 1311, 1338, 1351, 1363, 1544, 1617, 1669, 1811, 1919, 1929, 2066, 2073, 2125, 2420, 2559, 2666, 2710, 2803, 2984, 3000, 3009, 3023, 3123, 3154]. Robustness [1320, 1346, 1713, 1909, 2273]. Robustness-aware [1713]. RoCEv2 [2179]. ROEN [2963]. Role [1298, 1585, 2369, 2741, 2819]. Role-based [2819]. roles [720, 2652]. RoNS [1218]. root [1402, 2580]. Roots [619]. Roots-tracing [619]. Rosenblatt [1442]. rotation [2494]. Route [237, 286, 714, 1098, 1231]. RouteChain [1302]. router [196, 3010]. Routers [156, 408, 1392, 2464, 2637]. routes [886]. Routing [30, 77, 95, 136, 150, 187, 283, 286, 292, 346, 348, 349, 417, 447, 461, 470, 478, 530, 531, 544, 625, 627, 673, 743, 746, 753, 756, 759, 790, 854, 1031, 1056, 1096, 1107, 1129, 1151, 1157, 1164, 1192, 1194, 1271, 1275, 1299, 1302, 1317, 1323, 1324, 1347, 1386, 1391, 1397, 1403, 1418, 1463, 1468, 1487, 1535, 1538, 1539, 1569, 1600, 1632, 1639, 1681, 1698, 1787, 1794, 1872, 1898, 1952, 1994–1996, 2084, 2099, 2195, 2235, 2406, 2430, 2437, 2464, 2509, 2541, 2556, 2597, 2627, 2647, 2702, 2714, 2717, 2730, 2752, 2759, 2931, 2934, 2942, 2957, 3027, 3108, 3145, 3178]. routing [37, 218, 252, 520, 552, 660, 668, 679, 766, 799, 867, 1004, 1116, 1134, 1253, 1666, 1757, 1832, 1918, 2151, 2234, 2321, 2606, 2619, 2679, 2699, 2852, 2944]. ROV [3145]. ROVReco [3145]. RPL [157, 445, 809, 1326, 2502, 2534, 2559, 2633, 2752, 2804]. RPL-based [157, 809, 2633]. RR [2730]. RRH [365, 1210]. RRH-Sector [1210]. RRH-Sector-BBU [1210]. RSMA [2898, 2921]. RSSI [1543, 2044].

RSSI-based [1543]. RT [1826]. RTC [923]. RTIFed [1970]. RTQS [3010]. RTS [900]. RTS/CTS [900]. Rule [534, 1917, 2608, 2656]. rule-caching [534]. Rumors [661]. running [1949]. runtime [677, 1497, 3142]. rural [1052, 2816].

S [1095, 1423, 1719]. S-FoT [1719]. S-HIDRA [1423]. S2H [899]. S2S [2229]. S2V [2229]. S2X [1835]. SA [657, 3098]. Safe [496, 2941]. safeguarding [3151]. safer [2242]. safety [131, 1157, 2593]. SAFL [2935]. SAGIN [2467, 3187]. SAGTM [2822]. SAIDE [420]. Salience [1817]. Salience-based [1817]. Sample [2579, 2674, 2768]. sampling [1773, 2033, 2642, 2876]. Sanitization [556]. SAPWSN [1387]. SARM [2556]. SARS [2166]. Satellite [79, 150, 163, 410, 734, 1107, 1191, 1198, 1241, 1294, 1535, 1536, 1666, 1687, 1835, 1856, 1911, 1928, 2009, 2171, 2246, 2324, 2398, 2422, 2471, 2538, 2546, 2655, 2657, 2677, 2706, 2711, 2741, 2782, 2862, 2867, 2875, 2982, 2987, 3002, 3108, 3109, 3164]. satellite- [2862]. satellite-aerial [2706]. Satellite-assisted [1856]. Satellite-based [1241, 2246, 2398, 2422, 3109]. satellite-terrestrial [1536, 1666, 2875]. satellites [2762, 3074]. Satisfaction [133, 451, 2004, 3158]. Satisfaction-based [451]. satisfying [326]. saturation [80, 963, 1158]. Save [1162]. saving [1977, 2364, 2712, 2724]. SBERT [1405]. SBPG [846]. SBS [102]. SCADA [464]. ScaIR [2606]. scalability [768, 790, 1244, 2634]. Scalable [35, 56, 99, 152, 229, 285, 400, 536, 571, 603, 849, 916, 927, 960, 1014, 1055, 1126, 1141, 1288, 1454, 1474, 1611, 1619, 1641, 1718, 1917, 1986, 1998, 2191, 2332, 2388, 2559, 2606, 2797, 2839, 2877, 2970, 3065]. Scale [56, 278, 386, 394, 427, 456, 768, 868, 938, 1001, 1116, 1189, 1319, 1435, 1504, 1512, 1514, 1559, 1622, 1881, 1964, 2009, 2127, 2197, 2248, 2391, 2437, 2465, 2502, 2713, 2800, 2845, 2860, 2865, 2934, 3003, 3161]. scale-free [868, 1514, 2197]. ScaleDRL [571]. Scaling [15, 59, 353, 931, 2123, 2450, 3129]. scanner [72]. scanning [943]. scans [1668]. scarce [3130]. ScasDK [2895]. SCDFL [3170]. scenario [782, 1382, 1548, 1729, 1851, 1928, 3070]. scenarios [38, 293, 432, 594, 908, 1185, 1370, 1492, 1719, 1961, 2402, 2485, 2600, 2744, 2846, 2961, 3130, 3134]. SCEP [3177]. SCEP-TI [3177]. Schedulability [2996]. schedule [143, 760, 2936]. Scheduled [205]. scheduler [154, 271, 2086]. Schedulers [2551]. Scheduling [14, 73, 93, 99, 143, 190, 192, 249, 279, 335, 402, 404, 432, 436, 441, 459, 465, 485, 551, 587, 638, 655, 747, 789, 889, 892, 1003, 1020, 1036, 1088, 1096, 1113, 1139, 1161, 1178, 1199, 1211, 1330, 1337, 1357, 1401, 1426, 1430, 1466, 1470, 1494, 1521, 1527, 1554, 1629, 1696, 1698, 1713, 1785, 1794, 1795, 1846, 1858, 1869, 1888, 1910, 1918, 1949, 1960, 1968, 2000, 2005, 2009, 2052, 2070, 2097, 2110, 2124, 2150, 2154, 2164, 2228, 2299, 2310, 2314, 2373, 2375, 2408, 2437, 2550, 2570, 2572, 2723, 2755, 2761, 2777, 2782, 2837, 2860, 2928, 2937, 2949, 2993, 3003, 3010, 3031, 3042, 3092, 3093, 3149]. Scheme [65, 76, 101, 187, 190, 212, 230, 242, 277, 294, 353, 375, 393, 395, 415, 447, 475, 490, 494, 545, 589, 604, 627, 701, 707, 739, 840, 866, 875, 887, 900, 949, 964, 1072, 1103, 1133, 1157, 1159, 1180, 1187, 1192, 1200, 1212, 1222, 1240, 1260, 1267, 1272, 1283, 1287, 1300, 1436, 1486, 1508, 1623, 1632, 1637, 1691, 1718, 1776, 1789, 1812, 1826, 1837, 1906, 2004, 2047, 2067, 2127, 2128, 2167, 2176, 2191, 2239, 2298, 2341, 2366, 2367, 2376, 2393, 2424, 2443, 2456, 2480, 2506, 2536, 2601, 2744, 2748, 2754, 2777, 2780, 2806, 2814, 2822, 2827, 2893, 2970, 2999, 3065, 3068, 3090, 3131, 3134, 3135, 3159, 3175]. scheme [40, 122, 153, 485, 520, 542, 622, 670, 702, 885, 959, 966, 987, 1013, 1252, 1429, 1458, 1503, 1668, 1818, 1834, 1851, 1868, 1900, 2013, 2031, 2223, 2257, 2322, 2431, 2535, 2575, 2617, 2636, 2641, 2689, 2710, 2721, 2767, 2810, 2828, 2847, 2866, 2885, 2936, 3006, 3058, 3063]. Schemes [10, 267, 672, 2145, 2695, 2735, 2887, 3027]. Schnorr [277]. Schnorr-like [277]. science [2010]. scientific [8, 274, 335]. SCM [963]. SCMA [2955]. SCNs [2277]. scoping [1639]. score [2297]. scores [2103]. scoring [374, 1365, 2066, 3157]. scripting [11]. SD [499, 975, 1403, 2071, 2244, 2463, 2601]. SD-CDN [975]. SD-IoT [2601]. SD-MDN-TM [2463]. SD-WAN [499, 2071, 2244]. SD-WANs [1403]. SDBGPChain [1646]. SDDC [2244, 2970]. SDFog [1100]. SDFog-Mesh [1100]. SDHW [543]. SDHW-IoT [543]. SDLoRe [2665]. SDM [177, 1434, 3027]. SDM-EON [177]. SDM-EONs [1434, 3027]. SDN [19, 32, 35, 40, 86, 98, 101, 110, 118, 250, 254, 261, 282, 325, 347, 378, 420, 446, 469, 473, 486, 489, 506, 528, 571, 586, 597, 601, 610, 615, 642, 647, 655, 656, 659, 685, 704, 734, 764, 793, 797, 854, 872, 904, 987, 989, 1017, 1051, 1055, 1088, 1158, 1166, 1177, 1180, 1200, 1246, 1250, 1301, 1346, 1423, 1436, 1525, 1571, 1615, 1646, 1701, 1724, 1732, 1735, 1753, 1837, 1868, 1901, 1908, 1910, 1927, 1929, 1952, 1981, 1995, 2202, 2253, 2271, 2289, 2400, 2403, 2707, 2718, 2941, 2957, 2968, 3059]. SDN-aware [1166]. SDN-based [610, 647, 655, 685, 734, 1436, 1724, 1910, 2202, 2289, 2400, 2707]. SDN-driven [2253]. SDN-enabled [110, 1088, 1732]. SDN-IoV [642]. SDN-MTD [2968]. SDN/NFV [282, 854, 989, 1837]. SDN/NFV-enabled [989, 1837]. SDN/non [597]. SDN/non-SDN [597]. SDNs [550, 800]. SDNSandbox [1364]. SDONs [1477]. SDP [2692]. SDPC [2692]. SDR [19, 334, 1972]. SDR-based [334]. SDR-LoRa [1972]. SDTMA [8]. SDTMA-NDN [8]. SDWSN [654]. SEAF [2175]. SEAF-IoD [2175]. Seamless [489, 1143, 1198, 2185, 2354]. SEANet [2284]. Search [265, 679, 1226, 1552, 1880, 2239, 2469, 2688, 2888]. searchable [875, 1851, 2265]. searching [783]. SecFHome [1027]. SECHO [1623]. SECM [2256]. second [1460, 2860]. secPEFL [3180]. Secrecy [1480, 2831, 3131]. secret [277, 2031, 2810]. section [1407]. SecTopo [3075]. Sector [578, 1210]. sectorization [578]. Secure [56, 122, 242, 254, 258, 263, 265, 475, 476, 503, 505, 521, 599, 612, 647, 673, 707, 708, 721, 749, 765, 770, 831, 837, 840, 858, 875, 882, 885, 965, 1004, 1016, 1019, 1027, 1036, 1087, 1105, 1107, 1115, 1134, 1185, 1197, 1272, 1287, 1300, 1302, 1329, 1334, 1335, 1387, 1437, 1486, 1551, 1556, 1623, 1681, 1683, 1684, 1718, 1757, 1764, 1878, 1887, 1900, 1914, 1951, 2017, 2085, 2119, 2143, 2175, 2215, 2224, 2246, 2257, 2298, 2312, 2338, 2371, 2394, 2396, 2398, 2431, 2451, 2456, 2505, 2506, 2521, 2523, 2557, 2587, 2692, 2708, 2744, 2748, 2776, 2822, 2832, 2837, 2885, 3058, 3065, 3068, 3079, 3083, 3126, 3165, 3168]. SECURE [2195]. Secure-aware [476]. secured [512, 845, 2261, 2397, 2543, 2714, 2925, 3030]. SecureEI [2514]. Securely [2256]. Securing [407, 656, 768, 780, 1093, 1733, 2055, 2454, 3115]. Security [81, 82, 84, 155, 264, 276, 332, 355, 412, 424, 450, 458, 467, 492, 527, 532, 565, 590, 606, 617, 620, 624, 634, 641, 662, 754, 771, 774, 809, 833, 846, 947, 980, 986, 992, 1007, 1014, 1017, 1022, 1026, 1046, 1093, 1111, 1133, 1148, 1166, 1171, 1241, 1258, 1313, 1316, 1417, 1425, 1509, 1530, 1603, 1615, 1639, 1724, 1763, 1819, 1934, 1975, 1981, 2024, 2067, 2112, 2146, 2259, 2378, 2390, 2400, 2452, 2472, 2490, 2634, 2639, 2702, 2718, 2725, 2752, 2757, 2769, 2867, 2871, 2875, 2895, 2939, 2945, 2946, 2977, 3044, 3069, 3137, 3151, 3156, 3157, 3180]. security [287, 842, 884, 1429, 1514, 2027, 2222, 2392, 2508, 2631, 2677, 2721, 2804, 2853, 2882, 2892, 3070, 3093]. Security-aware [1258, 2508, 2631, 2752]. security-enhanced [992, 2721]. security-related [2024]. Security-reliability [2875]. SED [1718]. SED-SGC [1718]. segment [417, 867, 1116, 2665, 2852, 2942].

Segmentation [268, 494, 1186, 2195, 2227]. segmented [1236, 2546]. SEI2RS [2208]. select [308]. selected [1195, 1647, 2710]. selecting [1419, 1459, 2662]. Selection [22, 53, 119, 144, 158, 182, 266, 302, 315, 349, 384, 413, 445, 456, 481, 524, 568, 654, 753, 791, 800, 819, 837, 857, 894, 904, 974, 1034, 1079, 1089, 1178, 1184, 1210, 1225, 1229, 1231, 1272, 1303, 1321, 1434, 1445, 1485, 1533, 1542, 1611, 1686, 1701, 1722, 1770, 1787, 1837, 1889, 1895, 1954, 1977, 1994, 2015, 2071, 2101, 2237, 2245, 2302, 2344, 2456, 2472, 2579, 2589, 2656, 2755, 2791, 2841, 2863, 2868, 3021, 3036, 3067, 3085]. Selective [533, 610, 1741, 2214, 3005, 3133]. selective-edge [533]. Self [717, 737, 828, 944, 1002, 1051, 1062, 1297, 1299, 1361, 1444, 1547, 1645, 1714, 1722, 1768, 1770, 1825, 2038, 2212, 2495, 2815, 2831, 2977, 2985, 3002]. Self-adaptive [1051, 1547, 2815, 2977]. self-adjusting [1768]. self-attention [1770]. Self-attentive [737]. Self-backhauling [1297]. self-coordination [717]. self-driven [944]. Self-healing [1299, 2495]. self-knowledge [2985]. Self-organizing [717, 828, 1714, 1722]. Self-play [1002]. self-protection [1444]. Self-similar [3002]. self-sovereign [1361]. self-stabilizing [2038]. self-supervised [2212]. self-sustainable [1825]. self-tallying [2831]. Self-verifiable [1645]. Selfied [2573]. selfish [379]. Semantic [575, 691, 2227, 2268, 2343, 3020, 3095, 3185]. Semi [4, 538, 872, 1025, 1100, 1567, 2062, 2133, 2329, 2935, 3061]. Semi-asynchronous [2935]. semi-parallel [2133]. semi-permanent [1100]. semi-supervised [4, 538, 872, 1025, 2062, 2329, 3061]. semi-synthetic [1567]. Sensing [54, 55, 192, 308, 336, 444, 600, 735, 924, 1455, 1744, 2129, 2199, 2358, 2528, 2587, 2641, 2694, 2705, 2746, 2820, 2824, 3074, 3089, 3096, 3117, 3158]. Sensing-based [3089]. sensitive [337, 486, 1265, 1426, 1497, 1554, 1658, 1696, 1754, 1794, 1851, 2026, 2049, 2091, 2210, 2219, 2268, 2287, 2331, 2334, 2339, 2393, 2527, 2582, 2681, 2761, 2765, 2955, 2996]. Sensor [10, 28, 33, 37, 41, 43, 83, 104, 120, 180, 202, 239, 252, 262, 275, 289, 298, 340, 352, 481, 508, 621, 679, 711, 713, 725, 743, 759, 810, 868, 897, 927, 960, 1031, 1038, 1039, 1079, 1081, 1103, 1129, 1134, 1164, 1212, 1231, 1270, 1272, 1277, 1282, 1299, 1319, 1349, 1354, 1387, 1424, 1437, 1463, 1546, 1681, 1689, 1764, 1806, 1813, 1831, 1832, 1872, 2015, 2068, 2254, 2320, 2321, 2476, 2509, 2619, 2621, 2717, 2727, 2728, 2757, 2759, 2783, 2795, 2824, 2887, 3077]. sensors [322, 626, 783, 1592, 2278, 2988]. sentence [585]. sentence-level [585]. sentinel [3137]. September [270, 741, 1204, 1571, 1726, 2347, 3033]. sequences [503, 1399, 2603]. sequential [96, 223, 1700, 2549]. serendipity [372]. serial [1247, 2602]. serialized [1902]. Series [152, 1319, 1534, 1948, 2050, 2315, 2976]. serve [1752]. served [1803]. server [231, 328, 411, 471, 484, 814, 1065, 1118, 1782, 2138, 2219, 2224, 2302, 2364, 2386, 2396, 2590, 2961, 2992, 3067]. server-side [1782]. Serverless [145, 1090, 2738]. servers [1642, 2708]. Service [15, 16, 21, 148, 160, 164, 168, 190, 250, 253, 260, 311, 319, 371, 378, 395, 584, 676, 680, 692, 814, 907, 919, 950, 960, 974, 983, 1016, 1018, 1051, 1080, 1121, 1126, 1138, 1139, 1155, 1161, 1227, 1230, 1251, 1258, 1268, 1368, 1371, 1395, 1405, 1417, 1422, 1428, 1444, 1469, 1484, 1502, 1526, 1540, 1579, 1654, 1735, 1740, 1783, 1791, 1795, 1833, 1929, 1936, 1980, 1982, 1983, 1988, 2014, 2047, 2053, 2064, 2091, 2144, 2146, 2172, 2173, 2206, 2260, 2280, 2288, 2304, 2399, 2406, 2417, 2438, 2442, 2448, 2462, 2475, 2515, 2572, 2599, 2619, 2640, 2691, 2715, 2826, 2917, 2939, 3112]. service [66, 118, 363, 510, 558, 593, 650, 1106, 1318, 1430, 1456, 1457, 1459, 1491, 1515, 1536, 1703, 2021, 2026, 2211, 2270, 2426, 2508, 2610, 2631, 2635, 2840, 2864, 2891, 2932, 2954]. Service-aware [168, 2173, 2619]. Service-Based [983, 1018]. Service-centric [584]. service-oriented [160]. Services [16, 78, 175, 248, 357, 364, 475, 727, 818, 918, 1087, 1111, 1159, 1177, 1218, 1263, 1385, 1439, 1524, 1571, 1659, 1924, 1936, 1982, 2079, 2152, 2219, 2354, 2383, 2526, 2626, 2697, 2771, 2853, 2861, 2933, 3184, 3185]. serving [266]. session [189, 1584, 2687]. sessions [374, 2052]. Set [1072, 1347, 1789, 1811, 2329, 2493, 2513]. SETCAP [1018]. SETL [2385]. setter [899]. setting [277]. setup [2256]. SFACIF [2593]. SFC [2228, 2276, 2335, 3127, 3183]. SFCAG [2426]. SFCs [1803, 2183, 2707]. SGC [1718]. SGD [734]. SGs [2681]. SHA [768]. SHA-256 [768]. Shadowsocks [940]. shallow [2264]. Shannon [575]. SHAP [3118]. SHAPARAK [400]. shape [1597, 3097]. shaped [2188]. shapelet [152]. shaper [1794, 2996]. Shapley [3158]. Shapley-value [3158]. SharAcc [2634]. sharding [2095, 2450, 2634, 2980]. sharding-based [2634]. share [92, 2374]. Shared [297, 484, 566, 836, 1270, 1840, 1853, 2211]. Sharing [12, 20, 123, 215, 223, 383, 392, 405, 765, 875, 958, 1048, 1094, 1105, 1141, 1495, 1622, 1844, 1851, 2016, 2031, 2065, 2119, 2160, 2389, 2398, 2523, 2535, 2536, 2696, 2748, 2845, 2869, 2876, 2885, 2916, 3083, 3085, 3126, 3159, 3175]. SHARP [599]. shelf [600]. Shell [1743]. shift [2829]. Shining [1054]. ship [2207]. shipping [505]. shop [587]. Shore [2229]. Shore-to-Shore [2229]. Shore-to-Vessel [2229]. Short [658, 1778, 2229, 2278, 2498, 2990]. Short-range [2229]. shot [779, 1377, 1599, 1628, 2683, 3080]. should [767]. shuffling [842]. shutdown [2364]. SIC [1131, 1736]. Side [719, 805, 807, 984, 1041, 1065, 1239, 1782, 3051, 3177]. Side-channel [719, 805, 807, 1041, 1239, 3051, 3177]. sided [674]. sidelink [64, 2218, 2440, 2561, 2907]. sidelink-based [64]. SIDS [1820]. SIEM [641, 2918]. SIEVE [2918]. sightings [1602]. sigmoid [1869]. sigmoid-based [1869]. signal [832, 920, 1208, 2109, 2668, 2902]. signaling [755, 1530]. signals [935, 963, 1617, 2182, 2470]. Signature [1185, 1486, 2483, 2530, 2575, 2810]. signature-based [2810]. signatures [245, 1849]. Signcryption [1103, 1790, 2246, 2536, 3006, 3048]. signed [3030]. silent [1196]. Silo [3015]. SIM [1965, 2217, 2635]. similar [3002]. Similarity [895, 1806, 1984, 2163, 2420, 2469, 2899]. similarity-aware [2163]. SIMO [1675, 1759]. SIMO-Underwater [1675, 1759]. simple [31]. Simulation [528, 1316, 1771, 1911, 2196, 2412, 2481, 2635]. simulations [536, 2532, 2574]. simulator [149, 234, 1896, 2499]. Simulators [1752, 1860, 2821]. Simultaneous [158, 1430, 2358]. sine [1164]. single [651, 674, 832, 1541, 2574, 2664, 2672]. single-output [651]. single-queue [2664]. single-sided [674]. single-user [2574]. sink [1437, 1764, 2759]. sinkhole [809]. SINR [2830]. SINR-aware [2830]. SIP [189, 947, 1947]. SIP-based [1947]. SIP-enabled [947]. Site [11, 1422, 2244, 2663]. Site-trained [2663]. situation [882]. six [1848]. six-layer [1848]. Size [402, 1789, 2188]. Size-based [402]. sizing [2657]. SK [2794]. SK-CFR [2794]. skeleton [2084]. Sketch [683, 890, 1276, 1563, 1865, 2188]. skew [10]. Skipping [204]. Skipping-based [204]. SlaaS [1818]. Slardar [1199]. slate [1360]. sleep [315, 893, 1629]. sleeping [231, 1400]. Slice [1202, 1331, 1510, 1515, 1976, 2152, 2549, 2592, 3093]. sliceable [871]. sliced [39]. slices [114, 529, 731, 1228, 1817, 1821, 2960]. Slicing [19, 32, 228, 517, 667, 859, 907, 1215, 1265, 1267, 1545, 1709, 1735, 1755, 1816, 1821, 1889, 2070, 2173, 2233, 2577, 2580, 2707, 2715, 2771, 2936, 3017]. sliding [1080, 1563]. slope [462]. slot [1887]. slot-pairing [1887]. slots [698, 2278]. Slotted [818, 1244, 1748]. slotted-ALOHA [818]. Slow [164]. Slowloris [2943]. SM9 [3006]. small [128, 415, 562, 888, 1019, 1622, 1844, 3019]. small-cell [1844, 3019]. Smart [46, 47, 58, 62, 82, 95, 129, 140, 209, 321, 487, 505, 512, 548, 577, 585, 587, 611, 625, 641, 652, 703, 721, 749, 787, 831, 946, 990, 1027, 1074, 1100, 1104, 1115, 1141, 1222, 1269, 1383, 1601, 1625, 1646, 1742, 1777, 1783, 1853, 1923, 2011, 2013, 2043, 2082, 2232, 2253, 2257, 2453, 2455, 2468, 2513, 2529, 2603, 2636, 2676, 2747, 2753, 2856, 2877, 3084, 3137, 3157]. SmartDefense [1060]. SmartFCT [272]. SmartNIC [1082, 2478]. SmartNIC-embedded [2478]. SmartNICs [2817]. smartphone [608, 903, 1916]. smartphones [206]. SmartSBD [1853]. SmartTRO [1320]. SMASH [2968]. SMOTE [235]. SMS [2558]. SMUSAC [3173]. SNARKs [2030]. sniffing [18]. Snort [1163]. Snowflakes [2887]. Social [9, 22, 89, 96, 183, 188, 246, 338, 379, 405, 418, 444, 552, 661, 762, 873, 894, 990, 1102, 1118, 1274, 1356, 1383, 1409, 1562, 1820, 1867, 1884, 2088, 2239, 2311, 2330, 2369, 2453, 2662, 2691, 2787, 2815, 2945, 3080, 3171]. Social-aware [9, 188, 405, 1867]. Socially [2818]. Society [1078]. SOFI [1234]. Soft [2361, 2786, 2794, 3164]. SoftCaching [1787]. Software [8, 80, 84, 95, 156, 164, 218, 222, 263, 325, 356, 408, 423, 437, 442, 470, 476, 508, 516, 520, 536, 537, 543, 593, 609, 623, 624, 662, 673, 690, 745, 790, 849, 981, 1037, 1069, 1100, 1105, 1121, 1165, 1183, 1254, 1313, 1329, 1540, 1610, 1640, 1688, 1749, 1787, 1805, 1875, 1901, 1908, 1909, 1972, 1994, 2083, 2108, 2112, 2206, 2284, 2290, 2293, 2370, 2488, 2524, 2533, 2730, 2951, 2991, 3004, 3114, 3116]. Software-centric [2370]. Software-Defined [80, 84, 95, 222, 442, 470, 476, 508, 520, 593, 609, 623, 662, 690, 849, 1100, 1121, 1183, 1254, 1540, 1640, 1749, 1787, 1805, 1875, 1908, 1909, 2108, 2206, 2284, 2290, 2730, 3004, 3114]. Software-Defined-Radios [1972]. softwarization [1991, 2564]. Softwarized [251, 301, 539, 1359, 2081, 2427]. solar [796, 2673]. solar-aware [796]. solar-powered [2673]. solution [63, 67, 205, 286, 362, 463, 581, 586, 990, 1097, 1146, 1782, 2080, 2114, 2216, 2483, 2488, 2626, 3167]. Solutions [78, 82, 357, 412, 696, 704, 1124, 1241, 1357, 1409, 1442, 1496, 1751, 1850, 2088, 2702, 2832, 2946, 3159]. Solving [284, 348, 621, 2177]. SOREC [1714]. SØREN [1451]. soundscape [711]. Source [30, 132, 149, 192, 647, 923, 1163, 1200, 1499, 1621, 1972, 1982, 2019, 2117, 2250, 2887, 2951, 3163, 3171]. sourced [1232, 1747]. sources [129, 903]. sovereign [1361]. Sovereignty [2583]. SP [2133, 2299]. SP-DG [2299]. SP-PoR [2133]. SP2F [512]. Space [170, 177, 346, 528, 772, 1013, 1056, 1239, 1324, 1393, 1458, 1488, 1703, 1799, 1963, 2075, 2076, 2099, 2148, 2223, 2234, 2467, 2537, 2549, 2617, 2658, 2659, 3127, 3136]. space-aerial-terrestrial [170]. space-air-ground [2076, 2467, 2658]. Space-code [2148]. space-division [1324, 1393, 1703]. space-efficient [772]. Space-Ground [528, 1013]. space-terrestrial [2659]. spaces [1704]. Spam [409, 835, 1021, 2558]. span [2187]. spanners [2623]. Sparse [1025, 1189, 2190, 2641, 2682, 2820]. sparsification [1626]. Spatial [177, 468, 598, 778, 1278, 1491, 1556, 1940, 1974, 2033, 2059, 2136, 2277, 2381, 2695, 2830, 2954, 3133, 3139]. Spatial-Temporal [2059, 2381]. spatially [1894, 2028]. Spatio [113, 332, 1792, 2061, 2226, 2274, 2365, 2820, 2864, 3060]. Spatio-temporal [113, 332, 1792, 2061, 2226, 2274, 2365, 2820, 2864, 3060]. spatiotemporal [561, 589, 1256, 2432, 2641]. SPC [933]. SPDTM [3158]. SPEAR [641]. Special [662, 1407, 1583, 1647, 3012]. Specific [746, 1978, 3107]. specification [260, 1875]. Spectral [310, 560, 1308, 1637, 1913, 2754, 3170]. Spectral-Energy [310]. spectrally [871, 1894, 2028]. spectrally-spatially [1894, 2028]. Spectrum [12, 111, 136, 176, 177, 224, 230, 232, 308, 334, 346, 348, 392, 405, 440, 481, 491, 545, 652, 701, 724, 731, 753, 933, 1048, 1056, 1323, 1396, 1477, 1511, 1667, 1704, 1739, 1858, 2129, 2234, 2389, 2480, 2584, 2739, 2760, 2845, 2944, 2974, 3027, 3074, 3109, 3159]. SpectrumChain [1477]. speculative [498]. speech [2103]. speed [381, 404, 494, 644, 700, 1668, 1865, 1892, 1948, 2675, 2890, 3026]. Speeding [1012]. speeds [2401]. speedtest [427]. Sphinx [700]. Split [1264, 1315, 1801, 3067, 3104]. SplitFed [3046]. SplitPath [1031]. Splittable [1193]. Splitting [137, 275, 705, 1308, 1678, 1968, 2005, 2035, 2538, 2592]. Spoiler [2251]. spoke [352]. Spoofed [368]. Spoofing [112, 1232, 2463, 2987]. spray [552]. spray-and-wait [552]. spread [3109]. spreader [723]. spreaders [1865]. spreading [384]. Spring [1072]. SQHCP [476]. SQP [2799]. SQP-driven [2799]. squares [28]. Squatting [39]. SR [867]. SRE [846]. SRLG [678]. SRLG-aware [678]. SRRS [1362]. SRv6 [469, 1741]. SSAPPIDENTIFY [940]. SSDI [881]. ST [2864]. ST-RDP [2864]. Stability [1004, 1812, 3111]. Stability-aware [1812]. stabilizing [2038]. Stable [445, 1781, 2903]. stack [1154, 1619]. stacked [1025, 2708]. Stackelberg [1744, 2107, 2147, 2294, 2805]. Stackelberg-evolutionary [2107]. stacking [431]. stage [190, 437, 441, 868, 1032, 1330, 1706, 1711, 1774, 2281, 2289, 2894, 2992]. staged [1581]. Stake [822, 2018]. stakeholder [424]. stakeholder-oriented [424]. stale [1084]. staleness [2115]. stalling [2416]. stand [329]. stand-alone [329]. standalone [1985, 2250]. standard [1597, 1642]. standardisation [410]. standards [1361, 2116]. STAR [2605, 2629, 3039, 3111]. STAR-RIS [2629, 3039]. STAR-RIS-aided [2605]. STAR-RIS-assisted [3111]. StarCross [2398]. Start [1950]. Start-Time [1950]. starvation [2856]. State [391, 399, 458, 471, 537, 615, 646, 802, 1007, 1227, 1263, 1347, 1430, 1488, 1531, 2027, 2549, 2556, 2775, 2873, 3012]. State-Aware [2556]. State-of-art [1488]. State-of-the-Art [391, 458, 537, 615, 646, 1007, 1531, 2027, 3012]. Stateful [7, 423, 2307]. stateless [2302]. States [429]. static [901, 984, 1005]. Station [5, 231, 632, 818, 1691, 1772, 1951, 2074, 2724, 2771, 3001, 3107]. stationary [300]. stations [109, 110]. statistic [2876]. statistical [1737]. Status [958]. Steerable [988, 1073]. Steerable-Beam [1073]. steering [30, 1968, 2707]. stegomalware [607]. Step [165, 778, 782, 1303, 1970, 2465]. STM32 [2024]. Stochastic [364, 370, 573, 2025, 2074, 2252, 2764, 2967, 3015]. stock [639]. storage [650, 966, 1132, 1252, 1431, 1672, 1776, 1840, 2199, 2265, 2371, 2423, 2499, 2698, 3030, 3087, 3134]. store [411]. stores [2817]. straight [2669]. straight-line [2669]. strategies [172, 210, 829, 1002, 1316, 1355, 1531, 1589, 1638, 1716, 2050, 2230, 2260, 2263, 2344, 2353, 2387, 2455, 2461, 2462, 2562, 2648, 2760, 2931, 2938]. Strategy [23, 93, 118, 249, 257, 339, 509, 654, 663, 712, 796, 807, 831, 834, 841, 1040, 1085, 1134, 1135, 1242, 1309, 1400, 1522, 1536, 1642, 1701, 1710, 1720, 1742, 1859, 1887, 1898, 1903, 1945, 2076, 2098, 2135, 2226, 2272, 2324, 2372, 2437, 2438, 2464, 2476, 2534, 2590, 2597, 2609, 2643, 2684, 2713, 2727, 2804, 2850, 3049, 3084, 3093, 3127, 3154]. Stratum [2457]. strawberry [47]. Stream [341, 561, 804, 1606, 2210, 2217, 2527]. stream-based [561]. Streaming [171, 455, 598, 603, 865, 1131, 1332, 1349, 1360, 1388, 1616, 1767, 1782, 1845, 1999, 2079, 2087, 2198, 2317, 2416, 2498, 2630, 2898, 3018, 3029]. Streams [341, 890, 1033, 1277, 1865, 1892, 2588, 2644, 2860, 3177]. strength [920]. Strengthening [3180]. strict [30, 2302]. stricter [515]. stricter-than-ICNIRP [515]. STRIDE [1026]. STRIDE-based [1026]. Strike [1921]. strong [2018, 2681, 3009]. strongly [885]. structural [936]. structure [31, 259, 572, 684, 949, 1763, 1806, 1899, 2698]. Structured [891, 966, 1395, 1721, 2560]. structures [674, 1978]. Study [9, 47, 161, 162, 386, 532, 629, 783, 886, 1033, 1051, 1104, 1170, 1355, 1383, 1418, 1529, 1689, 1783, 1861, 2036, 2044, 2123, 2147, 2249, 2278, 2316, 2546, 2600, 2633, 2745, 2824, 2833]. Studying [1719, 1791]. sub [334, 388, 783, 1265, 1335, 1409, 1492, 1751, 2052, 2267, 2783, 3087]. sub-6 [334, 783]. sub-action [1265]. sub-blocks [3087]. sub-network [2783]. sub-terahertz [388]. sub-THz [2908]. sub-tree [2052]. Subchannel [43]. subject [33]. subjective [2103]. subliminal [277]. submarine [2007]. subscribe [840, 939, 1360, 1651, 1679]. Subscriber [676]. subscribers [2442]. subscription [1545]. substations [356, 438]. substitution [3036]. substrate [1541]. success [1345]. successful [142]. suitability [606]. suitable [308, 1554, 2666]. Sum [1190, 1809, 2092, 2604, 3077, 3084]. sum-rate [2092]. summarization [2217]. SUNETs [3173]. Super [871, 1327, 1863, 1865, 2198]. super-channels [871]. Super-resolution [2198]. super-spreaders [1865]. superlattice [1718]. supervised [4, 538, 872, 1025, 2062, 2212, 2329, 3061]. supplies [965]. supply [355, 873]. Support [225, 771, 1071, 1141, 1202, 1294, 1303]. supported [2996]. Supporting [241, 314, 380, 539, 913, 1082, 1451, 1747, 1771, 2691]. surface [1147, 1375, 2035]. surfaces [1986, 2048, 2196, 2444, 3168]. Surgery [1923, 2740]. Suricata [1163]. surrogate [1349]. surrogate-assisted [1349].

Surveillance

[573, 881, 1473, 1532, 1544, 2468, 2546, 3072]. Survey [32, 58, 81, 109, 189, 198, 263, 306, 336, 367, 410, 467, 493, 527, 548, 581, 615, 618, 620, 646, 682, 690, 717, 720, 727, 742, 754, 828, 852, 901, 930, 980, 995, 1007, 1015, 1017, 1026, 1028, 1048, 1069, 1119, 1120, 1128, 1154, 1179, 1372, 1425, 1496, 1509, 1513, 1538, 1585, 1615, 1643, 1705, 1715, 1716, 1752, 1788, 1815, 1825, 1828, 1860, 1882, 1932, 1938, 2008, 2100, 2184, 2197, 2344, 2359, 2415, 2434, 2446, 2461, 2473, 2484, 2583, 2649, 2650, 2652, 2741, 2757, 2769, 2795, 2807, 2858, 2871, 2884, 2946, 3037, 3044, 3050, 3059, 3098, 3147, 3169]. survey [11, 20, 170, 298, 387, 412, 423, 438, 474, 596, 605, 624, 696, 760, 761, 827, 884, 893, 1066, 1142, 1235, 1241, 1417, 1563, 1597, 1638, 1802, 1848, 1939, 2230, 2311, 2340, 2389, 2648, 2651, 2793, 2991, 3095, 3122, 3136, 3151]. Surveying [2378]. Survivability [1107, 1129, 1164, 1407, 2022]. survivable [678, 2234]. suspicious [374]. Sustainability [516, 3122]. Sustainable [251, 372, 827, 1197, 1825, 2349, 2719, 3028, 3176]. Sustaining [904]. SUTSEC [254]. SVC [598]. SVM [151, 610, 1205]. Swarm [670, 736, 811, 880, 978, 1299, 1657, 1663, 2254, 2295, 2959, 3082]. swarms [46]. sweeping [2480]. SWIPT [373, 1281, 2092, 2563, 2587]. SWIPT-empowered [1281]. SWIPT-enabled [373, 2092]. switch [987, 1183, 1217, 1701, 1739, 2441]. switch-migration [1183]. switch-off [1739]. switch-to-controller [1701]. Switches [45, 261, 537, 784, 1026, 1037, 1498, 1643, 2019, 2355, 2401, 2588, 2920, 2922]. switching [65, 275, 632, 674, 760, 792, 850, 1782, 1901, 1968, 2277, 2616]. Switzerland [1245]. Sybil [822, 1499, 1577, 2573]. symbiosis [2678]. Symmetric [326, 794, 1334]. symmetric-key [794, 1334]. symposium [1195]. Symptom [1234]. Symptom-fault [1234]. SYN [1656, 2083]. synchronisation [1624]. synchronization [10, 148, 291, 471, 1462, 1566, 1774, 2393, 2399, 2422]. synchronized [1604]. Synchronizing [1250]. synchronous [1325, 1348]. SYNCOP [473]. synergistic [2308]. syntax [2680]. Synthesis [2337, 2496, 3141]. synthetic [1009, 1567]. SYNTROPY [2083]. System [52, 57, 86, 141, 182, 374, 437, 442, 464, 467, 527, 574, 579, 602, 609, 610, 626, 630, 641, 686, 706, 792, 858, 863, 881, 883, 916, 924, 932, 939, 940, 992, 1052, 1104, 1115, 1206, 1216, 1361, 1473, 1542, 1580, 1595, 1630, 1675, 1717, 1759, 1809, 1824, 1835, 1838, 1871, 1911, 1923, 1929, 2035, 2085, 2102, 2130, 2132, 2154, 2197, 2281, 2306, 2325, 2334, 2376, 2405, 2433, 2443, 2507, 2525, 2552, 2563, 2590, 2620, 2638, 2655, 2673, 2736, 2749, 2894, 2921, 2936, 2945, 2950, 2952, 2964, 3028, 3055, 3087, 3111, 3119, 3146]. System-level [1911, 2405]. Systematic [284, 472, 774, 894, 978, 1188, 1716, 1934, 2273, 2872, 2929, 2991]. Systematizing [1981]. Systems [62, 72, 113, 145, 215, 255, 264, 343, 380, 392, 438, 452, 461, 463, 466, 519, 553, 582, 749, 751, 769, 778, 806, 833, 837, 946, 948, 978, 1041, 1063, 1066, 1071, 1131, 1197, 1229, 1291, 1316, 1319, 1335, 1366, 1409, 1480, 1483, 1489, 1547, 1556, 1598, 1609, 1621, 1633, 1679, 1685, 1705, 1715, 1751, 1776, 1778, 1847, 1848, 1864, 2008, 2025, 2168, 2208, 2230, 2238, 2242, 2258, 2267, 2283, 2309, 2351, 2386, 2390, 2407, 2412, 2510, 2562, 2577, 2587, 2615, 2666, 2701, 2709, 2725, 2754, 2833, 2838, 2878, 2884, 2929, 2955, 2956, 2979, 2987, 3012, 3016, 3081, 3088, 3161]. systems [73, 119, 129, 322, 345, 588, 662, 733, 966, 1095, 1235, 1295, 1412, 1445, 1472, 1576, 1658, 1661, 1680, 1850, 1947, 1954, 2021, 2101, 2166, 2247, 2367, 2629, 2733, 2776, 2829, 2908, 2993, 3039].

T [2223, 2637]. TABDeep [2052]. table [156, 2019, 2171, 3134]. table-based [156]. tables [50, 156, 1908]. Tabu [265]. tabular [2482]. Tactile [1161, 1652, 1923]. tag [821, 2716, 3161]. tags [1300]. tail [767, 2598]. tailed [2975]. tailored [517, 2900]. Tailoring [1729]. Tally [2215]. tallying [2831]. TAMBUS [407]. Taming [2172]. tampering [2366]. tangle [819]. TAO [3042]. Target [28, 810, 1562, 2366]. targets [401, 3106]. Task [133, 211, 279, 309, 343, 366, 398, 401, 588, 611, 637, 638, 689, 695, 696, 789, 815, 832, 839, 869, 1003, 1036, 1064, 1085, 1088, 1123, 1179, 1193, 1248, 1337, 1466, 1472, 1475, 1485, 1491, 1494, 1536, 1568, 1599, 1633, 1711, 1742, 1744, 1758, 1760, 1831, 1906, 1974, 1975, 2040, 2113, 2135, 2155, 2232, 2262, 2288, 2296, 2310, 2313, 2331, 2334, 2448, 2461, 2466, 2471, 2518, 2576, 2643, 2679, 2681, 2685, 2686, 2726, 2732, 2777, 2799, 2806, 2835, 2840, 2878, 2888, 2901, 2938, 2950, 2979, 2992, 3011, 3031, 3042, 3049, 3078, 3081, 3099, 3158, 3182, 3187]. task-dependency [2835]. task-processing [637]. tasks [73, 432, 1359, 1394, 1470, 1475, 1579, 1677, 1731, 1910, 1949, 2070, 2223, 2937, 2949, 2976]. tat [1181]. taxonomical [2477]. Taxonomy [32, 260, 472, 557, 672, 754, 1209, 2652, 2793]. TbDd [2095]. TCN [2839]. TCP [162, 196, 402, 428, 518, 803, 1119, 1500, 1729, 1928, 2083, 2182, 2362]. TCP/AQM [196]. TD [1403]. TD3 [2314, 2552]. TD3-based [2314, 2552]. TDD [1680]. TDMA [124, 736, 3089]. TE [35, 1052]. TEA [1916]. TEA-RFFI [1916]. team [444]. tech [631]. Technical [234, 372, 1585]. technique [159, 315, 446, 630, 762, 833, 1201, 1277, 1303, 2039, 2187]. Techniques [62, 163, 529, 581, 716, 876, 893, 995, 1007, 1033, 1179, 1231, 1643, 1788, 2014, 2087, 2112, 2214, 2368, 2389, 2473, 2651, 2741, 2871, 2991, 3050, 3053, 3109, 3169]. technological [3147]. technologies [194, 362, 577, 581, 629, 1066, 1580, 1777, 1882, 2116, 2653, 2818, 2858, 2989, 3091]. Technology [646, 768, 1518, 1528, 1735, 1770, 1850, 2261, 2405]. telecom [667]. telecommunication [994, 1575, 3122]. Telegram [2915]. Telemetry [307, 483, 493, 1228, 1233, 1392, 1471, 1586, 1988, 2033, 2082, 2292, 2611, 2983, 3114]. Temperature [1916]. Temporal [113, 332, 468, 934, 1018, 1328, 1476, 1792, 2059, 2061, 2226, 2274, 2365, 2381, 2581, 2682, 2713, 2772, 2792, 2820, 2864, 3060, 3139, 3153]. temporal- [1328]. temporal-frequency [2682]. tenancy [232]. tenant [66, 353, 1258, 1267, 2577, 2857, 2949]. tenant-driven [1267]. tensor [651, 860]. tensor-information [651]. terabit [2401, 2857]. Terahertz [288, 292, 388, 935, 1187, 2470]. TeraNova [288]. Term [658, 1469, 1761, 2245, 3111]. termination [1708]. ternary [2190]. Terrestrial [170, 1536, 1666, 1670, 2164, 2336, 2359, 2405, 2547, 2656, 2657, 2659, 2677, 2867, 2875, 2982]. terrestrial-satellite [2677]. territories [2777]. test [1779, 2895, 3051]. testbed [288, 653, 1749]. testbeds [820, 1604, 1860]. tester [2637]. testing [334, 466, 1410, 2123, 2222, 2882]. text [585, 1236, 2720, 2781]. their [114, 1005]. Themis [2517]. Theorem [2827]. theoretic [215, 785, 943, 1097, 2124, 2522]. theoretical [617, 1151, 2745]. theoretical-plus-experimental [2745]. Theory [257, 654, 696, 835, 1242, 1309, 1338, 1401, 1516, 1859, 2033, 2076, 2353, 2417, 2495, 2755, 2919, 2975]. theory-based [654, 2417, 2495, 2755]. thermal [1507, 2550]. thermal-aware [1507, 2550]. Things [20, 157, 160, 161, 175, 183, 191, 207, 246, 249, 258, 276, 306, 412, 433, 445, 447, 450, 453, 465, 487, 495, 517, 518, 521, 548, 582, 584, 633, 660, 765, 768, 770, 787, 795, 864, 876, 885, 894, 939, 941, 960, 964, 1005, 1039, 1059, 1128, 1156, 1248, 1296, 1315, 1320, 1326, 1333, 1398, 1436, 1513, 1538, 1583, 1652, 1676, 1697, 1721, 1787, 1793, 1802, 1820, 1829, 1854, 1856, 1912, 1991, 2100, 2119, 2147, 2176, 2209, 2219, 2246, 2284, 2311, 2328, 2330, 2338, 2471, 2525, 2610, 2633, 2741, 2777, 2815, 2850, 2896, 2991, 3046]. third [36]. Thomas [102]. Thought [2720]. thousands [2860]. Threat [566, 609, 754, 882, 1453, 1495, 1602, 1702, 2122, 2297, 2374, 2434, 2504, 2968, 3076, 3125]. Threats [82, 100, 986, 1093, 1241, 1417, 1429, 1567, 1639, 2392, 2453, 2832, 2892, 2945, 3151]. three [230, 754, 1274, 1569, 1806, 2128, 2176, 2283, 2294, 2642]. three-dimensional [754, 1806, 2642]. three-factor [2176]. three-layer [2294]. three-phase [230]. three-tier [1274, 2128, 2283]. Threshold [793, 1677, 2724, 2809]. Threshold-based [2724]. Throughput [2, 83, 482, 745, 786, 969, 1010, 1031, 1306, 1358, 1446, 1691, 1809, 2084, 2231, 2802, 2977, 2993, 3077]. THz [2908]. TI [3177]. tie [1884]. tie-driven [1884]. tier [219, 225, 369, 814, 1274, 1277, 1280, 2128, 2283, 2577, 3129]. Tiger [2215]. tile [598]. Time [10, 124, 152, 177, 207, 214, 275, 379, 486, 619, 725, 726, 850, 889, 957, 1141, 1178, 1181, 1194, 1219, 1239, 1303, 1319, 1349– 1351, 1398, 1426, 1488, 1497, 1514, 1534, 1547, 1595, 1604, 1624, 1658, 1669, 1696, 1699, 1737, 1745, 1754, 1794, 1831, 1948, 1950, 2003, 2049, 2050, 2075, 2091, 2154, 2166, 2173, 2186, 2210, 2217, 2219, 2231, 2247, 2299, 2315, 2320, 2334, 2356, 2393, 2527, 2538, 2541, 2582, 2644, 2718, 2761, 2765, 2831, 2855, 2856, 2860, 2954, 2955, 2958, 2969, 2976, 2993, 2996, 3010, 3042, 3100, 3119, 3160, 3174, 3186]. time-aware [1794]. time-based [2718]. time-bounded [2831]. time-constrained [1398]. time-critical [2299]. time-delay [1754]. Time-efficient [2186]. time-frequency [1488]. time-of-arrival [214]. time-of-departure [214]. time-sensitive [1426, 1497, 1658, 1754, 1794, 2049, 2091, 2210, 2334, 2393, 2527, 2582, 2761, 2765, 2955, 2996]. Time-series [1319, 1534, 1948, 2315]. time-space [1239]. time-switching [275]. time-synchronized [1604]. Time-triggered [2154, 2210, 2860]. time-varying [2538, 2993]. Timely [1409]. timeout [2539]. times [127, 540, 1682]. timescale [800, 1874, 2091, 2448]. timestamp [2393]. timestamp-free [2393]. Timing [536, 1165, 1688, 1717, 2430]. TIoT [1652]. Tip [819]. tit [1181]. tit-for-tat-based [1181]. title [3177]. TLS [1385, 2046, 2192, 2548, 2941]. TLS-handshake [2046]. TM [2463, 2479]. TMAGRL [2091]. TMLpSA [2917]. TMLpSA-MEC [2917]. TND} [2894]. TNN [1871]. TNN-IDS [1871]. Tolerance [545, 798, 963, 1906, 3173]. tolerant [7, 73, 187, 552, 825, 1078, 1129, 1268, 1299, 1611, 1785, 2333, 2486, 2488, 2501, 3082, 3116]. tone [2542]. TONTA [671]. tool [149, 154, 271, 425, 466, 923]. toolkit [227]. tools [62, 255, 677, 2117, 2796]. top [723, 966, 1640]. top- [723, 1640]. top-down [966]. topological [1806, 3082]. topologies [329, 597, 1153, 1245, 1347, 1971]. Topology [101, 124, 209, 239, 260, 352, 714, 849, 868, 994, 1074, 1320, 1347, 1463, 1605, 1608, 1797, 1911, 1946, 2022, 2252, 2273, 2382, 2403, 2486, 2528, 2607, 2669, 2797, 2826, 2963, 3019, 3075]. topology-adaptive [2963]. Topology-aware [1463, 3019]. topology-based [2669]. TOPSIS [259]. Tor [1993, 3143, 3148]. TORU [2732]. Torus [2797]. TorVIA [3148]. touch [301, 1121, 1815, 1875, 2057]. Tour [2747]. tourism [2747]. Tower [1289]. TPE [2376]. TPE-BFL [2376]. TPM [287]. trace [801, 1355]. traceability [1838, 1883, 2748, 2822]. traceable [1551, 2443]. Traceback [370, 610, 2463]. Tracer [499]. traceroute [1797]. traces [997]. tracing [157, 607, 619, 1604, 1838, 2114, 2322, 2430]. Tracking [28, 36, 153, 196, 499, 757, 811, 1322, 1560, 1876, 1992, 2007]. TrackSign [1560]. TrackSign-labeled [1560]. tractable [256]. trade [278, 560, 985, 1003]. trade-off [560, 1003]. trade-offs [278, 985]. tradeoff [310, 617, 1308, 2685]. tradeoff-based [1308]. tradeoffs [137]. trades [1208]. trading [602, 1516, 2333, 2876]. Traditional [428, 1372]. Traffic [77, 137, 208, 217, 223, 290, 314, 340, 344, 345, 354, 380, 461, 530, 550, 571, 589, 671, 734, 758, 800, 843, 853, 891, 926, 931, 975, 997, 1006, 1009, 1052, 1117, 1189, 1257, 1279, 1311, 1324, 1369, 1428, 1438, 1504, 1533, 1540, 1554, 1581, 1584, 1596, 1610, 1663, 1672, 1762, 1773, 1792, 1810, 1811, 1902, 1915, 1948, 1953, 1958, 2005, 2045, 2058, 2097, 2118, 2120, 2128, 2169, 2210, 2227, 2263, 2271, 2286, 2329, 2350, 2365, 2402, 2432, 2460, 2479, 2497, 2513, 2524, 2536, 2581, 2599, 2601, 2612, 2642, 2682, 2770, 2772, 2798, 2803, 2883, 2886, 2893, 2905, 2973, 2975, 3002, 3022, 3042, 3064, 3094, 3107, 3143]. traffic [3, 18, 27, 30, 31, 38, 51, 63, 118, 158, 163, 267, 358, 368, 425, 491, 509, 555, 576, 624, 657, 737, 870, 872, 874, 877, 882, 903, 908, 923, 940, 957, 967, 982, 993, 1008, 1033, 1067, 1150, 1217, 1271, 1301, 1493, 1529, 1599, 1737, 1772, 1786, 1893, 1899, 1921, 1928, 1954, 1968, 1995, 2009, 2059, 2137, 2159, 2192, 2218, 2231, 2255, 2280, 2291, 2306, 2315, 2381, 2411, 2425, 2493, 2588, 2618, 2651, 2735, 2852, 2891, 2936, 2942, 2963, 3013, 3027, 3045, 3060, 3108, 3110, 3130, 3139, 3150, 3179]. Traffic-aware [1540, 1610]. traffic-processing [27]. traffics [1921]. trained [2618, 2663]. Training [509, 905, 1467, 1870, 2376, 2553, 2731, 3078, 3149]. trajectories [1001, 1077]. Trajectory [290, 293, 561, 815, 917, 1312, 1337, 1408, 1573, 1627, 1758, 1873, 2201, 2204, 2268, 2489, 2507, 2552, 2563, 2673, 2790, 2829, 2955, 2964, 3092, 3105]. trajectory-resource [1627]. Tram [859]. TranBiLSTM [1793]. transaction [127, 1178]. transactions [590, 639, 909, 2133, 2319]. transceivers [1979]. transcoding [2317]. Transfer [103, 195, 648, 856, 1570, 2570, 2680, 2897]. transferability [2865]. transferable [1428, 2460]. transferring [8]. transfers [829, 1798]. transformation [2281, 2915]. Transformer [1871, 2381, 2391, 2893, 2917, 3022, 3118, 3132]. Transformer-based [2893, 2917, 3118, 3132].

Transformers [2618, 2781, 3037]. transient [426]. transients [1207]. Transition [172, 2352]. transitions [1143]. Translation [3123]. Transmission [108, 331, 342, 504, 713, 770, 850, 966, 1089, 1095, 1213, 1420, 1445, 1446, 1556, 1607, 1608, 1727, 1826, 1881, 1887, 2030, 2084, 2085, 2332, 2419, 2429, 2446, 2456, 2582, 2706, 2754, 2765, 2767, 2913, 2936, 3003, 3071, 3148, 3168]. transmission-aware [504]. transmissions [1352, 1591, 2490]. transmit [1770]. transmitters [2500]. transmitting [1411]. TransMUSE [1428]. Transparent [92, 1775]. transponder [871]. Transport [162, 283, 700, 880, 1695, 1822, 2111, 2271, 2615]. Transportation [345, 380, 946, 1412, 2102]. transports [644]. TransURL [2391]. traversal [162]. Tree [69, 544, 679, 966, 1153, 1217, 1646, 1682, 1702, 1768, 1963, 2052, 2508, 2654, 2999]. Tree-structured [966]. Trend [671, 2272]. Trend-based [671]. Trends [100, 1531, 1715, 1735, 2359, 2428, 2652]. trials [1069]. triangle [708]. Trickle [725]. trigger [2003]. triggered [2154, 2210, 2860]. trilateration [1079]. trilemma [2773]. Triple [1970, 2906]. Triple-step [1970]. troubleshooting [197, 1365, 1433]. Truck [458]. true [288]. TruSD [253]. Trust [246, 253, 254, 325, 434, 444, 582, 809, 837, 861, 930, 944, 1133, 1144, 1164, 1296, 1298, 1336, 1398, 1757, 1854, 1901, 1975, 2095, 2122, 2216, 2297, 2449, 2636, 2806, 2815, 2869, 2877, 2896, 2910, 3000, 3014, 3016, 3140, 3146, 3157]. Trust-Aware [444, 837, 1901, 1975]. Trust-based [2095, 2877]. Trusted [101, 381, 414, 643, 1335, 1722, 1730, 1851, 2064, 2153, 2295, 2310, 2369, 2692]. TrustedBaaS [414]. Trustless [2374, 2670]. trustworthiness [1890, 2311, 3155]. trustworthy [875, 2088, 2832]. truth [425, 3090]. Truthful [1285, 1532, 2327, 2466, 2475, 2726]. TSBFT [1454]. TSCH [747]. TSCL [1239].

TSCRNN [589]. TSN

[1636, 1960, 2375, 2723]. TSN-based [1636]. TSNMRA [1558]. TT&C [2223]. TTL [1984]. TTS [2310]. tune [805, 1539]. tuned [117]. tuning [857, 2781]. tunnel [761, 2612, 3088]. tuple [921]. Turing [2945]. Tutorial [234, 1259]. TVWS [111]. Twin [1284, 1654, 1815, 1989, 2036, 2100, 2330, 2549, 2699, 2738, 3046, 3079, 3132]. twin-assisted [1654]. twin-enabled [1815]. TwinPeaks [191]. twins [775, 2343, 3130]. Twitter [2477]. Twitter/X [2477]. Two [142, 219, 242, 287, 465, 701, 838, 1032, 1181, 1255, 1277, 1303, 1330, 1478, 1778, 1887, 1936, 2023, 2052, 2091, 2281, 2289, 2448, 2465, 2483, 2494, 2529, 2695, 2894, 3022, 3057, 3079]. two-class [1181]. two-dimensional [1255]. two-factor [242, 838, 2494, 3057]. two-layer [2529]. two-level [2023, 2052, 2483]. Two-phase [287, 465, 701, 3079]. Two-stage [1032, 2281, 2289, 2894]. two-step [1303, 2465]. two-tier [219, 1277]. Two-timescale [2091, 2448]. Two-way [142, 1778, 1887]. type [174, 388, 832, 1282, 2871, 3038]. types [114, 1669, 1978].

U [698, 797]. UA [2285]. UAP [526].

UASNs [986, 1757]. UAV [46, 130, 186, 293, 404, 453, 638, 730, 815, 818, 830, 851, 880, 897, 917, 985, 1036, 1039, 1040, 1077, 1143, 1145, 1337, 1340, 1386, 1413, 1472, 1503, 1517, 1526, 1573, 1618, 1627, 1631, 1657, 1663, 1691, 1746, 1758, 1760, 1778, 1796, 1815, 1861, 1873, 1887, 1934, 1956, 1957, 1997, 2036, 2037, 2130, 2142, 2181, 2187, 2205, 2295, 2312, 2314, 2378, 2397, 2408, 2409, 2433, 2459, 2489, 2507, 2510, 2552, 2563, 2578, 2584, 2638, 2667, 2673, 2689, 2709, 2714, 2727, 2771, 2777, 2784, 2790, 2825, 2829, 2862, 2878, 2921, 2936, 2955, 2964, 2979, 3020, 3057, 3092, 3105, 3111, 3131, 3144, 3184]. UAV-aided [1036, 1337, 1778]. UAV-and-RIS-aided [2433]. UAV-assisted [404, 730, 1039, 1040, 1413, 1526, 1760, 1796, 1957, 2037, 2205, 2408, 2489, 2552, 2667, 2727, 2979, 3020, 3144]. UAV-based [1077, 1934]. UAV-BS [2771]. UAV-CFmMIMOs [1631]. UAV-enabled [1145, 1472, 1627, 1758, 2187, 2507, 2878, 3111]. UAV-ground [2584]. UAV-IOT [917, 2936]. UAV-mounted [2510]. UAV-RIS-assisted [2921]. UAV-to-ground [1861]. UAV-to-UAV [1618]. UAV-UGV-assisted [2709]. UAVs [223, 367, 377, 432, 460, 715, 736, 775, 831, 881, 979, 1264, 1341, 1509, 1760, 2201, 2545, 2636, 3106]. UAVs-assisted [460, 979, 2201]. ubiquitous [1530]. UDNs [888]. UDP [1826]. UDP-based [1826]. UDP-RT [1826]. UE [562]. UE-VBS [562]. UGL [2784]. UGV [2709]. UINT [2627]. Ukraine [1487]. Ultra [204, 288, 322, 466, 562, 579, 730, 769, 893, 1300, 1332, 1400, 1739, 1823, 1896, 2262, 2944, 3040, 3104]. ultra-broadband [288, 2944]. Ultra-Dense [204, 562, 579, 730, 893, 1332, 1739, 2262]. ultra-lightweight [1300]. ultra-low [466, 1823, 3104]. ultra-reliable [769, 1896]. ultrabroadband [653]. ultralightweight [2367]. UltraLite [2263]. ultrasonic [121]. ultrasound [600]. unattended [1437, 1764]. unauthenticated [2483]. unauthorized [2824]. Unavailability [1935, 2609]. Unavailability-aware [1935]. Unbalanced [1257]. Unbiased [890]. Unboxing [1890]. uncertain [1510, 1669, 2113, 2280, 2579, 2808, 3163]. uncertainty [773, 810, 1112, 2458, 2984, 3004, 3154]. unclonable [412, 1641]. Uncoded [1012]. Uncoordinated [145]. underground [1689]. underlaid [1280]. underlay [2, 1281, 2845]. underlaying [265, 2345]. Understanding [85, 234, 248, 555, 803, 925, 1005, 2218, 2311]. Understood [1780]. Underwater [83, 120, 159, 759, 1134, 1272, 1391, 1463, 1538, 1675, 1759, 1969, 1992, 2006, 2015, 2034, 2072, 2131, 2167, 2207, 2284, 2320, 2321, 2332, 2368, 2456, 2602, 2619, 2783, 3012, 3072]. unified [526, 808, 1642, 1696]. Unifying [2997, 3095]. unit [1557]. Units [1518]. Universal [309, 1606, 1948, 2460, 2963]. Universally [179]. university [2075]. unknown [273, 566, 1657, 1915, 2269, 2803, 2873, 2893]. unknowns [2329]. Unlicensed [224, 232, 1045, 2000, 2480]. unlinkability [326, 2615]. Unlocking [1924, 2933]. Unmanned [322, 496, 512, 626, 714, 973, 1097, 1108, 1503, 2545, 2928]. Unobtrusive [1]. Unraveled [1567]. unsaturated [786]. unscented [2393]. unseen [1112, 2893, 2984]. unstable [292]. unstructured [197]. Unsupervised [312, 616, 1756, 2134, 2297, 2594, 2772]. untrusted [1887]. Unveiling [684]. unwittingly [914]. up-offloaded [1579]. updatable [1656]. update [1061, 1074, 1329, 1430, 1610, 2171, 2257, 2358, 2424]. Updates [238, 263, 915, 1276, 1399, 3063]. updating [1915, 2901]. UPF [1214, 1449]. upgraded [1556]. Uplink [373, 1142, 1211, 1384, 1445, 1479, 1627, 1907, 2092, 2130, 2706, 2875, 2913, 2955, 2967]. uplink/downlink [1479]. uploading [9, 1655]. upsampling [2630]. Urban [364, 711, 817, 1568, 1719, 1930, 2147, 2597, 2641, 2678, 3139]. URL [247, 1339, 2134, 2391, 2683, 2900, 3078, 3188]. URL-based [2900]. URLLC [174, 232, 255, 267, 818, 1445, 1961, 2230, 2562, 2629, 2756]. URLs [904]. Usability [162, 2227]. usage [502, 555, 1460, 2442, 2785]. usages [1256]. Use [131, 234, 578, 859, 1378, 1523]. used [1860, 2097, 2112]. User [5, 100, 102, 131, 133, 198, 233, 257, 258, 328, 354, 365, 374, 418, 419, 443, 553, 563, 628, 686, 735, 744, 755, 836, 1131, 1229, 1271, 1295, 1340, 1352, 1432, 1439, 1518, 1545, 1609, 1619, 1663, 1670, 1700, 1722, 1738, 1746, 1809, 1833, 1846, 1851, 1884, 1945, 1951, 2004, 2021, 2035, 2075, 2128, 2145, 2172, 2175, 2296, 2335, 2369, 2372, 2392, 2409, 2444, 2489, 2574, 2615, 2687, 2833, 2915, 3029, 3030, 3041, 3181]. user-aware [2172]. user-cell [2128]. User-Centric [102, 2335, 2833]. user-level [1619]. user-perceived [2687]. user-space [1963]. user-to-multiple [744]. users [107, 266, 308, 502, 583, 701, 834, 1049, 1133, 1206, 1562, 2037, 2114, 2710, 2915, 2930]. uses [488]. Using [8, 30, 32, 61, 116, 131, 135, 151, 176, 239, 264, 268, 274, 282, 291, 315, 354, 392, 436, 531, 538, 544, 562, 610, 630, 768, 778, 780, 816, 831, 857, 965, 1003, 1025, 1037, 1046, 1049, 1068, 1069, 1071, 1113, 1119, 1140, 1165, 1228, 1247, 1273, 1281, 1284, 1353, 1382, 1401, 1409, 1448, 1531, 1617, 1622, 1645, 1646, 1660, 1743, 1780, 1816, 1820, 1923, 2022, 2044, 2085, 2176, 2185, 2256, 2261, 2374, 2430, 2447, 2505, 2523, 2541, 2569, 2576, 2609, 2623, 2707, 2773, 2786, 2792, 2803, 2827, 2884, 2923, 2961, 2978, 3097, 3109, 3118, 3143, 3157]. using [34, 57, 117, 159, 212, 261, 302, 321, 385, 390, 419, 462, 481, 497, 505, 516, 529, 566, 587, 640, 645, 721, 743, 801, 809, 811, 835, 870, 896, 904, 967, 973, 1031, 1061, 1088, 1155, 1159, 1186, 1230, 1231, 1289, 1297, 1299, 1301, 1303, 1317, 1338, 1386, 1396, 1408, 1453, 1473, 1478, 1503, 1542, 1557, 1558, 1572, 1577, 1594, 1595, 1638, 1729, 1737, 1753, 1779, 1849, 1859, 1899, 1900, 2025, 2030, 2089, 2109, 2134, 2179, 2188, 2190, 2320, 2358, 2455, 2459, 2464, 2472, 2474, 2570, 2605, 2612, 2672, 2789, 2813, 2872, 2922, 2924, 2943, 3004, 3053, 3085, 3162, 3172]. using [1, 1149, 1150, 1493, 2087, 2111, 2234, 2245, 2425, 2716, 2774, 2988]. Utility [38, 133, 583, 1049, 1193, 1208, 1263, 1821, 2266, 2538, 2655]. utility-maximizing [583]. Utilization [138, 339, 1518, 2255, 2908]. Utilization-aware [339]. Utilized [254]. utilizing [718, 1989, 2798, 3069, 3164]. UTL DGA22 [1421]. UTXO [454, 2423]. UTXO-based [454]. UVLC [1675, 1759]. UW [1391]. UW-GRE [1391]. UWB [902]. UWSNs [2348]. v1 [716]. V2G [1460]. V2I [922]. V2V [111, 3166]. V2X [81, 1015, 1291, 1424, 1779, 1903, 2218, 2372, 2440, 2561]. vague [826]. vagus [116]. Validation [165, 962, 1695, 1801, 1847, 1986, 2400, 2583, 2904]. validations [1724]. validator [953]. VALNET [962]. Value [826, 1155, 2369, 2636, 2817, 3028, 3122, 3158]. value-added [1155]. value-driven [3028]. VANET [204]. VANETs [84, 371, 1157, 1185, 1468, 1499, 1551, 1790, 1812, 2085, 2143, 2298, 2530, 2545, 2575, 2810, 2909, 3006, 3058, 3138]. Variable [871, 1515, 1653, 2385]. variant [2781]. variation [150, 2190]. Variational [816, 2404]. VariBan [1653]. varied [2316]. various [1245, 2061, 2735]. varying [2538, 2993]. VB [339]. VBS [562]. vClos [2949]. VCMaker [865]. VEC [1466, 2635, 3081, 3154]. VEC-Sim [2635]. VECLLF [2909]. vector [31, 1303, 2854]. vector-coded [31]. vector-matrix [2854]. Vehicle [212, 457, 778, 821, 859, 1097, 1123, 1282, 1412, 1503, 1579, 1790, 1903, 2204, 2298, 2308, 2354, 2438, 2475, 2689, 2716, 2732, 2755, 2909, 2956, 3115]. Vehicle-Based [1123]. Vehicle-cluster-based [1412]. vehicle-edge [2308, 2909]. vehicle-related [2956]. vehicle-to-everything [212, 3115]. vehicle-to-infrastructure [457]. vehicle-to-vehicle [1790]. Vehicles [53, 263, 313, 322, 496, 509, 512, 626, 714, 831, 930, 973, 1108, 1212, 1232, 1329, 1470, 1568, 2055, 2080, 2216, 2366, 2526, 2536, 2544, 2545, 2570, 2640, 2744, 2776, 2873, 2928, 3070, 3140, 3175]. VEhicular [122, 140, 141, 379, 380, 383, 393, 426, 520, 542, 617, 655, 742, 762, 781, 839, 849, 946, 1049, 1094, 1151, 1246, 1287, 1371, 1530, 1632, 1731, 1771, 1781, 1789, 1836, 1858, 1874, 1895, 1913, 2012, 2039, 2040, 2046, 2055, 2132, 2141, 2181, 2186, 2336, 2349, 2358, 2365, 2383, 2445, 2474, 2540, 2544, 2545, 2614, 2632, 2635, 2653, 2685, 2689, 2693, 2746, 2755, 2769, 2799, 2836, 2841, 2868, 2953, 2985, 2999, 3053, 3065, 3079, 3089, 3104, 3118, 3135, 3165]. vendor [489, 2508, 2569, 2631]. ventilation [437]. verifiability [2750]. Verifiable [179, 861, 1166, 1645, 1878, 1900, 2189, 2394, 2733]. Verification [50, 101, 178, 238, 452, 964, 1611, 1779, 2636, 3178]. verifier [1252]. VeriTable [50]. versatile [653]. version [1780]. versus [1376, 2368]. vertex [679]. vertical [1331, 2506, 2834]. very [18, 64]. vessel [565, 1992, 2229]. VFL [2834, 3126]. via [63, 140, 174, 735, 897, 960, 1175, 1184, 1212, 1232, 1288, 1320, 1322, 1502, 1532, 1599, 1688, 1701, 1706, 1712, 1881, 1887, 1902, 1988, 2023, 2034, 2040, 2045, 2163, 2286, 2405, 2420, 2452, 2459, 2532, 2537, 2573, 2618, 2663, 2680, 2761, 2762, 2876, 2889, 2899, 2908, 2987, 2994, 3005, 3036, 3054, 3099, 3103, 3161]. vibration [686]. vibration-based [686]. Video [99, 204, 248, 364, 455, 573, 598, 603, 865, 1009, 1327, 1332, 1360, 1616, 1727, 1745, 1782, 1844, 1886, 1945, 2005, 2094, 2317, 2323, 2332, 2416, 2498, 2630, 2644, 2700, 2740, 3018, 3029, 3148, 3177]. Video-on-Demand [364, 2005]. videogame [634]. VideoTrain [1009]. View [574, 855, 1175, 1886, 2411, 2418, 2492, 2728, 2740, 3007, 3099]. views [942]. violation [446, 1184]. violations [982]. VIOS [1980]. Virtual [6, 56, 231, 280, 303, 330, 460, 562, 616, 699, 727, 739, 931, 989, 1063, 1187, 1226, 1242, 1309, 1352, 1375, 1393, 1484, 1507, 1541, 1603, 1665, 1734, 1908, 1933, 1935, 1983, 2074, 2150, 2152, 2170, 2236, 2260, 2569, 2707, 2712, 2715, 2863, 2908, 3094]. virtual/physical [2707]. Virtualization [21, 584, 1160, 1528, 2200]. Virtualized [391, 424, 662, 668, 691, 692, 899, 1082, 1266, 1816, 1869, 1885, 2087, 2173, 2318, 2555]. Visible [1675, 1759, 2039]. Vision [57, 2653, 3022, 3186]. visit [85]. visualization [906]. VLC [1903, 2004, 2372, 3021]. VLC-NOMA [2004]. VLC-RF [3021]. VLC/RF [1903, 2372]. VM [892, 899]. VNE [1818]. VNF [59, 90, 240, 329, 333, 472, 530, 844, 950, 1047, 1285, 1713, 1817, 2063, 2158, 2335, 2715, 2790, 2864, 3085]. VNF-AAPC [240]. VNFs [7]. voice [3109]. void [759, 1569]. VoIP [751]. Volatility [2199]. Volatility-based [2199]. Volume [357, 527, 1571, 1764, 1893]. vote [826]. Voting [179, 305, 2708, 2831]. voting-based [2708]. VPLS [727]. VPN [1810, 2652]. VPNs [1389]. VR [1727, 2681, 2684]. VR-SGs [2681]. vs [515, 630, 985, 1857]. VSI [1407]. VSI-networks [1407]. vSwitch [534]. vulnerabilities [1425, 2378]. vulnerability [994, 1017, 1026, 1269, 1590, 1852, 2011, 3152]. Vulnerable [131, 2145].

WAIC [2978]. wait [552]. waiting [1350, 1682]. wake [322, 3101]. wake-up [322, 3101]. Walking [1936]. wallet [2801]. WAN [499, 1194, 2071, 2244, 2582]. WANs [623, 1403, 1994]. WAOA [2509]. WAPS [1826]. war [1487]. wardens [2249]. warning [319, 379]. WatchEDGE [2043]. Water [113, 202, 577, 2264]. watermarking [1351]. Wave [14, 428, 786, 798, 963, 985, 1073, 1749, 2008]. waveform [2543]. wavelength [2277]. way [142, 1778, 1887, 2242]. WBAN [863, 1489]. WBAN-IoT [863]. WBANs [524, 1629, 3182]. weak [69]. weakly [1348]. Wearable [646, 2377, 2988]. weather [920, 2044, 3108]. web [206, 555, 684, 788, 1059, 1127, 1389, 1438, 1560, 2138, 2416, 2523, 2858, 2962]. Web3 [1577, 2801]. WebGPU [3054]. WebGPU-based [3054]. webpage [1236, 3078, 3188]. WebRTC [2103]. WebRTC-QoE [2103]. webshell [2138]. website [538, 779, 874, 1236, 1377, 1628, 1662, 1993, 2839]. websites [85, 2900]. Weight [845, 987, 1448, 2533, 2654]. weighted [544, 718, 2667]. welfare [1274]. WGAN [1257]. whale [743, 2509]. whale-ant [2509]. Which [297, 1163]. White [1704]. Who [1574, 1821]. wholly [1803]. Wi [94, 698, 797, 998, 1446, 1617, 1940, 2000, 2274, 2385, 2694, 2824, 3008, 3147]. Wi-Fi [94, 698, 797, 998, 1446, 1617, 1940, 2000, 2274, 2385, 2694, 2824, 3008, 3147]. Wi-Gait [1617]. Wide [307, 943, 1310, 1639, 1651, 1909, 2123, 2202, 2337, 3114]. wide-band [1310]. width [1957]. Wifi [38, 173, 281, 321, 359, 394, 506, 773, 783, 922, 1328, 1425, 1448, 2109, 2229, 2309, 2672]. Wifi-based [321]. wild [675, 922, 2007]. wildcard [347]. will [1597]. Window [294, 428, 1080, 1563, 1729, 2278, 2385]. winged [2737]. wire [2905]. wired [597]. wired/wireless [597]. Wireless [10, 33, 41, 43, 103, 104, 124, 142, 198, 228, 239, 242, 262, 289, 327, 369, 455, 508, 540, 543, 544, 551, 573, 625, 678, 705, 711, 713, 725, 750, 757, 820, 837, 868, 897, 901, 927, 948, 1020, 1079, 1100, 1105, 1129, 1161, 1164, 1190, 1212, 1249, 1270, 1299, 1387, 1442, 1508, 1526, 1546, 1561, 1627, 1664, 1684, 1705, 1752, 1777, 1813, 1827, 1830, 1863, 1872, 1920, 1992, 2006, 2124, 2149, 2187, 2213, 2292, 2331, 2343, 2377, 2388, 2407, 2429, 2439, 2446, 2481, 2484, 2507, 2509, 2510, 2649, 2673, 2694, 2717, 2720, 2808, 2816, 2824, 2854, 2870, 2938, 2970, 2989, 3071, 3077, 3113, 3129]. wireless [12, 28, 34, 37, 54, 174, 180, 195, 202, 252, 288, 298, 385, 474, 597, 632, 679, 743, 1031, 1065, 1102, 1103, 1187, 1211, 1231, 1344, 1354, 1407, 1414, 1463, 1485, 1556, 1566, 1578, 1681, 1689, 1699, 1733, 1806, 1809, 1831, 1832, 1875, 1989, 2068, 2230, 2254, 2314, 2321, 2389, 2476, 2543, 2549, 2602, 2604, 2684, 2699, 2727, 2757, 2759, 2761, 2783, 2795, 2863, 2887, 3034, 3069, 3132]. wireless-powered [12, 1485, 2149]. wise [1350, 2554, 2768, 2851]. within [715, 899, 2794, 2971]. without [1682, 1733, 2049, 2373, 2375, 2407, 2826]. WLAN [413, 732, 763, 774, 1083]. WLANs [65, 302, 712, 1125, 1260, 1353, 1518]. WLANs-based [1518]. WNQA [1827]. wolf [37]. work [481, 631, 1172, 1288, 1293, 2453, 3055]. worker [119, 409, 1290, 3090]. worker-selection [119]. workers [535]. workflow [197]. workflows [274, 335, 889, 938, 2837]. Workforce [1433]. workgroup [488]. Workload [1534, 1669, 1710, 1785, 1885, 2971]. workload-dependent [1669]. workloads [169]. world [675, 1993]. worldwide [886]. worm [3171]. WoT [2889]. WoTemu [1059]. WPCN [2725]. WPCNs [2, 195]. WPT [2829, 2950, 3181]. WPT-MEC [2829, 3181]. wrapper [144]. WSN [468, 536, 726, 796, 845, 1074, 1420, 2638]. WSNs [113, 278, 473, 842, 1182, 1558, 1813, 2501]. WUWNet [3012].

X [1638, 2477]. X.509 [1144].

X.509-compatible [1144]. XAInomaly [2770]. XG [2106]. XG-PON [2106]. XGBoost [2601, 2791]. XR [2402]. XSS [11, 2303, 3152].

YANG [1452]. Yardstick [2805].

Yardstick-Stackelberg [2805]. YB [1903]. Yes [2251]. Yinker [1446]. YouTube [2933].

Zeek [1163]. Zero [301, 312, 944, 1121, 1133, 1298, 1596, 1815, 1834, 1875, 1975, 2057, 2122, 2203, 2259, 2297, 2960, 3014, 3016, 3084, 3157]. zero-day [312, 2203, 2960]. zero-knowledge [1834]. Zero-relabelling [1596]. Zero-touch [301, 1121, 1815, 1875, 2057]. zero-trust [944, 1133, 3157]. ZET-ROS [3157]. zk [2030]. zk-SNARKs [2030]. zkRPC [2670]. zone [393, 1544]. zones [1821]. Zoom [2580]. Zoom-inRCL [2580]. zooming [1400]. ZTA [3014].

References

Zhang:2020:URH

[1] Qian Zhang, Run Zhao, Dong Li, and Dong Wang. Unobtrusive and robust human identification using COTS RFID. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861831079X.

Yuan:2020:OTF

[2] Lina Yuan, Suzhi Bi, Xiaohui Lin, and Hui Wang. Optimizing throughput fairness of cluster-based cooperation in underlay cognitive WPCNs. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311514.

Saddoud:2020:RRM

[3] Ahlem Saddoud, Wael Doghri, Emna Charfi, and Lamia Chaari Fourati. 5G radio resource management approach for multi-traffic IoT communications. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618303876.

Noferesti:2020:AAS

[4] Morteza Noferesti and Rasool Jalili. ACoPE: an adaptive semi-supervised learning approach for complex-policy enforcement in high-bandwidth networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304074.

Kazi:2020:CMC

[5] Baha Uddin Kazi and Gabriel Wainer. Coordinated multi-cell cooperation with user centric dynamic coordination station. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304335.

He:2020:DEE

[6] Mengyang He, Lei Zhuang, Shuaikui Tian, Guoqing Wang, and Kunli Zhang. DROI: Energy-efficient virtual network embedding algorithm based on dynamic regions of interest. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304438.

Yuan:2020:FTP

[7] Guochang Yuan, Zichuan Xu, Binxu Yang, Weifa Liang, Wei Koong Chai, Daphné Tuncer, Alex Galis, George Pavlou, and Guowei Wu. Fault tolerant placement of stateful VNFs and dynamic fault recovery in cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300970.

Alhowaidi:2020:ESN

[8] Mohammad Alhowaidi, Byrav Ramamurthy, Brian Bockelman, and David Swanson. Enhancing the SDTMA-NDN architecture for transferring the scientific data software using named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302087.

Liu:2020:SAD

[9] Xiaolan Liu, Bin Yang, Xiaohong Jiang, Lisheng Ma, and Shikai Shen. On social-aware data uploading study of D2D-enabled cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307145.

Huan:2020:PIP

[10] Xintao Huan and Kyeong Soo Kim. On the practical implementation of propagation delay and clock skew compensated high-precision time synchronization schemes with resourceconstrained sensor nodes in multi-hop wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619300398.

Rodriguez:2020:CSS

[11] Germán E. Rodríguez, Jenny G. Torres, Pamela Flores, and Diego E. Benavides. Cross-site scripting (XSS) attacks and mitigation: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): ??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311247.

Tang:2020:BAC

[12] Kun Tang and Shaowei Liao. Bufferaided cooperative spectrum sharing with full-duplex wireless-powered relay. Computer Networks (Amsterdam, Netherlands: 1999), 166(??): ??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306966.

Li:2020:EMB

[13] Kuanyang Li, Lilan Tu, and Lang Chai. Ensemble-model-based link prediction of complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308710.

Ma:2020:QOJ

[14] Zhongyu Ma, Bo Li, Zhongjiang Yan, and Mao Yang. QoS-Oriented joint optimization of resource allocation and concurrent scheduling in 5G millimeter-wave network. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306668.

Subramanya:2020:MLD

[15] Tejas Subramanya, Davit Harutyunyan, and Roberto Riggio. Machine learning-driven service function chain placement and scaling in MEC-enabled 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310254.

Khanouche:2020:FQA

[16] Mohamed Essaid Khanouche, Hania Gadouche, Zoubeyr Farah, and Abdelkamel Tari. Flexible QoS-aware services composition for service computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307169.

Sharifi:2020:CAE

[17] Seyed Ali Sharifi and Seyed Morteza Babamir. The clustering algorithm for efficient energy management in mobile ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302865.

Roquero:2020:PEC

[18] P. Roquero, E. Magaña, R. Leira, and J. Aracil. Performance evaluation of client-based traffic sniffing for very large populations. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302336.

Barmpounakis:2020:NSE

[19] Sokratis Barmpounakis, Nikolaos Maroulis, Michael Papadakis, George Tsiatsios, Dimitrios Soukaras, and Nancy Alonistioti. Network slicing — enabled RAN management for 5G: Cross layer control based on SDN and SDR. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309776.

deMatos:2020:CIS

[20] Everton de Matos, Ramão Tiago Tiburski, Carlos Roberto Moratelli, Sergio Johann Filho, Leonardo Albernaz Amaral, Gowri Ramachandran, Bhaskar Krishnamachari, and Fabiano Hessel. Context information sharing for the Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 166 (??):??, January 15, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310400.

Pham:2020:MAR

[21] Tuan-Minh Pham, Serge Fdida, Thi-Thuy-Lien Nguyen, and Hoai-Nam Chu. Modeling and analysis of robust service composition for network functions virtualization. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305080.

Mao:2020:JHS

[22] Zhifei Mao, Yuming Jiang, Xiaoqiang Di, and Yordanos Woldeyohannes. Joint head selection and airtime allocation for data dissemination in mobile social networks. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303780.

Hu:2020:DPC

[23] Xiaoyan Hu, Shaoqi Zheng, Guoqiang Zhang, Lixia Zhao, Guang Cheng, Jian Gong, and Ruidong Li. An on-demand off-path cache exploration based multipath forwarding strategy. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619304979.

Melodia:2020:ED

[24] Tommaso Melodia and Antonio Iera. Editorial — December 2019. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619316615.

Anonymous:2020:EBa

[25] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619316652.

Anonymous:2020:Ja

[26] Anonymous. 15 January 2020. Computer Networks (Amsterdam, Netherlands: 1999), 166(??):??, January 15, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yu:2020:WQM

[27] Li Yu, Zongpeng Li, Yucun Zhong, Zhenzhou Ji, and Jiangchuan Liu. When QoE meets learning: a distributed traffic-processing framework for elastic resource provisioning in HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618312829.

Wang:2020:TLT

[28] Tao Wang, Xiang Wang, Wei Shi, Zongmin Zhao, Zhenxue He, and Tongsheng Xia. Target localization and tracking based on improved Bayesian enhanced least-squares algorithm in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306406.

Pan:2020:PML

[29] Lujia Pan, Jianfeng Zhang, Patrick P. C. Lee, Marcus Kalander, Junjian Ye, and Pinghui Wang. Proactive microwave link anomaly detection in cellular data networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307662.

Dominicini:2020:KTS

[30] Cristina K. Dominicini, Gilmar L. Vassoler, Rodolfo Valentim, Rodolfo S. Villaca, Moisés R. N. Ribeiro, Magnos Martinello, and Eduardo Zambon. KeySFC: Traffic steering using strict source routing for dynamic and efficient network orchestration. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930194X.

Soylu:2020:BVC

[31] Tuncay Soylu, Oguzhan Erdem, and Aydin Carus. Bit vector-coded simple CART structure for low latency traffic classification on FPGAs. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619304323.

Barakabitze:2020:NSU

[32] Alcardo Alex Barakabitze, Arslan Ahmad, Rashid Mijumbi, and Andrew Hines. 5G network slicing using SDN and NFV: a survey of taxonomy, architectures and future challenges. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619304773.

Lattanzi:2020:EEI

[33] Emanuele Lattanzi, Paolo Capellacci, and Valerio Freschi. Experimental evaluation of the impact of packet length on wireless sensor networks subject to interference. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619303755.

Raschella:2020:DAP

[34] Alessandro Raschellà, Faycal Bouhafs, Michael Mackay, Qi Shi, Jorge Ortín, José Ramón Gállego, and Maria Canales. A dynamic access point allocation algorithm for dense wireless LANs using potential game. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302439.

Jaglarz:2020:ELD

[35] Piotr Jaglarz, Piotr Boryło, Andrzej Szymański, and Piotr Chołda. Enhanced Lagrange Decomposition for multi-objective scalable TE in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618308260.

Cozza:2020:HLD

[36] Federico Cozza, Alfonso Guarino, Francesco Isernia, Delfina Malandrino, Antonio Rapuano, Raffaele Schiavone, and Rocco Zaccagnino. Hybrid and lightweight detection of third party tracking: Design, implementation, and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311685.

Xiu-wu:2020:CRA

[37] Y. U. Xiu-wu, Y. U. Hao, Liu Yong, and Xiao Ren-rong. A clustering routing algorithm based on wolf pack algorithm for heterogeneous wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308679.

Swain:2020:NEA

[38] Siba Narayan Swain and C. Siva Ram Murthy. A novel energy-aware utility maximization for efficient deviceto-device communication in LTE-WiFi networks under mixed traffic scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308035.

El-mekkawi:2020:SKM

[39] Ahmed El-mekkawi, Xavier Hesselbach, and Jose Ramon Piney. Squatting and kicking model evaluation for prioritized sliced resource management. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305766.

Yang:2020:FMS

[40] Ze Yang and Kwan L. Yeung. Flow monitoring scheme design in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618306431.

Liu:2020:RDC

[41] Ren-Shiou Liu and Yen-Chen Chen. Robust data collection for energyharvesting wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128618310594.

Alam:2020:MNF

[42] Shahid Alam, Soltan Abed Alharbi, and Serdar Yildirim. Mining nested flow of dominant APIs for detecting Android malware. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309855.

Liu:2020:SRA

[43] Zhixin Liu, Mingye Zhao, Yazhou Yuan, and Xinping Guan. Subchannel and resource allocation in cognitive radio sensor network with wireless energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309934.

Al-Musawi:2020:IOL

[44] Bahaa Al-Musawi, Philip Branch, Mohammed Falih Hassan, and Shiva Raj Pokhrel. Identifying OSPF LSA falsification attacks through non-linear analysis. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310333.

Singh:2020:FEI

[45] Deepak Singh, Bryan Ng, Yuan-Cheng Lai, Ying-Dar Lin, and Winston K. G. Seah. Full encapsulation or internal buffering in OpenFlow based hardware switches? Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308424.

Mukherjee:2020:DAP

[46] Anandarup Mukherjee, Sudip Misra, Anumandala Sukrutha, and Narendra Singh Raghuwanshi. Distributed aerial processing for IoTbased edge UAV swarms in smart farming. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305328.

Angelopoulos:2020:KDE

[47] Constantinos Marios Angelopoulos, Gabriel Filios, Sotiris Nikoletseas, and Theofanis P. Raptis. Keeping data at the edge of smart irrigation networks: a case study in strawberry greenhouses. Computer Networks (Amsterdam, Netherlands: 1999), 167 (??):??, February 11, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305195.

Anonymous:2020:EBb

[48] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128619317529.

Anonymous:2020:Fa

[49] Anonymous. 11 February 2020. Computer Networks (Amsterdam, Netherlands: 1999), 167(??):??, February 11, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2020:VFE

[50] Yaoqing Liu, Garegin Grigoryan, Jun Li, Guchuan Sun, and Tony Tauber. VeriTable: Fast equivalence verification of multiple large forwarding tables. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314105.

Sun:2020:ERE

[51] Xiuwen Sun, Hao Li, Dan Zhao, Xingxing Lu, Zheng Peng, and Chengchen Hu. Efficient regular expression matching over compressed traffic. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618311939.

Wang:2020:MMC

[52] Wenjie Wang, Donghai Tian, Weizhi Meng, Xiaoqi Jia, Runze Zhao, and Rui Ma. MSYM: a multichannel communication system for Android devices. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309983.

Gohar:2020:CEM

[53] Ali Gohar and Sanghwan Lee. A cost efficient multi remote driver selection for remote operated vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307728.

Zhang:2020:LBJ

[54] Ping Zhang, Jianxin Wang, and Wenjun Li. A learning based joint compressive sensing for wireless sensing networks. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310217.

Konstantinidis:2020:MCB

[55] Dimitrios Konstantinidis, Vasileios Argyriou, Tania Stathaki, and Nikolaos Grammalidis. A modular CNNbased building detector for remote sensing images. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304475.

Perakis:2020:CFP

[56] Konstantinos Perakis, Fenareti Lampathaki, Konstantinos Nikas, Yiannis Georgiou, Oskar Marko, and Jarissa Maselyne. CYBELE — Fostering precision agriculture and livestock farming through secure access to large-scale HPC enabled virtual industrial experimentation environments fostering scalable big data analytics. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305353.

Wan:2020:FRC

[57] Shaohua Wan and Sotirios Goudos. Faster R-CNN for multi-class fruit detection using a robotic vision system. Computer Networks (Amsterdam, Netherlands: 1999), 168(??): ??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306978.

Glaroudis:2020:SCR

[58] Dimitrios Glaroudis, Athanasios Iossifides, and Periklis Chatzimisios. Survey, comparison and research challenges of IoT application protocols for smart farming. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306942.

Yao:2020:FAV

[59] Yifu Yao, Songtao Guo, Pan Li, Guiyan Liu, and Yue Zeng. Forecasting assisted VNF scaling in NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619305705.

Heimovski:2020:FAH

[60] Gustavo B. Heimovski, Rogério C. Turchetti, Juliano A. Wickboldt, Lisandro Z. Granville, and Elias P. Duarte, Jr. FT-Aurora: a highly available IaaS cloud manager based on replication. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861831346X.

Elmasry:2020:EDL

[61] Wisam Elmasry, Akhan Akbulut, and Abdul Halim Zaim. Evolving deep learning architectures for network intrusion detection using a double PSO metaheuristic. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930800X.

Sarigiannidis:2020:BDE

[62] Panagiotis Sarigiannidis, Thomas Lagkas, Konstantinos Rantos, and Paolo Bellavista. The Big Data era in IoT-enabled smart farming: Redefining systems, tools, and techniques. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316251.

Sciancalepore:2020:PEE

[63] Savio Sciancalepore, Omar Adel Ibrahim, Gabriele Oligeri, and Roberto Di Pietro. PiNcH: an effective, efficient, and robust solution to drone detection via network traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311764.

Castagno:2020:LSB

[64] Paolo Castagno, Vincenzo Mancuso, Matteo Sereno, and Marco Ajmone Marsan. Limitations and sidelinkbased extensions of 3GPP cellular access protocols for very crowded environments. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305377.

Ali:2020:CLA

[65] Asad Ali and Faisal Ahmed Khan. Condition and location-aware channel switching scheme for multi-hop multiband WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306863.

Vuletic:2020:LNS

[66] Pavle Vuletić, Bartosz Bosak, Marinos Dimolianis, Pascal Mérindol, David Schmitz, and Henrik Wessing. Localization of network service performance degradation in multi-tenant networks. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619303548.

Jie:2020:CBO

[67] Yu Jie, Ahmad Alsharoa, Ahmed E. Kamal, and Mohammad Alnuem. A CoMP-Based outage compensation solution for heterogeneous femtocell networks. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930060X.

Cerda-Alabern:2020:GNC

[68] Llorenç Cerdà-Alabern, Roger Baig, and Leandro Navarro. On the Guifi.net community network economics. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309077.

Madhja:2020:EAT

[69] Adelina Madhja, Sotiris Nikoletseas, and Alexandros A. Voudouris. Energy-aware tree network formation among computationally weak nodes. Computer Networks (Amsterdam, Netherlands: 1999), 168 (??):??, February 26, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309533.

Anonymous:2020:EBc

[70] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620300505.

Anonymous:2020:Fb

[71] Anonymous. 26 February 2020. Computer Networks (Amsterdam, Netherlands: 1999), 168(??):??, February 26, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sakib:2020:MAM

[72] Muhammad N. Sakib, Chin-Tser Huang, and Ying-Dar Lin. Maximizing accuracy in multi-scanner malware detection systems. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619302518.

Zhou:2020:OAS

[73] Chongyu Zhou, Chen-Khong Tham, and Mehul Motani. Online auction for scheduling concurrent delay tolerant tasks in crowdsourcing systems. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930163X.

Zhong:2020:HNN

[74] Ying Zhong, Wenqi Chen, Zhiliang Wang, Yifan Chen, Kai Wang, Yahui Li, Xia Yin, Xingang Shi, Jiahai Yang, and Keqin Li. HELAD: a novel network anomaly detection model based on heterogeneous ensemble learning. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304086.

Abduljabbar:2020:IMO

[75] Dhuha Abdulhadi Abduljabbar, Siti Zaiton Mohd Hashim, and Roselina Sallehuddin. An improved multiobjective evolutionary algorithm for detecting communities in complex networks with graphlet measure. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912861931182X.

Swain:2020:NCA

[76] Siba Narayan Swain and C. Siva Ram Murthy. A novel collision aware network assisted device discovery scheme empowering massive D2D communications in 3GPP LTE-A networks. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302130.

Ghazvini:2020:LBM

[77] Seyedeh Mina Hosseini Ghazvini, Akbar Ghaffarpour Rahbar, and Behrooz Alizadeh. Load balancing, multipath routing and adaptive modulation with traffic grooming in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309223.

Akkari:2020:MMS

[78] Nadine Akkari and Nikos Dimitriou. Mobility management solutions for 5G networks: Architecture and services. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306346. See

corrigendum [357].

Kafiloglu:2020:MMS

[79] S. Sinem Kafiloglu, Gürkan Gür, and Fatih Alagöz. A Markovian model for satellite integrated cognitive and D2D HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309235.

Xu:2020:ESA

[80] Jianfeng Xu, Liming Wang, and Zhen Xu. An enhanced saturation attack and its mitigation mechanism in software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619307625.

Ghosal:2020:SIC

[81] Amrita Ghosal and Mauro Conti. Security issues and challenges in V2X: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305857.

Gunduz:2020:CSS

[82] Muhammed Zekeriya Gunduz and Resul Das. Cyber-security on smart grid: Threats and potential solutions. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311235.

Liu:2020:TOM

[83] Linfeng Liu, Ran Wang, Gaoxi Xiao, and Dongyue Guo. On the throughput optimization for message dissemination in opportunistic underwater sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618305711.

BenJaballah:2020:SDR

[84] Wafa Ben Jaballah, Mauro Conti, and Chhagan Lal. Security and design requirements for software-defined VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619306553.

Zhuang:2020:ULV

[85] Shuying Zhuang, Jessie Hui Wang, Pei Zhang, and Jilong Wang. Understanding the latency to visit websites in China: an infrastructure perspective. Computer Networks (Amsterdam, Netherlands: 1999), 169 (??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308722.

Torres:2020:BSR

[86] Pedro R. Torres, Jr., Alberto García-Martínez, Marcelo Bagnulo, and Eduardo Parente Ribeiro. Bartolomeu: an SDN rebalancing system across multiple interdomain paths. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310084.

Anonymous:2020:EBd

[87] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128620301341.

Anonymous:2020:Ma

[88] Anonymous. 14 March 2020. Computer Networks (Amsterdam, Netherlands: 1999), 169(??):??, March 14, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lai:2020:AER

[89] Wei Kuang Lai, Yi Uan Chen, and Tin-Yu Wu. Analysis and evaluation of random-based message propagation models on the social networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305602.

Xie:2020:OVC

[90] An Xie, Huawei Huang, Xiaoliang Wang, Zhuzhong Qian, and Sanglu Lu. Online VNF chain deployment on resource-limited edges by exploiting peer edge devices. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311454.

Siris:2020:PAM

[91] Vasilios A. Siris. Popularity-aware mobile content management: When is knowing less better? Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309120.

Chen:2020:BEB

[92] Kang Chen and Ning Yang. Bw-Share: Efficient bandwidth guarantee in cloud with transparent share adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931120X.

Li:2020:ESS

[93] Chunlin Li, Yihan Zhang, Zhiqiang Hao, and Youlong Luo. An effective scheduling strategy based on hypergraph partition in geographically distributed datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930684X.

Camacho:2020:LAC

[94] José Camacho, Chris McDonald, Ron Peterson, Xia Zhou, and David Kotz. Longitudinal analysis of a campus Wi-Fi network. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308187.

Malik:2020:SRT

[95] Ali Malik, Benjamin Aziz, Mo Adda, and Chih-Heng Ke. Smart routing: Towards proactive fault handling of software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619300271.

Bourahla:2020:PPS

[96] Safia Bourahla, Maryline Laurent, and Yacine Challal. Privacy preservation for social networks sequential publishing. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305547.

Chao:2020:ACH

[97] Chih-Min Chao, Chia-Tsun Chen, and Hsin-Chung Huang. An adjustable channel hopping algorithm for multiradio cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307704.

Castillo:2020:IAI

[98] Edwin Ferney Castillo, Oscar Mauricio Caicedo Rendon, Armando Ordonez, and Lisandro Zambenedetti Granville. IPro: an approach for intelligent SDN monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304608.

Zhang:2020:ELC

[99] Guoqiang Zhang, Yue Wu, Xu Han, Qian Gao, and Jiasi Chen. Exploiting the layer correlation to improve DASH scheduling with scalable video coding. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619307777.

Wang:2020:UAM

[100] Chen Wang, Yan Wang, Yingying Chen, Hongbo Liu, and Jian Liu. User authentication on mobile devices: Approaches, threats and trends. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128618312799.

Huang:2020:TTE

[101] Xinli Huang, Peng Shi, Yufei Liu, and Fei Xu. Towards trusted and efficient SDN topology discovery: a lightweight topology verification scheme. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619307558.

Ullah:2020:PAU

[102] Arif Ullah, Ziaul Haq Abbas, Ghulam Abbas, Fazal Muhammad, and Lei Jiao. Performance analysis of usercentric SBS deployment with load balancing in heterogeneous cellular networks: a Thomas cluster process approach. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308874.

Ding:2020:OCP

[103] Xingjian Ding, Yongcai Wang, Guodong Sun, Chuanwen Luo, Deying Li, Wenping Chen, and Qian Hu. Optimal charger placement for wireless power transfer. Computer Networks (Amsterdam, Netherlands: 1999), 170 (??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619307947.

Ma:2020:GDP

[104] Xiaoyuan Ma, Peilin Zhang, Oliver Theel, and Jianming Wei. Gathering data with packet-in-packet in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619306255.

Anonymous:2020:EBe

[105] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620303364.

Anonymous:2020:Aa

[106] Anonymous. 7 April 2020. Computer Networks (Amsterdam, Netherlands: 1999), 170(??):??, April 7, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Martin-Lopo:2020:LRI

[107] Miguel M. Martín-Lopo, Jaime Boal, and Álvaro Sánchez-Miralles. A literature review of IoT energy platforms aimed at end users. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931271X.

Zubow:2020:PCM

[108] Anatolij Zubow, Ahmad Rostami, and Suzan Bayhan. On practical cooperative multi point transmission for 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861830803X.

Ahmad:2020:SRP

[109] Abrar Ahmad, Adnan Ahmad Cheema, and Dewar Finlay. A survey of radio propagation channel modelling for low altitude flying base stations. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310692.

Bozkaya:2020:SED

[110] Elif Bozkaya and Berk Canberk. SDNenabled deployment and path planning of aerial base stations. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313179.

Midya:2020:QAD

[111] Sadip Midya, Asmita Roy, Koushik Majumder, and Santanu Phadikar. QoS aware distributed dynamic channel allocation for V2V communication in TVWS spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618309228.

Yin:2020:ASM

[112] Wei Yin, Peizhao Hu, Jiahui Wen, and Hongjian Zhou. ACK spoofing on MAC-layer rate control: Attacks and defenses. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311387.

Ayadi:2020:STC

[113] Aya Ayadi, Oussama Ghorbel, M. S. BenSalah, and Mohamed Abid. Spatiotemporal correlations for damages identification and localization in water pipeline systems based on WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309946.

Kotulski:2020:NTI

[114] Zbigniew Kotulski, Tomasz W. Nowak, Mariusz Sepczuk, and Marcin A. Tunia. 5G networks: Types of isolation and their parameters in RAN and CN slices. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304797.

Imputato:2020:RDC

[115] Pasquale Imputato, Stefano Avallone, Mohit P. Tahiliani, and Gautam Ramakrishnan. Revisiting design choices in queue disciplines: the PIE case. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313775.

Donohoe:2020:DBD

[116] Michael Donohoe, Brendan Jennings, and Sasitharan Balasubramaniam. Deep brain drug-delivery control using vagus nerve communications. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930982X.

Vasan:2020:IIB

[117] Danish Vasan, Mamoun Alazab, Sobia Wassan, Hamad Naeem, Babak Safaei, and Qin Zheng. IMCFN: Image-based malware classification using fine-tuned convolutional neural network architecture. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619304736.

Wang:2020:MSD

[118] Yaomin Wang, Xia Wang, Haiyan Li, Yi Dong, Qing Liu, and Xinling Shi. A multi-service differentiation traffic management strategy in SDN cloud data center. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309740.

Wang:2020:WSI

[119] Yingjie Wang, Yang Gao, Yingshu Li, and Xiangrong Tong. A workerselection incentive mechanism for optimizing platform-centric mobile crowdsourcing systems. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311557.

Junior:2020:CDC

[120] Eduardo P. M. Câmara Júnior, Luiz F. M. Vieira, and Marcos A. M. Vieira. CAPTAIN: a data collection algorithm for underwater opticalacoustic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309016.

Sciacca:2020:IRA

[121] Elisabetta C. Sciacca and Laura Galluccio. Impulse response analysis of an ultrasonic human body channel. Computer Networks (Amsterdam, Netherlands: 1999), 171 (??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314938.

Wang:2020:GLA

[122] Danxin Wang, Chuanhe Huang, Xieyang Shen, and Naixue Xiong. A general location-authentication based secure participant recruitment scheme for vehicular crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310618.

Mokryn:2020:OCP

[123] Osnat Mokryn, Adi Akavia, and Yossi Kanizo. Optimal cache placement with local sharing: an ISP guide to the benefits of the sharing economy. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619301562.

Dragonas:2020:FAT

[124] Vasileios Dragonas, Georgios Tsoumanis, George Koufoudakis, Asterios Papamichail, Konstantinos Oikonomou, and Ioannis Stavrakakis. A Fairness-Aware topology independent TDMA MAC policy in time constrained wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316093.

Anonymous:2020:EBf

[125] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862030387X.

Anonymous:2020:Ab

[126] Anonymous. 22 April 2020. Computer Networks (Amsterdam, Netherlands: 1999), 171(??):??, April 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Erdin:2020:BPN

[127] Enes Erdin, Mumin Cebe, Kemal Akkaya, Senay Solak, Eyuphan Bulut, and Selcuk Uluagac. A Bitcoin payment network with reduced transaction fees and confirmation times. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308850.

Hou:2020:CRA

[128] Ronghui Hou, Kaiwen Huang, Huilin Xie, King-Shan Lui, and Hongyan Li. Caching and resource allocation in small cell networks. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309351.

Lytos:2020:TSF

[129] Anastasios Lytos, Thomas Lagkas, Panagiotis Sarigiannidis, Michalis Zervakis, and George Livanos. Towards smart farming: Systems, frameworks and exploitation of multiple sources. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301201.

Radoglou-Grammatikis:2020:CUA

[130] Panagiotis Radoglou-Grammatikis, Panagiotis Sarigiannidis, Thomas Lagkas, and Ioannis Moscholios. A compilation of UAV applications for precision agriculture. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030116X.

Barmpounakis:2020:CAU

[131] Sokratis Barmpounakis, George Tsiatsios, Michael Papadakis, Evangelos Mitsianis, Nikolaos Koursioumpas, and Nancy Alonistioti. Collision avoidance in 5G using MEC and NFV: the vulnerable road user safety use case. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315816.

Bosch:2020:AMI

[132] Patrick Bosch, Steven Latré, and Chris Blondia. An analytical model for IEEE 40802.11 with non-IEEE 40802.11 interfering source. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931148X.

Yucel:2020:USA

[133] Fatih Yucel and Eyuphan Bulut. User satisfaction aware maximum utility task assignment in mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311697.

Boz:2020:HAQ

[134] Eren Boz and Jukka Manner. A hybrid approach to QoS measurements in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314240.

Abbasian:2020:CNA

[135] Amir Abbasian and Masoumeh Safkhani. CNCAA: a new anti-collision algorithm using both collided and noncollided parts of information. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619312241.

Adhikari:2020:IBF

[136] Devlina Adhikari, Debasish Datta, and Raja Datta. Impact of BER in fragmentation-aware routing and spectrum assignment in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311776.

Abbas:2020:PAT

[137] Nadine Abbas, Sanaa Sharafeddine, Hazem Hajj, and Zaher Dawy. Priceaware traffic splitting in D2D HetNets with cost-energy-QoE tradeoffs. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312307.

Hajisami:2020:ERP

[138] Abolfazl Hajisami, Tuyen X. Tran, Ayman Younis, and Dario Pompili. Elastic resource provisioning for increased energy efficiency and resource utilization in cloud-RANs. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316019.

Xu:2020:GBN

[139] Lingwei Xu, Tianqi Quan, Jingjing Wang, T. Aaron Gulliver, and Khoa N. Le. GR and BP neural networkbased performance prediction of dualantenna mobile communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309545.

Li:2020:SVC

[140] Xiaotong Li, Ruiting Zhou, Ying-Jun Angela Zhang, Lei Jiao, and Zongpeng Li. Smart vehicular communication via 5G mmWaves. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313672.

Zhou:2020:DCI

[141] Man Zhou, Lansheng Han, Hongwei Lu, and Cai Fu. Distributed collaborative intrusion detection system for vehicular ad hoc networks based on invariant. Computer Networks (Amsterdam, Netherlands: 1999), 172 (??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311879.

Nguyen:2020:WEH

[142] Tan N. Nguyen, Phuong T. Tran, and Miroslav Voznak. Wireless energy harvesting meets receiver diversity: a successful approach for twoway half-duplex relay networks over block Rayleigh fading channel. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314689.

Abbasloo:2020:SSW

[143] Soheil Abbasloo, Yang Xu, and H. Jonathan Chao. To schedule or not to schedule: When no-scheduling can beat the best-known flow scheduling algorithm in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619309788.

Khammassi:2020:NLW

[144] Chaouki Khammassi and Saoussen Krichen. A NSGA2-LR wrapper approach for feature selection in network intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315270.

Cicconetti:2020:UAS

[145] Claudio Cicconetti, Marco Conti, and Andrea Passarella. Uncoordinated access to serverless computing in MEC systems for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313684.

Anonymous:2020:EBg

[146] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304722.

Anonymous:2020:Mb

[147] Anonymous. 8 May 2020. Computer Networks (Amsterdam, Netherlands: 1999), 172(??):??, May 8, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Park:2020:AEP

[148] Jeman Park, Manar Mohaisen, Dae-Hun Nyang, and Aziz Mohaisen. Assessing the effectiveness of pulsing denial of service attacks under realistic network synchronization assumptions. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313118.

Martiradonna:2020:ASO

[149] Sergio Martiradonna, Alessandro Grassi, Giuseppe Piro, and Gennaro Boggia. 5G-air-simulator: an opensource tool modeling the 5G air interface. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317359.

Geng:2020:ODR

[150] Sunyue Geng, Sifeng Liu, Zhigeng Fang, and Su Gao. An optimal delay routing algorithm considering delay variation in the LEO satellite communication network. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313040.

Hosseini:2020:NHM

[151] Soodeh Hosseini and Behnam Mohammad Hasani Zade. New hybrid method for attack detection using combination of evolutionary algorithms, SVM, and ANN. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619302191.

Zhao:2020:RSL

[152] Huiyun Zhao, Zhisong Pan, and Wei Tao. Regularized shapelet learning for scalable time series classification. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309181.

Tu:2020:TAL

[153] Shanshan Tu, Muhammad Waqas, Qiangqiang Lin, Sadaqat Ur Rehman, Muhammad Hanif, Chuangbai Xiao, M. Majid Butt, and Chin-Chen Chang. Tracking area list allocation scheme based on overlapping community algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309326.

Sanchez:2020:DDS

[154] P. A. Sánchez, S. Luna-Ramírez, M. Toril, C. Gijón, and J. L. Bejarano-Luque. A data-driven scheduler performance model for QoE assessment in a LTE radio network planning tool. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/

                                        article/pii/S1389128619309338. See

corrigendum [271].

Kondo:2020:NDN

[155] Daishi Kondo, Vassilis Vassiliades, Thomas Silverston, Hideki Tode, and Tohru Asami. The named data networking flow filter: Towards improved security over information leakage attacks. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619316524.

Takemasa:2020:DPF

[156] Junji Takemasa, Yuki Koizumi, and Toru Hasegawa. Data prefetch for fast NDN software routers based on hash table-based forwarding tables. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931196X.

Suhail:2020:PEP

[157] Sabah Suhail, Rasheed Hussain, Mohammad Abdellatif, Shashi Raj Pandey, Abid Khan, and Choong Seon Hong. Provenance-enabled packet path tracing in the RPL-based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619316123.

Wu:2020:CMB

[158] Zheng Wu, Yu ning Dong, Hua-Liang Wei, and Wei Tian. Consistency measure based simultaneous feature selection and instance purification for multimedia traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311624.

Li:2020:UCM

[159] Zhangyu Li, Soham Desai, Vaishnendr D. Sudev, Pu Wang, Jinsong Han, and Zhi Sun. Underwater cooperative MIMO communications using hybrid acoustic and magnetic induction technique. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300165.

Huang:2020:ESO

[160] Mingfeng Huang, Anfeng Liu, Neal N. Xiong, Tian Wang, and Athanasios V. Vasilakos. An effective serviceoriented networking management architecture for 5G-enabled Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311442.

Rastogi:2020:NIT

[161] Eshita Rastogi, Navrati Saxena, Abhishek Roy, and Dong Ryeol Shin. Narrowband Internet of Things: a comprehensive study. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313593.

Barik:2020:UTP

[162] Runa Barik, Michael Welzl, Gorry Fairhurst, Ahmed Elmokashfi, Thomas Dreibholz, and Stein Gjessing. On the usability of transport protocols other than TCP: a home gateway and Internet path traversal study. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931446X.

Pacheco:2020:FCH

[163] Fannia Pacheco, Ernesto Exposito, and Mathieu Gineste. A framework to classify heterogeneous Internet traffic with machine learning and deep learning techniques for satellite communications. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313544.

Pascoal:2020:SDS

[164] Túlio A. Pascoal, Iguatemi E. Fonseca, and Vivek Nigam. Slow denialof-service attacks on software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312927.

Ma:2020:AFS

[165] Lin Ma, Kai Bu, Ningchao Wu, Tianxiang Luo, and Kui Ren. Atlas: a first step toward multipath validation. Computer Networks (Amsterdam, Netherlands: 1999), 173 (??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310667.

Anonymous:2020:EBh

[166] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620305363.

Anonymous:2020:Mc

[167] Anonymous. 22 May 2020. Computer Networks (Amsterdam, Netherlands: 1999), 173(??):??, May 22, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2020:SAO

[168] Ran Zhang, Jiang Liu, Renchao Xie, Tao Huang, F. Richard Yu, and Yunjie Liu. Service-aware optimal caching placement for named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311193.

Carrega:2020:CEE

[169] Alessandro Carrega and Matteo Repetto. Coupling energy efficiency and quality for consolidation of cloud workloads. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316226.

Zhang:2020:SSA

[170] Shunliang Zhang, Dali Zhu, and Yongming Wang. A survey on spaceaerial-terrestrial integrated 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314045.

Hassan:2020:NBN

[171] Muhammad Hassan, Mauro Conti, and Chhagan Lal. NC based DAS– NDN: Network coding for robust dynamic adaptive streaming over NDN. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312551.

Mazzenga:2020:ESG

[172] Franco Mazzenga, Romeo Giuliano, and Francesco Vatalaro. Effective strategies for gradual copperto-fiber transition in access networks. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301183.

Zhao:2020:IMM

[173] Yi Zhao, Ke Xu, Yifeng Zhong, Xiang-Yang Li, Ning Wang, Hui Su, Meng Shen, and Ziwei Li. Incentive mechanisms for mobile data offloading through operator-owned WiFi access points. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316858.

Nouri:2020:MTW

[174] Parisa Nouri, Hirley Alves, Mikko A. Uusitalo, Onel Alcaraz López, and Matti Latva-aho. Machine-type wireless communications enablers for beyond 5G: Enabling URLLC via diversity under hard deadlines. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315361.

Curado:2020:EEI

[175] Marilia Curado, Giacomo Tanganelli, Antonio A. F. Loureiro, and Eirini Eleni Tsiropoulou. Editorial: 5G-Enabled Internet of Things, applications and services. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304333.

Calvo-Palomino:2020:ECR

[176] Roberto Calvo-Palomino, Héctor Cordobés, Markus Engel, Markus Fuchs, Pratiksha Jain, Marc Liechti, Sreeraj Rajendran, Matthias Schäfer, Bertold Van den Bergh, Sofie Pollin, Domenico Giustiniano, and Vincent Lenders. Electrosense+: Crowdsourcing radio spectrum decoding using IoT receivers. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314471.

Yousefi:2020:FTA

[177] Faezeh Yousefi, Akbar Ghaffarpour Rahbar, and Amin Ghadesi. Fragmentation and time aware algorithms in spectrum and spatial assignment for space division multiplexed elastic optical networks (SDM-EON). Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931597X.

Hofer-Schmitz:2020:TFV

[178] Katharina Hofer-Schmitz and Branka Stojanović. Towards formal verification of IoT protocols: a review. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317116.

Dimitriou:2020:ECF

[179] Tassos Dimitriou. Efficient, coercionfree and universally verifiable Blockchainbased voting. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317414.

Nguyen:2020:MBN

[180] Linh Nguyen and Hoc T. Nguyen. Mobility based network lifetime in wireless sensor networks: a review. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315865.

Galal:2020:PBP

[181] Akram Galal and Xavier Hesselbach. Probability-based path discovery protocol for electromagnetic nano-networks. Computer Networks (Amsterdam, Netherlands: 1999), 174 (??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619308801.

Zhou:2020:BEI

[182] Yuyang Zhou, Guang Cheng, Shanqing Jiang, and Mian Dai. Building an efficient intrusion detection system based on feature selection and ensemble classifier. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314203.

Marche:2020:HES

[183] Claudio Marche, Luigi Atzori, Virginia Pilloni, and Michele Nitti. How to exploit the Social Internet of Things: Query generation model and device profiles’ dataset. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861931730X.

Anonymous:2020:EBi

[184] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128620307386.

Anonymous:2020:Jb

[185] Anonymous. 19 June 2020. Computer Networks (Amsterdam, Netherlands: 1999), 174(??):??, June 19, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shi:2020:QAU

[186] Lin Shi, Shoukun Xu, Haoyu Liu, and Zhongxu Zhan. QoS-Aware UAV Coverage path planning in 5G mmWave network. Computer Networks (Amsterdam, Netherlands: 1999), 175 (??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313015.

Sharma:2020:AFT

[187] Kapil Sharma and Ram Narayan Yadav. An adaptive, fault tolerant, flow-level routing scheme for data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309107.

Qian:2020:MMM

[188] Shenshen Qian, Bowen Wang, Song Li, Yanjing Sun, Yi Yu, and Jingjing Wang. Many-to-many matching for social-aware minimized redundancy caching in D2D-enabled cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 175 (??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316299.

Naeem:2020:SRH

[189] Makhdoom Muhammad Naeem, Intesab Hussain, and Malik Muhammad Saad Missen. A survey on registration hijacking attack consequences and protection for session initiation protocol (SIP). Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312332.

Shen:2020:FSA

[190] Gengbiao Shen, Qing Li, Yong Jiang, Yu Wu, and Jianhui Lv. A fourstage adaptive scheduling scheme for service function chain in NFV. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314276.

Cho:2020:TAC

[191] Eunsang Cho, Jeongnyeo Kim, Minkyung Park, Hyeonmin Lee, Chorom Hamm, Soobin Park, Sungmin Sohn, Minhyeok Kang, and Ted Taekyoung Kwon. TwinPeaks: an approach for certificateless public key distribution for the Internet and Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128618314051.

Yang:2020:JSC

[192] Yang Yang, Songtao Guo, Guiyan Liu, and Quyuan Wang. Joint source coding rate allocation and flow scheduling for data aggregation in collaborative sensing networks. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314707.

Mollahasani:2020:DAM

[193] Shahram Mollahasani, Alperen Eroglu, Ilker Demirkol, and Ertan Onur. Density-aware mobile networks: Opportunities and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309910.

Antevski:2020:INM

[194] Kiril Antevski, Carlos J. Bernardos, Luca Cominardi, Antonio de la Oliva, and Alain Mourad. On the integration of NFV and MEC technologies: architecture analysis and benefits for edge robotics. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300797.

Zheng:2020:RWP

[195] Yuan Zheng, Suzhi Bi, Xiaohui Lin, and Hui Wang. Reusing wireless power transfer for backscatterassisted relaying in WPCNs. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309399.

Ma:2020:CTC

[196] Lujuan Ma, Xiaoping Liu, Huanqing Wang, and Xiaoping Deng. Congestion tracking control for multirouter TCP/AQM network based on integral backstepping. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300013.

Zhou:2020:LCW

[197] Pengpeng Zhou, Yang Wang, Zhenyu Li, Gareth Tyson, Hongtao Guan, and Gaogang Xie. Logchain: Cloud workflow reconstruction and troubleshooting with unstructured logs. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316731.

BenHassouna:2020:MUD

[198] Asma Ben Hassouna, Hend Koubaa, and Leila Azouz Saidane. Multiuser diversity wireless multicast: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 175 (??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308564.

Yin:2020:FFA

[199] Wei Yin, Peizhao Hu, Wenbo Wang, Jiahui Wen, and Hongjian Zhou. FA-SUS: a fast association mechanism for 802.11ah networks. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312083.

Anonymous:2020:EBj

[200] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620307957.

Anonymous:2020:Jc

[201] Anonymous. 5 July 2020. Computer Networks (Amsterdam, Netherlands: 1999), 175(??):??, July 5, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Olatinwo:2020:EEM

[202] Segun O. Olatinwo and Trudi-H. Joubert. Energy efficiency maximization in a wireless powered IoT sensor network for water quality monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308709.

Dwivedi:2020:ECG

[203] Shubhra Dwivedi, Manu Vardhan, and Sarsij Tripathi. An effect of chaos grasshopper optimization algorithm for protection of network infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310175.

Costa:2020:SBH

[204] Allan Costa, Lucas Pacheco, Denis Rosário, Leandro Villas, Antonio A. F. Loureiro, Susana Sargento, and Eduardo Cerqueira. Skipping-based handover algorithm for video distribution over ultra-dense VANET. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302188.

AlKhansa:2020:HSG

[205] Rasha Al Khansa, Hassan A. Artail, Mohamad Assaad, and Karim Kabalan. A Hybrid Scheduled and group-based random access solution for massive MTC networks. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310679.

Alami:2020:EQE

[206] Abdul Jawad Alami and Hassan Artail. Energy and quality of experience aware web browsing for smartphones. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317013.

Darabkh:2020:LAG

[207] Khalid A. Darabkh, Wafa’a K. Kassab, and Ala’ F. Khalifeh. LiM-AHP-G-C: Life time maximizing based on analytical hierarchal process and genetic clustering protocol for the Internet of Things environment. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316111.

Dong:2020:CCE

[208] Cong Dong, Chen Zhang, Zhigang Lu, Baoxu Liu, and Bo Jiang. CETAnalytics: Comprehensive effective traffic information analytics for encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309466.

Desikan:2020:TCF

[209] K. E. Srinivasa Desikan, Vijeth J. Kotagi, and C. Siva Ram Murthy. Topology control in fog computing enabled IoT networks for smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620301468.

Stamou:2020:CAH

[210] Adamantia Stamou, Nikos Dimitriou, Kimon Kontovasilis, and Symeon Papavassiliou. Context-aware handover management for HetNets: Performance evaluation models and comparative assessment of alternative context acquisition strategies. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315075.

Xia:2020:NOL

[211] Qiufen Xia, Zheng Lou, Wenzheng Xu, and Zichuan Xu. Near-optimal and learning-driven task offloading in a 5G multi-cell mobile edge cloud. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620301699.

Rasheed:2020:PPS

[212] Iftikhar Rasheed, Lin Zhang, and Fei Hu. A privacy preserving scheme for vehicle-to-everything communications using 5G mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315658.

Kaltenberger:2020:ODI

[213] Florian Kaltenberger, Aloizio P. Silva, Abhimanyu Gosain, Luhan Wang, and Tien-Thinh Nguyen. OpenAirInterface: Democratizing innovation in the 5G Era. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314410.

Rezar:2020:ITD

[214] Matija Rezar and Fabio Ricciato. On the impact of time-of-departure knowledge on the accuracy of timeof-arrival localization. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313532.

Cao:2020:RSE

[215] Zhi Cao, Honggang Zhang, Benyuan Liu, and Bo Sheng. Revenue sharing in edge-cloud systems: a gametheoretic perspective. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931028X.

Trevisan:2020:ERE

[216] Martino Trevisan, Ali Safari Khatouni, and Danilo Giordano. ER-RANT: Realistic emulation of radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 176 (??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301420.

Favale:2020:CTL

[217] Thomas Favale, Francesca Soro, Martino Trevisan, Idilio Drago, and Marco Mellia. Campus traffic and e-Learning during COVID-19 pandemic. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620306046.

Zhao:2020:FAT

[218] Zhipeng Zhao, Weidong Yang, and Bin Wu. Flow aggregation through dynamic routing overlaps in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311818.

Liu:2020:RRA

[219] Zhixin Liu, Guochen Hou, Yazhou Yuan, Kit Yan Chan, Kai Ma, and Xinping Guan. Robust resource allocation in two-tier NOMA heterogeneous networks toward 5G. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314422.

Anonymous:2020:EBk

[220] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620308987.

Anonymous:2020:Jd

[221] Anonymous. 20 July 2020. Computer Networks (Amsterdam, Netherlands: 1999), 176(??):??, July 20, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sun:2020:MEC

[222] Penghao Sun, Zehua Guo, Gang Wang, Julong Lan, and Yuxiang Hu. MARVEL: Enabling controller load balancing in software-defined networks with multi-agent reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931566X.

Alioua:2020:UTM

[223] Ahmed Alioua, Houssem eddine Djeghri, Mohammed Elyazid Tayeb Cherif, Sidi-Mohammed Senouci, and Hichem Sedjelmaci. UAVs for traffic monitoring: a sequential gamebased computation offloading/sharing approach. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315798.

Campos:2020:AHN

[224] Pablo Campos, Ángela Hernández-Solana, and Antonio Valdovinos-Bardají. Analysis of hidden node problem in LTE networks deployed in unlicensed spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300396.

Santos:2020:MTF

[225] Hugo Santos, Derian Alencar, Rodolfo Meneguette, Denis Rosário, Jéferson Nobre, Cristiano Both, Eduardo Cerqueira, and Torsten Braun. A multi-tier fog content orchestrator mechanism with quality of experience support. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304448.

Wu:2020:BAA

[226] Lijuan Wu, Chaoqin Gan, Hubao Qiao, Jianqiang Hui, and Zhongsen Xu. Bandwidth allocation algorithm based on differential BRP models in Ethernet PON. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315993.

Pejovic:2020:RDM

[227] Veljko Pejović, Ivan Majhen, Miha Janez, and Blaz Zupan. RICER-CANDO: Data mining toolkit for mobile broadband measurements. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620301535.

Marinova:2020:EEN

[228] Simona Marinova, Thomas Lin, Hadi Bannazadeh, and Alberto Leon-Garcia. End-to-end network slicing for future wireless in multi-region cloud platforms. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316081.

Gharib:2020:ASH

[229] Mohammed Gharib, Ahmad Foroozani, Shahbaz Rezaei, Ali Mohammad Afshin Hemmatyar, and Ali Movaghar. An area-scalable human-based mobility model. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619308667.

Kim:2020:HNS

[230] Sungwook Kim. Heterogeneous network spectrum allocation scheme based on three-phase bargaining game. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309727.

Sigwele:2020:EEC

[231] Tshiamo Sigwele, Yim Fun Hu, and Misfa Susanto. Energy-efficient 5G cloud RAN with virtual BBU server consolidation and base station sleeping. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301742.

Zaki-Hindi:2020:MTU

[232] Ayat Zaki-Hindi, Salah-Eddine Elayoubi, and Tijani Chahed. Multitenancy and URLLC on unlicensed spectrum: Performance and design. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300220.

Cheng:2020:JUA

[233] Zhipeng Cheng, Ning Chen, Bang Liu, Zhibin Gao, Lianfen Huang, Xiaojiang Du, and Mohsen Guizani. Joint user association and resource allocation in HetNets based on user mobility prediction. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030298X.

Martiradonna:2020:UAS

[234] Sergio Martiradonna, Alessandro Grassi, Giuseppe Piro, and Gennaro Boggia. Understanding the 5G-air-simulator: a tutorial on design criteria, technical components, and reference use cases. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317347.

Zhang:2020:ECN

[235] Hongpo Zhang, Lulu Huang, Chase Q. Wu, and Zhanbo Li. An effective convolutional neural network based on SMOTE and Gaussian mixture model for intrusion detection in imbalanced dataset. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300712.

Sai:2020:EOA

[236] Yinghui Sai, Xiaotong Li, Ruiting Zhou, and Zongpeng Li. An efficient online auction for resource leasing in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912861930790X.

Chatterjee:2020:DBR

[237] Bijoy Chand Chatterjee and Eiji Oki. Defragmentation based on route partitioning in 1 + 1 protected elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310138.

Li:2020:ARV

[238] Yahui Li, Zhiliang Wang, Xia Yin, Xingang Shi, Jianping Wu, Fangdan Ye, Jiangyuan Yao, and Han Zhang. Assisting reachability verification of network configurations updates with NUV. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314458.

Fu:2020:TOA

[239] Xiuwen Fu, Pasquale Pace, Gianluca Aloi, Lin Yang, and Giancarlo Fortino. Topology optimization against cascading failures on wireless sensor networks using a memetic algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317888.

Sharma:2020:VAA

[240] Gourav Prateek Sharma, Wouter Tavernier, Didier Colle, and Mario Pickavet. VNF-AAPC: Acceleratoraware VNF placement and chaining. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030058X.

Madureira:2020:SID

[241] André Luiz R. Madureira, Francisco Renato C. Araújo, and Leobino N. Sampaio. On supporting IoT data aggregation through programmable data planes. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301195.

Fotouhi:2020:LST

[242] Mahdi Fotouhi, Majid Bayat, Ashok Kumar Das, Hossein Abdi Nasib Far, S. Morteza Pournaghi, and M. A. Doostari. A lightweight and secure two-factor authentication scheme for wireless body area networks in healthcare IoT. Computer Networks (Amsterdam, Netherlands: 1999), 177 (??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316457.

Anonymous:2020:EBl

[243] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128620310811.

Anonymous:2020:Ac

[244] Anonymous. 4 August 2020. Computer Networks (Amsterdam, Netherlands: 1999), 177(??):??, August 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lopes:2020:GAP

[245] Ana Paula Golembiouski Lopes and Paulo R. L. Gondim. Group authentication protocol based on aggregated signatures for D2D communication. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313076.

Jafarian:2020:DAT

[246] Besat Jafarian, Nasser Yazdani, and Mohammad Sayad Haghighi. Discrimination-aware trust management for social Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128619316743.

Wei:2020:AFU

[247] Wei Wei, Qiao Ke, Jakub Nowak, Marcin Korytkowski, Rafał Scherer, and Marcin Woźniak. Accurate and fast URL phishing detector: a convolutional neural network approach. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301109.

Hu:2020:CCU

[248] Yuhan Hu, Ye Tian, Wen Yang, Xiaodong Wang, and Xinming Zhang. Content to cash: Understanding and improving crowdsourced live video broadcasting services with monetary donations. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311545.

Tarek:2020:NSP

[249] Dina Tarek, Abderrahim Benslimane, M. Darwish, and Amira M. Kotb. A new strategy for packets scheduling in cognitive radio Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313994.

Joshi:2020:PLP

[250] Kalpana D. Joshi and Kotaro Kataoka. pSMART: a lightweight, privacy-aware service function chain orchestration in multi-domain NFV/SDN. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311181.

Uniyal:2020:ETS

[251] Navdeep Uniyal, Abubakar Siddique Muqaddas, Dimitrios Gkounis, Anderson Bravalheri, Shadi Moazzeni, Fragkiskos Sardis, Mischa Dohler, Reza Nejabati, and Dimitra Simeonidou. 5GUK exchange: Towards sustainable end-to-end multi-domain orchestration of softwarized 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619316287.

Wang:2020:EED

[252] Denghui Wang, Jian Liu, and Dezhong Yao. An energy-efficient distributed adaptive cooperative routing based on reinforcement learning in wireless multimedia sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313568.

Kalkan:2020:TTF

[253] Kübra Kalkan and Kasper Rasmussen. TruSD: Trust framework for service discovery among IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619311636.

Kalkan:2020:SSU

[254] Kübra Kalkan. SUTSEC: SDN utilized trust based secure clustering in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310783.

Karamyshev:2020:FAA

[255] A. Karamyshev, E. Khorov, A. Krasilov, and I. F. Akyildiz. Fast and accurate analytical tools to estimate network capacity for URLLC in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316445.

Banditwattanawong:2020:FOA

[256] Thepparit Banditwattanawong and Masawee Masdisornchote. On formulation of online algorithm and framework of near-optimally tractable eviction for nonuniform caches. Computer Networks (Amsterdam, Netherlands:

1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301870.

Zhou:2020:PCO

[257] Shuchen Zhou and Waqas Jadoon. The partial computation offloading strategy based on game theory for multiuser in mobile edge computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316482.

Trivedi:2020:DSA

[258] Hiral S. Trivedi and Sankita J. Patel. Design of secure authentication protocol for dynamic user addition in distributed Internet-of-Things. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316949.

Lu:2020:NIE

[259] Mengke Lu. Node importance evaluation based on neighborhood structure hole and improved TOPSIS. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861931031X.

Fulber-Garcia:2020:NST

[260] Vinicius Fulber-Garcia, Elias P. Duarte, Alexandre Huff, and Carlos R. P. dos Santos. Network service topology: Formalization, taxonomy and the CUSTOM specification model. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619312460.

Wang:2020:PLB

[261] Haibo Wang, Hongli Xu, Chen Qian, Juncheng Ge, Jianchun Liu, and He Huang. PrePass: Load balancing with data plane resource constraints using commodity SDN switches. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620305272.

Cui:2020:CDA

[262] Jin Cui, Khaled Boussetta, and Fabrice Valois. Classification of data aggregation functions in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620307143.

Halder:2020:SAS

[263] Subir Halder, Amrita Ghosal, and Mauro Conti. Secure over-the-air software updates in connected vehicles: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314963.

Benil:2020:CBS

[264] T. Benil and J. Jasper. Cloud based security on outsourcing using blockchain in e-health systems. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316986.

Ashtiani:2020:PAR

[265] Amirhossein Feizi Ashtiani and Samuel Pierre. Power allocation and resource assignment for secure D2D communication underlaying cellular networks: a tabu search approach. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314392.

Sarieddine:2020:FMR

[266] Khaled Sarieddine, Malak Charaf, Mohammad Ayad, and Hassan Artail. A framework for mobile relay node selection for serving outdoor cell edge users. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302164.

Weerasinghe:2020:PBI

[267] Thilina N. Weerasinghe, Indika A. M. Balapuwaduge, and Frank Y. Li. Priority-based initial access for URLLC traffic in massive IoT networks: Schemes and performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 178(??): ??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317967.

Gupta:2020:CSA

[268] Surbhi Gupta, Sangeeta R, Ravi Shankar Mishra, Gaurav Singal, Tapas Badal, and Deepak Garg. Corridor segmentation for automatic robot navigation in indoor environment using edge devices. Computer Networks (Amsterdam, Netherlands: 1999), 178 (??):??, September 4, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620306782.

Anonymous:2020:EBm

[269] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620311385.

Anonymous:2020:S

[270] Anonymous. 4 September 2020. Computer Networks (Amsterdam, Netherlands: 1999), 178(??):??, September 4, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sanchez:2020:CDD

[271] P. A. Sánchez, S. Luna-Ramírez, M. Toril, C. Gijón, and J. L. Bejarano-Luque. Corrigendum to “A datadriven scheduler performance model for QoE assessment in a LTE radio network planning tool” Computer Networks 173 (2020) 107186. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304321. See

[154].

Sun:2020:SIP

[272] Penghao Sun, Zehua Guo, Sen Liu, Julong Lan, Junchao Wang, and Yuxiang Hu. SmartFCT: Improving powerefficiency for data center networks with deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300384.

Sun:2020:CUP

[273] Fanghui Sun, Shen Wang, Chunrui Zhang, and Hongli Zhang. Clustering of unknown protocol messages based on format comparison. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030445X.

Asghari:2020:CRM

[274] Ali Asghari, Mohammad Karim Sohrabi, and Farzin Yaghmaee. A cloud resource management framework for multiple online scientific workflows using cooperative reinforcement learning agents. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305791.

Singh:2020:ATS

[275] Vikash Singh, Roshan Kumar, and Zhao Wei. Adaptive time-switching and power-splitting protocols for energy harvesting sensor networks with multiple relays. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912861930862X.

Sicari:2020:ITE

[276] Sabrina Sicari, Alessandra Rizzardi, and Alberto Coen-Porisini. 5G in the Internet of Things era: an overview on security and privacy challenges. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300827.

Krzywiecki:2020:SLI

[277] Lukasz Krzywiecki, Adam Bobowski, Marta Słowik, Marcin Słowik, and Patryk Kozieł. Schnorr-like identification scheme resistant to malicious subliminal setting of ephemeral secret. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620301560.

Jeske:2020:DTO

[278] Marlon Jeske, Valério Rosset, and Mariá C. V. Nascimento. Determining the trade-offs between data delivery and energy consumption in largescale WSNs by multi-objective evolutionary optimization. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620303741.

Abdelmoneem:2020:MAT

[279] Randa M. Abdelmoneem, Abderrahim Benslimane, and Eman Shaaban. Mobility-aware task scheduling in cloud-fog IoT-based healthcare architectures. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619313106.

Kibalya:2020:NDP

[280] Godfrey Kibalya, Joan Serrat, Juan-Luis Gorricho, Haipeng Yao, and Peiying Zhang. A novel dynamic programming inspired algorithm for embedding of virtual networks in future networks. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317980.

Huang:2020:REW

[281] Tao Huang, Bin Tang, Baoliu Ye, Zhihao Qu, and Sanglu Lu. Rateless802.11: Extending WiFi applicability in extremely poor channels. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): ??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619316202.

Abdulqadder:2020:MLI

[282] Ihsan H. Abdulqadder, Shijie Zhou, Deqing Zou, Israa T. Aziz, and Syed Muhammad Abrar Akber. Multilayered intrusion detection and prevention in the SDN/NFV enabled cloud of 5G networks using AI-based defense mechanisms. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310205.

Palmieri:2020:RLA

[283] Francesco Palmieri. A reliability and latency-aware routing framework for 5G transport infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620305132.

Hussain:2020:APE

[284] Anwar Hussain, Shah Nazir, Sulaiman Khan, and Ayaz Ullah. Analysis of PMIPv6 extensions for identifying and assessing the efforts made for solving the issues in the PMIPv6 domain: a systematic review. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619315889.

deOliveira:2020:BRB

[285] Marcela T. de Oliveira, Lúcio H. A. Reis, Dianne S. V. Medeiros, Ricardo C. Carrano, Sílvia D. Olabarriaga, and Diogo M. F. Mattos. Blockchain reputation-based consensus: a scalable and resilient mechanism for distributed mistrusting applications. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300360.

Seyfollahi:2020:LLB

[286] Ali Seyfollahi and Ali Ghaffari. A lightweight load balancing and route minimizing solution for routing protocol for low-power and lossy networks. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619310357.

Zhang:2020:CFS

[287] Qianying Zhang and Shijun Zhao. A comprehensive formal security analysis and revision of the two-phase key exchange primitive of TPM 2.0. Computer Networks (Amsterdam, Netherlands: 1999), 179(??): ??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619317918.

Sen:2020:TPI

[288] Priyangshu Sen, Dimitris A. Pados, Stella N. Batalama, Erik Einarsson, Jonathan P. Bird, and Josep M. Jornet. The TeraNova platform: an integrated testbed for ultra-broadband wireless communications at true Terahertz frequencies. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304473.

Bhat:2020:OBL

[289] Soumya J. Bhat and Santhosh K. Venkata. An optimization based localization with area minimization for heterogeneous wireless sensor networks in anisotropic fields. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620306083.

Li:2020:ATR

[290] Ding Li, Wenzhong Li, Xiaoliang Wang, Cam-Tu Nguyen, and Sanglu Lu. App trajectory recognition over encrypted Internet traffic based on deep neural network. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620302140.

Alizadeh:2020:CCK

[291] Faezeh Alizadeh and Amir Jalaly Bidgoly. Cipher chaining key resynchronization in LPWAN IoT network using a deep learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620301006.

Aliouat:2020:ERP

[292] Lina Aliouat, Hakim Mabed, and Julien Bourgeois. Efficient routing protocol for concave unstable terahertz nanonetworks. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304205.

Demiane:2020:OUT

[293] Freddy Demiane, Sanaa Sharafeddine, and Omar Farhat. An optimized UAV trajectory planning for localization in disaster scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313234.

Kumar:2020:OCC

[294] Sumit Kumar and Rajeev Tiwari. Optimized content centric networking for future Internet: Dynamic popularity window based caching scheme. Computer Networks (Amsterdam, Netherlands: 1999), 179 (??):??, October 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311233.

Anonymous:2020:EBn

[295] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620311671.

Anonymous:2020:Oa

[296] Anonymous. 9 October 2020. Computer Networks (Amsterdam, Netherlands: 1999), 179(??):??, October 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chouayakh:2020:LSA

[297] Ayman Chouayakh, Aurélien Bechler, Isabel Amigo, Loutfi Nuaymi, and Patrick Maillé. Licensed shared access for 5G: Which auction mechanism to choose? Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619311491.

Shahraki:2020:COW

[298] Amin Shahraki, Amir Taherkordi, Øystein Haugen, and Frank Eliassen. Clustering objectives in wireless sensor networks: a survey and research direction analysis. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620303121.

Li:2020:PPB

[299] Xionglve Li, Zhiping Cai, Bingnan Hou, Ning Liu, Fang Liu, and Jieren Cheng. ProbInfer: Probability-based AS path inference from multigraph perspective. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302036.

Panigrahy:2020:NCD

[300] Nitish K. Panigrahy, Jian Li, Don Towsley, and C. V. Hollot. Network cache design under stationary requests: Exact analysis and Poisson approximation. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619313271.

Bonati:2020:CZT

[301] Leonardo Bonati, Salvatore D’Oro, Lorenzo Bertizzolo, Emrecan Demirors, Zhangyu Guan, Stefano Basagni, and Tommaso Melodia. CellOS: Zerotouch softwarized open cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862030503X.

Lopez-Raventos:2020:CDC

[302] Álvaro López-Raventós and Boris Bellalta. Concurrent decentralized channel allocation and access point selection using multi-armed bandits in multi BSS WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314823.

Garcia-Aviles:2020:AFF

[303] Gines Garcia-Aviles, Carlos Donato, Marco Gramaglia, Pablo Serrano, and Albert Banchs. ACHO: a framework for flexible re-orchestration of virtual network functions. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619317165.

Sasabe:2020:MEA

[304] Masahiro Sasabe. Mathematical epidemiological analysis of dynamics of delay attacks on pull-based competitive information diffusion. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619310904.

Vergados:2020:LVN

[305] Dimitrios J. Vergados, Katina Kralevska, Yuming Jiang, and Angelos Michalas. Local voting: a new distributed bandwidth reservation algorithm for 6TiSCH networks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128619314628.

Muratkar:2020:BLI

[306] Tushar S. Muratkar, Ankit Bhurane, and Ashwin Kothari. Battery-less Internet of Things — a survey. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620303984.

Lin:2020:NTD

[307] Yunsenxiao Lin, Yu Zhou, Zhengzheng Liu, Ke Liu, Yangyang Wang, Mingwei Xu, Jun Bi, Ying Liu, and Jianping Wu. NetView: Towards ondemand network-wide telemetry in the data center. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620302449.

Rajaguru:2020:HSS

[308] R. Rajaguru, K. Vimala Devi, and P. Marichamy. A hybrid spectrum sensing approach to select suitable spectrum band for cognitive users. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620304850.

Zhang:2020:TCT

[309] Quanxin Zhang, Yuhang Zhao, Yajie Wang, Thar Baker, Jian Zhang, and Jingjing Hu. Towards cross-task universal perturbation against blackbox object detectors in autonomous driving. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862030606X.

Sasikumar:2020:NMO

[310] Syama Sasikumar and J. Jayakumari. A novel method for the optimization of spectral-energy efficiency tradeoff in 5G heterogeneous cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620302115.

Kasim:2020:ERD

[311] Ömer Kasim. An efficient and robust deep learning based network anomaly detection against distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620304114.

Blaise:2020:DZD

[312] Agathe Blaise, Mathieu Bouet, Vania Conan, and Stefano Secci. Detection of zero-day attacks: an unsupervised port-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620300761.

Giannone:2020:OHM

[313] Francesco Giannone, Pantelis A. Frangoudis, Adlen Ksentini, and Luca Valcarenghi. Orchestrating heterogeneous MEC-based applications for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620301997.

Chousainov:2020:AFC

[314] Iskanter-Alexandros Chousainov, Ioannis Moscholios, Panagiotis Sarigiannidis, Alexandros Kaloxylos, and Michael Logothetis. An analytical framework of a C-RAN supporting random, quasirandom and bursty traffic. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620306484.

Alostad:2020:DPR

[315] Jasem M. Alostad. Design of power and resource management in OFDMA networks using sleep mode selection technique. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619309600.

Hu:2020:PAB

[316] Zhiqun Hu, Hang Qi, Xiangming Wen, Zhaoming Lu, and Wenpeng Jing.

Performance analysis based Markov chain in random access heterogeneous MIMO networks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619305341.

deSouza:2020:HAI

[317] Cristiano Antonio de Souza, Carlos Becker Westphall, Renato Bobsin Machado, João Bosco Mangueira Sobral, and Gustavo dos Santos Vieira. Hybrid approach to intrusion detection in fog-based IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619315439.

Xu:2020:IMA

[318] Jin Xu, Zbigniew Dziong, Yan Luxin, Zhe Huang, Ping Xu, and Adnane Cabani. Intelligent multi-agent based C-RAN architecture for 5G radio resource management. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128619314872.

Ma:2020:EWD

[319] Lisheng Ma, Wei Su, Bin Wu, Bin Yang, and Xiaohong Jiang. Early warning disaster-aware service protection in geo-distributed data centers. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620300074.

Spinnewyn:2020:DCN

[320] Bart Spinnewyn, Steven Latré, and Juan Felipe Botero. Delay-constrained NFV orchestration for heterogeneous cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311099.

Nassar:2020:WBL

[321] Mohamed A. Nassar, Mahmud Hasan, Md Khan, Mirza Sultana, Md Hasan, Len Luxford, Peter Cole, Giles Oatley, and Polychronis Koutsakis. Wifibased localisation datasets for No-GPS open areas using smart bins. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
acompnetamsterdam2020rticle/pii/
S1389128620311117.

Trotta:2020:BDU

[322] Angelo Trotta, Marco Di Felice, Luca Perilli, Eleonora Franchi Scarselli, and Tullio Salmon Cinotti. BEE-DRONES: Ultra low-power monitoring systems based on unmanned aerial vehicles and wake-up radio ground sensors. Computer Networks (Amsterdam, Netherlands: 1999), 180 (??):??, October 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311142.

Anonymous:2020:EBo

[323] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128620311956.

Anonymous:2020:Ob

[324] Anonymous. 24 October 2020. Computer Networks (Amsterdam, Netherlands: 1999), 180(??):??, October 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Aliyu:2020:TMF

[325] Aliyu Lawal Aliyu, Adel Aneiba, Mohammad Patwary, and Peter Bull. A trust management framework for Software Defined Network (SDN) controller and network applications. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311105.

Braeken:2020:SKB

[326] An Braeken. Symmetric key based 5G AKA authentication protocol satisfying anonymity and unlinkability. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): ??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311130.

Karmakar:2020:OLA

[327] Raja Karmakar, Samiran Chattopadhyay, and Sandip Chakraborty. An online learning approach for auto linkconfiguration in IEEE 802.11ac wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311154.

Babamir:2020:MCU

[328] Faezeh Sadat Babamir and Mürvet Kirci. A multibiometric cryptosystem for user authentication in clientserver networks. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311166.

VanRossem:2020:VPM

[329] Steven Van Rossem, Wouter Tavernier, Didier Colle, Mario Pickavet, and Piet Demeester. VNF performance modelling: From stand-alone to chained topologies. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311178.

Martins:2020:VNF

[330] Rafael de Jesus Martins, Cristiano Bonato Both, Juliano Araújo Wickboldt, and Lisandro Zambenedetti Granville. Virtual network functions migration cost: from identification to prediction. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031118X.

Hu:2020:CPM

[331] Chia-Cheng Hu, Jeng-Shyang Pan, Chin-Feng Lai, and Yueh-Min Huang. Content placement for minimizing transmission cost in multiple cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311191.

Chen:2020:SMM

[332] Yuanyuan Chen, Jing Qiu, Xiaojiang Du, Lihua Yin, and Zhihong Tian. Security of mobile multimedia data: The adversarial examples for spatiotemporal data. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031121X.

Bunyakitanon:2020:AAV

[333] Monchai Bunyakitanon, Aloizio Pereira da Silva, Xenofon Vasilakos, Reza Nejabati, and Dimitra Simeonidou. Auto-3P: an autonomous VNF performance prediction and placement framework based on machine learning. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311221.

Bertizzolo:2020:AAS

[334] Lorenzo Bertizzolo, Leonardo Bonati, Emrecan Demirors, Amani Alshawabka, Salvatore D’Oro, Francesco Restuccia, and Tommaso Melodia. Arena: a 64-antenna SDR-based ceiling grid testing platform for sub-6 GHz 5G-and-beyond radio spectrum research. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311257.

Pasdar:2020:HSS

[335] Amirmohammad Pasdar, Young Choon Lee, and Khaled Almi’ani. Hybrid scheduling for scientific workflows on hybrid clouds. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311269.

Bai:2020:ABS

[336] Yang Bai, Li Lu, Jerry Cheng, Jian Liu, Yingying Chen, and Jiadi Yu. Acoustic-based sensing and applications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311282.

Celdran:2020:PTP

[337] Alberto Huertas Celdrán, Manuel Gil Pérez, Izidor Mlakar, Jose M. Alcaraz Calero, Félix J. García Clemente, Gregorio Martínez Pérez, and Zakirul A. Bhuiyan. PROTECTOR: Towards the protection of sensitive data in Europe and the US. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311300.

Yu:2020:PEP

[338] Genghua Yu, Zhigang Chen, Jia Wu, and Jian Wu. Predicted encounter probability based on dynamic programming proposed probability algorithm in opportunistic social network. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311440.

Mahapatra:2020:UAV

[339] Byomakesh Mahapatra, Ashok Kumar Turuk, Sanket Kumar Panda, and Sarat Kumar Patra. Utilization-aware VB migration strategy for inter-BBU load balancing in 5G cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311725.

Ali:2020:MSB

[340] Hamida Qumber Ali and Sayeed Ghani. Multi-sensor based M_{K} /Hyper_{K} /1/ M queuing model for heterogeneous traffic. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311774.

Chen:2020:FNH

[341] Lin Chen, Haipeng Dai, Lei Meng, and Jihong Yu. Finding needles in a hay stream: On persistent item lookup in data streams. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311804.

Park:2020:AAC

[342] Eunjeong Park, Myung-Sup Lee, Hyung-Sin Kim, and Saewoong Bahk. AdaptaBLE: Adaptive control of data rate, transmission power, and connection interval in Bluetooth low energy. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311816.

Gorlatova:2020:CTC

[343] Maria Gorlatova, Hazer Inaltekin, and Mung Chiang. Characterizing task completion latencies in multipoint multi-quality fog computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311853.

Li:2020:TMO

[344] Junfei Li, Penghao Sun, and Yuxiang Hu. Traffic modeling and optimization in datacenters with graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311865.

Boukerche:2020:MLB

[345] Azzedine Boukerche and Jiahao Wang. Machine learning-based traffic prediction models for Intelligent Transportation Systems. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311877.

Jafari-Beyrami:2020:DFA

[346] Mohammad Jafari-Beyrami, Akbar Ghaffarpour Rahbar, and Soheil Hosseini.

On-demand fragmentation-aware spectrum allocation in space division multiplexed elastic optical networks with minimized crosstalk and multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311889.

Alimohammadi:2020:CNW

[347] Hamed Alimohammadi and Mahmood Ahmadi. Common non-wildcard portion-based partitioning approach to SDN many-field packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311890.

Araujo:2020:SCR

[348] Carlos M. Araújo, João Marcos P. Silva, Anand Subramanian, and Iguatemi E. Fonseca. On solving the capacitated routing and spectrum allocation problem for flexgrid optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311907.

Anees:2020:DAE

[349] Junaid Anees, Hao-Chun Zhang, Bachirou Guene Lougou, Sobia Baig, and Yabibal Getahun Dessie. Delay aware energy-efficient opportunistic node selection in restricted routing. Computer Networks (Amsterdam, Netherlands: 1999), 181(??): ??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311919.

Pal:2020:NNF

[350] Amitangshu Pal and Krishna Kant. NFMI: Near Field Magnetic Induction based communication. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312007.

Tanaka:2020:EAL

[351] Kyosuke Tanaka, Hayato Yamaki, Shinobu Miwa, and Hiroki Honda. Evaluating architecture-level optimization in packet processing caches. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312020.

Karatas:2020:MOB

[352] Mumtaz Karatas. A multi-objective bi-level location problem for heterogeneous sensor networks with hubspoke topology. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312019.

Hu:2020:PAS

[353] Guangwu Hu, Qing Li, Shuo Ai, Tan Chen, Jingpu Duan, and Yu Wu. A proactive auto-scaling scheme with latency guarantees for multi-tenant NFV cloud. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312032.

Labayen:2020:OCU

[354] Víctor Labayen, Eduardo Magaña, Daniel Morató, and Mikel Izal. Online classification of user activities using machine learning on network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312081.

Safkhani:2020:IMP

[355] Masoumeh Safkhani, Samad Rostampour, Ygal Bendavid, and Nasour Bagheri. IoT in medical and pharmaceutical: Designing lightweight RFID security protocols for ensuring supply chain integrity. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312068.

Leal:2020:RPS

[356] Alexánder Leal, Mario Durán, and Juan Felipe Botero. Reliability provision in software defined power substations communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312093.

Akkari:2020:CMM

[357] Nadine Akkari and Nikos Dimitriou. Corrigendum to “Mobility Management Solutions for 5G: Architecture and Services” Computer Networks Volume 169 (2020) 107082. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031210X. See

[78].

Vikhrova:2020:GBD

[358] Olga Vikhrova, Sara Pizzi, Antonella Molinaro, Antonio Iera, Konstantin Samouylov, and Giuseppe Araniti. Group-based delivery of critical traffic in cellular IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312111.

Fan:2020:LEO

[359] Wenhao Fan, Junting Han, Le Yao, Fan Wu, and Yuan’an Liu. Latencyenergy optimization for joint WiFi and cellular offloading in mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 181 (??):??, November 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312159.

Anonymous:2020:EBp

[360] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620312433.

Anonymous:2020:N

[361] Anonymous. 9 November 2020. Computer Networks (Amsterdam, Netherlands: 1999), 181(??):??, November 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ayoub:2020:MIS

[362] Wael Ayoub, Abed Ellatif Samhat, Fabienne Nouvel, Mohamad Mroue, Hassan Jradi, and Jean-Christophe Prévotet. Media independent solution for mobility management in heterogeneous LPWAN technologies. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311129.

Zhao:2020:MMA

[363] Xuhui Zhao, Yan Shi, and Shanzhi Chen. MAESP: Mobility aware edge service placement in mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311245.

Begin:2020:DVD

[364] Thomas Begin, Anthony Busson, Isabelle Guérin Lassous, and Azzedine Boukerche. Delivering video-on-demand services with IEEE 802.11p to major non-urban roads: a stochastic performance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311270.

Taleb:2020:FDA

[365] Hussein Taleb, Kinda Khawam, Samer Lahoud, Melhem El Helou, and Steven Martin. A fully distributed approach for joint user association and RRH clustering in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311294.

Zhang:2020:JTO

[366] Ni Zhang, Songtao Guo, Yifan Dong, and Defang Liu. Joint task offloading and data caching in mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311312.

Mishra:2020:SCC

[367] Debashisha Mishra and Enrico Natalizio. A survey on cellular-connected UAVs: Design challenges, enabling 5G/B5G innovations, and experimental advancements. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311324.

Muller:2020:STI

[368] Lucas Müller, Matthew Luckie, Bradley Huffaker, kc claffy, and Marinho Barcellos. Spoofed traffic inference at IXPs: Challenges, methods and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311336.

Khaki:2020:IMM

[369] Mahmood Khaki and Abdorasoul Ghasemi. The impact of mobility model on handover rate in heterogeneous multi-tier wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311348.

Fazio:2020:PMM

[370] Peppino Fazio, Mauro Tropea, Miroslav Voznak, and Floriano De Rango. On packet marking and Markov modeling for IP traceback: a deep probabilistic and stochastic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311439.

Mchergui:2020:BBS

[371] Abir Mchergui, Tarek Moulahi, and Salem Nasri. BaaS: Broadcast as a service cross-layer learningbased approach in cloud assisted VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311464.

Tuysuz:2020:SSG

[372] Mehmet Fatih Tuysuz and Ramona Trestian. From serendipity to sustainable green IoT: Technical, industrial and political perspective. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311452.

Baidas:2020:RAS

[373] Mohammed W. Baidas, Emad Alsusa, and Yao Shi. Resource allocation for SWIPT-enabled energyharvesting downlink/uplink clustered NOMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311476.

Lalouani:2020:MOR

[374] Wassila Lalouani and Mohamed Younis. Multi-observable reputation scoring system for flagging suspicious user sessions. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311506.

Iqbal:2020:DGK

[375] Salman Iqbal, Miss Laiha Mat Kiah, Aqeel ur Rehman, Zahid Abbas, and Babak Daghighi. DM-GKM: a key management scheme for dynamic group based applications. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031152X.

Daldoul:2020:PEO

[376] Yousri Daldoul, Djamal-Eddine Meddour, and Adlen Ksentini. Performance evaluation of OFDMA and MU-MIMO in 802.11ax networks. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311531.

Ullah:2020:UJO

[377] Zaib Ullah, Fadi Al-Turjman, Uzair Moatasim, Leonardo Mostarda, and Roberto Gagliardi. UAVs joint optimization problems and machine learning to improve the 5G and beyond communication. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311518.

Polverini:2020:IDS

[378] Marco Polverini, Jaime Galán-Jiménez, Francesco G. Lavacca, Antonio Cianfrani, and Vincenzo Eramo. Improving dynamic service function chaining classification in NFV/SDN networks through the offloading concept. Computer Networks (Amsterdam, Netherlands: 1999), 182(??): ??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311543.

Ullah:2020:RTD

[379] Noor Ullah, Xiangjie Kong, Limei Lin, Mubarak Alrashoud, Amr Tolba, and Feng Xia. Real-time dissemination of emergency warning messages in 5G enabled selfish vehicular social networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311555.

Boukerche:2020:AIB

[380] Azzedine Boukerche, Yanjie Tao, and Peng Sun. Artificial intelligence-based vehicular traffic flow prediction methods for supporting intelligent transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311567.

Tseng:2020:RBT

[381] Lewis Tseng, Yingjian Wu, Haochen Pan, Moayad Aloqaily, and Azzedine Boukerche. Reliable broadcast with trusted nodes: Energy reduction, resilience, and speed. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311579.

Dehkordi:2020:ENR

[382] Mohammad Jafari Dehkordi and Babak Sadeghiyan. An effective node-removal method against P2P botnets. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311592.

Chouikhi:2020:GNE

[383] Samira Chouikhi, Lyes Khoukhi, Moez Esseghir, and Leila Merghem-Boulahia. Generalized Nash equilibrium approach for radio resource sharing and power allocation in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311580.

Loubany:2020:AAS

[384] Ali Loubany, Samer Lahoud, and Rida El Chall. Adaptive algorithm for spreading factor selection in Lo-RaWAN networks with multiple gateways. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311609.

Pacheco-Paramo:2020:DAD

[385] Diego Pacheco-Paramo and Luis Tello-Oquendo. Delay-aware dynamic access control for mMTC in wireless networks using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311610.

Candela:2020:ICP

[386] Massimo Candela, Valerio Luconi, and Alessio Vecchio. Impact of the COVID-19 pandemic on the Internet latency: a large-scale study. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311622.

Shakarami:2020:SCO

[387] Ali Shakarami, Mostafa Ghobaei-Arani, and Ali Shahidinejad. A survey on the computation offloading approaches in mobile edge computing: a machine learning-based perspective. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311634.

Sabuj:2020:ANO

[388] Saifur Rahman Sabuj, A. M. Musa Shakib Khan, and Masanori Hamamura. Application of non-orthogonal multiple access for machine type communication in sub-terahertz band. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311737.

Shahryari:2020:EED

[389] Om-Kolsoom Shahryari, Hossein Pedram, Vahid Khajehvand, and Mehdi Dehghan TakhtFooladi. Energy-efficient and delay-guaranteed computation offloading for fog-based IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311762.

Mei:2020:RMB

[390] Lifan Mei, Runchen Hu, Houwei Cao, Yong Liu, Zifan Han, Feng Li, and Jin Li. Realtime mobile bandwidth prediction using LSTM neural network and Bayesian fusion. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311798.

Bonati:2020:OPV

[391] Leonardo Bonati, Michele Polese, Salvatore D’Oro, Stefano Basagni, and Tommaso Melodia. Open, programmable, and virtualized 5G networks: State-of-the-art and the road ahead. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311786.

Briones-Reyes:2020:MES

[392] Arturo Briones-Reyes, Luis Alberto Vásquez-Toledo, Alfonso Prieto-Guerrero, and Rafael Aguilar-Gonzalez. Mathematical evaluation of spectrum sharing in cognitive radio networks for 5G systems using Markov processes. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620311828.

Gallego-Madrid:2020:EZE

[393] Jorge Gallego-Madrid, Ramon Sanchez-Iborra, Jose Santa, and Antonio Skarmeta. Evaluation of a zone encryption scheme for vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031183X.

Al-shawabka:2020:MSQ

[394] Amani Al-shawabka, Francesco Restuccia, Salvatore D’Oro, and Tommaso Melodia. Massive-scale I/Q datasets for WiFi radio fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312123.

Roy:2020:AEM

[395] Palash Roy, Sujan Sarker, Md. Abdur Razzaque, Mohammad Mehedi Hassan, Salman A. AlQahtani, Gianluca Aloi, and Giancarlo Fortino. AIenabled mobile multimedia service instance placement scheme in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 182 (??):??, December 9, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312160.

Anonymous:2020:EBq

[396] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128620312779.

Anonymous:2020:Da

[397] Anonymous. 9 December 2020. Computer Networks (Amsterdam, Netherlands: 1999), 182(??):??, December 9, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mashhadi:2020:OAD

[398] Farshad Mashhadi, Sergio A. Salinas Monroy, Arash Bozorgchenani, and Daniele Tarchi. Optimal auction for delay and energy constrained task offloading in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620311841.

Zhang:2020:TAI

[399] Shunliang Zhang and Dali Zhu. Towards artificial intelligence enabled 6G: State of the art, challenges, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031207X.

Hajian:2020:SSH

[400] R. Hajian, S. ZakeriKia, S. H. Erfani, and M. Mirabi. SHAPARAK: Scalable healthcare authentication protocol with attack-resilience and anonymous key-agreement. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312147.

Jiang:2020:CFL

[401] Hui Jiang, Min Liu, Bo Yang, Qingxiang Liu, Jizhong Li, and Xiaobing Guo. Customized federated learning for accelerated edge computing with heterogeneous task targets. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312135.

Marin:2020:SBS

[402] Andrea Marin, Sabina Rossi, and Carlo Zen. Size-based scheduling for TCP flows: Implementation and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312172.

Sun:2020:EFM

[403] Penghao Sun, Julong Lan, Junfei Li, Zehua Guo, Yuxiang Hu, and Tao Hu. Efficient flow migration for NFV with graph-aware deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312184.

Ye:2020:OOE

[404] Weidu Ye, Junzhou Luo, Feng Shan, Wenjia Wu, and Ming Yang. Offspeeding: Optimal energy-efficient flight speed scheduling for UAV-assisted edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312196.

Llerena:2020:SAS

[405] Yarisley Peña Llerena and Paulo R. L. Gondim. Social-aware spectrum sharing for D2D communication by artificial bee colony optimization. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312202.

Liu:2020:PPP

[406] Yuxian Liu, Fagui Liu, Hao-Tian Wu, Xinglin Zhang, Bowen Zhao, and Xingfu Yan. PriDPM: Privacypreserving dynamic pricing mechanism for robust crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312226.

Bernieri:2020:TNA

[407] Giuseppe Bernieri, Stefano Cecconello, Mauro Conti, and Gianluca Lain. TAMBUS: a novel authentication method through covert channels for securing industrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312214.

Redzovic:2020:IPC

[408] Hasan Redzović, Aleksandra Smiljanić, and Mihailo Vesović. Implementation and performance comparison of highcapacity software routers. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312238.

Kuang:2020:SWD

[409] Li Kuang, Huan Zhang, Ruyi Shi, Zhifang Liao, and Xiaoxian Yang. A spam worker detection approach based on heterogeneous network embedding in crowdsourcing platforms. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312251.

Guidotti:2020:ASP

[410] A. Guidotti, S. Cioni, G. Colavolpe, M. Conti, T. Foggi, A. Mengali, G. Montorsi, A. Piemontese, and A. Vanelli-Coralli. Architectures, standardisation, and procedures for 5G satellite communications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): ??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031224X.

Chodak:2020:HLC

[411] Grzegorz Chodak, Grazyna Suchacka, and Yash Chawla. HTTP-level ecommerce data based on server access logs for an online store. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312263.

Shamsoshoara:2020:SPU

[412] Alireza Shamsoshoara, Ashwija Korenda, Fatemeh Afghah, and Sherali Zeadally. A survey on physical unclonable function (PUF)-based security solutions for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312275.

Abdulshakoor:2020:OAM

[413] Ahmed I. Abdulshakoor, M. G. Anany, and Mahmoud M. Elmesalawy. Outage-aware matching game approach for cell selection in LTE/ WLAN multi-RAT HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312317.

Ma:2020:TBE

[414] Zhaofeng Ma, Weizhe Zhao, Shoushan Luo, and Lingyun Wang. TrustedBaaS: Blockchain-enabled distributed and higher-level trusted platform. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620312330.

Calhan:2020:HSS

[415] Ali Çalhan and Murtaza Cicioglu. Handover scheme for 5G small cell networks with non-orthogonal multiple access. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312342.

Dai:2020:JAB

[416] Xiangwen Dai and Jinsong Gui. Joint access and backhaul resource allocation for D2D-assisted dense mmWave cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312354.

Li:2020:IFM

[417] Xiaoqian Li and Kwan L. Yeung. ILP formulations for monitoringcycle design based on segment routing. Computer Networks (Amsterdam, Netherlands: 1999), 183(??): ??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312378.

Sun:2020:PEP

[418] Heli Sun, Shan Zhang, Jianbin Huang, Liang He, Xiaolong Jiang, and Zhangtian Duan. A personalized eventparticipant arrangement framework based on user interests in social network. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031238X.

Mahdizadeh:2020:RFP

[419] Mohammad Saleh Mahdizadeh and Behnam Bahrak. A regression framework for predicting user’s next location using Call Detail Records. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312470.

Hu:2020:SEA

[420] Tao Hu, Peng Yi, Yuxiang Hu, Julong Lan, Zhen Zhang, and Ziyong Li. SAIDE: Efficient application interference detection and elimination in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 183 (??):??, December 24, 2020. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312494.

Anonymous:2020:EBr

[421] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620313116.

Anonymous:2020:Db

[422] Anonymous. 24 December 2020. Computer Networks (Amsterdam, Netherlands: 1999), 183(??):??, December 24, 2020. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2021:SSD

[423] Xiaoquan Zhang, Lin Cui, Kaimin Wei, Fung Po Tso, Yangyang Ji, and Weijia Jia. A survey on stateful data plane in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312305.

Suraci:2021:SOS

[424] Chiara Suraci, Giuseppe Araniti, Andrea Abrardo, Giuseppe Bianchi, and Antonio Iera. A stakeholder-oriented security analysis in virtualized 5G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312366.

Zaki:2021:GGG

[425] Faiz Zaki, Abdullah Gani, Hamid Tahaei, Steven Furnell, and Nor Badrul Anuar. Grano-GT: a granular ground truth collection tool for encrypted browser-based Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 184(??): ??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312482.

Amadeo:2021:DIC

[426] Marica Amadeo, Giuseppe Ruggeri, Claudia Campolo, and Antonella Molinaro. Diversity-improved caching of popular transient contents in Vehicular Named Data Networking. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312548.

Midoglu:2021:LSS

[427] Cise Midoglu, Konstantinos Kousias, Özgü Alay, Andra Lutu, Antonios Argyriou, Michael Riegler, and Carsten Griwodz. Large scale “speedtest” experimentation in Mobile Broadband Networks. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312561.

Abdelsalam:2021:TWL

[428] A. Abdelsalam, M. Luglio, N. Patriciello, C. Roseti, and F. Zampognaro. TCP Wave over Linux: a disruptive alternative to the traditional TCP window approach. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312585.

Sviridov:2021:LDR

[429] German Sviridov, Marco Bonola, Angelo Tulumello, Paolo Giaccone, Andrea Bianco, and Giuseppe Bianchi. LOcAl DEcisions on Replicated States (LOADER) in programmable dataplanes: Programming abstraction and experimental evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312597.

Bai:2021:CAF

[430] Yude Bai, Zhenchang Xing, Duoyuan Ma, Xiaohong Li, and Zhiyong Feng. Comparative analysis of feature representations and machine learning methods in Android family classification. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312603.

Zeng:2021:FDD

[431] Yuwei Zeng, Xiaochun Yun, Xunxun Chen, Boquan Li, Haiwei Tsang, Yipeng Wang, Tianning Zang, and Yongzheng Zhang. Finding disposable domain names: a linguisticsbased stacking approach. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312627.

Rottondi:2021:SET

[432] Cristina Rottondi, Francesco Malandrino, Andrea Bianco, Carla Fabiana Chiasserini, and Ioannis Stavrakakis. Scheduling of emergency tasks for multiservice UAVs in post-disaster scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312639.

Balci:2021:MCM

[433] Abdullah Balci and Radosveta Sokullu. Massive connectivity with machine learning for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1389128620312652.

Qureshi:2021:ADT

[434] Kashif Naseer Qureshi, Gwanggil Jeon, and Francesco Piccialli. Anomaly detection and trust authority in artificial intelligence and cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312664.

Alonso:2021:MOO

[435] Rodney Martinez Alonso, David Plets, Margot Deruyck, Luc Martens, Glauco Guillen Nieto, and Wout Joseph. Multi-objective optimization of cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031269X.

Chen:2021:LSB

[436] Ning Chen, Sheng Zhang, Jie Wu, Zhuzhong Qian, and Sanglu Lu. Learning scheduling bursty requests in Mobile Edge Computing using DeepLoad. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312706.

Mahbub:2021:CEI

[437] Mobasshir Mahbub, M. Mofazzal Hossain, and Md. Shamrat Apu Gazi. Cloud-enabled IoT-based embedded system and software for intelligent indoor lighting, ventilation, early stage fire detection and prevention. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312858.

Quincozes:2021:SID

[438] Silvio E. Quincozes, Célio Albuquerque, Diego Passos, and Daniel Mossé. A survey on intrusion detection and prevention systems in digital substations. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312895.

Chao:2021:LAA

[439] Chih-Min Chao and Wei-Che Lee. Load-aware anti-jamming channel hopping design for cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312901.

Zeb:2021:JPS

[440] Jehan Zeb, Aamir Hassan, and Muhammad Danish Nisar. Joint power and spectrum allocation for D2D communication overlaying cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312913.

Zeng:2021:SCM

[441] Yue Zeng, Baoliu Ye, Bin Tang, Songtao Guo, and Zhihao Qu. Scheduling coflows of multi-stage jobs under network resource constraints. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312937.

Yazdinejadna:2021:KBI

[442] Abbas Yazdinejadna, Reza M. Parizi, Ali Dehghantanha, and Mohammad S. Khan. A kangaroo-based intrusion detection system on software-defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312949.

Palumbo:2021:CAC

[443] Fabio Palumbo, Giuseppe Aceto, Alessio Botta, Domenico Ciuonzo, Valerio Persico, and Antonio Pescapé.

Characterization and analysis of cloudto-user latency: the case of Azure and AWS. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312962.

Liu:2021:TAS

[444] Xiuwen Liu, Jianming Fu, Yanjiao Chen, Weichen Luo, and Zihan Tang. Trust-aware sensing quality estimation for team crowdsourcing in social IoT. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312986.

Pushpalatha:2021:RRP

[445] M. Pushpalatha, T. Anusha, T. Rama Rao, and Revathi Venkataraman. L-RPL: RPL powered by Laplacian energy for stable path selection during link failures in an Internet of Things network. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312998.

Ibrar:2021:PFM

[446] Muhammad Ibrar, Lei Wang, Gabriel-Miro Muntean, Aamir Akbar, Nadir Shah, and Kaleem Razzaq Malik. PrePass-Flow: a machine learning based technique to minimize ACL policy violation due to links failure in hybrid SDN. Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313025.

Adil:2021:CFO

[447] Muhammad Adil. Congestion free opportunistic multipath routing load balancing scheme for Internet of Things (IoT). Computer Networks (Amsterdam, Netherlands: 1999), 184 (??):??, January 15, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313049.

Anonymous:2021:EBa

[448] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620313517.

Anonymous:2021:Ja

[449] Anonymous. 15 January 2021. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):??, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jha:2021:LBS

[450] Rakesh Kumar Jha, Puja, Haneet Kour, Manoj Kumar, and Shubha Jain. Layer based security in Narrow Band Internet of Things (NB-IoT). Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312299.

Bhooanusas:2021:SBD

[451] Nuntanut Bhooanusas, Sok-Ian Sou, and Kai-Chun Cheng. Satisfactionbased dynamic bandwidth reallocation for multipath mobile data offloading. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312287.

Asari:2021:NPH

[452] Amirhossein Asari, Mahdi R. Alagheband, Majid Bayat, and Maryam Rajabzadeh Asaar. A new provable hierarchical anonymous certificateless authentication protocol with aggregate verification in ADS-B systems. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312329.

Wang:2021:IUB

[453] Xiaoding Wang, Sahil Garg, Hui Lin, Georges Kaddoum, Jia Hu, and Mohammed F. Alhamid. An intelligent UAV based data aggregation algorithm for 5G-enabled Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031255X.

Wang:2021:ICR

[454] Xu Wang, Yanjiao Chen, and Qian Zhang. Incentivizing cooperative relay in UTXO-based blockchain network. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312573.

Afzal:2021:MMB

[455] Samira Afzal, Christian Esteve Rothenberg, Vanessa Testoni, Prakash Kolan, and Imed Bouazizi. Multipath MMTbased approach for streaming high quality video over multiple wireless access networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312615.

Huang:2021:MCA

[456] Lisheng Huang, Jinye Ran, Wenyong Wang, Tan Yang, and Yu Xiang. A multi-channel anomaly detection method with feature selection and multi-scale analysis. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312640.

Ndashimye:2021:MCB

[457] Emmanuel Ndashimye, Nurul I. Sarkar, and Sayan Kumar Ray. A multicriteria based handover algorithm for vehicle-to-infrastructure communications. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312688.

Ghosal:2021:TPS

[458] Amrita Ghosal, Sang Uk Sagong, Subir Halder, Kalana Sahabandu, Mauro Conti, Radha Poovendran, and Linda Bushnell. Truck platoon security: State-of-the-art and road ahead. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862031272X.

Eksert:2021:IIC

[459] M. Levent Eksert, Hamdullah Yücel, and Ertan Onur. Intra- and intercluster link scheduling in CUPS-based ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620312718.

Cao:2021:TIV

[460] Haotong Cao, Yue Hu, and Longxiang Yang. Towards intelligent virtual resource allocation in UAVs-assisted 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312731.

Geng:2021:FLE

[461] Nan Geng, Yuan Yang, and Mingwei Xu. Flow-level and efficient traffic engineering in conventional routing systems. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312834.

Shehzad:2021:PAC

[462] Khurram Shehzad, Noor M. Khan, and Junaid Ahmed. Performance analysis of coverage-centric heterogeneous cellular networks using dualslope path loss model. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312846.

Noura:2021:ERD

[463] Hassan N. Noura, Reem Melki, Ali Chehab, and Javier Hernandez Fernandez. Efficient and robust data availability solution for hybrid PLC/ RF systems. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031286X.

Sheng:2021:CPM

[464] Chuan Sheng, Yu Yao, Qiang Fu, and Wei Yang. A cyber-physical model for SCADA system and its intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312883.

Shadroo:2021:TPS

[465] Shabnam Shadroo, Amir Masoud Rahmani, and Ali Rezaee. The two-phase scheduling based on deep learning in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312925.

Park:2021:FFB

[466] Taejune Park, Seungwon Shin, Insik Shin, and Kilho Lee. Formullar: an FPGA-based network testing tool for flexible and precise measurement of ultra-low latency networking systems. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312950.

Khormali:2021:DNS

[467] Aminollah Khormali, Jeman Park, Hisham Alasmary, Afsah Anwar, Muhammad Saad, and David Mohaisen. Domain name system security and privacy: a contemporary survey. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313001. See

corrigendum [527].

Yemeni:2021:RST

[468] Zaid Yemeni, Haibin Wang, Waleed M. Ismael, Yanan Wang, and Zhengming Chen. Reliable spatial and temporal data redundancy reduction approach for WSN. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313013.

Mayer:2021:PMH

[469] Andrea Mayer, Pierpaolo Loreti, Lorenzo Bracciale, Paolo Lungaroni, Stefano Salsano, and Clarence Filsfils. Performance monitoring with H^{2}: Hybrid Kernel/eBPF data plane for SRv6 based hybrid SDN. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313037.

Sun:2021:QGI

[470] Weifeng Sun, Zun Wang, and Guanghao Zhang. A QoS-guaranteed intelligent routing mechanism in softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313050.

Moll:2021:QBS

[471] Philipp Moll, Selina Isak, Hermann Hellwagner, and Jeff Burke. A quadtree-based synchronization protocol for inter-server game state synchronization. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313177.

Schardong:2021:NRA

[472] Frederico Schardong, Ingrid Nunes, and Alberto Schaeffer-Filho. NFV resource allocation: a systematic review and taxonomy of VNF forwarding graph embedding. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313189.

Tahmasebi:2021:SEM

[473] Shirin Tahmasebi, Nayereh Rasouli, Amir Hosein Kashefi, Elmira Rezabeyk, and Hamid Reza Faragardi. SYNCOP: an evolutionary multiobjective placement of SDN controllers for optimizing cost and network performance in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313190.

Terzi:2021:GWD

[474] Caglar Terzi and Ibrahim Korpeoglu. 60 GHz wireless data center networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313207.

Ali:2021:CRS

[475] Zeeshan Ali, Shehzad Ashraf Chaudhry, Khalid Mahmood, Sahil Garg, Zhihan Lv, and Yousaf Bin Zikria. A clogging resistant secure authentication scheme for fog computing services. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313219.

Yi:2021:SSA

[476] Peng Yi, Tao Hu, Yuxiang Hu, Julong Lan, Zhen Zhang, and Ziyong Li. SQHCP: Secure-aware and QoSguaranteed heterogeneous controller placement for software-defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 185 (??):??, February 11, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313268.

Wu:2021:DLP

[477] Yulei Wu, Yuxiang Ma, Hong-Ning Dai, and Hao Wang. Deep learning for privacy preservation in autonomous moving platforms enhanced 5G heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031327X.

Amorim:2021:ACO

[478] Kelvin Santos Amorim and Gustavo Sousa Pavani. Ant colony optimization-based distributed multilayer routing and restoration in IP/ MPLS over optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031330X.

Anonymous:2021:EBb

[479] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128620313682.

Anonymous:2021:Fa

[480] Anonymous. 11 February 2021. Computer Networks (Amsterdam, Netherlands: 1999), 185(??):??, February 11, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sekaran:2021:ISS

[481] Ramesh Sekaran, Surya Narayana Goddumarri, Suresh Kallam, Manikandan Ramachandran, Rizwan Patan, and Deepak Gupta. 5G integrated spectrum selection and spectrum access using AI-based frame work for IoT based sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312676.

Haile:2021:EEC

[482] Habtegebreil Haile, Karl-Johan Grinnemo, Simone Ferlin, Per Hurtig, and Anna Brunstrom. End-to-end congestion control approaches for high throughput and low delay in 4G/5G cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620312974.

AlGhadhban:2021:FFD

[483] Amer AlGhadhban. F4Tele: FSO for data center network management and packet telemetry. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313062.

Fujita:2021:SBR

[484] Risa Fujita, Fujun He, and Eiji Oki. Shared backup resource assignment for middleboxes considering server protection capabilities. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313220.

Triantafyllou:2021:LFL

[485] Anna Triantafyllou, Panagiotis Sarigiannidis, Thomas Lagkas, Ioannis D. Moscholios, and Antonios Sarigiannidis. Leveraging fairness in LoRaWAN: a novel scheduling scheme for collision avoidance. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313232.

Balasubramanian:2021:SAT

[486] Venkatraman Balasubramanian, Moayad Aloqaily, and Martin Reisslein. An SDN architecture for time sensitive industrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313256.

Chen:2021:CNN

[487] Chen Chen, Qiang Hui, Wenxuan Xie, Shaohua Wan, Yang Zhou, and Qingqi Pei. Convolutional neural networks for forecasting flood process in Internet-of-Things enabled smart city. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313281.

Ramirez:2021:ING

[488] Pedro Luis González Ramírez, Jaime Lloret, Jesús Tomás, and Mikel Hurtado. IoT-networks group-based model that uses AI for workgroup allocation. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313293.

Neto:2021:SMM

[489] Emidio P. Neto, Felipe S. Dantas Silva, Lucas M. Schneider, Augusto V. Neto, and Roger Immich. Seamless MANO of multi-vendor SDN controllers across federated multi-domains. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313311.

Choudhury:2021:HLS

[490] Hiten Choudhury. HashXor: a lightweight scheme for identity privacy of IoT devices in 5G mobile network. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313335.

Majumdar:2021:NBS

[491] Prasanta Majumdar and Tanmay De. A non-backtracking spectrum allocation algorithm in a large EON under dynamic traffic grooming. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313359.

Jradi:2021:OMR

[492] Hassan Jradi, Abed Ellatif Samhat, Fabienne Nouvel, Mohamad Mroue, and Jean-Christophe Prévotet. Overview of the mobility related security challenges in LPWANs. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313384.

Tan:2021:BNT

[493] Lizhuang Tan, Wei Su, Wei Zhang, Jianhui Lv, Zhenyi Zhang, Jingying Miao, Xiaoxi Liu, and Na Li. In-band network telemetry: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313396.

Talati:2021:ADO

[494] Shreya Talati, Darshan Vekaria, Aparna Kumari, and Sudeep Tanwar. An AI-driven object segmentation and speed control scheme for autonomous moving platforms. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313566.

Ge:2021:TDL

[495] Mengmeng Ge, Naeem Firdous Syed, Xiping Fu, Zubair Baig, and Antonio Robles-Kelly. Towards a deep learningdriven intrusion detection approach for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031358X.

Malboubi:2021:PFI

[496] Mehdi Malboubi, Abhijeet Bhorkar, and Frank Jiang. PAveMENT: a framework for the intelligent and safe navigation of unmanned aerial vehicles over optimal PAths in MobilE NeTworks. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313591.

Rios:2021:DRQ

[497] Vinícius de Miranda Rios, Pedro R. M. Inácio, Damien Magoni, and Mário M. Freire. Detection of reduction-of-quality DDoS attacks using fuzzy logic and machine learning algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128620313633.

Li:2021:CLB

[498] Chunlin Li, Mingyang Song, Qingchuan Zhang, and Youlong Luo. Cluster load based content distribution and speculative execution for geographically distributed cloud environment. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000025.

Dayal:2021:SWF

[499] Neelam Dayal and Shashank Srivastava. SD-WAN flood tracer: Tracking the entry points of DDoS attack flows in WAN. Computer Networks (Amsterdam, Netherlands: 1999), 186 (??):??, February 26, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000050.

Anonymous:2021:EBc

[500] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621000293.

Anonymous:2021:Fb

[501] Anonymous. 26 February 2021. Computer Networks (Amsterdam, Netherlands: 1999), 186(??):??, February 26, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Walelgne:2021:CPD

[502] Ermias Andargie Walelgne, Alemnew Sheferaw Asrese, Jukka Manner, Vaibhav Bajpai, and Jörg Ott. Clustering and predicting the data usage patterns of geographically diverse mobile users. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313244.

Navas:2021:PRI

[503] Renzo E. Navas, Frédéric Cuppens, Nora Boulahia Cuppens, Laurent Toutain, and Georgios Z. Papadopoulos. Physical resilience to insider attacks in IoT networks: Independent cryptographically secure sequences for DSSS anti-jamming. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313323.

Rezaee:2021:QTA

[504] Arash Rezaee, Omid Akbari Sheikhabad, and Lotfollah Beygi. Quality of transmission-aware control plane performance analysis for elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313347.

Yang:2021:DAI

[505] Jiapie Yang, Prosanta Gope, Yongqiang Cheng, and Li Sun. Design, analysis and implementation of a smart next generation secure shipping infrastructure using autonomous robot. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313554.

Shi:2021:PFF

[506] Zhou Shi, Ye Tian, Xiaodong Wang, Jiangyu Pan, and Xinming Zhang. Po-fi: Facilitating innovations on WiFi networks with an SDN approach. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313578.

Xue:2021:RCD

[507] Gang Xue, Shuiguang Deng, Di Liu, and Zeming Yan. Reaching consensus in decentralized coordination of distributed microservices. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313608.

Liu:2021:CBF

[508] Qingzhi Liu, Long Cheng, Renan Alves, Tanir Ozcelebi, Fernando Kuipers, Guixian Xu, Johan Lukkien, and Shanzhi Chen. Cluster-based flow control in hybrid software-defined wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862031361X.

Wang:2021:NPM

[509] Jiahao Wang and Azzedine Boukerche. Non-parametric models with optimized training strategy for vehicles traffic flow prediction. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313621.

Toumi:2021:CDS

[510] Nassima Toumi, Olivier Bernier, Djamal-Eddine Meddour, and Adlen Ksentini. On cross-domain Service Function Chain orchestration: an architectural framework. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000013.

Malik:2021:MNC

[511] Hirah Malik, Cédric Adjih, Claudio Weidmann, and Michel Kieffer. MICN: a network coding protocol for ICN with multiple distinct interests per generation. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000074.

Kumar:2021:SSP

[512] Randhir Kumar, Prabhat Kumar, Rakesh Tripathi, Govind P. Gupta, Thippa Reddy Gadekallu, and Gautam Srivastava. SP2F: a secured privacy-preserving framework for smart agricultural unmanned aerial vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000086.

Hassani:2021:QPA

[513] Hamid Hassani, Francesco Gringoli, and Douglas J. Leith. Quick and plenty: Achieving low delay and high rate in 802.11ac edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000098.

Thomdapu:2021:DCM

[514] Srujan Teja Thomdapu, Palash Katiyar, and Ketan Rajawat. Dynamic cache management in content delivery networks. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000116.

Galan-Jimenez:2021:MII

[515] Jaime Galán-Jiménez and Luca Chiaraviglio. Measuring the impact of IC-NIRP vs. stricter-than-ICNIRP exposure limits on QoS and EMF from cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000128.

Shah:2021:IMI

[516] Sayed Qaiser Ali Shah, Farrukh Zeeshan Khan, and Muneer Ahmad. The impact and mitigation of ICMP based economic denial of sustainability attack in cloud computing environment using software defined network. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100013X.

Al-Makhadmeh:2021:ITN

[517] Zafer Al-Makhadmeh and Amr Tolba. Independent and tailored networkslicing architecture for leveraging industrial Internet of Things job processing. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000153.

Morawski:2021:GMT

[518] Michał Morawski and Przemysław Ignaciuk. A green multipath TCP framework for industrial Internet of Things applications. Computer Networks (Amsterdam, Netherlands: 1999), 187 (??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000189.

Nikkhah:2021:LLA

[519] Fahimeh Nikkhah and Masoumeh Safkhani. LAPCHS: a lightweight authentication protocol for cloudbased health-care systems. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000190.

Zhao:2021:IFB

[520] Liang Zhao, Zhenguo Bi, Mingwei Lin, Ammar Hawbani, Junling Shi, and Yunchong Guan. An intelligent fuzzybased routing scheme for softwaredefined vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000219.

Deng:2021:SBD

[521] Yanlin Deng, Quanzhong Li, Qi Zhang, Liang Yang, and Jiayin Qin. Secure beamforming design in MIMO NOMA networks for Internet of Things with perfect and imperfect CSI. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000220.

Anonymous:2021:EBd

[522] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621000505.

Anonymous:2021:Ma

[523] Anonymous. 14 March 2021. Computer Networks (Amsterdam, Netherlands: 1999), 187(??):??, March 14, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Touijer:2021:ICC

[524] Bethaina Touijer, Yann Ben Maissa, and Salma Mouline. IEEE 802.15.6 CSMA/CA access method for WBANs:

Performance evaluation and new backoff counter selection procedure. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128620313372.

Kazemifard:2021:MDF

[525] Nasim Kazemifard and Vahid Shah-Mansouri. Minimum delay function placement and resource allocation for Open RAN (O-RAN) 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000037.

Hendaoui:2021:UUA

[526] Fatma Hendaoui, Hamdi Eltaief, and Habib Youssef. UAP: a unified authentication platform for IoT environment. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000049.

Khormali:2021:CDN

[527] Aminollah Khormali, Jeman Park, Hisham Alasmary, Afsah Anwar, Muhammad Saad, and David Mohaisen. Corrigendum to “Domain name system security and privacy: a contemporary survey” Computer Networks Volume 185 (2020) 107699. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000062. See

[467].

Zhang:2021:RIS

[528] Yu Zhang, Bo Wang, Bingli Guo, Yabo Yuan, Tao Dong, Jie Yin, Kexin Li, Xiang Guo, and Shanguo Huang. A research on integrated space-ground information network simulation platform based on SDN. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000104.

Salhab:2021:NSR

[529] Nazih Salhab, Rami Langar, and Rana Rahim. 5G network slices resource orchestration using machine learning techniques. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000165.

Golkarifard:2021:DVP

[530] Morteza Golkarifard, Carla Fabiana Chiasserini, Francesco Malandrino, and Ali Movaghar. Dynamic VNF placement, resource allocation and traffic routing in 5G. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000177.

Moriano:2021:UBA

[531] Pablo Moriano, Raquel Hill, and L. Jean Camp. Using bursty announcements for detecting BGP routing anomalies. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000207.

Kilincer:2021:MLM

[532] Ilhan Firat Kilincer, Fatih Ertam, and Abdulkadir Sengur. Machine learning methods for cyber security intrusion detection: Datasets and comparative study. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000141.

Yang:2021:DSE

[533] Lei Yang, Liang Liu, Zuchao Ma, and Youwei Ding. Detection of selectiveedge packet attack based on edge reputation in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000232.

Gao:2021:OCE

[534] Peixuan Gao, Yang Xu, and H. Jonathan Chao. OVS-CAB: Efficient rulecaching for Open vSwitch hardware offloading. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000244.

Xu:2021:DBW

[535] Chuan Xu, Giovanni Neglia, and Nicola Sebastianelli. Dynamic backup workers for parallel machine learning. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000256.

Volnes:2021:IAT

[536] Espen Volnes, Stein Kristiansen, and Thomas Plagemann. Improving the accuracy of timing in scalable WSN simulations with communication software execution models. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000335.

Zhang:2021:PBS

[537] Tianzhu Zhang, Leonardo Linguaglossa, Paolo Giaccone, Luigi Iannone, and James Roberts. Performance benchmarking of state-of-the-art software switches for NFV. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000372.

Faroughi:2021:TWD

[538] Azadeh Faroughi, Andrea Morichetta, Luca Vassio, Flavio Figueiredo, Marco Mellia, and Reza Javidan. Towards website domain name classification using graph based semi-supervised learning. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000384.

Benedetti:2021:SME

[539] Paolo Benedetti, Giuseppe Piro, and Luigi Alfredo Grieco. A softwarized and MEC-enabled protocol architecture supporting consumer mobility in information-centric networks. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000396.

Doost-Mohammady:2021:GTW

[540] Rahman Doost-Mohammady, Oscar Bejarano, and Ashutosh Sabharwal. Good times for wireless research. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000414.

Amaizu:2021:CED

[541] G. C. Amaizu, C. I. Nwakanma, S. Bhardwaj, J. M. Lee, and D. S. Kim. Composite and efficient DDoS attack detection framework for B5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000438.

Wang:2021:DFG

[542] Tao Wang, Li Kang, and Jiang Duan. Dynamic fine-grained access control scheme for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000402.

Kumar:2021:OAO

[543] Rohit Kumar, Venkanna U., and Vivek Tiwari. Opt-ACM: an optimized load balancing based admission control mechanism for software defined hybrid wireless based IoT (SDHW-IoT) network. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000542.

Derakhshanfard:2021:ORW

[544] Nahideh Derakhshanfard and Reza Soltani. Opportunistic routing in wireless networks using bitmap-based weighted tree. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100058X.

Takeda:2021:MPS

[545] Kenta Takeda, Takehiro Sato, Ryoichi Shinkuma, and Eiji Oki. Multipath provisioning scheme for fault tolerance to minimize required spectrum resources in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 188 (??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000591.

Anonymous:2021:EBe

[546] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000980.

Anonymous:2021:Aa

[547] Anonymous. 7 April 2021. Computer Networks (Amsterdam, Netherlands: 1999), 188(??):??, April 7, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Farrokhi:2021:AIT

[548] Alireza Farrokhi, Reza Farahbakhsh, Javad Rezazadeh, and Roberto Minerva. Application of Internet of Things and artificial intelligence for smart fitness: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000360.

Lin:2021:NLG

[549] Jie Lin, Lin Huang, Hanlin Zhang, Xinyu Yang, and Peng Zhao. A novel latency-guaranteed based resource double auction for marketoriented edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000426.

Huang:2021:ITC

[550] Xiaohong Huang, Man Zeng, and Kun Xie. Intelligent traffic control for QoS optimization in hybrid SDNs. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000451.

Fawaz:2021:RLA

[551] Hassan Fawaz, Melhem El Helou, Samer Lahoud, and Kinda Khawam. A reinforcement learning approach to queue-aware scheduling in fullduplex wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000578.

Wu:2021:AMS

[552] Libing Wu, Shuqin Cao, Yanjiao Chen, Jianqun Cui, and Yanan Chang. An adaptive multiple spray-and-wait routing algorithm based on social circles in delay tolerant networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000633.

Rahdari:2021:QAP

[553] Farhad Rahdari, Naser Movahhedinia, Mohammad Reza Khayyambashi, and Shahrokh Valaee. QoEaware power control and user grouping in cognitive radio OFDM-NOMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000657.

Zakariya:2021:ODM

[554] Ahmed Y. Zakariya, Sherif I. Rabia, and W. K. Zahra. Optimal decision making in multi-channel RF-powered cognitive radio networks with ambient backscatter capability. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000669.

Morichetta:2021:UWP

[555] Andrea Morichetta, Martino Trevisan, Luca Vassio, and Julia Krickl. Understanding web pornography usage from traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000670.

Ahmed:2021:MLM

[556] Usman Ahmed, Gautam Srivastava, and Jerry Chun-Wei Lin. A machine learning model for data sanitization. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000700.

Chuprikov:2021:FTC

[557] Pavel Chuprikov, Alex Davydow, Kirill Kogan, Sergey I. Nikolenko, and Alexander Sirotkin. Formalization and taxonomy of compute-aggregate problems for cloud computing applications. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000694.

Zhong:2021:CSC

[558] Shijie Zhong, Songtao Guo, Hongyan Yu, and Quyuan Wang. Cooperative service caching and computation offloading in multi-access edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000566.

Baidas:2021:RAO

[559] Mohammed W. Baidas. Resource allocation for offloading-efficiency maximization in clustered NOMAenabled mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000712.

Martinez:2021:ESE

[560] Cristiane A. Pendeza Martinez, Taufik Abrão, and André Luís Machado Martinez. Energy and spectral efficiency trade-off in OCDMA-PON assisted by non-linear programming methods. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000724.

Feng:2021:DRC

[561] Yue Feng, Weiqing Huang, Siye Wang, Yanfang Zhang, and Shang Jiang. Detection of RFID cloning attacks: a spatiotemporal trajectory data streambased practical approach. Computer Networks (Amsterdam, Netherlands:

1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000736.

Venkateswararao:2021:UUV

[562] Kuna Venkateswararao, Pravati Swain, Christophoros Christophorou, and Andreas Pitsillides. Using UE-VBS for dynamic virtual small cells deployment and backhauling in 5G ultradense networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000748.

Biswas:2021:LBU

[563] Soumadip Biswas, Arobinda Gupta, and Sandip Chakraborty. Loadbalanced user associations in dense LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000761.

Lu:2021:ECD

[564] Ning Lu, Dan Li, Wenbo Shi, Pandi Vijayakumar, Francesco Piccialli, and Victor Chang. An efficient combined deep neural network based malware detection framework in 5G environment. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000785.

Enoch:2021:NSM

[565] Simon Yusuf Enoch, Jang Se Lee, and Dong Seong Kim. Novel security models, metrics and security assessment for maritime vessel networks. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000797.

Shang:2021:DUA

[566] Longkang Shang, Dong Guo, Yuede Ji, and Qiang Li. Discovering unknown advanced persistent threat using shared features mined by neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000803.

Davydow:2021:CBM

[567] Alex Davydow, Pavel Chuprikov, Sergey I. Nikolenko, and Kirill Kogan. Competitive buffer management for packets with latency constraints. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000839.

Modeas:2021:ADN

[568] Ioannis Modeas, Alexandros Kaloxylos, Lazaros Merakos, and Dimitris Tsolkas. An adaptive and distributed network selection mechanism for 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 189 (??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000840.

Anonymous:2021:EBf

[569] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621001341.

Anonymous:2021:Ab

[570] Anonymous. 22 April 2021. Computer Networks (Amsterdam, Netherlands: 1999), 189(??):??, April 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sun:2021:SSD

[571] Penghao Sun, Zehua Guo, Julong Lan, Junfei Li, Yuxiang Hu, and Thar Baker. ScaleDRL: a scalable deep reinforcement learning approach for traffic engineering in SDN with pinning control. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000554.

Li:2021:ADD

[572] Huichun Li, Chengli Zhao, Yangyang Liu, and Xue Zhang. Anomaly detection by discovering bipartite structure on complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000621.

Huang:2021:SJR

[573] Guanglun Huang, Baoxian Zhang, Zheng Yao, and Cheng Li. Stochastic joint rate control and resource allocation for wireless video surveillance. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000645.

Chen:2021:MEC

[574] Xincheng Chen. Mobile edge computing resource allocation: an operating system view. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100075X.

Strinati:2021:NBS

[575] Emilio Calvanese Strinati and Sergio Barbarossa. 6G networks: Beyond Shannon towards semantic and goaloriented communications. Computer Networks (Amsterdam, Netherlands:

1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000773.

Shaer:2021:EEP

[576] Ibrahim Shaer, Greg Sidebottom, Anwar Haque, and Abdallah Shami. Efficient execution plan for egress traffic engineering. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000815.

Lalle:2021:CTS

[577] Yandja Lalle, Mohamed Fourati, Lamia Chaari Fourati, and João Paulo Barraca. Communication technologies for Smart Water Grid applications: Overview, opportunities, and research directions. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000827.

Sundberg:2021:LET

[578] Simon Sundberg and Johan Garcia. Locating eNodeBs through sectorization inference-sector fitting evaluated on a railway use case. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000852.

Adebayo:2021:PEU

[579] Adekunle Adebayo, Ibhadode Osagie, Adekunle Oluwaseun Tope, and Ibhadode Akhere Pauline. Performance evaluation of ultra-dense macrofemto cellular networks based on efficient channel resource allocation system. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000876.

Femminella:2021:EDT

[580] Mauro Femminella, Matteo Pergolesi, and Gianluca Reali. 5G experiment design through Blueprint. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000864.

Akbar:2021:NET

[581] Aamina Akbar, Sobia Jangsher, and Farrukh A. Bhatti. NOMA and 5G emerging technologies: a survey on issues and solution techniques. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000888.

Chen:2021:ATM

[582] Guozhu Chen, Fanping Zeng, Jian Zhang, Tingting Lu, Jingfei Shen, and Wenjuan Shu. An adaptive trust model based on recommendation filtering algorithm for the Internet of Things systems. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100089X.

Li:2021:DAD

[583] Yu Li and Wenjian Xu. D-AdFeed: a diversity-aware utility-maximizing advertising framework for mobile users. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000906.

Alqahtani:2021:ECR

[584] Fayez Alqahtani, Mohammed Al-Maitah, Khaldoun Besoul, and S. K. Elagan. Elastic computing resource virtualization method for a service-centric industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000918.

Dong:2021:SLT

[585] Jialiang Dong, Zhitao Guan, Longfei Wu, Xiaojiang Du, and Mohsen Guizani. A sentence-level text adversarial attack algorithm against IIoT based smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100092X.

Silva:2021:HSS

[586] Rui Silva, David Santos, Flávio Meneses, Daniel Corujo, and Rui L. Aguiar. A hybrid SDN solution for mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621000931.

Wang:2021:DJS

[587] Libing Wang, Xin Hu, Yin Wang, Sujie Xu, Shijun Ma, Kexin Yang, Zhijun Liu, and Weidong Wang. Dynamic job-shop scheduling in smart manufacturing using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001031.

Velusamy:2021:FTA

[588] Nandagopal Velusamy, Fadi Al-Turjman, Rajagopal Kumar, and Jothilakshmi Ramakrishnan. A framework for task allocation in IoToriented industrial manufacturing systems. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001043.

Lin:2021:TNC

[589] Kunda Lin, Xiaolong Xu, and Honghao Gao. TSCRNN: a novel classification scheme of encrypted traffic based on flow spatiotemporal features for efficient management of IIoT. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001067.

Attarian:2021:ACP

[590] Reyhane Attarian and Sattar Hashemi. An anonymity communication protocol for security and privacy of clients in IoT-based mobile health transactions. Computer Networks (Amsterdam, Netherlands: 1999), 190 (??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001079.

Anonymous:2021:EBg

[591] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001602.

Anonymous:2021:Mb

[592] Anonymous. 8 May 2021. Computer Networks (Amsterdam, Netherlands: 1999), 190(??):??, May 8, 2021.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zaballa:2021:TMH

[593] Eder Ollora Zaballa, David Franco, Signe Erdman Thomsen, Marivi Higuero, Henrik Wessing, and Michael S. Berger. Towards monitoring hybrid next-generation software-defined and service provider MPLS networks. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000943.

Cominelli:2021:ICR

[594] Marco Cominelli, Felix Kosterhon, Francesco Gringoli, Renato Lo Cigno, and Arash Asadi. IEEE 802.11 CSI randomization to preserve location privacy: an empirical evaluation in different scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100102X.

Ming:2021:NNB

[595] Zhongxing Ming and Mingwei Xu. NBA: a name-based approach to device mobility in Industrial IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001055.

Wen:2021:ACB

[596] Yujuan Wen, Fengyuan Lu, Yufei Liu, and Xinli Huang. Attacks and countermeasures on blockchains: a survey from layering perspective. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001080.

Martinez-Yelmo:2021:EEH

[597] Isaias Martinez-Yelmo, Joaquin Alvarez-Horcajo, Juan Antonio Carral, and Diego Lopez-Pajares. eHDDP: Enhanced Hybrid Domain Discovery Protocol for network topologies with both wired/wireless and SDN/non-SDN devices. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001110.

Zhang:2021:ELS

[598] Guoqiang Zhang, Chenyi Wu, and Qian Gao. Exploiting layer and spatial correlations to enhance SVC and tile based 360-degree video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001122.

Vasudev:2021:SPP

[599] Harsha Vasudev and Debasis Das. P^{2}-SHARP: Privacy preserving secure hash based authentication and revelation protocol in IoVs. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001146.

Zhai:2021:TPU

[600] Shuangjiao Zhai, Guixin Ye, Zhanyong Tang, Jie Ren, Dingyi Fang, Baoying Liu, and Zheng Wang. Towards practical 3D ultrasound sensing on commercial-off-the-shelf mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001134.

Zhong:2021:APP

[601] Hong Zhong, Jinpeng Fan, Jie Cui, Yan Xu, and Lu Liu. Assessing profit of prediction for SDN controllers load balancing. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001158.

Hu:2021:BBT

[602] Donghui Hu, Yifan Li, Lixuan Pan, Meng Li, and Shuli Zheng. A blockchain-based trading system for big data. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100116X.

Ergiz:2021:JMF

[603] Yamac Ergiz, Ali Murat Demirtas, and Tolga Girici. Joint multipath flow and layer allocation for scalable video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001171.

Chaudhry:2021:GIC

[604] Shehzad Ashraf Chaudhry, Khalid Yahya, Marimuthu Karuppiah, Rupak Kharel, Ali Kashif Bashir, and Yousaf Bin Zikria. GCACS-IoD: a certificate based generic access control scheme for Internet of drones. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001195.

Wu:2021:DRL

[605] Yulei Wu, Zehua Wang, Yuxiang Ma, and Victor C. M. Leung. Deep reinforcement learning for blockchain in Industrial IoT: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001213.

Brotsis:2021:SBP

[606] Sotirios Brotsis, Konstantinos Limniotis, Gueltoum Bendiab, Nicholas Kolokotronis, and Stavros Shiaeles. On the suitability of blockchain platforms for IoT applications: Architectures, security, privacy, and performance. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001225.

Caviglione:2021:KLT

[607] Luca Caviglione, Wojciech Mazurczyk, Matteo Repetto, Andreas Schaffhauser, and Marco Zuppelli. Kernel-level tracing for detecting stegomalware and covert channels in Linux environments. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001249.

Faghihi:2021:RDC

[608] Farnood Faghihi and Mohammad Zulkernine. RansomCare: Data-centric detection and mitigation against smartphone crypto-ransomware. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001250.

Yurekten:2021:CCT

[609] Ozgur Yurekten and Mehmet Demirci. Citadel: Cyber threat intelligence assisted defense system for softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 191 (??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001262.

Hadem:2021:SBI

[610] Pynbianglut Hadem, Dilip Kumar Saikia, and Soumen Moulik. An SDN-based intrusion detection system using SVM with selective logging for IP traceback. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001274.

Aazam:2021:TOE

[611] Mohammad Aazam, Sherali Zeadally, and Eduardo Feo Flushing. Task offloading in edge computing for machine learning-based smart healthcare. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001298.

Hei:2021:MMA

[612] Yiming Hei, Jianwei Liu, Hanwen Feng, Dawei Li, Yizhong Liu, and Qianhong Wu. Making MA-ABE fully accountable: a blockchain-based approach for secure digital right management. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001389.

Anonymous:2021:EBh

[613] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001870.

Anonymous:2021:Mc

[614] Anonymous. 22 May 2021. Computer Networks (Amsterdam, Netherlands: 1999), 191(??):??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Khorsandroo:2021:HSE

[615] Sajad Khorsandroo, Adrián Gallego Sánchez, Ali Saman Tosun, JM Arco, and Roberto Doriguzzi-Corin. Hybrid SDN evolution: a comprehensive survey of the state-of-the-art. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001109.

Spiekermann:2021:UPB

[616] Daniel Spiekermann and Jörg Keller. Unsupervised packet-based anomaly detection in virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621001286.

Sun:2021:CLT

[617] Zemin Sun, Yanheng Liu, Jian Wang, Rundong Yu, and Dongpu Cao. Crosslayer tradeoff of QoS and security in vehicular ad hoc networks: a game theoretical approach. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001390.

Jiang:2021:WML

[618] Huiling Jiang, Qing Li, Yong Jiang, GengBiao Shen, Richard Sinnott, Chen Tian, and Mingwei Xu. When machine learning meets congestion control: a survey and comparison. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001407.

Liu:2021:RTC

[619] Meili Liu, Xiaogang Qi, Lifang Liu, and Hao Pan. Roots-tracing of communication network alarm: a realtime processing framework. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001432.

Babun:2021:SIP

[620] Leonardo Babun, Kyle Denney, Z. Berkay Celik, Patrick McDaniel, and A. Selcuk Uluagac. A survey on IoT platforms: Communication, security, and privacy perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001444.

Yakici:2021:SMO

[621] Ertan Yakici and Mumtaz Karatas. Solving a multi-objective heterogeneous sensor network location problem with genetic algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001456.

Tomassilli:2021:DRP

[622] Andrea Tomassilli, Giuseppe Di Lena, Frédéric Giroire, Issam Tahiri, Damien Saucez, Stéphane Pérennes, Thierry Turletti, Ruslan Sadykov, François Vanderbeck, and Chidung Lac. Design of robust programmable networks with bandwidth-optimal failure recovery scheme. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001468.

Dou:2021:MMN

[623] Songshi Dou, Guochun Miao, Zehua Guo, Chao Yao, Weiran Wu, and Yuanqing Xia. Matchmaker: Maintaining network programmability for software-defined WANs under multiple controller failures. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100147X.

Priyadarsini:2021:SDN

[624] Madhukrishna Priyadarsini and Padmalochan Bera. Software defined networking architecture, traffic management, security, and placement: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001481.

Ashraf:2021:RQA

[625] Usman Ashraf, Adnan Ahmed, Mohammed Al-Naeem, and Unsa Masood. Reliable and QoS aware routing metrics for wireless neighborhood area networking in smart grids. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100150X.

Yang:2021:DAS

[626] Jiachen Yang, Shan Liu, Hansong Su, and Ying Tian. Driving assistance system based on data fusion of multisource sensors for autonomous unmanned ground vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001511.

Cong:2021:DRL

[627] Peizhuang Cong, Yuchao Zhang, Zheli Liu, Thar Baker, Hissam Tawfik, Wendong Wang, Ke Xu, Ruidong Li, and Fuliang Li. A deep reinforcement learning-based multi-optimality routing scheme for dynamic IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001535.

Jain:2021:UAR

[628] Akshay Jain, Elena Lopez-Aguilera, and Ilker Demirkol. User association and resource allocation in 5G (AURA-5G): a joint optimization framework. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001547.

Li:2021:MSD

[629] Fuliang Li, Zhenbei Guo, Bocheng Liang, Xiushuang Yi, Xingwei Wang, Weichao Li, and Yi Wang. A measurement study on device-to-device communication technologies for IIoT. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100164X.

Gupta:2021:LIH

[630] Neha Gupta, Vinita Jindal, and Punam Bedi. LIO-IDS: Handling class imbalance using LSTM and improved one-vs-one technique in intrusion detection system. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001675.

deVos:2021:CMF

[631] Martijn de Vos and Johan Pouwelse. ConTrib: Maintaining fairness in decentralized big tech alternatives by accounting work. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001705.

Vallero:2021:BSS

[632] Greta Vallero, Margot Deruyck, Michela Meo, and Wout Joseph. Base station switching and edge caching optimisation in high energy-efficiency wireless access network. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100181X.

Guo:2021:EFM

[633] Yinghao Guo, Zichao Zhao, Ke He, Shiwei Lai, Junjuan Xia, and Lisheng Fan. Efficient and flexible management for industrial Internet of Things: a federated learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001961.

Bryant:2021:EVN

[634] Blake Bryant and Hossein Saiedian. An evaluation of videogame network architecture performance and security. Computer Networks (Amsterdam, Netherlands: 1999), 192 (??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002000.

Anonymous:2021:EBi

[635] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S138912862100222X.

Anonymous:2021:Jb

[636] Anonymous. 19 June 2021. Computer Networks (Amsterdam, Netherlands: 1999), 192(??):??, June 19, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yu:2021:OEE

[637] Li Yu, Zongpeng Li, Jiangchuan Liu, and Ruiting Zhou. Online and energy-efficient task-processing for distributed edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100044X.

Tang:2021:JGL

[638] Yuan Tang, Yiming Miao, Ahmed Barnawi, Bander Alzahrani, Reem Alotaibi, and Kai Hwang. A joint global and local path planning optimization for UAV task scheduling towards crowd air monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621000682.

Bansal:2021:NGS

[639] Gaurang Bansal, Vinay Chamola, Georges Kaddoum, Md. Jalil Piran, and Mubarak Alrashoud. Next generation stock exchange: Recurrent neural learning model for distributed ledger transactions. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001183.

Shamsoshoara:2021:AIP

[640] Alireza Shamsoshoara, Fatemeh Afghah, Abolfazl Razi, Liming Zheng, Peter Z.

Fulé, and Erik Blasch. Aerial imagery pile burn detection using deep learning: the FLAME dataset. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001201.

Radoglou-Grammatikis:2021:SSS

[641] Panagiotis Radoglou-Grammatikis, Panagiotis Sarigiannidis, Eider Iturbe, Erkuden Rios, Saturnino Martinez, Antonios Sarigiannidis, Georgios Eftathopoulos, Yannis Spyridis, Achilleas Sesis, Nikolaos Vakakis, Dimitrios Tzovaras, Emmanouil Kafetzakis, Ioannis Giannoulakis, Michalis Tzifas, Alkiviadis Giannakoulias, Michail Angelopoulos, and Francisco Ramos. SPEAR SIEM: a security information and event management system for the smart grid. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001237.

Li:2021:MEC

[642] Bo Li, Xiaoheng Deng, and Yiqin Deng. Mobile-edge computing-based delay minimization controller placement in SDN-IoV. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001493.

Zhang:2021:TCF

[643] Qingyang Zhang, Hong Zhong, Weisong Shi, and Lu Liu. A trusted and collaborative framework for deep learning in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001523.

Huang:2021:MCM

[644] Shan Huang, Dezun Dong, Zejia Zhou, and Xiangke Liao. MP-CREDIT: Multi-path credit for highspeed data center transports. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001560.

Somesula:2021:CDA

[645] Manoj Kumar Somesula, Rashmi Ranjan Rout, and D. V. L. N. Somayajulu. Contact duration-aware cooperative cache placement using genetic algorithm for mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001559.

Ometov:2021:SWT

[646] Aleksandr Ometov, Viktoriia Shubina, Lucie Klus, Justyna Skibińska, Salwa Saafi, Pavel Pascacio, Laura Flueratoru, Darwin Quezada Gaibor, Nadezhda Chukhno, Olga Chukhno, Asad Ali, Asma Channa, Ekaterina Svertoka, Waleed Bin Qaim, Raúl Casanova-Marqués, Sylvia Holcer, Joaquín Torres-Sospedra, Sven Casteleyn, Giuseppe Ruggeri, Giuseppe Araniti, Radim Burget, Jiri Hosek, and Elena Simona Lohan. A survey on wearable technology: History, state-of-the-art and current challenges. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001651.

Zhou:2021:DLF

[647] Qizhao Zhou, Junqing Yu, and Dong Li. A dynamic and lightweight framework to secure source addresses in the SDN-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001663.

Wang:2021:TRL

[648] Ke Wang, Chien-Ming Chen, M. Shamim Hossain, Ghulam Muhammad, Sachin Kumar, and Saru Kumari. Transfer reinforcement learning-based road object detection in next generation IoT domain. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001687.

Vanderhallen:2021:RAA

[649] Stien Vanderhallen, Jo Van Bulck, Frank Piessens, and Jan Tobias Mühlberg. Robust authentication for automotive control networks through covert channels. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001699.

Wang:2021:DEC

[650] Jiansi Wang, Haopeng Chen, Fuxiao Zhou, Meng Sun, Ziang Huang, and Zhengtong Zhang. A-DECS: Enhanced collaborative edge-edge data storage service for edge computing with adaptive prediction. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001730.

Lu:2021:NTI

[651] Liangfu Lu, Xiaohan Ren, Chenwei Cui, Zhiyuan Tan, Yulei Wu, and Zhizhen Qin. A novel tensorinformation bottleneck method for multi-input single-output applications. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001742.

Lopez-Benitez:2021:SSR

[652] Miguel López-Benítez, Alessandro Raschellà, Sara Pizzi, Li Wang, Marco Di Felice, and Kaushik Roy Chowdhury. Smart spectrum and radio resource management for future 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001754.

Sen:2021:VET

[653] Priyangshu Sen, Viduneth Ariyarathna, Arjuna Madanayake, and Josep M. Jornet. A versatile experimental testbed for ultrabroadband communication networks above 100 GHz. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001778.

Shiny:2021:RGT

[654] S. Suja Golden Shiny, S. Sathya Priya, and K. Murugan. Repeated game theory-based reducer selection strategy for energy management in SD-WSN. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100178X.

Sharma:2021:PCP

[655] Abhilasha Sharma and Lalit Kumar Awasthi. Pr-CAI: Priority basedcontext aware information scheduling for SDN-based vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001791.

Bu:2021:SMP

[656] Kai Bu, Yutian Yang, Zixuan Guo, Yuanyuan Yang, Xing Li, and Shigeng Zhang. Securing middlebox policy enforcement in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001808.

Yang:2021:NTF

[657] Hanyu Yang, Xutao Li, Wenhao Qiang, Yuhan Zhao, Wei Zhang, and Chang Tang. A network traffic forecasting method based on SA optimized ARIMA-BP neural network. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001821.

Eramo:2021:ALS

[658] Vincenzo Eramo, Francesco Giacinto Lavacca, Tiziana Catena, and Paul Jaime Perez Salazar. Application of a long short term memory neural predictor with asymmetric loss function for the resource allocation in NFV network architectures. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001833.

Song:2021:FFC

[659] Huangfei Song, Songtao Guo, Pan Li, and Guiyan Liu. FCNR: Fast and consistent network reconfiguration with low latency for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001912.

Yarinezhad:2021:EER

[660] Ramin Yarinezhad and Sadoon Azizi. An energy-efficient routing protocol for the Internet of Things networks based on geographical location and link quality. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001936.

Yao:2021:RCM

[661] Xiaopeng Yao, Guangxian Liang, Chonglin Gu, and Hejiao Huang. Rumors clarification with minimum credibility in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621001973.

Valenza:2021:GES

[662] Fulvio Valenza, Matteo Repetto, and Stavros Shiaeles. Guest editorial: Special issue on novel cyber-security paradigms for software-defined and virtualized systems. Computer Networks (Amsterdam, Netherlands: 1999), 193 (??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001997.

Liu:2021:GBR

[663] Zhixin Liu, Xiaopin Li, Yazhou Yuan, Yi Yang, and Xinping Guan. Game based robust power allocation strategy with QoS guarantee in D2D communication network. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002012.

Anonymous:2021:EBj

[664] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002371.

Anonymous:2021:Jc

[665] Anonymous. 5 July 2021. Computer Networks (Amsterdam, Netherlands: 1999), 193(??):??, July 5, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Panicker:2021:AEO

[666] Ashwin Panicker, Ozgur Ozdemir, Mihail L. Sichitiu, Ismail Guvenc, Rudra Dutta, Vuk Marojevic, and Brian Floyd. AERPAW emulation overview and preliminary performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001717.

Bruschi:2021:MNS

[667] Roberto Bruschi, Jane Frances Pajo, Franco Davoli, and Chiara Lombardo. Managing 5G network slicing and edge computing with the MATILDA telecom layer platform. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001766.

Xie:2021:OAM

[668] Yanghao Xie, Sheng Wang, Binbin Wang, Shizhong Xu, Xiong Wang, and Jing Ren. Online algorithm for migration aware virtualized network function placing and routing in dynamic 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001924.

Yu:2021:PPB

[669] Lian Yu, Jingtao Dong, Lihao Chen, Mengyuan Li, Bingfeng Xu, Zhao Li, Lin Qiao, Lijun Liu, Bei Zhao, and Chen Zhang. PBCNN: Packet bytesbased convolutional neural network for network intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001948.

Wubben:2021:NRR

[670] Jamie Wubben, Francisco Fabra, Carlos T. Calafate, Juan-Carlos Cano, and Pietro Manzoni. A novel resilient and reconfigurable swarm management scheme. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100195X.

Shahraki:2021:TTB

[671] Amin Shahraki, Amir Taherkordi, and Øystein Haugen. TONTA: Trendbased online network traffic analysis in ad-hoc IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621001985.

El-Hajj:2021:TPS

[672] Mohammed El-Hajj, Ahmad Fadlallah, Maroun Chamoun, and Ahmed Serhrouchni. A taxonomy of PUF schemes with a novel arbiter-based PUF resisting machine learning attacks. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002036.

Ren:2021:MRR

[673] Quan Ren, Tao Hu, Jiangxing Wu, Yuxiang Hu, Lei He, and Julong Lan. Multipath resilient routing for endogenous secure software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002048.

Glabowski:2021:MSS

[674] Mariusz Głabowski, Maciej Sobieraj, and Maciej Stasiak. Modelling of single-sided multiservice switching structures. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100205X.

Palama:2021:ICW

[675] Ivan Palamà, Francesco Gringoli, Giuseppe Bianchi, and Nicola Blefari-Melazzi. IMSI catchers in the wild: a real world 4G/5G assessment. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002061.

Ozdem:2021:SAD

[676] Mehmet Özdem and Mustafa Alkan. Subscriber aware dynamic service function chaining. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002073.

Caiazza:2021:MDD

[677] Chiara Caiazza, Claudio Cicconetti, Valerio Luconi, and Alessio Vecchio. Measurement-driven design and runtime optimization in edge computing: Methodology and tools. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002085.

Gupta:2021:OPD

[678] Jitendra Gupta, Md Shahbaz Akhtar, Aneek Adhya, and Sudhan Majhi. Optimal planning and design of SRLGaware survivable LR-PON for wireless and FTTx networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002097.

Yigit:2021:BFS

[679] Yasin Yigit, Vahid Khalilpour Akram, and Orhan Dagdeviren. Breadth-first search tree integrated vertex cover algorithms for link monitoring and routing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002103.

Maia:2021:IMO

[680] Adyson M. Maia, Yacine Ghamri-Doudane, Dario Vieira, and Miguel Franklin de Castro. An improved multiobjective genetic algorithm with heuristic initialization for service placement and load distribution in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002127.

Polverini:2021:EDL

[681] Marco Polverini, Juan Luis Herrera, Pierpaolo Salvo, and Jaime Galán-Jiménez. Early detection of link failures through the modeling of the hardware deterioration process. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002115.

Navidan:2021:GAN

[682] Hojjat Navidan, Parisa Fard Moshiri, Mohammad Nabati, Reza Shahbazian, Seyed Ali Ghorashi, Vahid Shah-Mansouri, and David Windridge. Generative adversarial networks (GANs) in networking: a comprehensive survey and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002139.

Wang:2021:DSA

[683] Rui Wang, Hongchao Du, Zhaoyan Shen, and Zhiping Jia. DAP-Sketch: an accurate and effective network measurement sketch with Deterministic Admission Policy. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002164.

Magan-Carrion:2021:UIW

[684] Roberto Magán-Carrión, Alberto Abellán-Galera, Gabriel Maciá-Fernández, and Pedro García-Teodoro. Unveiling the I2P web structure: a connectivity analysis. Computer Networks (Amsterdam, Netherlands: 1999), 194 (??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002188.

Cai:2021:DAP

[685] Yun-Zhan Cai, Yu-Ting Wang, and Meng-Hsun Tsai. Dynamic adjustment for proactive flow installation mechanism in SDN-based IoT. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002279.

Xing:2021:VBM

[686] Tianzhang Xing, Chase Wu, Jie Wang, Fei Shang, Ruilin Li, and Xiaojiang Chen. A vibration-based multiuser concurrent communication system with commercial devices. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002280.

Anonymous:2021:EBk

[687] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002577.

Anonymous:2021:Jd

[688] Anonymous. 20 July 2021. Computer Networks (Amsterdam, Netherlands: 1999), 194(??):??, July 20, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yao:2021:KKM

[689] Yingbiao Yao, Yuancheng Qin, Wei Feng, Pei Li, Xiaorong Xu, Xin Xu, and Xuesong Liang. KFTO: Kuhn–Munkres based fair task offloading in fog networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002024.

Mohamed:2021:SDN

[690] Arwa Mohamed, Mosab Hamdan, Suleman Khan, Ahmed Abdelaziz, Sharief F. Babiker, Muhammad Imran, and M. N. Marsono. Software-defined networks for resource allocation in cloud computing: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002152.

elhoudaNouar:2021:SVN

[691] Nour el houda Nouar, Sami Yangui, Noura Faci, Khalil Drira, and Saïd Tazi. A semantic virtualized network functions description and discovery model. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002140.

Magoula:2021:GAA

[692] Lina Magoula, Sokratis Barmpounakis, Ioannis Stavrakakis, and Nancy Alonistioti. A genetic algorithm approach for service function chain placement in 5G and beyond, virtualized edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002176.

Hortelano:2021:REC

[693] Diego Hortelano, Teresa Olivares, and M. Carmen Ruiz. Reducing the energy consumption of the friendship mechanism in Bluetooth mesh. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002309.

Steadman:2021:DED

[694] Jacob Steadman and Sandra Scott-Hayward. DNSxP: Enhancing data exfiltration protection through data plane programmability. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002310.

Huang:2021:MAR

[695] Binbin Huang, Xiao Liu, Shangguang Wang, Linxuan Pan, and Victor Chang. Multi-agent reinforcement learning for cost-aware collaborative task execution in energyharvesting D2D networks. Computer Networks (Amsterdam, Netherlands:

1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002334.

Saeik:2021:TOE

[696] Firdose Saeik, Marios Avgeris, Dimitrios Spatharakis, Nina Santi, Dimitrios Dechouniotis, John Violos, Aris Leivadeas, Nikolaos Athanasopoulos, Nathalie Mitton, and Symeon Papavassiliou. Task offloading in edge and cloud computing: a survey on mathematical, artificial intelligence and control theory solutions. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002322.

Chen:2021:EIC

[697] Miaojiang Chen, Wei Liu, Tian Wang, Anfeng Liu, and Zhiwen Zeng. Edge intelligence computing for mobile augmented reality with deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002425.

Kosek-Szott:2021:DCA

[698] Katarzyna Kosek-Szott, Alice Lo Valvo, Szymon Szott, Pierluigi Gallo, and Ilenia Tinnirello. Downlink channel access performance of NR-U: Impact of numerology and mini-slots on coexistence with Wi-Fi in the 5 GHz band. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002437.

Aguilar-Fuster:2021:NEF

[699] Christian Aguilar-Fuster and Javier Rubio-Loyola. A novel evaluation function for higher acceptance rates and more profitable metaheuristicbased online virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002449.

Li:2021:STP

[700] Junfeng Li, Dan Li, Wenfei Wu, K. K. Ramakrishnan, Jinkun Geng, Fanzhao Wang, and Kai Zheng. Sphinx: a transport protocol for high-speed and lossy mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002462.

Chakraborty:2021:DIN

[701] Tamal Chakraborty and Iti Saha Misra. Design and implementation of a novel two-phase spectrum handoff scheme for QoS aware mobile users in cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002450.

Tu:2021:ROM

[702] Shanshan Tu, Muhammad Waqas, Fengming Huang, Ghulam Abbas, and Ziaul Haq Abbas. A revocable and outsourced multi-authority attributebased encryption scheme in fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002474.

Gao:2021:EAB

[703] Honghao Gao, Walayat Hussain, Yuyu Yin, Wenbing Zhao, and Muddesar Iqbal. Editorial: AI-based mobile multimedia computing for datasmart processing. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002486.

Rojas:2021:CSH

[704] Elisa Rojas, Rashid Amin, Carmen Guerrero, Marco Savi, and Adib Rastegarnia. Challenges and solutions for hybrid SDN. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002498.

Liu:2021:REE

[705] Zhixin Liu, Meihua Zhou, Yanyan Shen, Yazhou Yuan, Kit Yan Chan, and Yi Yang. Robust energy efficient maximization in wireless powered CRNs based on power splitting. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002528.

Junior:2021:DRS

[706] Franklin Magalhães Ribeiro Junior and Carlos Alberto Kamienski. Data resilience system for fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 195 (??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002656.

Irshad:2021:SBO

[707] Azeem Irshad, Shehzad Ashraf Chaudhry, Anwar Ghani, and Muhammad Bilal. A secure blockchain-oriented data delivery and collection scheme for 5Genabled IoD environment. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002668.

Tian:2021:PSP

[708] Junfeng Tian and Haoning Wang. A provably secure and public auditing protocol based on the bell triangle for cloud data. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100267X.

Anonymous:2021:EBl

[709] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621003169.

Anonymous:2021:A

[710] Anonymous. 4 August 2021. Computer Networks (Amsterdam, Netherlands: 1999), 195(??):??, August 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Arce:2021:FBL

[711] Pau Arce, David Salvo, Gema Piñero, and Alberto Gonzalez. FIWARE based low-cost wireless acoustic sensor network for monitoring and classification of urban soundscape. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002504.

Chadda:2021:ACW

[712] Amel Chadda, Marija Stojanova, Thomas Begin, Anthony Busson, and Isabelle Guérin Lassous. Assigning channels in WLANs with channel bonding: a fair and robust strategy. Computer Networks (Amsterdam, Netherlands: 1999), 196(??): ??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002516.

Khalily-Dermany:2021:TPA

[713] M. Khalily-Dermany. Transmission power assignment in network-coding-based-multicast-wireless-sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100253X.

Pustokhina:2021:EEN

[714] Irina V. Pustokhina, Denis A. Pustokhin, E. Laxmi Lydia, Mohamed Elhoseny, and K. Shankar. Energy efficient neuro-fuzzy cluster based topology construction with metaheuristic route planning algorithm for unmanned aerial vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002632.

Khan:2021:BBD

[715] Ammar Ahmed Khan, Muhammad Mubashir Khan, Kashif Mehboob Khan, Junaid Arshad, and Farhan Ahmad. A blockchain-based decentralized machine learning framework for collaborative intrusion detection within UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002644.

Dhananjay:2021:PRV

[716] Aditya Dhananjay, Kai Zheng, Marco Mezzavilla, Lorenzo Iotti, Dennis Shasha, and Sundeep Rangan. Pi-Radio v1: Calibration techniques to enable fully-digital beamforming at 60 GHz. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002620.

Bayazeed:2021:SSC

[717] Adnan Bayazeed, Khaldoun Khorzom, and Mohamad Aljnidi. A survey of self-coordination in self-organizing network. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002681.

Wang:2021:DNL

[718] Dongsheng Wang, Prayag Tiwari, Mohammad Shorfuzzaman, and Ingo Schmitt. Deep neural learning on weighted datasets utilizing label disagreement from crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 196(??): ??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002711.

Paguada:2021:TPA

[719] Servio Paguada, Lejla Batina, and Igor Armendariz. Toward practical autoencoder-based side-channel analysis evaluations. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002735.

Nayak:2021:SRB

[720] Sabuzima Nayak, Ripon Patgiri, and Angana Borah. A survey on the roles of Bloom filter in implementation of the named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002747.

Paul:2021:BBS

[721] Rourab Paul, Nimisha Ghosh, Suman Sau, Amlan Chakrabarti, and Prasant Mohapatra. Blockchain based secure smart city architecture using low resource IoTs. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002759.

Sun:2021:CCD

[722] Rong Sun, Huihui Zheng, and Jingwei Liu. Coded caching design for fogaided networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002760.

Bruschi:2021:FHP

[723] Valerio Bruschi, Salvatore Pontarelli, Jerome Tollet, Dave Barach, and Giuseppe Bianchi. FlowFight: High performance-low memory top-k spreader detection. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002772.

Behera:2021:PMB

[724] Sadananda Behera and Goutam Das. Penalty-method based impairmentaware performance of elastic optical networks with spectrum conversion. Computer Networks (Amsterdam, Netherlands: 1999), 196(??): ??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002784.

Aghaei:2021:TNA

[725] Abdollah Aghaei, Javad Akbari Torkestani, Hamidreza Kermajani, and Abbas Karimi. LA-Trickle: a novel algorithm to reduce the convergence time of the wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002796.

Achroufene:2021:MCC

[726] Achour Achroufene, Mourad Chelik, and Nassima Bouadem. Modified CSMA/CA protocol for realtime data fusion applications based on clustered WSN. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002802.

Gaur:2021:SVP

[727] Kuntal Gaur, Anshuman Kalla, Jyoti Grover, Mohammad Borhani, Andrei Gurtov, and Madhusanka Liyanage. A survey of virtual private LAN services (VPLS): Past, present and future. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002814.

Tseng:2021:LLR

[728] Yuh-Min Tseng, Jian-Lun Chen, and Sen-Shan Huang. A lightweight leakage-resilient identity-based mutual authentication and key exchange protocol for resource-limited devices. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002826.

Zhang:2021:RAM

[729] Bibo Zhang, Francesco Devoti, Ilario Filippini, and Danilo De Donno. Resource allocation in mmWave 5G IAB networks: a reinforcement learning approach based on column generation. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002838.

Chen:2021:DQN

[730] Xin Chen, Xu Liu, Ying Chen, Libo Jiao, and Geyong Min. Deep Q-Network based resource allocation for UAV-assisted ultra-dense networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100284X.

Zambianco:2021:IMB

[731] Marco Zambianco and Giacomo Verticale. Intelligent multi-branch allocation of spectrum slices for inter-numerology interference minimization. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002875.

Uddin:2021:DPA

[732] Md. Forkan Uddin. Downlink performance analysis of a CSMA based WLAN under physical interference model. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002887.

Wan:2021:JCO

[733] Zhilan Wan, Ding Xu, Dahu Xu, and Ishtiaq Ahmad. Joint computation offloading and resource allocation for NOMA-based multi-access mobile edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002863.

Aurizzi:2021:SBT

[734] Matteo Maria Aurizzi, Tommaso Rossi, Emanuele Raso, Ludovico Funari, and Ernestina Cianca. An SDN-based traffic handover control procedure and SGD management logic for EHF satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002905.

Tang:2021:UIE

[735] Yutao Tang, Yue Li, Qun Li, Kun Sun, Haining Wang, and Zhengrui Qin. User input enrichment via sensing devices. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002917.

Touati:2021:CRD

[736] Haifa Touati, Amira Chriki, Hichem Snoussi, and Farouk Kamoun. Cognitive radio and dynamic TDMA for efficient UAVs swarm communications. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002929.

Xie:2021:SAD

[737] Guorui Xie, Qing Li, and Yong Jiang. Self-attentive deep learning method for online traffic classification and its interpretability. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002930.

Krishnamurthy:2021:AFA

[738] Prashanth Krishnamurthy, Farshad Khorrami, and Rahul Kumar. An approximate factorization approach to multi-jammer location and range estimation from peer-to-peer connectivity measurements. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002942.

Kiani:2021:NAP

[739] Mohsen Kiani and Mohammad Reza Khayyambashi. A network-aware and power-efficient virtual machine placement scheme in cloud datacenters based on chemical reaction optimization. Computer Networks (Amsterdam, Netherlands: 1999), 196 (??):??, September 4, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002954.

Anonymous:2021:EBm

[740] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621003650.

Anonymous:2021:S

[741] Anonymous. 4 September 2021. Computer Networks (Amsterdam, Netherlands: 1999), 196(??):??, September 4, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dziyauddin:2021:COC

[742] Rudzidatul Akmam Dziyauddin, Dusit Niyato, Nguyen Cong Luong, Ahmad Ariff Aizuddin Mohd Atan, Mohd Azri Mohd Izhar, Marwan Hadri Azmi, and Salwani Mohd Daud. Computation offloading and content caching and delivery in Vehicular Edge Network: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002723.

SureshKumar:2021:EER

[743] K. SureshKumar and P. Vimala. Energy efficient routing protocol using exponentially-ant lion whale optimization algorithm in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002851.

Dinh:2021:DUM

[744] Thi Ha Ly Dinh, Megumi Kaneko, Keisuke Wakao, Kenichi Kawamura, Takatsune Moriyama, Hirantha Abeysekera, and Yasushi Takatori. Distributed user-to-multiple access points association through deep learning for beyond 5G. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002899.

Yang:2021:AHR

[745] Xuwei Yang, Hongli Xu, Jianchun Liu, Chen Qian, Xingpeng Fan, He Huang, and Haibo Wang. Achieving high reliability and throughput in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002966.

Rahman:2021:AAS

[746] Md. Rashedur Rahman, Md. Motaharul Islam, Ahmed Iqbal Pritom, and Yazed Alsaawy. ASRPH: Application specific routing protocol for health care. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621002978.

Boucetta:2021:LRB

[747] Chérifa Boucetta, Boubakr Nour, Michel Sortais, and Hassine Moungla. A Latin rectangles-based TSCH scheduling and interference mitigation design. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100298X.

Belhamra:2021:ECN

[748] Mohamed Amine Belhamra and El Mamoun Souidi. Error correcting network codes. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621002991.

Al-Marridi:2021:RLA

[749] Abeer Z. Al-Marridi, Amr Mohamed, and Aiman Erbad. Reinforcement learning approaches for efficient and secure blockchain-powered smart health systems. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003005.

Breen:2021:PPO

[750] Joe Breen, Andrew Buffmire, Jonathon Duerig, Kevin Dutt, Eric Eide, Anneswa Ghosh, Mike Hibler, David Johnson, Sneha Kumar Kasera, Earl Lewis, Dustin Maas, Caleb Martin, Alex Orange, Neal Patwari, Daniel Reading, Robert Ricci, David Schurig, Leigh B. Stoller, Allison Todd, Jacobus Van der Merwe, Naren Viswanathan, Kirk Webb, and Gary Wong. Powder: Platform for open wireless data-driven experimental research. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003017.

Alvares:2021:DAL

[751] Christabelle Alvares, Dristi Dinesh, Syed Alvi, Tannish Gautam, Maheen Hasib, and Ali Raza. Dataset of attacks on a live enterprise VoIP network for machine learning based intrusion detection and prevention systems. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003030.

Shakouri:2021:PRL

[752] Saba Shakouri and Akbar Ghaffarpour Rahbar. Proactive restoration with low re-provisioning in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003042.

Dinarte:2021:RSA

[753] Henrique A. Dinarte, Bruno V. A. Correia, Daniel A. R. Chaves, and Raul C. Almeida. Routing and spectrum assignment: a metaheuristic for hybrid ordering selection in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003054.

Madi:2021:NSS

[754] Taous Madi, Hyame Assem Alameddine, Makan Pourzandi, and Amine Boukhtouta. NFV security survey in 5G networks: a three-dimensional threat taxonomy. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003066.

Thang:2021:JBU

[755] Ha Duc Thang, Lila Boukhatem, Megumi Kaneko, and Nhan Nguyen-Thanh. Joint beamforming and user association with reduced CSI signaling in mobile environments: a Deep Q-learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003078.

Shao:2021:PRI

[756] Xiaozhe Shao and Lixin Gao. Policyrich interdomain routing with local coordination. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100308X.

Campeon:2021:JPC

[757] Juan Pablo Grisales Campeón and Pablo I. Fierens. Joint position and clock tracking of wireless nodes. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003091.

Fang:2021:CCB

[758] Yong Fang, Kai Li, Rongfeng Zheng, Shan Liao, and Yue Wang. A communication-channel-based method for detecting deeply camouflaged malicious traffic. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100311X.

Khan:2021:QLB

[759] Zahoor Ali Khan, Obaida Abdul Karim, Shahid Abbas, Nadeem Javaid, Yousaf Bin Zikria, and Usman Tariq. Q-learning based energyefficient and void avoidance routing protocol for underwater acoustic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003212.

Tang:2021:EFS

[760] Yinan Tang, Tongtong Yuan, Bo Liu, and Chuangbai Xiao. Effective *- flow schedule for optical circuit switching based data center networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003224.

Wang:2021:CSD

[761] Yue Wang, Anmin Zhou, Shan Liao, Rongfeng Zheng, Rong Hu, and Lei Zhang. A comprehensive survey on DNS tunnel detection. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003248.

Vegni:2021:MMB

[762] Anna Maria Vegni, Carlos Borrego Iglesias, and Valeria Loscrí. MOVES: a MemOry-based VEhicular Social forwarding technique. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003236.

Gringoli:2021:LDH

[763] Francesco Gringoli and Douglas J. Leith. Low-delay high-rate operation of 802.11ac WLAN downlink: Nonlinear controller analysis and design. Computer Networks (Amsterdam, Netherlands: 1999), 197(??): ??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003261.

Chai:2021:NCO

[764] Rong Chai, Xizheng Yang, Chunling Du, and Qianbin Chen. Network cost optimization-based capacitated controller deployment for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003273.

Miao:2021:FDR

[765] Qinyang Miao, Hui Lin, Xiaoding Wang, and Mohammad Mehedi Hassan. Federated deep reinforcement learning based secure data sharing for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003285.

Yin:2021:IDC

[766] Sheng lin Yin, Xing lan Zhang, and Shuo Liu. Intrusion detection for capsule networks based on dual routing mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003327.

Lu:2021:CCM

[767] Yifei Lu, Xu Ma, and Zhengzhi Xu. Choose a correct marking position: ECN should be freed from tail mark. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003297.

Fotohi:2021:SCB

[768] Reza Fotohi and Fereidoon Shams Aliee. Securing communication between things using blockchain technology based on authentication and SHA-256 to improving scalability in large-scale IoT. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003303.

Lezzar:2021:OUR

[769] Mohamed Yacine Lezzar and Mustafa Mehmet-Ali. Optimization of ultrareliable low-latency communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003339.

Li:2021:SBN

[770] Quanzhong Li and Liang Yang. Secure beamforming with nonorthogonal multiple access transmission in cooperative CR networks for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003340.

Seoane:2021:PEC

[771] Victor Seoane, Carlos Garcia-Rubio, Florina Almenares, and Celeste Campo. Performance evaluation of CoAP and MQTT with security support for IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003364.

Wei:2021:ASE

[772] Guiyi Wei, Xiaohang Mao, Rongxing Lu, Jun Shao, Yunguo Guan, and Genhua Lu. Achieve spaceefficient key management in lightning network. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003406.

Bhooanusas:2021:PMM

[773] Nuntanut Bhooanusas and Sok-Ian Sou. Performance modeling of multipath mobile data offloading in cellular/WiFi networks with bandwidth uncertainty. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003431.

Lindroos:2021:SMC

[774] Saku Lindroos, Antti Hakkala, and Seppo Virtanen. A systematic methodology for continuous WLAN abundance and security analysis. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003479.

Lv:2021:BDT

[775] Zhihan Lv, Dongliang Chen, Hailing Feng, Ranran Lou, and Huihui Wang. Beyond 5G for digital twins of UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 197 (??):??, October 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003534.

Anonymous:2021:EBn

[776] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621004060.

Anonymous:2021:Oa

[777] Anonymous. 9 October 2021. Computer Networks (Amsterdam, Netherlands: 1999), 197(??):??, October 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gharrad:2021:FSD

[778] Hana Gharrad, Nafaa Jabeur, Ansar Ul-Haque Yasar, Stephane Galland, and Mohammed Mbarki. A fivestep drone collaborative planning approach for the management of distributed spatial events and vehicle notification using multi-agent systems and firefly algorithms. Computer Networks (Amsterdam, Netherlands:

1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003029.

Chen:2021:FSW

[779] Mantun Chen, Yongjun Wang, Hongzuo Xu, and Xiatian Zhu. Few-shot website fingerprinting attack. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003108.

He:2021:TSD

[780] Lin He, Peng Kuang, Ying Liu, Gang Ren, and Jiahai Yang. Towards securing duplicate address detection using P4. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100325X.

Busacca:2021:DML

[781] Fabio Busacca, Giuseppe Faraci, Christian Grasso, Sergio Palazzo, and Giovanni Schembra. Designing a multi-layer edge-computing platform for energy-efficient and delayaware offloading in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003315.

Mao:2021:MMS

[782] Beifeng Mao, Jing Liu, Yingxu Lai, and Motong Sun. MIF: a multi-step attack scenario reconstruction and attack chains extraction method based on multi-information fusion. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003376.

Rea:2021:BSM

[783] Maurizio Rea, Domenico Giustiniano, Pablo Jiménez Mateo, Yago Lizarribar, and Joerg Widmer. Beam searching for mmWave networks with sub-6 GHz WiFi and inertial sensors inputs: an experimental study. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100339X.

Yan:2021:PIF

[784] Fulong Yan, Hugo Meyer, Changshun Yuan, Xuwei Xue, Bitao Pan, Xiaotao Guo, Nicola Calabretta, and Chongjin Xie. On the performance investigation of a fast optical switches based optical high performance computing infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003418.

Roostaei:2021:GTJ

[785] Razie Roostaei, Zahra Dabiri, and Zeinab Movahedi. A game-theoretic joint optimal pricing and resource allocation for Mobile Edge Computing in NOMA-based 5G networks and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100342X.

Parsa:2021:TAM

[786] Mahshad Parsa and Seyed Ahmad Motamedi. Throughput analysis in a millimeter-wave network for a hybrid unsaturated RCAP/CTAP MAC protocol. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003443.

Ali:2021:SCO

[787] Zaiwar Ali, Ziaul Haq Abbas, Ghulam Abbas, Abdullah Numani, and Muhammad Bilal. Smart computational offloading for mobile edge computing in next-generation Internet of Things networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003467.

Perdices:2021:NLP

[788] Daniel Perdices, Javier Ramos, José L. García-Dorado, Iván González, and Jorge E. López de Vergara. Natural language processing for web browsing analytics: Challenges, lessons learned, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003455.

Cai:2021:FRD

[789] Lingfeng Cai, Xianglin Wei, Changyou Xing, Xia Zou, Guomin Zhang, and Xiulei Wang. Failure-resilient DAG task scheduling in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003480.

Nayyer:2021:LBH

[790] Amit Nayyer, Aman Kumar Sharma, and Lalit Kumar Awasthi. Learningbased hybrid routing for scalability in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003492.

Qi:2021:PPS

[791] Xiaogang Qi, Jiahui Li, Zhiping Wang, and Lifang Liu. Probabilistic probe selection algorithm for fault diagnosis in communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003522.

Zheng:2021:ADP

[792] Ling Zheng, Weitao Pan, Ya Gao, Huan Liu, and Jing Jiang. Architecture design and performance analysis of a novel memory system for high-bandwidth onboard switching fabric. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003509.

Mokhtar:2021:MLT

[793] Hamza Mokhtar, Xiaoqiang Di, Ying Zhou, Alzubair Hassan, Ziyi Ma, and Shafiu Musa. Multiple-level threshold load balancing in distributed SDN controllers. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003510.

Munilla:2021:ESK

[794] Jorge Munilla, Mike Burmester, and Raquel Barco. An enhanced symmetric-key based 5G-AKA protocol. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003546.

Xu:2021:GAC

[795] Rui Xu, Xu Wang, and Kirill Morozov. Group authentication for cloudto-things computing: Review and improvement. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100356X.

Ma:2021:ASA

[796] Dongchao Ma, Xiaofu Huang, Yuekun Hu, Pengyu Wang, Mingwei Xu, and Li Ma. An adaptive solar-aware framework and strategy for outdoor deployment of WSN. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003583.

Zilberman:2021:HSC

[797] Aviram Zilberman, Yoram Haddad, Sefi Erlich, Yossi (Joseph) Peretz, and Amit Dvir. Heterogeneous SDN controller placement problem — the Wi-Fi and 4G LTE-U case. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003558.

Liu:2021:BTR

[798] Yuchen Liu and Douglas M. Blough. Blockage tolerance in roadside millimeterwave backhaul networks. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003571.

daCosta:2021:QFI

[799] Luis Antonio L. F. da Costa, Rafael Kunst, and Edison Pignaton de Freitas. Q-FANET: Improved Qlearning based routing protocol for FANETs. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003595.

Feng:2021:MTM

[800] Li Feng, Yiru Yao, Liangmin Wang, and Geyong Min. Multi-timescale and multi-centrality layered node selection for efficient traffic monitoring in SDNs. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003601.

Tian:2021:MNM

[801] Donghai Tian, Qianjin Ying, Xiaoqi Jia, Rui Ma, Changzhen Hu, and Wenmao Liu. MDCHD: a novel malware detection method in cloud using hardware trace and deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003728.

Alam:2021:BBI

[802] Shadab Alam, Mohammed Shuaib, Wazir Zada Khan, Sahil Garg, Georges Kaddoum, M. Shamim Hossain, and Yousaf Bin Zikria. Blockchain-based initiatives: Current state and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100373X.

Ding:2021:UCI

[803] Lan Ding, Ye Tian, Tong Liu, Zhongxiang Wei, and Xinming Zhang. Understanding commercial 5G and its implications to (multipath) TCP. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003765.

Hou:2021:DTF

[804] Changsheng Hou, Bingnan Hou, Tongqing Zhou, and Zhiping Cai. DMatrix: Toward fast and accurate queries in graph stream. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003777.

Rioja:2021:ATP

[805] Unai Rioja, Lejla Batina, Jose Luis Flores, and Igor Armendariz. Autotune POIs: Estimation of distribution algorithms for efficient side-channel analysis. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003789.

Esfahani:2021:ECA

[806] Mahdi Esfahani, Hadi Soleimany, and Mohammad Reza Aref. Enhanced cache attack on AES applicable on ARM-based devices with new operating systems. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003790.

Jin:2021:ESC

[807] Shan Jin and Riccardo Bettati. Efficient side-channel attacks beyond divide-and-conquer strategy. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003807.

Khaturia:2021:FUM

[808] Meghna Khaturia, Pranav Jha, and Abhay Karandikar. 5G-Flow: a unified multi-RAT RAN architecture for beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003820.

Prathapchandran:2021:TAS

[809] K. Prathapchandran and T. Janani. A trust aware security mechanism to detect sinkhole attack in RPL-based IoT environment using random forest — RFTRUST. Computer Networks (Amsterdam, Netherlands: 1999), 198 (??):??, October 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003832.

Eriskin:2021:PCH

[810] Levent Eriskin. Point coverage with heterogeneous sensor networks: a robust optimization approach under target location uncertainty. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003844.

Shokouhifar:2021:SIR

[811] Mohammad Shokouhifar. Swarm intelligence RFID network planning using multi-antenna readers for asset tracking in hospital environments. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003911.

Anonymous:2021:EBo

[812] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621004278.

Anonymous:2021:Ob

[813] Anonymous. 24 October 2021. Computer Networks (Amsterdam, Netherlands: 1999), 198(??):??, October 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ahat:2021:OSS

[814] Betül Ahat, Ahmet Cihat Baktır, Necati Aras, İ. Kuban Altınel, Atay Özgövde, and Cem Ersoy. Optimal server and service deployment for multi-tier edge cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 199(??): ??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003716.

Liu:2021:TOO

[815] Peng Liu, Han He, Tingting Fu, Huijuan Lu, Abdulhameed Alelaiwi, and Md Wasif Islam Wasi. Task offloading optimization of cruising UAV with fixed trajectory. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003741.

Baarli:2021:DMU

[816] Eirik Molde Bårli, Anis Yazidi, Enrique Herrera Viedma, and Hårek Haugerud. DoS and DDoS mitigation using variational autoencoders. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003753.

Popli:2021:GNA

[817] Sakshi Popli, Rakesh Kumar Jha, and Sanjeev Jain. Green NOMA assisted NB-IoT based urban farming in multistory buildings. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003819.

Salehi:2021:IUB

[818] Shavbo Salehi and Behdis Eslamnour. Improving UAV base station energy efficiency for Industrial IoT URLLC services by irregular repetition slotted-ALOHA. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003856.

Halgamuge:2021:OFB

[819] Malka N. Halgamuge. Optimization framework for best approver selection method (BASM) and best tip selection method (BTSM) for IOTA tangle network: Blockchain-enabled next generation Industrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003868.

Kohli:2021:OAF

[820] Manav Kohli, Tingjun Chen, Mahmood Baraani Dastjerdi, Jackson Welles, Ivan Seskar, Harish Krishnaswamy, and Gil Zussman. Openaccess full-duplex wireless in the OR-BIT and COSMOS testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100387X.

Qin:2021:CAM

[821] Hua Qin, Weihong Chen, Weimin Chen, Ni Li, Min Zeng, and Yang Peng. A collision-aware mobile tag reading algorithm for RFID-based vehicle localization. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003881.

Platt:2021:SAI

[822] Moritz Platt and Peter McBurney. Sybil attacks on identity-augmented proof-of-stake. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003893.

Xiang:2021:RCM

[823] Bin Xiang, Jocelyne Elias, Fabio Martignon, and Elisabetta Di Nitto. Resource calendaring for mobile edge computing: Centralized and decentralized optimization approaches. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862100390X.

Liu:2021:CEA

[824] Jianchun Liu, Hongli Xu, Yang Xu, Zhenguo Ma, Zhiyuan Wang, Chen Qian, and He Huang. Communicationefficient asynchronous federated learning in resource-constrained edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003923.

Arnold:2021:CDB

[825] Rachel Arnold and Dave Longley. Continuity: a deterministic Byzantine fault tolerant asynchronous consensus algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003935.

Liu:2021:FDD

[826] Yong Liu and Guangxia Xu. Fixed degree of decentralization DPoS consensus mechanism in blockchain based on adjacency vote and the average fuzziness of vague value. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003947.

Popli:2021:CSG

[827] Sakshi Popli, Rakesh Kumar Jha, and Sanjeev Jain. A comprehensive survey on green ICT with 5G-NB-IoT: Towards sustainable planet. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003959.

Fourati:2021:CSS

[828] Hasna Fourati, Rihab Maaloul, Lamia Chaari, and Mohamed Jmaiel. Comprehensive survey on self-organizing cellular network approaches applied to 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003960.

Alhowaidi:2021:CML

[829] Mohammad Alhowaidi, Deepak Nadig, Boyang Hu, Byrav Ramamurthy, and Brian Bockelman. Cache management for large data transfers and multipath forwarding strategies in named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003972.

Wei:2021:COM

[830] Dawei Wei, Jianfeng Ma, Linbo Luo, Yunbo Wang, Lei He, and Xinghua Li. Computation offloading over multi-UAV MEC network: a distributed deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003984.

Deng:2021:SDC

[831] Qingyong Deng, Shaobo Huang, Zhetao Li, Bin Guo, Liyao Xiang, and Rong Ran. A secure data collection strategy using mobile vehicles joint UAVs in smart city. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003996.

Jagannath:2021:DMC

[832] Anu Jagannath and Jithin Jagannath. Dataset for modulation classification and signal type classification for multitask and single task learning. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100400X.

Nait-Hamoud:2021:CPK

[833] Othmane Nait-Hamoud, Tayeb Kenaza, and Yacine Challal. Certificateless public key systems aggregation: an enabling technique for 5G multidomain security management and delegation. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004011.

Arana:2021:NQA

[834] Oscar Arana, Hector Benítez-Pérez, Javier Gomez, and Miguel Lopez-Guerrero. Never query alone: a distributed strategy to protect Internet users from DNS fingerprinting attacks. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004023.

Sharaff:2021:SMD

[835] Aakanksha Sharaff, Chandramani Kamal, Siddhartha Porwal, Surbhi Bhatia, Kuljeet Kaur, and Mohammad Mehendi Hassan. Spam message detection using Danger theory and Krill herd optimization. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004102.

Sheshjavani:2021:CCS

[836] Abdollah Ghaffari Sheshjavani, Ahmad Khonsari, Seyed Pooya Shariatpanahi, and Masoumeh Moradian. Content caching for shared medium networks under heterogeneous users’ behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004114.

Liao:2021:TAB

[837] Xuening Liao, Zhenqiang Wu, Yuanyu Zhang, and Xiaohong Jiang. Trustaware buffer-aided relay selection for secure communications in cooperative wireless systems. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004126.

Sadri:2021:ATF

[838] Mohammad Javad Sadri and Maryam Rajabzadeh Asaar. An anonymous two-factor authentication protocol for IoT-based applications. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004151.

Vemireddy:2021:FRL

[839] Satish Vemireddy and Rashmi Ranjan Rout. Fuzzy reinforcement learning for energy efficient task offloading in vehicular fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004163.

Amanlou:2021:LSA

[840] Sanaz Amanlou, Mohammad Kamrul Hasan, and Khairul Azmi Abu Bakar. Lightweight and secure authentication scheme for IoT network based on publish-subscribe fog computing model. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004175.

Liu:2021:FAP

[841] Juncai Liu, Jessie Hui Wang, Chenghao Rong, Yuedong Xu, Tao Yu, and Jilong Wang. FedPA: an adaptively partial model aggregation strategy in federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004199.

Velasco:2021:LMS

[842] Francisco Alcaraz Velasco, Jose Manuel Palomares, and Joaquin Olivares. Lightweight method of shuffling overlapped data-blocks for data integrity and security in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004205.

Cheng:2021:MLN

[843] Jin Cheng, Yulei Wu, Yuepeng E, Junling You, Tong Li, Hui Li, and Jingguo Ge. MATEC: a lightweight neural network for online encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004217.

Araujo:2021:HOM

[844] Samuel M. A. Araújo, Fernanda S. H. de Souza, and Geraldo R. Mateus. A hybrid optimization-machine learning approach for the VNF placement and chaining problem. Computer Networks (Amsterdam, Netherlands: 1999), 199 (??):??, November 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004229.

Mall:2021:CPB

[845] Priyanka Mall, Ruhul Amin, Mohammad S. Obaidat, and Kuei-Fang Hsiao. CoMSeC++: PUF-based secured lightweight mutual authentication protocol for drone-enabled WSN. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004230.

Samaila:2021:PES

[846] Musa G. Samaila, Carolina Lopes, Édi Aires, João B. F. Sequeiros, Tiago Simões, Mário M. Freire, and Pedro R. M. Inácio. Performance evaluation of the SRE and SBPG components of the IoT hardware platform security advisor framework. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004370.

Anonymous:2021:EBp

[847] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621004679.

Anonymous:2021:N

[848] Anonymous. 9 November 2021. Computer Networks (Amsterdam, Netherlands: 1999), 199(??):??, November 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Aljeri:2021:EES

[849] Noura Aljeri and Azzedine Boukerche. EDiPSo: an efficient scalable topology discovery protocol for softwaredefined vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621003388.

Jung:2021:RTD

[850] SungHoon Jung and SaeWoong Bahk. Real-time dual-link transmission control for minimizing power and link switching cost. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004187.

Montero:2021:PRQ

[851] Emanuel Montero, Carlos Rocha, Helder Oliveira, Eduardo Cerqueira, Paulo Mendes, Aldri Santos, and Denis Rosário. Proactive radio- and QoSaware UAV as BS deployment to improve cellular operations. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100431X.

Robinson:2021:SCC

[852] Peter Robinson. Survey of crosschain communications protocols. Computer Networks (Amsterdam, Netherlands:

1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004321.

Garrett:2021:HAL

[853] Thiago Garrett, Luis C. E. Bona, and Elias P. Duarte. A holistic approach for locating traffic differentiation in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004333.

Moosavi:2021:EET

[854] Reza Moosavi, Saeedeh Parsaeefard, Mohammad Ali Maddah-Ali, Vahid Shah-Mansouri, Babak Hossein Khalaj, and Mehdi Bennis. Energy efficiency through joint routing and function placement in different modes of SDN/NFV networks. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004345.

Wang:2021:DGV

[855] Shengling Wang, Yu Pu, Hongwei Shi, Jianhui Huang, and Yinhao Xiao. A differential game view of antagonistic dynamics for cybersecurity. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004357.

Pillai:2021:BCA

[856] Babu Pillai, Kamanashis Biswas, Zhé Hóu, and Vallipuram Muthukkumarasamy. Burn-to-claim: an asset transfer protocol for blockchain interoperability. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004369.

Batchu:2021:GML

[857] Raj Kumar Batchu and Hari Seetha. A generalized machine learning model for DDoS attacks detection using hybrid feature selection and hyperparameter tuning. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004394.

Zaabar:2021:HSB

[858] Bessem Zaabar, Omar Cheikhrouhou, Faisal Jamil, Meryem Ammi, and Mohamed Abid. HealthBlock: a secure blockchain-based healthcare data management system. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004382.

Tamang:2021:ARS

[859] Dinesh Tamang, Sergio Martiradonna, Andrea Abrardo, Gianluca Mandó, Gabriele Roncella, and Gennaro Boggia. Architecting 5G RAN slicing for location aware vehicle to infrastructure communications: the Autonomous Tram use case. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004400.

Streit:2021:NAD

[860] Ananda Streit, Gustavo H. A. Santos, Rosa M. M. Leão, Edmundo de Souza e Silva, Daniel Menasché, and Don Towsley. Network anomaly detection based on tensor decomposition. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004412.

Pal:2021:BBT

[861] Shantanu Pal, Ambrose Hill, Tahiry Rabehaja, and Michael Hitchens. A blockchain-based trust management framework with verifiable interactions. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004436.

Yuan:2021:CHR

[862] Peiyan Yuan, Saike Shao, Lijuan Geng, and Xiaoyan Zhao. Caching hit ratio maximization in mobile edge computing with node cooperation. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004424.

Sharmila:2021:EIA

[863] A. Helen Sharmila and N. Jaisankar. Edge intelligent agent assisted hybrid hierarchical blockchain for continuous healthcare monitoring and recommendation system in 5G WBAN-IoT. Computer Networks (Amsterdam, Netherlands: 1999), 200(??): ??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004448.

Lekssays:2021:PBB

[864] Ahmed Lekssays, Luca Landa, Barbara Carminati, and Elena Ferrari. PAutoBotCatcher: a blockchain-based privacy-preserving botnet detector for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100445X.

Chen:2021:VCA

[865] Ning Chen, Sheng Zhang, Siyi Quan, Zhi Ma, Zhuzhong Qian, and Sanglu Lu. VCMaker: Content-aware configuration adaptation for video streaming and analysis in live augmented reality. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004461.

Ma:2021:ISC

[866] Huihui Ma, Du Xu, Yueyue Dai, and Qing Dong. An intelligent scheme for congestion control: When active queue management meets deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004485.

Zhang:2021:QSE

[867] Jianwei Zhang and Chenwei Zhao. Q-SR: an extensible optimization framework for segment routing. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004497.

Fu:2021:TRE

[868] Xiuwen Fu, Pasquale Pace, Gianluca Aloi, Wenfeng Li, and Giancarlo Fortino. Toward robust and energy-efficient clustering wireless sensor networks: a double-stage scale-free topology evolution model. Computer Networks (Amsterdam, Netherlands:

1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004527.

Su:2021:GBD

[869] Yi Su, Wenhao Fan, Yuan’an Liu, and Fan Wu. Game-based distributed pricing and task offloading in multi-cloud and multi-edge environments. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004539.

Montieri:2021:PLP

[870] Antonio Montieri, Giampaolo Bovenzi, Giuseppe Aceto, Domenico Ciuonzo, Valerio Persico, and Antonio Pescapè. Packet-level prediction of mobile-app traffic using multitask deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 200(??): ??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004564.

Ujjwal:2021:DAH

[871] Ujjwal, Neha Kumari, Jaisingh Thangaraj, and Chhandita Roy. Distance adaptive hybrid super-channels enabled by sliceable bandwidth variable transponder for spectrally efficient elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004588.

Wang:2021:BSS

[872] Pan Wang, Zixuan Wang, Feng Ye, and Xuejiao Chen. ByteSGAN: a semisupervised generative adversarial network for encrypted traffic classification in SDN edge gateway. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100459X.

Powell:2021:PPS

[873] Warwick Powell, Shoufeng Cao, Thomas Miller, Marcus Foth, Xavier Boyen, Barry Earsman, Santiago del Valle, and Charles Turner-Morris. From premise to practice of social consensus: How to agree on common knowledge in blockchain-enabled supply chains. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004606.

Zhan:2021:WFE

[874] Pengwei Zhan, Liming Wang, and Yi Tang. Website fingerprinting on early QUIC traffic. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004618.

Tang:2021:STM

[875] Xinyu Tang, Cheng Guo, Kim-Kwang Raymond Choo, Yining Liu, and Long Li. A secure and trustworthy medical record sharing scheme based on searchable encryption and blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100462X.

Wu:2021:TEF

[876] Yulei Wu, Laizhong Cui, Victor C. M. Leung, Tarik Taleb, and Sangheon Pack. Towards efficient and flexible management and interworking techniques for Industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 200 (??):??, December 9, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004631.

Shusterman:2021:CBC

[877] Anatoly Shusterman, Chen Finkelstein, Ofir Gruner, Yarin Shani, and Yossi Oren. Cache-based characterization: a low-infrastructure, distributed alternative to network-based traffic and application characterization. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004710.

Anonymous:2021:EBq

[878] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621005028.

Anonymous:2021:Da

[879] Anonymous. 9 December 2021. Computer Networks (Amsterdam, Netherlands: 1999), 200(??):??, December 9, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mao:2021:DLD

[880] Qian Mao, Lin Zhang, Fei Hu, Elizabeth Serena Bentley, and Sunil Kumar. Deep learning (DL)-based adaptive transport layer control in UAV swarm networks. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004473.

Akter:2021:CSC

[881] Rubina Akter, Van-Sang Doan, Jae-Min Lee, and Dong-Seong Kim. CNN-SSDI: Convolution neural network inspired surveillance system for UAVs detection and identification. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004503.

Zhao:2021:SIE

[882] Yuyu Zhao, Guang Cheng, Yu Duan, Zhouchao Gu, Yuyang Zhou, and Lu Tang. Secure IoT edge: Threat situation awareness based on network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004540.

Ma:2021:GAP

[883] Baolin Ma, Chao Yang, Mingzhe Chen, and Jianfeng Ma. GramMatch: an automatic protocol feature extraction and identification system. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004552.

Yu:2021:IAI

[884] Chuan Yu, Shuhui Chen, Fei Wang, and Ziling Wei. Improving 4G/ 5G air interface security: a survey of existing attacks on different LTE layers. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004576.

Zhang:2021:DSS

[885] Zhishuo Zhang and Shijie Zhou. A decentralized strongly secure attributebased encryption and authentication scheme for distributed Internet of mobile things. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004722.

Candela:2021:WSG

[886] Massimo Candela, Valerio Luconi, and Alessio Vecchio. A worldwide study on the geographic locality of Internet routes. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004734.

Palma:2021:PPC

[887] Noelia Pérez Palma, Falko Dressler, and Vincenzo Mancuso. Precise: Predictive content dissemination scheme exploiting realistic mobility patterns. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004746.

Torre:2021:PEM

[888] Roberto Torre, Muhammad Tayyab, George Koudouridis, Xavier Gelabert, Riccardo Bassoli, and Frank H. P. Fitzek. Power efficient mobile small cell placement for network-coded cooperation in UDNs. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004515.

Davami:2021:DSM

[889] Fatemeh Davami, Sahar Adabi, Ali Rezaee, and Amir Masoud Rahmani. Distributed scheduling method for multiple workflows with parallelism prediction and DAG prioritizing for time constrained cloud applications. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100476X.

Jie:2021:OUF

[890] Lu Jie, Chen Hongchang, Sun Penghao, Hu Tao, and Zhang Zhen. OrderSketch: an unbiased and fast sketch for frequency estimation of data streams. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004795.

Atif:2021:SNM

[891] Syed Muhammad Atif, Nicolas Gillis, Sameer Qazi, and Imran Naseem. Structured nonnegative matrix factorization for traffic flow estimation of large cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004771.

Wang:2021:EEC

[892] Bin Wang, Fagui Liu, Weiwei Lin, Zhenjiang Ma, and Dishi Xu. Energyefficient collaborative optimization for VM scheduling in cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004783.

Salahdine:2021:SSM

[893] Fatima Salahdine, Johnson Opadere, Qiang Liu, Tao Han, Ning Zhang, and Shaohua Wu. A survey on sleep mode techniques for ultra-dense networks in 5G and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004801.

Farhadi:2021:FSM

[894] Babak Farhadi, Amir Masoud Rahmani, Parvaneh Asghari, and Mehdi Hosseinzadeh. Friendship selection and management in social internet of things: a systematic review. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004813.

Garetto:2021:CPN

[895] Michele Garetto, Emilio Leonardi, and Giovanni Neglia. Content placement in networks of similarity caches. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004825.

Qian:2021:DMN

[896] Bing Qian and Shun Lu. Detection of mobile network abnormality using deep learning models on massive network measurement data. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004837.

Liang:2021:CUD

[897] Shuang Liang, Zhiyi Fang, Geng Sun, Chi Lin, Jiahui Li, Songyang Li, and Aimin Wang. Charging UAV deployment for improving charging performance of wireless rechargeable sensor networks via joint optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004849.

Braeken:2021:DDG

[898] An Braeken. Device-to-device group authentication compatible with 5G AKA protocol. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004850.

Yang:2021:SHS

[899] Ye Yang, Haiyang Jiang, Guangxing Zhang, Xin Wang, Yilong Lv, Xing Li, Serge Fdida, and Gaogang Xie. S2H: Hypervisor as a setter within Virtualized Network I/O for VM isolation on cloud platform. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004862.

Sari:2021:CRC

[900] T. Tolga Sari and Gökhan Seçinti. Chain RTS/CTS scheme with opportunistic channel allocation. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004874.

Abid:2021:SRC

[901] Khaled Abid, Hicham Lakhlef, and Abdelmadjid Bouabdallah. A survey on recent contention-free MAC protocols for static and mobile wireless decentralized networks in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004886.

Sharma:2021:POU

[902] Abhishek Sharma, Amik Garg, Sanjay Kumar Sharma, Vibhav Kumar Sachan, and Parvin Kumar. Performance optimization for UWB communication network under IEEE 802.15.4a channel conditions. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621004904.

Ren:2021:AIB

[903] Qiuning Ren, Chao Yang, and Jianfeng Ma. App identification based on encrypted multi-smartphone sources traffic fingerprints. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100493X.

Wazirali:2021:SAD

[904] Raniyah Wazirali, Rami Ahmad, and Ashraf Abdel-Karim Abu-Ein. Sustaining accurate detection of phishing URLs using SDN and feature selection approaches. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004941.

Liu:2021:PLP

[905] Yuan Liu, Yixiao Lan, Boyang Li, Chunyan Miao, and Zhihong Tian. Proof of learning (PoLe): Empowering neural network training with consensus building on blockchains. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004965.

Kumar:2021:DMP

[906] Sanjeev Kumar and B. Janet. Distinguishing malicious programs based on visualization and hybrid learning algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004953.

Shinde:2021:NOB

[907] Swapnil Sadashiv Shinde, Dania Marabissi, and Daniele Tarchi. A network operator-biased approach for multi-service network function placement in a 5G network slicing architecture. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004989.

Miuccio:2021:DBE

[908] Luciano Miuccio, Daniela Panno, and Salvatore Riolo. A DNN-based estimate of the PRACH traffic load for massive IoT scenarios in 5G networks and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005077.

Rodrigues:2021:ABI

[909] Carlo Kleber da Silva Rodrigues. Analyzing blockchain integrated architectures for effective handling of IoTecosystem transactions. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005089.

Dhooge:2021:HFB

[910] Laurens D’hooge, Miel Verkerken, Tim Wauters, Bruno Volckaert, and Filip De Turck. Hierarchical feature block ranking for data-efficient intrusion detection modeling. Computer Networks (Amsterdam, Netherlands: 1999), 201 (??):??, December 24, 2021. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005107.

Anonymous:2021:EBr

[911] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128621005272.

Anonymous:2021:Db

[912] Anonymous. 24 December 2021. Computer Networks (Amsterdam, Netherlands: 1999), 201(??):??, December 24, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Min:2022:BNM

[913] Xinping Min, Lanju Kong, Qingzhong Li, Yuan Liu, Baochen Zhang, Yongguang Zhao, Zongshui Xiao, and Bin Guo. Blockchain-native mechanism supporting the circulation of complex physical assets. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621004928.

Guarino:2022:AMP

[914] Alfonso Guarino, Delfina Malandrino, and Rocco Zaccagnino. An automatic mechanism to provide privacy awareness and control over unwittingly dissemination of online private information. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005090.

Horchulhack:2022:TFM

[915] Pedro Horchulhack, Eduardo K. Viegas, and Altair O. Santin. Toward feasible machine learning model updates in network-based intrusion detection. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005120.

Ruckel:2022:FIP

[916] Timon Rückel, Johannes Sedlmeir, and Peter Hofmann. Fairness, integrity, and privacy in a scalable blockchain-based federated learning system. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005132.

Banerjee:2022:EED

[917] Anuradha Banerjee, Abu Sufian, Krishna Keshob Paul, and Sachin Kumar Gupta. EDTP: Energy and delay optimized trajectory planning for UAV-IoT environment. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005156.

Aureli:2022:ADO

[918] Davide Aureli, Antonio Cianfrani, Marco Listanti, Marco Polverini, and Stefano Secci. Augmenting DiffServ operations with dynamically learned classes of services. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005144.

He:2022:JOS

[919] Yexiao He, Xiaoning Zhang, Zixiang Xia, Yutao Liu, Keshav Sood, and Shui Yu. Joint optimization of service chain graph design and mapping in NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005168.

Goldoni:2022:CBW

[920] Emanuele Goldoni, Pietro Savazzi, Lorenzo Favalli, and Anna Vizziello. Correlation between weather and signal strength in LoRaWAN networks: an extensive dataset. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100517X.

Zhang:2022:DDA

[921] Chunyang Zhang, Gaogang Xie, and Xin Wang. DynamicTuple: the dynamic adaptive tuple for high-performance packet classification. Computer Networks (Amsterdam, Netherlands: 1999), 202(??): ??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005193.

Yang:2022:RWO

[922] Furong Yang, Andrea Ferlini, Davide Aguiari, Davide Pesavento, Rita Tse, Suman Banerjee, Gaogang Xie, and Giovanni Pau. Revisiting WiFi offloading in the wild for V2I applications. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005211.

Perna:2022:ROS

[923] Gianluca Perna, Dena Markudova, Martino Trevisan, Paolo Garza, Michela Meo, and Maurizio M. Munafò. Retina: an open-source tool for flexible analysis of RTC traffic. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005235.

Nguyen:2022:CAO

[924] Viet Dung Nguyen, Phi Le Nguyen, Kien Nguyen, and Phan Thuan Do. Constant approximation for opportunistic sensing in mobile air quality monitoring system. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005314.

Zeng:2022:UIO

[925] Man Zeng, Xiaohong Huang, Pei Zhang, and Dandan Li. Understanding the impact of outsourcing mitigation against BGP prefix hijacking. Computer Networks (Amsterdam, Netherlands: 1999), 202 (??):??, January 15, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005338.

Lin:2022:MML

[926] Kunda Lin, Xiaolong Xu, and Fu Xiao. MFFusion: a multi-level features fusion model for malicious traffic detection based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005399.

Dutta:2022:DRL

[927] Hrishikesh Dutta and Subir Biswas. Distributed reinforcement learning for scalable wireless medium access in IoTs and sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005417.

Anonymous:2022:EBa

[928] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128621005788.

Anonymous:2022:Ja

[929] Anonymous. 15 January 2022. Computer Networks (Amsterdam, Netherlands: 1999), 202(??):??, January 15, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hbaieb:2022:STM

[930] Amal Hbaieb, Samiha Ayed, and Lamia Chaari. A survey of trust management in the Internet of vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004758.

Carpio:2022:SMR

[931] Francisco Carpio, Wolfgang Bziuk, and Admela Jukan. Scaling migrations and replications of Virtual Network Functions based on network traffic forecasting. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004898.

Zhang:2022:BBP

[932] Guipeng Zhang, Zhenguo Yang, and Wenyin Liu. Blockchain-based privacy preserving e-health system for healthcare data in cloud. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004916.

Kumar:2022:HSH

[933] Alok Kumar, Prabhat Thakur, Shweta Pandit, and Ghanshyam Singh. HSA-SPC: Hybrid spectrum access with spectrum prediction and cooperation for performance enhancement of multiuser cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621004977.

Wang:2022:LLT

[934] Zumin Wang, Jiyu Tian, Hui Fang, Liming Chen, and Jing Qin. Light-Log: a lightweight temporal convolutional network for log anomaly detection on the edge. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005119.

Bodet:2022:DST

[935] Duschia Bodet, Jacob Hall, Priyangshu Sen, Rachel Johnson, Isabelle Brandicourt, Xavier Cantos Roman, Omar Shoura, and Josep Miquel Jornet. Data signals for terahertz communications research. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005181.

Tang:2022:ILP

[936] Rui Tang, Xingshu Chen, Chuancheng Wei, Qindong Li, Wenxian Wang, Haizhou Wang, and Wei Wang. Interlayer link prediction based on multiple network structural attributes. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862100534X.

Momeni:2022:EEN

[937] M. Momeni and H. S. Shahhoseini. Energy efficient 3D network-on-chip based on approximate communication. Computer Networks (Amsterdam, Netherlands: 1999), 203(??): ??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005351.

Kathiravelu:2022:TIS

[938] Pradeeban Kathiravelu, Zachary Zaiman, Judy Gichoya, Luís Veiga, and Imon Banerjee. Towards an Internetscale overlay network for latencyaware decentralized workflows at the edge. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005363.

Gundogan:2022:MCP

[939] Cenk Gündogan, Peter Kietzmann, Thomas C. Schmidt, and Matthias Wählisch. A mobility-compliant publish-subscribe system for an informationcentric Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005375.

Wang:2022:SRS

[940] Suixing Wang, Chao Yang, Gang Guo, Mingzhe Chen, and Jianfeng Ma. SSAPPIDENTIFY: a robust system identifies application over Shadowsocks’s traffic. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005387.

Campos:2022:EFL

[941] Enrique Mármol Campos, Pablo Fernández Saura, Aurora González-Vidal, José L. Hernández-Ramos, Jorge Bernal Bernabé, Gianmarco Baldini, and Antonio Skarmeta. Evaluating federated learning for intrusion detection in Internet of Things: Review and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005405.

Xia:2022:PIE

[942] Qiufen Xia, Lizhen Zhou, Wenhao Ren, and Yi Wang. Proactive and intelligent evaluation of big data queries in edge clouds with materialized views. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005429.

Yang:2022:GTA

[943] Tao Yang, Bingnan Hou, Zhiping Cai, Kui Wu, Tongqing Zhou, and Chengyu Wang. 6Graph: a graphtheoretic approach to address pattern mining for Internet-wide IPv6 scanning. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005430.

Hireche:2022:DDP

[944] Othmane Hireche, Chafika Benzaïd, and Tarik Taleb. Deep data plane programming and AI for zero-trust self-driven networking in beyond 5G. Computer Networks (Amsterdam, Netherlands: 1999), 203(??): ??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005442.

Zhang:2022:OPB

[945] Kaiyuan Zhang, Xiaolin Gui, Dewang Ren, Tianjiao Du, and Xin He. Optimal pricing-based computation offloading and resource allocation for blockchain-enabled beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005478.

Baker:2022:BBF

[946] Thar Baker, Muhammad Asim, Hezekiah Samwini, Nauman Shamim, Mohammed M. Alani, and Rajkumar Buyya. A blockchain-based fogoriented lightweight framework for smart public vehicular transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100548X.

Febro:2022:ESS

[947] Aldo Febro, Hannan Xiao, Joseph Spring, and Bruce Christianson. Edge security for SIP-enabled IoT devices with P4. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005612.

Bonati:2022:EAE

[948] Leonardo Bonati, Stefano Basagni, and Tommaso Melodia. Editorial: Advances in experimental wireless platforms and systems. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005624.

Luo:2022:FIA

[949] Wanming Luo, Shihao Liu, Yihao Jia, Zhe Chen, and Sheng Jiang. Flexible IP: an adaptable IP address structure and its efficient addressing scheme. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005636.

Khatiri:2022:BRA

[950] Atefeh Khatiri, Ghasem Mirjalily, and Zhi-Quan Luo. Balanced resource allocation for VNF service chain provisioning in inter-datacenter elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 203 (??):??, February 11, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005727.

Anonymous:2022:EBb

[951] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622000123.

Anonymous:2022:Fa

[952] Anonymous. 11 February 2022. Computer Networks (Amsterdam, Netherlands: 1999), 203(??):??, February 11, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Saltini:2022:BRO

[953] Roberto Saltini. BigFooT: a robust optimal-latency BFT blockchain consensus protocol with dynamic validator membership. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100520X.

Griner:2022:CLP

[954] Chen Griner, Stefan Schmid, and Chen Avin. CacheNet: Leveraging the principle of locality in reconfigurable network design. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005326.

Wan:2022:PPB

[955] Yichen Wan, Youyang Qu, Longxiang Gao, and Yong Xiang. Privacypreserving blockchain-enabled federated learning for B5G-driven edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005454.

Guo:2022:EHC

[956] Mian Guo, Qirui Li, Zhiping Peng, Xiushan Liu, and Delong Cui. Energy harvesting computation offloading game towards minimizing delay for mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005491.

Salman:2022:TER

[957] Ola Salman, Imad H. Elhajj, Ali Chehab, and Ayman Kayssi. Towards efficient real-time traffic classifier: a confidence measure with ensemble deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005521.

Aski:2022:ANE

[958] Vidyadhar Jinnappa Aski, Vijaypal Singh Dhaka, Sunil Kumar, Sahil Verma, and Danda B. Rawat. Advances on networked ehealth information access and sharing: Status, challenges and prospects. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005545.

Wang:2022:OCO

[959] Jian Wang, Hongchang Ke, Xuejie Liu, and Hui Wang. Optimization for computational offloading in multiaccess edge computing: a deep reinforcement learning scheme. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005569.

Abubaker:2022:BSP

[960] Zain Abubaker, Nadeem Javaid, Ahmad Almogren, Mariam Akbar, Mansour Zuair, and Jalel Ben-Othman. Blockchained service provisioning and malicious node detection via federated learning in scalable Internet of Sensor Things networks. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005570.

Rey:2022:FLM

[961] Valerian Rey, Pedro Miguel Sánchez Sánchez, Alberto Huertas Celdrán, and Gérôme Bovet. Federated learning for malware detection in IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005582.

Sengupta:2022:VPP

[962] Binanda Sengupta. VALNET: Privacy-preserving multi-path validation. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005594.

Dien:2022:TSN

[963] Nguyen V. Dien, Nguyen V. Tuan, Le T. P. Mai, Nguyen V. Hieu, Vuong Q. Phuoc, Nguyen Q. N. Quynh, and Nguyen T. Hung. Tolerance of SCM Nyquist and OFDM signals for heterogeneous fiber-optic and millimeter-wave mobile backhaul links under the effect of power amplifier saturation induced clipping. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005600.

Hameed:2022:TFM

[964] Khizar Hameed, Saurabh Garg, Muhammad Bilal Amin, and Byeong Kang. Towards a formal modelling, analysis and verification of a clone node attack detection scheme in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005648.

Cheema:2022:BBS

[965] Muhammad Asaad Cheema, Rafay Iqbal Ansari, Nouman Ashraf, Syed Ali Hassan, Hassaan Khaliq Qureshi, Ali Kashif Bashir, and Christos Politis. Blockchain-based secure delivery of medical supplies using drones. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005661.

Zhou:2022:TSD

[966] Anan Zhou, Benshun Yi, and Laigan Luo. Tree-structured data placement scheme with cluster-aided topdown transmission in erasure-coded distributed storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005703.

Wu:2022:OMT

[967] Zheng Wu, Yu ning Dong, Xiaohui Qiu, and Jiong Jin. Online multimedia traffic classification from the QoS perspective using deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005715.

Aouini:2022:NFN

[968] Zied Aouini and Adrian Pekar. NF-Stream: a flexible network data analysis framework. Computer Networks (Amsterdam, Netherlands: 1999), 204 (??):??, February 26, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005739.

Zheng:2022:EPM

[969] Kechen Zheng, Haijiang Ge, Kaikai Chi, and Xiaoying Liu. Energy provision minimization of energy-harvesting cognitive radio networks with minimal throughput demands. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005740.

Mei:2022:RMB

[970] Lifan Mei, Jinrui Gou, Yujin Cai, Houwei Cao, and Yong Liu. Realtime mobile bandwidth and handoff predictions in 4G/5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005879.

Anonymous:2022:EBc

[971] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622000366.

Anonymous:2022:Fb

[972] Anonymous. 26 February 2022. Computer Networks (Amsterdam, Netherlands: 1999), 204(??):??, February 26, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Nasser:2022:LFL

[973] Nidal Nasser, Zubair Md Fadlullah, Mostafa M. Fouda, Asmaa Ali, and Muhammad Imran. A lightweight federated learning based privacy preserving B5G pandemic response network using unmanned aerial vehicles: a proof-of-concept. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005466.

Li:2022:CAS

[974] Wenhao Li, Hamid Faragardi, Mustafa Ozger, Cicek Cavdar, and Björn Skubic. Cost aware service selection in a mobile edge marketplace. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005508.

Kuk:2022:APA

[975] Ekber Çetin Kük and Müge Erel-Özçevik. Access protocol aware controller design for eMBB traffic in SD-CDN. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005533.

Bagies:2022:CDN

[976] Enas Bagies, Ahmed Barnawi, Saoucene Mahfoudh, and Neeraj Kumar. Content delivery network for IoT-based fog computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005557.

Cao:2022:APM

[977] Tuo Cao, Yibo Jin, Xiongfeng Hu, Sheng Zhang, Zhuzhong Qian, Baoliu Ye, and Sanglu Lu. Adaptive provisioning for mobile cloud gaming at edges. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100565X.

Nasir:2022:SII

[978] Muhammad Hassan Nasir, Salman A. Khan, Muhammad Mubashir Khan, and Mahawish Fatima. Swarm intelligence inspired intrusion detection systems — a systematic literature review. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005673.

Lin:2022:NLB

[979] Jie Lin, Lin Huang, Hanlin Zhang, Xinyu Yang, and Peng Zhao. A novel Lyapunov based dynamic resource allocation for UAVs-assisted edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005685.

Casar:2022:SBL

[980] Matthias Cäsar, Tobias Pawelke, Jan Steffan, and Gabriel Terhorst. A survey on Bluetooth low energy security and privacy. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005697.

Wang:2022:COR

[981] Zhongyu Wang, Tiejun Lv, and Zheng Chang. Computation offloading and resource allocation based on distributed deep learning and software defined mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005843.

deCarvalho:2022:JJT

[982] Márcio Barbosa de Carvalho and Lisandro Zambenedetti Granville. JurisNN: Judging traffic differentiations as network neutrality violations according to the regulation. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005880.

Aumayr:2022:SBA

[983] Erik Aumayr, Giuseppe Caso, Anne-Marie Bosneag, Almudena Diaz Zayas, Özgü Alay, Bruno Garcia, Konstantinos Kousias, Anna Brünstrom, Pedro Merino Gomez, and Harilaos Koumaras. Service-based analytics for 5G open experimentation platforms. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005892.

Salehi:2022:NMI

[984] Majid Salehi, Gilles De Borger, Danny Hughes, and Bruno Crispo. Nemesis-Guard: Mitigating interrupt latency side channel attacks with static binary rewriting. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005910.

Wang:2022:BDM

[985] Yang Wang, Marco Giordani, Xiangming Wen, and Michele Zorzi. On the beamforming design of millimeter wave UAV networks: Power vs. capacity trade-offs. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005922.

Liu:2022:EOD

[986] Linfeng Liu, Zhipeng Zhang, Jiagao Wu, and Jia Xu. Entropy optimization of degree distributions against security threats in UASNs. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005934.

Zhong:2022:PBD

[987] Hong Zhong, Jinshan Xu, Jie Cui, Xiuwen Sun, Chengjie Gu, and Lu Liu. Prediction-based dual-weight switch migration scheme for SDN load balancing. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005946.

Santos:2022:TOO

[988] Ricardo Santos, Nina Skorin-Kapov, Hakim Ghazzai, Andreas Kassler, and Gia Khanh Tran. Towards the optimal orchestration of steerable mmWave backhaul reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621005958.

Zahedi:2022:PEP

[989] Seyed Reza Zahedi, Shahram Jamali, and Peyman Bayat. A power-efficient and performance-aware online virtual network function placement in SDN/ NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005971.

Floris:2022:SIB

[990] Alessandro Floris, Simone Porcu, Luigi Atzori, and Roberto Girau. A social IoT-based platform for the deployment of a smart parking solution. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005995.

Cortesi:2022:NAB

[991] Eugenio Cortesi, Francesco Bruschi, Stefano Secci, and Sami Taktak. A new approach for Bitcoin pool-hopping detection. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621006009.

Liu:2022:MMB

[992] Liang Liu, Lin Liu, Beibei Li, Yi Zhong, Shan Liao, and Lei Zhang. MSCCS: a monero-based securityenhanced covert communication system. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621006010.

Wang:2022:NTA

[993] Chonghua Wang, Hao Zhou, Zhiqiang Hao, Shu Hu, Jun Li, Xueying Zhang, Bo Jiang, and Xuehong Chen. Network traffic analysis over clustering-based collective anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000019.

Liu:2022:CNT

[994] Meili Liu, Xiaogang Qi, and Hao Pan. Construction of network topology and geographical vulnerability for telecommunication network. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000032.

Mukta:2022:SDM

[995] Rahma Mukta, Hye young Paik, Qinghua Lu, and Salil S. Kanhere. A survey of data minimisation techniques in blockchain-based healthcare. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000044.

Fronteddu:2022:DHG

[996] Graziano Fronteddu, Simone Porcu, Alessandro Floris, and Luigi Atzori. A dynamic hand gesture recognition dataset for human-computer interfaces. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000172.

Barcelo-Armada:2022:AAT

[997] Rubén Barceló-Armada, Ismael Castell-Uroz, and Pere Barlet-Ros. Amazon Alexa traffic traces. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000184.

Pintor:2022:DLD

[998] Lucia Pintor and Luigi Atzori. A dataset of labelled device Wi-Fi probe requests for MAC address derandomization. Computer Networks (Amsterdam, Netherlands: 1999), 205 (??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000196.

Anonymous:2022:EBd

[999] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622000573.

Anonymous:2022:Ma

[1000] Anonymous. 14 March 2022. Computer Networks (Amsterdam, Netherlands: 1999), 205(??):??, March 14, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

deSousa:2022:PLS

[1001] Roniel S. de Sousa, Azzedine Boukerche, and Antonio A. F. Loureiro. On the prediction of large-scale roadnetwork constrained trajectories. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128621003352.

Wang:2022:SPL

[1002] Xiaoyang Wang, Jonathan D. Thomas, Robert J. Piechocki, Shipra Kapoor, Raúl Santos-Rodríguez, and Arjun Parekh. Self-play learning strategies for resource assignment in Open-RAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862100551X.

Hosseini:2022:OTS

[1003] Entesar Hosseini, Mohsen Nickray, and Shamsollah Ghanbari. Optimized task scheduling for cost-latency trade-off in mobile fog computing using fuzzy analytical hierarchy process. Computer Networks (Amsterdam, Netherlands:

1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862100596X.

Yang:2022:PSP

[1004] Yan Yang, Xingang Shi, Qiang Ma, Yahui Li, Xia Yin, and Zhiliang Wang. Path stability in partially deployed secure BGP routing. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000020.

Choi:2022:UIT

[1005] Jinchun Choi, Afsah Anwar, Abdulrahman Alabduljabbar, Hisham Alasmary, Jeffrey Spaulding, An Wang, Songqing Chen, DaeHun Nyang, Amro Awad, and David Mohaisen. Understanding Internet of Things malware by analyzing endpoints in their static artifacts. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000056.

Dahanayaka:2022:DTF

[1006] Thilini Dahanayaka, Guillaume Jourjon, and Suranga Seneviratne. Dissecting traffic fingerprinting CNNs with filter activations. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000068.

Ahanger:2022:SAS

[1007] Tariq Ahamed Ahanger, Abdullah Aljumah, and Mohammed Atiquzzaman. State-of-the-art survey of artificial intelligent techniques for IoT security. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200007X.

Raskar:2022:MEM

[1008] Charushila Raskar and Shikha Nema. Metaheuristic enabled modified hidden Markov model for traffic flow prediction. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000160.

Madarasingha:2022:VGB

[1009] Chamara Madarasingha, Shashika R. Muramudalige, Guillaume Jourjon, Anura Jayasumana, and Kanchana Thilakarathna. VideoTrain++: GANbased adaptive framework for synthetic video traffic generation. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000202.

Silva:2022:TLD

[1010] Giovanni Maciel Ferreira Silva and Taufik Abrão. Throughput and latency in the distributed Q-learning random access mMTC networks. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000214.

Yang:2022:BTB

[1011] Furong Yang, Qinghua Wu, Zhenyu Li, Yanmei Liu, Giovanni Pau, and Gaogang Xie. BBRv2+: Towards balancing aggressiveness and fairness with delay-based bandwidth probing. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000226.

Zhang:2022:SBP

[1012] Lihao Zhang, Taotao Wang, and Soung Chang Liew. Speeding up block propagation in Bitcoin network: Uncoded and coded designs. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000238.

Xiong:2022:BBP

[1013] Ting Xiong, Ran Zhang, Jiang Liu, Tao Huang, Yunjie Liu, and F. Richard Yu. A blockchain-based and privacy-preserved authentication scheme for inter-constellation collaboration in Space-Ground Integrated Networks. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200024X.

Sabur:2022:TSG

[1014] Abdulhakim Sabur, Ankur Chowdhary, Dijiang Huang, and Adel Alshamrani. Toward scalable graph-based security analysis for cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000251.

Brehon-Grataloup:2022:MEC

[1015] Lucas Bréhon-Grataloup, Rahim Kacimi, and André-Luc Beylot. Mobile edge computing for V2X architectures and applications: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000263.

Sciancalepore:2022:PPP

[1016] Savio Sciancalepore. PARFAIT: Privacypreserving, secure, and low-delay service access in fog-enabled IoT ecosystems. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000275.

Deb:2022:CSV

[1017] Raktim Deb and Sudipta Roy. A comprehensive survey of vulnerability and information security in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000299.

El-Zawawy:2022:SSB

[1018] Mohamed A. El-Zawawy, Alessandro Brighente, and Mauro Conti. SETCAP: Service-based energy-efficient temporal credential authentication protocol for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000305.

Vasudevan:2022:IAE

[1019] Vipindev Adat Vasudevan, Muhammad Tayyab, George P. Koudouridis, Xavier Gelabert, and Ilias Politis. An integrated approach for energy efficient handover and key distribution protocol for secure NC-enabled small cells. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000317.

Kim:2022:CFO

[1020] Hak-Jin Kim, Marie S. Kim, and Seung-Jae Han. Collision-free optimal packet scheduling algorithm for multi-hop wireless IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000408.

Magdy:2022:ESP

[1021] Safaa Magdy, Yasmine Abouelseoud, and Mervat Mikhail. Efficient spam and phishing emails filtering based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 206 (??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000469.

Sicari:2022:SIC

[1022] Sabrina Sicari, Alessandra Rizzardi, and Alberto Coen-Porisini. Security & privacy issues and challenges in NoSQL databases. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000470.

Anonymous:2022:EBe

[1023] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200086X.

Anonymous:2022:Aa

[1024] Anonymous. 7 April 2022. Computer Networks (Amsterdam, Netherlands: 1999), 206(??):??, April 7, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Aouedi:2022:HPL

[1025] Ons Aouedi, Kandaraj Piamrat, and Dhruvjyoti Bagadthey. Handling partially labeled network data: a semi-supervised approach using stacked sparse autoencoder. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128621005909.

AlSabeh:2022:SSA

[1026] Ali AlSabeh, Joseph Khoury, Elie Kfoury, Jorge Crichigno, and Elias Bou-Harb. A survey on security applications of P4 programmable switches and a STRIDE-based vulnerability assessment. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000287.

Guo:2022:SSR

[1027] Yimin Guo, Zhenfeng Zhang, and Yajun Guo. SecFHome: Secure remote authentication in fog-enabled smart home environment. Computer Networks (Amsterdam, Netherlands:

1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200041X.

Nassef:2022:SDM

[1028] Omar Nassef, Wenting Sun, Hakimeh Purmehdi, Mallik Tatipamula, and Toktam Mahmoodi. A survey: Distributed Machine Learning for 5G and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000421.

Makarem:2022:DEC

[1029] Nabil Makarem, Wafaa Bou Diab, Imad Mougharbel, and Naceur Malouch. On the design of efficient congestion control for the Constrained Application Protocol in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000457.

Huang:2022:PLF

[1030] Cheng Huang, Aimin Tang, Bangzhao Zhai, and Xudong Wang. Physical layer forwarding for 5G multihop Backhaul networks. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000482.

Ribeiro:2022:SHT

[1031] Nildo dos Santos Ribeiro, Marcos A. M.

Vieira, Luiz F. M. Vieira, and Omprakash Gnawali. SplitPath: High throughput using multipath routing in dual-radio Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000494.

Nduwayezu:2022:LEA

[1032] Maurice Nduwayezu and Ji-Hoon Yun. Latency and energy aware rate maximization in MC-NOMA-based multi-access edge computing: a two-stage deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000500.

Shahraki:2022:CSO

[1033] Amin Shahraki, Mahmoud Abbasi, Amir Taherkordi, and Anca Delia Jurcut. A comparative study on online machine learning techniques for network traffic streams analysis. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000512.

Alghafari:2022:DJR

[1034] Hadeel Alghafari and Mohammad Sayad Haghighi. Decentralized joint resource allocation and path selection in multi-hop integrated access backhaul 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000524.

Mondal:2022:ENC

[1035] Sourav Mondal, Goutam Das, and Elaine Wong. An economic and noncooperative load-balancing framework among federated cloudlets. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000615.

Wang:2022:JFS

[1036] Zuyan Wang, Jun Tao, Yang Gao, Yifan Xu, Weice Sun, Yu Gao, and Wenqiang Li. Joint flight scheduling and task allocation for secure data collection in UAV-aided IoTs. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000627.

Bouzidi:2022:DCS

[1037] EL Hocine Bouzidi, Abdelkader Outtagarts, Rami Langar, and Raouf Boutaba. Dynamic clustering of software defined network switches and controller placement using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000640.

Zhang:2022:CSN

[1038] Ke Zhang, Guang Zhang, Xiuwu Yu, Shaohua Hu, and Moxiao Li. Clustering the sensor networks based on energy-aware affinity propagation. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000652.

Pei:2022:UAC

[1039] Jiahui Pei, Hongbin Chen, and Lei Shu. UAV-assisted connectivity enhancement algorithms for multiple isolated sensor networks in agricultural Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000664.

Tang:2022:UAM

[1040] Qiang Tang, Lixin Liu, Caiyan Jin, Jin Wang, Zhuofan Liao, and Yuansheng Luo. An UAV-assisted mobile edge computing offloading strategy for minimizing energy consumption. Computer Networks (Amsterdam, Netherlands:

1999), 207(??):??, April 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000688.

Conti:2022:SCA

[1041] Mauro Conti, Eleonora Losiouk, Radha Poovendran, and Riccardo Spolaor. Side-channel attacks on mobile and IoT devices for cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200069X.

Hyland-Wood:2022:GEB

[1042] David Hyland-Wood and Sandra Johnson. Guest editorial: Blockchain consensus protocols. Computer Networks (Amsterdam, Netherlands: 1999), 207 (??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000718.

Anonymous:2022:EBf

[1043] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622001050.

Anonymous:2022:Ab

[1044] Anonymous. 22 April 2022. Computer Networks (Amsterdam, Netherlands: 1999), 207(??):??, April 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Campos:2022:MLH

[1045] Pablo Campos, Ángela Hernández-Solana, and Antonio Valdovinos-Bardají. Machine learning for hidden nodes detection in unlicensed LTE networks. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200072X.

Enoch:2022:ISH

[1046] Simon Yusuf Enoch, Júlio Mendonça, Jin B. Hong, Mengmeng Ge, and Dong Seong Kim. An integrated security hardening optimization for dynamic networks using security and availability modeling with multi-objective algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000731.

Qu:2022:PAV

[1047] Hua Qu, Ke Wang, and Jihong Zhao. Priority-awareness VNF migration method based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000755.

Bhattacharjee:2022:CRB

[1048] Sangeeta Bhattacharjee, Tamaghna Acharya, and Uma Bhattacharya. Cognitive radio based spectrum sharing models for multicasting in 5G cellular networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000779.

Nguyen:2022:UCP

[1049] Thi Thuy Nga Nguyen, Olivier Brun, and Balakrishna J. Prabhu. Using channel predictions for improved proportional-fair utility for vehicular users. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000780.

Enoch:2022:PFC

[1050] Simon Yusuf Enoch, Chun Yong Moon, Donghwan Lee, Myung Kil Ahn, and Dong Seong Kim. A practical framework for cyber defense generation, enforcement and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000810.

Gharbaoui:2022:ESL

[1051] M. Gharbaoui, C. Contoli, G. Davoli, D. Borsatti, G. Cuffaro, F. Paganelli, W. Cerroni, P. Cappanera, and B. Martini. An experimental study on latencyaware and self-adaptive service chaining orchestration in distributed NFV and SDN infrastructures. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000822.

Islam:2022:ITE

[1052] Nazrul Islam and Chris Phillips. Intelligent traffic engineering (TE) system for rural broadband. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000901.

Yin:2022:JAD

[1053] Wei Yin, Peizhao Hu, Hongjian Zhou, Guoqiang Xing, and Jiahui Wen. Jamming attacks and defenses for fast association in IEEE 802.11ah networks. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000913.

Bian:2022:SLD

[1054] Rui Bian, Shuai Hao, Haining Wang, and Chase Cotton. Shining a light on dark places: a comprehensive analysis of open proxy ecosystem. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000937.

Li:2022:PSR

[1055] Ziyong Li, Yuxiang Hu, Jiangxing Wu, and Jie Lu. P4Resilience: Scalable resilience for multi-failure recovery in SDN with programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 208 (??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000950.

Mahmoudi:2022:QAR

[1056] Arezou Mahmoudi, Akbar Ghaffarpour Rahbar, and Mohammad Jafari-Beyrami. QoS-aware routing, space, and spectrum assignment in space division multiplexing networks. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001116.

Anonymous:2022:EBg

[1057] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001281.

Anonymous:2022:Mb

[1058] Anonymous. 8 May 2022. Computer Networks (Amsterdam, Netherlands: 1999), 208(??):??, May 8, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mangas:2022:WEF

[1059] Andrés García Mangas, Francisco José Suárez Alonso, Daniel Fernando García Martínez, and Fidel Díez Díaz. WoTemu: an emulation framework for edge computing architectures based on the Web of Things. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000767.

Myneni:2022:SDD

[1060] Sowmya Myneni, Ankur Chowdhary, Dijiang Huang, and Adel Alshamrani. SmartDefense: a distributed deep defense against DDoS attacks with edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000792.

Somesula:2022:CCU

[1061] Manoj Kumar Somesula, Rashmi Ranjan Rout, and D. V. L. N. Somayajulu. Cooperative cache update using multi-agent recurrent deep reinforcement learning for mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000809.

Nacef:2022:MLB

[1062] Abdelhakim Nacef, Abdellah Kaci, Youcef Aklouf, and Diego Leonel Cadette Dutra. Machine learning based fast self optimized and life cycle management network. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000949.

Haris:2022:LMV

[1063] Raseena M. Haris, Khaled M. Khan, and Armstrong Nhlabatsi. Live migration of virtual machine memory content in networked systems. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000962.

Abbas:2022:JCC

[1064] Nadine Abbas, Sanaa Sharafeddine, Azzam Mourad, Chadi Abou-Rjeily, and Wissam Fawaz. Joint computing, communication and cost-aware task offloading in D2D-enabled Het-MEC. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000974.

Kaur:2022:FBJ

[1065] Jasneet Kaur and M. Arif Khan. Federation based joint client and server side machine learning for 5G and beyond Wireless Channel Estimation. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000986.

Sassi:2022:CSA

[1066] Mohamed Saifeddine Hadj Sassi and Lamia Chaari Fourati. Comprehensive survey on air quality monitoring systems based on emerging computing and communication technologies. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000998.

Pei:2022:PFL

[1067] Jiaming Pei, Kaiyang Zhong, Mian Ahmad Jan, and Jinhai Li. Personalized federated learning framework for network traffic anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001001. See

retraction notice [1762].

Kita:2022:PRL

[1068] Kentaro Kita, Yuki Koizumi, and Toru Hasegawa. Private retrieval of locationrelated content using k-anonymity and application to ICN. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001013.

Jagannath:2022:DFC

[1069] Jithin Jagannath, Anu Jagannath, Justin Henney, Tyler Gwin, Zackary Kane, Noor Biswas, and Andrew Drozd. Design of fieldable crosslayer optimized network using embedded software defined radios: Survey and novel architecture with field trials. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001098.

Zhan:2022:DDH

[1070] Mengqi Zhan, Yang Li, Guangxi Yu, Bo Li, and Weiping Wang. Detecting DNS over HTTPS based data exfiltration. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001104.

Katila:2022:AMS

[1071] Risto Katila, Tuan Nguyen Gia, and Tomi Westerlund. Analysis of mobility support approaches for edgebased IoT systems using high data rate Bluetooth Low Energy 5. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001141.

Chen:2022:EDH

[1072] Tianfei Chen, Shuaixin Hou, and Lijun Sun. An enhanced DV-hop positioning scheme based on spring model and reliable beacon node set. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001153.

Tang:2022:CPR

[1073] Kun Tang, Feiyu Jiao, Xuxun Liu, Wenquan Che, and Quan Xue. Coverage probability of relay-assisted NOMA millimeter wave networks with steerable-beam. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001177.

Serper:2022:CCB

[1074] Elif Zeynep Serper and Aysegül Altin-Kayhan. Coverage and connectivity based lifetime maximization with topology update for WSN in smart grid applications. Computer Networks (Amsterdam, Netherlands: 1999), 209 (??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001232.

Anonymous:2022:EBh

[1075] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200144X.

Anonymous:2022:Mc

[1076] Anonymous. 22 May 2022. Computer Networks (Amsterdam, Netherlands: 1999), 209(??):??, May 22, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yang:2022:DPP

[1077] Hailiang Yang, Rukhsana Ruby, and Kaishun Wu. Delay performance of priority-queue equipped UAV-based mobile relay networks: Exploring the impact of trajectories. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000676.

Ghosh:2022:BED

[1078] Timam Ghosh, Arijit Roy, and Sudip Misra. B2H: Enabling delay-tolerant blockchain network in healthcare for Society 5.0. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622000706.

Luomala:2022:ARB

[1079] Jari Luomala and Ismo Hakala. Adaptive range-based localization algorithm based on trilateration and reference node selection for outdoor wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622000743.

Baldini:2022:ODD

[1080] Gianmarco Baldini and Irene Amerini.

Online Distributed Denial of Service (DDoS) intrusion detection based on adaptive sliding window and morphological fractal dimension. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200113X.

Charr:2022:PCO

[1081] Jean-Claude Charr, Karine Deschinkel, Rania Haj Mansour, and Mourad Hakem. Partial coverage optimization under network connectivity constraints in heterogeneous sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001165.

Borromeo:2022:FAS

[1082] Justine Cris Borromeo, Koteswararao Kondepu, Nicola Andriolli, and Luca Valcarenghi. FPGA-accelerated Smart-NIC for supporting 5G virtualized radio access network. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001189.

Gupta:2022:DAI

[1083] Ankit Kumar Gupta and T. G. Venkatesh. Design and analysis of IEEE 802.11 based full duplex WLAN MAC protocol. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001190.

Beraldi:2022:ISI

[1084] Roberto Beraldi, Claudia Canali, Riccardo Lancellotti, and Gabriele Proietti Mattia. On the impact of stale information on distributed online load balancing protocols for edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001207.

Li:2022:TOS

[1085] Zhehao Li, Lei Shi, Yi Shi, Zhenchun Wei, and Yang Lu. Task offloading strategy to maximize task completion rate in heterogeneous edge computing environment. Computer Networks (Amsterdam, Netherlands:

1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001219.

Wang:2022:TEE

[1086] Yang Wang, Yutong Li, Ting Wang, and Gang Liu. Towards an energyefficient data center network based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001220.

Ogundoyin:2022:SPP

[1087] Sunday Oyinlola Ogundoyin and Ismaila Adeniyi Kamil. Secure and privacy-preserving D2D communication in fog computing services. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001244.

Sellami:2022:EAT

[1088] Bassem Sellami, Akram Hakiri, Sadok Ben Yahia, and Pascal Berthou. Energy-aware task scheduling and offloading using deep reinforcement learning in SDN-enabled IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001359.

Ioannou:2022:DAM

[1089] Iacovos Ioannou, Christophoros Christophorou, Vasos Vassiliou, and Andreas Pitsillides. A distributed AI/ML framework for D2D transmission mode selection in 5G and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001396.

Sabbioni:2022:DDD

[1090] Andrea Sabbioni, Lorenzo Rosa, Armir Bujari, Luca Foschini, and Antonio Corradi. DIFFUSE: a DIstributed and decentralized platForm enabling Function composition in Serverless Environments. Computer Networks (Amsterdam, Netherlands: 1999), 210 (??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200161X.

Anonymous:2022:EBi

[1091] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622001797.

Anonymous:2022:Jb

[1092] Anonymous. 19 June 2022. Computer Networks (Amsterdam, Netherlands: 1999), 210(??):??, June 19, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lacava:2022:SBL

[1093] Andrea Lacava, Valerio Zottola, Alessio Bonaldo, Francesca Cuomo, and Stefano Basagni. Securing Bluetooth low energy networking: an overview of security procedures and threats. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001335.

Gomides:2022:PCC

[1094] Thiago S. Gomides, Robson E. De Grande, Rodolfo I. Meneguette, Fernanda S. H. de Souza, and Daniel L. Guidoni. Predictive congestion control based on collaborative information sharing for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001347.

Uzunidis:2022:BPO

[1095] D. Uzunidis, C. Matrakidis, E. Kosmatos, A. Stavdas, and A. Lord. On the benefits of power optimization in the S, C and L-band optical transmission systems. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001360.

Sharma:2022:RSP

[1096] Gourav Prateek Sharma, Wouter Tavernier, Didier Colle, and Mario Pickavet. Routing and scheduling for 1 + 1 protected DetNet flows. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001372.

Mairaj:2022:GTS

[1097] Aakif Mairaj and Ahmad Y. Javaid. Game theoretic solution for an unmanned aerial vehicle network host under DDoS attack. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001384.

Bagnulo:2022:PRL

[1098] Marcelo Bagnulo, Alberto García-Martínez, Stefano Angieri, Andra Lutu, and Jinze Yang. Practicable route leak detection and protection with ASIRIA. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001402.

Ji:2022:JOR

[1099] Pengshuo Ji, Jie Jia, Jian Chen, Yunhe Xie, and Xingwei Wang. Joint optimization for RIS-assisted multicast D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001499.

Ali:2022:SMS

[1100] Shabir Ali, Mayank Pandey, and Neeraj Tyagi. SDFog-Mesh: a software-defined fog computing architecture over wireless mesh networks for semi-permanent smart environments. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001578.

Ioannou:2022:NDA

[1101] Iacovos Ioannou, Christophoros Christophorou, Vasos Vassiliou, and Andreas Pitsillides. A novel distributed AI framework with ML for D2D communication in 5G/6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200158X.

Ojha:2022:BWC

[1102] Tamoghna Ojha, Theofanis P. Raptis, Marco Conti, and Andrea Passarella. Balanced wireless crowd charging with mobility prediction and social awareness. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001591.

Niu:2022:BSS

[1103] Shufen Niu, Siwei Zhou, Lizhi Fang, Ying Hu, and Caifen Wang. Broadcast signcryption scheme based on certificateless in wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 211 (??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001621.

Garcia:2022:ESC

[1104] Rodrigo D. Garcia, Gowri Ramachandran, and Jó Ueyama. Exploiting smart contracts in PBFT-based blockchains: a case study in medical prescription system. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001669.

Hasan:2022:BBS

[1105] Khalid Hasan, Mohammad Jabed Morshed Chowdhury, Kamanashis Biswas, Khandakar Ahmed, Md. Saiful Islam, and Muhammad Usman. A blockchain-based secure data-sharing framework for Software Defined Wireless Body Area Networks. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001670.

Zhang:2022:EIR

[1106] Songli Zhang, Weijia Jia, Zhiqing Tang, Jiong Lou, and Wei Zhao. Efficient instance reuse approach for service function chain placement in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001700.

Li:2022:SRL

[1107] Hui Li, Dongcong Shi, Weizheng Wang, Dan Liao, Thippa Reddy Gadekallu, and Keping Yu. Secure routing for LEO satellite network survivability. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001712.

Vegni:2022:GER

[1108] Anna Maria Vegni, Chaker Abdelaziz Kerrache, Waleed Ejaz, Enrico Natalizio, Jiming Chen, and Houbing Song. Guest editorial: Recent advances in connected and autonomous unmanned aerial/ground vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001724.

Anonymous:2022:EBj

[1109] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002018.

Anonymous:2022:Jc

[1110] Anonymous. 5 July 2022. Computer Networks (Amsterdam, Netherlands: 1999), 211(??):??, July 5, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Qin:2022:JOS

[1111] Yudong Qin, Deke Guo, Lailong Luo, and Ming Xu. A joint orchestration of security and functionality services at network edge. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001323.

Maryam:2022:LQQ

[1112] Hafsa Maryam, Tania Panayiotou, and Georgios Ellinas. Learning quantile QoT models to address uncertainty over unseen lightpaths. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001608.

Giakoumakis:2022:SMN

[1113] Giannis Giakoumakis, Eva Papadogiannaki, Giorgos Vasiliadis, and Sotiris Ioannidis. Scheduling of multiple network packet processing applications using Pythia. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001682.

Wu:2022:RNC

[1114] Ke Wu, Dezun Dong, Cunlu Li, and Weixia Xu. Revisiting network congestion avoidance through adaptive packet-chaining reservation. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001694.

Chang:2022:SMD

[1115] Jinyong Chang, Qiaochuan Ren, Yanyan Ji, Maozhi Xu, and Rui Xue. Secure medical data management with privacy-preservation and authentication properties in smart healthcare system. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001736.

Luttringer:2022:DNO

[1116] Jean-Romain Luttringer, Thomas Alfroy, Pascal Mérindol, Quentin Bramas, François Clad, and Cristel Pelsser. Deploying near-optimal delayconstrained paths with Segment Routing in massive-scale networks. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001748.

Li:2022:TCC

[1117] Ying Li, Yi Huang, Suranga Seneviratne, Kanchana Thilakarathna, Adriel Cheng, Guillaume Jourjon, Darren Webb, David B. Smith, and Richard Yi Da Xu. From traffic classes to content: a hierarchical approach for encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200175X.

Tsipis:2022:JOS

[1118] Athanasios Tsipis and Konstantinos Oikonomou. Joint optimization of social interactivity and server provisioning for interactive games in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001840.

Gomez:2022:STE

[1119] Jose Gomez, Elie F. Kfoury, Jorge Crichigno, and Gautam Srivastava. A survey on TCP enhancements using P4-programmable devices. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001852.

Macas:2022:SDL

[1120] Mayra Macas, Chunming Wu, and Walter Fuertes. A survey on deep learning for cybersecurity: Progress, challenges, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001864.

Martini:2022:IBZ

[1121] B. Martini, M. Gharbaoui, and P. Castoldi. Intent-based zero-touch service chaining layer for software-defined edge cloud networks. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001876.

Bagnulo:2022:EEL

[1122] Marcelo Bagnulo and Alberto García-Martínez. An experimental evaluation of LEDBAT++. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001888.

Sarieddine:2022:OVB

[1123] Khaled Sarieddine, Hassan Artail, and Haidar Safa. An opportunistic vehiclebased task assignment for IoT offloading. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200189X.

Asaad:2022:CRI

[1124] Safar M. Asaad and Halgurd S. Maghdid. A comprehensive review of indoor/outdoor localization solutions in IoT era: Research challenges and future perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001918.

Wu:2022:LAC

[1125] Wenjia Wu, Yujing Liu, Jiazhi Yao, Xiaolin Fang, Feng Shan, Ming Yang, Zhen Ling, and Junzhou Luo. Learning-aided client association control for high-density WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200192X.

Chen:2022:DFS

[1126] Chen Chen, Lars Nagel, Lin Cui, and Fung Po Tso. Distributed federated service chaining: a scalable and cost-aware approach for multidomain networks. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001931.

Callejo:2022:MDW

[1127] Patricia Callejo, Marcelo Bagnulo, Jaime González Ruiz, Andra Lutu, Alberto García-Martínez, and Rubén Cuevas. Measuring DoH with web ads. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001943.

Boobalan:2022:FFL

[1128] Parimala Boobalan, Swarna Priya Ramu, Quoc-Viet Pham, Kapal Dev, Sharnil Pandya, Praveen Kumar Reddy Maddikunta, Thippa Reddy Gadekallu, and Thien Huynh-The. Fusion of Federated Learning and Industrial Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001955.

Mansour:2022:EAF

[1129] Romany F. Mansour, Suliman A. Alsuhibany, Sayed Abdel-Khalek, Randa Alharbi, Thavavel Vaiyapuri, Ahmed J. Obaid, and Deepak Gupta. Energy aware fault tolerant clustering with routing protocol for improved survivability in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001967.

Nakimuli:2022:ERA

[1130] Winnie Nakimuli, Jaime Garcia-Reinoso, J. Enrique Sierra-Garcia, and Pablo Serrano. Exploiting radio access information to improve performance of remote-controlled mobile robots in MEC-based 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002067.

Dao:2022:ABS

[1131] Nhu-Ngoc Dao, Duc-Nghia Vu, Woongsoo Na, Trong-Minh Hoang, Dinh-Thuan Do, and Sungrae Cho. Adaptive bitrate streaming in multi-user downlink NOMA edge caching systems with imperfect SIC. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002080.

Junfeng:2022:PDS

[1132] Tian Junfeng, Bai Wenqing, and Jia Haoyi. PGCE: a distributed storage causal consistency model based on partial geo-replication and cloudedge collaboration architecture. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002092.

Garcia-Teodoro:2022:NZT

[1133] P. García-Teodoro, J. Camacho, G. Maciá-Fernández, J. A. Gómez-Hernández, and V. J. López-Marín. A novel zero-trust network access control scheme based on the security profile of devices and users. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002109.

Uyan:2022:RSM

[1134] Osman Gokhan Uyan, Ayhan Akbas, and Vehbi Cagri Gungor. A reliable and secure multi-path routing strategy for underwater acoustic sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002122.

Li:2022:FHD

[1135] Chunlin Li, Yong Zhang, and Youlong Luo. Flexible heterogeneous data fusion strategy for object positioning applications in edge computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 212 (??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002201.

Anonymous:2022:EBk

[1136] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200233X.

Anonymous:2022:Jd

[1137] Anonymous. 20 July 2022. Computer Networks (Amsterdam, Netherlands: 1999), 212(??):??, July 20, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hegyi:2022:SDD

[1138] Péter Hegyi. Service deployment design in latency-critical multi-cloud environment. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001487.

Su:2022:DSP

[1139] Lina Su, Ne Wang, Ruiting Zhou, and Zongpeng Li. Dynamic service placement and request scheduling for edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622001633.

Cremonezi:2022:IAR

[1140] Bruno Cremonezi, Airton Ribeiro Gomes Filho, Edelberto Franco Silva, José Augusto Miranda Nacif, Alex Borges Vieira, and Michele Nogueira. Improving the attribute retrieval on ABAC using opportunistic caches for Fog-Based IoT Networks. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001657.

Vitor:2022:SAS

[1141] Gonçalo Vítor, Pedro Rito, Susana Sargento, and Filipe Pinto. A scalable approach for smart city data platform: Support of real-time processing and data sharing. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001839.

Shi:2022:SDU

[1142] Yao Shi, Emad Alsusa, and Mohammed W. Baidas. A survey on downlink-uplink decoupled access: Advances, challenges, and open problems. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622001906.

Tu:2022:RES

[1143] Wanqing Tu. Resource-efficient seamless transitions for high-performance multi-hop UAV multicasting. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622001979.

Saleem:2022:PXC

[1144] Tania Saleem, Muhammad Umar Janjua, Muhammad Hassan, Talha Ahmad, Filza Tariq, Khadija Hafeez, Muhammad Ahsan Salal, and Muhammad Danish Bilal. ProofChain: an x.509-compatible blockchain-based PKI framework with decentralized trust. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002110.

Basturk:2022:EEC

[1145] Ilhan Bastürk. Energy-efficient communication for UAV-enabled mobile relay networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002134.

Langer:2022:NCK

[1146] Martin Langer and Rainer Bermbach. NTS4PTP — a comprehensive key management solution for PTP networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002158.

Bonafini:2022:EEP

[1147] Stefano Bonafini, Claudio Sacchi, Riccardo Bassoli, Koteswararao Kondepu, Fabrizio Granelli, and Frank H. P. Fitzek. End-to-end performance assessment of a 3D network for 6G connectivity on Mars surface. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002171.

Luo:2022:MEO

[1148] Meng Luo, Yepeng Yao, Liling Xin, Zhengwei Jiang, Qiuyun Wang, and Wenchang Shi. Measurement for encrypted open resolvers: Applications and security. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002183.

Zahedinia:2022:FBC

[1149] Marzieh Sadat Zahedinia, Mohammad Reza Khayyambashi, and Ali Bohlooli. Fog-based caching mechanism for IoT data in information centric network using prioritization. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002195.

Zaki:2022:GGM

[1150] Faiz Zaki, Firdaus Afifi, Shukor Abd Razak, Abdullah Gani, and Nor Badrul Anuar. GRAIN: Granular multi-label encrypted traffic classification using classifier chain. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002213.

Mezher:2022:GGT

[1151] Ahmad Mohamad Mezher and Mónica Aguilar Igartua. G-3MRP: a game-theoretical multimedia multimetric map-aware routing protocol for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002225.

Liang:2022:PDC

[1152] Yihuai Liang, Yan Li, and Byeong-Seok Shin. Private decentralized crowdsensing with asynchronous blockchain access. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002237.

Souza:2022:CDA

[1153] Otavio A. de O. Souza, Olga Goussevskaia, and Stefan Schmid. CB-Net: Demand-aware tree topologies for Reconfigurable Datacenter Networks. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002249.

Illiano:2022:QIP

[1154] Jessica Illiano, Marcello Caleffi, Antonio Manzalini, and Angela Sara Cacciapuoti. Quantum Internet protocol stack: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002250.

Mian:2022:VAI

[1155] Adnan Noor Mian, Syed Waqas Haider Shah, Sanaullah Manzoor, Anwar Said, Kurtis Heimerl, and Jon Crowcroft. A value-added IoT service for cellular networks using federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002262.

Alioua:2022:IMC

[1156] Ahmed Alioua, Roumayssa Hamiroune, Oumayma Amiri, Manel Khelifi, Sidi-Mohammed Senouci, Mikael Gidlund, and Sarder Fakhrul Abedin. Incentive mechanism for competitive edge caching in 5G-enabled Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002274.

Abbas:2022:PBR

[1157] Ghulam Abbas, Sami Ullah, Muhammad Waqas, Ziaul Haq Abbas, and Muhammad Bilal. A positionbased reliable emergency message routing scheme for road safety in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002286.

Ran:2022:DSA

[1158] Longyan Ran, Yunhe Cui, Chun Guo, Qing Qian, Guowei Shen, and Huanlai Xing. Defending saturation attacks on SDN controller: a confusable instance analysis-based algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002298.

Nisha:2022:NPP

[1159] Nisha Nisha, Iynkaran Natgunanathan, Shang Gao, and Yong Xiang. A novel privacy protection scheme for location-based services using collaborative caching. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002377.

Catena:2022:DLB

[1160] Tiziana Catena, Vincenzo Eramo, Massimo Panella, and Antonello Rosato. Distributed LSTM-based cloud resource allocation in Network Function Virtualization Architectures. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002390.

Guo:2022:JWR

[1161] Mian Guo, Mithun Mukherjee, Jaime Lloret, Jiangtao Ou, and Chengyuan Fan. Joint wireless resource allocation and service function chaining scheduling for Tactile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002407.

Hernandez-Solana:2022:PDP

[1162] Á. Hernández-Solana, D. Perez-Diaz de Cerio, M. García-Lozano, A. Valdovinos, and J. L. Valenzuela. PSM-DMO: Power save mode and discontinuous BLE mesh operation. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002419.

Waleed:2022:WOS

[1163] Abdul Waleed, Abdul Fareed Jamali, and Ammar Masood. Which open-source IDS? Snort, Suricata or Zeek. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002420.

Hilal:2022:TAO

[1164] Anwer Mustafa Hilal, Amani Abdulrahman Albraikan, Sami Dhahbi, Saud S. Alotaibi, Rana Alabdan, Mesfer Al Duhayyim, Abdelwahed Motwakel, and Ishfaq Yaseen. Trust aware oppositional sine cosine based multihop routing protocol for improving survivability of wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002444.

Arsalan:2022:IBT

[1165] Ahmad Arsalan and Rana Asif Rehman. Interest Broadcasting and Timing Attack in IoV (IBTA-IoV): a novel architecture using Named Software Defined Network. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002456.

Bringhenti:2022:AVO

[1166] Daniele Bringhenti, Jalolliddin Yusupov, Alejandro Molina Zarca, Fulvio Valenza, Riccardo Sisto, Jorge Bernal Bernabe, and Antonio Skarmeta. Automatic, verifiable and optimized policybased security enforcement for SDNaware IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002468.

Singh:2022:EEF

[1167] Namrata Singh and Ayan Kumar Das.

Energy-efficient fuzzy data offloading for IoMT. Computer Networks (Amsterdam, Netherlands: 1999), 213 (??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002493.

Anonymous:2022:EBl

[1168] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622003024.

Anonymous:2022:Ac

[1169] Anonymous. 4 August 2022. Computer Networks (Amsterdam, Netherlands: 1999), 213(??):??, August 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jmila:2022:AML

[1170] Houda Jmila and Mohamed Ibn Khedher. Adversarial machine learning for network intrusion detection: a comparative study. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002146.

Ribeiro:2022:DAE

[1171] Rafael Hengen Ribeiro, Arthur Selle Jacobs, Luciano Zembruzki, Ricardo Parizotto, Eder John Scheid, Alberto Egon Schaeffer-Filho, Lisandro Zambenedetti Granville, and Burkhard Stiller. A deterministic approach for extracting network security intents. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002389.

Lasla:2022:GPE

[1172] Noureddine Lasla, Lina Al-Sahan, Mohamed Abdallah, and Mohamed Younis. Green-PoW: an energy-efficient blockchain Proof-of-Work consensus algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002432.

Chen:2022:AJC

[1173] Kuan-Wei Chen, Chih-Min Chao, Chih-Yu Lin, and Chun-Chao Yeh. Anti-jamming channel hopping protocol design based on channel occupancy probability for Cognitive Radio Networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200247X.

Huang:2022:FEC

[1174] Shan Huang, Dezun Dong, Zejia Zhou, Hanyi Shi, Wenxiang Yang, and Xiangke Liao. FastCredit: Expediting credit-based congestion control in datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002481.

Peng:2022:MVF

[1175] Songtao Peng, Jiaqi Nie, Xincheng Shu, Zhongyuan Ruan, Lei Wang, Yunxuan Sheng, and Qi Xuan. A multi-view framework for BGP anomaly detection via graph attention network. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200250X.

Yang:2022:DDD

[1176] Donghui Yang, Zhenyu Li, Haiyang Jiang, Gareth Tyson, Hongtao Li, Gaogang Xie, and Yu Zeng. A deep dive into DNS behavior and query failures. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002511.

Barakabitze:2022:SNQ

[1177] Alcardo Alex Barakabitze and Ray Walshe. SDN and NFV for QoEdriven multimedia services delivery: the road towards 6G and beyond networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002523. See

corrigendum [1571].

Luong:2022:JTS

[1178] Nguyen Cong Luong, Tran The Anh, Zehui Xiong, Dusit Niyato, and Dong In Kim. Joint time scheduling and transaction fee selection in blockchain-based RFpowered backscatter cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002535.

Kumari:2022:TOF

[1179] Nidhi Kumari, Anirudh Yadav, and Prasanta K. Jana. Task offloading in fog computing: a survey of algorithms and optimization techniques. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002547.

Fouladi:2022:DAD

[1180] Ramin Fadaei Fouladi, Orhan Ermis, and Emin Anarim. A DDoS attack detection and countermeasure scheme based on DWT and auto-encoder neural network for SDN. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002560.

Sasabe:2022:AMD

[1181] Masahiro Sasabe and Masaki Kiyomitsu. Analysis of minimum distribution time of two-class tit-for-tat-based P2P file distribution. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002572.

Mertens:2022:MFC

[1182] J. S. Mertens, L. Galluccio, and G. Morabito. MGM-4-FL: Combining federated learning and model gossiping in WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002584.

Zafar:2022:DDS

[1183] Samra Zafar, Zefeng Lv, Nizam Hussain Zaydi, Muhammad Ibrar, and Xiaopeng Hu. DSMLB: Dynamic switch-migration based load balancing for software-defined IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002596.

Habibi:2022:RBP

[1184] Mehdi Habibi and Hamzeh Beyranvand. Reducing blocking probability and QoT violation in dynamic elastic optical networks via load-aware margin selection. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002602.

Cao:2022:FSE

[1185] Yibo Cao, Shiyuan Xu, Xue Chen, Yunhua He, and Shuo Jiang. A forward-secure and efficient authentication protocol through lattice-based group signature in VANETs scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 214(??): ??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002626.

Li:2022:CEC

[1186] Nan Li, Alexandros Iosifidis, and Qi Zhang. Collaborative edge computing for distributed CNN inference acceleration using receptive field-based segmentation. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002638.

Picano:2022:HLV

[1187] Benedetta Picano and Romano Fantacci. Human-in-the-loop virtual reality offloading scheme in wireless 6G terahertz networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200264X.

deSouza:2022:IDP

[1188] Cristiano Antonio de Souza, Carlos Becker Westphall, Renato Bobsin Machado, Leandro Loffi, Carla Merkle Westphall, and Guilherme Arthur Geronimo. Intrusion detection and prevention in fog based IoT environments: a systematic literature review. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002651.

Duan:2022:FFG

[1189] Jian-Hui Duan, Wenzhong Li, Xiao Zhang, and Sanglu Lu. Forecasting fine-grained city-scale cellular traffic with sparse crowdsourced measurements. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002663.

Chen:2022:DBP

[1190] Wenchao Chen, Guanqun Shen, Kaikai Chi, Shubin Zhang, and Xiaolong Chen. DRL based partial offloading for maximizing sum computation rate of FDMA-based wireless powered mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002675.

Markovitz:2022:LSB

[1191] Oren Markovitz and Michael Segal. LEO satellite beam management algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002687.

Bomgni:2022:NNS

[1192] Alain Bertrand Bomgni, Miguel Landry Foko Sindjoung, Dhalil Kamdem Tchibonsou, Mthulisi Velempini, and Jean Frédéric Myoupo. NE-SEPRIN: a new scheme for energyefficient permutation routing in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002699.

Wang:2022:UMS

[1193] Jiacheng Wang, Jianhui Zhang, Liming Liu, Xuzhao Zheng, Hanxiang Wang, and Zhigang Gao. Utility maximization for splittable task offloading in IoT edge network. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002705.

Almasan:2022:EER

[1194] Paul Almasan, Shihan Xiao, Xiangle Cheng, Xiang Shi, Pere Barlet-Ros, and Albert Cabellos-Aparicio. EN-ERO: Efficient real-time WAN routing optimization with deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002717.

Tsiropoulou:2022:GEI

[1195] Eirini Eleni Tsiropoulou, Christos Douligeris, Luca Foschini, Gang Li, and Theofanis P. Raptis. Guest editorial: 26th IEEE symposium on computers and communications (ISCC 2021) selected papers. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002729.

Qian:2022:DDS

[1196] Mimi Qian, Lin Cui, Xiaoquan Zhang, Fung Po Tso, and Yuhui Deng. Drops: Detecting silent packet drops on programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002742.

Jurdak:2022:EBB

[1197] Raja Jurdak, Juan M. Corchado, Jong Hyuk Park, Chintan Bhatt, and Kapal Dev. Editorial: Blockchain based sustainable, secure healthcare systems. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200278X.

Wang:2022:RAO

[1198] Zhifei Wang, Xiangming Wen, Zhaoming Lu, Wenpeng Jing, and Yujing Zhang. Random access optimization for initial access and seamless handover for 5g-satellite network. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002791.

Dong:2022:SSI

[1199] Xiaodong Dong and Binlei Cai. Slardar: Scheduling information incomplete inter-datacenter deadline-aware coflows with a decentralized framework. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002808.

Eltaief:2022:FCF

[1200] Hamdi Eltaief. Flex-CC: a flexible connected chains scheme for multicast source authentication in dynamic SDN environment. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002766.

Fu:2022:GIE

[1201] Juan juan Fu and Xing lan Zhang. Gradient importance enhancement based feature fusion intrusion detection technique. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002754.

Wang:2022:ODO

[1202] Cen Wang, Noboru Yoshikane, Filippos Balasis, and Takehiro Tsuritani. OSDL: Dedicated optical slice provisioning in support of distributed deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 214 (??):??, September 4, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002857.

Anonymous:2022:EBm

[1203] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622003474.

Anonymous:2022:S

[1204] Anonymous. 4 September 2022. Computer Networks (Amsterdam, Netherlands: 1999), 214(??):??, September 4, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Alam:2022:ESB

[1205] Gowthul Alam MM, Jerald Nirmal Kumar S., Uma Mageswari R., and Michael Raj TF. An efficient SVM based DEHO classifier to detect DDoS attack in cloud computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002559.

Qu:2022:RAM

[1206] Bin Qu, Yan Bai, Yul Chu, Li e Wang, Feng Yu, and Xianxian Li. Resource allocation for MEC system with multi-users resource competition based on deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002778.

Bringhenti:2022:ODF

[1207] Daniele Bringhenti and Fulvio Valenza.

Optimizing distributed firewall reconfiguration transients. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200281X.

Zhang:2022:PUT

[1208] Jiang Zhang, Lillian Clark, Matthew Clark, Konstantinos Psounis, and Peter Kairouz. Privacy-utility trades in crowdsourced signal map obfuscation. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002833.

Costa:2022:MFC

[1209] Breno Costa, João Bachiega, Leonardo Rebouças Carvalho, Michel Rosa, and Aleteia Araujo. Monitoring fog computing: a review, taxonomy and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002845.

Mouawad:2022:RSS

[1210] Mostafa Mouawad, Firmin Mah, and Zbigniew Dziong. RRH-Sector selection and load balancing based on MDP and dynamic RRH-Sector-BBU mapping in C-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002869.

Wu:2022:CMU

[1211] Fan Wu, Chaonong Xu, Yutong Zhu, and Chao Li. Complexity of minimum uplink power scheduling with delay bound for backbone-assisted PD-NOMA wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002870.

Li:2022:MOO

[1212] Jiahui Li, Geng Sun, Aimin Wang, Ming Lei, Shuang Liang, Hui Kang, and Yanheng Liu. A many-objective optimization charging scheme for wireless rechargeable sensor networks via mobile charging vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002882.

Zeng:2022:OMQ

[1213] Hongxin Zeng, Lin Cui, Fung Po Tso, and Zhen Zhang. Optimizing multipath QUIC transmission over heterogeneous paths. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002894.

Leyva-Pupo:2022:DUP

[1214] Irian Leyva-Pupo, Cristina Cervelló-Pastor, Christos Anagnostopoulos, and Dimitrios P. Pezaros. Dynamic UPF placement and chaining reconfiguration in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002900.

Gedikli:2022:DRL

[1215] Ahmet Melih Gedikli, Mehmet Koseoglu, and Sevil Sen. Deep reinforcement learning based flexible preamble allocation for RAN slicing in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002912.

Wang:2022:PPF

[1216] Weilong Wang, Yingjie Wang, Yan Huang, Chunxiao Mu, Zice Sun, Xiangrong Tong, and Zhipeng Cai. Privacy protection federated learning system based on blockchain and edge computing in mobile crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002936.

Robin:2022:PPS

[1217] Debobroto Das Robin and Javed I. Khan. P4TE: PISA switch based traffic engineering in fat-tree data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200295X.

Tu:2022:RRN

[1218] Huaqing Tu, Gongming Zhao, Hongli Xu, Yangming Zhao, Yuhang Qiu, and Liusheng Huang. RoNS: Robust network function services in clouds. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622002961.

DiCicco:2022:OTR

[1219] Nicola Di Cicco, Federico Tonini, Valentina Cacchiani, and Carla Raffaelli. Optimization over time of reliable 5G-RAN with network function migrations. Computer Networks (Amsterdam, Netherlands: 1999), 215 (??):??, October 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002985.

Anonymous:2022:EBn

[1220] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622003802.

Anonymous:2022:Oa

[1221] Anonymous. 9 October 2022. Computer Networks (Amsterdam, Netherlands: 1999), 215(??):??, October 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pirayesh:2022:PHB

[1222] Jamshid Pirayesh, Alberto Giaretta, Mauro Conti, and Parviz Keshavarzi. A PLS-HECC-based device authentication and key agreement scheme for smart home networks. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200216X.

Beduneau:2022:MPC

[1223] Guillaume Béduneau, Ghada Jaber, and Bertrand Ducourthial. A method for predicting ITS cooperative applications performances. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002614.

Yan:2022:FLP

[1224] Lei Yan, Maode Ma, Dandan Li, Xiaohong Huang, Yan Ma, and Kun Xie. A flexible and lightweight privacypreserving handshake protocol based on DTLShps for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002730.

Khawam:2022:CSG

[1225] Kinda Khawam, Hassan Fawaz, Samer Lahoud, Odalric-Ambrym Maillard, and Steven Martin. A channel selection game for multi-operator Lo-RaWAN deployments. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002821.

Elkael:2022:MBR

[1226] Maxime Elkael, Massinissa Ait Aba, Andrea Araldo, Hind Castel-Taleb, and Badii Jouaber. Monkey business: Reinforcement learning meets neighborhood search for virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002924.

Lima:2022:GAS

[1227] Diogo Lima and Hugo Miranda. A geographical-aware state deployment service for fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002948.

Bhamare:2022:IBN

[1228] Deval Bhamare, Andreas Kassler, Jonathan Vestin, Mohammad Ali Khoshkholghi, Javid Taheri, Toktam Mahmoodi, Peter Öhlén, and Calin Curescu. IntOpt: In-band network telemetry optimization framework to monitor network slices using P4. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622002973.

Gharagezlou:2022:EEP

[1229] Abdolrasoul Sakhaei Gharagezlou, Mahdi Nangir, and Nima Imani. Energy efficient power allocation with joint antenna and user selection in massive MIMO systems. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003073.

Paolucci:2022:LCS

[1230] Francesco Paolucci, Davide Scano, Piero Castoldi, and Emiliano De Paoli. Latency control in service chaining using P4-based data plane programmability. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003085.

Vanitha:2022:OPR

[1231] C. N. Vanitha, S. Malathy, Rajesh Kumar Dhanaraj, and Anand Nayyar. Optimized Pollard route deviation and route selection using Bayesian machine learning techniques in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003097.

Oligeri:2022:GSD

[1232] Gabriele Oligeri, Savio Sciancalepore, Omar Adel Ibrahim, and Roberto Di Pietro. GPS spoofing detection via crowd-sourced information for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003103.

Xie:2022:FFB

[1233] Shengxu Xie, Guyu Hu, Changyou Xing, Jiachen Zu, and Yaqun Liu. FINT: Flexible in-band network telemetry method for data center network. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003115.

Vargas-Arcila:2022:SDS

[1234] Angela Vargas-Arcila, Juan Carlos Corrales, Araceli Sanchis, and Alvaro Rendón. SOFI dataset: Symptom-fault relationship for IP-network. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003127.

Wu:2022:SAB

[1235] Weiqi Wu, Xingfu Wang, Ammar Hawbani, Longzhi Yuan, and Wei Gong. A survey on ambient backscatter communications: Principles, systems, applications, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003139.

Sun:2022:ASW

[1236] Guoying Sun, Zhaoxin Zhang, Yanan Cheng, and Tingting Chai. Adaptive segmented webpage text based malicious website detection. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003140.

Niu:2022:DLO

[1237] Juncui Niu, Shubin Zhang, Kaikai Chi, Guanqun Shen, and Wei Gao. Deep learning for online computation offloading and resource allocation in NOMA. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003152.

Naha:2022:MLR

[1238] Ranesh Naha, Saurabh Garg, Sudheer Kumar Battula, Muhammad Bilal Amin, and Dimitrios Georgakopoulos. Multiple linear regression-based energy-aware resource allocation in the Fog computing environment. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003164.

Li:2022:TTS

[1239] Yanbin Li, Jiajie Zhu, Ming Tang, Shougang Ren, and Fusheng Wu. TSCL: a time-space crossing location for side-channel leakage detection. Computer Networks (Amsterdam, Netherlands: 1999), 216(??): ??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003176.

Sharma:2022:BME

[1240] Payal Sharma and B. R. Purushothama. BP-MGKM: an efficient multi-group key management scheme based on bivariate polynomial. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003188.

Tedeschi:2022:SBC

[1241] Pietro Tedeschi, Savio Sciancalepore, and Roberto Di Pietro. Satellitebased communications security: a survey of threats, solutions, and research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200319X.

Satpathy:2022:CEV

[1242] Anurag Satpathy, Manmath Narayan Sahoo, Arun Kumar Sangaiah, Chittaranjan Swain, and Sambit Bakshi. CoMap: an efficient virtual network re-mapping strategy based on coalitional matching theory. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003206.

Rajak:2022:EEM

[1243] Shaik Rajak, Poongundran Selvaprabhu, Sunil Chinnadurai, A. S. M. Sanwar Hosen, Aldosary Saad, and Amr Tolba. Energy efficient MIMO-NOMA aided IoT network in B5G communications. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003218.

Herreria-Alonso:2022:ELS

[1244] Sergio Herrería-Alonso, Andrés Suárez-González, Miguel Rodríguez-Pérez, and Cándido López-García. Enhancing LoRaWAN scalability with Longest First Slotted CSMA. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200322X.

Castellanos:2022:MOO

[1245] German Castellanos, Simon De Gheselle, Luc Martens, Niels Kuster, Wout Joseph, Margot Deruyck, and Sven Kuehn. Multi-objective optimisation of human exposure for various 5G network topologies in Switzerland. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003231.

Sharma:2022:OEO

[1246] Abhilasha Sharma and Lalit Kumar Awasthi. Ob-EID: Obstacle aware event information dissemination for SDN enabled vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003255.

Chen:2022:DSC

[1247] Junmei Chen, Hao Chen, and Zongpeng Li. A double serial concatenated code using CRC-aided error correction for highly reliable communication. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003279.

Zhang:2022:BEM

[1248] Jianfeng Zhang, Wensheng Zhang, and Jingdong Xu. Bandwidth-efficient multi-task AI inference with dynamic task importance for the Internet of Things in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003280.

Restuccia:2022:GWA

[1249] Francesco Restuccia, Salvatore D’Oro, Amani Al-Shawabka, Bruno Costa Rendon, Kaushik Chowdhury, Stratis Ioannidis, and Tommaso Melodia. Generalized wireless adversarial deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003292.

Febro:2022:SDD

[1250] Aldo Febro, Hannan Xiao, Joseph Spring, and Bruce Christianson. Synchronizing DDoS defense at network edge with P4, SDN, and Blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003310.

Batchu:2022:IAE

[1251] Raj Kumar Batchu and Hari Seetha. An integrated approach explaining the detection of distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003334.

Zhao:2022:CPA

[1252] Yubo Zhao and Jinyong Chang. Certificateless public auditing scheme with designated verifier and privacypreserving property in cloud storage. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):

??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003322.

Wang:2022:EEP

[1253] Yantong Wang and Vasilis Friderikos.

Energy-efficient proactive caching with multipath routing. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003346.

Miura:2022:PBD

[1254] Hideyoshi Miura, Kouji Hirata, and Takuji Tachibana. P4-based design of fast failure recovery for softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003358.

Khorasani:2022:NTD

[1255] Yaghoub Khorasani, Akbar Ghaffarpour Rahbar, and Mohammad Jafari-Beyrami. A novel two-dimensional metric for fragmentation evaluation in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200336X.

Blaise:2022:GAD

[1256] Agathe Blaise, Mathieu Bouet, Vania Conan, and Stefano Secci. Group anomaly detection in mobile app usages: a spatiotemporal convex hull methodology. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003371.

Li:2022:UNA

[1257] Kehong Li, Wengang Ma, Huawei Duan, Han Xie, Juanxiu Zhu, and Ruiqi Liu. Unbalanced network attack traffic detection based on feature extraction and GFDA-WGAN. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003401.

Gao:2022:RCS

[1258] Jing Gao, Lei Feng, Peng Yu, Fanqin Zhou, Zihao Wu, Xuesong Qiu, Jingchun Li, and Yifei Zhu. Resource consumption and security-aware multi-tenant service function chain deployment based on hypergraph matching. Computer Networks (Amsterdam, Netherlands: 1999), 216(??): ??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003516.

Silveira:2022:TCB

[1259] Lucas B. D. Silveira, Henrique C.

de Resende, Cristiano B. Both, Johann M. Marquez-Barja, Bruno Silvestre, and Kleber V. Cardoso. Tutorial on communication between access networks and the 5G core. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003528.

Lv:2022:ANN

[1260] Yunxin Lv, Zixin Liu, Meihua Bi, Hao Chi, Yanrong Zhai, Yang Lu, and Zhengfeng Qian. Artificial neural network assisted polling scheme for central coordinated low-latency WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 216 (??):??, October 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003541.

Anonymous:2022:EBo

[1261] Anonymous. Editorial board. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128622004066.

Anonymous:2022:Ob

[1262] Anonymous. 24 October 2022. Computer Networks (Amsterdam, Netherlands: 1999), 216(??):??, October 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2022:CUN

[1263] Kewei Wang, Changzhen Hu, and Chun Shan. Calculation of utility of network services based on state manifolds. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003267.

Mountaser:2022:GMO

[1264] Ghizlane Mountaser, Enric Pardo, and Toktam Mahmoodi. Graphical modelling and optimization of RAN function split deployed through UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003309.

Xiao:2022:SAA

[1265] Da Xiao, Shuo Chen, Wei Ni, Jie Zhang, Andrew Zhang, and Renping Liu. A sub-action aided deep reinforcement learning framework for latencysensitive network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003383.

Kaur:2022:RVI

[1266] Karamjeet Kaur, Veenu Mangat, and Krishan Kumar. A review on Virtualized Infrastructure Managers with management and orchestration features in NFV architecture. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003395.

Rago:2022:TDS

[1267] Arcangela Rago, Sergio Martiradonna, Giuseppe Piro, Andrea Abrardo, and Gennaro Boggia. A tenant-driven slicing enforcement scheme based on Pervasive Intelligence in the Radio Access Network. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003413.

Kang:2022:FTR

[1268] Rui Kang, Fujun He, and Eiji Oki. Fault-tolerant resource allocation model for service function chains with joint diversity and redundancy. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003425.

Zhang:2022:SCV

[1269] Lejun Zhang, Jinlong Wang, Weizheng Wang, Zilong Jin, Yansen Su, and Huiling Chen. Smart contract vulnerability detection combined with multi-objective detection. Computer Networks (Amsterdam, Netherlands:

1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003437.

Cavdar:2022:APS

[1270] Mustafa Can Çavdar, Ibrahim Korpeoglu, and Özgür Ulusoy. Application placement with shared monitoring points in multi-purpose IoT wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200353X.

Mesodiakaki:2022:REE

[1271] A. Mesodiakaki, E. Zola, and A. Kassler. Robust and energy-efficient user association and traffic routing in B5G HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003553.

Su:2022:SRS

[1272] Yishan Su, Yao Liu, Rong Fan, Luyu Li, Meng Han, and Hongliang Zhang. A secure relay selection scheme based on cooperative jamming for Underwater Acoustic Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003565.

Jiang:2022:AMA

[1273] Minghao Jiang, Zhen Li, Peipei Fu, Wei Cai, Mingxin Cui, Gang Xiong, and Gaopeng Gou. Accurate mobileapp fingerprinting using flow-level relationship with graph neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003577.

Li:2022:PBR

[1274] Yupeng Li, Mengjia Xia, Jingpu Duan, and Yang Chen. Pricing-based resource allocation in three-tier edge computing for social welfare maximization. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003589.

Kuang:2022:PAE

[1275] Ye Kuang, Dandan Li, Xiaohong Huang, Mo Zhou, and Weidu Wang. PoiEvent: an approach to extract the persistent and destructive routing events. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003590.

Mazziane:2022:ACM

[1276] Younes Ben Mazziane, Sara Alouf, and Giovanni Neglia. Analyzing count min sketch with conservative updates. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003607.

Buccafurri:2022:IPT

[1277] Francesco Buccafurri, Vincenzo De Angelis, and Gianluca Lax. An integritypreserving technique for range queries over data streams in two-tier sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003619.

Hou:2022:NID

[1278] Tianhao Hou, Hongyan Xing, Xinyi Liang, Xin Su, and Zenghui Wang. Network intrusion detection based on DNA spatial information. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003620.

Lu:2022:FAM

[1279] Tingting Lu and Junfeng Wang. F2DC: Android malware classification based on raw traffic and neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003632.

Abdollahi:2022:OOE

[1280] Mehran Pourmohammad Abdollahi, Javad Musevi Niya, and Mahdi Nangir. Optimizing the operator’s economy in multi-tier cellular networks with underlaid D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003644.

Chhea:2022:EEF

[1281] Kimchheang Chhea and Jung-Ryun Lee. Energy-efficient full-duplex D2D for SWIPT-empowered underlay cellular networks using a deep neural network. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003656.

Kolukisa:2022:DLA

[1282] Burak Kolukisa, Veli Can Yildirim, Bahadir Elmas, Cem Ayyildiz, and Vehbi Cagri Gungor. Deep learning approaches for vehicle type classification with 3-D magnetic sensor. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003668.

Kumari:2022:PQL

[1283] Swati Kumari, Maninder Singh, Raman Singh, and Hitesh Tewari. A post-quantum lattice based lightweight authentication and code-based hybrid encryption scheme for IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200367X.

Ferriol-Galmes:2022:BDT

[1284] Miquel Ferriol-Galmés, José Suárez-Varela, Jordi Paillissé, Xiang Shi, Shihan Xiao, Xiangle Cheng, Pere Barlet-Ros, and Albert Cabellos-Aparicio. Building a Digital Twin for network optimization using graph neural networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003681.

Wang:2022:TAM

[1285] Xueyi Wang, Xingwei Wang, Dongkuo Wu, Lianbo Ma, and Min Huang. Truthful auction mechanisms for VNF chain provisioning and allocation across geo-distributed datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003693.

Taneja:2022:EAR

[1286] Ashu Taneja, Shalli Rani, Sahil Garg, Mohammad Mehedi Hassan, and Salman A. AlQahtani. Energy aware resource control mechanism for improved performance in future green 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200370X.

Akram:2022:EES

[1287] Waseem Akram, Khalid Mahmood, Xiong Li, Mazhar Sadiq, Zhihan Lv, and Shehzad Ashraf Chaudhry. An energy-efficient and secure identity based RFID authentication scheme for vehicular cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003711.

Alrubei:2022:HHB

[1288] Subhi Alrubei, Edward Ball, and Jonathan Rigelsford. HDPoA: Honesty-based distributed proof of authority via scalable work consensus protocol for IoT-blockchain applications. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003723.

Violos:2022:IHA

[1289] John Violos, Stylianos Tsanakas, Theodoros Theodoropoulos, Aris Leivadeas, Konstantinos Tserpes, and Theodora Varvarigou. Intelligent horizontal autoscaling in edge computing using a double tower neural network. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003735.

Zhang:2022:ACB

[1290] Pengfei Zhang, Xiang Cheng, Sen Su, and Ning Wang. Area coveragebased worker recruitment under geo-indistinguishability. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003747.

Barmpounakis:2022:ADQ

[1291] Sokratis Barmpounakis, Nikolaos Maroulis, Nikolaos Koursioumpas, Apostolos Kousaridas, Angeliki Kalamari, Panagiotis Kontopoulos, and Nancy Alonistioti. AI-driven, QoS prediction for V2X communications in beyond 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003759.

Yang:2022:DPA

[1292] Yi Yang, Fenglei Li, Xinzhe Zhang, Zhixin Liu, and Kit Yan Chan. Dynamic power allocation in cellular network based on multi-agent double deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003760.

Duong:2022:FWH

[1293] Tho Minh Duong and Sungoh Kwon.

A frame work of handover analysis for randomly deployed heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003851.

Luglio:2022:SMB

[1294] M. Luglio, S. P. Romano, C. Roseti, and F. Zampognaro. Satellite multibeam multicast support for an efficient community-based CDN. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003863.

Tran:2022:CMR

[1295] Phuong T. Tran, Ba Cao Nguyen, Tran Manh Hoang, Le The Dung, and Nguyen Van Vinh. Combining multi-RIS and relay for performance improvement of multi-user NOMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003875.

Ren:2022:NMT

[1296] Junyu Ren and Tuanfa Qin. A novel multidimensional trust evaluation and fusion mechanism in fog-based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003887.

Selim:2022:SBF

[1297] Mohamed Y. Selim and Ahmed E. Kamal. Self-backhauling failure mitigation using 5G new radio. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003899.

Ramezanpour:2022:IZT

[1298] Keyvan Ramezanpour and Jithin Jagannath. Intelligent zero trust architecture for 5G/6G networks: Principles, challenges, and the role of machine learning in the context of O-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003929.

M:2022:SHO

[1299] Shyama M., Anju S. Pillai, and Alagan Anpalagan. Self-healing and optimal fault tolerant routing in wireless sensor networks using genetical swarm optimization. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003930.

Shariq:2022:EES

[1300] Mohd Shariq, Karan Singh, Chhagan Lal, Mauro Conti, and Tayyab Khan. ESRAS: an efficient and secure ultra-lightweight RFID authentication scheme for low-cost tags. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003942.

Shalini:2022:DDD

[1301] P. V. Shalini, V. Radha, and Sriram G. Sanjeevi. DOCUS-DDoS detection in SDN using modified CUSUM with flash traffic discrimination and mitigation. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003954.

Saad:2022:RTB

[1302] Muhammad Saad, Afsah Anwar, Ashar Ahmad, Hisham Alasmary, Murat Yuksel, and David Mohaisen. RouteChain: Towards blockchain-based secure and efficient BGP routing. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003966.

Masoudi-Sobhanzadeh:2022:RTI

[1303] Yosef Masoudi-Sobhanzadeh and Shabnam Emami-Moghaddam. A real-time IoT-based botnet detection method using a novel two-step feature selection technique and the support vector machine classifier. Computer Networks (Amsterdam, Netherlands: 1999), 217 (??):??, November 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622003991.

Anonymous:2022:EBp

[1304] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622004467.

Anonymous:2022:N

[1305] Anonymous. 9 November 2022. Computer Networks (Amsterdam, Netherlands: 1999), 217(??):??, November 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Diarra:2022:RRA

[1306] Mamoutou Diarra, Walid Dabbous, Amine Ismail, Brice Tetu, and Thierry Turletti. RAPID: a RAN-aware performance enhancing proxy for high throughput low delay flows in MECenabled cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003917.

Xie:2022:IEA

[1307] Pengshou Xie, Haoxuan Yang, Tao Feng, and Yan Yan. Implementing efficient attribute encryption in IoV under cloud environments. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622003978.

Gao:2022:CRS

[1308] Caina Gao, Jia Zhang, Linlin Guo, Ke Liu, Lili Meng, and Jiande Sun. Coordinated rate splitting and power allocation in energy-spectral efficiency tradeoff-based multicell networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200398X.

Dandachi:2022:RCT

[1309] Ghina Dandachi, Sophie Cerf, Yassine Hadjadj-Aoul, Abdelkader Outtagarts, and Eric Rutten. A robust controltheory-based exploration strategy in deep reinforcement learning for virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004005.

Azari:2022:ADL

[1310] Bahar Azari, Hai Cheng, Nasim Soltani, Haoqing Li, Yanyu Li, Mauro Belgiovine, Tales Imbiriba, Salvatore D’Oro, Tommaso Melodia, Yanzhi Wang, Pau Closas, Kaushik Chowdhury, and Deniz Erdoğmuş. Automated deep learning-based wide-band receiver. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004017.

Li:2022:RNT

[1311] Wenhao Li, Xiao-Yu Zhang, Huaifeng Bao, Qiang Wang, and Zhaoxuan Li. Robust network traffic identification with graph matching. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004029.

Zhang:2022:DDP

[1312] Ziming Zhang, Xiaolong Xu, and Fu Xiao. 5GMEC-DP: Differentially private protection of trajectory data based on 5G-based mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004108.

Bella:2022:DIA

[1313] Giampaolo Bella, Pietro Biondi, Gianpiero Costantino, and Ilaria Matteucci. Designing and implementing an AUTOSAR-based Basic Software Module for enhanced security. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200411X.

Ou:2022:OCC

[1314] Wei Ou, Shiying Huang, Jingjing Zheng, Qionglu Zhang, Guang Zeng, and Wenbao Han. An overview on cross-chain: Mechanism, platforms, challenges and advances. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004121.

Samikwa:2022:AAR

[1315] Eric Samikwa, Antonio Di Maio, and Torsten Braun. ARES: Adaptive resource-aware split learning for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004145.

Koita:2022:GLS

[1316] Moussa Koïta, Youssouf M. Diagana, Oumar Y. Maïga, and Mamadou K. Traore. A generic learning simulation framework to assess security strategies in cyber-physical production systems. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004157.

Liu:2022:AGA

[1317] Jianmin Liu, Qi Wang, and Yongjun Xu. AR-GAIL: Adaptive routing protocol for FANETs using generative adversarial imitation learning. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004169.

Zhang:2022:RDA

[1318] Junye Zhang, Peng Yu, Fanqin Zhou, Lei Feng, Wenjing Li, and Xuesong Qiu. Resource and delay aware finegrained service offloading in collaborative edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004170.

Alwan:2022:TSC

[1319] Ahmed A. Alwan, Allan J. Brimicombe, Mihaela Anca Ciupala, Seyed Ali Ghorashi, Andres Baravalle, and Paolo Falcarin. Timeseries clustering for sensor fault detection in large-scale cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004182.

Peng:2022:SOT

[1320] Yabin Peng, Caixia Liu, Shuxin Liu, Yuchen Liu, and Yiteng Wu. Smart-TRO: Optimizing topology robustness for Internet of Things via deep reinforcement learning with graph convolutional networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004194.

Parsa:2022:JPA

[1321] Ali Parsa, Neda Moghim, and Pouyan Salavati. Joint power allocation and MCS selection for energy-efficient link adaptation: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004200.

Lyu:2022:CTE

[1322] Minzhao Lyu, Hassan Habibi Gharakheili, and Vijay Sivaraman. Classifying and tracking enterprise assets via dual-grained network behavioral analysis. Computer Networks (Amsterdam, Netherlands: 1999), 218(??): ??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004212.

Sarraf-Maralaniyan:2022:MNM

[1323] Sarah Sarraf-Maralaniyan and Akbar Ghaffarpour Rahbar. MT^{3}A: a novel multicast routing, spectrum and modulation-level assignment in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004224.

Lopes:2022:PAT

[1324] Rafael Lopes, Denis Rosário, Eduardo Cerqueira, Helder Oliveira, and Sherali Zeadally. Priority-aware traffic routing and resource allocation mechanism for space-division multiplexing elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004236.

Wilhelmi:2022:AES

[1325] Francesc Wilhelmi, Lorenza Giupponi, and Paolo Dini. Analysis and evaluation of synchronous and asynchronous FLchain. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004248.

Nandhini:2022:LEE

[1326] P. S. Nandhini, S. Kuppuswami, S. Malliga, and R. DeviPriya. A lightweight energy-efficient algorithm for mitigation and isolation of internal rank attackers in RPL based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200425X.

Shan:2022:LBE

[1327] Minghui Shan, Sheng Zhang, Mingjun Xiao, and Yanchao Zhao. LENS: Bandwidth-efficient video analytics with adaptive super resolution. Computer Networks (Amsterdam, Netherlands: 1999), 218(??): ??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004261.

Uras:2022:MAR

[1328] Marco Uras, Enrico Ferrara, Raimondo Cossu, Antonio Liotta, and Luigi Atzori. MAC address de-randomization for WiFi device counting: Combining temporal- and content-based fingerprints. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004273.

Ghosal:2022:SAS

[1329] Amrita Ghosal, Subir Halder, and Mauro Conti. Secure over-the-air software update for connected vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004285.

Ranjan:2022:TSD

[1330] Shashi Ranjan, Pranav Jha, Abhay Karandikar, and Prasanna Chaporkar. Two stage downlink scheduling for balancing QoS in multihop IAB networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004297.

Fedrizzi:2022:AGB

[1331] Riccardo Fedrizzi, Rasoul Behravesh, Cristina Costa, and Fabrizio Granelli. An adaptive GA-based slice provisioning method for vertical industries in 5G and beyond networks. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004315.

Huang:2022:QDV

[1332] Guanglun Huang, Jianming Liu, Baoxian Zhang, and Cheng Li. Qualitydriven video streaming for ultradense OFDMA heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004327.

Ramineni:2022:PER

[1333] Chaturasan Ramineni, T. G. Venkatesh, and Lokesh Bommisetty. Performance evaluation of random access in narrow band Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004339.

Yadav:2022:IPS

[1334] Awaneesh Kumar Yadav, Manoj Misra, Pradumn Kumar Pandey, An Braeken, and Madhusanka Liyange. An improved and provably secure symmetric-key based 5G-AKA Protocol. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004340.

Rathee:2022:AIA

[1335] Geetanjali Rathee, Chaker Abdelaziz Kerrache, and Carlos T. Calafate. An ambient intelligence approach to provide secure and trusted Pub/Sub messaging systems in IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004352.

Tu:2022:BBT

[1336] Zhe Tu, Huachun Zhou, Kun Li, Haoxiang Song, and Yuzheng Yang. A blockchain-based trust and reputation model with dynamic evaluation mechanism for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004388.

Ye:2022:MOT

[1337] Weidu Ye, Junzhou Luo, Wenjia Wu, Feng Shan, and Ming Yang. MU-TAA: an online trajectory optimization and task scheduling for UAV-aided edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200439X.

Shahid:2022:DRC

[1338] Syed Maaz Shahid and Sungoh Kwon. Distributed robust channel allocation for clustered cognitive radio-based IoT networks using graph theory. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004406.

Alani:2022:PCB

[1339] Mohammed M. Alani and Hissam Tawfik. PhishNot: a cloud-based machinelearning approach to phishing URL detection. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004418.

Hadi:2022:JRA

[1340] Majid Hadi and Reza Ghazizadeh.

Joint resource allocation, user clustering and 3-d location optimization in multi-UAV-enabled mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 218 (??):??, December 9, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004546.

Tian:2022:RPB

[1341] Chuang Tian, Qi Jiang, Teng Li, Junwei Zhang, Ning Xi, and Jianfeng Ma. Reliable PUF-based mutual authentication protocol for UAVs towards multi-domain environment. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004558.

Anonymous:2022:EBq

[1342] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128622004753.

Anonymous:2022:Da

[1343] Anonymous. 9 December 2022. Computer Networks (Amsterdam, Netherlands: 1999), 218(??):??, December 9, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pinyoanuntapong:2022:ETN

[1344] Pinyarash Pinyoanuntapong, Prabhu Janakaraj, Ravikumar Balakrishnan, Minwoo Lee, Chen Chen, and Pu Wang. EdgeML: Towards networkaccelerated federated learning over wireless edge. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004303.

Halgamuge:2022:ESP

[1345] Malka N. Halgamuge. Estimation of the success probability of a malicious attacker on blockchain-based edge network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004364.

Ruchel:2022:ERS

[1346] Lucas V. Ruchel, Rogério C. Turchetti, and Edson T. de Camargo. Evaluation of the robustness of SDN controllers ONOS and ODL. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004376.

Mihailovic:2022:CMT

[1347] Andrej Mihailovic, Benyamin Abrishamchi, Mohammad Farhoudi, and A. Hamid Aghvami. A comprehensive multi-topology minimum set cover link-state routing approach for emerging random all-I access network topologies. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004522.

Huang:2022:LLW

[1348] Kaiwen Huang, Ronghui Hou, and Yingming Zeng. LWSBFT: Leaderless weakly synchronous BFT protocol. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004534.

Debski:2022:RTS

[1349] Roman Debski and Rafał Drezewski. Real-time surrogate-assisted preprocessing of streaming sensor data. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S138912862200456X.

Latah:2022:CDC

[1350] Majd Latah and Kübra Kalkan. CWT-DPA: Component-wise waiting time for BC-enabled data plane authentication. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004571.

Li:2022:HRN

[1351] Tengyao Li, Kaiyue Liu, Wangxin Feng, Chunfang Yang, and Xiangyang Luo. HeteroTiC: a robust network flow watermarking based on heterogeneous time channels. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004583.

Li:2022:FFM

[1352] Zhao Li, Bigui Zhang, Chengyu Liu, Zhixian Chang, Kang G. Shin, and Zheng Yan. Framed Fidelity MAC: Losslessly packing multi-user transmissions in a virtual point-to-point framework. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004595.

Lei:2022:QOM

[1353] Jianjun Lei, Lu Li, and Ying Wang. QoS-oriented media access control using reinforcement learning for next-generation WLANs. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004601.

Tripathy:2022:EAI

[1354] Pujasuman Tripathy and P. M. Khilar. An ensemble approach for improving localization accuracy in wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004613.

Attanasio:2022:DSR

[1355] Giulia Attanasio, Claudio Fiandrino, Marco Fiore, Joerg Widmer, Norbert Ludant, Bastian Bloessl, Konstantinos Kousias, Özgü Alay, Lise Jacquot, and Razvan Stanica. In-depth study of RNTI management in mobile networks: Allocation strategies and implications on data trace analysis. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004625.

Yi:2022:TPM

[1356] Yuzi Yi, Nafei Zhu, Jingsha He, Anca Delia Jurcut, and Bin Zhao. Toward pragmatic modeling of privacy information propagation in online social networks. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004637.

Sahni:2022:DRS

[1357] Yuvraj Sahni, Jiannong Cao, Lei Yang, and Shengwei Wang. Distributed resource scheduling in edge computing: Problems, solutions, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004649.

Hasselquist:2022:QTF

[1358] David Hasselquist, Christoffer Lindström, Nikita Korzhitskii, Niklas Carlsson, and Andrei Gurtov. QUIC throughput and fairness over dual connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004650.

Lia:2022:NPD

[1359] Gianmarco Lia, Marica Amadeo, Giuseppe Ruggeri, Claudia Campolo, Antonella Molinaro, and Valeria Loscrì. In-network placement of delayconstrained computing tasks in a softwarized intelligent edge. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004662.

Ashraf:2022:FCV

[1360] M. Wasim Abbas Ashraf, Chuanhe Huang, Arif Raza, Kashif Sharif, Md Monjurul Karim, and Shidong Huang. Forwarding and caching in video streaming over ICSDN: a clean-slate publish–subscribe approach. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004674.

Enge:2022:OMA

[1361] Alexander Enge, Abylay Satybaldy, and Mariusz Nowostawski. An offline mobile access control system based on self-sovereign identity standards. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004686.

Wang:2022:SBF

[1362] Hui Wang, Xiaowei Zhang, Ming Li, and Donghui Hu. SRRS: a blockchain fast propagation protocol based on non-Markovian process. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004698.

Hong:2022:RPN

[1363] Hsiang-Jen Hong, Wenjun Fan, Simeon Wuthier, Jinoh Kim, C. Edward Chow, Xiaobo Zhou, and Sang-Yoon Chang. Robust P2P networking connectivity estimation engine for permissionless Bitcoin cryptocurrency. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004704.

Solomon:2022:SEL

[1364] Yossi Solomon, Osnat Mokryn, and Tsvi Kuflik. SDNSandbox — enabling learning-based innovation in provider networks. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004807.

Navarro:2022:HRN

[1365] Jose M. Navarro, Alexis Huet, and Dario Rossi. Human readable network troubleshooting based on anomaly detection and feature scoring. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004819.

Grasso:2022:HHL

[1366] Christian Grasso, Raoul Raftopoulos, Giovanni Schembra, and Salvatore Serrano. H-HOME: a learning framework of federated FANETs to provide edge computing to future delayconstrained IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004832.

Fan:2022:EEI

[1367] Linna Fan, Lin He, Enhuan Dong, Jiahai Yang, Chenglong Li, Jinlei Lin, and Zhiliang Wang. EvoIoT: an evolutionary IoT and non-IoT classification model in open environments. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004844.

Khoshkholghi:2022:EIS

[1368] Mohammad Ali Khoshkholghi and Toktam Mahmoodi. Edge intelligence for service function chain deployment in NFV-enabled networks. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004856.

Guarino:2022:CCM

[1369] Idio Guarino, Giuseppe Aceto, Domenico Ciuonzo, Antonio Montieri, Valerio Persico, and Antonio Pescapè. Contextual counters and multimodal deep learning for activity-level traffic classification of mobile communication apps during COVID-19 pandemic. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004868.

Busacca:2022:MMD

[1370] F. Busacca, L. Galluccio, and S. Palazzo. A marketplace model for drone-assisted edge computing in 5G scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200487X.

Sindjoung:2022:MAB

[1371] Miguel Landry Foko Sindjoung, Mthulisi Velempini, and Alain Bertrand Bomgni. A MEC architecture for a better quality of service in an Autonomous Vehicular Network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004881.

Jagannath:2022:CSR

[1372] Anu Jagannath, Jithin Jagannath, and Prem Sagar Pattanshetty Vasanth Kumar. A comprehensive survey on radio frequency (RF) fingerprinting: Traditional approaches, deep learning, and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004893.

Voulgaridis:2022:IDC

[1373] Konstantinos Voulgaridis, Thomas Lagkas, Constantinos Marios Angelopoulos, and Sotiris E. Nikoletseas. IoT and digital circular economy: Principles, applications, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200490X.

Zheng:2022:IBN

[1374] Xiaoang Zheng, Aris Leivadeas, and Matthias Falkner. Intent Based Networking management with conflict detection and policy resolution in an enterprise network. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004911.

Mohammadi:2022:DSO

[1375] Marzieh Mohammadi, S. Ahmad Motamedi, and Saeed Sharifian. Drops on surface optimization (DSO): a new reinforcement learning based metaheuristic algorithm for virtual network functions placement in distributed cloud architectures. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004935.

Bagnulo:2022:WLM

[1376] Marcelo Bagnulo and Alberto García-Martínez. When less is more: BBR versus LEDBAT++. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004947.

Zou:2022:ECD

[1377] Hongcheng Zou, Jinshu Su, Ziling Wei, Shuhui Chen, and Baokang Zhao. An efficient cross-domain few-shot website fingerprinting attack with Brownian distance covariance. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004959.

Ayoub:2022:EAI

[1378] Omran Ayoub, Nicola Di Cicco, Fatima Ezzeddine, Federica Bruschetta, Roberto Rubino, Massimo Nardecchia, Michele Milano, Francesco Musumeci, Claudio Passera, and Massimo Tornatore. Explainable artificial intelligence in communication networks: a use case for failure identification in microwave networks. Computer Networks (Amsterdam, Netherlands: 1999), 219 (??):??, December 24, 2022. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200500X.

Anonymous:2022:EBr

[1379] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862200528X.

Anonymous:2022:Db

[1380] Anonymous. 24 December 2022. Computer Networks (Amsterdam, Netherlands: 1999), 219(??):??, December 24, 2022. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wan:2023:CPM

[1381] Zheng Wan and Xiaogang Dong. Computation power maximization for mobile edge computing enabled dense network. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622004923.

Avanzato:2023:DBA

[1382] Roberta Avanzato, Francesco Beritelli, and Corrado Rametta. An DL-based approach for packet error compensation using radio mobile network quality parameters in a rainfall scenario. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004972.

Anedda:2023:SSC

[1383] Matteo Anedda, Mauro Fadda, Roberto Girau, Giovanni Pau, and Daniele Giusto. A social smart city for public and private mobility: a real case study. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004984.

Baidas:2023:NCU

[1384] Mohammed W. Baidas, Ahmed M. AbdelGhaffar, and Emad Alsusa. Network-coded uplink clustered NOMA relay networks: Models and performance comparisons. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622004996.

Luxemburk:2023:FGT

[1385] Jan Luxemburk and Tomás Cejka. Fine-grained TLS services classification with reject option. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005011.

Sharvari:2023:CCC

[1386] N. P. Sharvari, Dibakar Das, Jyotsna Bapat, and Debabrata Das. Connectivity and collision constrained opportunistic routing for emergency communication using UAV. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005023.

Darbandeh:2023:SSA

[1387] Foroozan Ghosairi Darbandeh and Masoumeh Safkhani. SAPWSN: a secure authentication protocol for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005035.

Ji:2023:AQA

[1388] Xiaolan Ji, Biao Han, Cao Xu, Congxi Song, and Jinshu Su. Adaptive QoS-aware multipath congestion control for live streaming. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005047.

Perdices:2023:WBP

[1389] Daniel Perdices, Jorge E. López de Vergara, Iván González, and Luis de Pedro. Web browsing privacy in the deep learning era: Beyond VPNs and encryption. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005059.

Xu:2023:CDM

[1390] Zhilin Xu, Hao Sun, and Weibin Han. A collaboration-driven mechanism for AI diagnose with multiple requesters under incomplete information. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005060.

FeCastro:2023:UGU

[1391] Igor C. F.e Castro, Eduardo P. M. Câmara, Marcos A. M. Vieira, and Luiz F. M. Vieira. UW-GRE: Underwater greedy geographic routing by network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005072.

Foroughi:2023:AAD

[1392] Parisa Foroughi, Frank Brockners, and Jean-Louis Rougier. ADT: AIdriven network telemetry processing on routers. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005084.

Zhang:2023:LCV

[1393] Shengyu Zhang and Kwan L. Yeung. Location constrained virtual optical network embedding in spacedivision multiplexing elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005096.

Robles-Enciso:2023:MLG

[1394] Alberto Robles-Enciso and Antonio F.

Skarmeta. A multi-layer guided reinforcement learning-based tasks offloading in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005102.

Mehmood:2023:IDA

[1395] Kashif Mehmood, Katina Kralevska, and David Palma. Intent-driven autonomous network and service management in future cellular networks: a structured literature review. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005114.

Lira:2023:SAU

[1396] Clayton J. N. Lira, Raul C. Almeida, and Daniel A. R. Chaves. Spectrum allocation using multiparameter optimization in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005126.

Biomo:2023:NRP

[1397] Jean-Daniel Medjo Me Biomo, Thomas Kunz, and Marc St-Hilaire. A novel routing protocol for reducing packet delay with multi-beam antennas. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005138.

Qi:2023:LTB

[1398] Liuling Qi, Junfeng Tian, Mengjia Chai, and Hongyun Cai. LightPoW: a trust based time-constrained PoW for blockchain in Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200514X.

Ariemma:2023:LLS

[1399] Lorenzo Ariemma, Alessandro Dell’Orco, Simone Liotta, Massimo Candela, and Giuseppe Di Battista. Long-lasting sequences of BGP updates. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005151.

Liu:2023:JCZ

[1400] Zhixin Liu, Xi Chen, Yi Yang, Kit Yan Chan, and Yazhou Yuan. Joint cell zooming and sleeping strategy in ultra dense heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005163.

Kazemi:2023:OSF

[1401] Seyed Mojtaba Kazemi, Shamsollah Ghanbari, Manochehr Kazemi, and Mohamed Othman. Optimum scheduling in fog computing using the Divisible Load Theory (DLT) with linear and nonlinear loads. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005175.

Jiang:2023:AAP

[1402] Wenhao Jiang and Yuebin Bai.

APGNN: Alarm propagation graph neural network for fault detection and alarm root cause analysis. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005199.

Ma:2023:LDT

[1403] Dongchao Ma, Pengyu Wang, Lihua Song, Wenlong Chen, Li Ma, Mingwei Xu, and Laizhong Cui. A lightweight deployment of TD routing based on SD-WANs. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005205.

Venkateswararao:2023:NPC

[1404] Kuna Venkateswararao, Pravati Swain, Shashi Shekhar Jha, Iacovos Ioannou, and Andreas Pitsillides. A novel power consumption optimization framework in 5G heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005217.

Deng:2023:BSG

[1405] Hanqiang Deng, Jian Huang, Quan Liu, Cong Zhou, and Jialong Gao. BGSD: a SBERT and GAT-based service discovery framework for heterogeneous distributed IoT. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005229.

Wang:2023:CAF

[1406] Zhiyuan Wang, Hongli Xu, Yang Xu, Zhida Jiang, and Jianchun Liu. CoopFL: Accelerating federated learning with DNN partitioning and offloading in heterogeneous edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005242.

Sharma:2023:ISS

[1407] Rohit Sharma, Danda B. Rawat, Amiya Nayak, Sheng-Lung Peng, and Qin Xin. Introduction to the special section on survivability analysis of wireless networks with performance evaluation (VSI-networks survivability). Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005321.

Vrind:2023:FBA

[1408] Tushar Vrind, Chandan Kumar, Lalit Pathak, and Debabrata Das. Feedbackbased algorithm for aerial cell’s trajectory using deep learning for efficient performance in 6G. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005333.

Yu:2023:ESP

[1409] Jiayi Yu, Anfeng Liu, Neal N. Xiong, Shaobo Zhang, Tian Wang, and Mianxiong Dong. Employing social participants for timely data collection using pub/sub solutions in dynamic IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005357.

Bonati:2023:OGA

[1410] Leonardo Bonati, Michele Polese, Salvatore D’Oro, Stefano Basagni, and Tommaso Melodia. OpenRAN Gym: AI/ML development, data collection, and testing for O-RAN on PAWR platforms. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005369.

Esubonteng:2023:EIA

[1411] Paa Kwesi Esubonteng and Roberto Rojas-Cessa. Effect of the incident angle of a transmitting laser light on the coverage of a NLOS-FSO network. Computer Networks (Amsterdam, Netherlands: 1999), 220 (??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005382.

Sun:2023:VCB

[1412] Sheng Sun, Zengqi Zhang, Quyang Pan, Min Liu, and Zhongcheng Li. Vehicle-cluster-based opportunistic relays for data collection in intelligent transportation systems. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005436.

Xiong:2023:FTE

[1413] Runqun Xiong, Chuan Liang, Huajun Zhang, Xiangyu Xu, and Junzhou Luo. FlyingLoRa: Towards energy efficient data collection in UAVassisted LoRa networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862200545X.

Yan:2023:NMA

[1414] Yan Yan, Baoxian Zhang, and Cheng Li. A networked multi-agent reinforcement learning approach for cooperative FemtoCaching assisted wireless heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005473.

Anonymous:2023:EBa

[1415] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005795.

Anonymous:2023:Ja

[1416] Anonymous. January 2023. Computer Networks (Amsterdam, Netherlands: 1999), 220(??):??, January 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pattaranantakul:2023:SFC

[1417] Montida Pattaranantakul, Chalee Vorakulpipat, and Takeshi Takahashi. Service function chaining security survey: Addressing security challenges and threats. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005187.

Maier:2023:LMS

[1418] Markus Maier and Johanna Ullrich. In the loop: a measurement study of persistent routing loops on the IPv4/ IPv6 Internet. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005345.

Ibrahim:2023:DPS

[1419] Rita Ibrahim, Mohamad Assaad, and Berna Sayrac. A dynamic policy for selecting D2D mobile relays. Computer Networks (Amsterdam, Netherlands:

1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005394.

Bao:2023:MMC

[1420] Yu Bao, Yonggang Li, Liang Zhao, Aijuan Zhang, and Yutao Wang. MCEN: Maximum cooperative equilibrium WSN based on greedy prediction to reduce opposite transmission. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005400.

Tuan:2023:UDD

[1421] Tong Anh Tuan, Nguyen Viet Anh, Tran Thi Luong, and Hoang Viet Long. UTL DGA22 — a dataset for DGA botnet detection and classification. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005424.

Peng:2023:SPA

[1422] Chengzong Peng, Danyang Zheng, Yihan Zhong, and Xiaojun Cao. Offsite protection against service function forwarder failures in NFV. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005448.

Nunez-Gomez:2023:HBS

[1423] Carlos Núñez-Gómez, Carmen Carrión, Blanca Caminero, and Francisco M. Delicado. S-HIDRA: a blockchain and SDN domain-based architecture to orchestrate fog computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005461.

Giuliano:2023:MMS

[1424] Romeo Giuliano, Anna Maria Vegni, Valeria Loscrí, Eros Innocenti, Alessandro Vizzarri, and Franco Mazzenga. MuSLi: a multi sensor LiDAR detection for C-V2X networks. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005485.

Ramezanpour:2023:SPV

[1425] Keyvan Ramezanpour, Jithin Jagannath, and Anu Jagannath. Security and privacy vulnerabilities of 5G/ 6G and WiFi 6: Survey and research directions from a coexistence perspective. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005497.

Gong:2023:ESA

[1426] Kai Gong, Dong Yang, Weiting Zhang, and Jie Ren. An efficient scheduling approach for multi-level industrial chain flows in time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005503.

Du:2023:EFL

[1427] Haizhou Du, Yijian Chen, Xiaojie Feng, Qiao Xiang, and Haoyu Liu. An efficient federated learning framework for multi-channeled mobile edge network with layered gradient compression. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005515.

Xu:2023:TTT

[1428] Luyang Xu, Haoyu Liu, Junping Song, Rui Li, Yahui Hu, Xu Zhou, and Paul Patras. TransMUSE: Transferable Traffic Prediction in MUlti-Service Edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005527.

Waqas:2023:DSA

[1429] Muhammad Waqas, Shanshan Tu, Jialin Wan, Talha Mir, Hisham Alasmary, and Ghulam Abbas. Defense scheme against advanced persistent threats in mobile fog computing security. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005539.

Takahashi:2023:SMS

[1430] Tomoki Takahashi, Takehiro Sato, and Eiji Oki. Scheduling model for simultaneous update of multiple service function chains with state consistency. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005540.

Carvalho:2023:MPL

[1431] Rodrigo Carvalho, Noélia Correia, and Faroq Al-Tam. Mobility planning of LoRa gateways for edge storage of IoT data. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005552.

Liu:2023:OMU

[1432] Jiadi Liu, Songtao Guo, Quyuan Wang, Chengsheng Pan, and Li Yang. Optimal multi-user offloading with resources allocation in mobile edge cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005564.

Ramos:2023:WCP

[1433] Javier Ramos, José Luis García-Dorado, and Javier Aracil. Workforce capacity planning for proactive troubleshooting in the Network Operations Center. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005576.

Zhang:2023:RMF

[1434] Shengyu Zhang and Kwan L. Yeung.

Revisiting the modulation format selection problem in crosstalk-aware SDM-EONs. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005588.

Garrido-Hidalgo:2023:EOR

[1435] Celia Garrido-Hidalgo, Luis Roda-Sanchez, F. Javier Ramírez, Antonio Fernández-Caballero, and Teresa Olivares. Efficient online resource allocation in large-scale LoRaWAN networks: a multi-agent approach. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200559X.

Salehi:2023:ADC

[1436] Shahab Salehi, Hamed Farbeh, and Alireza Rokhsari. An adaptive data coding scheme for energy consumption reduction in SDN-based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862200562X.

Zhou:2023:SMA

[1437] Boqing Zhou, Sujun Li, Jianxin Wang, Yun Cheng, and Jie Wu. A secure model against mobile sink replication attacks in unattended sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005631. See

corrigendum [1764].

Luglio:2023:FWT

[1438] M. Luglio, M. Quadrini, C. Roseti, and F. Zampognaro. A flexible Web traffic generator for the dimensioning of a 5G backhaul in NPN. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005655.

Roy:2023:APM

[1439] Prasanta Kumar Roy and Ansuman Bhattacharya. An anonymitypreserving mobile user authentication protocol for global roaming services. Computer Networks (Amsterdam, Netherlands: 1999), 221 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005667.

Anonymous:2023:EBb

[1440] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623000129.

Anonymous:2023:Fa

[1441] Anonymous. February 2023. Computer Networks (Amsterdam, Netherlands: 1999), 221(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Farquhar:2023:MRF

[1442] Collin Farquhar, Swatantra Kafle, Kian Hamedani, Anu Jagannath, and Jithin Jagannath. Marconi–Rosenblatt framework for intelligent networks (MRiNet Gym): For rapid design and implementation of distributed multiagent reinforcement learning solutions for wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005230.

Liu:2023:GEB

[1443] Xiuwen Liu and Yanjiao Chen. Group effect-based privacy-preserving data aggregation for mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005412.

Benlloch-Caballero:2023:DDL

[1444] Pablo Benlloch-Caballero, Qi Wang, and Jose M. Alcaraz Calero. Distributed dual-layer autonomous closed loops for self-protection of 5G/6G IoT networks from distributed denial of service attacks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005606.

Shashin:2023:APS

[1445] Aleksei Shashin, Andrey Belogaev, Artem Krasilov, and Evgeny Khorov. Adaptive parameters selection for uplink grant-free URLLC transmission in 5G systems. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005618.

Xie:2023:YFB

[1446] Yi Xie, Xianliang Jiang, Guanghui Gong, Ziyi Jiang, Guang Jin, and Haiming Chen. Yinker: a flexible BBR to achieve the high-throughput and low-latency data transmission over Wi-Fi and 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005643.

Yu:2023:CPR

[1447] Mengxin Yu, Yibo Pi, Aimin Tang, and Xudong Wang. Coordinated parallel resource allocation for integrated access and backhaul networks. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005679.

Bak:2023:CAW

[1448] Charn doh Bak and Seung-Jae Han.

Contention alleviation in WiFi networks by using light-weight machine learning model. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005680.

Zhou:2023:CFA

[1449] Jianer Zhou, Zengxie Ma, Weijian Tu, Xinyi Qiu, Jingpu Duan, Zhenyu Li, Qing Li, Xinyi Zhang, and Weichao Li.

Cable: a framework for accelerating 5G UPF based on eBPF. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005692.

Jha:2023:JRP

[1450] Rakesh Kumar Jha, Mittal K. Pedhadiya, Anutusha Dogra, Haneet Kour, and Puja. Joint resource and power allocation for 5G enabled D2D networking with NOMA. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005709.

Li:2023:NSS

[1451] David Chunhu Li, Muhamad Rizka Maulana, and Li-Der Chou. NNSplit-SØREN: Supporting the model implementation of large neural networks in a programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005710.

Liu:2023:MEM

[1452] Yongbo Liu, Yongqiang Dong, Jun Shen, and Chong Feng. MAYA: Exploring multiform attributes of node to align YANG data models. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005722.

Seth:2023:DDD

[1453] Abhay Deep Seth, Santosh Biswas, and Amit Kumar Dhar. DAD-CNF: Diagnoser design for duplicate address detection threat using conjunctive normal form. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005734.

Tian:2023:TSE

[1454] Junfeng Tian, Jin Tian, and Hongwei Xu. TSBFT: a scalable and efficient leaderless Byzantine consensus for consortium blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005758.

Fu:2023:DCM

[1455] Yanming Fu, Xian Zhang, Xiaoqiong Qin, Qingwen Meng, and Bocheng Huang. Data collection of multiplayer cooperative game based on edge computing in mobile crowd sensing. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005850.

deNeira:2023:DDS

[1456] Anderson Bergamini de Neira, Burak Kantarci, and Michele Nogueira. Distributed denial of service attack prediction: Challenges, open issues and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005874.

Zhai:2023:MMS

[1457] Dong Zhai, Xiangru Meng, Zhenhua Yu, Hang Hu, and Yuan Liang. A migration method for service function chain based on failure prediction. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005886.

Zhang:2023:RES

[1458] Le Zhang and Ye Du. Resilience enhancement scheme for gateway placement in space information networks. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128622005898.

Wang:2023:CSA

[1459] Xueyi Wang, Xingwei Wang, Ying Shi, Dongkuo Wu, Lianbo Ma, and Min Huang. Core-selecting auctionbased mechanisms for service function chain provisioning and pricing in NFV markets. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000026.

Augello:2023:CBU

[1460] Alessandro Augello, Pierluigi Gallo, Eleonora Riva Sanseverino, Gaetano Sciabica, and Giuseppe Sciumè. Certifying battery usage for V2G and second life with a blockchain-based framework. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000038.

Yoshinaka:2023:DAL

[1461] Yutaro Yoshinaka, Junji Takemasa, Yuki Koizumi, and Toru Hasegawa. Design and analysis of lightweight anonymity protocol for host- and ASlevel anonymity. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300004X.

Hirsch:2023:DDG

[1462] Christian Hirsch, Luca Davoli, Radu Grosu, and Gianluigi Ferrari. Dyn-GATT: a dynamic GATT-based data synchronization protocol for BLE networks. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000051.

Nandyala:2023:QQL

[1463] Chandra Sukanya Nandyala, Hee-Won Kim, and Ho-Shin Cho. QTAR: a Q-learning-based topology-aware routing protocol for underwater wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 222 (??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000075.

Anonymous:2023:EBc

[1464] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623000336.

Anonymous:2023:Fb

[1465] Anonymous. February 2023. Computer Networks (Amsterdam, Netherlands: 1999), 222(??):??, February 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bi:2023:BAB

[1466] Xiang Bi, Xiaokai Sun, Zengwei Lyu, Benhong Zhang, and Xing Wei. A back adjustment based dependent task offloading scheduling algorithm with fairness constraints in VEC networks. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128622005862.

Zhu:2023:OTD

[1467] Konglin Zhu, Wentao Chen, Lei Jiao, Jiaxing Wang, Yuyang Peng, and Lin Zhang. Online training data acquisition for federated learning in cloud-edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000014.

Monfared:2023:DFD

[1468] Saleh Khalaj Monfared and Saeed Shokrollahi. DARVAN: a fully decentralized anonymous and reliable routing for VANets. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000063.

Qin:2023:SFC

[1469] Yudong Qin, Deke Guo, Lailong Luo, Jingyu Zhang, and Ming Xu. Service function chain migration with the long-term budget in dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 223 (??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000087.

Ju:2023:COT

[1470] Xiang Ju, Shengchao Su, Chaojie Xu, and Haoxuan Wang. Computation offloading and tasks scheduling for the Internet of Vehicles in edge computing: a deep reinforcement learningbased pointer network approach. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000178.

Sheng:2023:GDB

[1471] Siyuan Sheng, Qun Huang, and Patrick P. C. Lee. A general delta-based in-band network telemetry framework with extremely low bandwidth overhead. Computer Networks (Amsterdam, Netherlands: 1999), 223 (??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300018X.

Xu:2023:CTO

[1472] Dahu Xu and Ding Xu. Cooperative task offloading and resource allocation for UAV-enabled mobile edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000191.

Truong:2023:SPO

[1473] Truong Van Truong and Anand Nayyar. System performance and optimization in NOMA mobile edge computing surveillance network using GA and PSO. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000208.

Campolo:2023:SIE

[1474] Claudia Campolo, Giacomo Genovese, Gurtaj Singh, and Antonella Molinaro. Scalable and interoperable edge-based federated learning in IoT contexts. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300021X.

Aghapour:2023:TOR

[1475] Zahra Aghapour, Saeed Sharifian, and Hassan Taheri. Task offloading and resource allocation algorithm based on deep reinforcement learning for distributed AI execution tasks in IoT edge computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000221.

Li:2023:TDC

[1476] Shuyang Li, Gianluca Francini, and Enrico Magli. Temporal dynamics clustering for analyzing cell behavior in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000233.

Guler:2023:BEC

[1477] Evrim Guler, Murat Karakus, and Suleyman Uludag. Blockchainenhanced cross-ISP spectrum assignment framework in SDONs: SpectrumChain. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000245.

Patel:2023:LLD

[1478] Chintan Patel and Nishant Doshi. LDA-2IoT: a level dependent authentication using two factor for IoT paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000257.

Shi:2023:AUD

[1479] Yao Shi, Emad Alsusa, and Mohammed W. Baidas. On the application of uplink/downlink decoupled access in heterogeneous mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000385.

Hoang:2023:SAC

[1480] Tran Manh Hoang, Le The Dung, Ba Cao Nguyen, Nguyen Van Vinh, and Gia Thien Luu. Secrecy analysis of cooperative NOMA-FDR systems with imperfect CSI and colluding eavesdroppers. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000397.

Anonymous:2023:EBd

[1481] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000634.

Anonymous:2023:Ma

[1482] Anonymous. March 2023. Computer Networks (Amsterdam, Netherlands: 1999), 223(??):??, March 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Das:2023:CRM

[1483] Sree Krishna Das, Fatma Benkhelifa, Yao Sun, Hanaa Abumarshoud, Qammer H. Abbasi, Muhammad Ali Imran, and Lina Mohjazi. Comprehensive review on ML-based RISenhanced IoT systems: basics, research progress and future challenges. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000269.

Laroui:2023:SFC

[1484] Mohammed Laroui, Hatem Ibn Khedher, Hassine Moungla, and Hossam Afifi. Service function chains multiresource orchestration in virtual mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000270.

Shi:2023:JTP

[1485] KunLi Shi and Fei Wang. Joint task processing/offloading mode selection and resource-allocation for backscatter-aided and wireless-powered MEC. Computer Networks (Amsterdam, Netherlands: 1999), 224 (??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000294.

Chen:2023:AAQ

[1486] Xue Chen, Shiyuan Xu, Yibo Cao, Yunhua He, and Ke Xiao. AQRS: Anti-quantum ring signature scheme for secure epidemic control with blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000403.

Luconi:2023:IFM

[1487] Valerio Luconi and Alessio Vecchio.

Impact of the first months of war on routing and latency in Ukraine. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000415.

Li:2023:ROT

[1488] Mao Li, Wei Liu, and Jing Lei. A review on orthogonal time-frequency space modulation: State-of-art, hotspots and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000427.

Pirmoradian:2023:EEB

[1489] Fatemeh Pirmoradian, Masoumeh Safkhani, and Seyed Mohammad Dakhilalian. ECCPWS: an ECC-based protocol for WBAN systems. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000439.

Ahmad:2023:DCN

[1490] Rami Ahmad, Matti Hämäläinen, Raniyah Wazirali, and Tarik Abu-Ain. Digital-care in next generation networks: Requirements and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000440.

Zhou:2023:PLP

[1491] Menglun Zhou, Yifeng Zheng, Songlei Wang, Zhongyun Hua, Hejiao Huang, Yansong Gao, and Xiaohua Jia. PPTA: a location privacy-preserving and flexible task assignment service for spatial crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000452.

Longo:2023:DIA

[1492] Edoardo Longo and Alessandro E. C. Redondi. Design and implementation of an advanced MQTT broker for distributed pub/sub scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000464.

Zhu:2023:IIL

[1493] Wenbin Zhu, Xiuli Ma, Yanliang Jin, and Rui Wang. ILETC: Incremental learning for encrypted traffic classification using generative replay and exemplar. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000476.

Raju:2023:DEA

[1494] Mekala Ratna Raju and Sai Krishna Mothku. Delay and energy aware task scheduling mechanism for fogenabled IoT applications: a reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000488.

Jiang:2023:BBF

[1495] Tongtong Jiang, Guowei Shen, Chun Guo, Yunhe Cui, and Bo Xie. BFLS: Blockchain and federated learning for sharing threat detection models as cyber threat intelligence. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300049X.

Ali:2023:SHA

[1496] Yasir Ali and Habib Ullah Khan.

A survey on harnessing the applications of mobile computing in healthcare during the COVID-19 pandemic: Challenges and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000506.

Gartner:2023:FIR

[1497] Christoph Gärtner, Amr Rizk, Boris Koldehofe, René Guillaume, Ralf Kundel, and Ralf Steinmetz. Fast incremental reconfiguration of dynamic timesensitive networks at runtime. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000518.

Song:2023:DEA

[1498] Cha Hwan Song, Xin Zhe Khooi, Dinil Mon Divakaran, and Mun Choon Chan. DySO: Enhancing application offload efficiency on programmable switches. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300052X.

Chen:2023:MMS

[1499] Ye Chen, Yingxu Lai, Zhaoyi Zhang, Hanmei Li, and Yuhang Wang. MDFD: a multi-source data fusion detection framework for Sybil attack detection in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000531.

Bruhn:2023:PIT

[1500] Philipp Bruhn, Mirja Kühlewind, and Maciej Muehleisen. Performance and improvements of TCP CUBIC in lowdelay cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000543.

Toledo:2023:CFA

[1501] Kfir Toledo, David Breitgand, Dean Lorenz, and Isaac Keslassy. CloudPilot: Flow acceleration in the cloud. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000555.

Bakhshi:2023:MPN

[1502] Bahador Bakhshi, Josep Mangues-Bafalluy, and Jorge Baranda. Multiprovider NFV network service delegation via average reward reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000567.

Yu:2023:LUL

[1503] Sungjin Yu, Joonyoung Lee, Anil Kumar Sutrala, Ashok Kumar Das, and Youngho Park. LAKA-UAV: Lightweight authentication and key agreement scheme for cloud-assisted unmanned aerial vehicle using blockchain in flying ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000579.

Diao:2023:EGE

[1504] Zulong Diao, Gaogang Xie, Xin Wang, Rui Ren, Xuying Meng, Guangxing Zhang, Kun Xie, and Mingyu Qiao. EC-GCN: a encrypted traffic classification framework based on multi-scale graph convolution networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000592.

Huang:2023:EDA

[1505] Ying Huang, Lang Li, Ying Guo, Yu Ou, and Xiantong Huang. An efficient differential analysis method based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000671.

Arribas:2023:OFC

[1506] Edgar Arribas, Vincenzo Mancuso, and Vicent Cholvi. Optimizing fairness in cellular networks with mobile drone relays. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000683.

Aghasi:2023:DAM

[1507] Ali Aghasi, Kamal Jamshidi, Ali Bohlooli, and Bahman Javadi. A decentralized adaptation of model-free Q-learning for thermal-aware energyefficient virtual machine placement in cloud data centers. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000695.

Attir:2023:LAM

[1508] Azeddine Attir, Farid Naït-Abdesselam, and Kamel Mohamed Faraoun. Lightweight anonymous and mutual authentication scheme for wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000701.

Mekdad:2023:SSP

[1509] Yassine Mekdad, Ahmet Aris, Leonardo Babun, Abdeslam El Fergougui, Mauro Conti, Riccardo Lazzeretti, and A. Selcuk Uluagac. A survey on security and privacy issues of UAVs. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000713.

Liu:2023:PNSa

[1510] Jiayi Liu, Wenbin Yao, Chen Wang, and Qinghai Yang. Provisioning network slice for mobile content delivery in uncertain MEC environment. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000749.

Peng:2023:DQL

[1511] Bao Peng, Zhi Yao, Xin Liu, and Guofu Zhou. Deep Q-learning multiple networks based dynamic spectrum access with energy harvesting for green cognitive radio network. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000750.

Yu:2023:CNO

[1512] Jiachao Yu, Chao Zhai, Yujun Li, and Lina Zheng. Cooperative non-orthogonal multiple access with D2D communication and opportunistic relaying in large-scale cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000762.

Bothra:2023:HCA

[1513] Priyanka Bothra, Raja Karmakar, Sanjukta Bhattacharya, and Sayantani De. How can applications of blockchain and artificial intelligence improve performance of Internet of Things? — A survey. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000798.

Zhang:2023:DGA

[1514] Hengwei Zhang, Yan Mi, Xiaohu Liu, Yuchen Zhang, Jindong Wang, and Jinglei Tan. A differential game approach for real-time security defense decision in scale-free networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000804.

Tokuda:2023:NSR

[1515] Kairi Tokuda, Takehiro Sato, and Eiji Oki. Network slice reconfiguration with deep reinforcement learning under variable number of service function chains. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000816.

Feng:2023:TPP

[1516] Zhenni Feng, Sijia Yu, and Yanmin Zhu. Towards personalized privacy preference aware data trading: a contract theory based approach. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000828.

Akram:2023:BBP

[1517] Muhammad Arslan Akram, Hira Ahmad, Adnan Noor Mian, Anca Delia Jurcut, and Saru Kumari. Blockchainbased privacy-preserving authentication protocol for UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300083X.

Brahmi:2023:TFA

[1518] Saloua Brahmi and Mohand Yazid. Towards a fair allocation and effective utilization of resource units in multi-user WLANs-based OFDMA technology. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000841.

Anonymous:2023:EBe

[1519] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001196.

Anonymous:2023:Aa

[1520] Anonymous. April 2023. Computer Networks (Amsterdam, Netherlands: 1999), 224(??):??, April 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Li:2023:PRA

[1521] Ziyong Li, Yuxiang Hu, Le Tian, and Zhao Lv. Packet rank-aware active queue management for programmable flow scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000774.

Hamnache:2023:JLB

[1522] Mohamed Hamnache, Rahim Kacimi, and André-Luc Beylot. Joint loadbalancing and power control strategy to maximize the data extraction rate of LoRaWAN networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000786.

Iqbal:2023:UQA

[1523] Faheem Iqbal, Moneeb Gohar, Hanen Karamti, Walid Karamti, Seok-Joo Koh, and Jin-Ghoo Choi. Use of QUIC for AMQP in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000853.

Wang:2023:GQA

[1524] Hua Wang, Adrián Pérez Aguilar, Almudena Díaz Zayas, Germán Corrales Madueño, Changming Zhang, Nan Hao, and Xianbin Yu. A general QoE assessment framework for applications and services. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000865.

Zhou:2023:CPD

[1525] Hongliang Zhou, Yifeng Zheng, Xiaohua Jia, and Jiangang Shu. Collaborative prediction and detection of DDoS attacks in edge computing: a deep learning-based approach with distributed SDN. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000877.

Jin:2023:ESP

[1526] Nansen Jin, Jinsong Gui, and Xinran Zhou. Equalizing service probability in UAV-assisted wireless powered mmWave networks for postdisaster rescue. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000890.

Chen:2023:BBA

[1527] Hao Chen, Hua Qin, Weimin Chen, Ni Li, Tao Wang, Jianxin He, Gelan Yang, and Yang Peng. BMS: Bandwidth-aware multi-interface scheduling for energy-efficient and delayconstrained gateway-to-device communications in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000907.

Valsamas:2023:VTB

[1528] Polychronis Valsamas, Sotiris Skaperas, Lefteris Mamatas, and Luis M. Contreras. Virtualization Technology Blending for resource-efficient edge clouds. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000919.

Malekghaini:2023:DLE

[1529] Navid Malekghaini, Elham Akbari, Mohammad A. Salahuddin, Noura Limam, Raouf Boutaba, Bertrand Mathieu, Stephanie Moteau, and Stephane Tuffin. Deep learning for encrypted traffic classification in the face of data drift: an empirical study. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000932.

Mabrouk:2023:IDG

[1530] Abdelfettah Mabrouk and Assia Naja.

Intrusion detection game for ubiquitous security in vehicular networks: a signaling game based approach. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000944.

Gupta:2023:FLU

[1531] Rajni Gupta and Juhi Gupta. Federated learning using game strategies: State-of-the-art and future trends. Computer Networks (Amsterdam, Netherlands: 1999), 225 (??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000956.

Jung:2023:TPO

[1532] Soyi Jung, Jae-Hyun Kim, David Mohaisen, and Joongheon Kim. Truthful and performance-optimal computation outsourcing for aerial surveillance platforms via learning-based auction. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623000968.

Dai:2023:GGL

[1533] Jianbang Dai, Xiaolong Xu, and Fu Xiao. GLADS: a global-local attention data selection model for multimodal multitask encrypted traffic classification of IoT. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300097X.

Ullah:2023:ITS

[1534] Farman Ullah, Muhammad Bilal, and Su-Kyung Yoon. Intelligent timeseries forecasting framework for nonlinear dynamic workload and resource prediction in cloud. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000981.

Markovitz:2023:DIR

[1535] Oren Markovitz and Michael Segal. Demand Island routing for LEO satellite constellations. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001007.

Sun:2023:JSS

[1536] Jiayu Sun, Huiqiang Wang, Lili Nie, Guangsheng Feng, Zhibo Zhang, and Jingyao Liu. A joint strategy for service deployment and task offloading in satellite-terrestrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001019.

Drainakis:2023:CFL

[1537] Georgios Drainakis, Panagiotis Pantazopoulos, Konstantinos V. Katsaros, Vasilis Sourlas, Angelos Amditis, and Dimitra I. Kaklamani. From centralized to federated learning: Exploring performance and end-to-end resource consumption. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001020.

Jiang:2023:SOR

[1538] Jinfang Jiang, Guangjie Han, and Chuan Lin. A survey on opportunistic routing protocols in the Internet of Underwater Things. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001032.

Hong:2023:ATE

[1539] Hsiang-Jen Hong, Sang-Yoon Chang, and Xiaobo Zhou. Auto-tune: an efficient autonomous multi-path payment routing algorithm for Payment Channel Networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001044.

Goscien:2023:TAS

[1540] Róza Goścień. Traffic-aware service relocation in software-defined and intentbased elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001056.

G:2023:NNR

[1541] Keerthan Kumar T. G., Sourav Kanti Addya, Anurag Satpathy, and Shashidhar G. Koolagudi. NORD: NOde Ranking-based efficient virtual network embedding over single Domain substrate networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001068.

Syed:2023:FCB

[1542] Naeem Firdous Syed, Mengmeng Ge, and Zubair Baig. Fog-cloud based intrusion detection system using Recurrent Neural Networks and feature selection for IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300107X.

Carpi:2023:EAR

[1543] Fabrizio Carpi, Marco Martalò, Luca Davoli, Antonio Cilfone, Yingjie Yu, Yi Wang, and Gianluigi Ferrari. Experimental analysis of RSSI-based localization algorithms with NLOS premitigation for IoT applications. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001081.

Deebak:2023:IDA

[1544] B. D. Deebak and Seong Oun Hwang. Intelligent drone-assisted robust lightweight multi-factor authentication for military zone surveillance in the 6G era. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001093.

Vidal:2023:USM

[1545] José-Ramón Vidal, Luis Guijarro, and Vicent Pla. A user subscription model in mobile radio access networks with network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300110X.

Chu:2023:ILD

[1546] Yi Chu, Paul Mitchell, David Grace, Jonathan Roberts, Dominic White, and Tautvydas Mickus. IRIS: a low duty cycle cross-layer protocol for long-range wireless sensor networks with low power budget. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001111.

Jo:2023:SAE

[1547] Hyun-Chul Jo, Hyun-Wook Jin, and Joongheon Kim. Self-adaptive endto-end resource management for realtime monitoring in cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 225 (??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001147.

Borgohain:2023:LDA

[1548] Ponjit Borgohain and Hiten Choudhury. A lightweight D2D authentication protocol for relay coverage scenario in 5G mobile network. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300124X.

Anonymous:2023:EBf

[1549] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001536.

Anonymous:2023:Ab

[1550] Anonymous. April 2023. Computer Networks (Amsterdam, Netherlands: 1999), 225(??):??, April 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Liu:2023:PNSb

[1551] Xiaoxue Liu, Yichuan Wang, Yanping Li, and Hao Cao. PTAP: a novel secure privacy-preserving and traceable authentication protocol in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000889.

Alqarni:2023:OBO

[1552] Manal Alqarni, Asma Cherif, and Entisar Alkayyal. ODM-BCSA: an offloading decision-making framework based on binary cuckoo search algorithm for mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000920.

Hassanpour:2023:PPE

[1553] Seyedeh Bahereh Hassanpour, Ahmad Khonsari, Masoumeh Moradian, and Seyed Pooya Shariatpanahi. Privacy-preserving edge caching: a probabilistic approach. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000993.

Akar:2023:PFS

[1554] Nail Akar and Ezhan Karasan. Is proportional fair scheduling suitable for age-sensitive traffic? Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001135.

Huang:2023:RFF

[1555] Xiaohong Huang, Lu Han, Dandan Li, Kun Xie, and Yong Zhang. A reliable and fair federated learning mechanism for mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001238.

Rengaraj:2023:SUS

[1556] Ramaprabha Rengaraj, Lavanya Dhamodharan Loganathan, and Gunaseelan Konganathan. Secure upgraded spatial modulation for wireless transmission systems. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001251.

Mushtaq:2023:KHA

[1557] Earum Mushtaq, Aneela Zameer, and Rubina Nasir. Knacks of a hybrid anomaly detection model using deep auto-encoder driven gated recurrent unit. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001263.

Singh:2023:OLA

[1558] Prabhjot Singh, Parulpreet Singh, Nitin Mittal, Urvinder Singh, and Supreet Singh. An optimum localization approach using hybrid TSN-MRA in 2D WSNs. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001275.

Jeon:2023:HNC

[1559] So-Eun Jeon, Sun-Jin Lee, and Il-Gu Lee. Hybrid in-network computing and distributed learning for large-scale data processing. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001317.

Castell-Uroz:2023:TLW

[1560] Ismael Castell-Uroz and Pere Barlet-Ros. TrackSign-labeled web tracking dataset. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001329.

Fazio:2023:NPA

[1561] Peppino Fazio, Miralem Mehic, Miroslav Voznak, Floriano De Rango, and Mauro Tropea. A novel predictive approach for mobility activeness in mobile wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001342.

Zheng:2023:DCT

[1562] Jiaxing Zheng and Li Pan. Dominant coverage for target users at the lowest cost under competitive propagation in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300138X.

Zeng:2023:SSW

[1563] Zijie Zeng, Lin Cui, Mimi Qian, Zhen Zhang, and Kaimin Wei. A survey on sliding window sketch for network measurement. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300141X.

Anonymous:2023:EBg

[1564] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300186X.

Anonymous:2023:Mb

[1565] Anonymous. May 2023. Computer Networks (Amsterdam, Netherlands: 1999), 226(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhang:2023:DSB

[1566] Hang Zhang, Dongqi Yan, Yanxi Zhang, Jiamu Liu, and Mingwu Yao. Distributed synchronization based on model-free reinforcement learning in wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001159.

Myneni:2023:USS

[1567] Sowmya Myneni, Kritshekhar Jha, Abdulhakim Sabur, Garima Agrawal, Yuli Deng, Ankur Chowdhary, and Dijiang Huang. Unraveled — a semisynthetic dataset for advanced persistent threats. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001330.

Shao:2023:MTD

[1568] Sujie Shao, Lili Su, Qinghang Zhang, Shuang Wu, Shaoyong Guo, and Feng Qi. Multi task dynamic edgeend computing collaboration for urban Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001354.

Gupta:2023:ABV

[1569] Naveen Kumar Gupta, Rama Shankar Yadav, Rajendra Kumar Nagaria, Deepak Gupta, Achyut Mani Tripathi, and Om Jee Pandey. Anchor-based void detouring routing protocol in three dimensional IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001366.

Zhang:2023:CDN

[1570] Chunrui Zhang, Gang Wang, Shen Wang, Dechen Zhan, and Mingyong Yin. Cross-domain network attack detection enabled by heterogeneous transfer learning. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001378.

Barakabitze:2023:CSN

[1571] Alcardo Alex Barakabitze and Ray Walshe. Corrigendum to “SDN and NFV for QoE-driven multimedia services delivery: the road towards 6G and beyond networks” COMPNW, Volume 214, September 2022, pp. 1–25, 109133. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001391. See

[1177].

Liu:2023:LBI

[1572] Wai-Xi Liu, Jun Cai, Ying-Hao Zhu, Jun-Ming Luo, and Jin Li. Load balancing inside programmable data planes based on network modeling prediction using a GNN with network behaviors. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001408.

Ji:2023:RLB

[1573] Pengshuo Ji, Jie Jia, Jian Chen, Liang Guo, An Du, and Xingwei Wang. Reinforcement learning based joint trajectory design and resource allocation for RIS-aided UAV multicast networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001421.

Hu:2023:WEB

[1574] Xinxin Hu, Hongchang Chen, Shuxin Liu, Haocong Jiang, Kai Wang, and Yahui Wang. Who are the evil backstage manipulators: Boosting graph attention networks against deep fraudsters. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001433.

Zidic:2023:APP

[1575] Dinko Zidic, Toni Mastelic, Ivana Nizetic Kosovic, Mario Cagalj, and Josip Lorincz. Analyses of ping-pong handovers in real 4G telecommunication networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001445.

Yuan:2023:AAA

[1576] Jingling Yuan, Nana Wang, Siqi Cai, Mincheng Chen, and Xinping Li. AMLFN-AD: Adaptive multilevel integrated fusion attack detection framework for intelligent building systems. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001457.

Stokkink:2023:WSA

[1577] Quinten Stokkink, Can Umut Ileri, Dick Epema, and Johan Pouwelse. Web3 Sybil avoidance using network latency. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001469.

Ugurlu:2023:APA

[1578] Onur Ugurlu. ADA-PC: an asynchronous distributed algorithm for minimizing pairwise connectivity in wireless multi-hop networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001482.

Guo:2023:AES

[1579] Hui Guo, Rui chang Shi, Ping li Gu, Jia lu Li, and Shu long Wang. Allocating edge service resources to the up-offloaded vehicle tasks in ICV environment. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001603.

Vizziello:2023:IBC

[1580] Anna Vizziello, Maurizio Magarini, Pietro Savazzi, and Laura Galluccio. Intra-body communications for nervous system applications: Current technologies and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001639.

Lu:2023:NTP

[1581] Yufei Lu, Qian Ning, Linyu Huang, and Bingcai Chen. A network traffic prediction model based on reinforced staged feature interaction and fusion. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001640.

Sarah:2023:RAM

[1582] Annisa Sarah, Gianfranco Nencioni, and Md. Muhidul I. Khan. Resource allocation in multi-access edge computing for 5G-and-beyond networks. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001652.

Anonymous:2023:ISIa

[1583] Anonymous. Introduction to the special issue on machine learning and artificial intelligence for the Internet of Things, 5G, and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001676.

Liu:2023:MNM

[1584] Jian Liu, Qingsai Xiao, Liling Xin, Qiuyun Wang, Yepeng Yao, and Zhengwei Jiang. M3F: a novel multisession and multi-protocol based malware traffic fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001688.

Mathur:2023:SRB

[1585] Shikha Mathur, Anshuman Kalla, Gürkan Gür, Manoj Kumar Bohra, and Madhusanka Liyanage. A survey on role of blockchain for IoT: Applications and technical aspects. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001718.

Alhamed:2023:PTC

[1586] Faris Alhamed, Davide Scano, Piero Castoldi, Juan Jose Vegas Olmos, Ilya Vershkov, Francesco Paolucci, and Filippo Cugini. P4 Telemetry collector. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300172X.

Anonymous:2023:EBh

[1587] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002098.

Anonymous:2023:Mc

[1588] Anonymous. May 2023. Computer Networks (Amsterdam, Netherlands: 1999), 227(??):??, May 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hasslinger:2023:OAM

[1589] Gerhard Hasslinger, Mahshid Okhovatzadeh, Konstantinos Ntougias, Frank Hasslinger, and Oliver Hohlfeld. An overview of analysis methods and evaluation results for caching strategies. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623000282.

Fei:2023:VEA

[1590] Teng Fei and Wenye Wang. The vulnerability and enhancement of AKA protocol for mobile authentication in LTE/5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001305.

Guo:2023:DOR

[1591] Zhongfu Guo, Xinsheng Ji, Wei You, Mingyan Xu, Yu Zhao, Zhimo Cheng, and Deqiang Zhou. Delay optimal for reliability-guaranteed concurrent transmissions with raptor code in multi-access 6G edge network. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001615.

Yao:2023:OAM

[1592] Pei Yao, Longkun Guo, Peng Li, and Jiawei Lin. Optimal algorithm for minmax line barrier coverage with mobile sensors on 2-dimensional plane. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001627.

Sun:2023:MDA

[1593] Yongxia Sun, Weijin Jiang, Ying Yang, Haoran Zhu, and Yirong Jiang. Multidomain authorization and decisionmaking method of access control in the edge environment. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001664.

Yang:2023:EID

[1594] Run Yang, Hui He, Yixiao Xu, Bangzhou Xin, Yulong Wang, Yue Qu, and Weizhe Zhang. Efficient intrusion detection toward IoT networks using cloud-edge collaboration. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300169X.

Quezada:2023:RTB

[1595] Vicente Quezada, Fabian Astudillo-Salinas, Luis Tello-Oquendo, and Paul Bernal. Real-time bot infection detection system using DNS fingerprinting and machine-learning. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001706.

Jiang:2023:ZRM

[1596] Minghao Jiang, Mingxin Cui, Chang Liu, Gaopeng Gou, Gang Xiong, and Zhen Li. Zero-relabelling mobileapp identification over drifted encrypted network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001731.

Roy:2023:GBR

[1597] Debashri Roy, Batool Salehi, Stella Banou, Subhramoy Mohanti, Guillem Reus-Muns, Mauro Belgiovine, Prashant Ganesh, Chris Dick, and Kaushik Chowdhury. Going beyond RF: a survey on how AI-enabled multimodal beamforming will shape the NextG standard. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001743.

Boudermine:2023:DLB

[1598] Antoine Boudermine, Rida Khatoun, and Jean-Henri Choyer. Dynamic logicbased attack graph for risk assessment in complex computer systems. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001755.

Yang:2023:FSE

[1599] Chen Yang, Gang Xiong, Qing Zhang, Junzheng Shi, Gaopeng Gou, Zhen Li, and Chang Liu. Fewshot encrypted traffic classification via multi-task representation enhanced meta-learning. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001767.

Marinho:2023:CEA

[1600] Ricardo Pagoto Marinho, Luiz F. M. Vieira, Marcos A. M. Vieira, and Antonio A. F. Loureiro. CAIN: an energy-aware and intelligent increasing coverage area routing protocol for future 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001780.

Zhang:2023:DAP

[1601] XiaoXu Zhang, Run hua Shi, WanPeng Guo, PengBo Wang, and WeiYang Ke. A dual auditing protocol for finegrained access control in the edgecloud-based smart home. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001809.

Lin:2023:CCT

[1602] Po-Ching Lin, Wen-Hao Hsu, Ying-Dar Lin, Ren-Hung Hwang, Hsiao-Kuang Wu, Yuan-Cheng Lai, and Chung-Kuan Chen. Correlation of cyber threat intelligence with sightings for intelligence assessment and augmentation. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001810.

Bringhenti:2023:NAS

[1603] Daniele Bringhenti, Riccardo Sisto, and Fulvio Valenza. A novel abstraction for security configuration in virtual networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001901.

Harms:2023:GLC

[1604] Laura Harms, Christian Richter, and Olaf Landsiedel. Grace: Lowcost time-synchronized GPIO tracing for IoT testbeds. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001913.

Herrera:2023:PAC

[1605] Manuel Herrera, Manu Sasidharan, Stephen Cassidy, and Ajith Kumar Parlikad. Performance assessment of a communication infrastructure with redundant topology: a complex network approach. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001925.

Wu:2023:CFU

[1606] Xiaocan Wu, He Huang, Yang Du, Yu-E Sun, and Shigang Chen. Coupon Filter: a universal and lightweight filter framework for more accurate data stream processing. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001937.

Dursun:2023:GNB

[1607] Yunus Dursun, Muhammed Burak Goktas, and Zhiguo Ding. Green NOMA based MU-MIMO transmission for MEC in 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001949.

Liu:2023:TCT

[1608] Yaqun Liu, Jun Xie, Changyou Xing, and Shengxu Xie. Topology construction and topology adjustment in flying ad hoc networks for relay transmission. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001986.

Nguyen:2023:PRA

[1609] Ba Cao Nguyen, Nguyen Thu Phuong, Xuan Nam Tran, and Pham Thanh Hiep. On performance of RIS-aided ground-to-air and air-to-ground communications in multi-user NOMA systems. Computer Networks (Amsterdam, Netherlands: 1999), 228 (??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001998.

Li:2023:TAE

[1610] Pan Li, Guiyan Liu, Songtao Guo, and Yue Zeng. Traffic-aware efficient consistency update in NFV-enabled software defined networking. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002001.

Zhang:2023:SFT

[1611] Qian Zhang, Shangping Wang, Duo Zhang, Jifang Wang, and Yinjuan Deng. Scalable and faulttolerant selection method of verification and accounting nodes for permissionless blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002025.

Garg:2023:DRL

[1612] Sahil Garg, Jia Hu, Giancarlo Fortino, Laurence T. Yang, Mohsen Guizani, Xianjun Deng, and Danda B. Rawat. Deep reinforcement learning for nextgeneration IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002050.

Anonymous:2023:EBi

[1613] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002347.

Anonymous:2023:Jb

[1614] Anonymous. June 2023. Computer Networks (Amsterdam, Netherlands: 1999), 228(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Javanmardi:2023:SPI

[1615] Saeed Javanmardi, Mohammad Shojafar, Reza Mohammadi, Mamoun Alazab, and Antonio M. Caruso. An SDN perspective IoT-Fog security: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001779.

Chkirbene:2023:DIL

[1616] Zina Chkirbene, Ridha Hamila, Aiman Erbad, Serkan Kiranyaz, and Nasser Al-Emadi. D2DLive: Iterative live video streaming algorithm for D2D networks. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001792.

Jiang:2023:WGP

[1617] Jianguo Jiang, Shang Jiang, Yi Liu, Siye Wang, Yanfang Zhang, Yue Feng, and Ziwen Cao. Wi-Gait: Pushing the limits of robust passive personnel identification using Wi-Fi signals. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001962.

Lau:2023:GOP

[1618] Wei Jian Lau, Joanne Mun-Yee Lim, Chun Yong Chong, Nee Shen Ho, and Thomas Wei Min Ooi. General outage probability model for UAV-to-UAV links in multi-UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001974.

Li:2023:LCH

[1619] Junfeng Li, Dan Li, Huiyou Jiang, Du Lin, Jinkun Geng, Yukai Huang, K. K. Ramakrishnan, and Kai Zheng. Light: a compatible, highperformance and scalable user-level network stack. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002013.

Seyoum:2023:DLB

[1620] Yemane Teklay Seyoum, Syed Maaz Shahid, Eun Seon Cho, and Sungoh Kwon. Distributed load balancing algorithm considering QoS for next generation multi-RAT HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002037.

Kress:2023:COS

[1621] Fabian Kreß, Vladimir Sidorenko, Patrick Schmidt, Julian Hoefer, Tim Hotfilter, Iris Walter, Tanja Harbaum, and Jürgen Becker. CN-NParted: an open source framework for efficient Convolutional Neural Network inference partitioning in embedded systems. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002049.

Almstedt:2023:CRA

[1622] Lennart Almstedt, Kai Bleeke, Mohammad Mahhouk, Leander Jehl, Rüdiger Kapitza, and Lars Wolf. ContractBox: Realizing accountable data sharing on the edge using a small scale blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300213X.

Cheng:2023:SDR

[1623] Zhimo Cheng, Xinsheng Ji, Wei You, Yu Zhao, and Zhongfu Guo. SE-CHO: a deep reinforcement learningbased scheme for secure handover in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002141.

Nande:2023:QET

[1624] Swaraj Shekhar Nande, Marius Paul, Stefan Senk, Marian Ulbricht, Riccardo Bassoli, Frank H. P. Fitzek, and Holger Boche. Quantum enhanced time synchronisation for communication network. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002177.

Rivas:2023:RIG

[1625] Angel Esteban Labrador Rivas and Taufik Abrão. Raptor-IRSA grantfree random access protocol for smart grids applications. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002207.

Yang:2023:DJD

[1626] Cheng Yang, Yangming Zhao, Gongming Zhao, and Hongli Xu. DFS: Joint data formatting and sparsification for efficient communication in Distributed Machine Learning. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002220.

Hu:2023:JTR

[1627] Bo Hu, Liangyu Chen, and Shanzhi Chen. Joint trajectory-resource optimization for UAV-enabled uplink communication networks with wireless backhaul. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002244.

Zhou:2023:FSW

[1628] Qiang Zhou, Liangmin Wang, Huijuan Zhu, and Tong Lu. Fewshot website fingerprinting attack with cluster adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002256.

Mohammadi:2023:REE

[1629] Razieh Mohammadi and Zahra Shirmohammadi. RLS^{2}: an energy efficient reinforcement learning-based sleep scheduling for energy harvesting WBANs. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002268.

Lastovicka:2023:POS

[1630] Martin Lastovicka, Martin Husák, Petr Velan, Tomás Jirsík, and Pavel Celeda. Passive operating system fingerprinting revisited: Evaluation and current challenges. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300227X.

Khalil:2023:DLB

[1631] Alvi Ataur Khalil, Mohamed Y. Selim, and Mohammad Ashiqur Rahman. Deep learning-based energy harvesting with intelligent deployment of RIS-assisted UAV-CFmMIMOs. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002293.

Ali:2023:RPB

[1632] Zainab H. Ali, Noha A. Sakr, Nora El-Rashidy, and Hesham A. Ali. A reliable position-based routing scheme for controlling excessive data dissemination in vehicular ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300230X.

Wang:2023:TRM

[1633] Jian Wang, Yuping Yan, and Guosheng Zhao. Task recommendation method combining multimodal cognition and collaboration in mobile crowdsensing systems. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002414.

Anonymous:2023:EBj

[1634] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002645.

Anonymous:2023:Jc

[1635] Anonymous. June 2023. Computer Networks (Amsterdam, Netherlands: 1999), 229(??):??, June 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garroppo:2023:DMI

[1636] Rosario Giuseppe Garroppo and Maria Grazia Scutellà. Design model of an IEEE 802.11ad infrastructure for TSN-based industrial applications. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002165.

Kim:2023:RRM

[1637] Sungwook Kim. Radio resource management scheme in radar and communication spectral coexistence platform. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623002189.

Qin:2023:HCD

[1638] Xingsheng Qin, Frank Jiang, Mingcan Cen, and Robin Doss. Hybrid cyber defense strategies using Honey-X: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002219.

Kowalski:2023:TMR

[1639] Mikołaj Kowalski and Wojciech Mazurczyk. Toward the mutual routing security in wide area networks: a scoping review of current threats and countermeasures. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002232.

Shu:2023:NCM

[1640] Zhaogang Shu, Haoxian Feng, Tarik Taleb, and Zhifang Zhang. A novel combinatorial multi-armed bandit game to identify online the changing top-K flows in softwaredefined networks. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002281.

Kurt:2023:SRA

[1641] Isil Kurt, Fatih Alagöz, and Mete Akgün. Scalable RFID authentication protocol based on physically unclonable functions. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002384.

Li:2023:DEC

[1642] Pengmiao Li, Yuchao Zhang, Huahai Zhang, Wendong Wang, Ke Xu, and Zhili Zhang. A delayed eviction caching replacement strategy with unified standard for edge servers. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002396.

Mazloum:2023:SRT

[1643] Ali Mazloum, Elie Kfoury, Jose Gomez, and Jorge Crichigno. A survey on rerouting techniques with P4 programmable data plane switches. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002402.

Tang:2023:RRB

[1644] Tong Tang, Yingxu Lai, and Yipeng Wang. Relational reasoning-based approach for network protocol reverse engineering. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002426.

Fotiou:2023:SVC

[1645] Nikos Fotiou, Yannis Thomas, Vasilios A. Siris, George Xylomenos, and George C. Polyzos. Selfverifiable content using decentralized identifiers. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300244X.

Kayathri:2023:SDL

[1646] T. L. Kayathri, N. Kumaresan, and R. Vijayabhasker. SDBGPChain: a decentralized low complexity framework to detect and prevent the BGPattacks using SDN with smart contract based Dendrimer tree blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002451.

Anonymous:2023:ISIb

[1647] Anonymous. Introduction to special issue on selected papers from IFIP networking conference 2022. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002517.

Cerquitelli:2023:MLE

[1648] Tania Cerquitelli, Michela Meo, Marilia Curado, Lea Skorin-Kapov, and Eirini Eleni Tsiropoulou. Machine learning empowered computer networks. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002529.

Anonymous:2023:EBk

[1649] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002773.

Anonymous:2023:Jd

[1650] Anonymous. July 2023. Computer Networks (Amsterdam, Netherlands: 1999), 230(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chang:2023:TNA

[1651] Hyunseok Chang, Fang Hao, Murali Kodialam, T. V. Lakshman, Sarit Mukherjee, and Matteo Varvello. Towards network-assisted publish– subscribe over wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623001470.

Alnajar:2023:TIF

[1652] Omar Alnajar and Ahmed Barnawi. Tactile internet of federated things: Toward fine-grained design of FL-based architecture to meet TIoT demands. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623001573.

Milanez:2023:VVB

[1653] Guilherme A. S. Milanez, Marcos A. M. Vieira, Luiz F. M. Vieira, and José Augusto M. Nacif. VariBan: a variable bandwidth channel allocation algorithm for IEEE 802.15.4ebased networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002190.

Bozkaya:2023:DTA

[1654] Elif Bozkaya. Digital twin-assisted and mobility-aware service migration in Mobile Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002438.

Yang:2023:MME

[1655] Huan Yang, Sheng Sun, Min Liu, Qiuping Zhang, and Yuwei Wang. MJOA-MU: End-to-edge collaborative computation for DNN inference based on model uploading. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002463.

Sahin:2023:CCP

[1656] Mehmet Emin Sahin and Mehmet Demirci. ConPoolUBF: Connection pooling and updatable Bloom filter based SYN flood defense in programmable data planes. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002475.

Moorthy:2023:SUN

[1657] Sabarish Krishna Moorthy, Nicholas Mastronarde, Scott Pudlewski, Elizabeth Serena Bentley, and Zhangyu Guan. Swarm UAV networking with collaborative beamforming and automated ESN learning in the presence of unknown blockages. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002499.

Zhao:2023:AMP

[1658] Zhe Zhao, Lu Gao, Weitao Pan, Zhiliang Qiu, Xu Guo, Ya Gao, and Ling Zheng. Access mechanism for period flows of non-deterministic end systems for time-sensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002505.

Baena:2023:MEK

[1659] Carlos Baena, O. S. Peñaherrera-Pulla, Raquel Barco, and Sergio Fortes. Measuring and estimating Key Quality Indicators in Cloud Gaming services. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002530.

Alatram:2023:DMI

[1660] Alaa Alatram, Leslie F. Sikos, Mike Johnstone, Patryk Szewczyk, and James Jin Kang. DoS/DDoS-MQTT-IoT: a dataset for evaluating intrusions in IoT networks using the MQTT protocol. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002542.

Wang:2023:IFB

[1661] Yang Wang, Yawen Chen, Zhaoming Lu, and Xiangming Wen. Intelligent flying-beamformer for hybrid mmWave systems: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002554.

Attarian:2023:EWF

[1662] Reyhane Attarian and Alireza Keshavarz-Haddad. Effective website fingerprinting attack based on the first packet direction only. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002566.

Ma:2023:UAC

[1663] Bo Ma, Heng Kuang, Shiyuan Liu, and Chuanhuang Li. UAV assisted cellular network traffic offloading: Joint swarm, 3D deployment, and user allocation optimization based on a data-aware method. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002578.

Leon:2023:MLB

[1664] Juan Pablo Astudillo León, Luis J.

de la Cruz Llopis, and Francisco J. Rico-Novella. A machine learning based Distributed Congestion Control Protocol for multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300258X.

Troia:2023:PCP

[1665] Sebastian Troia, Marco Savi, Giulia Nava, Ligia Maria Moreira Zorello, Thomas Schneider, and Guido Maier. Performance characterization and profiling of chained CPU-bound Virtual Network Functions. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002608.

Yin:2023:JDR

[1666] Yabo Yin, Chuanhe Huang, Neal N. Xiong, Dong-Fang Wu, and Shidong Huang. Joint dynamic routing and resource allocation in satelliteterrestrial integrated networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002682.

Almasaeid:2023:MCS

[1667] Hisham M. Almasaeid. Minimum cost spectrum allocation with QoS guarantees in multi-interface multi-hop dynamic spectrum access networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002694.

Wu:2023:PCP

[1668] Hua Wu, Ziling Shao, Fuhao Yang, Guang Cheng, Xiaoyan Hu, Jing Ren, and Wei Wang. PD-CPS: a practical scheme for detecting covert port scans in high-speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002700.

Zhu:2023:RFD

[1669] Mengfei Zhu and Eiji Oki. Robust function deployment against uncertain recovery time in different protection types with workloaddependent failure probability. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002712.

Birabwa:2023:MAD

[1670] Denise Joanitah Birabwa, Daniel Ramotsoela, and Neco Ventura. Multiagent deep reinforcement learning for user association and resource allocation in integrated terrestrial and non-terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002724.

Li:2023:MAA

[1671] Peng Li, Junzuo Lai, Dehua Zhou, Ye Yang, Wei Wu, and Junbin Fang. Multi-authority anonymous authentication with public accountability for incentive-based applications. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002736.

Hu:2023:MTC

[1672] Chia-Cheng Hu, Wen-Wu Liu, and Jeng-Shyang Pan. Minimizing traffic cost of content distribution and storage allocation in cloud radio access networks. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002815.

Anonymous:2023:EBl

[1673] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003043.

Anonymous:2023:Je

[1674] Anonymous. July 2023. Computer Networks (Amsterdam, Netherlands: 1999), 231(??):??, July 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ali:2023:SUV

[1675] Mohammad Furqan Ali and Dushantha Nalin K. Jayakody. SIMO-Underwater Visible Light Communication (UVLC) system. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623001950. See

corrigendum [1759].

Zeba:2023:MPH

[1676] Ousmane Zeba, Shunsuke Saruwatari, and Takashi Watanabe. MAC protocols for high capacity and increased reliability in bistatic backscatter networks for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002153.

Mu:2023:DRL

[1677] Liting Mu, Bin Ge, Chenxing Xia, and Cai Wu. Deep reinforcement learning based adaptive threshold multi-tasks offloading approach in MEC. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002487.

Jung:2023:OFS

[1678] Sehun Jung and Hyang-Won Lee.

Optimization framework for splitting DNN inference jobs over computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002591.

Dong:2023:EEE

[1679] Yongpeng Dong, Shiyou Qian, Wanghua Shi, Jian Cao, and Guangtao Xue. EEM: an elastic event matching framework for content-based publish/subscribe systems. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002827.

Wang:2023:ECR

[1680] You-Chiun Wang and Chia-Wei Chou. Efficient coordination of radio frames to mitigate cross-link interference in 5G D-TDD systems. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002852.

Biswas:2023:MRP

[1681] Kamanashis Biswas, Vallipuram Muthukkumarasamy, Mohammad Jabed Morshed Chowdhury, Xin-Wen Wu, and Kalvinder Singh. A multipath routing protocol for secure energy efficient communication in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002876.

Wang:2023:CCB

[1682] Zhujun Wang, Lejun Zhang, Ran Guo, Guopeng Wang, Jing Qiu, Shen Su, Yuan Liu, Guangxia Xu, and Zhihong Tian. A covert channel over blockchain based on label tree without long waiting times. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002888.

Chen:2023:ASM

[1683] Jie Chen, Jiaxu Niu, Hao Lei, Li Lin, and Yunhao Ling. Adaptively secure multi-authority attribute-based broadcast encryption in fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 232 (??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300289X.

Dung:2023:EED

[1684] Le The Dung and Taejoon Kim. Exploring the effects of directional antennas on the secure connectivity and hop count of secure multi-hop adhoc wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002906.

Dao:2023:ONI

[1685] Thi-Nga Dao, Duc Van Le, and Xuan Nam Tran. Optimal network intrusion detection assignment in multi-level IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002918.

Guda:2023:AAO

[1686] Sruthi Guda and Sreenivasa Rao Duggirala. Adaptive Aquila optimized relay selection and power allocation in CR-NOMA model. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002931.

Valente:2023:OBC

[1687] Francesco Valente, Vincenzo Eramo, and Francesco G. Lavacca. Optimal bandwidth and computing resource allocation in low earth orbit satellite constellation for earth observation applications. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002943.

Yigit:2023:NFT

[1688] Beytüllah Yigit, Gürkan Gür, Fatih Alagöz, and Bernhard Tellenbach. Network fingerprinting via timing attacks and defense in software defined networks. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002955.

Zhao:2023:FSL

[1689] Guozheng Zhao, Kaiqiang Lin, and Tong Hao. A feasibility study of LoRaWAN-based wireless underground sensor networks for underground monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 232 (??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002967.

Duong:2023:CMA

[1690] Tho Minh Duong, Seung-Ho Hyun, and Sungoh Kwon. Channel modeling and achievable data rate for 1-D molecular communication in bounded environments. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002979.

Abubakar:2023:CTA

[1691] Attai Ibrahim Abubakar, Michael S.

Mollel, Oluwakayode Onireti, Metin Ozturk, Iftikhar Ahmad, Syed Muhammad Asad, Yusuf Sambo, Ahmed Zoha, Sajjad Hussain, and Muhammad Ali Imran. Coverage and throughput analysis of an energy efficient UAV base station positioning scheme. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002992.

Garcia-Martin:2023:ECA

[1692] Juan Pablo García-Martín and Antonio Torralba. Energy consumption analytical modeling of NB-IoT devices for diverse IoT applications. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003006.

Anonymous:2023:EBm

[1693] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003316.

Anonymous:2023:Ac

[1694] Anonymous. August 2023. Computer Networks (Amsterdam, Netherlands: 1999), 232(??):??, August 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bagnulo:2023:DIV

[1695] Marcelo Bagnulo, Alberto García-Martínez, Anna Maria Mandalari, Praveen Balasubramanian, Daniel Havey, and Gabriel Montenegro. Design, implementation and validation of a receiver-driven less-than-best-effort transport. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623002864.

Yao:2023:UFS

[1696] Mingwu Yao, Jiamu Liu, Jing Du, Dongqi Yan, Yanxi Zhang, Wei Liu, and Anthony Man-Cho So. A unified flow scheduling method for time sensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300292X.

Lekssays:2023:MBB

[1697] Ahmed Lekssays, Barbara Carminati, and Elena Ferrari. Mal-Con: a blockchain-based malware containment framework for Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623002980.

Ren:2023:CAD

[1698] Xiangyu Ren, Lin Cai, Pu Yang, and Jiequ Ji. Congestion-aware delayguaranteed scheduling and routing with renewal optimization. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003080.

Zhang:2023:CRT

[1699] Menghan Zhang, Xianliang Jiang, Guang Jin, Penghui Li, and Haiming Chen. CapRadar: Real-time adaptive bandwidth prediction for dynamic wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003109.

Li:2023:PEC

[1700] Chen Li, Xiaoyu Wang, Tongyu Zong, Houwei Cao, and Yong Liu. Predictive edge caching through deep mining of sequential patterns in user content retrievals. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003110.

Zhou:2023:EPL

[1701] Yuwen Zhou, Bangbang Ren, Junjie Xie, Lailong Luo, Deke Guo, and Xiaobo Zhou. Enable the proactively load-balanced control plane for SDN via intelligent switch-to-controller selection strategy. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003122.

Ruiz-Villafranca:2023:MII

[1702] Sergio Ruiz-Villafranca, José Roldán-Gómez, Javier Carrillo-Mondéjar, Juan Manuel Castelo Gómez, and José Miguel Villalón. A MEC-IIoT intelligent threat detector based on machine learning boosted tree algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003134.

Zhang:2023:EES

[1703] Shengyu Zhang and Kwan L. Yeung.

Efficient embedding of service function chains in space-division multiplexing elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003146.

Alonso:2023:WSP

[1704] Rodney Martinez Alonso, David Plets, Luc Martens, Wout Joseph, Ernesto Fontes Pupo, and Glauco Guillen Nieto. White spaces pattern finding and inference based on machine learning for multi-frequency spectrum footprints. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300316X.

Damsgaard:2023:ACB

[1705] Hans Jakob Damsgaard, Aleksandr Ometov, Md Munjure Mowla, Adam Flizikowski, and Jari Nurmi. Approximate computing in B5G and 6G wireless systems: a survey and future outlook. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003171.

Lyu:2023:PPE

[1706] Minzhao Lyu, Hassan Habibi Gharakheili, and Vijay Sivaraman. PEDDA: Practical and effective detection of distributed attacks on enterprise networks via progressive multi-stage inference. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003183.

Naqvi:2023:IID

[1707] Haidlir Achmad Naqvi, Muhammad Hafizhuddin Hilman, and Bayu Anggorojati. Implementability improvement of deep reinforcement learning based congestion control in cellular network. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003195.

Satoh:2023:AFT

[1708] Daisuke Satoh. Analysis for flow termination in the decentralized and centralized pre-congestion notification architecture. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003213.

Oliveira:2023:TGM

[1709] Afonso Oliveira and Teresa Vazão. Towards green machine learning for resource allocation in beyond 5G RAN slicing. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003225.

Li:2023:DAW

[1710] Yinlong Li, Siyao Cheng, Hao Zhang, and Jie Liu. Dynamic adaptive workload offloading strategy in mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003237.

Li:2023:PBM

[1711] Xinhao Li, Yingjie Wang, Bingyi Xie, Lingkang Meng, Zhaowei Liu, Xiangrong Tong, Ao Zhou, and Zhipeng Cai. Planning-based mobile crowdsourcing bidirectional multi-stage online task assignment. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003249.

Wong:2023:FGH

[1712] KaKei Wong and Lin Cui. Finegrained HTTP/3 prioritization via reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003250.

Fang:2023:RRA

[1713] Jin Fang, Gongming Zhao, Hongli Xu, Huaqing Tu, and Haibo Wang. Reveal: Robustness-aware VNF placement and request scheduling in edge clouds. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003274.

Dogan:2023:SSO

[1714] Orkun Dogan and Kübra Kalkan. SOREC: Self-organizing and resource efficient clustered blockchain network. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003389.

Li:2023:CSD

[1715] Qing Li, He Huang, Ruoyu Li, Jianhui Lv, Zhenhui Yuan, Lianbo Ma, Yi Han, and Yong Jiang. A comprehensive survey on DDoS defense systems: New trends and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003407.

N:2023:SSC

[1716] Pruthvi C. N., H. S. Vimala, and Shreyas J. A systematic survey on content caching in ICN and ICN-IoT: Challenges, approaches and strategies. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003419.

Megias:2023:EBT

[1717] Carlos Megías, Víctor Vázquez, Eduardo Ros, Mauro Cappelli, and Javier Díaz. Ethernet-based timing system for accelerator facilities: the IFMIF-DONES case. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003420.

Liu:2023:SSS

[1718] Jing Liu, Jianguo Xie, Biao Liu, Liwei Xu, Xiaoming Chen, and Huamin Feng. SED-SGC: a scalable, efficient, and distributed secure group communication scheme based on superlattice PUF. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003432.

Amador:2023:SIP

[1719] Oscar Amador, Ignacio Soto, Maria Calderon, and Manuel Urueña. Studying and improving the performance of ETSI ITS contention-based forwarding (CBF) in urban and highway scenarios: S-FoT+. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003444.

Jiang:2023:PPS

[1720] Wanchun Jiang, Renfu Yao, Kaiqin Liao, Yulong Yan, Jiawei Huang, Weiping Wang, and Jianxin Wang. Practical periodic strategy for 40/100 Gbps energy efficient Ethernet. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003468.

Sedrati:2023:IGS

[1721] Anass Sedrati, Abdellatif Mezrioui, and Aafaf Ouaddah. IoT-Gov: a structured framework for Internet of Things governance. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300347X.

Wang:2023:TUS

[1722] Jian Wang, Jia Liu, and Guosheng Zhao. Trusted user selection for fusion of multimodal cognition in selforganizing pattern of mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003481.

Liu:2023:DFF

[1723] Zunyi Liu, Dian Shen, Jiaang Bao, Fang Dong, and Jiong You. Deep-MetricCorr: Fast flow correlation for data center networks with deep metric learning. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003493.

Parra-Espin:2023:SBA

[1724] José Antonio Parra-Espín, Rafael Marín-López, Gabriel López-Millán, Fernando Pereñíguez-García, and Oscar Canovas. SDN-based automated rekey of IPsec security associations: Design and practical validations. Computer Networks (Amsterdam, Netherlands: 1999), 233 (??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300350X.

Anonymous:2023:EBn

[1725] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623003742.

Anonymous:2023:S

[1726] Anonymous. September 2023. Computer Networks (Amsterdam, Netherlands: 1999), 233(??):??, September 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Chen:2023:MMC

[1727] Luyao Chen, Yong Tang, Jingwen Xia, Siyang Chen, Chengyu Zheng, Hai Lin, and Wenyong Wang. Multi-MEC collaboration for VR video transmission: Architecture and cache algorithm design. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003092.

Sharara:2023:MEC

[1728] Mahdi Sharara, Francesca Fossati, Sahar Hoteit, Véronique Vèque, and Francesca Bassi. Minimizing energy consumption by joint radio and computing resource allocation in Cloud-RAN. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003158.

Talau:2023:ITP

[1729] Marcos Talau, Thiago A. Herek, Mauro Fonseca, and Emilio C. G. Wille. Improving TCP performance over a common IoT scenario using the Early Window Tailoring method. Computer Networks (Amsterdam, Netherlands:

1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003201.

Minmin:2023:TEC

[1730] Huang Minmin, Yuan Lingyun, Pan Xue, and Zhou Chuan. Trusted edge and cross-domain privacy enhancement model under multi-blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003262.

Cao:2023:RLB

[1731] Shaohua Cao, Di Liu, Congcong Dai, Chengqi Wang, Yansheng Yang, Weishan Zhang, and Danyang Zheng. Reinforcement learning based tasks offloading in vehicular edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003390.

Xu:2023:LAD

[1732] Chenglin Xu, Cheng Xu, Bo Li, Siqi Li, and Tao Li. Load-aware dynamic controller placement based on deep reinforcement learning in SDNenabled mobile cloud-edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003456.

Yu:2023:CJA

[1733] Kan Yu, Jiguo Yu, Zhiyong Feng, and Min Guo. Cooperative jamming aided securing wireless communications without CSI of eavesdroppers. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003511.

Wang:2023:MCE

[1734] Yang Wang, Qian Hu, Long Nguyen, and Maryam Jalalitabar. Minimumcost embedding of virtual networks: an iterative decomposition approach. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003523.

Park:2023:TTC

[1735] Kibeom Park, Sangmo Sung, Hokeun Kim, and Jae il Jung. Technology trends and challenges in SDN and service assurance for end-to-end network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003535.

Nguyen:2023:OAC

[1736] Tien-Tung Nguyen and Anh-Vinh Nguyen. Outage analysis of cognitive inspired NOMA networks in the presence of imperfect SIC, CCI, and non-ID fading channels. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003547.

Moure-Garrido:2023:RTD

[1737] Marta Moure-Garrido, Celeste Campo, and Carlos Garcia-Rubio. Real time detection of malicious DoH traffic using statistical analysis. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003559.

TedT:2023:HDI

[1738] Ted “Taekyoung” Kwon, Junghwan Song, Heeyoung Jung, Selin Chun, Hyunwoo Lee, Minhyeok Kang, Minkyung Park, and Eunsang Cho. How to decentralize the Internet: a focus on data consolidation and user privacy. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003560.

Rezaei:2023:ESE

[1739] Zahra Rezaei and Behrouz Shahgholi Ghahfarokhi. Energy and spectrum efficient cell switch-off with channel and power allocation in ultradense networks: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003572.

Antevski:2023:ABC

[1740] Kiril Antevski and Carlos J. Bernardos. Applying blockchain consensus mechanisms to network service federation: Analysis and performance evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003584.

Liu:2023:SDN

[1741] Jiayi Liu, Xiangjie Shi, Yizhi Huang, and Qinghai Yang. Selective and ondemand network measurement with SRv6 and INT. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003596.

Qu:2023:ETO

[1742] Xiaofeng Qu and Huiqiang Wang. Emergency task offloading strategy based on cloud-edge-end collaboration for smart factories. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003602.

K:2023:KSI

[1743] Venkatakrishna Rao. K. and C. Ravindranath Chowdary. K++ Shell: Influence maximization in multilayer networks using community detection. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003614.

Fu:2023:PPM

[1744] Yanming Fu, Bocheng Huang, Xiao Liu, Jiayuan Chen, and Shenglin Lu. Privacy-preserving mobile crowd sensing task assignment with Stackelberg game. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003626.

Liu:2023:ERT

[1745] Libin Liu, Jingzong Li, Hong Xu, Kaiwen Xue, and Jason Chun Xue. Efficient real-time video conferencing with adaptive frame delivery. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003638.

Zeng:2023:JOM

[1746] Yaoping Zeng, Dongyang Lu, and Jianbo Du. Joint optimized multiuser access and UAV deployments based on heterogeneous revenue in IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300364X.

Allegretta:2023:CSC

[1747] Mauro Allegretta, Giuseppe Siracusano, Roberto Gonzalez, and Marco Gramaglia. Are crowd-sourced CTI datasets ready for supporting anticybercrime intelligence? Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003651.

Fernandez-Veiga:2023:IRS

[1748] Manuel Fernández-Veiga, M. E. Sousa-Vieira, Ana Fernández-Vilas, and Rebeca P. Díaz-Redondo. Irregular repetition slotted Aloha with multiuser detection: a density evolution analysis. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003663.

Chen:2023:OAM

[1749] Tingjun Chen, Prasanthi Maddala, Panagiotis Skrimponis, Jakub Kolodziejski, Abhishek Adhikari, Hang Hu, Zhihui Gao, Arun Paidimarri, Alberto Valdes-Garcia, Myung Lee, Sundeep Rangan, Gil Zussman, and Ivan Seskar. Open-access millimeter-wave software-defined radios in the PAWR COSMOS testbed: Design, deployment, and experimentation. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003675.

Kar:2023:DPD

[1750] Snigdhaswin Kar, Prabodh Mishra, and Kuang-Ching Wang. Dynamic packet duplication for reliable low latency communication under mobility in 5G NR-DC networks. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003687.

Manzoni:2023:EPS

[1751] Pietro Manzoni, Claudio Palazzi, Flavia Delicato, Erika Rosas, and Spyridon Mastorakis. Editorial: Pub/ sub solutions for interoperable and dynamic IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003705.

Ergun:2023:SHN

[1752] Serap Ergun, Ibrahim Sammour, and Gerard Chalhoub. A survey on how network simulators serve reinforcement learning in wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003791.

Phu:2023:DSA

[1753] Anh Tuan Phu, Bo Li, Faheem Ullah, Tanvir Ul Huque, Ranesh Naha, Muhammad Ali Babar, and Hung Nguyen. Defending SDN against packet injection attacks using deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003808.

Luo:2023:IAD

[1754] Feng Luo, Zitong Wang, and Baoyin Zhang. Impact analysis and detection of time-delay attacks in time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300381X.

Boutiba:2023:ORR

[1755] Karim Boutiba, Miloud Bagaa, and Adlen Ksentini. Optimal radio resource management in 5G NR featuring network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003821.

Onsu:2023:HCM

[1756] Murat Arda Onsu, Burak Kantarci, and Azzedine Boukerche. How to cope with malicious federated learning clients: an unsupervised learningbased approach. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003833.

Zhu:2023:DEE

[1757] Rongxin Zhu, Azzedine Boukerche, Xiangdang Huang, and Qiuling Yang. DESLR: Energy-efficient and secure layered routing based on channel-aware trust model for UASNs. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003845.

Li:2023:EET

[1758] Bin Li, Wenshuai Liu, Wancheng Xie, and Xiaohui Li. Energy-efficient task offloading and trajectory planning in UAV-enabled mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003857.

Ali:2023:CSU

[1759] Mohammad Furqan Ali and Dushantha Nalin K. Jayakody. Corrigendum to “SIMO-Underwater Visible Light Communication (UVLC) system” Computer Networks. 2023 May 9:109750. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003870. See

[1675].

Shen:2023:JTO

[1760] Bo Shen, Qi Gu, and Gang Yang. Joint task offloading and UAVs deployment for UAV-assisted mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003882.

Su:2023:IDL

[1761] Lina Su and Zongpeng Li. Incentivedriven long-term optimization for hierarchical federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003894.

Pei:2023:RNP

[1762] Jiaming Pei, Kaiyang Zhong, Mian Ahmad Jan, and Jinhai Li. Retraction notice to ‘Personalized Federated Learning Framework for Network Traffic Anomaly Detection’. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003900. See

[1067].

Qu:2023:HNS

[1763] Lianwei Qu, Yong Wang, Jing Yang, and Meng Zhao. A heterogeneous network structure publishing security framework based on cloud-edge collaboration. Computer Networks (Amsterdam, Netherlands: 1999), 234 (??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003924.

Zhou:2023:CSM

[1764] Boqing Zhou, Sujun Li, Jianxin Wang, Yun Cheng, and Jie Wu. Corrigendum to “A secure model against mobile sink replication attacks in unattended sensor networks” COMPNW, Volume 221, February 2023, 109529. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004000. See

[1437].

Anonymous:2023:EBo

[1765] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300419X.

Anonymous:2023:O

[1766] Anonymous. October 2023. Computer Networks (Amsterdam, Netherlands: 1999), 234(??):??, October 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cao:2023:NAI

[1767] Chenhong Cao, Miaoling Dai, Bonan Shen, Guobing Zou, and Wei Dong. Neural adaptive IoT streaming analytics with RL-Adapt. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003699.

Souza:2023:RDS

[1768] Otavio A. de O. Souza, Caio A. Caldeira, and Olga Goussevskaia. Randomized distributed self-adjusting tree networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003869.

Ma:2023:GGB

[1769] Zhaorui Ma, Shicheng Zhang, Na Li, Tianao Li, Xinhao Hu, Hao Feng, Qinglei Zhou, Fenlin Liu, Xiaowen Quan, Hongjian Wang, Guangwu Hu, Shubo Zhang, Yaqi Zhai, Shuaibin Chen, and Shuaiwei Zhang. GraphNEI: a GNNbased network entity identification method for IP geolocation. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003912.

Kumar:2023:OTA

[1770] G. V. Pradeep Kumar, V. V. Satyanarayana Tallapragada, and N. Alivelu Manga. Optimized transmit antenna selection and self-attention based convolutional resource allocation model for massive MIMO technology. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003936.

Araujo:2023:CCS

[1771] Guilherme Araujo, Maycon Peixoto, and Leobino Sampaio. A comprehensive and configurable simulation environment for supporting vehicular named-data networking applications. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003948.

Perifanis:2023:FLB

[1772] Vasileios Perifanis, Nikolaos Pavlidis, Remous-Aris Koutsiamanis, and Pavlos S. Efraimidis. Federated learning for 5G base station traffic forecasting. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300395X.

Campazas-Vega:2023:AIS

[1773] Adrián Campazas-Vega, Ignacio Samuel Crespo-Martínez, Ángel Manuel Guerrero-Higueras, Claudia Álvarez-Aparicio, Vicente Matellán, and Camino Fernández-Llamas. Analyzing the influence of the sampling rate in the detection of malicious traffic on flow data. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003961.

Wang:2023:MSD

[1774] Kaiyu Wang, Zhiying Tu, Zhenzhou Ji, and Shufan He. Multistage data synchronization for public blockchain in complex network environment. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623003973.

Alkwai:2023:TCM

[1775] Lamia Alkwai, Abdelfettah Belghith, Achraf Gazdar, and Saad AlAhmadi. Transparent consumer mobility management in named data networking under the push communication mode. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003985.

Chen:2023:CRS

[1776] Junmei Chen, Zongpeng Li, Guang Fang, Yeqiao Hou, and Xianglong Li. A comprehensive repair scheme for distributed storage systems. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623003997.

Leon:2023:EPL

[1777] Juan Pablo Astudillo León, Carlos Lester Duenas Santos, Ahmad Mohamad Mezher, Julián Cárdenas Barrera, Julian Meng, and Eduardo Castillo Guerra. Exploring the potential, limitations, and future directions of wireless technologies in smart grid networks: a comparative analysis. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004012.

Nguyen:2023:TWU

[1778] Tien-Tung Nguyen, Anh-Vinh Nguyen, Tan-Loc Nguyen, and Van-Hung Le. Two-way UAV-aided systems with short packet communications. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004024.

Xu:2023:IBQ

[1779] Yaqi Xu, Yan Shi, Yuming Ge, Shanzhi Chen, and Longxiang Wang. Informerbased QoS prediction for V2X communication: a method with verification using reality field test data. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004036.

Garcia-Luna-Aceves:2023:ANC

[1780] J. J. Garcia-Luna-Aceves and Dylan Cirimelli-Low. ALOHA-NUI: a collision-free version of ALOHA using a neighborhood-understood index. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004048.

Mukhtaruzzaman:2023:SDF

[1781] Mohammad Mukhtaruzzaman and Mohammed Atiquzzaman. Stable dynamic feedback-based predictive clustering protocol for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300405X.

Belda:2023:DSS

[1782] Román Belda, Pau Arce, Juan Carlos Guerri, and Ismael de Fez. A DASH server-side delay-based representation switching solution to improve the quality of experience for lowlatency live video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004061.

Santana:2023:ALR

[1783] Juan Ramón Santana, Pablo Sotres, Jesús Pérez, Luis Sánchez, Jorge Lanza, and Luis Muñoz. Assessing LoRaWAN radio propagation for smart parking service: an experimental study. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004073.

Bine:2023:NAC

[1784] Lailla M. S. Bine, Azzedine Boukerche, Linnyer B. Ruiz, and Antonio A. F. Loureiro. A novel ant colony-inspired coverage path planning for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004085.

Abbasi:2023:FTS

[1785] Felor Beikzadeh Abbasi, Ali Rezaee, Sahar Adabi, and Ali Movaghar. Fault-tolerant scheduling of graphbased loads on fog/cloud environments with multi-level queues and LSTMbased workload prediction. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004097.

Xu:2023:FLM

[1786] Yuwei Xu, Jie Cao, Kehui Song, Qiao Xiang, and Guang Cheng. FastTraffic: a lightweight method for encrypted traffic fast classification. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004103. See

corrigendum [1893].

Rafique:2023:SFC

[1787] Wajid Rafique, Abdelhakim Senhaji Hafid, and Soumaya Cherkaoui. Soft-Caching: a framework for caching node selection and routing in Software-Defined Information Centric Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004115.

Kumar:2023:CSH

[1788] E. Praveen Kumar and S. Priyanka.

A comprehensive survey on hardwareassisted malware analysis and primitive techniques. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004127.

Naderi:2023:HLE

[1789] Mohammad Naderi, Jacob Chakareski, and Mohammad Ghanbari. Hierarchical Q-learning-enabled neutrosophic AHP scheme in candidate relay set size adaption in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004139.

Shao:2023:BAC

[1790] Huishuang Shao, Yurong Xia, and Changhao Piao. Blockchain-assisted certificateless signcryption for vehicleto-vehicle communication in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623004140.

Batalla:2023:SAR

[1791] Jordi Mongay Batalla, Sławomir Sujecki, Jan M. Kelner, Piotr Śliwka, and Dariusz Zmysłowski. On studying active radio measurements estimating the mobile network quality of service for the Regulatory Authority’s purposes. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004255.

Jin:2023:MAN

[1792] Zilong Jin, Jun Qian, Zhixiang Kong, and Chengsheng Pan. A mobility aware network traffic prediction model based on dynamic graph attention spatio-temporal network. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004267.

Wang:2023:RTI

[1793] Shiyu Wang, Wenxiang Xu, and Yiwen Liu. Res-TranBiLSTM: an intelligent approach for intrusion detection in the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004279.

Min:2023:RLB

[1794] Junhong Min, Yongjun Kim, Moonbeom Kim, Jeongyeup Paek, and Ramesh Govindan. Reinforcement learning based routing for time-aware shaper scheduling in time-sensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004280.

He:2023:RLM

[1795] Rui He, Bangbang Ren, Junjie Xie, Deke Guo, Yuwen Zhou, Laiping Zhao, and Yong Li. A reinforcement learning method for scheduling service function chains with multi-resource constraints. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004309.

Liu:2023:MDG

[1796] Lingling Liu, Aimin Wang, Geng Sun, and Jiahui Li. Maximizing data gathering and energy efficiency in UAV-assisted IoT: a multi-objective optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004310.

Liu:2023:RLB

[1797] Ning Liu, Chunbo Jia, Bingnan Hou, Changsheng Hou, Yingwen Chen, and Zhiping Cai. 6Search: a reinforcement learning-based traceroute approach for efficient IPv6 topology discovery. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004322.

Dong:2023:BBD

[1798] Xiaodong Dong and Binlei Cai. Balanar: Balancing deadline guarantee and Jain’s fairness for inter-datacenter transfers. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004334.

Wang:2023:PDR

[1799] Ting Wang, Xi Fan, Kai Cheng, Xiao Du, Haibin Cai, and Yang Wang. Parameterized deep reinforcement learning with hybrid action space for energy efficient data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004346.

Wang:2023:CED

[1800] Han Wang, Zhangguo Tang, Huanzhou Li, Jian Zhang, Shuangcheng Li, and Junfeng Wang. CI GRU: an efficient DGA botnet classification model based on an attention recurrence plot. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004371.

Erazo-Agredo:2023:ERE

[1801] Cristian C. Erazo-Agredo, Luis Diez, Ramón Agüero, Mario Garza-Fabre, and Javier Rubio-Loyola. Enabling realistic experimentation of disaggregated RAN: Design, implementation, and validation of a modular multisplit eNodeB. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004383.

Ullah:2023:SBE

[1802] Syed Sajid Ullah, Vladimir Oleshchuk, and Harsha S. Gardiyawasam Pussewalage. A survey on blockchain envisioned attribute based access control for Internet of Things: Overview, comparative analysis, and open research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004395.

Lin:2023:NOD

[1803] Xu Lin, Chuchu Liu, Lailong Luo, Deke Guo, and Ming Xu. NEST: Optimal deploying DAG-SFCs to maximize the flows wholly served in the network edge. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004401.

Hidouri:2023:QIL

[1804] Abdelhak Hidouri, Haifa Touati, Mohamed Hadded, Nasreddine Hajlaoui, Paul Muhlethaler, and Samia Bouzefrane. Q-ICAN: a Q-learning based cache pollution attack mitigation approach for named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004437.

Zhu:2023:DRL

[1805] Xiaojuan Zhu, Tianhao Zhang, Jinwei Zhang, Bao Zhao, Shunxiang Zhang, and Cai Wu. Deep reinforcement learning-based edge computing offloading algorithm for softwaredefined IoT. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004516.

Wang:2023:DHO

[1806] Xiang Wang, Yang Liu, Jiahui Ma, Yunling Dong, Jiaxing Chen, and Zhihua Liu. A DV-Hop optimization localization algorithm based on topological structure similarity in three-dimensional wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 235 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004589.

Anonymous:2023:EBp

[1807] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623004711.

Anonymous:2023:Na

[1808] Anonymous. November 2023. Computer Networks (Amsterdam, Netherlands: 1999), 235(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2023:STO

[1809] Jing Xu, Jiarun Tang, Yuze Zou, Ruikai Wen, Wei Liu, and Jianhua He. Sum throughput optimization of wireless powered IRS-assisted multi-user MISO system. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004292.

Lv:2023:ADO

[1810] Sicai Lv, Chao Wang, Zibo Wang, Shuo Wang, Bailing Wang, and Yongzheng Zhang. AAE-DSVDD: a one-class classification model for VPN traffic identification. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004358.

Dahanayaka:2023:ROS

[1811] Thilini Dahanayaka, Yasod Ginige, Yi Huang, Guillaume Jourjon, and Suranga Seneviratne. Robust openset classification for encrypted traffic fingerprinting. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300436X.

Zafar:2023:DRB

[1812] Waseeq Ul Islam Zafar, Muhammad Atif Ur Rehman, Farhana Jabeen, Rehmat Ullah, Ghulam Abbas, and Abid Khan. Decentralized receiver-based link stability-aware forwarding scheme for NDN-based VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004413.

Khah:2023:DML

[1813] Sahar Ahmadi Khah, Ali Barati, and Hamid Barati. A dynamic and multi-level key management method in wireless sensor networks (WSNs). Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004425.

Cotter:2023:CAP

[1814] Jamie Cotter, Ignacio Castiñeiras, Donna O’Shea, and Victor Cionca. A comparative analysis of proactive and reactive methods for privacy-aware interleaved DNN offloading. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004449.

McManus:2023:DTE

[1815] Maxwell McManus, Yuqing Cui, Josh (Zhaoxi) Zhang, Jiangqi Hu, Sabarish Krishna Moorthy, Nicholas Mastronarde, Elizabeth Serena Bentley, Michael Medley, and Zhangyu Guan. Digital twin-enabled domain adaptation for zero-touch UAV networks: Survey and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623004450.

Camargo:2023:DSR

[1816] Juan Sebastian Camargo, Estefanía Coronado, Wilson Ramirez, Daniel Camps, Sergi Sánchez Deutsch, Jordi Pérez-Romero, Angelos Antonopoulos, Oscar Trullols-Cruces, Sergio Gonzalez-Diaz, Borja Otura, and Giovanni Rigazzi. Dynamic slicing reconfiguration for virtualized 5G networks using ML forecasting of computing capacity. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004462.

Hirayama:2023:SBV

[1817] Takahiro Hirayama, Masahiro Jibiki, and Ved P. Kafle. Salience-based VNF placement method for evolving multiple network slices. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004474.

Wang:2023:EVP

[1818] Ying Wang, Jiang Liu, Mingwei Cui, Weihong Wu, and Tao Huang. ED-VNE: a profit-oriented VNE optimization scheme of energy and delay in 5G SlaaS. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004486.

Mohammadzad:2023:MMA

[1819] Maryam Mohammadzad, Jaber Karimpour, and Farnaz Mahan. MAGD: Minimal attack graph generation dynamically in cyber security. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004498.

Amiri-Zarandi:2023:SFL

[1820] Mohammad Amiri-Zarandi, Rozita A.

Dara, and Xiaodong Lin. SIDS: a federated learning approach for intrusion detection in IoT using Social Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004504.

deMattos:2023:SWS

[1821] Ekler Paulino de Mattos, Augusto C. S. A. Domingues, Fabrício A. Silva, Heitor S. Ramos, and Antonio A. F. Loureiro. Slicing who slices: Anonymization quality evaluation on deployment, privacy, and utility in mix-zones. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004528.

Khan:2023:RAT

[1822] Fátima Khan, Cristina Hervella, Luis Diez, Fátima Fernández, Néstor J. Hernández Marcano, Rune Hylsberg Jacobsen, and Ramón Agüero. Realistic assessment of transport protocols performance over LEO-based communications. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300453X.

Lu:2023:EUL

[1823] Yuan Lu, Guoyuan Yuan, Yang Bai, Dezun Dong, and Renjie Zhou. EagerCC: an ultra-low latency congestion control mechanism in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004541.

deOliveira:2023:FNN

[1824] Jonathas A. de Oliveira, Vinícius P. Gonçalves, Rodolfo I. Meneguette, Rafael T. de Sousa, Daniel L. Guidoni, José C. M. Oliveira, and Geraldo P. Rocha Filho. F-NIDS — a network intrusion detection system based on federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004553.

Banotra:2023:EHS

[1825] Atul Banotra, Sarbani Ghose, Deepak Mishra, and Sudhakar Modem. Energy harvesting in self-sustainable IoT devices and applications based on cross-layer architecture design: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004565.

Lu:2023:URU

[1826] Qiuyu Lu, June Li, Kai Yuan, Kaipei Liu, Ming Ni, and Jianbo Luo. UDP-RT: a UDP-based reliable transmission scheme for power WAPS. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004577.

Liu:2023:PWP

[1827] Pei Liu, Bing Qian, Qi Sun, and Longgang Zhao. Prompt-WNQA: a prompt-based complex question answering for wireless network over knowledge graph. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004590.

Chaudhary:2023:DAI

[1828] Shubhankar Chaudhary and Pramod Kumar Mishra. DDoS attacks in Industrial IoT: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004607.

M:2023:DRL

[1829] Jesús F. Cevallos M., Alessandra Rizzardi, Sabrina Sicari, and Alberto Coen Porisini. Deep reinforcement learning for intrusion detection in Internet of Things: Best practices, lessons learnt, and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004619.

Amin:2023:PAD

[1830] Md. Ruhul Amin, Md. Shohrab Hossain, and Mohammed Atiquzzaman. Performance analysis of distributed coordination function with early collision detection in in-band full-duplex wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004620.

Okhovvat:2023:JTE

[1831] Morteza Okhovvat, Mohammad Taghi Kheirabadi, Mohammad Reza Okhovvat, and Ali Nodehi. Joint time and energy-optimal approach to allocate task to actors in wireless sensor actor networks. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004632.

Wang:2023:CCM

[1832] Zhendong Wang, Liwei Shao, Shuxin Yang, Junling Wang, and Dahai Li. CRLM: a cooperative model based on reinforcement learning and metaheuristic algorithms of routing protocols in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004644.

Liang:2023:MAM

[1833] Yu Liang and Sheng Zhang. Mobilityaware multi-user service placement and resource allocation in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004656.

Wang:2023:LZK

[1834] Zhangquan Wang, Jiaxuan Huang, Kelei Miao, Xiaowen Lv, Yourong Chen, Bing Su, Liyuan Liu, and Meng Han. Lightweight zero-knowledge authentication scheme for IoT embedded devices. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004668.

Awad:2023:EED

[1835] Sumaya D. Awad, A. Sali, Mohanad M.

Al-Wani, Ali M. Al-Saegh, J. S. Mandeep, and R. S. A. Raja Abdullah. End-to-end DVB-S2X system design with DL-based channel estimation over satellite fading channels at Ka-band. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300467X.

Xiao:2023:DPB

[1836] Guangbing Xiao, Haibo Zhang, Zhiyi Huang, and Yawen Chen. Decentralized piggybacking-based dissemination of Cooperative Awareness Messages in vehicular ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004772.

Li:2023:PSS

[1837] Man Li, Shuangxing Deng, Huachun Zhou, and Yajuan Qin. A path selection scheme for detecting malicious behavior based on deep reinforcement learning in SDN/NFVenabled network. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004796.

Li:2023:GDA

[1838] Kun Li, Huachun Zhou, Zhe Tu, Ouyang Liu, and Hongke Zhang. AT-GCN: a DDoS attack path tracing system based on attack traceability knowledge base and GCN. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004814.

Garcia-Rodriguez:2023:PPA

[1839] Jesús García-Rodríguez and Antonio Skarmeta. A privacy-preserving attribute-based framework for IoT identity lifecycle management. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862300484X.

Liu:2023:BBI

[1840] Zhenpeng Liu, Shuo Wang, and Yi Liu.

Blockchain-based integrity auditing for shared data in cloud storage with file prediction. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004851.

Tian:2023:PEB

[1841] He Tian, Kaihong Guo, Ran Zhang, and Shiliang Shao. Prediction of evolution behavior of Internet bottleneck delay based on improved logistic equation. Computer Networks (Amsterdam, Netherlands: 1999), 236 (??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004863.

Anonymous:2023:EBq

[1842] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623005121.

Anonymous:2023:Nb

[1843] Anonymous. November 2023. Computer Networks (Amsterdam, Netherlands: 1999), 236(??):??, November 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mahboob:2023:CVC

[1844] Shadab Mahboob, Koushik Kar, Jacob Chakareski, and Md Ibrahim Alam. CLoSER: Video caching in small-cell edge networks with local content sharing. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004784.

Fernandez:2023:SAH

[1845] Yaime Fernández, Javier E. Soto, Sofía Vera, Yasmany Prieto, Cecilia Hernández, and Miguel Figueroa. A streaming algorithm and hardware accelerator to estimate the empirical entropy of network flows. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004802.

Leyva-Pupo:2023:ISU

[1846] Irian Leyva-Pupo and Cristina Cervelló-Pastor. An intelligent scheduling for 5G user plane function placement and chaining reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004826.

Omri:2023:EDN

[1847] Aymen Omri, Javier Hernandez Fernandez, and Roberto Di Pietro. Extending device noise measurement capacity for OFDM-based PLC systems: Design, implementation, and onfield validation. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004838.

Sartayeva:2023:SIP

[1848] Yerkezhan Sartayeva, Henry C. B.

Chan, Yik Him Ho, and Peter H. J. Chong. A survey of indoor positioning systems based on a six-layer model. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004875.

Liang:2023:LNB

[1849] Weibiao Liang, Lin You, and Gengran Hu. LRS PKI: a novel blockchainbased PKI framework using linkable ring signatures. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004887.

Taleb:2023:AMB

[1850] Tarik Taleb, Chafika Benzaïd, Rami Akrem Addad, and Konstantinos Samdanis. AI/ML for beyond 5G systems: Concepts, technology enablers and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004899.

Wang:2023:MKS

[1851] Miaomiao Wang, Lanlan Rui, Siya Xu, Zhipeng Gao, Huiyong Liu, and Shaoyong Guo. A multi-keyword searchable encryption sensitive data trusted sharing scheme in multi-user scenario. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004905.

Arat:2023:NMV

[1852] Ferhat Arat and Sedat Akleylek. A new method for vulnerability and risk assessment of IoT. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623004917.

Dong:2023:SSS

[1853] Enhuan Dong, Peng Gao, Yuan Yang, Mingwei Xu, Xiaoming Fu, and Jiahai Yang. SmartSBD: Smart shared bottleneck detection for efficient multipath congestion control over heterogeneous networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004929.

Wang:2023:BEF

[1854] Houtian Wang, Taotao Wang, Long Shi, Naijin Liu, and Shengli Zhang. A blockchain-empowered framework for decentralized trust management in Internet of Battlefield Things. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004930.

S:2023:DED

[1855] Ahmad Salehi S., Runchao Han, Carsten Rudolph, and Marthie Grobler. DACP: Enforcing a dynamic access control policy in cross-domain environments. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004942.

Zhu:2023:SAE

[1856] Yan Zhu and Dawei Lu. Satelliteassisted edge computing management based on deep reinforcement learning in Industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004954.

Bozorgchenani:2023:NMO

[1857] Arash Bozorgchenani, Charilaos C. Zarakovitis, Su Fong Chien, Tiew On Ting, Qiang Ni, and Wissam Mallouli. Novel modeling and optimization for joint Cybersecurity-vs-QoS Intrusion Detection Mechanisms in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004966.

Zhang:2023:CVN

[1858] Guoan Zhang, Wei Cao, and Xiaohui Gu. Cooperative vehicular networks over Nakagami-m fading: Joint power control and spectrum scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004978.

Liu:2023:OIS

[1859] Yujie Liu, Shangping Wang, Duo Zhang, Qian Zhang, and Jifang Wang. Optimal incentive strategy in blockchain-based mobile crowdsensing using game theory. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300498X.

Gomez:2023:SNS

[1860] Jose Gomez, Elie F. Kfoury, Jorge Crichigno, and Gautam Srivastava. A survey on network simulators, emulators, and testbeds used for research and education. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623004991.

Zhang:2023:ESL

[1861] Yuandi Zhang, Jiawangnan Lu, Hongtao Zhang, Ziyi Huang, César Briso-Rodríguez, and Lei Zhang. Experimental study on low-altitude UAVto-ground propagation characteristics in campus environment. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005005.

Wang:2023:GGB

[1862] Xiao Wang and Xudong Wang. Graph-PowerNet: Graph-based power consumption profiling for mobile phone applications. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005017.

Ahmad:2023:ISR

[1863] Muneeb Ahmad, Tanzeela Shakeel, and Soo Young Shin. Image super resolution based channel estimation for future wireless communication. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005029.

Robitzsch:2023:PAM

[1864] Sebastian Robitzsch, Marco Centenaro, Nicola di Pietro, Luis Cordeiro, André S. Gomes, Peter Sanders, and Arif Ishaq. Prospects on the adoption of a microservice-based architecture in 5G systems and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005030.

Zhang:2023:GSE

[1865] Quanwei Zhang, Qingjun Xiao, and Yuexiao Cai. A generic sketch for estimating super-spreaders and per-flow cardinality distribution in high-speed data streams. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005042.

Rochman:2023:CAC

[1866] Muhammad Iqbal Rochman, Vanlin Sathya, Damian Fernandez, Norlen Nunez, Ahmed S. Ibrahim, William Payne, and Monisha Ghosh. A comprehensive analysis of the coverage and performance of 4G and 5G deployments. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005054.

Ahmadi:2023:SAE

[1867] Mojtaba Ahmadi, Sahar Kianian, and Zahra Shirmohammadi. Social-aware energy management in D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005066.

Zhu:2023:FFF

[1868] Zhengbin Zhu, Hong Yu, Qinrang Liu, Dongpei Liu, and Bo Mei. FFRLI: Fast fault recovery scheme based on link importance for data plane in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005078.

Chang:2023:DES

[1869] Wei-Teng Chang and Ben-Jye Chang.

Dynamic extended sigmoid-based scheduling with virtualized RB allocation for maximizing frequency numerology efficiency in 5G/B5G NR networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300508X.

Luo:2023:NCC

[1870] Pengcheng Luo, Yuan Liu, Zekun Wang, Jian Chu, and Genke Yang. A novel Congestion Control algorithm based on inverse reinforcement learning with parallel training. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005169.

Ullah:2023:TIT

[1871] Safi Ullah, Jawad Ahmad, Muazzam A. Khan, Mohammed S. Alshehri, Wadii Boulila, Anis Koubaa, Sana Ullah Jan, and M Munawwar Iqbal Ch. TNN-IDS: Transformer neural network-based intrusion detection system for MQTTenabled IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005170.

Hajar:2023:RMA

[1872] Muhammad Shadi Hajar, Harsha Kumara Kalutarage, and M. Omar Al-Kadri. 3R: a reliable multi agent reinforcement learning based routing protocol for wireless medical sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005182.

Abdelhakam:2023:CCT

[1873] Mostafa M. Abdelhakam, Mahmoud M. Elmesalawy, Ibrahim I. Ibrahim, and Samir G. Sayed. Collaborative CoMP and trajectory optimization for energy minimization in multi-UAV-assisted IoT networks with QoS guarantee. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005194.

Liu:2023:JMT

[1874] Jingyao Liu, Guangsheng Feng, Zhibo Zhang, Liying Zheng, Huiqiang Wang, and Jyri Hämäläinen. Joint mixedtimescale optimization of content caching and delivery policy in NOMAbased vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005200.

Moorthy:2023:OHS

[1875] Sabarish Krishna Moorthy, Nicholas Mastronarde, Elizabeth Serena Bentley, Michael Medley, and Zhangyu Guan. OSWireless: Hiding specification complexity for zero-touch software-defined wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128623005212.

Wu:2023:MAC

[1876] Lisheng Wu, Xiaoming You, and Sheng Liu. Multi-ant colony algorithm based on cooperative game and dynamic path tracking. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005224.

Kurihara:2023:ODA

[1877] Jun Kurihara, Toshiaki Tanaka, and Takeshi Kubo. μODNS: a distributed approach to DNS anonymization with collusion resistance. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005236.

Wu:2023:LVS

[1878] Qiyu Wu, Fucai Zhou, Jian Xu, and Da Feng. Lightweight and verifiable secure aggregation for multi-dimensional data in edge-enhanced IoT. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005248.

Ren:2023:NNA

[1879] Qingqing Ren, Shuyong Zhu, Lu Lu, Zhiqiang Li, Guangyu Zhao, and Yujun Zhang. NetShield: an in-network architecture against Byzantine failures in distributed deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005261.

Sahin:2023:NBS

[1880] Emre Sahin, Mustafa Alper Akkas, and Orhan Dagdeviren. Nanonetworkbased search and rescue operations in debris areas. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005273.

Chen:2023:ELS

[1881] Ciyuan Chen, Junzhou Luo, Zhuqing Xu, Runqun Xiong, Dian Shen, and Zhimeng Yin. Enabling large-scale lowpower LoRa data transmission via multiple mobile LoRa gateways. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005285.

Jawad:2023:CSB

[1882] Aqeel Thamer Jawad, Rihab Maaloul, and Lamia Chaari. A comprehensive survey on 6G and beyond: Enabling technologies, opportunities of machine learning and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005303.

Li:2023:ABA

[1883] Peng Li, Junzuo Lai, Ye Yang, Meng Sun, Chi Wu, Wei Wu, and Xiaowei Yuan. Attribute-based anonymous credential: Delegation, traceability, and revocation. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005315.

Gan:2023:STD

[1884] Chenquan Gan, Anqi Liu, Qingyi Zhu, Ye Zhu, Yong Xiang, and Jun Liu. Social tie-driven coupling propagation of user awareness and information in device-to-device communications. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005327.

Awad:2023:MRP

[1885] Mirna Awad, Aris Leivadeas, and Abir Awad. Multi-resource predictive workload consolidation approach in virtualized environments. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005339.

Gonzalez:2023:QEM

[1886] N. González, M. Solera, F. Ruiz, C. Gijón, and M. Toril. A quality of experience model for live video in first-person-view drone control in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005340.

Wang:2023:STW

[1887] Wei Wang, Weiping Zhao, Dingcheng Yang, An Li, Yapeng Wang, and Lin Xiao. Secure two-way transmission via untrusted UAV relay: Joint path design and slot-pairing strategy. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005352.

Li:2023:DAM

[1888] Xia Li, Xiaoli Wang, and Yuping Wang. DRL assisted multi-objective algorithm for multicast scheduling in elastic optical network. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005364.

Dangi:2023:FSB

[1889] Ramraj Dangi and Praveen Lalwani. Feature selection based machine learning models for 5G network slicing approximation. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005388.

Zaballos:2023:UTT

[1890] Agustín Zaballos, Adrià Mallorquí, and Joan Navarro. Unboxing trustworthiness through quantum internet. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300539X.

Ebrahim:2023:PAL

[1891] Maad Ebrahim and Abdelhakim Hafid. Privacy-aware load balancing in fog networks: a reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005406.

Xiao:2023:FRP

[1892] Qingjun Xiao, Yifei Li, and Yeke Wu. Finding recently persistent flows in high-speed packet streams based on cuckoo filter. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300542X.

Xu:2023:CFL

[1893] Yuwei Xu, Jie Cao, Kehui Song, Qiao Xiang, and Guang Cheng. Corrigendum to “FastTraffic: a lightweight method for encrypted traffic fast classification” [Computer Networks, Volume 235, November 2023, 109965]. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005431. See

[1786].

Takeda:2023:LPM

[1894] Kenta Takeda, Takehiro Sato, Bijoy Chand Chatterjee, and Eiji Oki. Lightpath provisioning model considering crosstalk-derived fragmentation in spectrally-spatially elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005443.

Liu:2023:FJC

[1895] Su Liu, Peiyuan Guan, Jiong Yu, and Amir Taherkordi. FedSSC: Joint client selection and resource management for communication-efficient federated vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005455.

Bolla:2023:MGS

[1896] Raffaele Bolla, Roberto Bruschi, Chiara Lombardo, Alireza Mohammadpour, Riccardo Trivisonno, and Wint Yi Poe. A 5G multi-gNodeB simulator for ultra-reliable 0.5–100 GHz communication in indoor Industry 4.0 environments. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005480.

Chaudhary:2023:EIC

[1897] Pankaj Chaudhary, Neminath Hubballi, and Sameer G. Kulkarni. eN-Cache: Improving content delivery with cooperative caching in Named Data Networking. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005492.

Tan:2023:NRO

[1898] Haining Tan, Tao Ye, Sadaqat ur Rehman, Obaid ur Rehman, Shanshan Tu, and Jawad Ahmad. A novel routing optimization strategy based on reinforcement learning in perception layer networks. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005509.

Quan:2023:FOC

[1899] Yu xuan Quan, Yu ning Dong, Yang Xiang, Shan shan Chen, Zai jian Wang, and Jiong Jin. Fast online classification of network traffic using new feature-embedded hierarchical structure. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005510.

Yang:2023:LPH

[1900] Xu Yang, Jiahe Yu, Saiyu Qi, Qiuhao Wang, Jianfeng Wang, Yanan Qiao, and Yong Qi. Less payment and higher efficiency: a verifiable, fair and forward-secure range query scheme using blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005534.

Karmakar:2023:TAO

[1901] Kallol Krishna Karmakar, Vijay Varadharajan, Michael Hitchens, Uday Tupakula, and Prajna Sariputra. A trustaware OpenFlow switching framework for software defined networks (SDN). Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005546.

Hu:2023:FGE

[1902] Xiaoyan Hu, Zhuozhuo Shu, Zhongqi Tong, Guang Cheng, Ruidong Li, and Hua Wu. Fine-grained Ethereum behavior identification via encrypted traffic analysis with serialized backward inference. Computer Networks (Amsterdam, Netherlands: 1999), 237 (??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005558.

Chen:2023:VYB

[1903] Yitong Chen, Chaoqin Gan, Shibao Wu, Xiaoqi Wang, and Yixin Chen. (Vehicle’s YB)-based LPB prediction model and anti-LPB (communication link)-reconstruction strategy in the (VLC/RF)-V2X network. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005571.

Anonymous:2023:EBr

[1904] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128623005716.

Anonymous:2023:D

[1905] Anonymous. December 2023. Computer Networks (Amsterdam, Netherlands: 1999), 237(??):??, December 2023. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Premalatha:2024:OEE

[1906] B. Premalatha and P. Prakasam. Optimal energy-efficient resource allocation and fault tolerance scheme for task offloading in IoT-FoG computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300525X.

Wu:2024:QBR

[1907] Yutong Wu, Jianyue Zhu, Xiao Chen, Yu Zhang, Yao Shi, and Yaqin Xie. QoS-based resource allocation for uplink NOMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005297.

Xiong:2024:EAL

[1908] Bing Xiong, Jing Wu, Qiaorong Huang, Jinyuan Zhao, Qiang Tang, Jin Zhang, Kun Yang, and Keqin Li. Elastically accelerating lookup on virtual SDN flow tables for softwaredefined cloud gateways. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005376.

Dou:2024:MIC

[1909] Songshi Dou, Li Qi, and Zehua Guo. Mitigating the impact of controller failures on QoS robustness for softwaredefined wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005418.

Al-hammadi:2024:CCO

[1910] Ikhlas Al-hammadi, Mingchu Li, Sardar M. N. Islam, and Esmail Al-Mosharea. Collaborative computation offloading for scheduling emergency tasks in SDN-based mobile edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005467.

Wang:2024:ETP

[1911] Ke Wang, Xiaojuan Ma, Heng Kang, Zheng Lyu, Baorui Feng, Wenliang Lin, Zhongliang Deng, and Yun Zou. An efficient topology partitioning algorithm for system-level parallel simulation of mega satellite constellation communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005479.

Kashif:2024:EEP

[1912] Muhammad Kashif and Kubra Kalkan.

EPIoT: Enhanced privacy preservation based blockchain mechanism for Internet-of-Things. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005522.

Alam:2024:JES

[1913] Amjad Alam, Kamran Ali, Ramona Trestian, Purav Shah, and Glenford Mapp. Joint energy and spectral optimization in heterogeneous vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300556X.

Islam:2024:PIP

[1914] SK Hafizul Islam, Krittibas Parai, and Daya Sagar Gupta. PF-IBDA: Provably secure and pairing-free identitybased deniable authentication protocol for MANET environments. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005583.

Le:2024:ACU

[1915] Siqi Le, Yingxu Lai, Yipeng Wang, and Huijie He. An adaptive classification and updating method for unknown network traffic in open environments. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005595.

Gu:2024:TRT

[1916] Xiaolin Gu, Wenjia Wu, Yusen Zhou, Aibo Song, Ming Yang, Zhen Ling, and Junzhou Luo. TEA-RFFI: Temperature adjusted radio frequency fingerprint-based smartphone identification. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005601.

Lin:2024:SPC

[1917] Hsin-Tsung Lin and Pi-Chung Wang. Scalable packet classification based on rule categorization and crossproducting. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005613.

Yang:2024:RBO

[1918] Sijin Yang, Lei Zhuang, Julong Lan, Jianhui Zhang, and Bingkui Li. Reusebased online joint routing and scheduling optimization mechanism in deterministic networks. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005625.

Zhang:2024:PIP

[1919] Ze Zhang, Chingfang Hsu, Man Ho Au, Lein Harn, Jianqun Cui, Zhe Xia, and Zhuo Zhao. PRLAP-IoD: a PUF-based robust and lightweight authentication protocol for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005637.

Kankane:2024:CRW

[1920] Bhawna Kankane, Sandeep Sharma, and Rajesh Mishra. κ-coverage reliability for wireless multihop network incorporating boundary effect. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005649.

Yang:2024:PPA

[1921] Xiaodu Yang, Sijie Ruan, Yinliang Yue, and Bo Sun. PETNet: Plaintextaware encrypted traffic detection network for identifying Cobalt Strike HTTPS traffics. Computer Networks (Amsterdam, Netherlands: 1999), 238 (??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005650.

Wang:2024:MCM

[1922] Rui Wang, Gang Tian, and Shi Ying. MicroCM: a cloud monitoring architecture for microservice invocation. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005662.

Kumar:2024:SCB

[1923] Neeraj Kumar and Rifaqat Ali. A smart contract-based robotic surgery authentication system for healthcare using 6G-Tactile Internet. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005789.

Chakour:2024:UQP

[1924] Imane Chakour, Sajida Mhammedi, Cherki Daoui, and Mohamed Baslam. Unlocking QoS potential: Integrating IoT services and Monte Carlo control for heterogeneous IoT device management in gateways. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005790.

Anonymous:2024:EBa

[1925] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005960.

Anonymous:2024:Ja

[1926] Anonymous. January 2024. Computer Networks (Amsterdam, Netherlands: 1999), 238(??):??, January 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Sousa:2024:FSF

[1927] Bruno Sousa and Carolina Gonçalves.

FedAAA-SDN: Federated authentication, authorization and accounting in SDN controllers. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005753.

Ramos:2024:ETC

[1928] Daniel Ramos, Oscar Esparza, Jorge Mata-Díaz, and Juanjo Alins. Evaluation of TCP congestion control algorithms with traffic control policies in a PEP-based geosynchronous satellite scenario. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005765.

Chen:2024:GNN

[1929] Hongyang Chen, Pengfei Chen, Benran Wang, Xian Yu, Xiaofan Chen, Dandan Ma, and Zibin Zheng. Graph neural network based robust anomaly detection at service level in SDN driven microservice system. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005807.

Bine:2024:CID

[1930] Lailla M. S. Bine, Azzedine Boukerche, Linnyer B. Ruiz, and Antonio A. F. Loureiro. Connecting Internet of Drones and urban computing: Methods, protocols and applications. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005819.

Gonzalez-Soto:2024:DCM

[1931] Martín González-Soto, Rebeca P. Díaz-Redondo, Manuel Fernández-Veiga, Bruno Fernández-Castro, and Ana Fernández-Vilas. Decentralized and collaborative machine learning framework for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005820.

Cen:2024:RED

[1932] Mingcan Cen, Frank Jiang, Xingsheng Qin, Qinghong Jiang, and Robin Doss. Ransomware early detection: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005832.

Duan:2024:TLB

[1933] Zhonglei Duan and Ting Wang. Towards learning-based energy-efficient online coordinated virtual network embedding framework. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005844.

Mohammed:2024:CSR

[1934] Ahmed Burhan Mohammed, Lamia Chaari Fourati, and Ahmed M. Fakhrudeen. Comprehensive systematic review of intelligent approaches in UAV-based intrusion detection, blockchain, and network security. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005856.

Kita:2024:UAA

[1935] Nozomi Kita and Eiji Oki. Unavailabilityaware allocation of backup resources considering failures of virtual and physical machines. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005868.

Cao:2024:WTL

[1936] Tuo Cao, Qinhui Wang, Yuhan Zhang, Zhuzhong Qian, Yue Zeng, Mingtao Ji, Hesheng Sun, and Baoliu Ye. Walking on two legs: Joint service placement and computation configuration for provisioning containerized services at edges. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005893.

Yu:2024:EEN

[1937] Ya-Ju Yu and Shan-Yao Lo. Energyefficient non-anchor channel allocation in NB-IoT cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862300590X.

Bimpas:2024:LPC

[1938] Athanasios Bimpas, John Violos, Aris Leivadeas, and Iraklis Varlamis. Leveraging pervasive computing for ambient intelligence: a survey on recent advancements, applications and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006011.

Trabelsi:2024:IMB

[1939] Nessrine Trabelsi, Lamia Chaari Fourati, and Chung Shue Chen. Interference management in 5G and beyond networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006047.

Imputato:2024:BWF

[1940] Pasquale Imputato, Stefano Avallone, Malcolm Smith, David Nunez, and Boris Bellalta. Beyond Wi-Fi 7: Spatial reuse through multi-AP coordination. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006059.

Doriguzzi-Corin:2024:IPL

[1941] Roberto Doriguzzi-Corin, Luis Augusto Dias Knob, Luca Mendozzi, Domenico Siracusa, and Marco Savi. Introducing packet-level analysis in programmable data planes to advance Network Intrusion Detection. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006072.

Tu:2024:PDD

[1942] Huaqing Tu, Gongming Zhao, Hongli Xu, and Chunming Qiao. Programmable device deployment for efficient network function offloading. Computer Networks (Amsterdam, Netherlands: 1999), 239 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006084.

Anonymous:2024:EBb

[1943] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624000094.

Anonymous:2024:Fa

[1944] Anonymous. February 2024. Computer Networks (Amsterdam, Netherlands: 1999), 239(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Firouzjaee:2024:UPA

[1945] Mostafa Taghizade Firouzjaee, Kamal Jamshidi, Neda Moghim, and Sachin Shetty. User preference-aware content caching strategy for video delivery in cache-enabled IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862300587X.

Griner:2024:BMD

[1946] Chen Griner, Chen Avin, and Gil Einziger. Beyond matchings: Dynamic multi-hop topology for demandaware datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005881.

Younes:2024:HDL

[1947] Osama S. Younes. A hybrid deep learning model for detecting DDoS flooding attacks in SIP-based systems. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623005911.

Koumar:2024:NEI

[1948] Josef Koumar, Karel Hynek, Jaroslav Pesek, and Tomás Cejka. Net-TiSA: Extended IP flow with timeseries features for universal bandwidthconstrained high-speed network traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623005923.

Seifhosseini:2024:MOC

[1949] Seyyedamin Seifhosseini, Mirsaeid Hosseini Shirvani, and Yaser Ramzanpoor. Multi-objective cost-aware bagof-tasks scheduling optimization model for IoT applications running on heterogeneous fog environment. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006060.

Song:2024:IIA

[1950] Junda Song, Jiajin Liu, Peixuan Gao, Guyue Liu, and H. Jonathan Chao. Inversion impact of approximate PIFO to Start-Time Fair Queueing. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006096.

Chang:2024:BSG

[1951] Sang-Yoon Chang, Arijet Sarker, Simeon Wuthier, Jinoh Kim, Jonghyun Kim, and Xiaobo Zhou. Base station gateway to secure user channel access at the first hop edge. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128623006102.

Pathan:2024:PBE

[1952] Md. Naimul Pathan, Maisha Muntaha, Selina Sharmin, Sajeeb Saha, Md. Ashraf Uddin, Fernaz Narin Nur, and Sunil Aryal. Priority based energy and load aware routing algorithms for SDN enabled data center network. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006114.

Li:2024:DLB

[1953] Shuyang Li, Enrico Magli, Gianluca Francini, and Giorgio Ghinamo. Deep learning based prediction of traffic peaks in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006126.

Wadatkar:2024:JMO

[1954] Prachi Vinod Wadatkar, Rosario Giuseppe Garroppo, Gianfranco Nencioni, and Marco Volpi. Joint multi-objective MEH selection and traffic path computation in 5G-MEC systems. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128623006138.

Zhai:2024:NEH

[1955] Chao Zhai, Jiachao Yu, Kun Du, Hao Dai, Xinhua Wang, and Yujun Li. Nonlinear energy harvesting based alternate cooperative nonorthogonal multiple access with adaptive interference cancellation. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400001X.

Zhang:2024:RBR

[1956] Yiguo Zhang, Wei Wang, and Feiyang Shi. Reputation-based Raft-Poa layered consensus protocol converging UAV network. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000021.

Arif:2024:CJA

[1957] Mohammad Arif and Abdul Shakoor. Clustered jamming and antenna beamwidth fluctuations for UAV-assisted cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000033.

Hu:2024:INT

[1958] Xueyan Hu, Wei Liu, and Hua Huo. An intelligent network traffic prediction method based on Butterworth filter and CNN-LSTM. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000045.

Garcia-Martinez:2024:BCM

[1959] Alberto García-Martínez, Pedro Rodrigues Torres, and Marcelo Bagnulo. BGP convergence in an MRAI-free Internet. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400015X.

Zheng:2024:NLL

[1960] Xiao Zheng, Yan Liu, Shuangping Zhan, Yao Xin, and Yi Wang. A novel low-latency scheduling approach of TSN for multi-link rate networking. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000161.

Cai:2024:ACL

[1961] Yinan Cai, Wei Li, Xiangxu Meng, Wenqi Zheng, Chuhao Chen, and Zhuoxuan Liang. Adaptive contrastive learning based network latency prediction in 5G URLLC scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000173.

Qiao:2024:BBM

[1962] Yuncheng Qiao, Qiujun Lan, Shiyu Jia, Xianhua Kuang, Qingyue Xiong, and Chaoqun Ma. Blockchain-based multiinstitution collaborative credit evaluation modeling method. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000197.

Hara:2024:PKU

[1963] Takanori Hara and Masahiro Sasabe. Practicality of in-kernel/user-space packet processing empowered by lightweight neural network and decision tree. Computer Networks (Amsterdam, Netherlands: 1999), 240 (??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000203.

Pi:2024:MCI

[1964] Yibo Pi, Sugih Jamin, and Yichen Wei. Measuring congestion-induced performance imbalance in Internet load balancing at scale. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000215.

Magoula:2024:SCI

[1965] Lina Magoula, Nikolas Koursioumpas, Theodora Panagea, Nancy Alonistioti, Chaima Ghribi, and Joshua Shakya. SIM+: a comprehensive implementation-agnostic information model assisting AI-driven optimization for beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000227.

Anonymous:2024:EBc

[1966] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624000598.

Anonymous:2024:Fb

[1967] Anonymous. February 2024. Computer Networks (Amsterdam, Netherlands: 1999), 240(??):??, February 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pieska:2024:LDC

[1968] Marcus Pieska, Alexander Rabitsch, Anna Brunstrom, Andreas Kassler, Markus Amend, and Eckard Bogenfeld. Low-delay cost-aware multipath scheduling over dynamic links for access traffic steering, switching, and splitting. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000185.

Teglasy:2024:AUA

[1969] Bálint Z. Téglásy, Emil Wengle, John R. Potter, and Sokratis Katsikas. Authentication of underwater assets. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000239.

Wen:2024:RRB

[1970] Tian Wen, Hanqing Zhang, Han Zhang, Huixin Wu, Danxin Wang, Xiuwen Liu, Weishan Zhang, Yuwei Wang, and Shaohua Cao. RTIFed: a Reputation based Triple-step Incentive mechanism for energy-aware Federated learning over battery-constricted devices. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000240.

Gulzari:2024:CAM

[1971] Usman Ali Gulzari, Zoran Salcic, Waqar Farooq, Sheraz Anjum, Sarzamin Khan, Muhammad Sajid, and Frank Sill Torres. Comparative analysis of 2D mesh topologies with additional communication links for onchip networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000252.

Busacca:2024:SLO

[1972] Fabio Busacca, Stefano Mangione, Sergio Palazzo, Francesco Restuccia, and Ilenia Tinnirello. SDR-LoRa, an opensource, full-fledged implementation of LoRa on Software-Defined-Radios: Design and potential exploitation. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000264.

Liu:2024:QAB

[1973] Ying Liu, Jiawang Zhi, Xiaoyu Xia, Yuzheng Han, Changsheng Zhang, and Bin Zhang. QoE-aware budgeted edge data caching online: a primal-dual approach. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000276.

Lin:2024:PPB

[1974] Yuming Lin, Youjia Jiang, You Li, and Ya Zhou. Privacy-preserving batchbased task assignment over spatial crowdsourcing platforms. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000288.

Ali:2024:IZT

[1975] Belal Ali, Mark A. Gregory, Shuo Li, and Omar Amjad Dib. Implementing zero trust security with dual fuzzy methodology for trust-aware authentication and task offloading in Multiaccess Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400029X.

Gohar:2024:OCM

[1976] Ali Gohar and Gianfranco Nencioni. An online cost minimization of the slice broker based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000306.

Lei:2024:RES

[1977] Wenli Lei, Yang Lei, Hongbo Guo, and Kun Jia. Research on energysaving relay selection algorithm based on olive forwarding area in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000318.

Song:2024:QIS

[1978] Jiahui Song. Quantifying the importance of specific types of higher-order structures in complex networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400032X.

Chen:2024:MAD

[1979] Lin Chen, Weijia Wang, Yichuan Song, Jihong Yu, Kehao Wang, Shurong Zhang, Weihua Yang, and Celimuge Wu. Meet in the air: Distributed neighbor discovery in 3D networks with directional transceivers. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000331.

Moazzeni:2024:VTN

[1980] Shadi Moazzeni, Konstantinos Katsaros, Nasim Ferdosian, Konstantinos Antonakoglou, Mark Rouse, Dritan Kaleshi, Adriana Fernández-Fernández, Miguel Catalan-Cid, Constantinos Vrontos, Reza Nejabati, and Dimitra Simeonidou. 5G-VIOS: Towards next generation intelligent interdomain network service orchestration and resource optimisation. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000343.

Kim:2024:ESS

[1981] Jinwoo Kim, Minjae Seo, Seungsoo Lee, Jaehyun Nam, Vinod Yegneswaran, Phillip Porras, Guofei Gu, and Seungwon Shin. Enhancing security in SDN: Systematizing attacks and defenses from a penetration perspective. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000355.

Ozaki:2024:MSM

[1982] Shintaro Ozaki, Takehiro Sato, and Eiji Oki. Multi-source multicast service chaining model guaranteeing reliability of network services. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000367.

Liu:2024:ESR

[1983] Dongquan Liu, Zhengyan Zhou, Dong Zhang, Kaiwei Guo, Yanni Wu, and Chunming Wu. Efficient service reconfiguration with partial virtual network function migration. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000379.

BenMazziane:2024:TML

[1984] Younes Ben Mazziane, Sara Alouf, Giovanni Neglia, and Daniel S. Menasche. TTL model for an LRU-based similarity caching policy. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000380.

Kousias:2024:EPA

[1985] Konstantinos Kousias, Mohammad Rajiullah, Giuseppe Caso, Ozgu Alay, Anna Brunstrom, Usman Ali, Luca De Nardis, Marco Neri, and Maria-Gabriella Di Benedetto. Empirical performance analysis and ML-based modeling of 5G non-standalone networks. Computer Networks (Amsterdam, Netherlands: 1999), 241 (??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000392.

Rossanese:2024:DVS

[1986] Marco Rossanese, Placido Mursia, Andres Garcia-Saavedra, Vincenzo Sciancalepore, Arash Asadi, and Xavier Costa-Perez. Design and validation of scalable reconfigurable intelligent surfaces. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000409.

Singh:2024:PPQ

[1987] Nisha Singh, Mukesh Kumar, and Bhaskar Biswas. PQCLP: Parameterized quantum circuit based link prediction in dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000422.

Ji:2024:IPB

[1988] Mingtao Ji, Chenwei Su, Yitao Fan, Yibo Jin, Zhuzhong Qian, Yuting Yan, Yu Chen, Tuo Cao, Sheng Zhang, and Baoliu Ye. INTaaS: Provisioning in-band network telemetry as a service via online learning. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000434.

Thomas:2024:DTA

[1989] Yannis Thomas, Stavros Toumpis, and Nikolaos Smyrnioudis. Digital twin approach to estimating and utilizing the capacity region of wireless ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000458.

Chong:2024:MPU

[1990] Kah Meng Chong and Amizah Malip. May the privacy be with us: Correlated differential privacy in location data for ITS. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400046X.

Zormati:2024:RAR

[1991] Mohamed Ali Zormati, Hicham Lakhlef, and Sofiane Ouni. Review and analysis of recent advances in intelligent network softwarization for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000471.

Lowes:2024:PPA

[1992] Gavin James Lowes and Jeffrey Neasham. PADAL — Passive Acoustic Detection and Localisation: Low energy underwater wireless vessel tracking network. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000483.

Yuan:2024:HEW

[1993] Xiaoyun Yuan, Tengyao Li, Lingling Li, Ruixiang Li, Zhiquan Wang, and Xiangyang Luo. HSWF: Enhancing website fingerprinting attacks on Tor to address real world distribution mismatch. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000495.

Fan:2024:DRN

[1994] Chenyu Fan, Xiaoning Zhang, Yangming Zhao, Yexiao He, and Yi Yang. Dynamic relay node selection and routing for cloud-native Software Defined WANs. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000513.

Pei:2024:EER

[1995] Xinglong Pei, Penghao Sun, Yuxiang Hu, Dan Li, Bo Chen, and Le Tian. Enabling efficient routing for traffic engineering in SDN with Deep Reinforcement Learning. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000525.

Liu:2024:ARR

[1996] Yantao Liu and Na Qin. Achievable rate regions for fractional routing. Computer Networks (Amsterdam, Netherlands: 1999), 241 (??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000537.

Wang:2024:EEM

[1997] Zhongyu Wang, Yashuai Cao, Zheng Chang, Tiejun Lv, and Wei Ni. Energy efficiency maximization in UAV communication networks with nonlinear energy harvesting. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000549.

Kumari:2024:HCP

[1998] Deepa Kumari, Abhirath Singh Parmar, Harshvadhan Sunil Goyal, Kushal Mishra, and Subhrakanta Panda. HealthRec-Chain: Patient-centric blockchain enabled IPFS for privacy preserving scalable health data. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000550.

Miano:2024:ANA

[1999] Sebastiano Miano, Giuseppe Lettieri, Gianni Antichi, and Gregorio Procissi. Accelerating network analytics with an on-NIC streaming engine. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400063X.

Yarinezhad:2024:NSA

[2000] Ramin Yarinezhad and Eylem Ekici.

A novel scheduling algorithm for LTE on unlicensed bands to ensure fair coexistence with Wi-Fi. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000641.

Anonymous:2024:EBd

[2001] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000963.

Anonymous:2024:Ma

[2002] Anonymous. March 2024. Computer Networks (Amsterdam, Netherlands: 1999), 241(??):??, March 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Huang:2024:RBT

[2003] Chung-Ming Huang and Chia-Han Hsieh. Registered-backoff-time (RBT) based channel access with grouping control for the trigger RAW mode of IEEE 802.11ah IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000410.

Liu:2024:JUG

[2004] Huanlin Liu, Xilin Yuan, Yong Chen, Yang Hu, Huiling Zhou, and Wenjie Song. Joint user grouping and resource allocation scheme based on user satisfaction management for indoor multicolor VLC-NOMA network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000446.

Amend:2024:CEM

[2005] Markus Amend and Veselin Rakocevic. Cost-efficient multipath scheduling of video-on-demand traffic for the 5G ATSSS splitting function. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000501.

Zhao:2024:LAU

[2006] Xueyuan Zhao, Zhuoran Qi, and Dario Pompili. Link adaptation in Underwater Wireless Optical Communications based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000653.

Livadariu:2024:TSC

[2007] Ioana Livadariu, Ahmed Elmokashfi, and Georgios Smaragdakis. Tracking submarine cables in the wild. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000665.

Chen:2024:SMW

[2008] Weilin Chen, Wei Yang, and Wei Gong.

A survey of millimeter wave backscatter communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000677.

Zhu:2024:LGJ

[2009] Yan Zhu and Dawei Lu. Latency guaranteed joint optimal traffic intelligent scheduling in large-scale satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000689.

Vega:2024:MLC

[2010] Christian Vega, Elie F. Kfoury, Jose Gomez, Jorge E. Pezoa, Miguel Figueroa, and Jorge Crichigno. Machine learning controller for data rate management in science DMZ networks. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000690.

Zhen:2024:GSC

[2011] Zixian Zhen, Xiangfu Zhao, Jinkai Zhang, Yichen Wang, and Haiyue Chen. DA-GNN: a smart contract vulnerability detection method based on dual attention graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000707.

Baganal-Krishna:2024:FLA

[2012] Nehal Baganal-Krishna, Ralf Lübben, Eirini Liotou, Konstantinos V. Katsaros, and Amr Rizk. A federated learning approach to QoS forecasting in cellular vehicular communications: Approaches and empirical evidence. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000719.

Yang:2024:BBC

[2013] Huan Yang, Yajun Guo, and Yimin Guo. Blockchain-based cloud-fog collaborative smart home authentication scheme. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000720.

Mohamed:2024:SFC

[2014] Ramy Mohamed, Marios Avgeris, Aris Leivadeas, Ioannis Lambadaris, John Chinneck, Todd Morris, and Petar Djukic. Service function chain network planning through offline, online and infeasibility restoration techniques. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000732.

Dai:2024:MHR

[2015] Jun Dai, Xinbin Li, Song Han, Zhixin Liu, Haihong Zhao, and Lei Yan. Multi-hop relay selection for underwater acoustic sensor networks: a dynamic combinatorial multi-armed bandit learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000744.

Alvarez-Horcajo:2024:MDP

[2016] Joaquin Alvarez-Horcajo, Isaias Martinez-Yelmo, Elisa Rojas, Juan A. Carral, and Victoria Noci-Luna. MuHoW: Distributed protocol for resource sharing in collaborative edge-computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000756.

Miao:2024:SEC

[2017] Junfeng Miao, Zhaoshun Wang, Mei Wang, Sahil Garg, M. Shamim Hossain, and Joel J. P. C. Rodrigues. Secure and efficient communication approaches for Industry 5.0 in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000768.

Du:2024:BPS

[2018] Zhiqiang Du, Liangxin Liu, Muhong Huang, Yanfang Fu, and Wendong Zhang. Bulwark: a proofof-stake protocol with strong consistency and liveness. Computer Networks (Amsterdam, Netherlands:

1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400077X.

Robin:2024:OSP

[2019] Debobroto Das Robin and Javed I.

Khan. An open-source P4_{16} compiler backend for reconfigurable matchaction table switches: Making networking innovation accessible. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000781.

Wei:2024:CCA

[2020] Zhenchun Wei, Yang Zhao, Zengwei Lyu, Xiaohui Yuan, Yu Zhang, and Lin Feng. Cooperative caching algorithm for mobile edge networks based on multi-agent meta reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000793.

Yu:2024:HLB

[2021] Hao Yu and Qiang Zhang. Hybrid learning based service migration for cost minimization with deadlines in multi-user mobile edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000811.

Sari:2024:UCB

[2022] T. Tolga Sari and Gökhan Seçinti. Using centrality based topology control for FANET survivability against jamming. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000823.

Feng:2024:MDC

[2023] Huifen Feng, Weiting Zhang, Ying Liu, Chuan Zhang, Chenhao Ying, Jian Jin, and Zhenzhen Jiao. Multi-domain collaborative two-level DDoS detection via hybrid deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000835.

Rzepka:2024:PER

[2024] Karol Rzepka, Przemysław Szary, Krzysztof Cabaj, and Wojciech Mazurczyk. Performance evaluation of Raspberry Pi 4 and STM32 Nucleo boards for security-related operations in IoT environments. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624000847.

Raj:2024:SME

[2025] Raina Raj. Stochastic modeling of enhancing energy efficiency for HAP in HAP-LEO systems using Markov regenerative process. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000859.

Wang:2024:CDS

[2026] Yuantao Wang, Zhaogang Shu, Shuwu Chen, Jiaxiang Lin, and Zhenchang Zhang. A cost and demand sensitive adjustment algorithm for service function chain in data center network. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000860.

Zhang:2024:AIE

[2027] Shunliang Zhang, Dali Zhu, and Yinlong Liu. Artificial intelligence empowered physical layer security for 6G: State-of-the-art, challenges, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000872.

Ahmed:2024:MAC

[2028] Imran Ahmed, Eiji Oki, and Bijoy Chand Chatterjee. Modeling and analysis of crosstalk-avoided and crosstalk-aware approaches in spectrally-spatially elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 242 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000884.

Scott:2024:MPD

[2029] Ben A. Scott, Michael N. Johnstone, Patryk Szewczyk, and Steven Richardson. Matrix profile data mining for BGP anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000896.

Soler:2024:PPK

[2030] David Soler, Carlos Dafonte, Manuel Fernández-Veiga, Ana Fernández Vilas, and Francisco J. Nóvoa. A privacy-preserving key transmission protocol to distribute QRNG keys using zk-SNARKs. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000914.

Yang:2024:CBP

[2031] Zhimei Yang, Changgen Peng, Chongyi Zhong, and Yangyang Long. Consortium blockchain private key protection scheme based on rational secret sharing and blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000926.

Shakeri:2024:REC

[2032] Ali Shakeri, Tianliang Zhang, Shunmugapriya Ramanathan, Miguel Razo, Marco Tacca, and Andrea Fumagalli. Reliable edge-to-core optical networks: an optimal algorithm for maximal path diversity. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001002.

Polverini:2024:RBN

[2033] Marco Polverini, Stefania Sardellitti, Sergio Barbarossa, Antonio Cianfrani, Paolo Di Lorenzo, and Marco Listanti. Reducing the In band Network Telemetry overhead through the spatial sampling: Theory and experimental results. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001014.

Pelekanakis:2024:CEC

[2034] Konstantinos Pelekanakis, Pietro Paglierani, Alberto Alvarez, and João Alves. Comparison of error correction codes via optimal channel replay of high north underwater acoustic channels. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001026.

K:2024:PAI

[2035] Aswini K. and Surendar M. Performance analyses of intelligent reflecting surface aided downlink multiuser rate-splitting multiple access system for 6G applications. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001038.

Alaez:2024:DTM

[2036] D. Aláez, P. Lopez-Iturri, M. Celaya-Echarri, L. Azpilicueta, F. Falcone, J. Villadangos, and J. J. Astrain. Digital twin modeling of open category UAV radio communications: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001087.

Luo:2024:UAF

[2037] Xijian Luo, Jun Xie, Liqin Xiong, Zhen Wang, and Yaqun Liu. UAVassisted fair communications for multipair users: a multi-agent deep reinforcement learning method. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001099.

Cai:2024:SSA

[2038] Binlei Cai, Xiaoli Wang, Bin Wang, Meihong Yang, Ying Guo, and Qin Guo. A self-stabilizing and autoprovisioning orchestration for microservices in edge-cloud continuum. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001117.

Mayahi:2024:IEI

[2039] Meysam Mayahi, Valeria Loscrí, and Anna Maria Vegni. INVISIBLE 2.0: an enhanced interference-based handover technique for visible light communications in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001129.

Liu:2024:QAT

[2040] Jingxian Liu, Yitian Wang, Duotao Pan, and Decheng Yuan. QoS-aware task offloading and resource allocation optimization in vehicular edge computing networks via MADDPG. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001142.

Anonymous:2024:EBe

[2041] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400149X.

Anonymous:2024:Aa

[2042] Anonymous. April 2024. Computer Networks (Amsterdam, Netherlands: 1999), 242(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Maier:2024:WSN

[2043] Guido Maier, Antonino Albanese, Michele Ciavotta, Nicola Ciulli, Stefano Giordano, Elisa Giusti, Alfredo Salvatore, and Giovanni Schembra. WatchEDGE: Smart networking for distributed AI-based environmental control. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400080X.

Guerra:2024:FLR

[2044] Renata Rojas Guerra, Anna Vizziello, Pietro Savazzi, Emanuele Goldoni, and Paolo Gamba. Forecasting LoRaWAN RSSI using weather parameters: a comparative study of ARIMA, artificial intelligence and hybrid approaches. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624000902.

Gu:2024:LGH

[2045] Zheyuan Gu, Gaopeng Gou, Chang Liu, Chen Yang, Xiyuan Zhang, Zhen Li, and Gang Xiong. Let gambling hide nowhere: Detecting illegal mobile gambling apps via heterogeneous graphbased encrypted traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001105.

Kilic:2024:THP

[2046] Ahmet Kilic. TLS-handshake for Plug and Charge in vehicular communications. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001130.

Batham:2024:ICF

[2047] Deepak Batham and Vandana Vikas Thakare. An improved cost functionbased class of service provisioning scheme for elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 243 (??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001154.

Malandrino:2024:EIR

[2048] Francesco Malandrino, Alessandro Nordio, and Carla Fabiana Chiasserini. Eavesdropping with intelligent reflective surfaces: Nearoptimal configuration cycling. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001166.

Zhao:2024:DIF

[2049] Zhe Zhao, Zhiliang Qiu, Weitao Pan, Hui Li, Ling Zheng, and Ya Gao. Design and implementation of a frame preemption model without guard bands for time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001178.

DiMauro:2024:HLS

[2050] Mario Di Mauro, Giovanni Galatro, Fabio Postiglione, Wei Song, and Antonio Liotta. Hybrid learning strategies for multivariate time series forecasting of network quality metrics. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400118X.

Vila:2024:REB

[2051] I. Vilà, O. Sallent, and J. Pérez-Romero. Relay-empowered beyond 5G radio access networks with edge computing capabilities. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001191.

Li:2024:TTL

[2052] Xia Li and Yuping Wang. TABDeep:

a two-level action branch architecturebased deep reinforcement learning for distributed sub-tree scheduling of online multicast sessions in EON. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001208.

Chintapalli:2024:EED

[2053] Venkatarami Reddy Chintapalli, Rajat Partani, Bheemarjuna Reddy Tamma, and Siva Ram Murthy C. Energy efficient and delay aware deployment of parallelized service function chains in NFV-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400121X.

Gudepu:2024:DHF

[2054] Venkateswarlu Gudepu, Venkatarami Reddy Chintapalli, Piero Castoldi, Luca Valcarenghi, Bheemarjuna Reddy Tamma, and Koteswararao Kondepu. The drift handling framework for open radio access networks: an experimental evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001221.

ElGhafour:2024:SCV

[2055] Mohamed G. Abd El Ghafour, Ahmed H. Abd El-Malek, Ola E. Hassan, and Mohammed Abo-Zahhad. Securing cooperative vehicular networks amid obstructing vehicles and mixed fading channels. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001233.

Larsen:2024:EMN

[2056] Line M. P. Larsen, Henrik L. Christiansen, Sarah Ruepp, and Michael S. Berger. The evolution of mobile network operations: a comprehensive analysis of open RAN adoption. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001245.

ElRajab:2024:ZTN

[2057] Mirna El Rajab, Li Yang, and Abdallah Shami. Zero-touch networks: Towards next-generation network automation. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001269.

Gallego-Madrid:2024:MLP

[2058] Jorge Gallego-Madrid, Irene Bru-Santa, Alvaro Ruiz-Rodenas, Ramon Sanchez-Iborra, and Antonio Skarmeta. Machine learning-powered traffic processing in commodity hardware with eBPF. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001270.

Peng:2024:NTP

[2059] Yufei Peng, Yingya Guo, Run Hao, and Chengzhe Xu. Network traffic prediction with Attention-based Spatial-Temporal Graph Network. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001282.

Valko:2024:RCE

[2060] Danila Valko and Daniel Kudenko.

Reducing CO_{2} emissions in a peerto-peer distributed payment network: Does geography matter in the lightning network? Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001294.

Li:2024:LBC

[2061] Zhiyuan Li and Xiaoping Xu. L2-BiTCN-CNN: Spatio-temporal features fusion-based multi-classification model for various internet applications identification. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001300.

Jin:2024:HIH

[2062] Yanliang Jin, Jiahao Zhou, and Yuan Gao. HSGAN-IoT: a hierarchical semi-supervised generative adversarial networks for IoT device classification. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001312.

Liao:2024:CFB

[2063] Zhuofan Liao, Wenqiang Deng, Shiming He, and Qiang Tang. Collaborative Filtering-based Fast Delay-aware algorithm for joint VNF deployment and migration in edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001324.

Buccafurri:2024:HDT

[2064] Francesco Buccafurri, Vincenzo De Angelis, Maria Francesca Idone, and Cecilia Labrini. A hierarchical distributed trusted location service achieving location k-anonymity against the global observer. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001336.

Chen:2024:RSC

[2065] Shiyou Chen, Lanlan Rui, Zhipeng Gao, Yang Yang, Xuesong Qiu, and Shaoyong Guo. Resource sharing for collaborative edge learning: a privacyaware incentive mechanism combined with demand prediction. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001348.

Jovanovic:2024:RIB

[2066] Zorka Jovanovic, Zhe Hou, Kamanashis Biswas, and Vallipuram Muthukkumarasamy. Robust integration of blockchain and explainable federated learning for automated credit scoring. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400135X.

He:2024:BBC

[2067] Yunhua He, Shuang Wang, Zhihao Zhou, Ke Xiao, Anke Xie, and Bin Wu. A blockchain-based carbon emission security accounting scheme. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001361.

Song:2024:EEF

[2068] Yu Song, Shilong Zhang, and Shubin Wang. An energy efficient fusing data gathering protocol in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001373.

Haj-Hassan:2024:DMP

[2069] Ali Haj-Hassan, Youcef Imine, Antoine Gallais, and Bruno Quoitin. Detecting malicious proxy nodes during IoT network joining phase. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001403.

Bi:2024:CSB

[2070] Meihua Bi, Chenghao Yan, Rui Lin, Yuanyuan Hong, and Jun Li. Client scheduling and bandwidth slicing for multiple federated learning tasks over multiple passive optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001415.

Botta:2024:AOS

[2071] Alessio Botta, Roberto Canonico, Annalisa Navarro, Giovanni Stanco, and Giorgio Ventre. Adaptive overlay selection at the SD-WAN edges: a reinforcement learning approach with networked agents. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001427.

Ardizzon:2024:RBA

[2072] Francesco Ardizzon, Paolo Casari, and Stefano Tomasin. A RNN-based approach to physical layer authentication in underwater acoustic networks with mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001439.

Zhou:2024:TRP

[2073] Hongliang Zhou, Yifeng Zheng, and Xiaohua Jia. Towards robust and privacy-preserving federated learning in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001531.

Wu:2024:ABD

[2074] Yunyi Wu, Kai Zhang, and Yongbing Zhang. An accelerated Benders decomposition approach for virtual base station formation in stochastic Cloud-RANs. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001609.

Ruiz:2024:STU

[2075] Olga Ruiz, Juan Sánchez-González, Jordi Pérez-Romero, Oriol Sallent, and Irene Vilà. Space and time user distribution measurements dataset in a university campus. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001610.

Liu:2024:ECO

[2076] Liang Liu, Wuping Mao, Wenwei Li, Jie Duan, Guanyu Liu, and Bingchuan Guo. Edge computing offloading strategy for space-air-ground integrated network based on game theory. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001634.

Anonymous:2024:EBf

[2077] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001981.

Anonymous:2024:Ab

[2078] Anonymous. April 2024. Computer Networks (Amsterdam, Netherlands: 1999), 243(??):??, April 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Marin:2024:QML

[2079] Andrea Marin, Marco Ajmone Marsan, Michela Meo, and Matteo Sereno. Queuing models of links carrying streaming and elastic services. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001385.

Zhu:2024:ECC

[2080] Sifeng Zhu, Xiaohua Tian, Hao Chen, Hai Zhu, and Rui Qiao. Edge collaborative caching solution based on improved NSGA II algorithm in Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001397.

Bellin:2024:MBA

[2081] Arturo Bellin, Fabrizio Granelli, and Daniele Munaretto. A measurementbased approach to analyze the power consumption of the softwarized 5G core. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001440.

deAlmeida:2024:DDE

[2082] Leandro C. de Almeida, Washington Rodrigo Dias da Silva, Thiago C. Tavares, Rafael Pasquini, Chrysa Papagianni, and Fábio L. Verdi. DESiRED — Dynamic, Enhanced, and Smart iRED: a P4-AQM with deep reinforcement learning and in-band network telemetry. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001580.

Shirsath:2024:STS

[2083] Vaishali A. Shirsath, Madhav M. Chandane, Chhagan Lal, and Mauro Conti. SYNTROPY: TCP SYN DDoS attack detection for Software Defined Network based on Rényi entropy. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001592.

Toorchi:2024:DRL

[2084] Niloofar Toorchi, Weiqiang Lyu, Linsheng He, Jiamiao Zhao, Iftikhar Rasheed, and Fei Hu. Deep reinforcement learning enhanced skeleton based pipe routing for highthroughput transmission in flying adhoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001622.

Feraudo:2024:DDB

[2085] Angelo Feraudo, Nicolò Romandini, Carlo Mazzocca, Rebecca Montanari, and Paolo Bellavista. DIVA: a DID-based reputation system for secure transmission in VANETs using IOTA. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001646.

Zhao:2024:MSB

[2086] Baosen Zhao, Wanghong Yang, Wenji Du, Yongmao Ren, Jianan Sun, Qinghua Wu, and Xu Zhou. A multipath scheduler based on crosslayer information for low-delay applications in 5G edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001658.

delRio:2024:MGS

[2087] Alberto del Rio, Javier Serrano, David Jimenez, Luis M. Contreras, and Federico Alvarez. Multisite gaming streaming optimization over virtualized 5G environment using deep reinforcement learning techniques. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400166X.

Mlika:2024:BST

[2088] Fatma Mlika, Wafa Karoui, and Lotfi Ben Romdhane. Blockchain solutions for trustworthy decentralization in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001683.

Mortagua:2024:EAI

[2089] Duarte Mortágua, André Zúquete, and Paulo Salvador. Enhancing 802.1X authentication with identity providers using EAP-OAUTH and OAuth 2.0. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001695.

Liang:2024:PPM

[2090] Junjie Liang, Lei Zhang, Can Bu, Guangyu Yang, Hao Wu, and Aiguo Song. Plug-and-play multi-dimensional attention module for accurate Human Activity Recognition. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001701.

Du:2024:OTT

[2091] An Du, Jie Jia, Jian Chen, Liang Guo, and Xingwei Wang. Online twotimescale service placement for timesensitive applications in MEC-assisted network: a TMAGRL approach. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001713.

Baidas:2024:NSR

[2092] Mohammed W. Baidas, Ahmed M.

AbdelGhaffar, and Emad Alsusa. Network sum-rate maximization for network-coded clustered uplink NOMA networks with SWIPT-enabled relays. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624001725.

Faticanti:2024:OOC

[2093] Francescomaria Faticanti and Giovanni Neglia. Optimistic online caching for batched requests. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001737.

Li:2024:ORA

[2094] Yue Li, Yanjun Li, Yuzhe Chen, Jiahui Tong, Xianzhong Tian, and Kaikai Chi. Online resolution adaptation and resource allocation for edgeassisted video analytics. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001749.

Zhang:2024:TNT

[2095] Zixu Zhang, Guangsheng Yu, Caijun Sun, Xu Wang, Ying Wang, Ming Zhang, Wei Ni, Ren Ping Liu, Andrew Reeves, and Nektarios Georgalas. TbDd: a new trustbased, DRL-driven framework for blockchain sharding in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001750.

Wengle:2024:EAJ

[2096] Emil Wengle, Elias Strandell Erstorp, Viktor Lidström, Damiano Varagnolo, and Hefeng Dong. Experimental assessment of a JANUSbased consensus protocol. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001774.

Iqbal:2024:IQO

[2097] Muhammad Shahid Iqbal and Chien Chen. Instant queue occupancy used for automatic traffic scheduling in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001786.

Bilir:2024:HCH

[2098] Emin Bilir, Murtaza Cicioglu, and Ali Çalhan. Hybrid cell handover strategy for O-RAN-based campus networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001798.

Hosseini:2024:EEM

[2099] Soheil Hosseini, Ignacio de Miguel, Noemí Merayo, Ramón de la Rosa, Rubén M. Lorenzo, and Ramón J. Durán Barroso. Energy efficient multipath routing in space division multiplexed elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001816.

Hakiri:2024:CSD

[2100] Akram Hakiri, Aniruddha Gokhale, Sadok Ben Yahia, and Nedra Mellouli. A comprehensive survey on digital twin for future networks and emerging Internet of Things industry. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001828.

Neto:2024:FFL

[2101] Helio N. Cunha Neto, Jernej Hribar, Ivana Dusparic, Natalia C. Fernandes, and Diogo M. F. Mattos. FedSBS: Federated-learning participant-selection method for Intrusion Detection Systems. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400183X.

Kumar:2024:BEA

[2102] Neeraj Kumar and Rifaqat Ali. Blockchain-enabled authentication framework for Maritime Transportation System empowered by 6G-IoT. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001853.

Bingol:2024:WQD

[2103] Gülnaziye Bingöl, Simone Porcu, Alessandro Floris, and Luigi Atzori. WebRTC-QoE: a dataset of QoE assessment of subjective scores, network impairments, and facial and speech features. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624001889.

Anonymous:2024:EBg

[2104] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002433.

Anonymous:2024:Mb

[2105] Anonymous. May 2024. Computer Networks (Amsterdam, Netherlands: 1999), 244(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Garima:2024:DLB

[2106] Garima, Vivekanand Jha, and Rakesh Kumar Singh. A deep learning based dynamic bandwidth allocation method for XG-PON based mobile fronthaul for CRAN. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001762.

Kim:2024:HAO

[2107] Sungwook Kim. Hierarchical aerial offload computing algorithm based on the Stackelberg-evolutionary game model. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624001804.

Guo:2024:MCO

[2108] Yinzhi Guo, Xiaolong Xu, and Fu Xiao.

MADRLOM: a computation offloading mechanism for software-defined cloudedge computing power network. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001841.

Peng:2024:LLI

[2109] Min Peng, Xianxin Fu, Haiyang Zhao, Yu Wang, and Caihong Kai. LiKey: Location-independent keystroke recognition on numeric keypads using WiFi signal. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001865.

Guo:2024:CEL

[2110] Feng Guo, Shidong Sun, Junjie Hu, Ning Zhang, and Zhiqiang Lv. CIPO: Efficient, lightweight and programmable packet scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001877.

Zhang:2024:RDT

[2111] Xuhui Zhang, Qing Li, Feixue Han, and Yong Jiang. A receiver-driven transport protocol using differentiated algorithms for differential congestion in datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001890.

Kazmi:2024:SFL

[2112] Syed Hussain Ali Kazmi, Faizan Qamar, Rosilah Hassan, Kashif Nisar, and Mohammed Azmi Al-Betar. Security of federated learning in 6G era: a review on conceptual techniques and software platforms used for research and analysis. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001907.

Deng:2024:TOO

[2113] Tao Deng, Zhanwei Yu, and Di Yuan.

Task offloading optimization in mobile edge computing under uncertain processing cycles and intermittent communications. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001919.

Hu:2024:RSA

[2114] Jianwei Hu, Kaiqi Huang, Genqing Bian, and Yanpeng Cui. Redact4Trace: a solution for auditing the data and tracing the users in the redactable blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001920.

Wilhelmi:2024:IDB

[2115] Francesc Wilhelmi, Nima Afraz, Elia Guerra, and Paolo Dini. The implications of decentralization in blockchained federated learning: Evaluating the impact of model staleness and inconsistencies. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001932.

Dao:2024:RNT

[2116] Nhu-Ngoc Dao, Ngo Hoang Tu, Trong-Dai Hoang, Tri-Hai Nguyen, Luong Vuong Nguyen, Kyungchun Lee, Laihyuk Park, Woongsoo Na, and Sungrae Cho. A review on new technologies in 3GPP standards for 5G access and beyond. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002020.

Centofanti:2024:IPC

[2117] Carlo Centofanti, José Santos, Venkateswarlu Gudepu, and Koteswararao Kondepu. Impact of power consumption in containerized clouds: a comprehensive analysis of open-source power measurement tools. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002032.

Han:2024:GDE

[2118] Xinbo Han, Guizhong Xu, Meng Zhang, Zheng Yang, Ziyang Yu, Weiqing Huang, and Chen Meng. DE-GNN: Dual embedding with graph neural network for fine-grained encrypted traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002044.

Pei:2024:PRE

[2119] Hongmei Pei, Peng Yang, Weihao Li, Miao Du, and Zhongjian Hu. Proxy re-encryption for secure data sharing with blockchain in Internet of Medical Things. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002056.

Cerasuolo:2024:MNA

[2120] Francesco Cerasuolo, Alfredo Nascita, Giampaolo Bovenzi, Giuseppe Aceto, Domenico Ciuonzo, Antonio Pescapè, and Dario Rossi. MEMENTO: a novel approach for class incremental learning of encrypted traffic. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002068.

Liu:2024:MTE

[2121] Libin Liu and Xiuting Xu. Marvel: Towards efficient federated learning on IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400207X.

Zhang:2024:ABA

[2122] Jingci Zhang, Jun Zheng, Zheng Zhang, Tian Chen, Yu an Tan, Quanxin Zhang, and Yuanzhang Li. ATT&CK-based Advanced Persistent Threat attacks risk propagation assessment model for zero trust networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002081.

Stergiopoulos:2024:SPP

[2123] George Stergiopoulos, Panagiotis Kotzanikolaou, Konstantinos Adamos, and Lilian Mitrou. Scaling private proximity testing protocols for geofenced information exchange: a metropolitan-wide case study. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002135.

Ge:2024:HIP

[2124] Fei Ge and Liansheng Tan. HITLinQ: Improving path capacity through hybrid information-theoretic link scheduling in multi-hop wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624002147.

Zhang:2024:RPP

[2125] Fan Zhang, Hui Huang, Zhixiong Chen, and Zhenjie Huang. Robust and privacy-preserving federated learning with distributed additive encryption against poisoning attacks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002159.

Zorello:2024:BBO

[2126] Ligia Maria Moreira Zorello, Laurens Bliek, Sebastian Troia, Guido Maier, and Sicco Verwer. Black-box optimization for anticipated baseband-function placement in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002160.

Li:2024:MNA

[2127] Beibei Li and Wei Hu. A multinode attack scheme based on community partitioning in large scale infrastructure networks. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002184.

Han:2024:CAU

[2128] Xiaoli Han and Bo Hu. A congestionaware user-cell association scheme to assist with traffic offloading in a threetier heterogeneous network. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002202.

Liu:2024:AMA

[2129] Xiaoying Liu, Xiaodong Li, Kechen Zheng, and Jia Liu. AoI minimization of ambient backscatterassisted EH-CRN with cooperative spectrum sensing. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002214.

Wang:2024:RLB

[2130] Yi Wang, Yu Deng, Ling Kang, Fang Jiang, and Fulin Jiang. Reinforcement learning-based energy efficiency optimization for RIS-assisted UAV hybrid uplink and downlink system. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002226.

Kvasic:2024:DRC

[2131] Igor Kvasić, Derek Orbaugh Antillon, – Dula Na– d, Christopher Walker, Iain Anderson, and Nikola Mišković. Diver-robot communication dataset for underwater hand gesture recognition. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S138912862400224X.

Li:2024:QQR

[2132] Junhao Li, Qiang Nong, and Ziyu Liu. QBDD: Quantum-resistant blockchain-assisted deep data deduplication protocol for vehicular crowdsensing system. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002251.

Wang:2024:SIB

[2133] Kaiyu Wang, Guangsheng Feng, Zhenzhou Ji, Zhiying Tu, and Shufan He. SP-PoR: Improve blockchain performance by semi-parallel processing transactions. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002263.

Rashid:2024:PUD

[2134] Fariza Rashid, Ben Doyle, Soyeon Caren Han, and Suranga Seneviratne. Phishing URL detection generalisation using Unsupervised Domain Adaptation. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624002305.

Xie:2024:OTO

[2135] Mande Xie, Xiangquan Su, Hao Sun, and Guoping Zhang. Online task offloading algorithm based on multiobjective optimization caching strategy. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624002329.

Lavernelle:2024:ASI

[2136] Jean de Bonfils Lavernelle, Pierre-François Bonnefoi, Benoît Gonzalvo, and Damien Sauveron. Assessment of spatial isolation in Jailhouse: Towards a generic approach. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002342.

Wang:2024:CII

[2137] Xiangbin Wang, Qingjun Yuan, Yongjuan Wang, Gaopeng Gou, Chunxiang Gu, Gang Yu, and Gang Xiong. Combine intra- and interflow: a multimodal encrypted traffic classification model driven by diverse features. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002354.

Feng:2024:GGL

[2138] Pengbin Feng, Dawei Wei, Qiaoyang Li, Qin Wang, Youbing Hu, Ning Xi, and Jianfeng Ma. GlareShell: Graph learning-based PHP webshell detection for web server of industrial internet. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400238X.

Anonymous:2024:EBh

[2139] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002718.

Anonymous:2024:Mc

[2140] Anonymous. May 2024. Computer Networks (Amsterdam, Netherlands: 1999), 245(??):??, May 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Umar:2024:CON

[2141] Amara Umar, Syed Ali Hassan, Haejoon Jung, Sahil Garg, M. Shamim Hossain, and Mohsen Guizani. Computation offloading in NOMA-MEC-enabled aerial-vehicular networks exploiting mmWave capabilities. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001671.

Zhang:2024:IAE

[2142] Shuang Zhang, Huilong Jin, and Pingkang Guo. IRS-assisted energy efficient communication for UAV mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002196.

Awais:2024:PSF

[2143] Syed Muhammad Awais, Wu Yucheng, Khalid Mahmood, Hafiz Muhammad Sanaullah Badar, Rupak Kharel, and Ashok Kumar Das. Provably secure fog-based authentication protocol for VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002238.

Chu:2024:OOS

[2144] Weibo Chu, Xiaoyan Zhang, Xinming Jia, John C. S. Lui, and Zhiyong Wang. Online optimal service caching for multi-access edge computing: a constrained Multi-Armed Bandit optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002275.

Hejazi:2024:MEC

[2145] Hamdan Hejazi and László Bokor.

Modeling and evaluation of cooperative vulnerable road user protection schemes in realistic C-ITS environments. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002287.

Pittala:2024:LDP

[2146] Gaetano Francesco Pittalà, Lorenzo Rinieri, Amir Al Sadi, Gianluca Davoli, Andrea Melis, Marco Prandini, and Walter Cerroni. Leveraging data plane programmability to enhance service orchestration at the edge: a focus on industrial security. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002299.

Jin:2024:CAS

[2147] Jiahui Jin, Zhendong Guo, Wenchao Bai, Biwei Wu, Xiang Liu, and Weiwei Wu. Congestion-aware Stackelberg pricing game in urban Internet-of-Things networks: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002378.

Li:2024:SCD

[2148] Zhengnan Li, Diego A. Cuji, and Milica Stojanovic. Space-code division multiple access for broadband acoustic networks. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002391.

Ai:2024:JOC

[2149] Pengjie Ai and Fei Wang. Joint optimization for computation offloading and 3C resource allocations over wireless-powered and NOMA-enabled multi-access MEC. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002470.

Liu:2024:LLV

[2150] Zhiwei Liu, Zhaogang Shu, Shuwu Chen, Yiwen Zhong, and Jiaxiang Lin. Low-latency virtual network function scheduling algorithm based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002500.

Wang:2024:DDR

[2151] Shaocong Wang, Xiaoyun Zhang, Changhong Wang, Ke Wu, Cunlu Li, and Dezun Dong. DRLAR: a deep reinforcement learning-based adaptive routing framework for networkon-chips. Computer Networks (Amsterdam, Netherlands: 1999), 246 (??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002512.

Silva:2024:MBI

[2152] Felipe S. Dantas Silva, Sérgio N.

Silva, Lucileide M. D. da Silva, Ayuri Bessa, Samuel Ferino, Pablo Paiva, Marcos Medeiros, Lucas Silva, José Neto, Kevin Costa, Charles Santos, Eduardo Aranha, Allan Martins, Uirá Kulesza, Roger Immich, Augusto V. Neto, Ramon Fontes, Vicente Sousa, and Marcelo A. C. Fernandes. ML-based inter-slice load balancing control for proactive offloading of virtual services. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002548.

Xiang:2024:ABK

[2153] Axin Xiang, Hongfeng Gao, Youliang Tian, Linjie Wang, and Jinbo Xiong. Attribute-based key management for patient-centric and trusted data access in blockchain-enabled IoMT. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002573.

Hotescu:2024:TTS

[2154] Oana Hotescu, Katia Jaffrès-Runser, and Jean-Luc Scharbarg. Timetriggered scheduling of mixed-critical flows at end-system in asynchronous AFDX avionic network. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002597.

Chu:2024:IRA

[2155] Weibo Chu, Xinming Jia, Zhiwen Yu, John C. S. Lui, and Yi Lin. On incentivizing resource allocation and task offloading for cooperative edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002603.

Anonymous:2024:EBi

[2156] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003189.

Anonymous:2024:Jb

[2157] Anonymous. June 2024. Computer Networks (Amsterdam, Netherlands: 1999), 246(??):??, June 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zahedi:2024:HMV

[2158] Seyed Reza Zahedi and Shahram Jamali. A hybrid model for VNF deployment capable of responding to online requests at network edge. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002172.

Luan:2024:MWM

[2159] Zeyu Luan, Qing Li, Yong Jiang, Jingpu Duan, Ruobin Zheng, Dingding Chen, and Shaoteng Liu. MATE: When multi-agent Deep Reinforcement Learning meets Traffic Engineering in multi-domain networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002317.

Regan:2024:BDP

[2160] R. Regan, R. Josphineleela, Mohammad Khamruddin, and R. Vijay. Balancing data privacy and sharing in IIoT: Introducing the GFL-LFF aggregation algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002330.

Yu:2024:MCC

[2161] Zhanwei Yu, Tao Deng, Yi Zhao, and Di Yuan. Multi-cell content caching: Optimization for cost and information freshness. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002524.

Tao:2024:DPL

[2162] Yuan Tao, Taochun Wang, Yong Qiang, Leilei Shen, Fulong Chen, and Chuanxin Zhao. A differential privacy location protect approach with intelligence data collection paradigm for MCS. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002536.

Chen:2024:NAD

[2163] Wenqi Chen, Zhiliang Wang, Liyuan Chang, Kai Wang, Ying Zhong, Dongqi Han, Chenxin Duan, Xia Yin, Jiahai Yang, and Xingang Shi. Network anomaly detection via similarity-aware ensemble learning with ADSim. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400255X.

Li:2024:AAA

[2164] Xianglong Li, Kaiwei Mo, Yeqiao Hou, Zongpeng Li, Hong Xu, and Chun Jason Xue. An auction approach to aircraft bandwidth scheduling in non-terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002561.

Li:2024:PFP

[2165] Sensen Li, Yicai Huang, and Bin Yu. A practical and flexible PUFbased end-to-end anonymous authentication protocol for IoT. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002585.

Wang:2024:STM

[2166] Xiaoyu Wang, Sun Xu, and Yangming Zhao. SARS: Towards minimizing average Coflow Completion time in MapReduce systems. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002615.

Montanari:2024:AMA

[2167] Antonio Montanari, Filippo Campagnaro, and Michele Zorzi. An adaptive MAC-agnostic ranging scheme for underwater acoustic mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 247 (??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002627.

Chaurasia:2024:DIE

[2168] Bhavana Chaurasia, Anshul Verma, and Pradeepika Verma. An in-depth and insightful exploration of failure detection in distributed systems. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002640.

Gorshenin:2024:MNT

[2169] Andrey Gorshenin, Anastasia Kozlovskaya, Sergey Gorbunov, and Irina Kochetkova. Mobile network traffic analysis based on probability-informed machine learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002652.

Moreira:2024:ERL

[2170] Cristiano L. Moreira, Carlos A. Kamienski, and Reinaldo A. C. Bianchi. 5G and edge: a reinforcement learning approach for Virtual Network Embedding with cost optimization and improved acceptance rate. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002664.

Zhang:2024:MTP

[2171] Jin Zhang, Daoye Wang, Kai Liu, and Jianhua Lu. Multi-Table Programmable Parser with online flowlevel update consistency for satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002676.

Centofanti:2024:TLE

[2172] Carlo Centofanti, Walter Tiberti, Andrea Marotta, Fabio Graziosi, and Dajana Cassioli. Taming latency at the edge: a user-aware service placement approach. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002767.

Tsourdinis:2024:SAR

[2173] Theodoros Tsourdinis, Ilias Chatzistefanidis, Nikos Makris, Thanasis Korakis, Navid Nikaein, and Serge Fdida. Service-aware real-time slicing for virtualized beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002779.

AlHidaifi:2024:TCR

[2174] Saleh Mohamed AlHidaifi, Muhammad Rizwan Asghar, and Imran Shafique Ansari. Towards a Cyber Resilience Quantification Framework (CRQF) for IT infrastructure. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002780.

Tanveer:2024:SIS

[2175] Muhammad Tanveer, Abdallah Aldosary, Neeraj Kumar, and Saud Alhajaj Aldossari. SEAF-IoD: Secure and efficient user authentication framework for the Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002810.

He:2024:NTF

[2176] Kaiyan He and Zhe Ren. A new three-factor authentication scheme using Chebyshev chaotic map for peer-to-peer Industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 247 (??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002822.

Fan:2024:NMS

[2177] Huilong Fan, Chongxiang Sun, Jun Long, Shangpeng Wang, and Fei Zeng. A novel method for solving the multi-commodity flow problem on evolving networks. Computer Networks (Amsterdam, Netherlands:

1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002834.

Jerabek:2024:CAD

[2178] Kamil Jerabek, Karel Hynek, and Ondrej Rysavy. Comparative analysis of DNS over HTTPS detectors. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002846.

Yan:2024:PRC

[2179] Dingyu Yan, Yaping Liu, Shuo Zhang, Binxing Fang, Feng Zhao, and Zhikai Yang. PCNP: a RoCEv2 congestion control using precise CNP. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002858.

Yungaicela-Naula:2024:MRA

[2180] Noe M. Yungaicela-Naula, Vishal Sharma, and Sandra Scott-Hayward. Misconfiguration in O-RAN: Analysis of the impact of AI/ML. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002871.

Hao:2024:AFR

[2181] Jialin Hao, Rola Naja, and Djamal Zeghlache. Adaptive federated reinforcement learning for critical realtime communications in UAV assisted vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002883.

Lu:2024:DDC

[2182] Yifei Lu, Xu Ma, and Changjiang Cui. DCCS: a dual congestion control signals based TCP for datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002895.

Xia:2024:CFD

[2183] Dongyu Xia, Jiayi Liu, Chen Wang, and Weihua Wu. Comprehensive fault diagnosis for multiple coupled SFCs based on deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002901.

Bhandari:2024:LOC

[2184] Abhay Bhandari, Akhil Gupta, Sudeep Tanwar, Joel J. P. C. Rodrigues, Ravi Sharma, and Anupam Singh. Latency optimized C-RAN in optical backhaul and RF fronthaul architecture for beyond 5G network: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002913.

Jebri:2024:SAI

[2185] Sarra Jebri, Arij Ben Amor, and Salah Zidi. A seamless authentication for intra and inter metaverse platforms using blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002925.

Djaidja:2024:TED

[2186] Taki Eddine Toufik Djaidja, Bouziane Brik, Abdelwahab Boualouache, Sidi Mohammed Senouci, and Yacine Ghamri-Doudane. Time-efficient detection of false position attack in 5G and beyond vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002937.

Lee:2024:MDM

[2187] Ki-Hun Lee, Seungmin Lee, Jaedon Park, Howon Lee, and Bang Chul Jung. MUSCAT: Distributed multiagent Q-learning-based minimum span channel allocation technique for UAVenabled wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002949.

Liu:2024:EAF

[2188] Jindian Liu, Zhuo Li, Huipeng Du, Haodong Zhou, Leyang Li, Yi An, Yu Zhang, Kaihua Liu, and Qiang Li. An effective and accurate flow size measurement using funnelshaped sketch. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624002998.

Pang:2024:AVI

[2189] Zengyu Pang, Ligen Shi, and Hua Xiang. Asynchronous verifiable information dispersal protocol of achieving optimal communication. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003025.

Liu:2024:EFL

[2190] Yuanyuan Liu, Kexin Chen, and Lu Zhu. Efficient federated learning algorithm using sparse ternary compression based on layer variation classification. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003037.

Salim:2024:PPS

[2191] Mikail Mohammed Salim, Laurence Tianruo Yang, and Jong Hyuk Park. Privacy-preserving and scalable federated blockchain scheme for healthcare 4.0. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003049.

Yu:2024:TFE

[2192] Linxiao Yu, Jun Tao, Yifan Xu, Weice Sun, and Zuyan Wang. TLS fingerprint for encrypted malicious traffic detection with attributed graph kernel. Computer Networks (Amsterdam, Netherlands: 1999), 247 (??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003074.

Anonymous:2024:EBj

[2193] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003542.

Anonymous:2024:Jc

[2194] Anonymous. June 2024. Computer Networks (Amsterdam, Netherlands: 1999), 247(??):??, June 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Konduru:2024:SEE

[2195] Prasanthi Konduru and N. P. Nethravathi. SECURE and energy-efficient routing protocol based on microsegmentation and batch authentication. Computer Networks (Amsterdam, Netherlands: 1999), 248 (??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624001257.

Segata:2024:CFS

[2196] Michele Segata, Paolo Casari, Marios Lestas, Alexandros Papadopoulos, Dimitrios Tyrovolas, Taqwa Saeed, George Karagiannidis, and Christos Liaskos. CoopeRIS: a framework for the simulation of reconfigurable intelligent surfaces in cooperative driving environments. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002755.

Rahmatov:2024:ESF

[2197] Nematullo Rahmatov and Hoki Baek. Exploring scale-free networks: Survey on autonomous system dynamics and connectivity analysis. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400286X.

Filho:2024:SRP

[2198] Joao da Mata Liborio Filho, Jhonathan Oliveira, and Cesar A. V. Melo. Super-resolution with perceptual quality for improved live streaming delivery on edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002950.

Peng:2024:VBD

[2199] Jiaxin Peng, Siwang Zhou, Liubo Ouyang, and Xingting Liu. Volatilitybased diversity awareness for distributed data storage of mobile crowd sensing. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002986.

Tian:2024:OPM

[2200] Shu-Juan Tian, Ke-Ke Xu, Wen-Jian Ding, Yan-Chun Li, and De-Ze Zeng. An offloading and pricing mechanism based on virtualization in edge-cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003001.

Fan:2024:MAD

[2201] Chenchen Fan, Hongyu Xu, and Qingling Wang. Multi-agent deep reinforcement learning for trajectory planning in UAVs-assisted mobile edge computing with heterogeneous requirements. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003013.

Kazemi:2024:IBP

[2202] Nader Kazemi, Reza Ghaderi, and Soheila Nazari. An IoT-based packet aggregation mechanism for the SDNbased wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003062.

Toure:2024:FDZ

[2203] Almamy Touré, Youcef Imine, Alexis Semnont, Thierry Delot, and Antoine Gallais. A framework for detecting zero-day exploits in network flows. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003086.

Wang:2024:LBG

[2204] Jiaqin Wang, Kai Liu, and Hantao Li. LSTM-based graph attention network for vehicle trajectory prediction. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003098.

Wei:2024:PDP

[2205] Xingchen Wei, Laixian Peng, Renhui Xu, Aijing Li, Xingyue Yu, and Hai Wang. 3D position deployment and performance optimization of mmWave UAV-assisted HetIoT under jamming condition. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003104.

Martin:2024:SDC

[2206] Raul Martin, Ivan Vidal, and Francisco Valera. A software-defined connectivity service for multi-cluster cloud native applications. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003116.

Alexandri:2024:DCS

[2207] Talmon Alexandri and Roee Diamant. Detection and characterization of ship underwater radiated narrowband noise. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003128.

Nithya:2024:DDA

[2208] D. Nithya, V Madhusudanan, B. S. N. Murthy, R. Geetha, Nguyen Xuan Mung, Nhu-Ngoc Dao, and Sungrae Cho. Delayed dynamics analysis of SEI2RS malware propagation models in cyber-physical systems. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400313X.

Osman:2024:IMT

[2209] Radwa Ahmed Osman. Internet of Medical Things (IoMT) optimization for healthcare: a deep learning-based interference avoidance model. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003232.

Chen:2024:DSP

[2210] Zhuoxing Chen, Yiqin Lu, Haihan Wang, Jiancheng Qin, Meng Wang, and Weiqiang Pan. Dynamic stream partitioning for time-triggered traffic in Time-Sensitive Networking. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003244.

Zheng:2024:TRO

[2211] Danyang Zheng, He Fang, Shaohua Cao, Yihan Zhong, and Xiaojun Cao. Towards resources optimization in deploying service function chains with shared protection. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003268.

Xu:2024:ASS

[2212] Renjie Xu, Guangwei Wu, Weiping Wang, Xing Gao, An He, and Zhengpeng Zhang. Applying selfsupervised learning to network intrusion detection for network flows with graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400327X.

Dursun:2024:WPN

[2213] Yunus Dursun, Suhaib Al Basit, and Zhiguo Ding. Wireless powered NOMA-based cognitive radio for 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003293.

Garcia-Rodriguez:2024:BSD

[2214] Jesús García-Rodríguez, Stephan Krenn, Jorge Bernal Bernabe, and Antonio Skarmeta. Beyond selective disclosure: Extending distributed p-ABC implementations by commitand-prove techniques. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400330X.

Zhao:2024:TTS

[2215] Liushun Zhao, Deke Guo, Lailong Luo, Junjie Xie, Yulong Shen, and Bangbang Ren. Tiger Tally: a secure IoT data management approach based on redactable blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003323.

Rathee:2024:TMS

[2216] Geetanjali Rathee, Akshay Kumar, Chaker Abdelaziz Kerrache, and Carlos T. Calafate. A trust management solution for 5G-based future generation Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003335.

Li:2024:SFR

[2217] Zhuo Li, Shuaijun Liu, Jindian Liu, Yu Zhang, Teng Liang, and Kaihua Liu. SIM: a fast real-time graph stream summarization with improved memory efficiency and accuracy. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003347.

Rolich:2024:UIP

[2218] Alexey Rolich, Ion Turcanu, Alexey Vinel, and Andrea Baiocchi. Understanding the impact of persistence and propagation on the age of information of broadcast traffic in 5G NR-V2X sidelink communications. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003359.

Imanaka:2024:PTS

[2219] Shoya Imanaka, Akio Kawabata, Bijoy Chand Chatterjee, and Eiji Oki. Polynomial-time server allocation algorithm in delay-sensitive Internet-of-Things monitoring services. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003360.

Anonymous:2024:EBk

[2220] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003815.

Anonymous:2024:Jd

[2221] Anonymous. June 2024. Computer Networks (Amsterdam, Netherlands: 1999), 248(??):??, June 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2024:BAS

[2222] Jitao Wang, Bo Zhang, Kai Wang, Yuzhou Wang, and Weili Han. BFT-Diagnosis: an automated security testing framework with malicious behavior injection for BFT protocols. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002366.

Wang:2024:BEA

[2223] Chuanhua Wang, Quan Zhang, Xin Xu, and Huimin Wang. AS-T^{3}BP: an efficient assignment scheme for space TT&C tasks with bidirectional privacy-preserving under blockchain architecture. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002962.

Xu:2024:ESE

[2224] Yihang Xu, Yuxing Mao, Jian Li, Xueshuo Chen, and Shunxin Wu. Edge server enhanced secure and privacy preserving federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624002974.

Li:2024:HDE

[2225] Yang Li, Chunhe Xia, Chunyan Li, Yuan Zhao, Chen Chen, and Tianbo Wang. HL-DPoS: an enhanced anti-long-range attack DPoS algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003050.

Lian:2024:LCC

[2226] Linming Lian, Ningjiang Chen, Xuemei Yuan, and Jianbo Lu. Lowcomplexity collaborative caching strategy based on spatio-temporal graph convolutional model. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003220.

Mezina:2024:RUS

[2227] Anzhelika Mezina, Radim Burget, and Aleksandr Ometov. Reinterpreting usability of semantic segmentation approach for Darknet traffic analysis. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003256.

Chen:2024:CEF

[2228] Weihan Chen, Zhiliang Wang, Han Zhang, Xia Yin, and Xingang Shi. Cost-efficient flow migration for SFC dynamical scheduling in geo-distributed clouds. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003281.

dOrey:2024:ASR

[2229] Pedro M. d’Orey, Miguel Gutiérrez Gaitán, Pedro M. Santos, Manuel Ribeiro, João Borges de Sousa, and Luís Almeida. Assessing shortrange Shore-to-Shore (S2S) and Shore-to-Vessel (S2V) WiFi communications. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003372.

Shaik:2024:AMA

[2230] Razeena Begum Shaik, Prabagarane Nagaradjane, Iacovos Ioannou, Vitawat Sittakul, Vasos Vasiliou, and Andreas Pitsillides. AI/ML-aided capacity maximization strategies for URLLC in 5G/ 6G wireless systems: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003384.

Song:2024:DDL

[2231] Tailai Song, Paolo Garza, Michela Meo, and Maurizio M. Munafò. DeX: Deep learning-based throughput prediction for real-time communications with emphasis on traffic eXtremes. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003396.

Cao:2024:FLS

[2232] Manxia Cao, Qingshan Wang, and Qi Wang. Federated learning in smart home: a dynamic contractbased incentive approach with task preferences. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003426.

Yadav:2024:ACP

[2233] Rashmi Yadav, Rashmi Kamran, Pranav Jha, and Abhay Karandikar. An architecture for control plane slicing in beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003438.

deLima:2024:CFA

[2234] Leandro Alvarez de Lima and Gustavo Sousa Pavani. Crosstalk- and fragmentation-aware survivable routing, modulation, space, and spectrum assignment using ant colony optimization. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003451.

Lin:2024:JRC

[2235] Na Lin, Jinjiao Huang, Ammar Hawbani, Liang Zhao, Hailun Tang, Yunchong Guan, and Yunhe Sun. Joint routing and computation offloading based deep reinforcement learning for Flying Ad hoc Networks. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003463.

Fan:2024:BLG

[2236] Wentao Fan, Jun Yao, Shiyuan Cui, Yan Wang, Shuo Xu, Yuehui Tan, Fan Yang, and Weihong Wu. Bi-LSTM/ GRU-based anomaly diagnosis for virtual network function instance. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003475.

Su:2024:CCA

[2237] Dongyuan Su, Yipeng Zhou, Laizhong Cui, and Quan Z. Sheng. Communication cost-aware client selection in online federated learning: a Lyapunov approach. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003499.

Huso:2024:DIL

[2238] Ingrid Huso, Marco Olivieri, Leonardo Galgano, Adnan Rashid, Giuseppe Piro, and Gennaro Boggia. Design and implementation of a lookingforward Lawful Interception architecture for future mobile communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003505.

Sun:2024:MKR

[2239] Jin Sun, Wenjuan Su, Lu Wang, and Kexin Ye. Multi-keyword ranked search scheme for privacy protection in social networks. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003669.

Anonymous:2024:EBl

[2240] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004183.

Anonymous:2024:Je

[2241] Anonymous. July 2024. Computer Networks (Amsterdam, Netherlands: 1999), 249(??):??, July 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Charpentier:2024:PWT

[2242] Vincent Charpentier, Nina Slamnik-Kriještorac, Giada Landi, Matthias Caenepeel, Olivier Vasseur, and Johann M. Marquez-Barja. Paving the way towards safer and more efficient maritime industry with 5G and beyond edge computing systems. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003311.

Bakar:2024:FNH

[2243] Rana Abu Bakar, Lorenzo De Marinis, Filippo Cugini, and Francesco Paolucci. FTG-Net-E: a hierarchical ensemble graph neural network for DDoS attack detection. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003402.

Wang:2024:SWH

[2244] Junjie Wang, Michael Bewong, and Lihong Zheng. SD-WAN: Hybrid edge cloud network between multisite SDDC. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003414.

Zhu:2024:CSF

[2245] Konglin Zhu, Fuchun Zhang, Lei Jiao, Bowei Xue, and Lin Zhang. Client selection for federated learning using combinatorial multi-armed bandit under long-term energy constraint. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400344X.

Farrea:2024:PSO

[2246] Kerry Anne Farrea, Zubair Baig, Robin Ram Mohan Doss, and Dongxi Liu. Provably secure optimal homomorphic signcryption for satellitebased Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003487.

Wang:2024:LAC

[2247] Qianxing Wang, Wei Li, and Amin Mohajer. Load-aware continuous-time optimization for multi-agent systems: toward dynamic resource allocation and real-time adaptability. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400358X.

ElBouanani:2024:FAL

[2248] Houssam ElBouanani, Chadi Barakat, Walid Dabbous, and Thierry Turletti.

Fidelity-aware large-scale distributed network emulation. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003633.

Jiao:2024:SCR

[2249] Jingsen Jiao, Ranran Sun, Yizhi Cao, Qifeng Miao, Yanchun Zuo, and Weidong Yang. Study on covert rate in the D2D networks with multiple non-colluding wardens. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003645.

Bhattacharyya:2024:EED

[2250] Abhishek Bhattacharyya, Shunmugapriya Ramanathan, Andrea Fumagalli, and Koteswararao Kondepu. An end-to-end DPDK-integrated opensource 5G standalone Radio Access Network: a proof of concept. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003657.

Rashid:2024:NNS

[2251] Adnan Rashid and Tommaso Pecorella. Is 6LoWPAN-ND necessary? (Spoiler alert: Yes). Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003670.

Chen:2024:APC

[2252] Huiling Chen, Chunmei Zhang, and Zeyan Zhang. Adaptive pinning control to topology identification of output stochastic complex dynamical network. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003682.

Raza:2024:LGB

[2253] Arif Raza, Salabat Khan, Shivanshu Shrivastava, M. Wasim Abbas Ashraf, Ting Wang, Kaishun Wu, and Lu Wang. A lightweight group-based SDN-driven encryption protocol for smart home IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003694.

Simionato:2024:SIH

[2254] Giada Simionato and Mario G. C. A.

Cimino. Swarm intelligence for hole detection and healing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003700.

Stepanov:2024:FTA

[2255] E. P. Stepanov, R. L. Smeliansky, A. V. Plakunov, A. V. Borisov, Xia Zhu, Jianing Pei, and Zhen Yao. On fair traffic allocation and efficient utilization of network resources based on MARL. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003724.

Guo:2024:SSE

[2256] Xingyu Guo, Guo Chen, Xiaoning Zhan, Ting Qu, and Zhaojiao Han. SECM: Securely and efficiently connections setup using RDMA-CM. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003736.

Wu:2024:SCA

[2257] Caihong Wu and Jihua Liu. Smart contract assisted secure aggregation scheme for model update in federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003748.

Voutsas:2024:MAF

[2258] Fotios Voutsas, John Violos, and Aris Leivadeas. Mitigating alert fatigue in cloud monitoring systems: a machine learning perspective. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400375X.

Haddad:2024:EPS

[2259] Zaher Haddad. Enhancing privacy and security in 5G networks with an anonymous handover protocol based on blockchain and Zero Knowledge Proof. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003761.

Giarre:2024:EES

[2260] Federico Giarrè and Yassine Hadjadj-Aoul. Exploring the effectiveness of service migration strategies for virtual network embedding. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003852.

Kumar:2024:ISL

[2261] Vipin Kumar, Rifaqat Ali, and Pawan Kumar Sharma. IoEPM+: a secured and lightweight 6g-enabled pollution monitoring authentication framework using IoT and blockchain technology. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003864.

Pang:2024:ITO

[2262] Shanchen Pang, Teng Wang, Haiyuan Gui, Xiao He, and Lili Hou. An intelligent task offloading method based on multi-agent deep reinforcement learning in ultra-dense heterogeneous network with mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003876.

Cao:2024:GUM

[2263] Jie Cao, Yuwei Xu, Enze Yu, Qiao Xiang, Kehui Song, Liang He, and Guang Cheng. GateKeeper: an UltraLite malicious traffic identification method with dual-aspect optimization strategies on IoT gateways. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003888.

Busacca:2024:CAP

[2264] F. Busacca, L. Galluccio, S. Palazzo, and A. Panebianco. A comparative analysis of predictive channel models for real shallow water environments. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400389X.

Shen:2024:LBM

[2265] Xiajiong Shen, Xiaoran Li, Hongjian Yin, Chaoyang Cao, and Lei Zhang. Lattice-based multi-authority ciphertextpolicy attribute-based searchable encryption with attribute revocation for cloud storage. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003918.

Chen:2024:PUE

[2266] Rui Chen, Xiaoyu Chen, and Jing Zhao. Private and utility enhanced intrusion detection based on attack behavior analysis with local differential privacy on IoV. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400392X.

Dong:2024:OOA

[2267] Yongpeng Dong, Shiyou Qian, Wanghua Shi, Junshen Li, Jian Cao, and Guangtao Xue. OEM: an operation-aware event matching algorithm for contentbased Pub/Sub systems. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003931.

Xing:2024:TPP

[2268] Ling Xing, Bing Li, Lulu Liu, Yuanhao Huang, Honghai Wu, Huahong Ma, and Xiaohui Zhang. Trajectory privacy protection method based on sensitive semantic location replacement. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003943.

Li:2024:MIU

[2269] Zhaoxuan Li, Ziming Zhao, Rui Zhang, Haoyang Lu, Wenhao Li, Fan Zhang, Siqi Lu, and Rui Xue. metaNet: Interpretable unknown mobile malware identification with a novel metafeatures mining algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003955.

Xie:2024:DRL

[2270] Mande Xie, Jiefeng Ye, Guoping Zhang, and Xueping Ni. Deep reinforcement learning-based computation offloading and distributed edge service caching for Mobile Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003967.

Adanza:2024:ETF

[2271] Daniel Adanza, Lluis Gifre, Pol Alemany, Juan-Pedro Fernández-Palacios, Oscar González de Dios, Raul Muñoz, and Ricard Vilalta. Enabling traffic forecasting with cloud-native SDN controller in transport networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003979.

Pan:2024:ANC

[2272] Chengsheng Pan, Xiaosong Cui, Chen Zhao, Yingzhi Wang, and Yuyue Wang. An adaptive network congestion control strategy based on the change trend of average queue length. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003980.

Changazi:2024:ONT

[2273] Sabir Ali Changazi, Asim Dilawar Bakhshi, Muhammad Yousaf, Syed Muhammad Mohsin, Syed Muhammad Abrar Akber, Mohammed Abazeed, and Mohammed Ali. Optimization of network topology robustness in IoTs: a systematic review. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004006.

Shaabanzadeh:2024:STP

[2274] Seyedeh Soheila Shaabanzadeh and Juan Sánchez-González. A spatiotemporal prediction methodology based on deep learning and real Wi-Fi measurements. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004018.

Garcia-Penas:2024:HDA

[2275] Rodolfo García-Peñas, Rafael A. Rodríguez-Gómez, and Gabriel Maciá-Fernández. HoDiNT: Distributed architecture for collection and analysis of Internet Background Radiation. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400402X.

Vieira:2024:MAS

[2276] Juan Lucas Vieira, Evandro L. C. Macedo, Anselmo L. E. Battisti, Julia Noce, Paulo F. Pires, Débora C. Muchaluat-Saade, Ana C. B. Oliveira, and Flavia C. Delicato. Mobilityaware SFC migration in dynamic 5G-Edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004031.

Zheng:2024:OSC

[2277] Weichang Zheng, Jinke Li, Ming Ke, Mingcong Yang, Yu Zheng, Chenxiao Zhang, Qiang Liu, Yongbing Zhang, and Kun Yang. Optimizing spatial channel networks (SCNs) in hierarchical optical crossconnect (HOXC) architectures: Impact of wavelength switching granularity on performance. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004043.

Zazhigina:2024:ASR

[2278] Elizaveta Zazhigina, Ruslan Yusupov, Andrey Lyakhov, and Evgeny Khorov. Analytical study of Restricted Access Window with short slots for fast and reliable data delivery from energy-harvesting sensors. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004055.

Das:2024:DND

[2279] Manoj Das, Madhurima Buragohain, and Sukumar Nandi. Distributed neighbor discovery with mobility adaptive probing for beyond 5G D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004067.

Ren:2024:LDS

[2280] Wenhao Ren, Zichuan Xu, Weifa Liang, Haipeng Dai, Omer F. Rana, Pan Zhou, Qiufen Xia, Haozhe Ren, Mingchu Li, and Guowei Wu. Learningdriven service caching in MEC networks with bursty data traffic and uncertain delays. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004079.

Chen:2024:MLE

[2281] Zhiyan Chen, Murat Simsek, Burak Kantarci, Mehran Bagheri, and Petar Djukic. Machine learning-enabled hybrid intrusion detection system with host data transformation and an advanced two-stage classifier. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004080.

Wu:2024:ANA

[2282] Xiaodong Wu, Zhigang Jin, Junyi Zhou, Kai Liu, and Zepei Liu. Against network attacks in renewable power plants: Malicious behavior defense for federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004092.

Kuswiradyo:2024:OEC

[2283] Primatar Kuswiradyo, Binayak Kar, and Shan-Hsiang Shen. Optimizing the energy consumption in three-tier cloud-edge-fog federated systems with omnidirectional offloading. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004109.

Unal:2024:DPE

[2284] Deniz Unal, Sara Falleni, Kerem Enhos, Emrecan Demirors, Stefano Basagni, and Tommaso Melodia. Design and performance evaluation of SEANet, a software-defined networking platform for the Internet of Underwater Things. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004110.

Szabo:2024:TAN

[2285] Géza Szabó. Towards the automatic network resource management of OPC UA in 5G private networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004134.

Cai:2024:IET

[2286] Wei Cai, Chengshang Hou, Mingxin Cui, Bingxu Wang, Gang Xiong, and Gaopeng Gou. Incremental encrypted traffic classification via contrastive prototype networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004237.

Li:2024:ELS

[2287] Wenzhi Li, Lin Cui, and Xiaoquan Zhang. Enabling locality-sensitive machine learning towards low predictive overhead in flow classification. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004249.

Li:2024:LGD

[2288] Nianxin Li, Linbo Zhai, Zeyao Ma, Xiumin Zhu, and Yumei Li. Lyapunovguided Deep Reinforcement Learning for service caching and task offloading in Mobile Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004250.

Wang:2024:ENE

[2289] You-Chiun Wang and Cheng-Yan Wu. EC-NTD: Efficient countermeasure against DrDoS attacks with NAPT and two-stage detection in SDN-based networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004262.

Palama:2024:PSD

[2290] Ivan Palamà, Yago Lizarribar, Lorenzo Maria Monteforte, Giuseppe Santaromita, Stefania Bartoletti, Domenico Giustiniano, Giuseppe Bianchi, and Nicola Blefari Melazzi. 5G positioning with software-defined radios. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004274.

Zang:2024:EMT

[2291] Xiaodong Zang, Tongliang Wang, Xinchang Zhang, Jian Gong, Peng Gao, and Guowei Zhang. Encrypted malicious traffic detection based on natural language processing and deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004304.

Ismail:2024:PFB

[2292] Layal Ismail, Domenico Uomo, Andrea Sgambelluri, Faris Alhamed, and Francesco Paolucci. P4 FANET In-band Telemetry (FINT) for AIassisted wireless link failure forecasting and recovery. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004316.

Ma:2024:IAB

[2293] Yurui Ma, Jue Chen, Wenjing Lv, Xihe Qiu, Yue Zhang, and Wanxiao Liu. An improved artificial bee colony algorithm to minimum propagation latency and balanced load for controller placement in Software Defined Network. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004328.

Chen:2024:ORA

[2294] Wen Chen, Yuxiao Yang, Sibin Liu, and Wenjing Hu. Optimizing resource allocation for cluster D2D-assisted fog computing networks: a three-layer Stackelberg game approach. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400433X.

Han:2024:DDH

[2295] Pengbin Han, Xinfeng Wu, and Aina Sui. DTPBFT:A dynamic and highly trusted blockchain consensus algorithm for UAV swarm. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004341.

An:2024:JTO

[2296] Xiaobei An, Yanjun Li, Yuzhe Chen, and Tingting Li. Joint task offloading and resource allocation for multiuser collaborative mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004365.

Heino:2024:JUE

[2297] Jenny Heino, Christian Jalio, Antti Hakkala, and Seppo Virtanen. JAPPI: an unsupervised endpoint application identification methodology for improved Zero Trust models, risk score calculations and threat detection. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004389.

Cao:2024:SEV

[2298] Chengliang Cao, Fenghe Wang, Han Xiao, and Ye Wang. Secure and efficient vehicle data downloading scheme with privacy-preserving in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 250 (??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004420.

Dai:2024:SDP

[2299] Bin Dai, Hetao Li, and Yifan Wang. SP-DG: a programmable packet-level scheduling for queuing delay guarantees in time-critical networks. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004468.

Anonymous:2024:EBm

[2300] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004778.

Anonymous:2024:Ac

[2301] Anonymous. August 2024. Computer Networks (Amsterdam, Netherlands: 1999), 250(??):??, August 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Mancuso:2024:EDS

[2302] Vincenzo Mancuso, Leonardo Badia, Paolo Castagno, Matteo Sereno, and Marco Ajmone Marsan. Effectiveness of distributed stateless network server selection under strict latency constraints. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624003906.

Wang:2024:DEX

[2303] Wenbo Wang, Peng Yi, and Huikai Xu. DoubleR: Effective XSS attacking reality detection. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624003992.

Siasi:2024:PFR

[2304] N. Siasi, M. A. Jasim, and N. Ghani.

Post-fault restoration of service function chains in fog networks. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004122.

Xu:2024:MIN

[2305] Ke Xu, Donghong Jiang, Yanbiao Li, Xin Wang, Dafang Zhang, and Gaogang Xie. MaP: Increasing node capacity of programmable cloud gateways. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004225.

Rustam:2024:FNM

[2306] Furqan Rustam, Wajdi Aljedaani, Mahmoud Said Elsayed, and Anca Delia Jurcut. FAMTDS: a novel MFO-based fully automated malicious traffic detection system for multi-environment networks. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004353.

Guo:2024:SBB

[2307] Jiaxing Guo, Chunxiang Gu, Xi Chen, Xieli Zhang, Kai Tian, and Ji Li. Stateful black-box fuzzing for encryption protocols and its application in IPsec. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004377.

Zheng:2024:DAV

[2308] Yuxin Zheng, Lin Cui, Fung Po Tso, Zhetao Li, and Weijia Jia. DNN acceleration in vehicle edge computing with mobility-awareness: a synergistic vehicle-edge and edge-edge framework. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004390.

Lu:2024:CAW

[2309] Xinyue Lu, Yifan Yang, and Wei Gong. Challenges of ambient WiFi backscatter systems in healthcare applications. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004407.

Li:2024:BTB

[2310] Jianbin Li, Hengyang Zhang, Shike Li, Long Cheng, Yiguo Guo, and Sixing Wu. BD-TTS: a blockchain and DRL-based framework for trusted task scheduling in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004419.

Sagar:2024:UTM

[2311] Subhash Sagar, Adnan Mahmood, Quan Z. Sheng, Wei Emma Zhang, Yang Zhang, and Jitander Kumar Pabani. Understanding the trustworthiness management in the social Internet of Things: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004432.

Xiong:2024:LLB

[2312] Runqun Xiong, Qing Xiao, Zhoujie Wang, Zhuqing Xu, and Feng Shan. Leveraging lightweight blockchain for secure collaborative computing in UAV ad-hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004444.

Chen:2024:OQB

[2313] Zhenping Chen, Miaomiao Xu, and Chunxia Su. Online quality-based privacy-preserving task allocation in mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004456.

Zhao:2024:CCO

[2314] Liang Zhao, Yujun Yao, Jianmeng Guo, Qingjun Zuo, and Victor C. M. Leung. Collaborative computation offloading and wireless charging scheduling in multi-UAV-assisted MEC networks: a TD3-based approach. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400447X.

Sivaroopan:2024:NNT

[2315] Nirhoshan Sivaroopan, Dumindu Bandara, Chamara Madarasingha, Guillaume Jourjon, Anura P. Jayasumana, and Kanchana Thilakarathna. Net-Diffus: Network traffic generation by diffusion models through time-series imaging. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004481.

Pekar:2024:EMB

[2316] Adrian Pekar and Richard Jozsa.

Evaluating ML-based anomaly detection across datasets of varied integrity: a case study. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004493.

Ma:2024:CLC

[2317] Yunxiao Ma, Changqiao Xu, Zhonghui Wu, Renjie Ding, Han Xiao, Lujie Zhong, Yirong Zhuang, and Gabriel-Miro Muntean. CA-Live360: Crowdassisted transcoding and delivery for live 360-degree video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400450X.

Shlingbaum:2024:VNP

[2318] Erez Shlingbaum, Raz Ben Yehuda, Michael Kiperberg, and Nezer Jacob Zaidenberg. Virtualized network packet inspection. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004511.

Liu:2024:CBP

[2319] Yanxiu Liu, Linpeng Jia, Xin Wang, Huawei Huang, Qinglin Zhao, Zhongcheng Li, and Yi Sun. Coral: a blockchain protocol for handling transactions with deadline constraints. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624004523.

McEachern:2024:RTP

[2320] Jordin McEachern, Ehsan Malekshahi, Mojgan Mirzaei Hotkani, and Jean-François Bousquet. Real-time passive underwater localization using a compact acoustic sensor array. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004535.

Zhu:2024:RCB

[2321] Rongxin Zhu, Azzedine Boukerche, Yanxia Chen, and Qiuling Yang. A reliable cluster-based opportunistic routing protocol for underwater wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004547.

Zhang:2024:ADA

[2322] Qianhui Zhang, Lingyun Yuan, Tianyu Xie, and Han Chen. Auditable and dynamic access control scheme with behavior and identity tracing. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004559.

Chen:2024:RRA

[2323] Tao Chen, Yanling Bu, Yue Zeng, Lei Xie, and Sanglu Lu. RegionFilter: Region-aware video filtering mechanism on resource-constrained edge nodes. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004560.

Valente:2024:PID

[2324] Francesco Valente, Francesco G.

Lavacca, Tiziana Fiori, and Vincenzo Eramo. Proposal and investigation of a distributed learning strategy in Orbital Edge Computingendowed satellite networks for Earth Observation applications. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004572.

Rajab:2024:ARF

[2325] Abubakarsidiq Makame Rajab and Bang Wang. Autonomous and reliable fingerprint map maintenance for indoor positioning system. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004584.

Dinarte:2024:MOO

[2326] Henrique A. Dinarte, Karcius D. R. Assis, Daniel A. R. Chaves, Raul C. Almeida, and Raouf Boutaba. Multiobjective optimization of asymmetric bit rate partitioning for multipath protection in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004596.

Wu:2024:DPT

[2327] Dongkuo Wu, Xingwei Wang, Xueyi Wang, Rongfei Zeng, and Min Huang. Differentially private and truthful auction-based resource procurement for budget-constrained DAG applications in clouds. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004602.

deDiego:2024:CCI

[2328] Santiago de Diego, Cristina Regueiro, and Gabriel Maciá-Fernández. Collaborative credentials for the Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004614.

Chen:2024:KUN

[2329] Rui Chen, Lailong Luo, Xiaodong Wang, Bangbang Ren, Deke Guo, and Shi Zhu. Knowing the unknowns: Network traffic detection with open-set semi-supervised learning. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004626.

Girau:2024:DIC

[2330] Roberto Girau, Matteo Anedda, Roberta Presta, Silvia Corpino, Pietro Ruiu, Mauro Fadda, Chan-Tong Lam, and Daniele Giusto. Definition and implementation of the cloud infrastructure for the integration of the Human Digital Twin in the Social Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400464X.

Li:2024:TEL

[2331] Long Li, Gaochao Xu, Ziqi Liu, Jiaqi Ge, Wenchao Jiang, and Jiahui Li. Task execution latency minimization for energy-sensitive IoTs in wireless powered mobile edge computing: a DRL-based method. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004651.

Qi:2024:RAR

[2332] Zhuoran Qi, Roberto Petroccia, and Dario Pompili. RD-ASVTuw: Receiver-Driven Adaptive Scalable Video Transmission in underwater acoustic networks. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004663.

Ye:2024:LAB

[2333] Jin Ye, Huilin Hu, Jiahua Liang, Linfei Yin, and Jiawen Kang. Lightweight adaptive Byzantine fault tolerant consensus algorithm for distributed energy trading. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004675.

Xu:2024:BTC

[2334] Zhilin Xu, Hao Sun, and Weibin Han. Boosting task completion rate for timesensitive MCS system. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004687.

Taktak:2024:DAO

[2335] Wiem Taktak, Mohamed Escheikh, and Kamel Barkaoui. DRL approach for online user-centric QoSaware SFC embedding with dynamic VNF placement. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004699.

Dutta:2024:FRO

[2336] Ashit Kumar Dutta, Nuha Alruwais, Eatedal Alabdulkreem, Noha Negm, Abdulbasit A. Darem, Mesfer Al Duhayyim, Wali Ullah Khan, and Ali Nauman. Fair resource optimization for cooperative non-terrestrial vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004717.

Dai:2024:IID

[2337] Yuqi Dai, Hua Zhang, Jingyu Wang, Jianxin Liao, and Penghui Zhang. INCS: Intent-driven network-wide configuration synthesis based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004729.

Li:2024:SAI

[2338] Pengyong Li, Jiaqi Xia, Qian Wang, Yujie Zhang, and Meng Wu. Secure architecture for industrial edge of Things(IEoT): a hierarchical perspective. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004730.

Wu:2024:PEP

[2339] Zuodong Wu, Dawei Zhang, Yong Li, Chaoyang Li, and Xu Han. PRSD: Efficient protocol for privacy-preserving retrieval of sensitive data based on labeled PSI. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400481X.

Yamansavascilar:2024:ACS

[2340] Baris Yamansavascilar, Atay Ozgovde, and Cem Ersoy. Air computing: a survey on a new generation computation paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004857.

Bai:2024:IIS

[2341] Yang Bai, Gaojie Xing, Hongyan Wu, Zhihong Rao, Chengzong Peng, Yutang Rao, Wentao Yang, Chuan Ma, Jiani Li, and Yimin Zhou. ISPPFL: an incentive scheme based privacypreserving federated learning for avatar in metaverse. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004869.

Qin:2024:CBC

[2342] Xingsheng Qin, Frank Jiang, Xingguo Qin, Lina Ge, Meiqu Lu, and Robin Doss. CGAN-based cyber deception framework against reconnaissance attacks in ICS. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004870.

Bertalanic:2024:ESW

[2343] Blaz Bertalanic, Vid Hanzel, and Carolina Fortuna. Explainable semantic wireless anomaly characterization for digital twins. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004924.

Li:2024:CSC

[2344] Jian Li, Tongbao Chen, and Shaohua Teng. A comprehensive survey on client selection strategies in federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 251 (??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400495X.

Xu:2024:EER

[2345] Jun Xu and Dejun Yang. Energyefficient resource allocation for D2D communication underlaying cellular networks with incomplete CSI. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624004961.

Anonymous:2024:EBn

[2346] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624005346.

Anonymous:2024:S

[2347] Anonymous. September 2024. Computer Networks (Amsterdam, Netherlands: 1999), 251(??):??, September 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhu:2024:DAR

[2348] Rongxin Zhu, Azzedine Boukerche, Deshun Li, and Qiuling Yang. Delayaware and reliable medium access control protocols for UWSNs: Features, protocols, and classification. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004638.

Aljeri:2024:NNE

[2349] Noura Aljeri and Azzedine Boukerche. NEMa: a novel energy-efficient mobility management protocol for 5G/ 6G-enabled sustainable vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004705.

Hu:2024:ONT

[2350] Yahui Hu, Ziqian Zeng, Junping Song, Luyang Xu, and Xu Zhou. Online network traffic classification based on external attention and convolution by IP packet header. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004882.

Fazio:2024:NCM

[2351] Peppino Fazio, Miralem Mehic, and Miroslav Voznak. Next-cell and mobility prediction in new generation cellular systems based on convolutional neural networks and encoding mobility data as images. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004894.

Ashfaq:2024:EEI

[2352] Khuram Ashfaq and Ghazanfar Ali Safdar. Energy efficiency and interoperability through O-RAN Rapid Transition Protocol (ORTP). Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004900.

Ma:2024:ASM

[2353] Zhongyu Ma, Xueyao Zhang, Shunbao Zhang, Jianbing Pu, Xianghong Lin, and Qun Guo. Access strategies in mmWave cell-free network: a matching and auction theory based approach. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004912.

Bi:2024:SGH

[2354] Ye Bi, Kai Fan, Zhilin Zeng, Kan Yang, Hui Li, and Yintang Yang. Seamless group handover authentication protocol for vehicle networks: Services continuity. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624004936.

Lin:2024:PAR

[2355] Jing Lin, Weiwei Lin, Hang Lin, Longlong Zhu, Dong Zhang, and Chunming Wu. P4Rex: Accelerating regular expression matching with programmable switches. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004948.

Jia:2024:MAD

[2356] Kunkun Jia, Hui Xia, Rui Zhang, Yue Sun, and Kai Wang. Multi-agent DRL for edge computing: a real-time proportional compute offloading. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004973.

Tang:2024:ELA

[2357] Dan Tang, Hongbo Cao, Jiliang Zhang, Zheng Qin, Wei Liang, and Xiaopu Ma. EXCLF: a LDoS attack detection and mitigation model based on programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004985.

Xie:2024:SUS

[2358] Mangang Xie, Baozhen An, Xiangdong Jia, Meng Zhou, and Jintian Lu. Simultaneous update of sensing and control data using free-ride codes in vehicular networks: an age and energy perspective. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624004997.

Chiti:2024:SNT

[2359] Francesco Chiti, Roberto Picchi, and Laura Pierucci. A survey on nonterrestrial quantum networking: Challenges and trends. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005000.

Li:2024:CDD

[2360] Chao Li, Yanan Cheng, Zhaoxin Zhang, and ZunDong Zhang. Censorship data-driven DNS resolution anomaly detection: an ensemble algorithm model with multivariate feature fusion. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005012.

Anbazhagan:2024:NGR

[2361] S. Anbazhagan and R. K. Mugelan. Next-gen resource optimization in NB-IoT networks: Harnessing soft actorcritic reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005024.

Pokhrel:2024:MTI

[2362] Shiva Raj Pokhrel, Jonathan Kua, Brenton Fleming, Sebnem Ozer, Jeff Howe, and Anwar Walid. Multipath TCP implementation under FreeBSD-13 for pluggable machine learning models. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005036.

Zheng:2024:HRD

[2363] Yulu Zheng, Tiansheng Gu, Shangdong Liu, Sisi Shao, Qian Yang, Hongyu Du, Yuchen Shi, Yijun Nie, Zongkai Ji, and Yimu Ji. Highly reliable DHR-based polar compilation code communication method. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400505X.

Gomez:2024:LLA

[2364] Blas Gómez, Suzan Bayhan, Estefanía Coronado, José Villalón, and Antonio Garrido. LESS-ON: Load-aware edge server shutdown for energy saving in cellular networks. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005073.

Song:2024:STG

[2365] Chao Song, Jie Wu, Kunyang Xian, Jianfeng Huang, and Li Lu. Spatiotemporal graph learning: Traffic flow prediction of mobile edge computing in 5G/6G vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005085.

Li:2024:DTD

[2366] Mianjie Li, Qihan Pei, Chun Shan, Shen Su, Yuan Liu, and Zhihong Tian. Data tampering detection and recovery scheme based on multi-branch target extraction for Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005097.

Shariq:2024:ARU

[2367] Mohd Shariq, Mauro Conti, Karan Singh, Chhagan Lal, Ashok Kumar Das, Shehzad Ashraf Chaudhry, and Mehedi Masud. Anonymous and reliable ultralightweight RFID-enabled authentication scheme for IoT systems in cloud computing. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005103.

Busacca:2024:AVP

[2368] F. Busacca, L. Galluccio, S. Palazzo, A. Panebianco, Z. Qi, and D. Pompili. Adaptive versus predictive techniques in underwater acoustic communication networks. Computer Networks (Amsterdam, Netherlands:

1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005115.

Wang:2024:SUR

[2369] Jian Wang, Xin Cheng, and Guosheng Zhao. Social user role value analysis and trusted user autonomous diffusion for participatory crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005127.

Alberti:2024:DAS

[2370] Antônio M. Alberti, Diego G. S. Pivoto, Tibério T. Rezende, Alexis V. A. Leal, Cristiano B. Both, Michelle S. P. Facina, Rodrigo Moreira, and Flávio de Oliveira Silva. Disruptive 6G architecture: Software-centric, AIdriven, and digital market-based mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005140.

Li:2024:SEL

[2371] Wenxian Li, Yong Feng, Nianbo Liu, Yingna Li, Xiaodong Fu, and YongTao Yu. A secure and efficient log storage and query framework based on blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005152.

Chen:2024:UBB

[2372] Yitong Chen, Chaoqin Gan, Shibao Wu, Xiaoqi Wang, and Yixin Chen. (User behavior)-based LS prediction model and anti-LS reverse-leadership strategy in the (VLC/RF)-V2X network. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005164.

Liu:2024:MMS

[2373] Danyang Liu, Yuanqing Xia, and Yongkang Wang. MSDQ: Multi-Scheduling Dual-Queues coflow scheduling without prior knowledge. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005176.

Dunnett:2024:PSP

[2374] Kealan Dunnett, Shantanu Pal, Zahra Jadidi, Volkan Dedeoglu, and Raja Jurdak. Priv-Share: a privacy-preserving framework for differential and trustless delegation of cyber threat intelligence using blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005188.

Feng:2024:ATF

[2375] Xingbo Feng, Yi Wang, Jiashuo Lin, Weichao Li, Shuangping Zhan, Yan Liu, Jin Zhang, and Jianping Wang. Advancing TSN flow scheduling: an efficient framework without flow isolation constraint. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005206.

Shen:2024:TBTa

[2376] Fanfan Shen, Qiwei Liang, Lijie Hui, Bofan Yang, Chao Xu, Jun Feng, and Yanxiang He. TPE-BFL: Training Parameter Encryption scheme for Blockchain based Federated Learning system. Computer Networks (Amsterdam, Netherlands: 1999), 252 (??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005231.

Chaudhary:2024:RGC

[2377] Sameeksha Chaudhary, Anirudh Agarwal, Deepak Mishra, and Santosh Shah. A review on green communication for wearable and implantable wireless body area networks. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005255.

Kumar:2024:SCV

[2378] Naveen Kumar and Ankit Chaudhary. Surveying cybersecurity vulnerabilities and countermeasures for enhancing UAV security. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005279.

Anonymous:2024:EBo

[2379] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005723.

Anonymous:2024:O

[2380] Anonymous. October 2024. Computer Networks (Amsterdam, Netherlands: 1999), 252(??):??, October 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Qin:2024:DTD

[2381] Xin Qin, Wenwu Zhu, Qian Hu, Zexi Zhou, Yi Ding, Xia Gao, and Rentao Gu. DenseNet-Transformer: a deep learning method for spatialtemporal traffic prediction in optical fronthaul network. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005061.

Palmieri:2024:INT

[2382] Luigi Palmieri, Chiara Boldrini, Lorenzo Valerio, Andrea Passarella, and Marco Conti. Impact of network topology on the performance of Decentralized Federated Learning. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005139.

Parastar:2024:RME

[2383] Paniz Parastar, Giuseppe Caso, Jesus Alberto Omaña Iglesias, Andra Lutu, and Ozgu Alay. Rethinking the mobile edge for vehicular services. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400519X.

Cianca:2024:LBN

[2384] Ernestina Cianca, Syed Junaid Nawaz, Carla Amatetti, Tommaso Rossi, and Mauro De Sanctis. LEObased network-centric localization in 6G: Challenges and future perspectives. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005218.

Tu:2024:ESM

[2385] Yi-Hao Tu, En-Cheng Lin, Chih-Heng Ke, and Yi-Wei Ma. Enhanced-SETL: a multi-variable deep reinforcement learning approach for contention window optimization in dense Wi-Fi networks. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400522X.

Lai:2024:JOA

[2386] Pan Lai, Yiran Tao, Jun Qin, Yuanai Xie, Shihua Zhang, Shanjiang Tang, Qirui Huang, and Shengquan Liao. Joint optimization of application placement and resource allocation for enhanced performance in heterogeneous multi-server systems. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005243.

Lyu:2024:IEC

[2387] Zengwei Lyu, Yu Zhang, Xiaohui Yuan, Zhenchun Wei, Yu Fu, Lin Feng, and Haodong Zhou. Innovative edge caching: a multi-agent deep reinforcement learning approach for cooperative replacement strategies. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005267.

Lei:2024:ASC

[2388] Lei Lei, Aimin Tang, and Xudong Wang. Achieving scalable capacity in wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005280.

Patil:2024:CSS

[2389] Anita Patil, Sridhar Iyer, Onel L. A.

López, Rahul J. Pandya, Krishna Pai, Anshuman Kalla, and Rakhee Kallimani. A comprehensive survey on spectrum sharing techniques for 5G/B5G intelligent wireless networks: Opportunities, challenges and future research directions. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005292.

Pham:2024:RRL

[2390] Van-Hau Pham, Hien Do Hoang, Phan Thanh Trung, Van Dinh Quoc, Trong-Nghia To, and Phan The Duy. Raiju: Reinforcement learning-guided post-exploitation for automating security assessment of network systems. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005383.

Liu:2024:TIM

[2391] Ruitong Liu, Yanbin Wang, Zhenhao Guo, Haitao Xu, Zhan Qin, Wenrui Ma, and Fan Zhang. TransURL:

Improving malicious URL detection with multi-layer Transformer encoding and multi-scale pyramid features. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005395.

Xavier:2024:PMD

[2392] Bruno Missi Xavier, Merim Dzaferagic, Magnos Martinello, and Marco Ruffini. Performance measurement dataset for open RAN with user mobility and security threats. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005425.

Ji:2024:CTF

[2393] Ruoyu Ji, Fangmin Xu, Shihui Duan, Jinyu Zhu, Yiwen Tao, Meihui Liu, Peng Lv, and Chenglin Zhao. A cooperative timestamp-free clock synchronization scheme based on fast unscented Kalman filtering for timesensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005437.

Pei:2024:BAV

[2394] Hongmei Pei, Peng Yang, Miao Du, Zengyu Liang, and Zhongjian Hu. Blockchain-assisted verifiable secure multi-party data computing. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005449.

Du:2024:NQO

[2395] Bin Du, He Zhang, Xiangle Cheng, and Lei Zhang. Neural quantile optimization for edge-cloud networking. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005450.

Sojdeh:2024:SMS

[2396] Mohammad Javad Sojdeh, Mehdi Letafati, Seyed Pooya Shariatpanahi, and Babak Hossein Khalaj. Secure multi-server coded caching. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005474.

Chandran:2024:PSL

[2397] Indu Chandran and Kizheppatt Vipin.

A PUF secured lightweight mutual authentication protocol for multi-UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005498.

Du:2024:SRB

[2398] Ruizhong Du, Tianyang Chen, Jin Tian, and Tao Shang. StarCross: Redactable blockchain-based secure and lightweight data sharing framework for satellite-based IoT. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005504.

Liu:2024:ALC

[2399] Yanbo Liu, Peng Yu, Junye Zhang, Zhe Xiao, and Wenjing Li. Adaptive and low-cost resource synchronization based on data distribution service in high dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005516.

Parra-Espin:2024:EIS

[2400] José Antonio Parra-Espín, Rafael Marín-López, and Gabriel López-Millán. Establishment of IPsec Security Associations with Diffie–Hellman following a SDN-based framework: Analysis and practical validation. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005528.

Yoshinaka:2024:LAP

[2401] Yutaro Yoshinaka, Mio Kochiyama, Yuki Koizumi, Junji Takemasa, and Toru Hasegawa. A lightweight anonymity protocol at terabit speeds on programmable switches. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400553X.

Koutlia:2024:IOR

[2402] Katerina Koutlia and Sandra Lagén.

On the impact of Open RAN Fronthaul Control in scenarios with XR traffic. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005541.

Elzoghbi:2024:DGE

[2403] Mahmoud Elzoghbi and Hui He. DIS-Guard: Enhancing SDN resilience to topology and RCO attacks. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005553.

Li:2024:IDM

[2404] Zhengfa Li, Chuanhe Huang, and Wanyu Qiu. An intrusion detection method combining variational auto-encoder and generative adversarial networks. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005565.

Pugliese:2024:ITN

[2405] Daniele Pugliese, Mattia Quadrini, Domenico Striccoli, Cesare Roseti, Francesco Zampognaro, Giuseppe Piro, Luigi Alfredo Grieco, and Gennaro Boggia. Integrating terrestrial and non-terrestrial networks via IAB technology: System-level design and evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005589.

Liu:2024:LEA

[2406] Ximeng Liu, Shizhen Zhao, Xinbing Wang, and Chenghu Zhou. LPulse: an efficient algorithm for service function chain placement and routing with delay guarantee. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005607.

Gong:2024:MBC

[2407] Bobin Gong, Gaofei Huang, and Wanqing Tu. Minimize BER without CSI for dynamic RIS-assisted wireless broadcast communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005619.

Wang:2024:JRS

[2408] Tianle Wang, Xiuwen Fu, and Antonio Guerrieri. Joint resource scheduling and flight path planning of UAVassisted IoTs in response to emergencies. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005632.

Guo:2024:JOE

[2409] Lili Guo, Shibing Zhang, and Xiaodong Ji. Joint optimization for energy efficient full-duplex UAV relaying with multiple user pairs. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005644.

Calsi:2024:IML

[2410] Davide Li Calsi, Paul Kohl, JinHyeock Choi, and Janis Nötzel. The impact of message losses and retransmissions on quantum cryptographic protocols. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400567X.

Zheng:2024:MVM

[2411] Yu Zheng, Zhangxuan Dang, Xinglin Lian, Chunlei Peng, and Xinbo Gao. Multi-view multi-label network traffic classification based on MLP-Mixer neural network. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005784.

Ramonet:2024:PIO

[2412] Alberto Gallegos Ramonet, Tommaso Pecorella, Benedetta Picano, and Kazuhiko Kinoshita. Perspectives on IoT-oriented network simulation systems. Computer Networks (Amsterdam, Netherlands: 1999), 253 (??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005814.

Anonymous:2024:EBp

[2413] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624006170.

Anonymous:2024:N

[2414] Anonymous. November 2024. Computer Networks (Amsterdam, Netherlands: 1999), 253(??):??, November 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Caleffi:2024:DQC

[2415] Marcello Caleffi, Michele Amoretti, Davide Ferrari, Jessica Illiano, Antonio Manzalini, and Angela Sara Cacciapuoti. Distributed quantum computing: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005048.

Niu:2024:VSI

[2416] Dandan Niu, Guang Cheng, Zihan Chen, and Xing Qiu. Video stalling identification for web live streaming under HTTP-FLV. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005462.

Shi:2024:SRJ

[2417] Jianzhi Shi, Rou Rao, Yang Song, Xingwei Wang, Bo Yi, Qiang He, Chao Zeng, Min Huang, and Sajal K. Das. Service recommendation in JointCloud environments: an efficient regret theory-based QoS-aware approach. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005486.

Rago:2024:MLN

[2418] Arcangela Rago, Alessandro Guidotti, Giuseppe Piro, Ernestina Cianca, Alessandro Vanelli-Coralli, Simone Morosi, Giuseppe Virone, Fabrizio Brasca, Martina Troscia, Marina Settembre, Laura Pierucci, Francesco Matera, Mauro De Sanctis, Sara Pizzi, and Luigi Alfredo Grieco. Multi-layer NTN architectures toward 6G: the ITA-NTN view. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005577.

Amiri:2024:QAD

[2419] Zuhra Amiri, Shahram Dehdashti, Kareem H. El-Safty, Igor Litvin, Pere Munar-Vallespir, Janis Nötzel, and Simon Sekavčnik. Quantum advantages for data transmission in future networks: an overview. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005590.

Zhu:2024:BRF

[2420] Tengteng Zhu, Zehua Guo, Chao Yao, Jiaxin Tan, Songshi Dou, Wenrun Wang, and Zhenzhen Han. Byzantinerobust Federated Learning via Cosine Similarity Aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005620.

Akbar:2024:EEJ

[2421] Aamina Akbar, Ashfaq Ahmed, Adnan Zafar, and Sobia Jangsher. An energyefficient JT-CoMP enabled framework with adaptive OMA/NOMA in HetNets. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005656.

Nande:2024:SBP

[2422] Swaraj Shekhar Nande, Tommaso Rossi, Muhammad Idham Habibie, Mohamed Barhoumi, Krishna Palaparthy, Wassim Mansouri, Ashwin Raju, Riccardo Bassoli, Ernestina Cianca, Frank H. P. Fitzek, and Mauro De Sanctis. Satellite-based positioning enhanced by quantum synchronization. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005668.

Zhang:2024:DDB

[2423] Jingyu Zhang, Jin Yang, Liyao Li, Qifeng Nian, Lailong Luo, and Deke Guo. DBUP: Dynamic blockchain UTXO processing for storage efficiency optimization. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005760.

Li:2024:LPC

[2424] Zhuo Li, Nan Zhang, Hao Xun, Jindian Liu, Peng Luo, Yu Zhang, Teng Liang, Kaihua Liu, Wang Zhang, and Wanli Zhao. LearningTuple: a packet classification scheme with high classification and high update. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005772.

Yu:2024:NAA

[2425] Jaehak Yu, Yangseo Choi, Kijong Koo, and Daesung Moon. A novel approach for application classification with encrypted traffic using BERT and packet headers. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005796.

Zhou:2024:DSS

[2426] Deqiang Zhou, Xinsheng Ji, Wei You, Hang Qiu, Yu Zhao, and Mingyan Xu. DDQN-SFCAG: a service function chain recovery method against network attacks in 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005802.

Arzo:2024:SCM

[2427] Sisay Tadesse Arzo, Domenico Scotece, Riccardo Bassoli, Michael Devetsikiotis, Luca Foschini, and Frank H. P. Fitzek. Softwarized and containerized microservices-based network management analysis with MSN. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005826.

Ullah:2024:OMM

[2428] Waheed Ullah, Fengfan Yang, and Dushantha Nalin K. Jayakody. OTFS modulated massive MIMO with 5G NR LDPC coding: Trends, challenges and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005838.

Jeon:2024:AAR

[2429] Yu-Ran Jeon, Jung-Hwa Ryu, and Il-Gu Lee. ART: Adaptive relay transmission for highly reliable communications in next-generation wireless LANs. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400584X.

Dedu:2024:RTR

[2430] Eugen Dedu and Masoud Asghari.

Ray tracing routing using packet reception timing in dense nanonetworks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005851.

Zhu:2024:BNS

[2431] Qiang Zhu, Lin You, Geng-Ran Hu, and Wei-Nan Liu. BIAS: a novel secure and efficient biometric-based anonymous authentication scheme. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128624005863.

Byun:2024:GSI

[2432] Gyurin Byun, Huigyu Yang, Syed M. Raza, Moonseong Kim, Min Young Chung, and Hyunseung Choo. Generative spatiotemporal image exploitation for datacenter traffic prediction. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005875.

Liu:2024:LEM

[2433] Jianxin Liu, Zhiguo Xu, Rui Fan, and Zhigang Wen. Learning efficiency maximization in UAV-and-RIS-aided mobile edge learning system. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005887.

Gong:2024:GBI

[2434] Yiru Gong, Susu Cui, Song Liu, Bo Jiang, Cong Dong, and Zhigang Lu. Graph-based insider threat detection: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005899.

Ge:2024:PDM

[2435] Huanhuan Ge, Shiva Raj Pokhrel, Zhenyu Liu, Jinlong Wang, and Gang Li. PFL-DKD: Modeling decoupled knowledge fusion with distillation for improving personalized federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005905.

Zhang:2024:INR

[2436] Yan Zhang, Haopeng Huang, Qingqing Huang, and Yan Han. 6TiSCH IIoT network: a review. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005917.

Liu:2024:CLH

[2437] Gaosai Liu, Xinglong Jiang, Huawang Li, Zhenhua Zhang, Siyue Sun, and Guang Liang. Collaborative lasthop scheduling strategy for large-scale LEO constellation routing. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005929.

Chen:2024:LLI

[2438] Yishan Chen, Shumei Ye, Jie Wu, Wei Li, and Jiyuan Wang. Lowlatency intelligent service combination caching strategy with density peak clustering algorithm in vehicle edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005930.

Frank:2024:IRA

[2439] Lucas R. Frank, Antonino Galletta, Lorenzo Carnevale, Alex B. Vieira, and Edelberto Franco Silva. Intelligent resource allocation in wireless networks: Predictive models for efficient access point management. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005942.

Todisco:2024:FDB

[2440] Vittorio Todisco, Claudia Campolo, Antonella Molinaro, Antoine O. Berthet, Richard A. Stirling-Gallacher, and Alessandro Bazzi. Full duplex based collision detection to enhance the V2X sidelink autonomous mode. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005954.

Huang:2024:MEM

[2441] Chengyuan Huang, Tianfan Zhang, Li Wang, Yibo Xiao, Chao Yang, Chen Tian, Xiaoliang Wang, Dong Zhang, Bingheng Yan, Ahmed M. Abdelmoniem, Wanchun Dou, and Guihai Chen. MpScope: Enabling multi-pipeline monitoring inside a switch. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005966.

Baig:2024:ECB

[2442] Roger Baig, Llorenç Cerdà-Alabern, Pedro Vílchez, and Efraín Foglia. eXO: a commons-based community network Internet Service Provider, description and subscribers usage patterns. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624005978.

Liu:2024:ETA

[2443] Xiao Liu, Zhenyang Wei, Gaoxiang Li, and Jining Chen. An enhanced traceable access control scheme based on multi-authority CP-ABE for cloudassisted e-health system. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400598X.

Nguyen:2024:PEN

[2444] Ba Cao Nguyen, Pham Thanh Hiep, Xuan Nam Tran, and Nguyen Thu Phuong. Performance enhancement of NOMA multi-user networks with aerial reconfigurable intelligent surfaces and aerial relay. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624005991.

M:2024:LDA

[2445] Sameera K. M., Vinod P., Rafidha Rehiman K. A., and Mauro Conti. LFGurad: a defense against label flipping attack in federated learning for vehicular network. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006005.

Guo:2024:EDT

[2446] Wenjing Guo, Yingao Hou, Yanglan Gan, and Wenli Guo. Efficient data transmission mechanisms in energy harvesting wireless body area networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006017.

Hu:2024:AAG

[2447] Xiaoyan Hu, Di Li, Miao Li, Guang Cheng, Ruidong Li, and Hua Wu. AHDom: Algorithmically generated domain detection using attribute heterogeneous graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006029.

Li:2024:TTJ

[2448] Yafei Li, Huiqiang Wang, Jiayu Sun, Hongwu Lv, Wenqi Zheng, and Guangsheng Feng. Two-timescale joint service caching and resource allocation for task offloading with edge-cloud cooperation. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006030.

Mehjabin:2024:PPE

[2449] Suhee Sanjana Mehjabin, Mohamed Younis, Ali Tekeoglu, Mohammad Ebrahimabadi, Tamim Sookoor, and Naghmeh Karimi. PETIT: PUFenabled trust evaluation framework for IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006042.

Tian:2024:PSB

[2450] Junfeng Tian, Caishi Jing, Jin Tian, and Yaze Li. PartChain: Scaling blockchain through account-based partitioned sharding. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006054.

Li:2024:DDB

[2451] Ying Li, Jiuqi Wei, Ziyu Fei, Yufan Fu, and Xiaodong Lee. DiS-Auth: a DNS-based secure authorization framework for protecting data decoupled from applications. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006066.

Foruhandeh:2024:AOI

[2452] Mahsa Foruhandeh, Hanchao Yang, Xiang Cheng, Angelos Stavrou, Haining Wang, and Yaling Yang. All in one: Improving GPS accuracy and security via crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006078.

Lu:2024:DCP

[2453] Qiuyu Lu, Jun’e Li, Zhao Peng, Libing Wu, Ming Ni, and Jianbo Luo. Detecting the cyber-physical-social cooperated APTs in high-DER-penetrated smart grids: Threats, current work and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400608X.

Fu:2024:SIB

[2454] Yufan Fu, Xiaodong Lee, Jiuqi Wei, Ying Li, and Botao Peng. Securing the Internet’s backbone: a blockchainbased and incentive-driven architecture for DNS cache poisoning defense. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006091.

Tekin:2024:LMI

[2455] Nazli Tekin, Bilge Kagan Dedeturk, and Vehbi Cagri Gungor. Lifetime maximization of IoT-enabled smart grid applications using error control strategies. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006108.

Pan:2024:JHS

[2456] Guoqing Pan, Luyu Li, Rong Fan, and Yishan Su. A joint helper selection and power allocation scheme for secure underwater acoustic transmission. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400611X.

Soto:2024:DNI

[2457] Paola Soto, Miguel Camelo, Ginés García-Avilés, Esteban Municio, Marco Gramaglia, Evangelos Kosmatos, Nina Slamnik-Kriještorac, Danny De Vleeschauwer, Antonio Bazco-Nogueras, Lidia Fuentes, Joaquin Ballesteros, Andra Lutu, Luca Cominardi, Ivan Paez, Sergi Alcalá-Marín, Livia Elena Chatzieleftheriou, Andrés García-Saavedra, and Marco Fiore. Designing the Network Intelligence Stratum for 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006121.

Das:2024:LQU

[2458] Siddharth Das, Stefan Krause, Kay-Uwe Giering, Ricardo J. B. Pousa, Riccardo Bassoli, and Frank H. P. Fitzek. Leveraging quantum uncertainty: Quantum randomness through the lens of classical communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006133.

Shamsoshoara:2024:JPP

[2459] Alireza Shamsoshoara, Fatemeh Lotfi, Sajad Mousavi, Fatemeh Afghah, and Ismail Güvenç. Joint path planning and power allocation of a cellularconnected UAV using apprenticeship learning via deep inverse reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006212.

Ding:2024:TUT

[2460] Ruiyang Ding, Lei Sun, Weifei Zang, Leyu Dai, Zhiyi Ding, and Bayi Xu. Towards universal and transferable adversarial attacks against network traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006224.

Dong:2024:TOS

[2461] Shi Dong, Junxiao Tang, Khushnood Abbas, Ruizhe Hou, Joarder Kamruzzaman, Leszek Rutkowski, and Rajkumar Buyya. Task offloading strategies for mobile edge computing: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006236.

Beschastnyi:2024:CEC

[2462] Vitalii Beschastnyi, Darya Ostrikova, Dmitri Moltchanov, Yuliya Gaidamaka, Yevgeni Koucheryavy, and Konstantin Samouylov. Comparison of energy conservation strategies for 5G NR RedCap service in industrial environment. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006248.

Wang:2024:SMT

[2463] Suyue Wang, Hua Wu, Guang Cheng, Xiaoyan Hu, and Jing Ren. SD-MDN-TM: a traceback and mitigation integrated mechanism against DDoS attacks with IP spoofing. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400625X.

Nakamura:2024:AFI

[2464] Ryo Nakamura and Noriaki Kamiyama. On the aggregation of FIBs at ICN routers using routing strategy. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006261.

Sripotchanart:2024:TSL

[2465] Romtham Sripotchanart, Weisheng Si, Rodrigo N. Calheiros, Qing Cao, and Tie Qiu. A two-step linear programming approach for repeater placement in large-scale quantum networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006273.

Liu:2024:TMJ

[2466] Xi Liu, Jun Liu, and Weidong Li.

Truthful mechanism for joint resource allocation and task offloading in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006285.

Umar:2024:HSH

[2467] Amara Umar, Syed Ali Hassan, Haejoon Jung, Sahil Garg, Georges Kaddoum, and M. Shamim Hossain. HAC-SAGIN: High-altitude computing enabled space-air-ground integrated networks for 6G. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006297.

Wahida:2024:AML

[2468] Farah Wahida, M. A. P. Chamikara, Ibrahim Khalil, and Mohammed Atiquzzaman. An adversarial machine learning based approach for privacy preserving face recognition in distributed smart city surveillance. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006303.

Li:2024:FMD

[2469] Tun Li, Peng Shou, Xin Wan, Qian Li, Rong Wang, Chaolong Jia, and Yunpeng Xiao. A fast malware detection model based on heterogeneous graph similarity search. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006315.

Bodet:2024:DSD

[2470] Duschia Bodet, Jacob Hall, Ahmad Masihi, Ngwe Thawdar, Tommaso Melodia, Francesco Restuccia, and Josep M. Jornet. Data signals for deep learning applications in terahertz communications. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006327.

Wang:2024:MOJ

[2471] Houpeng Wang, Suzhi Cao, Huanjing Li, Lei Yan, Zhonglin Guo, and Yu’e Gao. Multi-objective joint optimization of task offloading based on MADRL in Internet of Things assisted by satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006339.

Truong:2024:ESO

[2472] Truong Van Truong and Anand Nayyar. Enhancing security offloading performance in NOMA heterogeneous MEC networks using access point selection and meta-heuristic algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006340.

Shaikh:2024:SFM

[2473] Shahzaib Shaikh and Manar Jammal. Survey of fault management techniques for edge-enabled distributed metaverse applications. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006352.

Marwein:2024:ELD

[2474] Phibadeity S. Marwein, Samarendra Nath Sur, and Debdatta Kandar. Efficient load distribution in heterogeneous vehicular networks using hierarchical controllers. Computer Networks (Amsterdam, Netherlands:

1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006376.

Liu:2024:TDA

[2475] Xi Liu, Jun Liu, and Weidong Li. A truthful double auction mechanism for resource provisioning and elastic service in vehicle computing. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006388.

Wang:2024:GMS

[2476] Jun Wang, Ning Wang, Haoju Wang, Kerang Cao, and Ahmed M. El-Sherbeeny. GCP: a multi-strategy improved wireless sensor network model for environmental monitoring. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400639X.

Javed:2024:TFB

[2477] Danish Javed, NZ Jhanjhi, Navid Ali Khan, Sayan Kumar Ray, Alanoud Al Mazroa, Farzeen Ashfaq, and Shampa Rani Das. Towards the future of bot detection: a comprehensive taxonomical review and challenges on Twitter/X. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006406.

Patetta:2024:LRB

[2478] Mario Patetta, Stefano Secci, and Sami Taktak. Line rate botnet detection with SmartNIC-embedded feature extraction. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006418.

Li:2024:NTP

[2479] Feng Li, Wei Nie, Kwok-Yan Lam, and Li Wang. Network traffic prediction based on PSO-LightGBM-TM. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400642X.

Rastogi:2024:EBS

[2480] Eshita Rastogi, Abhishek Roy, Ayush Rastogi, Jaehoon (Paul) Jeong, and Navrati Saxena. An efficient beam sweeping scheme with backup paging occasions in NR-Unlicensed Spectrum. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006431.

Imputato:2024:AWN

[2481] Pasquale Imputato, Thomas Henderson, Giovanni Nardini, and Michele Polese. Advances in wireless networks simulation. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006443.

Ahsan:2024:LTG

[2482] Muhammad Ahtazaz Ahsan, Amna Arshad, and Adnan Noor Mian. Leveraging tabular GANs for malicious address classification in Ethereum network. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006455.

Yu:2024:PUM

[2483] Chuan Yu, Shuhui Chen, Qianqian Xing, and Ziling Wei. Protecting unauthenticated messages in LTE/5G mobile networks: a two-level Hierarchical Identity-Based Signature (HIBS) solution. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006467.

Madureira:2024:NMW

[2484] André Luiz R. Madureira and Leobino N. Sampaio. NDN multicast over wireless networks: a survey on fundamentals, challenges, and open issues. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006479.

Zheng:2024:BRI

[2485] Songsong Zheng, Jinyao Liu, Xu Yan, Ziyang Xing, Xiaoqiang Di, and Hui Qi. BBR-R: Improving BBR performance in multi-flow competition scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006480.

Luo:2024:FTT

[2486] Xijian Luo, Jun Xie, Liqin Xiong, Yaqun Liu, and Yuan He. Faulttolerant topology construction and down-link rate maximization in airground integrated networks. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006492.

Seth:2024:CCR

[2487] Sayanta Seth, Debashri Roy, and Murat Yuksel. Cognisseum: Cognitive radios on Colosseum facing adversaries. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006509.

Ahmed:2024:INI

[2488] Waqas Ahmed, Nadir Shah, and Gabriel-Miro Muntean. An innovative NSGA-II-based Byzantine Fault Tolerant solution for software defined network environments. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006510.

Wei:2024:JAT

[2489] Xingchen Wei, Laixian Peng, Renhui Xu, and Hai Wang. Jamming avoidance trajectory planning and load balancing user association in mmWave UAV-assisted HetECN. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006522.

Benamor:2024:PLS

[2490] Amani Benamor, Oussama Habachi, Jean-Pierre Cances, and Vahid Meghdadi. Physical layer security for confidential transmissions in frequency hopping-based downlink NOMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006534.

Wang:2024:LDP

[2491] Jie Wang, Zhiju Zhang, Jing Tian, and Hongtao Li. Local differential privacy federated learning based on heterogeneous data multi-privacy mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006546.

Palena:2024:EDC

[2492] Marco Palena, Tania Cerquitelli, and Carla Fabiana Chiasserini. Edgedevice collaborative computing for multi-view classification. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006558.

Zhao:2024:OSI

[2493] Xingwen Zhao, Han Zhang, Hui Li, and Xuangui Chen. Open set identification of malicious encrypted traffic based on multi-feature fusion. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400656X.

Rui:2024:MAH

[2494] Lanlan Rui, Liangchen Zhao, Zilong Guo, Zihan Wang, Xuesong Qiu, and Shaoyong Guo. Mobile ad hoc network access authentication mechanism based on rotation election and two-factor aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006583.

Bayazeed:2024:TCS

[2495] Adnan Bayazeed, Khaldoun Khorzom, and Mohamad Aljnidi. Towards coordinating self-healing instances: Policybased and non-cooperative game theory-based approaches with location awareness. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006595.

M:2024:AAS

[2496] Jesús F. Cevallos M., Alessandra Rizzardi, Sabrina Sicari, and Alberto Coen-Porisini. ASAP: Automatic Synthesis of Attack Prototypes, an online-learning, end-to-end approach. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006601.

Guan:2024:BFF

[2497] Zhong Guan, Chang Liu, Gaopeng Gou, Zhen Li, Gang Xiong, Yangyang Ding, and Chengshang Hou. A blind flow fingerprinting and correlation method against disturbed anonymous traffic based on pattern reconstruction. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006637.

Dai:2024:AAD

[2498] Bin Dai, Hetao Li, and Yifan Wang.

aCroSS: AI-driven cross-layer adaptive streaming for short video applications. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006649.

Gupta:2024:NQD

[2499] Jit Gupta, Sourav Das, and Krishna Kant. NeSt: a QoS differentiating end-to-end networked storage simulator. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006650.

Almasaeid:2024:EME

[2500] Hisham M. Almasaeid. Efficient multichannel energy harvesting with dedicated energy transmitters in CR-IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006662.

Temene:2024:FTN

[2501] Natalie Temene, Andreas Naoum, Charalampos Sergiou, Chryssis Georgiou, and Vasos Vassiliou. A fault tolerant node placement algorithm for WSNs and IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006674.

Albishari:2024:FDL

[2502] Mohammed Albishari, Mingchu Li, Majid Ayoubi, Ala Alsanabani, and Jiyu Tian. Federated deep learning models for detecting RPL attacks on large-scale hybrid IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006698.

Chen:2024:RIM

[2503] Honghong Chen, Jie Yang, Zhanjun Hao, Tian Qi, and TingTing Liu. Research on indoor multi-floor positioning method based on LoRa. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006704.

Tang:2024:CTI

[2504] Hailiang Tang, Dawei Lin, Wanyu Li, Wenxiao Zhang, and Jun Zhao. Cyber threat indicators extraction based on contextual knowledge prompt. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624006716.

Hong:2024:SEA

[2505] Hsiang-Jen Hong, Sang-Yoon Chang, Wenjun Fan, Simeon Wuthier, and Xiaobo Zhou. Secure and efficient authentication using linkage for permissionless Bitcoin network. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006728.

Abdulqadder:2024:DBA

[2506] Ihsan H. Abdulqadder, Israa T. Aziz, and Deqing Zou. DT-Block: Adaptive vertical federated reinforcement learning scheme for secure and efficient communication in 6G. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400673X.

Li:2024:ITO

[2507] Yuchen Li, Hongwei Ding, Zhijun Yang, Bo Li, and Zhuguan Liang. Integrated trajectory optimization for UAV-enabled wireless powered MEC system with joint energy consumption and AoI minimization. Computer Networks (Amsterdam, Netherlands:

1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006741.

Zheng:2024:PES

[2508] Danyang Zheng, Huanlai Xing, Li Feng, and Xiaojun Cao. Provably efficient security-aware service function tree composing and embedding in multi-vendor networks. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006753.

Kumar:2024:WHW

[2509] Navneet Kumar, Karan Singh, and Jaime Lloret. WAOA: a hybrid whaleant optimization algorithm for energyefficient routing in wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006777.

Hadi:2024:UMI

[2510] Majid Hadi and Reza Ghazizadeh. UAV-mounted IRS assisted wireless powered mobile edge computing systems: Joint beamforming design, resource allocation and position optimization. Computer Networks (Amsterdam, Netherlands: 1999), 254 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006789.

Anonymous:2024:EBq

[2511] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624007540.

Anonymous:2024:Da

[2512] Anonymous. December 2024. Computer Networks (Amsterdam, Netherlands: 1999), 254(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Brenner:2024:MCS

[2513] Bernhard Brenner, Joachim Fabini, Magnus Offermanns, Sabrina Semper, and Tanja Zseby. Malware communication in smart factories: a network traffic data set. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006364.

Li:2024:SPI

[2514] Peihao Li, Jie Huang, Shuaishuai Zhang, and Chunyang Qi. SecureEI: Proactive intellectual property protection of AI models for edge intelligence. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006571.

Chen:2024:NTE

[2515] Jie Chen, Jian Luo, and Kai Gao. NetBoost: Towards efficient distillation and service differentiation of network information exposure. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006613.

Chong:2024:LDP

[2516] Kah Meng Chong and Amizah Malip. Local differential privacy for correlated location data release in ITS. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006625.

Xiao:2024:TPA

[2517] Jingyu Xiao, Qing Li, Dan Zhao, Xudong Zuo, Wenxin Tang, and Yong Jiang. Themis: a passive-active hybrid framework with in-network intelligence for lightweight failure localization. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006686.

Amanatidis:2024:ART

[2518] Petros Amanatidis, Dimitris Karampatzakis, Georgios Michailidis, Thomas Lagkas, and George Iosifidis. Adaptive reverse task offloading in edge computing for AI processes. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006765.

Wang:2024:MMC

[2519] Yabo Wang, Ruizhi Xiao, Jiakun Sun, and Shuyuan Jin. MC-Det: Multi-channel representation fusion for malicious domain name detection. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006790.

Zhao:2024:CCB

[2520] Liushun Zhao, Deke Guo, Lailong Luo, Yulong Shen, Bangbang Ren, Shi Zhu, and Fangliao Yang. Concordit: a credit-based incentive mechanism for permissioned redactable blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006807.

Gurung:2024:PAE

[2521] Dev Gurung, Shiva Raj Pokhrel, and Gang Li. Performance analysis and evaluation of postquantum secure blockchained federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006819.

Deng:2024:CRA

[2522] Yingzhuo Deng, Zicheng Hu, Weihao Xu, Ningning Han, and Haibin Cai. Collaborative resource allocation in computing power networks: a game-theoretic double auction perspective. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006820.

Duan:2024:SCE

[2523] Pengfei Duan, Hongmin Gao, Yushi Shen, Zhetao Guo, Zhaofeng Ma, Tian Tian, and Yuqing Zhang. Secure collaborative EHR sharing using multiauthority attribute-based proxy reencryption in Web 3.0. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006832.

Ashraf:2024:TES

[2524] Usman Ashraf, Adnan Ahmed, Stefano Avallone, and Pasquale Imputato. Traffic evolution in Software Defined Networks. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006844.

Mfogo:2024:ALB

[2525] Volviane Saphir Mfogo, Alain Zemkoho, Laurent Njilla, Marcellin Nkenlifack, and Charles Kamhoua. Adaptive learning-based hybrid recommender system for deception in Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006856.

Casetti:2024:AMB

[2526] Claudio Casetti, Carla Fabiana Chiasserini, Falko Dressler, Agon Memedi, Diego Gasco, and Elad Michael Schiller. AI/ML-based services and applications for 6G-connected and autonomous vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006868.

Meyer:2024:NAD

[2527] Philipp Meyer, Timo Häckel, Sandra Reider, Franz Korf, and Thomas C. Schmidt. Network anomaly detection in cars: a case for time-sensitive stream filtering and policing. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400687X.

Zhu:2024:TSF

[2528] Zaixing Zhu, Tao Hu, Di Wu, Chengcheng Liu, Siwei Yang, and Zhifu Tian. Topology sensing of FANET under missing data. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006881.

Alkhaldi:2024:ESC

[2529] Tareq M. Alkhaldi, Abdulbasit A. Darem, Asma A. Alhashmi, Tawfik Al-Hadhrami, and Azza Elneil Osman. Enhancing smart city IoT communication: a two-layer NOMA-based network with caching mechanisms and optimized resource allocation. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624006893.

Xu:2024:CEC

[2530] Sheng wei Xu, Shu han Yu, Zi-Yan Yue, and Yi-Long Liu. CLLS: Efficient certificateless lattice-based signature in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400690X.

Oueslati:2024:LGF

[2531] Ibtissem Oueslati, Oussama Habachi, Jean-Pierre Cances, and Vahid Meghdadi. LoCoNOMA: a grant-free resource allocation for massive MTC. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624006911.

Bao:2024:MCN

[2532] Yubing Bao, Xin Du, Zhihui Lu, Jirui Yang, Shih-Chia Huang, Jianfeng Feng, and Qibao Zheng. Mitigating critical nodes in brain simulations via edge removal. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006923.

Li:2024:CBM

[2533] Qiyuan Li, Yumeng Wang, Donghai Tian, Chong Yuan, and Changzhen Hu. Component-based modeling of cascading failure propagation in directed dual-weight software networks. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006935.

Verma:2024:EHC

[2534] Himanshu Verma, Naveen Chauhan, and Lalit Kumar Awasthi. Enhanced hybrid congestion mitigation strategy for ‘6LoWPAN-RPL based patientcentric IoHT’. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006947.

Yang:2024:EHA

[2535] Guangcan Yang, Peixuan Li, Yang Xin, Yunhua He, Chao Wang, and Xiubo Chen. An efficient hierarchical attribute-based encryption scheme with cross-domain data sharing. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006959.

Liu:2024:IVR

[2536] Xueyan Liu, Xin Xiong, Jia Wang, and Yujiao Qi. An Internet of Vehicles road traffic data sharing scheme based on signcryption and editable blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006960.

Yang:2024:ODS

[2537] Lei Yang, Juan A. Fraire, Kanglian Zhao, Ruhai Wang, Wenfeng Li, and Hong Yang. Optimizing deep-space DTN congestion control via deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006972.

Zhang:2024:UOC

[2538] Ruidong Zhang, Jiadong Zhang, Xue Wang, and Wenxiao Shi. Utility optimization for computation offloading and splitting in time-varying HAP and LEO satellite integrated MEC networks. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006984.

Hu:2024:TFG

[2539] Jinbin Hu, Ruiqian Li, Ying Liu, and Jin Wang. Towards fine-grained load balancing with dynamical flowlet timeout in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624006996.

Lai:2024:GCB

[2540] Yingxu Lai, Jingwen Wei, and Ye Chen. Gradient correlation based detection of adversarial attacks on vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400700X.

GhorbanniaDelavar:2024:FMH

[2541] Arash GhorbanniaDelavar and Zahra Jormand. FMORT: the meta-heuristic routing method by integrating index parameters to optimize energy consumption and real execution time using FANET. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007011.

Gupta:2024:MAB

[2542] Anjali Gupta and Abhishek Dixit. Mathematical analysis of busy tone in full-duplex optical MAC for hidden node mitigation. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007023.

Mohammed:2024:EOS

[2543] Hesham Mohammed and Dola Saha. Encrypted-OFDM: a secured wireless waveform. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007035.

Pan:2024:JIO

[2544] Gang Pan, Xin Guan, Haiyang Jiang, Yongnan Liu, Huayang Wu, Hongyang Chen, Tomoaki Ohtsuki, and Zhu Han. Joint intelligent optimizing economic dispatch and electric vehicles charging in 5G vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007047.

Ashraf:2024:IUA

[2545] Muhammad Mansoor Ashraf, Saadi Boudjit, Sherali Zeadally, Nour El Houda Bahloul, and Nouman Bashir. Integrating Unmanned Aerial Vehicles (UAVs) with vehicular ad-hoc NETworks (VANETs): Architectures, applications, opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007059.

Fraire:2024:QAS

[2546] Juan A. Fraire, Santiago Henn, Gregory Stock, Robin Ohs, Holger Hermanns, Felix Walter, Lynn Van Broock, Gabriel Ruffini, Federico Machado, Pablo Serratti, and Jose Relloso. Quantitative analysis of segmented satellite network architectures: a maritime surveillance case study. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007060.

Aygul:2024:MLD

[2547] Mehmet Ali Aygul, Halise Turkmen, Hakan Ali Cirpan, and Huseyin Arslan. Machine learning-driven integration of terrestrial and non-terrestrial networks for enhanced 6G connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007072.

Blanco-Romero:2024:EIM

[2548] Javier Blanco-Romero, Vicente Lorenzo, Florina Almenares, Daniel Díaz-Sánchez, Carlos García Rubio, Celeste Campo, and Andrés Marín. Evaluating integration methods of a quantum random number generator in OpenSSL for TLS. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007096.

Ojijo:2024:SAC

[2549] Mourice Otieno Ojijo, Daniel Ramotsoela, and Ruth A. Oginga. Slice admission control in 5G wireless communication with multi-dimensional state space and distributed action space: a sequential twin actor-critic approach. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007102.

Rehman:2024:NMM

[2550] Rameesha Rehman, Saif Ur Rehman Malik, Shahida Hafeezan Qureshi, and Syed Atif Moqurrab. A novel multi-modal federated learning based thermal-aware job scheduling framework. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007114.

Naimi:2024:ILB

[2551] Mohamed Rabie Naimi, Chakib Zouaoui, Mohamed Elbahri, and Abdenacer Bounoua. Inventory of load balancing parameters in MPTCP schedulers. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007126.

Shen:2024:TBTb

[2552] Fanfan Shen, Bofan Yang, Jun Zhang, Chao Xu, Yong Chen, and Yanxiang He. TD3-based trajectory optimization for energy consumption minimization in UAV-assisted MEC system. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400714X.

Liu:2024:MPP

[2553] Wai xi Liu, Dao xiao Chen, Miao quan Tan, Kong yang Chen, Yue Yin, Wen-Li Shang, Jin Li, and Jun Cai. Model parameter prediction method for accelerating distributed DNN training. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007151.

Yang:2024:GCE

[2554] Duo Yang, Yunqi Gao, Bing Hu, A-Long Jin, Wei Wang, and Yang You. GWPF: Communication-efficient federated learning with Gradient-Wise Parameter Freezing. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007187.

Borylo:2024:AOV

[2555] Piotr Borylo, Wojciech Stasiak, Piotr Piwowar, and Nashid Shahriar. Automated orchestration of virtualized deployments of mobile core networks. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007205.

Zheng:2024:SNS

[2556] Tianshuai Zheng, Jinglei Tan, Xuesong Wu, Ruiqin Hu, Qifang Chen, Zhiquan Liu, and Ye Du. SARM: a network State-Aware Adaptive Routing Mutation method for power IoT. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007217.

Zhou:2024:SFR

[2557] Jie Zhou, Jiao Deng, Shengke Zeng, Mingxing He, and Xingwei Liu. Secure fuzzy retrieval protocol for multiple datasets. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007230.

Hosseinpour:2024:CNB

[2558] Shaghayegh Hosseinpour and Hadi Shakibian. Complex-network based model for SMS spam filtering. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624007242.

Wang:2024:REE

[2559] Ying Wang, Yuanyuan Li, Jianjun Lei, and Fengjun Shang. Robust and energy-efficient RPL optimization algorithm with scalable deep reinforcement learning for IIoT. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007266.

Pan:2024:PPL

[2560] Yingying Pan, Lanxiang Chen, and Gaolin Chen. Privacy-preserving local clustering coefficient query on structured encrypted graphs. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007278.

Trabelsi:2024:DRL

[2561] Nessrine Trabelsi, Lamia Chaari Fourati, and Wael Jaafar. Deep reinforcement learning for autonomous SideLink radio resource management in platoon-based C-V2X networks: an overview. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007333.

Ivanova:2024:PMC

[2562] Daria Ivanova, Elena Zhbankova, Ekaterina Markova, Yuliya Gaidamaka, and Konstantin Samouylov. Performance modeling and comparison of URLLC and eMBB coexistence strategies in 5G new radio systems. Computer Networks (Amsterdam, Netherlands: 1999), 255 (??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007369.

Chen:2024:MUT

[2563] Lu Chen and Zhijie Wang. Multi-UAV trajectory planning for RISassisted SWIPT system under connectivity preservation. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007382.

Troia:2024:ENS

[2564] Sebastian Troia, Marco Savi, Christian Grasso, Frank Slyne, Lihua Ruan, and Michael Seufert. Edge network softwarization and intelligence: Challenges and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007424.

Anonymous:2024:EB

[2565] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624007783.

Anonymous:2024:Db

[2566] Anonymous. December 2024. Computer Networks (Amsterdam, Netherlands: 1999), 255(??):??, December 2024. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2025:Ja

[2567] Anonymous. January 2025. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2025:EBa

[2568] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008004.

Prieto:2025:UVF

[2569] Yasmany Prieto, Christian Vega, and Miguel Figueroa. Using virtual function replacement to mitigate 0-day attacks in a multi-vendor NFV-based network. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007345.

Liu:2025:ASI

[2570] Xiuwen Liu, Shuo Wang, and Yanjiao Chen. Adaptive scheduling for Internet of Vehicles using deconfounded graph transfer learning. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400731X.

Yellas:2025:FPN

[2571] Nour-El-Houda Yellas, Bernardetta Addis, Selma Boumerdassi, Roberto Riggio, and Stefano Secci. Function placement for in-network federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007321.

Ji:2025:CML

[2572] Mingtao Ji, Ji Qi, Lei Jiao, Gangyi Luo, Hehan Zhao, Xin Li, Baoliu Ye, and Zhuzhong Qian. CPN meets learning: Online scheduling for inference service in computing power network. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007357.

Hou:2025:SSD

[2573] Ruomu Hou, Haifeng Yu, and Yucheng Sun. Selfied: Sybil defense in permissionless blockchains via in-protocol bandwidth consumption. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007229.

Bojovic:2025:CSU

[2574] Biljana Bojović and Sandra Lagén.

3GPP-compliant single-user MIMO model for high-fidelity mobile network simulations. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007448.

Xiong:2025:NCE

[2575] Wanjun Xiong, Yujue Wang, and Yongzhuang Wei. NTRU-CLS: Efficient quantum-resistant NTRU latticebased certificateless signature scheme for VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007175.

Jamil:2025:UIR

[2576] Syed Usman Jamil, M. Arif Khan, Muhammad Ali Paracha, and Abdul Rasheed. Using intelligence in resource allocation and task off-loading for the IoE-edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007163.

Hwang:2025:DRA

[2577] Ren-Hung Hwang, Jia-You Lin, Yen Chuang, and Ben-Jye Chang. Dynamic resource allocation and offloading optimization for network slicing in B5G multi-tier multi-tenant systems. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007084.

Coletta:2025:UAA

[2578] Andrea Coletta, Flavio Giorgi, Gaia Maselli, Matteo Prata, Domenicomichele Silvestri, Jonathan Ashdown, and Francesco Restuccia. A ^{2}- UAV: Application-aware resilient edgeassisted UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007199.

Zhang:2025:DDP

[2579] Heng-Ru Zhang, Ke-Xiong Wang, Xiang-Yu Liang, and Yi-Fan Yu. DUPS: Data poisoning attacks with uncertain sample selection for federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007412.

Tan:2025:ZIF

[2580] Yawen Tan, Jiajia Liu, and Jiadai Wang. Zoom-inRCL: Fine-grained root cause localization for B5G/6G network slicing. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007254.

Lu:2025:ITC

[2581] Jingwen Lu, Chaowei Tang, Zhengchuan Chen, Jiayuan Guo, Aobo Zou, Wen Yang, and Chenxi Tang. Intraflow temporal correlation-based network traffic prediction. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400745X.

Wang:2025:FPF

[2582] Mengyu Wang, Shuyong Zhu, and Yujun Zhang. FastDet: Providing faster deterministic transmission for timesensitive flows in WAN. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007138.

Chagnon:2025:SPV

[2583] Dorine Chagnon, Kevin Thiry-Atighehchi, and Gérard Chalhoub. A survey on path validation: Towards digital sovereignty. Computer Networks (Amsterdam, Netherlands: 1999), 256 (??):??, January 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007370.

Li:2025:OSE

[2584] Jingyi Li, Jiajia Liu, and Jiadai Wang. Optimizing spectrum and energy efficiency in IRS-enabled UAVground communications. Computer Networks (Amsterdam, Netherlands: 1999), 256(??):??, January 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007436.

Anonymous:2025:Fa

[2585] Anonymous. February 2025. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2025:EBb

[2586] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128625000192.

Nguyen:2025:ODI

[2587] Xuan-Xinh Nguyen and Ha Hoang Kha. Optimized design for integrated sensing and communication in secure MIMO SWIPT systems. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400762X.

Sharma:2025:POO

[2588] Sidharth Sharma, Yuan-Cheng Lai, Ashwin Gumaste, and Ying-Dar Lin. P4+NFV: Optimal offloading from P4 switches to NFV for diverse traffic streams. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007394.

Zhou:2025:JCS

[2589] Sheng Zhou, Liangmin Wang, Weihua Wu, and Li Feng. Joint client selection and resource allocation for federated edge learning with imperfect CSI. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007461.

Li:2025:CBS

[2590] Xin Li, Meiyan Teng, Yanling Bu, Jianjun Qiu, Xiaolin Qin, and Jie Wu. Cooperation-based server deployment strategy in mobile edge computing system. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007643.

Scott:2025:BAD

[2591] Ben A. Scott, Michael N. Johnstone, Patryk Szewczyk, and Steven Richardson. BGP anomaly detection as a group dynamics problem. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007588.

Sen:2025:SAB

[2592] Nabhasmita Sen and Antony Franklin A. Slice aware baseband function splitting and placement in disaggregated 5G radio access network. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007400.

Liu:2025:SSF

[2593] Kaixiang Liu, Yongfang Xie, Yuqi Chen, Shiwen Xie, Xin Chen, Dongliang Fang, and Limin Sun. SFACIF: a safety function attack and anomaly industrial condition identified framework. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400759X.

Geng:2025:DDG

[2594] Haijun Geng, Qi Ma, Haotian Chi, Zhi Zhang, Jing Yang, and Xia Yin. DUdetector: a dual-granularity unsupervised model for network anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007692.

Cui:2025:ARE

[2595] Zixi Cui, Yuxiang Hu, Le Tian, Peng Yi, Saifeng Hou, and Hongchang Chen.

Altair: Resource-efficient optimization and deployment for data plane programs. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007497.

Zeng:2025:LLP

[2596] Bo Zeng, Tian Wu, Fangchao Yu, Geying Yang, Kai Zhao, and Lina Wang. LPPAC: Lightweight privacypreserving distributed payments with access control. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007680.

Jin:2025:RRS

[2597] Zilong Jin, Huajian Xu, Zhixiang Kong, and Chengsheng Pan. A resilient routing strategy based on deep reinforcement learning for urban emergency communication networks. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007308.

Zhang:2025:RTL

[2598] Yuxiang Zhang, Lin Cui, Fung Po Tso, and Xiaolin Lei. Reducing tail latency for multi-bottleneck in datacenter networks: a compound approach. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007631.

Lei:2025:HEH

[2599] Fei Lei, Xianliang Jiang, Guang Jin, and Dingxin Yu. HawkEye: an endhost method to detect the low-rate denial-of-service attack of cross-traffic over bottleneck links. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007837.

Hu:2025:CMS

[2600] Jiahai Hu, Lin Wang, Jing Wu, Qiangyu Pei, Fangming Liu, and Bo Li. A comparative measurement study of cross-layer 5G performance under different mobility scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007849.

Jisi:2025:ESC

[2601] Chandroth Jisi, Byeong hee Roh, and Jehad Ali. An effective scheme for classifying imbalanced traffic in SD-IoT, leveraging XGBoost and active learning. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007710.

Xing:2025:JPA

[2602] Fangyuan Xing, Yutian Tan, Zhenduo Wang, Yaxing Yue, and Hongxi Yin. Joint power allocation for multiple nodes in serial relaying networks over underwater wireless optical channels. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007734.

Li:2025:EEA

[2603] Peiqiang Li, Guojun Wang, Xiaofei Xing, Jinyao Zhu, Wanyi Gu, and Yuheng Zhang. EAOS: Exposing attacks in smart contracts through analyzing opcode sequences with operands. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007916.

Shen:2025:SCR

[2604] Guanqun Shen, Xinchen Wei, Kaikai Chi, Fayez Alqahtani, and Amr Tolba. Sum computation rate maximization for wireless powered OFDMAbased mobile edge computing network. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400793X.

Song:2025:SRA

[2605] Boxuan Song, Fei Wang, and Yujie Su. STAR-RIS-aided NOMA communication for mobile edge computing using hybrid deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007928.

Zhang:2025:SSI

[2606] Jing Zhang, Jianfeng Guan, Kexian Liu, Yizhong Hu, Ao Shen, and Yuyin Ma. ScaIR: Scalable Intelligent Routing based on distributed graph reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007473.

Gulzari:2025:OCP

[2607] Usman Ali Gulzari, Waqar Farooq, Syed Nasir Mehmood Shah, Iftikhar Ahmed Khan, Sheraz Anjum, Zoran Salcic, and Hessam Sarjoughian. The octagonal-cross-by-pass-mesh topology design for the on-chip-communication. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007655.

Coscia:2025:FFR

[2608] Antonio Coscia, Antonio Maci, and Nicola Tamma. FROG: a firewall rule order generator for faster packet filtering. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007941.

Junior:2025:MUE

[2609] Paulo J. S. Júnior, Lucas R. Costa, André C. Drummond, and Marcelo A. Marotta. Minimizing unavailability in elastic optical networks: Preprovisioning and provisioning protection strategy using DLP and DPP. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007606.

Zhang:2025:JSC

[2610] Yue Zhang, Zhenyu Na, Zihao Wen, Arumugam Nallanathan, and Weidang Lu. Joint service caching, computation offloading and resource allocation for dual-layer aerial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008065.

Zhao:2025:POD

[2611] Deyu Zhao, Guang Cheng, Xuan Chen, Yuyu Zhao, Wei Zhang, Lu Lu, Siyuan Zhou, and Yuexia Fu. Probe-Optimizer: Discovering important nodes for proactive in-band network telemetry to achieve better probe orchestration. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007679.

Liu:2025:EDO

[2612] Mengyan Liu, Gaopeng Gou, Gang Xiong, Junzheng Shi, Zhong Guan, Hanwen Miao, and Yang Li. Enhanced detection of obfuscated HTTPS tunnel traffic using heterogeneous information network. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008077.

Wu:2025:NRE

[2613] Wenhao Wu, Xinsheng Ji, Jie Yang, Wei You, Buqing Xue, Deqiang Zhou, and Ran Li. 5G network resilience evaluation mechanism based on Petri nets. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007825.

Ma:2025:FAV

[2614] Junhao Ma, Yantao Yu, Guojin Liu, and Tiancong Huang. Freshness aware vehicular crowdsensing with multi-agent reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008107.

Zolotavkin:2025:OOA

[2615] Yevhen Zolotavkin, Yurii Baryshev, Jannik Mähn, Vitalii Lukichov, and Stefan Köpsell. Optimal obfuscation of awareness messages: Improving users’ unlinkability in intelligent transport systems. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008041.

Wan:2025:DDF

[2616] Ying Wan, Haoyu Song, Yu Jia, Yi Wang, Ling Qian, and Tian Pan. DeFlow: Differential flowlet switching for load balancing in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008120.

Wang:2025:AAP

[2617] Shanshan Wang, Chuan Xu, Guofeng Zhao, Zhenzhen Han, Rui Hu, and Shui Yu. APQA: an anonymous post quantum access authentication scheme based on lattice for space ground integrated network. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008119.

Zhan:2025:EET

[2618] Mingming Zhan, Jin Yang, Dongqing Jia, and Geyuan Fu. EAPT: an encrypted traffic classification model via adversarial pre-trained transformers. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008053.

Zhang:2025:SAL

[2619] Shuyun Zhang, Huifang Chen, and Lei Xie. Service-aware Q-learning-based routing protocol in the underwater acoustic sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008181.

Bermad:2025:PRS

[2620] Nabila Bermad, Salah Zemmoudj, Louiza Bouallouche-Medjkoune, and Lyes Khoukhi. Pseudonym revocation system for IoT-based medical applications. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008156.

Otoshi:2025:CMP

[2621] Tatsuya Otoshi and Masayuki Murata. Coordinated multi-point by distributed hierarchical active inference with sensor feedback. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008211.

Wu:2025:DRL

[2622] Jinru Wu, Ruizhong Du, and Ziyuan Wang. Deep reinforcement learning with dual-Q and Kolmogorov– Arnold networks for computation offloading in Industrial IoT. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008193.

Rozenberg:2025:USI

[2623] Guy Rozenberg and Michael Segal. Using spanners to improve network performance. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008089.

Bernardez:2025:GPG

[2624] Guillermo Bernárdez, José Suárez-Varela, Xiang Shi, Shihan Xiao, Xiangle Cheng, Pere Barlet-Ros, and Albert Cabellos-Aparicio. GraphCC: a practical graph learning-based approach to congestion control in datacenters. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008132.

Zhao:2025:FMP

[2625] Kai Zhao, Lina Wang, Fangchao Yu, Bo Zeng, and Zhi Pang. FedMP: a multi-pronged defense algorithm against Byzantine poisoning attacks in federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008223.

Liu:2025:CPS

[2626] Xiaorong Liu, Danyang Zheng, Huanlai Xing, Li Feng, Chengzong Peng, and Xiaojun Cao. A cost-provable solution for reliable in-network computingenabled services deployment. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128624008296.

Ma:2025:UIB

[2627] Huijie Ma, Yuxiang Ma, and Yulei Wu. UINT: an intent-based adaptive routing architecture. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008235.

Chaudhary:2025:PPB

[2628] Pankaj Chaudhary and Neminath Hubballi. PeNCache: Popularity based cooperative caching in named data networks. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008272.

Yang:2025:JRA

[2629] Shuhui Yang, Jian Chen, Jie Jia, Liang Guo, and Xingwei Wang. Joint resource allocation and blocklength assignment in STAR-RIS and NOMAassisted URLLC systems. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008338.

Lu:2025:AAB

[2630] Yichen Lu, Ji Qi, Sheng Zhang, Gangyi Luo, Andong Zhu, Jie Wu, and Zhuzhong Qian. ABUV: Adaptive bitrate and upsampling for video streaming on mobile devices. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008260.

Wang:2025:TCOa

[2631] Chao Wang, Danyang Zheng, Xiaorong Liu, Wenyi Tang, Honghui Xu, and Xiaojun Cao. Towards cost optimization in security-aware service function chaining and embedding over multivendor edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400834X.

Zhang:2025:NCM

[2632] Fuxin Zhang, Sihan Yao, Wei Liu, and Liang Qi. Non-cooperative multiagent deep reinforcement learning for channel resource allocation in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008387.

Albinali:2025:RAR

[2633] Hussah Albinali and Farag Azzedin.

Replay attacks in RPL-based Internet of Things: Comparative and empirical study. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008284.

Ding:2025:SES

[2634] Yuqing Ding, Zhongcheng Wu, Yongchun Miao, and Manyu Ding. SharAcc: Enhancing scalability and security in attribute-based access control with sharding-based blockchain and full decentralization. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008247.

Wu:2025:VSS

[2635] Fan Wu, Xiaolong Xu, Muhammad Bilal, Xiangwei Wang, Hao Cheng, and Siyu Wu. VEC-Sim: a simulation platform for evaluating service caching and computation offloading policies in vehicular edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400817X.

Yang:2025:HHV

[2636] Guangrong Yang, An He, Guangwei Wu, Jianing Zhao, Jinhuan Zhang, and Anfeng Liu. HVVU: a hash value verification joint UAVs scheme for trust data collection in smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008375.

Al-hamadani:2025:MRC

[2637] Ahmed Al-hamadani and Gábor Lencse. Maptperf: an RFC 8219 compliant tester for benchmarking MAP-T border relay routers. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008442.

Hydher:2025:UDW

[2638] Hassaan Hydher, Dushantha Nalin K.

Jayakody, Kasun T. Hemachandra, and Tharaka Samarasinghe. UAV deployment in WSN system for emergency/ remote area applications. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008090.

Mazloum:2025:SAP

[2639] Ali Mazloum, Ali AlSabeh, Elie Kfoury, and Jorge Crichigno. Security applications in P4: Implementation and lessons learned. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008430.

Hua:2025:ISM

[2640] Kaifeng Hua, Shengchao Su, and Yiwang Wang. Intelligent service migration for the Internet of Vehicles in edge computing: a mobility-aware deep reinforcement learning framework. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008533.

Wei:2025:DLS

[2641] Zihao Wei, Yantao Yu, Guojin Liu, and Yucheng Wu. A deep learning sparse urban sensing scheme based on spatiotemporal correlations. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008478.

Li:2025:DTM

[2642] Minyue Li, Yan Qiao, Rongyao Hu, Pei Zhao, Junjie Wang, Zhenchun Wei, Xuesen Ma, and Wenjing Li. 3DDPS: a traffic matrix estimation method based on three-dimensional diffusion posterior sampling. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008399.

Zhang:2025:IDT

[2643] Yang Zhang, Shukui Zhang, Qi Zhang, and Jianxi Fan. An intention-driven task offloading strategy based on imitation learning in pervasive edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008302.

Elhachi:2025:ERT

[2644] Hana Elhachi, Mohamed-Aymen Labiod, Farouk Boumehrez, and Salah Redadaa. Enhancing real-time mobile health video streams: a cross-layer regionof-interest based approach. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008466.

Wang:2025:NIG

[2645] Xinye Wang, Dong Zhao, Xinran Liu, Zhaoxin Zhang, and Tianzi Zhao. NeighborGeo: IP geolocation based on neighbors. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400728X.

Torres-Perez:2025:MAN

[2646] Claudia Torres-Pérez, Estefanía Coronado, Cristina Cervelló-Pastor, Javier Palomares, Estela Carmona-Cejudo, and Muhammad Shuaib Siddiqui. Minimizing active nodes in MEC environments: a distributed learningdriven framework for application placement. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008405.

Huang:2025:AAE

[2647] Longxia Huang, Changzhi Huo, Chengzhi Ge, and Mengmeng Yang. AERO: an adaptive and efficient routing for off-chain payment channel networks. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008417.

Xiong:2025:SRA

[2648] Xinning Xiong, Wenyong Wang, Ting Yang, and Gang Mao. A survey on RFID anti-collision strategies: From avoidance to recovery. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007291.

Lei:2025:FHI

[2649] Jiale Lei, Piao Jiang, Linghe Kong, Chi Xu, Chenren Xu, Kai Lin, Yueping Cai, Yanzhao Su, Weiping Ding, Zhen Wang, Bangyu Li, Xiaoguang Chen, Feng Gao, Weibo Wang, and Jiadi Yu. Fusion of heterogeneous industrial wireless networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624007618.

Bidabadi:2025:PCE

[2650] S. Bidabadi, M. A. Ouameur, M. Bagaa, D. Massicotte, F. D. Figueiredo, and A. Chaaban. Physically-consistent EM models-aware RIS-aided communication — a survey. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007953.

Sharma:2025:SEN

[2651] Adit Sharma and Arash Habibi Lashkari. A survey on encrypted network traffic: a comprehensive survey of identification/classification techniques, challenges, and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008168.

Li:2025:SVT

[2652] Jianhua Li, Bohao Feng, and Hui Zheng. A survey on VPN: Taxonomy, roles, trends and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007965.

Lang:2025:TVN

[2653] Ping Lang, Daxin Tian, Xu Han, Peiyu Zhang, Xuting Duan, Jianshan Zhou, and Victor C. M. Leung. Towards 6G vehicular networks: Vision, technologies, and open challenges. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007485.

Stypinski:2025:WEA

[2654] Miłosz Stypiński and Marcin Niemiec. Weight equalization algorithm for tree parity machines. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400820X.

Wu:2025:DBR

[2655] Shuang Wu, Zhigeng Fang, Chenchen Hua, Liangyan Tao, and Jingru Zhang. A DRL-based RAQ-GERT dynamic resource allocation algorithm considering utility for multibeam satellite system. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007722.

Korikawa:2025:PSM

[2656] Tomohiro Korikawa, Chikako Takasaki, and Kyota Hattori. A path selection method based on rule prediction in non-terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007904.

Zamacola:2025:JSP

[2657] Samuel Martínez Zamacola, Ramón Martínez Rodríguez-Osorio, and Miguel A. Salas-Natera. Joint satellite platform and constellation sizing for instantaneous beam-hopping in 5G/6G non-terrestrial networks. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007746.

Zhang:2025:MCF

[2658] Shuhang Zhang. Model collaboration framework design for space-air-ground integrated networks. Computer Networks (Amsterdam, Netherlands: 1999), 257(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008454.

Maldonado:2025:ELF

[2659] Diego Maldonado, Leonardo S. Cardoso, Juan A. Fraire, Alexandre Guitton, Oana Iova, Megumi Kaneko, and Hervé Rivano. Enhanced LR-FHSS receiver for headerless frame recovery in space-terrestrial integrated IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 257 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008508.

Anonymous:2025:Fb

[2660] Anonymous. February 2025. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2025:EBc

[2661] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S138912862500074X.

Farhadi:2025:NRB

[2662] Babak Farhadi, Parvaneh Asghari, Ebrahim Mahdipour, and Hamid Haj Seyyed Javadi. A novel recommendation-based framework for reconnecting and selecting the efficient friendship path in the heterogeneous social IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862400848X.

Wang:2025:ACE

[2663] Jie Wang, Meles G. Weldegebriel, and Neal Patwari. Augmenting channel estimation via loss field: Sitetrained Bayesian modeling and comparative analysis. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008259.

Faltelli:2025:CCN

[2664] Marco Faltelli, Giacomo Belocchi, Francesco Quaglia, and Giuseppe Bianchi. COREC: Concurrent nonblocking single-queue receive driver for low latency networking. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008144.

Li:2025:SLR

[2665] Shibao Li, Longfei Li, Wei Dou, Yunwu Zhang, Chengzhi Wang, Xuerong Cui, and Lianghai Li. SDLoRe: a loss recovery algorithm based on segment detection in lossy RDMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862400851X.

Servati:2025:EEB

[2666] Mohammad Reza Servati, Masoumeh Safkhani, Amir Masoud Rahmani, and Mehdi Hosseinzadeh. ERAS-MIS: an ECC-based robust authentication protocol suitable for medical IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007709.

Li:2025:UAD

[2667] Hongjian Li, Hengyu Li, Cheng Zhang, Dongjun Li, Xiaolin Duan, and Jiutong Liu. A UAV-assisted dynamic offloading based on maximum clique algorithm with weighted graphs in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008600.

Rahman:2025:ADI

[2668] Saifur Rahman, Shantanu Pal, Ahsan Habib, Lei Pan, and Chandan Karmakar. Attack-data independent defence mechanism against adversarial attacks on ECG signal. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008594.

Wang:2025:ODH

[2669] Liming Wang, Xuanzhi Zhao, Di Yang, Zengli Liu, Wlodek J. Kulesza, Jingmin Tang, and Wen Zhang. Optimizing DV-Hop localization through topology-based straight-line distance estimation. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008570.

Chen:2025:ZTB

[2670] Kundu Chen and Jie Luo. zkRPC:

Trustless bitcoin data availability network. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624007898.

Jin:2025:MLB

[2671] MeiYan Jin, HongBo Tang, Hang Qiu, and Jie Yang. Machine learningbased co-resident attack detection for 5G clouded environments. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008648.

Wu:2025:CWB

[2672] Weiqi Wu and Wei Gong. Concurrent WiFi backscatter communication using a single receiver in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008612.

Hao:2025:JOR

[2673] Conghui Hao, Yueyun Chen, Guang Chen, and Liping Du. Joint optimization of resource allocation, trajectory and altitude for solar-powered UAV assisted wireless charging MEC system. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008491.

Xie:2025:SAM

[2674] Jiang Xie, Shuhao Li, Xiaochun Yun, Chengxiang Si, and Tao Yin. Sample analysis and multi-label classification for malicious sample datasets. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008314.

Han:2025:EFM

[2675] Yifan Han, He Huang, Yu-E Sun, Jia Liu, and Shigang Chen. Expiration filter: Mining recent heavy flows in high-speed networks. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008429.

Barbosa:2025:FFC

[2676] Matheus T. M. Barbosa, Eric B. C. Barros, Vinícius F. S. Mota, Dionisio M. Leite Filho, Leobino N. Sampaio, Bruno T. Kuehne, Bruno G. Batista, Damla Turgut, and Maycon L. M. Peixoto. FOCCA: Fog-cloud continuum architecture for data imputation and load balancing in smart grids. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008636.

Zhang:2025:PLS

[2677] Peng Zhang and Qing Guo. Physical layer security in cognitive terrestrialsatellite networks with LEO relaying. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000015.

Almeida:2025:EDS

[2678] Ana Almeida, Pedro Rito, Susana Brás, Filipe Cabral Pinto, and Susana Sargento. Exploring the dynamic symbiosis of urban mobility and 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008569.

Wu:2025:ORM

[2679] Jia Wu, Tingyi Dai, Peiyuan Guan, Ziru Chen, Fangfang Gou, and Amir Taherkordi. Opportunistic routing for mobile edge computing: a community detected and task priority aware approach. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008326.

Zhao:2025:PSR

[2680] Yanyang Zhao, Zhengxiong Luo, Kai Liang, Feifan Wu, Wenlong Zhang, Heyuan Shi, and Yu Jiang. Protocol syntax recovery via knowledge transfer. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008545.

Sun:2025:TSC

[2681] Yunqi Sun, Hesheng Sun, Tuo Cao, Mingtao Ji, Zhuzhong Qian, Lingkun Meng, Dongxu Wang, and Xiangyu Li. Towards strong continuous consistency in edge-assisted VR-SGs: Delaydifferences sensitive online task redistribution. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008351.

Ma:2025:MMS

[2682] Haopeng Ma, Xiaoying Huang, Ke Ruan, Zehua Hu, and Yongqing Zhu. MST-FCAN: Multiscale sparse temporalfrequency cross attention network for traffic prediction. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000039.

Rashid:2025:LOS

[2683] Fariza Rashid, Nishavi Ranaweera, Ben Doyle, and Suranga Seneviratne. LLMs are one-shot URL classifiers and explainers. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008363.

Wang:2025:DBL

[2684] Jieru Wang, Hui Xia, Lijuan Xu, Rui Zhang, and Kunkun Jia. DRLbased latency-energy offloading optimization strategy in wireless VR networks with edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000027.

Long:2025:ELT

[2685] Yuxuan Long, Zhenyu Wang, Shizhan Lan, Rui Zhang, and Kai Xu. Energylatency tradeoff for task offloading and resource allocation in vehicular edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008582.

Zhai:2025:TOM

[2686] Linbo Zhai, Ping Zhao, Kai Xue, Yumei Li, and Chen Cheng. Task offloading and multi-cache placement in multi-access mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008624.

Yoon:2025:DDA

[2687] Jonghun Yoon, Yunbae Kim, Hyeyeon Kwon, and Seungkeun Park. Datadriven active session identification for LTE user-perceived QoS analysis. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000106.

Yang:2025:EES

[2688] Yang Yang, Jiaxing Zhang, Yanjiao Chen, Qing Huang, Fei Chen, and Jing Chen. An efficient encrypted search with owner-level and attributelevel access controls. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000076.

Wang:2025:MVD

[2689] Wenming Wang, Zhiquan Liu, Lingyan Xue, Haiping Huang, and Nageswara Rao Lavuri. Malicious vehicle detection scheme based on UAV and vehicle cooperative authentication in vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000052.

Pupo:2025:MRM

[2690] Ernesto Fontes Pupo, Claudia Carballo González, Eneko Iradier, Jon Montalban, Pablo Angueira, and Maurizio Murroni. Multi-rate multicasting aided NOMA for addressing the multiuser diversity in beyond 5G. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500009X.

Sciddurlo:2025:MCB

[2691] Giancarlo Sciddurlo, Pietro Camarda, Domenico Striccoli, Ilaria Cianci, Giuseppe Piro, and Gennaro Boggia. Markov chain-based analytical model supporting service provisioning and network design in the social Internet of everything. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000088.

Zhang:2025:SST

[2692] Zheng Zhang, Quan Ren, Jie Lu, Yuxiang Hu, and Hongchang Chen. ES-SDPC: a secure and trusted SDP framework. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000064.

Linn:2025:HMM

[2693] Nyi Nyi Linn, Yaodong Ma, Kai Liu, and Xiangfen Wang. A hybrid multichannel MAC protocol with low local delay for vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000374.

Haque:2025:BRW

[2694] Khandaker Foysal Haque, Milin Zhang, Francesca Meneghello, and Francesco Restuccia. BeamSense: Rethinking wireless sensing with MU-MIMO Wi-Fi beamforming feedback. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128624008521.

Zhu:2025:TES

[2695] Deqing Zhu, Lidong Wang, Genmei Pan, and Shenji Luan. Two enhanced schemes for coordinated spatial reuse in IEEE 802.11be: Adaptive and distributed approaches. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000283.

Yagcioglu:2025:MLB

[2696] Mert Yagcioglu. Machine learning based dynamic resource sharing and frequency reuse in 5G hetnets with dronecells. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000143.

Wang:2025:DMR

[2697] Kewei Wang, Changzhen Hu, and Chun Shan. Defining and measuring the resilience of network services. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000040.

Wang:2025:CCF

[2698] Shi Wang, Dayan Cao, Xiaoying Zhu, Han Jiang, and Mingyu Wang. A congestion control framework for heterogeneous storage structure IoT node. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500026X.

Xiao:2025:LBR

[2699] Linjie Xiao, Shining Li, Qin Wen, Xiao Liang, Yiming Li, Wanbao Wang, and Yuntao Fu. Load balancing routing algorithm of industrial wireless network for digital twin. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000271.

Qian:2025:FDR

[2700] Zhen Qian, Guanghui Li, Tao Qi, and Chenglong Dai. Federated deep reinforcement learning-based costefficient proactive video caching in energy-constrained mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 258 (??):??, February 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000301.

Belenguer:2025:RFL

[2701] Aitor Belenguer, Jose A. Pascual, and Javier Navaridas. A review of federated learning applications in intrusion detection systems. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128624008557.

Naik:2025:EIS

[2702] Azmera Chandu Naik, Lalit Kumar Awasthi, Priyanka R., T. P. Sharma, and Aryan Verma. Enhancing IoT security: a comprehensive exploration of privacy, security measures, and advanced routing solutions. Computer Networks (Amsterdam, Netherlands: 1999), 258(??):??, February 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000131.

Anonymous:2025:Ma

[2703] Anonymous. March 2025. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2025:EBd

[2704] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001033.

Huang:2025:FDC

[2705] Hui Huang, Di Xiao, Mengdi Wang, Jia Liang, Min Li, Lvjun Chen, and Yanqi Liu. Fog-driven communicationefficient and privacy-preserving federated learning based on compressed sensing. Computer Networks (Amsterdam, Netherlands: 1999), 259 (??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000118.

Ni:2025:UTS

[2706] Zihan Ni, Haoxing Zhang, Min Yan, Xiaqing Miao, Gaofeng Pan, Shuai Wang, and Jianping An. On the uplink transmission of satellite-aerial FSO links in presence of random optical interference. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000350.

Chang:2025:ODS

[2707] Ben-Jye Chang and Yu-Ting Lin. Optimizing differentiated SFCs of slicing using virtual/physical network resources for SDN-based flowing and steering in B5G/5G networking. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000337.

Zhang:2025:EDH

[2708] ZunDong Zhang, Yanan Cheng, Haiyan Xu, Zhaoxin Zhang, and Chao Li. Efficient DNS over HTTPS servers discovery method: a voting-based stacked ensemble model with secure connection metadata. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000416.

Fu:2025:LAD

[2709] Xiuwen Fu, Chang Deng, and Antonio Guerrieri. Low-AoI data collection in integrated UAV-UGV-assisted IoT systems based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500012X.

Wang:2025:RPP

[2710] Xiaoming Wang, Zhiquan Liu, and Binrui Huang. Robust and privacypreserving federated learning scheme based on ciphertext-selected users. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128625000404.

Zhang:2025:COD

[2711] Li Zhang, Zishuai Guo, Haoru Su, and Wei Zhao. CAAST: Optimizing data communications in satellite networks through cache and anycast. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000349.

Mazzenga:2025:VAM

[2712] Franco Mazzenga, Alessandro Vizzarri, and Romeo Giuliano. Virtual assistant mediated communications for radio resources saving. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000362.

Zhu:2025:DBD

[2713] Difeng Zhu, Yafei Zhang, Xuan Gong, Hai miao Hu, and Yi Gao. DTS-BWpredictor: Dual-scale temporal strategy based bandwidth prediction in highly dynamic links. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000398.

Chandran:2025:CLP

[2714] Indu Chandran and Kizheppatt Vipin. A cross-layer PUF-secured energy and congestion-aware on-demand routing for multi-UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000441.

P:2025:DNS

[2715] Vidhya P., Subashini K., Sathishkannan R., and Gayathri S. Dynamic network slicing based resource management and service aware Virtual Network Function (VNF) migration in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000325.

Zhang:2025:PPM

[2716] Guiqi Zhang, Jun Shen, Jiangtao Li, Wutao Qin, and Yufeng Li. Payload processor: Message authentication for in-vehicle CAN bus using data compression and tag filling. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000295.

Kavitha:2025:KBA

[2717] M. Joselin Kavitha, M. R. Geetha, and R. Isaac Sajan. Knowledgebased adaptive routing for energy efficiency and attack detection in ad hoc wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000544.

Sinha:2025:DES

[2718] Mitali Sinha, Padmalochan Bera, Manoranjan Satpathy, Kshira Sagar Sahoo, and Joel J. P. C. Rodrigues. DDoSBlocker: Enhancing SDN security with time-based address mapping and AI-driven approach. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000465.

Wang:2025:TSD

[2719] Fei Wang, Lei Jiao, Konglin Zhu, Lingjun Pu, and Lin Zhang. Toward sustainable diffusion-based AIGC: Design and online orchestration in distributed edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000453.

Li:2025:ECT

[2720] Binglei Li, Tong Liu, Jian Ma, Jingwei Dai, and Peizhen Liao. Evolving chain-of-thought and automatic text annotation based intent classification method for wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000428.

Ye:2025:SEC

[2721] Feixiang Ye and Wanqing Wu. A security-enhanced certificateless conditional privacy-preserving authentication scheme with collusion-resistance for IoV. Computer Networks (Amsterdam, Netherlands: 1999), 259 (??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000520.

Zhang:2025:CBR

[2722] Yuchen Zhang, Jinyong Chang, Yangxi Chen, and Ronghua Xu. Certificatebased remote auditing protocol with privacy protection and deduplication functions for cloud-assisted applications. Computer Networks (Amsterdam, Netherlands: 1999), 259 (??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000519.

Liu:2025:FFS

[2723] Meihui Liu, Renhe Yan, Shihui Duan, Jinyu Zhu, Fangmin Xu, and Chenglin Zhao. Flexible flow scheduling for industrial TSN: a hierarchical factory network scheduling approach. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000489.

Vallero:2025:TBN

[2724] Greta Vallero, Michela Meo, Wout Joseph, and Margot Deruyck. Thresholdbased 5G NR base station management for energy saving. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000386.

Liu:2025:PLS

[2725] Zhixin Liu, Haiyang Cao, Jiawei Su, Yazhou Yuan, and Xinping Guan. Physical layer security in double RIS-aided WPCN systems based on non-cooperative game. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500043X.

Wu:2025:PPT

[2726] Dongkuo Wu, Xingwei Wang, Xueyi Wang, Siyu Yan, Qiang He, and Min Huang. Privacy-preserving and truthful auction-based resource allocation mechanisms for task offloading in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000477.

Sun:2025:MAB

[2727] Hao Sun, Juntao Fei, Guoping Zhang, and Mande Xie. A multi-agent-based dynamic charging strategy for UAVassisted wireless rechargeable sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000490.

Liu:2025:PVC

[2728] Zhimin Liu, Shukun Liu, Weijia Jia, Guojun Wang, and Mingfeng Su. Prediction of k-view coverage for objects in camera sensor networks under complex environments. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000532.

Yilmaz:2025:NHA

[2729] Ali Yilmaz and Resul Das. A novel hybrid approach combining GCN and GAT for effective anomaly detection from firewall logs in campus networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000507.

Bai:2025:GRG

[2730] Luxin Bai, Hailong Ma, Yiming Jiang, Zinuo Yin, Huiqing Wan, and Hongguang Wang. GRL-RR: a graph reinforcement learning-based resilient routing framework for software-defined LEO mega-constellations. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500057X.

Wang:2025:RRD

[2731] Mingyao Wang, Keqiang He, Peirui Cao, Jiong Duan, Dongliang Lv, Zehao Yu, Yanqing Chen, Chengyuan Huang, Wanchun Dou, Guihai Chen, and Chen Tian. Reunion: Receiverdriven network load balancing mechanism in AI training clusters. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000568.

Akyildiz:2025:MTL

[2732] Oguzhan Akyildiz, Feyza Yildirim Okay, Ibrahim Kök, and Suat Özdemir. Mx-TORU: Location-aware multi-hop task offloading and resource optimization protocol for connected vehicle networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000623.

Xiong:2025:PED

[2733] Chenxi Xiong, Ting Yang, Yu Wang, and Bing Dong. PoVF: Empowering decentralized blockchain systems with verifiable function consensus. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500060X.

Wubete:2025:MLA

[2734] Bruck W. Wubete, Babak Esfandiari, and Thomas Kunz. Machine learning approaches for predicting link failures in production networks. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000660.

Wang:2025:CET

[2735] Jingzhou Wang, Gongming Zhao, Hongli Xu, Chunming Qiao, and He Huang. A cost-efficient traffic engineering framework with various pricing schemes in clouds. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000581.

Jin:2025:FDC

[2736] Zhigang Jin, Yu Ding, Xiaodong Wu, Xuyang Chen, Zepei Liu, and Gen Li. Federated dual correction intrusion detection system: Efficient aggregation for heterogeneous data. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000842.

Sun:2025:RFL

[2737] Haibin Sun and Shuai Yang. Rangefree localization algorithm based on modified distance and improved blackwinged kite algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000593.

Wermann:2025:KSK

[2738] Alexandre Gustavo Wermann and Juliano Araujo Wickboldt. KTWIN: a serverless Kubernetes-based Digital Twin platform. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000635.

Martian:2025:TOR

[2739] Alexandru Martian, Razvan-Florentin Trifan, Teodora-Cristina Stoian, Marius-Constantin Vochin, and Frank Y. Li. Towards Open RAN in beyond 5G networks: Evolution, architectures, deployments, spectrum, prototypes, and performance assessment. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000556.

Alekseeva:2025:EER

[2740] Daria Alekseeva, Anzhelika Mezina, Radim Burget, Otso Arponen, Elena Simona Lohan, and Aleksandr Ometov. Enhancing extended reality assisted surgery through a field-of-view video delivery optimization. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000611.

Choquenaira-Florez:2025:RML

[2741] Alexander Y. Choquenaira-Florez, Juan A. Fraire, Hannah B. Pasandi, and Hervé Rivano. On the role of machine learning in satellite Internet of Things: a survey of techniques, challenges, and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 259(??):??, March 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000313.

Anonymous:2025:Aa

[2742] Anonymous. April 2025. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Anonymous:2025:EBe

[2743] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001367.

Liu:2025:SDF

[2744] Chao Liu, Fulong Chen, Taochun Wang, Chuanxin Zhao, Dong Xie, and Peng Hu. A secure and dynamic fusion addressing scheme for Internet of Vehicles scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000805.

Lai:2025:ACN

[2745] Yurong Lai, Xinhui Han, Xueer Wang, Tse-Tin Chan, and Haoyuan Pan. Age of collection with non-orthogonal multiple access: a theoretical-plus-experimental study. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000702.

Luo:2025:IMJ

[2746] Gaoyu Luo, Shanhao Zhan, Chenyi Liang, Zhibin Gao, Yifeng Zhao, and Lianfen Huang. An incentive mechanism for joint sensing and communication vehicular crowdsensing by deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000672.

Esper:2025:TRP

[2747] João Paulo Esper, Luciano de S. Fraga, Aline C. Viana, Kleber Vieira Cardoso, and Sand Luz Correa. +Tour: Recommending personalized itineraries for smart tourism. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000866.

Luo:2025:ESC

[2748] Wei Luo, Ziyi Lv, Chengzhe Lai, and Tengfei Yang. Efficient and secure cross-domain data sharing scheme with traceability for Industrial Internet. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000854.

Chen:2025:LDM

[2749] Zhen Chen, Jianqiang Yu, Shuang Fan, Jing Zhao, and Dianlong You. Latent diffusion model-based data poisoning attack against QoS-aware cloud API recommender system. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500088X.

Wei:2025:EDE

[2750] Qianjin Wei, Xuanjing Wu, Zhiquan Liu, and Gang Rao. EVFL-DCs: Enhancing verifiability of federated learning by double commitments based on blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000659.

Xiao:2025:PBG

[2751] Siyi Xiao, Lejun Zhang, Zhihong Tian, Shen Su, Jing Qiu, and Ran Guo. Pheromone-based graph embedding algorithm for Ethereum phishing detection. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500091X.

Arat:2025:SAR

[2752] Ferhat Arat and Sedat Akleylek.

Security-aware RPL: Designing a novel objective function for risk-based routing with rank evaluation. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000908.

Fadi:2025:SCA

[2753] Oumaima Fadi, Adil Bahaj, Karim Zkik, Abdellatif El Ghazi, Mounir Ghogho, and Mohammed Boulmalf. Smart contract anomaly detection: the contrastive learning paradigm. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000891.

Omri:2025:SEE

[2754] Aymen Omri, Javier Hernandez Fernandez, and Roberto Di Pietro. A spectral and energy efficient transmission scheme for OFDM-based communication systems. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000696.

Gao:2025:GTB

[2755] Tianqi Gao, Yuanzhi Ni, Hongfeng Tao, Zhuocheng Du, and Zhenshu Zhu. Game theory-based vehicle selection and channel scheduling for federated learning in vehicular edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000799.

Souza:2025:DRA

[2756] Caio Souza, Marcos Falcão, Andson Balieiro, Elton Alves, and Tarik Taleb. Dynamic resource allocation for URLLC and eMBB in MEC-NFV 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000957.

Moslehi:2025:ECS

[2757] Mari M. Moslehi. Exploring coverage and security challenges in wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 260(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000647.

Yang:2025:IAA

[2758] Wei Yang, Qianyi Wang, and Yu Yao.

IPv6 active address detection model based on diffusion model. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000155.

Singh:2025:EEA

[2759] Parvinder Singh and Rajinder Vir. Enhanced energy-aware routing protocol with mobile sink optimization for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000684.

Alkadamani:2025:SET

[2760] Mohamad Alkadamani, Colin Brown, Kareem Baddour, Mathieu Châteauvert, Janaki Parekh, and Adrian Florea. Spectrum efficiency through data: a methodology for evaluating local licensing strategies. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000830.

Wang:2025:TWT

[2761] Xiaolin Wang, Jinglong Zhang, Xuanzhao Lu, Fangfei Li, Cailian Chen, and Xinping Guan. Towards wireless time-sensitive networking: Multilink deterministic scheduling via deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000878.

Xu:2025:BCC

[2762] Shuang Xu, Xutong Li, Jiayu Zhang, Feng Li, Yuanyuan Liang, and Shufeng Hao. Bilateral collaborative computing offloading via LEO satellites for remote network applications. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000921.

Fan:2025:FEP

[2763] Tingyu Fan, Xiaojun Chen, Xudong Chen, Ye Dong, Weizhan Jing, and Zhendong Zhao. FedShelter: Efficient privacy-preserving federated learning with poisoning resistance for resourceconstrained IoT network. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000933.

Wang:2025:SGA

[2764] Yulei Wang, Li Feng, Yalin Liu, and Zhongjie Li. Stochastic geometry analysis for information integration and communication in cellular and D2Dbased heterogeneous IoT. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000945.

Li:2025:EDT

[2765] Binqi Li, Yuan Zhu, Xiangxi Yao, Chenming Jiang, Ke Lu, and Zhipeng Sun. Enabling deterministic transmission for DDS by leveraging IEEE 802.1Qbv time-sensitive networking. Computer Networks (Amsterdam, Netherlands: 1999), 261 (??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000969.

Zhao:2025:AAI

[2766] Zirun Zhao, Zhaowen Lin, and Yi Sun. ADPF: Anti-inference differentially private protocol for federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000982.

Zhu:2025:FGC

[2767] Changpeng Zhu, Bincheng Fan, Nan Xiang, Bo Han, and Tian Zhou. A fine-grained compression scheme for block transmission acceleration over IPFS network. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625000994.

Yuan:2025:DPF

[2768] Jiangyong Yuan and Yong Chen. Differential private federated learning with per-sample adaptive clipping and layer-wise gradient perturbation. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001070.

Borah:2025:SDA

[2769] Abinash Borah, Anirudh Paranjothi, and Johnson P. Thomas. A survey on distributed approaches for security enhancement in vehicular adhoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001082.

Basaran:2025:XEI

[2770] Osman Tugay Basaran and Falko Dressler. XAInomaly: Explainable and interpretable Deep Contractive Autoencoder for O-RAN traffic anomaly detection. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001136.

Mishra:2025:NSA

[2771] Debashisha Mishra, Emiliano Traversi, Angelo Trotta, Prasanna Raut, Boris Galkin, Marco Di Felice, and Enrico Natalizio. Network slicing in aerial base station (UAV-BS) towards coexistence of heterogeneous 5G services. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001148.

Ahmed:2025:MAM

[2772] Mukhtar Ahmed, Jinfu Chen, Ernest Akpaku, and Rexford Nii Ayitey Sosu. MTCR-AE: a multiscale temporal convolutional recurrent autoencoder for unsupervised malicious network traffic detection. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500115X.

Far:2025:NLF

[2773] Saeed Banaeian Far, Mohammad Reza Chalak Qazani, Seyed Mojtaba Hosseini Bamakan, Azadeh Imani Rad, and Ahad Zareravasan. A novel layer 2 framework for breaking the blockchain trilemma problem using MPC-in-the-Head. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001161.

Zhao:2025:HAC

[2774] Guosheng Zhao, Zhiwen Li, and Jian Wang. Hybrid attacks collaborative defense model using an ensemble honey badger algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001173.

Meddeb:2025:QIB

[2775] Aref Meddeb. Quantum Internet building blocks state of research and development. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001197.

Shariq:2025:DPS

[2776] Mohd Shariq, Ismail Taha Ahmed, Mehedi Masud, Aymen Dia Eddine Berini, and Norziana Jamil. Design of a provable secure lightweight privacy-preserving authentication protocol for autonomous vehicles in IoT systems. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001239.

Khaleel:2025:CCE

[2777] Mustafa Ibrahim Khaleel. Collaborative cloud-edge task scheduling scheme in the networked UAV Internet of Battlefield Things (IoBT) territories based on deep reinforcement learning model. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001240.

Anonymous:2025:EBf

[2778] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001616.

Anonymous:2025:Ab

[2779] Anonymous. April 2025. Computer Networks (Amsterdam, Netherlands: 1999), 261(??):??, April 2025.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Dong:2025:LBP

[2780] Jisheng Dong, Qingni Shen, Junkai Liang, Cong Li, Xinyu Feng, and Yuejian Fang. A lattice-based privacypreserving decentralized multi-party payment scheme. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625000970.

Le:2025:DFT

[2781] Thi-Thu-Huong Le, Shinwook Heo, Jaehan Cho, and Howon Kim. DDoS-BERT: Fine-tuning variant text classification bidirectional encoder representations from transformers for DDoS detection. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001185.

Li:2025:DPS

[2782] Xianglong Li, Kaiwei Mo, Zongpeng Li, and Hong Xu. Dynamic pricing and scheduling in LEO satellite networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001203.

Nandyala:2025:AAI

[2783] Chandra Sukanya Nandyala and Ho-Shin Cho. AUV-aided isolated subnetwork prevention for reliable data collection by underwater wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001227.

Du:2025:UCH

[2784] Ying Du, Yilong Li, Pu Cheng, Zhijie Han, and Yanan Wang. UGL: a comprehensive hybrid model integrating GCN and LSTM for enhanced intrusion detection in UAV controller area networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001252.

Calagna:2025:EEC

[2785] Antonio Calagna, Stefano Ravera, and Carla Fabiana Chiasserini. Enabling efficient collection and usage of network performance metrics at the edge. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001264.

Ghosh:2025:SFD

[2786] Subhendu Ghosh and Aneek Adhya.

Soft failure detection and identification in optical networks using cascaded deep learning model. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001276.

Alavizadeh:2025:SNB

[2787] Hooman Alavizadeh, Ahmad Salehi S., A. S. M. Kayes, Wenny Rahayu, and Tharam Dillon. Social network botnet attack mitigation model for cloud. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001288.

Yin:2025:CIC

[2788] Zinuo Yin, Hongchang Chen, Hailong Ma, Tao Hu, and Luxin Bai. CAEAID: an incremental contrast learning-based intrusion detection framework for IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500129X.

Loevenich:2025:DEA

[2789] Johannes Loevenich, Erik Adler, Tobias Hürten, and Roberto Rigolin F. Lopes. Design and evaluation of an Autonomous Cyber Defence agent using DRL and an augmented LLM. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001306.

Liang:2025:JOV

[2790] Junbin Liang and Qiao He. Joint optimization of VNF deployment and UAV trajectory planning in multi-UAV-enabled mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001318.

Wang:2025:HLP

[2791] Zekang Wang, Fuxiang Yuan, Ruixiang Li, Meng Zhang, and Xiangyang Luo. Hidden AS link prediction based on random forest feature selection and GWO-XGBoost model. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500132X.

Cai:2025:MMA

[2792] Yongqing Cai, Dianjie Lu, Jing Chen, and Guijuan Zhang. MAID: Mobility-aware information dissemination in mobile IoT using temporal point processes. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001410.

Toopchinezhad:2025:MLA

[2793] Mohammad Parsa Toopchinezhad and Mahmood Ahmadi. Machine learning approaches for active queue management: a survey, taxonomy, and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001422.

Zhang:2025:SCR

[2794] Lianming Zhang, Shuqiang Peng, and Pingping Dong. SK-CFR: Rerouting critical flows through discrete soft actor–critic within the KP-GNN framework. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001434.

Najjar:2025:DRW

[2795] Salma Najjar, Michael David, William Derigent, and Ahmed Zouinkhi. Dynamic reconfiguration of wireless sensor networks: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001446.

Pinto:2025:RID

[2796] Daniela Pinto, Ivone Amorim, Eva Maia, and Isabel Praça. A review on intrusion detection datasets: tools, processes, and features. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001458.

Gulzari:2025:SOC

[2797] Usman Ali Gulzari, Waqar Farooq, Syed Nasir Mehmood Shah, Naveed Ahmad, Hessam Sarjoughian, Iftikhar Ahmed Khan, and Sheraz Anjum. The Scalable Octagonal-Cross-By-Pass-Torus topology for the on-chip-communication. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500146X.

Liu:2025:DTD

[2798] Wei Liu, Wentao Cui, Bin Wang, Heng Pan, Wei She, and Zhao Tian. Decentralized traffic detection utilizing blockchain-federated learning with quality-driven aggregation. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001471.

Mustafa:2025:DRL

[2799] Ehzaz Mustafa, Junaid Shuja, Faisal Rehman, Abdallah Namoun, Muhammad Bilal, and Kashif Bilal. Deep Reinforcement Learning and SQP-driven task offloading decisions in vehicular edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001483.

Perez-Gomariz:2025:LHN

[2800] Mario Pérez-Gomariz, Fernando Cerdán-Cartagena, and Jess García. LM-Hunter: an NLP-powered graph method for detecting adversary lateral movements in APT cyberattacks at scale. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001495.

Cruz:2025:AAB

[2801] E. M. Cruz, J. R. D. S. Júnior, Y. H. J. Souza, G. L. S. S. Jesus, and M. L. M. Peixoto. ArchW3: an adaptive blockchain wallet architecture for Web3 applications. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001501.

Zhu:2025:PHP

[2802] Fei Zhu, Lin You, Jixiang Wang, and Lei Li. P-HotStuff: Parallel BFT algorithm with throughput insensitive to propagation delay. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001513.

Akpaku:2025:RDU

[2803] Ernest Akpaku, Jinfu Chen, Mukhtar Ahmed, Francis Kwadzo Agbenyegah, and William Leslie Brown-Acquaye. RAGN: Detecting unknown malicious network traffic using a robust adaptive graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001525.

Yilmaz:2025:EIS

[2804] Selim Yılmaz. Enhancing IoT security:

a competitive coevolutionary strategy for detecting RPL attacks in challenging attack environments. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001537.

Yu:2025:YSP

[2805] Qianhui Yu, Hai Xue, Celimuge Wu, Ya Liu, and Wunan Guo. Yardstick-Stackelberg pricing-based incentive mechanism for federated learning in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001549.

Kang:2025:TBB

[2806] Yunchuan Kang, Houbing Herbert Song, Tian Wang, Shaobo Zhang, Mianxiong Dong, and Anfeng Liu. A trust and bundling-based task allocation scheme to enhance completion rate and data quality for mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001574.

Fathalizadeh:2025:SFO

[2807] Amir Fathalizadeh, Vahideh Moghtadaiee, and Mina Alishahi. A survey and future outlook on indoor location fingerprinting privacy preservation. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001677.

Jang:2025:CAM

[2808] Woo-Hyeok Jang and Seung-Jae Han. Collision avoidance by mitigating uncertain packet loss in multi-hop wireless IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001732.

Yang:2025:GLT

[2809] Fang Yang, Tao Ma, Chunlai Ma, Nina Shu, Chao Chang, Chunsheng Liu, Tao Wu, and Xingkui Du. GenNP: a lowthreshold and powerful network performance data generator. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500177X.

Yu:2025:QRR

[2810] Xiaoling Yu, Yuntao Wang, and Xin Huang. Quantum-resistant ring signature-based authentication scheme against secret key exposure for VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001811.

Anonymous:2025:EBg

[2811] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002178.

Anonymous:2025:Mb

[2812] Anonymous. May 2025. Computer Networks (Amsterdam, Netherlands: 1999), 262(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Tsai:2025:CNO

[2813] Meng-Jung Chloe Tsai and Leonard Kleinrock. Computer network optimization using the power metric for multiple flows: Part I. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001215.

Singh:2025:DDC

[2814] Saurabh Singh. Discrete dimensional coordinate scheme for distributed aggressive forwarding in mobile ad-hoc networks: an effective approach. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001409.

Moeinaddini:2025:NAS

[2815] Elham Moeinaddini, Eslam Nazemi, and Amin Shahraki. A new approach on self-adaptive trust management for social Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001550.

Islam:2025:DIA

[2816] Taimoor Ul Islam, Joshua Ofori Boateng, Md Nadim, Guoying Zu, Mukaram Shahid, Xun Li, Tianyi Zhang, Salil Reddy, Wei Xu, Ataberk Atalar, Vincent Lee, Yung-Fu Chen, Evan Gossling, Elisabeth Permatasari, Christ Somiah, Owen Perrin, Zhibo Meng, Reshal Afzal, Sarath Babu, Mohammed Soliman, Ali Hussain, Daji Qiao, Mai Zheng, Ozdal Boyraz, Yong Guan, Anish Arora, Mohamed Y. Selim, Arsalan Ahmad, Myra B. Cohen, Mike Luby, Ranveer Chandra, James Gross, Kate Keahey, and Hongwei Zhang. Design and implementation of ARA wireless living lab for rural broadband and applications. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001562.

Sun:2025:ODK

[2817] Shangyi Sun, Rui Zhang, Ming Yan, and Jie Wu. Offloading distributed key-value stores with offpath SmartNICs. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001653.

Xu:2025:SBM

[2818] Xiaolong Xu, Xuanhong Zhou, Muhammad Bilal, Sherali Zeadally, Jon Crowcroft, Lianyong Qi, and Shengjun Xue. Socially beneficial metaverse: Framework, technologies, applications, and challenges. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001665.

Kim:2025:RBF

[2819] Hyowon Kim, Gabin Heo, and Inshil Doh. Role-based federated learning exploiting IPFS for privacy enhancement in IoT environment. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001689.

Shi:2025:SMC

[2820] Yayong Shi, Haiyan Zhu, and Rui Wang. Sparse mobile crowd sensing in low-resource computing environments: an active spatio-temporal fracture approach. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001690.

Bel:2025:SQN

[2821] Oceane Bel and Mariam Kiran. Simulators for quantum network modeling: a comprehensive review. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001720.

Liang:2025:SSA

[2822] Xingyu Liang, Sen Wang, Ling Xiong, Zhicai Liu, and Neal N. Xiong. SAGTM: a secure authentication scheme with traceability for avatars in the Metaverse. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001744.

Zhai:2025:III

[2823] Jiqiang Zhai, Xinyu Wang, Zhonghui Zhai, Tao Xu, Zuming Qi, and Hailu Yang. Industrial IoT intrusion attack detection based on composite attention-driven multi-layer pyramid features. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001756.

Chu:2025:DCO

[2824] Zhiming Chu, Guyue Li, Qingchun Meng, Haobo Li, and Yuwei Zeng. Defeating CSI obfuscation mechanisms: a study on unauthorized Wi-Fi sensing in wireless sensor network. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001768.

Yuan:2025:TRC

[2825] Hao Yuan, Ruozhe Li, Laihe Wang, Bangbang Ren, Tao Chen, and Deke Guo. Toward resilient communication architecture: Online network reconfiguration for UAV failure. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001781.

Ai:2025:ATG

[2826] Yintan Ai, Hua Li, Hongwei Ruan, Hanlin Liu, and Xianrong Wang. Achieving topology generalization without retraining in service function chaining through cooperative multiagent reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001793.

Goswami:2025:LAK

[2827] Chandan Goswami, Aniket Basak, Rajeet Ghosh, Avishek Adhikari, and Pinaki Sarkar. Lightweight authenticated key agreement scheme for IoMT network using generalized Chinese Remainder Theorem. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500180X.

Wang:2025:BBMb

[2828] Jifang Wang, Shangping Wang, Duo Zhang, Qian Zhang, and Yinjuan Deng. Blockchain-based multiple auditing scheme against cheating owner in clouds. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001823.

Zhou:2025:LMI

[2829] Wenjie Zhou, Linbo Zhai, Zekun Lu, Kai Xue, and Tian Zhang. Latency minimization in IRS-UAV assisted WPT-MEC systems: an ID-AOPDDQN-based trajectory and phase shift optimization approach. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128625001835.

Chebbi:2025:ERA

[2830] Samar Chebbi, Oussama Habachi, Jean-Pierre Cances, Vahid Meghdadi, and Essaid Sabir. Efficient resource allocation in 5G massive MIMO-NOMA networks: Comparative analysis of SINR-aware power allocation and spatial correlation-based clustering. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001847.

Shi:2025:BBA

[2831] Yijie Shi, Kai Fan, Yuhan Bai, Chonglin Zhang, Kuan Zhang, Hui Li, and Yintang Yang. Blockchainbased anonymous and self-tallying voting with time-bounded ballot secrecy. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S1389128625001859.

Huang:2025:ETS

[2832] Mingfeng Huang, Qing Peng, Xiaoyu Zhu, Tan Deng, Ronghui Cao, and Wenzheng Liu. Ensuring trustworthy and secure IoT: Fundamentals, threats, solutions, and future hotspots. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001860.

Poyda:2025:ESU

[2833] Arseny Poyda, Andrey Tyarin, Kirill Glinskiy, Ilya Burtakov, Aleksey Kureev, and Evgeny Khorov. Experimental study of a user-centric RIS in existing cellular systems. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001872.

Xing:2025:MVF

[2834] Rui Xing, Zhenzhe Zheng, Qinya Li, Fan Wu, and Guihai Chen. MI-VFL: Feature discrepancy-aware distributed model interpretation for vertical federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001884.

Zhai:2025:JTO

[2835] Linbo Zhai, Zekun Lu, Jiande Sun, and Xiaole Li. Joint task offloading and computing resource allocation with DQN for task-dependency in multi-access edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001902.

Fan:2025:MAI

[2836] Na Fan, Yuxin Gao, Jialong Li, Zhiquan Liu, and Wenjun Fan. Multiattack identification and mitigation mechanism based on multi-agent collaboration in vehicular named data networking. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500194X.

Long:2025:DBS

[2837] Saiqin Long, Jianfeng Yang, Chongxi Rao, Qingyong Deng, and Kun Cao. D3QN-based secure scheduling of microservice workflows in cloud environments. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001951.

Cerasuolo:2025:AAN

[2838] Francesco Cerasuolo, Giampaolo Bovenzi, Domenico Ciuonzo, and Antonio Pescapè. Attack-adaptive network intrusion detection systems for IoT networks through class incremental learning. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001963.

Xie:2025:SPW

[2839] Lixia Xie, Hao Zhang, Hongyu Yang, Ze Hu, and Xiang Cheng. A scalable phishing website detection model based on dual-branch TCN and mask attention. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001987.

Tolba:2025:DEA

[2840] Bassant Tolba, Maha Elsabrouty, Mohammed Abo-Zahhad, Akira Uchiyama, and Ahmed H. Abd El-Malek. Delayenergy-aware joint multi-cell association, service caching, and task offloading in hybrid-task heterogeneous edge computing networks. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625001999.

Chen:2025:MAD

[2841] Dongyu Chen, Tao Deng, Juncheng Jia, Siwei Feng, and Di Yuan. Mobilityaware decentralized federated learning with joint optimization of local iteration and leader selection for vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002002.

Javan:2025:EED

[2842] Nastooh Taheri Javan, Elahe Zakizadeh Gharyeali, and Seyedakbar Mostafavi. An energy-efficient decentralized federated learning framework for mobile-IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002014.

Du:2025:AMA

[2843] Hongyu Du, Shouhui Zhang, Xi Xv, Yimu Ji, Sisi Shao, Fei Wu, and Shangdong Liu. AF-MCDC: Active feedback-based malicious client dynamic detection. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002026.

Gudepu:2025:GDG

[2844] Venkateswarlu Gudepu, Bhargav Chirumamilla, Venkatarami Reddy Chintapalli, Piero Castoldi, Luca Valcarenghi, Bheemarjuna Reddy Tamma, and Koteswararao Kondepu. GEN-DRIFT: Generative AIdriven drift handling for beyond 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002051.

Dai:2025:ECA

[2845] Hao Dai, Chao Zhai, Jiachao Yu, Jie Gao, and Jie Tian. Edgecaching assisted underlay spectrum sharing in large-scale cognitive radio networks. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002063.

Wan:2025:JDG

[2846] Zheng Wan, Shenglu Zhao, Cuifang Wang, Yifeng Tan, and Xiaogang Dong. Joint DRL and GCNbased Cloud–Edge–End collaborative cache optimization for metaverse scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002075.

Wang:2025:BBMa

[2847] Ziwen Wang, Yajun Guo, and Yimin Guo. A blockchain-based medical IoT authentication scheme resistant to combined attacks. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002099.

Anonymous:2025:EBh

[2848] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002609.

Anonymous:2025:Mc

[2849] Anonymous. May 2025. Computer Networks (Amsterdam, Netherlands: 1999), 263(??):??, May 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Xu:2025:DRL

[2850] Yanli Xu and Qianlian Yu. Deep reinforcement learning based computation offloading and resource allocation strategy for maritime Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001896.

Xu:2025:TLW

[2851] Yang Xu, Junhao Cheng, Hongli Xu, Changyu Guo, Yunming Liao, and Zhiwei Yao. Towards layer-wise quantization for heterogeneous federated clients. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001914.

Wang:2025:ATE

[2852] Linghao Wang, Miao Wang, Chungang Lin, and Yujun Zhang. Accelerating traffic engineering optimization for segment routing: a recommendation perspective. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001926.

Wang:2025:POS

[2853] Kewei Wang, Changzhen Hu, and Chun Shan. Process-oriented security assessment of network services. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001938.

Alphiya:2025:EQL

[2854] A. Alphiya and T. Latha. An efficient QC-LDPC channel encoder/ decoder architecture with parallel vector-matrix computations for 5G wireless networks on FPGA. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625001975.

Monaco:2025:RTL

[2855] Doriana Monaco, Alessio Sacco, Daniele Spina, Francesco Strada, Andrea Bottino, Tania Cerquitelli, and Guido Marchetto. Real-time latency prediction for cloud gaming applications. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002038.

Shi:2025:MCR

[2856] Huaguang Shi, Jian Huang, Hengji Li, Tianyong Ao, Wei Li, and Yi Zhou. Multi-channel real-time access with starvation avoidance for heterogeneous data in smart factories. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500204X.

Cheng:2025:MEP

[2857] Xuetan Cheng, Yingwen Chen, Xiangrui Yang, Huan Zhou, Lailong Luo, and Deke Guo. Memory-efficient programmable packet parsing for multitenant terabit networks. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002087.

Hassan:2025:ETW

[2858] Md Arif Hassan, Mohammad (Behdad) Jamshidi, Bui Duc Manh, Nam H. Chu, Chi-Hieu Nguyen, Nguyen Quang Hieu, Cong T. Nguyen, Dinh Thai Hoang, Diep N. Nguyen, Nguyen Van Huynh, Mohammad Abu Alsheikh, and Eryk Dutkiewicz. Enabling technologies for Web 3.0: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002105.

Han:2025:PPC

[2859] Lulu Han, Weiqi Luo, Anjia Yang, Yudan Cheng, Junzuo Lai, Fang Han, Jiaquan Shen, and Yongxin Zhang. Privacy-preserving cross-domain pointof-interests recommendation based on friendship in LBSs. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002117.

Geppert:2025:JSS

[2860] Heiko Geppert, Frank Dürr, Sukanya Bhowmik, and Kurt Rothermel. Just a second — scheduling thousands of time-triggered streams in largescale networks. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002129.

Tymplalexis:2025:DAS

[2861] Nikolaos Tymplalexis, Kostas Kolomvatsos, and Christos Anagnostopoulos. A demand aware services placement model in Pervasive Edge Computing. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002221.

Han:2025:DMN

[2862] Zihao Han, Ting Zhou, Tianheng Xu, Yuling Ouyang, Xianfu Chen, Tao Chen, and Honglin Hu. Delay minimization in NTNs: Deployment and caching optimization for satellite- and cache-aided UAV. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002257.

Oliveira:2025:OAS

[2863] Rafael Oliveira, Duarte Raposo, Miguel Luís, and Susana Sargento. Optimal action-selection optimization of wireless networks based on virtual representation. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002269.

Xiao:2025:SRD

[2864] Junbi Xiao, Qi Wang, and Yuhao Zhou. ST-RDP: a deep spatiotemporal network model for VNF resource demand prediction of Service Function Chains. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002282.

Li:2025:FDA

[2865] Gui-Hong Li, Heng-Ru Zhang, and Fan Min. Frequency-domain augmentation and multi-scale feature alignment for improving transferability of adversarial examples. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002294.

Zhao:2025:BEP

[2866] Feng Zhao, Benchang Yang, Zhaoyu Su, Chunhai Li, and Yong Ding. A blockchain-enabled privacy-preserving and incentive mechanism-driven federated learning scheme for IoV. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002300.

Teng:2025:SAN

[2867] Tao Teng, Yuanai Xie, and Jiawen Kang. Security analysis of NOMA integrated satellite–terrestrial relay networks with analog beamforming. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002312.

Li:2025:OPS

[2868] Mengge Li, Miao Ma, Liang Wang, and Bo Yang. Online participant selection incorporating coverage quality and participant ability for edgeaided vehicular crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002336.

Xiong:2025:ETC

[2869] Runqun Xiong, Jing Cheng, Jiahang Pu, Xirui Dong, Ciyuan Chen, and Zhuqing Xu. Enhancing trust and collaboration: a reputationdriven mechanism for cross-chain IoT data sharing. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002348.

Alwakeel:2025:EIP

[2870] Ahmed M. Alwakeel. Enhancing IoT performance in wireless and mobile networks through named data networking (NDN) and edge computing integration. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500235X.

Enwereonye:2025:PLS

[2871] Uchenna P. Enwereonye, Ahmad Salehi Shahraki, Hooman Alavizadeh, and A. S. M. Kayes. Physical layer security techniques for grant-free massive Machine-Type Communications in 5G and beyond: a survey, challenges, and future directions. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002361.

Pimentel:2025:DAF

[2872] Erivan Laranjeira Pimentel, Cristiano Antonio de Souza, and Carlos Becker Westphall. Detecting attacks in fog and cloud computing environments using Deep Learning: a systematic literature review. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002373.

Wang:2025:NRA

[2873] Jun Wang, Weibin Jiang, Haodong Xu, Jinsong Hu, Liang Wu, Feng Shu, and Zhou Fang. A novel resource allocation method based on hierarchical deep reinforcement learning for cognitive Internet of Vehicles with unknown channel state information. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002385.

Chan:2025:EFL

[2874] Yun-Hin Chan and Edith C. H. Ngai.

Exploiting features and logits in heterogeneous federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002397.

Zhang:2025:SRA

[2875] Peng Zhang and Qing Guo. Securityreliability analysis in uplink cognitive satellite-terrestrial networks with LEO relaying. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002403.

Yang:2025:ASQ

[2876] Fan Yang, Di Zhang, Junqing Le, and Jiali Wu. Approximate statistic query via sampling for IoT data trading and sharing. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002415.

Khashan:2025:TBF

[2877] Osama A. Khashan. Trust-based fogblockchain model for scalable authentication in smart cities. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002464.

Darchini-Tabrizi:2025:PEU

[2878] Mohsen Darchini-Tabrizi, Amirali Pakdaman-Donyavi, Reza Entezari-Maleki, and Leonel Sousa. Performance enhancement of UAV-enabled MEC systems through intelligent task offloading and resource allocation. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002488.

Anonymous:2025:EBi

[2879] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002907.

Anonymous:2025:Jb

[2880] Anonymous. June 2025. Computer Networks (Amsterdam, Netherlands: 1999), 264(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Wang:2025:CPG

[2881] Yabo Wang, Jiakun Sun, Ruizhi Xiao, Weilong Li, and Shuyuan Jin. Cache Periscope: Gain insights into the global epidemic of malicious domains through DNS Cache. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002130.

Yao:2025:MCB

[2882] Jiangyuan Yao, Weiyang Xin, Xia Yin, Xingang Shi, Zhiliang Wang, Li Zhou, and Ting Jin. A model checking-based framework for testing security properties of protocols under development. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002270.

M:2025:HID

[2883] Jesús F. Cevallos M., Alessandra Rizzardi, Sabrina Sicari, and Alberto Coen-Porisini. HERO: From H ighdimensional network traffic to zEROday attack detection. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002324.

Das:2025:SAB

[2884] Joydip Das, Syed Ashraf Al Tasin, Md. Forhad Rabbi, and Md Sadek Ferdous. A survey of attacks on blockchain systems using a layerbased approach. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002427.

Xu:2025:FFD

[2885] Feiyi Xu, Shihao Hu, Ying Sun, Xiaoxuan Hu, Jin Qi, Yanfei Sun, and Zhenjiang Dong. FDSS: Flight data sharing scheme based on blockchain with dynamic, secure and efficient consensus algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002439.

Baldoni:2025:HBN

[2886] Sara Baldoni and Federica Battisti. Histogram-based network traffic representation for anomaly detection through PCA. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002440.

Saginbekov:2025:SES

[2887] Sain Saginbekov and Dossay Oryspayev. Snowflakes: Efficient schemes for source location privacy in Wireless Sensor Networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002452.

Bao:2025:PEP

[2888] Haiyong Bao, Ronghai Xie, Zhehong Wang, Lu Xing, and Hong-Ning Dai. PDMA: Efficient and privacy-preserving dynamic task assignment with multi-attribute search in crowdsourcing. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002476.

Bellini:2025:CEM

[2889] Pierfrancesco Bellini, Edoardo Branchi, Enrico Collini, Luciano Alessandro Ipsaro Palesi, Paolo Nesi, and Gianni Pantaleo. Certifying entity models, entities and data messages on IoT/ WoT platforms via blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500249X.

Wang:2025:PPH

[2890] Jiaxuan Wang, Xiaofeng Wang, Wenzheng Liu, Qianqian Xing, Xiaoyong Tang, Tan Deng, Ronghui Cao, and Mingfeng Huang. A parallel and pipelined high speed Montgomery modular multiplier for IoT devices. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002506.

Wang:2025:MMF

[2891] Biying Wang, Baosheng Wang, Ziling Wei, Shuang Zhao, Shuhui Chen, Zhengpeng Li, and Minxin Wang. MFSI: Multi-flow based service identification for encrypted network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002518.

Wu:2025:BST

[2892] Zhijun Wu, Hanwen Xu, Meng Yue, and Yanrong Lu. Blockchain security threats: a comprehensive classification and impact assessment. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500252X.

Kwon:2025:EUT

[2893] Yongseok Kwon, Seyoung Ahn, Minho Cho, Yushin Kim, Soohyeong Kim, and Sunghyun Cho. Exploring the unseen: a transformer-based unknown traffic detection scheme with contextual feature representation. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002543.

Hu:2025:TTS

[2894] Zhichao Hu, Dewen Kong, Junzhong Miao, Qing Guo, Gang Du, Likun Liu, Lina Ma, and Xiangzhan Yu. TND: Two-stage non-invasive defense of intrusion detection system from adversarial attack. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002555.

Mancini:2025:SAO

[2895] Francesco Mancini, Riccardo Marzilli, Sara Da Canal, and Giuseppe Bianchi. ScasDK — an all-in-one test platform for security assurance in 5G core networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002646.

Ye:2025:CDT

[2896] Fang Ye, Zitao Zhou, Yifan Wang, Yibing Li, and Xiaoyu Geng. Cross-domain trust aggregation in blockchain-based Internet of Things with dual-layer incentive mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002658.

Chen:2025:TAB

[2897] Yan Chen, Lei Jiao, Tuo Cao, Ji Qi, Gangyi Luo, Sheng Zhang, Sanglu Lu, and Zhuzhong Qian. Toward auction-based edge AI: Orchestrating and incentivizing online transfer learning in edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002683.

Ahmed:2025:JCP

[2898] Fayshal Ahmed, The-Vinh Nguyen, Nam-Phuong Tran, Nhu-Ngoc Dao, and Sungrae Cho. Joint content popularity and audience retentionaware live streaming over RSMA edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002695.

Ren:2025:DCF

[2899] Tian Ren, Siyao Cheng, Hao Zhang, and Jie Liu. Dynamic clustered federated learning via adaptive distribution similarity computation. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002701.

Zhang:2025:AAU

[2900] Ziliang Zhang, Jinmin Wu, Ning Lu, Wenbo Shi, and Zhiquan Liu. Adapt-PUD: an accurate URL-based detection approach against tailored deceptive phishing websites. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002713.

Zhu:2025:ABC

[2901] Xiao Zhu, Wenming Ma, Xiagling Wu, Shengyang Sun, and Xiaoang Zhu. Activity-based capability updating method for task assignment in mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002725.

Hussienat:2025:ADA

[2902] Ayyoub Hussienat, Haitham Al-Obiedollah, Haythem Bany Salameh, and Yaser Jararweh. AI-driven approach for enhanced signal detection in future NOMA-enabled 6G IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002737.

Ren:2025:EEP

[2903] Haiyang Ren, Shiyou Qian, Zhonglong Zheng, Jiange Zhang, Zhengyu Liao, Hanwen Hu, Jian Cao, Guangtao Xue, and Minglu Li. EPC: an ensemble packet classification framework for efficient and stable performance. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002749.

Li:2025:PBI

[2904] Anqi Li, Quan Zhang, Xin Xu, and Yingbiao Yao. Practical Byzantine improved algorithm based on nodeindependent validation. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002750.

Delgado-Soto:2025:GWT

[2905] Javier Aday Delgado-Soto, Jorge E. López de Vergara, Iván González, Daniel Perdices, and Luis de Pedro. GPT on the wire: Towards realistic network traffic conversations generated with large language models. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002762.

Altaweel:2025:DRE

[2906] Ala Altaweel and Ahmed Bouridane. DR3DH: a DoS resistant extended triple Diffie–Hellman for mobile edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002774.

Chang:2025:RBC

[2907] Yawen Chang, Lei Lei, and Xudong Wang. Region-based coordination for distributed resource allocation in sidelink mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002828.

Sopin:2025:IFU

[2908] Eduard Sopin, Anastasia Daraseliya, Vyacheslav Begishev, Konstantin Samouylov, and Yevgeni Koucheryavy. Improving fairness and utilization in 5G/6G mmWave/sub-THz systems via virtual queuing. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002841.

Wang:2025:VVE

[2909] Yingqing Wang, Yanhua Liang, Yue Huang, and Guihe Qin. VECLLF: a vehicle-edge collaborative lifelong learning framework for anomaly detection in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002956.

Cheng:2025:MMF

[2910] Minglong Cheng, Wei Chen, Weidong Fang, Zehua Wang, Tingting Xu, Jueting Liu, and Victor C. M. Leung. MFTE: Multifactor and fuzzy trust evaluation for federated learning in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500307X.

Anonymous:2025:EBj

[2911] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003275.

Anonymous:2025:Jc

[2912] Anonymous. June 2025. Computer Networks (Amsterdam, Netherlands: 1999), 265(??):??, June 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Yin:2025:TDU

[2913] Yuan Yin, Yiyang Zhang, Jiaheng Wang, Ningxing Shen, Jianli Guo, and Kang Zheng. Transmission design for uplink multi-group MIMO– NOMA with finite blocklength. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002816.

Marzin:2025:CCE

[2914] Alon Marzin, Moshe Schwartz, and Michael Segal. Covert channel by exploiting error-correcting codes. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002786.

Shi:2025:CCU

[2915] Yiyang Shi, Xiangyang Luo, and Wenqi Shi. From crowds to coordinates: a user density-distance dynamic transformation method for Telegram users geolocalization. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500283X.

Liu:2025:BBR

[2916] Peng Liu, Zixu Zhang, Chenhao Ren, and Hailong You. A blockchainbased resource sharing incentivization mechanism for multi-to-multi in compute first networking. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002853.

Singha:2025:TMT

[2917] N. Rana Singha, Nityananda Sarma, and Dilip Kumar Saikia. TMLpSA-MEC: Transformer-based mobility aware periodic service assignment in mobile edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002968.

Artioli:2025:SGC

[2918] Pierpaolo Artioli, Vincenzo Dentamaro, Stefano Galantucci, Alessio Magrì, Gianluca Pellegrini, and Gianfranco Semeraro. SIEVE: Generating a cybersecurity log dataset collection for SIEM event classification. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500297X.

Chai:2025:RJG

[2919] Yuhao Chai, Yong Zhang, Zhenyu Zhang, Da Guo, and Yinglei Teng. Research on joint game theory and multi-agent reinforcement learningbased resource allocation in micro operator networks. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003007.

Chen:2025:MRE

[2920] Hongyang Chen, Pengfei Chen, Zilong He, Zibin Zheng, and Kaibin Fang. MoTor: Resource-efficient cloud-native network acceleration with programmable switches. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003020.

Mihertie:2025:RAU

[2921] Habtamu Demeke Mihertie and Zhengqiang Wang. Resource allocation for UAV-RIS-assisted RSMA system with hardware impairments. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003032.

Shekhawat:2025:MPE

[2922] Virendra Singh Shekhawat, Rakhee Kulshrestha, Pooja Yadav, Ajay Singh, and Farhaan Firdous. Modeling and performance evaluation of Open-Flow switches using a MAP/PH/1/n queueing model. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003056.

Gomez:2025:IFF

[2923] Jose Gomez, Elie F. Kfoury, Ali Mazloum, and Jorge Crichigno. Improving flow fairness in non-programmable networks using P4-programmable Data Planes. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003068.

Rivera-Dourado:2025:EFN

[2924] Martiño Rivera-Dourado, Christos Xenakis, Alejandro Pazos, and Jose Vázquez-Naya. EAP-FIDO: a novel EAP method for using FIDO2 credentials for network authentication. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003159.

Lee:2025:CLC

[2925] Chankyun Lee and Wonhyuk Lee. Cross-layering consideration for network and quantum resources-aware quantum-secured networking. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003172.

Anonymous:2025:EBk

[2926] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003573.

Anonymous:2025:Jd

[2927] Anonymous. July 2025. Computer Networks (Amsterdam, Netherlands: 1999), 266(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Jamshed:2025:EEA

[2928] Muhammad Ali Jamshed, Ali Nauman, Ayman A. Althuwayb, Haris Pervaiz, and Sung Won Kim. Electromagnetic emission-aware machine learning enabled scheduling framework for unmanned aerial vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002798.

Olthuis:2025:CDA

[2929] Jorrit J. Olthuis, Savio Sciancalepore, and Nicola Zannone. Cyberattacks and defenses for autonomous navigation systems: a systematic literature review. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002981.

Khan:2025:COM

[2930] Muhammad Saleem Khan, Sobia Jangsher, Junaid Qadir, and Hassaan Khaliq Qureshi. Cooperative offloading multi-access edge computing (COMEC) for cell-edge users in heterogeneous dense networks. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002993.

Navaridas:2025:IPD

[2931] Javier Navaridas and Jose A. Pascual. Improving the performance of Dragonfly networks through restrictive proxy routing strategies. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003019.

Zeng:2025:CSP

[2932] Chao Zeng, Xingwei Wang, Rongfei Zeng, Shining Zhang, Jianzhi Shi, and Min Huang. Containerized service placement and resource allocation at edge: a Hybrid Reinforcement Learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500310X.

Song:2025:UPM

[2933] Peixuan Song, JunKyu Lee, Ahmed M. Abdelmoniem, and Lev Mukhanov. Unlocking the power of 4G/5G mobile networks: an empirical dive into quality and energy efficiency in YouTube Edge services. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003111.

Bouchmal:2025:QAO

[2934] Oumayma Bouchmal, Bruno Cimoli, Ripalta Stabile, Juan Jose Vegas Olmos, and Idelfonso Tafur Monroy. Quantum Approximate Optimization Algorithm applied to multiobjective routing for large scale 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003123.

He:2025:DSS

[2935] Chunrong He, Songtao Guo, Guiyan Liu, and Wei Zhang. DP-SAFL:

Semi-asynchronous federated learning with differential privacy in heterogeneous edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003135.

Wang:2025:ERC

[2936] Hanwu Wang, Moshe Zukerman, Yingchao Zhao, and Ying Fu. An explicit rate control based traffic transmission and schedule scheme in UAV-IOT network slicing system. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003147.

Salimi:2025:HPA

[2937] Rezvan Salimi, Sadoon Azizi, and Javad Dogani. A hybrid priority-aware genetic algorithm and opposition-based learning for scheduling IoT tasks in green fog computing. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003160.

Liao:2025:PCT

[2938] Yang Liao, Huayang Zhou, Chengfeng Leng, Zhenlang Su, and Tuanfa Qin. Power control and task offloading strategies for high-density wireless body area networks based on deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003184.

Murcia:2025:BBA

[2939] José Manuel Bernabé Murcia, Alejandro M. Zarca, and Antonio Skármeta. BASTION: Beyond automated service and security orchestration for next-generation networks. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003196.

Amadeo:2025:ICP

[2940] Marica Amadeo, Claudia Campolo, Giuseppe Ruggeri, and Antonella Molinaro. Improving communication performance of Federated Learning: a networking perspective. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003202.

Buruaga:2025:HQS

[2941] Jaime S. Buruaga, Ruben B. Méndez, Juan P. Brito, and Vicente Martin. Hybrid quantum-safe integration of TLS in SDN networks. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003226.

Tian:2025:ATE

[2942] Ying Tian, Zhiliang Wang, Xia Yin, Xingang Shi, Jiahai Yang, and Han Zhang. Adaptive traffic engineering with segment routing through deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003238.

Ramirez-Martinez:2025:PAS

[2943] Erick D. Ramirez-Martinez, Jesús A.

Pérez-Díaz, and Noe M. Yungaicela-Naula. P4-Assisted Slowloris DDoS attack detection in IoT environments by using ML and DL. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003317.

Zheng:2025:JRB

[2944] Weichang Zheng, Mingcong Yang, Yu Zheng, Yongbing Zhang, and Kun Yang. Joint routing, band, and spectrum assignment in C+L band elastic optical networks towards ultra-broadband era. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003408.

Nadeem:2025:ASN

[2945] Muhammad Nadeem and Chen Hongsong. Advancing social network security with Magteon-Turing L3TM: a multi-layered defense system against cyber threats. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003421.

Kumar:2025:CSL

[2946] Ankit Kumar, Mikail Mohammed Salim, David Camacho, and Jong Hyuk Park. A comprehensive survey on large language models for multimedia data security: challenges and solutions. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003469.

Anonymous:2025:EBl

[2947] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500386X.

Anonymous:2025:J

[2948] Anonymous. July 2025. Computer Networks (Amsterdam, Netherlands: 1999), 267(??):??, July 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Han:2025:VNC

[2949] Xinchi Han, Shizhen Zhao, Yongxi Lv, Peirui Cao, Weihao Jiang, Qinwei Yang, Yunzhuo Liu, Shengkai Lin, Bo Jiang, Ximeng Liu, Yong Cui, Chenghu Zhou, and Xinbing Wang. vClos: Network contention aware scheduling for distributed machine learning tasks in multi-tenant GPU clusters. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002531.

Liu:2025:TOR

[2950] Ziqi Liu, Gaochao Xu, Bo Liu, Xu Xu, and Long Li. Task offloading and resource allocation in hybrid-powered WPT MEC system: an enhanced deep reinforcement learning method. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002804.

Liu:2025:JOS

[2951] Xin Liu, Haocheng Zhu, and Eylem Ekici. O-JRC: an open source software platform for mmWave joint radarcommunication development and experimentation. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003044.

Li:2025:RIF

[2952] Xuewei Li, Zengyang Zheng, Mankun Zhao, Yue Zhao, Lifeng Shi, and Baoliang Wang. RLFE-IDS: a framework of intrusion detection system based on retrieval augmented generation and large language model. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003081.

Patane:2025:VCC

[2953] Rosario Patanè, Nadjib Achir, Andrea Araldo, and Lila Boukhatem. Vehicular cloud computing: a costeffective alternative to edge computing in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003329.

Zou:2025:TFS

[2954] Zeguang Zou, Rentao Gu, Hongxiang Wang, Yejun Liu, and Lin Bai. Time–frequency–spatial fragmentation aware service provisioning in multiband elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003330.

Moges:2025:AIA

[2955] Teshager Hailemariam Moges, Thanh Phung Truong, Demeke Shumeye Lakew, Thien Ho Huong, Vinh Truong Hoang, Nhu-Ngoc Dao, and Sungrae Cho. Age of information-aware trajectory optimization for time-sensitive UAV systems in uplink SCMA networks. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003366.

Cho:2025:FIV

[2956] Seongbin Cho, Hojun Seong, Haein Kang, Seong je Cho, and Boo-Joong Kang. Forensic investigation of vehicle-related data in Android phones connected to in-vehicle infotainment systems. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003378.

S:2025:LDR

[2957] Lizeth P. Aguirre S., Yao Shen, and Minyi Guo. LCA: Deep reinforcement learning-based congestion avoidance routing model in SDN. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500338X.

Kaya:2025:NHA

[2958] Muhammed Onur Kaya, Mehmet Ozdem, and Resul Das. A new hybrid approach combining GCN and LSTM for real-time anomaly detection from dynamic computer network data. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003391.

Ju:2025:CDE

[2959] Bobo Ju, Yang Liu, Jing Liu, Peng Sun, and Liang Song. CNN-DAG-Editor: a convolutional neural network offloading analyzer with multiobjective dynamic adaptive resource competitive swarm optimization. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500341X.

Rasheed:2025:FLB

[2960] Iftikhar Rasheed and Hala Mostafa. Federated learning-based anomaly detection for zero-day attack prevention in 6G network slices. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003433.

Guo:2025:DOE

[2961] Yingya Guo and Cen Chen. DRLO: Optimizing edge server placement in dynamic MEC scenarios using deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003445.

Wu:2025:BED

[2962] Zhonghui Wu, Changqiao Xu, Mu Wang, Yunxiao Ma, Zicong Huang, Jingtian Liu, Han Xiao, Lujie Zhong, and Luigi Alfredo Grieco. Blockchainenabled dispersed computing paradigm in Web 3.0 metaverse. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003457.

Ren:2025:RUD

[2963] Linghao Ren, Sijia Wang, Shengwei Zhong, Yiyuan Li, and Bo Tang. ROEN: Universal dynamic topologyadaptive evolution model for multimodal mixed network traffic detection. Computer Networks (Amsterdam, Netherlands: 1999), 268 (??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003470.

Wang:2025:TDC

[2964] Zhijie Wang, Wei Zhang, Dingcheng Yang, Fahui Wu, Yu Xu, and Lin Xiao. Trajectory design of cellularconnected UAV patrol and mobile edge computing system: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003512.

Anonymous:2025:EBm

[2965] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004219.

Anonymous:2025:Ac

[2966] Anonymous. August 2025. Computer Networks (Amsterdam, Netherlands: 1999), 268(??):??, August 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Fadoul:2025:SGA

[2967] Moubachir Madani Fadoul, Mohammed S. M. Gismalla, and Hashim Elshafie. A stochastic geometry approach for relay-assisted uplink multicell networks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625002865.

dAmbrosio:2025:SSM

[2968] Nicola d’Ambrosio, Claudio Lista, Gaetano Perrone, and Simon Pietro Romano. SMASH: an SDN-MTD framework for efficient honeypot deployment and insider threat mitigation. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625002944.

Lacava:2025:DER

[2969] Andrea Lacava, Leonardo Bonati, Niloofar Mohamadi, Rajeev Gangula, Florian Kaltenberger, Pedram Johari, Salvatore D’Oro, Francesca Cuomo, Michele Polese, and Tommaso Melodia. dApps: Enabling real-time AIbased open RAN control. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003093.

Escheikh:2025:SLB

[2970] Mohamed Escheikh, Wiem Taktak, and Kamel Barkaoui. Scalable load balancing scheme for wireless distributed controllers in SDDC. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003214.

Guo:2025:WDR

[2971] Zhongfu Guo, Xinsheng Ji, Wei You, Hai Guo, Yang Zhang, Yu Zhao, Mingyan Xu, and Yi Bai. Workload distribution with rateless encoding: a low-latency computation offloading method within edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003482.

Li:2025:APC

[2972] Jingzong Li, Yik Hong Cai, Libin Liu, Yu Mao, Chun Jason Xue, and Hong Xu. Accelerating point cloud analytics on resource-constrained edge devices. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003494.

Chen:2025:BRR

[2973] Yan Chen, Xiaoyan Hu, Xinghai Chen, Guang Cheng, Ruidong Li, and Hua Wu. BTG-RF: Recognizing douyin payment behaviors based on behavioral traffic graph analysis. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003524.

Barbhuiya:2025:FFA

[2974] Dilwar Hussain Barbhuiya, Sanjib K. Deka, Nityananda Sarma, and Bijoy Chand Chatterjee. FARASP: Fragmentation-aware resource allocation with spectrum partitioning in elastic optical networks. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003536.

Fiorini:2025:QDA

[2975] Francesco Fiorini, Marco Cococcioni, and Michele Pagano. Quantitative delay analysis of GI/G/1 queues with heavy-tailed traffic by means of Alpha Theory. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003615.

Wehner:2025:EAT

[2976] Nikolas Wehner, Pedro Casas, Katharina Dietz, Stefan Geißler, Tobias Hoßfeld, and Michael Seufert. Exploring the application of Time Series Foundation Models to network monitoring tasks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003627.

Leduc:2025:DSA

[2977] Guilain Leduc, Sylvain Kubler, and Jean-Philippe Georges. DRAFTEE: a self-adaptive framework for BFTbased consensus to meet fluctuating throughput demands under security constraint. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003639.

Cheng:2025:EDD

[2978] Tao Cheng, Shining Li, Weiwei Dang, and Yan Pan. An effective data dissemination method in WAIC networks using enhanced CTC. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003664.

Liu:2025:OCT

[2979] Haolin Liu, Shi Yin, Tingrui Pei, Zhiquan Liu, Qingyong Deng, and Yanping Cheng. Optimizing cost through UAV deployment and task assignment in hybrid UAV-assisted MEC systems. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003676.

Dai:2025:DMD

[2980] Hui Dai, Lingyun Yuan, Jiaying Wu, Han Chen, and Haochen Bao. DBSRP-ML: Dynamic blockchain sharding reconfiguration protocol based on multilabel. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003688.

Yu:2025:DRL

[2981] Chengbo Yu, Yidan Sun, Xingxing Li, Fan Yang, Jie Huang, and Yingying Peng. Decentralized reinforcement learning-based resource management for dense-interference IoV. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500369X.

Zhang:2025:GRO

[2982] Hongxia Zhang, Ning Gao, Xiangxu Zhao, Yuanbo Zhao, and Peiying Zhang. GSDRL: Resource orchestration in integrated satellite– terrestrial network for a hybrid edge– cloud. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003706.

Zhang:2025:CIN

[2983] Penghui Zhang, Hua Zhang, Yuqi Dai, Yibo Pi, Jingyu Wang, and Jianxin Liao. Cache-INT: In-network cachingenabled In-band Network Telemetry. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003718.

Liu:2025:MUR

[2984] Xueting Liu, Xiaohan Wang, Chao Li, Joojo Walker, Wenxin Tai, Ting Zhong, Yong Wang, Fan Zhou, and Kai Chen. Mapping the unseen: Robust IP geolocation through the lens of uncertainty quantification. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500372X.

Liao:2025:PFL

[2985] Gengjian Liao, Yulin Yang, and Zhenni Feng. Personalized federated learning through self-knowledge distillation in vehicular edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003731.

Zhang:2025:LHP

[2986] Guannan Zhang, Dezun Dong, Tao Li, and Siyuan Yang. Libra: Highprecision congestion control for datacenter networks with in-network allreduce. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003743.

Wigchert:2025:DAS

[2987] Jos Wigchert, Savio Sciancalepore, and Gabriele Oligeri. Detection of aerial spoofing attacks to LEO satellite systems via deep learning. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003755.

Yu:2025:EEE

[2988] Jianglai Yu, Lei Zhang, Dongzhou Cheng, Can Bu, Liangdong Liu, Hao Wu, and Aiguo Song. Ensemble early exit network on human activity recognition using wearable sensors. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003767.

Kaltenberger:2025:DIW

[2989] Florian Kaltenberger, Tommaso Melodia, Irfan Ghauri, Michele Polese, Raymond Knopp, Tien Thinh Nguyen, Sakthivel Velumani, Davide Villa, Leonardo Bonati, Robert Schmidt, Sagar Arora, Mikel Irazabal, and Navid Nikaein. Driving innovation in 6G wireless technologies: the OpenAirInterface approach. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003779.

Lu:2025:LSF

[2990] Xiaojuan Lu, Liying Chen, Kai Wang, Pingping Dong, and Lianming Zhang. Long and short flow buffer management in data center networks. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003780.

Sarohe:2025:SCS

[2991] Sanehi Sarohe, Sandeep Harit, and Manish Kumar. A systematic and comprehensive survey of load balancing techniques in Software Defined Network based Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003792.

Wihidayat:2025:DAM

[2992] Endar Suprih Wihidayat, Sieteng Soh, Kwan-Wu Chin, and Duc-Son Pham. Delay-aware multi-stage edge server placement and task offloading with budget constraint. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003809.

Zhou:2025:OAS

[2993] Fangyu Zhou, Feng Shan, Weiwei Wu, Runqun Xiong, and Junzhou Luo. Optimal adaptive scheduling to maximize throughput for battery constrained time-varying RF-powered systems. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625003810.

Lee:2025:EDB

[2994] Jungmin Lee, Sungju Yun, Nahyun Kim, and Yeonjoon Lee. Enhanced DGA botnet domain detection and family classification via n-gram analysis and Hellinger distance. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003822.

Wang:2025:DPA

[2995] Yang Wang, Xiaofei Xing, Guojun Wang, Yuheng Zhang, and Peiqiang Li. Dikaios: Position-anchored group ordering with reputation for fair and efficient Byzantine consensus. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003901.

Luo:2025:SAT

[2996] Feng Luo, Lei Zhu, Zitong Wang, Haotian Gan, Yunpeng Li, Zhenyu Yang, and Dengcheng Liu. Schedulability analysis of time aware shaper with preemption supported in timesensitive networks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003913.

Soltanshahi:2025:MIU

[2997] Minoo Soltanshahi and Martin Maier.

Metaversal intelligence: Unifying human-AI interactions in human-in-the-loop AIB-Metaverse. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003925.

Liu:2025:AED

[2998] Yilong Liu, Hong Zhang, Miao Wang, Qiqi Xie, and Zhibo Sun. AntidoteFL: Enhancing defense against poisoning attacks in federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003949.

Liu:2025:EAS

[2999] Guijiang Liu, Hao Lu, Wenming Wang, and Haiping Huang. An efficient authentication scheme for vehicular networks based on Merkle tree. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003962.

Bai:2025:ADT

[3000] Yulong Bai, Ying Wang, Xiangrui Xu, Yuhang Yang, Hina Batool, Zahid Iqbal, and Jiuyun Xu. AsyncDefender: Dynamic trust adaptation and collaborative defense for Byzantinerobust asynchronous federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003974.

Li:2025:OBS

[3001] Jianpo Li, Jinjian Pang, Binfeng Jiang, Qi Xu, and Enyuan Zhang. Optimization of 5G base station deployment based on quantum genetic algorithm in outdoor 3D map. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003986.

Bie:2025:SST

[3002] Yuxia Bie, Xin Li, Ye Tian, and Yupeng Wang. Self-similar traffic prediction algorithm for satellite network based on dual decomposition and neural network. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003998.

Huang:2025:DDC

[3003] Xu Huang, Jia Chen, Deyun Gao, Shang Liu, Shangbing Qiao, and Hongke Zhang. DCDTS: Deterministic cross-domain transmission and scheduling for large-scale deterministic networks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004013.

Medjadba:2025:IDS

[3004] Yasmine Medjadba, Hamza Drid, and Mohamed Rahouti. Intrusion detection in software-defined networking using hybrid Bayesian model averaging for reliable uncertainty quantification. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004037.

Said:2025:ECB

[3005] Naina Said and Olaf Landsiedel. Edge-Boost: Confidence boosting for resource constrained inference via selective offloading. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004049.

Zhang:2025:CSS

[3006] Xiaohong Zhang, Heng Huang, and Shaojiang Zhong. Certificateless signcryption scheme incorporating dual blockchains and SM9 cryptographic algorithm in VANETs communication. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004050.

Tu:2025:MCA

[3007] YuKuan Tu, Tengyao Li, Meng Zhang, and Xiangyang Luo. MVC-Corr: an accurate and efficient flow correlation method based on multi-view fusion and contrast augmentation. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004062.

Zhao:2025:BBA

[3008] Jintao Zhao, Jialiang Yan, Siyao Cheng, and Jie Liu. BLAW: BLE assisted Wi-Fi in idle listening. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004074.

Li:2025:FRF

[3009] Yang Li, Jun Xu, and Dejun Yang.

FLCom: Robust federated learning against strong model poisoning attacks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004098.

Zhang:2025:RRT

[3010] Hong Zhang, Zhenge Xu, Feixue Han, Fuliang Li, and Qing Li. RTQS: Realtime flow fairness queue scheduling policy on router devices. Computer Networks (Amsterdam, Netherlands:

1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004104.

Chen:2025:DTO

[3011] Liqiong Chen, Xinyuan Yang, Huaiying Sun, Xiuchao Yu, and Kaiwen Zhi. Dependent task offloading in multi-access edge computing: a GCN augmented deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004128.

Casari:2025:SIE

[3012] Paolo Casari and Stefano Basagni. Special issue: Extended papers from the 16th International Conference on Underwater Networks & Systems “WUWNet 2022,” and Advances of the State-of-the-Art in Underwater Communications and Networking. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004141.

Carillo:2025:EFC

[3013] Raffaele Carillo, Francesco Cerasuolo, Giampaolo Bovenzi, Domenico Ciuonzo, and Antonio Pescapè. Explainable federated class incremental learning for Encrypted Network Traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004153.

Fu:2025:IZI

[3014] Yulong Fu, Yuanqi Xie, Wei Yi, Bikal Poudel, Jin Cao, and Hui Li. IPO-ZTA: an intelligent policy orchestration zero trust architecture for B5G and 6G. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004177.

Liu:2025:SLA

[3015] Hongze Liu, Junzhe Liu, Zhaojiacheng Zhou, Shijing Yuan, Jiong Lou, Chentao Wu, and Jie Li. A stochastic learning algorithm for multi-agent game in mobile network: a cross-silo federated learning perspective. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004256.

Lv:2025:AFL

[3016] Fei Lv, Hangyu Wang, Zhiwen Pan, Rongkang Sun, Shuaizong Si, Weidong Zhang, Shichao Lv, and Limin Sun. Asynchronous federated learning based zero trust architecture for the next generation industrial control systems. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004268.

Guo:2025:MOO

[3017] Wei Guo, Xiaoli Wang, and Yuping Wang. Multi-objective optimization model and algorithm for network slicing with demand exceeding resources. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500427X.

Li:2025:CCS

[3018] Chen Li, Tingwei Ye, Tongyu Zong, Liyang Sun, Houwei Cao, and Yong Liu. Coffee: Cost-effective edge caching for live 360 degree video streaming. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004281.

Shahid:2025:TAH

[3019] Syed Maaz Shahid, SungKyung Kim, and Sungoh Kwon. Topology-aware handover parameterization for radio link failures reduction in dynamic ON/ OFF small-cell networks. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004293.

Xie:2025:RAU

[3020] Wancheng Xie, Helin Yang, and Zehui Xiong. Resource allocation for UAV-assisted anti-jamming semantic D2D networks: a graph reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500430X.

Liu:2025:JAP

[3021] Huanlin Liu, Yangyang Hou, Yong Chen, Huiling Zhou, Haonan Chen, Yuxin Xie, and Wenjie Song. Joint access point selection and resource allocation for optimizing energy efficiency in aggregated VLC-RF network. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004347.

Yang:2025:TTT

[3022] Wen Yang, Chenxi Tang, Chaowei Tang, Jingwen Lu, Jing Si, Zhuo Zeng, and Wenyu Ma. Traffic through two lenses: a dual-branch vision transformer for IoT traffic classification. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004360.

Wu:2025:ALR

[3023] Jiahao Wu, Sanfeng Zhang, Hongxian Liu, and Wang Yang. AAR-Log: a robust log anomaly detection method resisting adversarial attacks. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004384.

Gharavi:2025:PPQ

[3024] Hadi Gharavi, Jorge Granjal, and Edmundo Monteiro. PQBFL: a Post-Quantum Blockchain-based protocol for Federated Learning. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004396.

Rezaei:2025:FFL

[3025] Hadiseh Rezaei, Rahim Taheri, and Mohammad Shojafar. FedLLMGuard: a federated large language model for anomaly detection in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004402.

Long:2025:PPH

[3026] Xingguo Long, Kun Huang, Rongwei Yang, Qingguo Dai, and Zhenyu Li.

Pegasus: a practical high-speed crossplatform NDN forwarder. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004414.

Sarraf-Maralaniyan:2025:RML

[3027] Sarah Sarraf-Maralaniyan and Akbar Ghaffarpour Rahbar. Routing, modulation-level, core and spectrum assignment schemes for dynamic multicast traffic in SDM-EONs. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004426.

Sciddurlo:2025:VDS

[3028] Giancarlo Sciddurlo, Federica de Trizio, Giuseppe Piro, and Gennaro Boggia. A value-driven system design framework for sustainable 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500444X.

Mehmeti:2025:CPR

[3029] Fidan Mehmeti, Serkut Ayvaşık, Furkan Kaynar, Thomas F. La Porta, and Wolfgang Kellerer. Constant playout rates: Resource allocation for improved user experience with live video streaming in 5G. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004487.

Matrouk:2025:ADL

[3030] Khaled M. Matrouk, Arunmozhi Selvi, Ahmad Yahiya Ahmad Bani Ahmad, and Dhurgadevi M. An adaptive deep learning-based user authentication with digitally signed optimal key-aided encryption for secured mobile edge computing data storage in blockchain. Computer Networks (Amsterdam, Netherlands: 1999), 269 (??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004499.

Shen:2025:DTO

[3031] Bo Shen, Qi Gu, and Gang Yang. Dynamic task offloading and online scheduling for edge-enabled IoT with a hierarchical framework. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004530.

Anonymous:2025:EBn

[3032] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S1389128625005031.

Anonymous:2025:S

[3033] Anonymous. September 2025. Computer Networks (Amsterdam, Netherlands: 1999), 269(??):??, September 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Neelima:2025:FAG

[3034] Mudadla Neelima, Munesh Singh, Kshira Sagar Sahoo, and Joel J. P. C. Rodrigues. Fingerprinting-assisted geometric approach for device-free localization in wireless network. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003500.

Hu:2025:BEB

[3035] Xingyu Hu, Dinghuang Hu, Dezun Dong, Cunlu Li, Zejia Zhou, and Guoyuan Yuan. BCN: Enhanced backpressure flow control with rapid notification in datacenter networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003937.

Zhan:2025:APF

[3036] Ziwei Zhan, Weijie Liu, Xiaoxi Zhang, Chee Wei Tan, Lei Xue, Haisheng Tan, and Xu Chen. Accelerating personalized federated learning via dynamic gradient substitution and client selection. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625003950.

Farias:2025:TMD

[3037] Euclides Peres Farias, Anderson Bergamini de Neira, Ligia Fracielle Borges, and Michele Nogueira. Transformers model for DDoS attack detection: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004001.

Ma:2025:GEG

[3038] Ronghao Ma, Jianhong Zhou, and Maode Ma. 5G-GRAKA: an efficient group based authentication and key agreement protocol for machine-type communication in 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004025.

Van:2025:ECM

[3039] Nguyen Thi Thanh Van, Hoang Le Hung, Nguyen Cong Luong, Huy T. Nguyen, Nguyen Tien Hoa, Ngo Manh Duy, and Ngo Manh Tien. Energy consumption minimization for robotic systems in intelligent factories with the assistance of STAR-RIS: a reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004086.

Braeken:2025:ULP

[3040] An Braeken. Ultra lightweight postquantum resistant 5G-AKA protocol. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004116.

Weedage:2025:BAD

[3041] L. Weedage, C. Stegehuis, and S. Bayhan. beam-align: Distributed user association for mmWave networks with multi-connectivity. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500413X.

Jiang:2025:TRT

[3042] Huaijie Jiang, Guang Cheng, and Li Deng. TAO: a real-time network traffic analysis task orchestration framework with optimized filtering and scheduling. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004165.

Xue:2025:ECI

[3043] Tao Xue, Ying Zhang, Yanbin Wang, Wenbo Wang, Shuailou Li, and Haibin Zhang. Edge computing for IoT: Novel insights from a comparative analysis of access control models. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004359.

Javanmardi:2025:IPS

[3044] Saeed Javanmardi, Alfredo Nascita, Antonio Pescapè, Giovanni Merlino, and Marco Scarpa. An integration perspective of security, privacy, and resource efficiency in IoT-Fog networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004372.

Sun:2025:JMF

[3045] Kai Sun, Bin Hu, Zhimeng Yin, Shuai Wang, Shuai Wang, Zhuqing Xu, and Tian He. Joint multidimensional features for LoRa reception in burst traffic. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004438.

Saha:2025:DDH

[3046] Himel Saha, Md Nur Ahmed, Palash Roy, Md. Abdur Razzaque, Nafis Fuad Tanvir, Mohammad Mehedi Hassan, and Md Zia Uddin. Device and data heterogeneity aware SplitFed learning for digital twin empowered industrial Internet of Things. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004451.

Kalakoti:2025:FLE

[3047] Rajesh Kalakoti, Sven Nõmm, and Hayretdin Bahsi. Federated learning of explainable AI(FedXAI) for deep learning-based intrusion detection in IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004463.

Yang:2025:LAS

[3048] Chen Yang, Lei Tian, Xiang Wang, Feilong Lin, Riheng Jia, Zhonglong Zheng, and Minglu Li. LAS: Lightweight aggregate signcryption for federated learning with blockchain in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004475.

Xie:2025:CAT

[3049] Mande Xie, Longchen Li, and Xueping Ni. Cache-assisted task offloading strategy based on multi-agent deep reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004505.

Haider:2025:SAD

[3050] Zeeshan Ali Haider, Asim Zeb, Taj Rahman, Sushil Kumar Singh, Rizwan Akram, Ali Arishi, and Inam Ullah. A survey on anomaly detection in IoT: Techniques, challenges, and opportunities with the integration of 6G. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004517.

Tsai:2025:LRC

[3051] Tung-Tso Tsai and Jung-Hao Yang. Leakage-resilient certificateless public key encryption with equality test resistant to side-channel attacks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004529.

Akbari:2025:MDA

[3052] Milad Akbari, Raffaele Bolla, Roberto Bruschi, Chiara Lombardo, Nicole Simone Martinelli, and Beatrice Siccardi. A management data analytics function for ethical 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004542.

Siriwardhana:2025:OHM

[3053] A. C. P. K. Siriwardhana, Jingling Yuan, and Zhishu Shen. Optimizing handover mechanism in vehicular networks using deep learning and optimization techniques. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004554.

Liu:2025:NTE

[3054] Jiawei Liu, Bing Dong, Weilong Wang, Muhan Yuan, Borui Li, Zhao-Dong Xu, and Shuai Wang. nnWeb: Towards efficient WebGPU-based DNN inference via automatic collaborative offloading. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004566.

Yu:2025:DDP

[3055] Shasha Yu, Yanan Qiao, Fan Yang, and Junge Bo. DPoW: a decentralized proof-of-work consensus mechanism for blockchain system. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004578.

Ma:2025:JRA

[3056] Jingya Ma and Hongyuan Gao.

Joint resource allocation for highmobility HCNs with D2D communications. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500458X.

Liu:2025:EEB

[3057] Yidan Liu, Liujia Cai, Haoyuan Xue, Bing Gao, Xiangyu Wang, Siqi Lu, and Yongjuan Wang. EUAV: an enhanced blockchain-based two-factor anonymous authentication key agreement protocol for UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004591.

Xie:2025:ESB

[3058] Qingqing Xie and Yifeng Zhao. Efficient and secure batch-aggregation scheme against poisoning attacks for federated learning in VANETs. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004608.

Djennadi:2025:ESA

[3059] Liticia Djennadi, Gladys Diaz, Khaled Boussetta, and Christophe Cerin. Exploring SDN architectures for IoT over low-power and lossy networks (LLNs): a survey. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500461X.

Yu:2025:EJE

[3060] Rongwei Yu, Zijian Zhou, Yuhao Zhang, Lina Wang, and Qiyun Shao. ESTK-JC: Encrypted malicious traffic detection fusing spatio-temporal features and JointCloud entity knowledge for JointCloud environment. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004621.

Xu:2025:LCA

[3061] Yang Xu, Qianshu Wang, Hongli Xu, Yunming Liao, Liusheng Huang, and Xin Hang. LOGO-CL: Accelerating semi-supervised federated learning in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004633.

Foka:2025:HLG

[3062] Gwladys Ornella Djuikom Foka, Razvan Stanica, and Diala Naboulsi. HGC-LSTM: a graph neural network-based model for HO forecasting in mobile networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004645.

Zhang:2025:IEP

[3063] Yuanhang Zhang, Xu An Wang, Xiaoxuan Xu, Weiwei Jiang, Wei Wei, Xiaoyuan Yang, and Hao Liu. Improved efficient public/private cloud auditing scheme with dynamic updates. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004657.

Li:2025:EDM

[3064] Zhiyuan Li and Yujie Jin. EncryptoVision: a dual-modal fusionbased multi-classification model for encrypted traffic recognition. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004669.

Li:2025:PBA

[3065] Juxia Li, Lejun Zhang, Chenglin Chen, Zuwei Li, Wenjie Long, Yuqian Wang, and Ran Guo. PBAS: a blockchainassisted priority-based batch authentication scheme for secure and scalable vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004670.

Caturano:2025:CCB

[3066] Francesco Caturano, Jordan Ciotola, Simon Pietro Romano, and Mario Varlese. A chit-chat between Llama 2 and ChatGPT for the automated creation of exploits. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004682.

Chen:2025:JCS

[3067] Yishan Chen, Xiangwei Zeng, Xiansong Luo, and Zhiquan Liu. Joint client–server selection and resource allocation based on split federated learning in edge-to-cloud computing environments. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004694.

Dong:2025:SLI

[3068] Wenlong Dong, Xiujun Wang, and Juyan Li. A secure lightweight identity authentication and key agreement scheme for Internet of Drones. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004700.

Song:2025:HAK

[3069] Yubo Song, Yuxin Lin, Hongyu Zhu, Han Zhao, Qingyue Yu, Yile Zhang, Dali Deng, and Zhefu Deng. Hybrid authenticated key agreement utilizing CSI for enhanced wireless security. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004712.

Zheng:2025:RDS

[3070] Guoli Zheng, Li Cao, Honglei Men, Yuanshuai Li, and Liang Chen. Research on data security delivery algorithm for crowdsourcing scenario in Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004724.

Kallitsis:2025:OFJ

[3071] Georgios Kallitsis, Vasileios Karyotis, and Symeon Papavassiliou. An optimization framework for joint wireless data-power transmission in distributed energy harvesting networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004736.

Liu:2025:DLB

[3072] Linfeng Liu, Xiangyu Yan, and Jia Xu. Distributed learning based message dissemination approach for underwater surveillance in OUSN. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004748.

Badjie:2025:MMG

[3073] Ansu Badjie, Grace Mupoyi Ntuala, Qi Xia, Jianbin Gao, and Hu Xia. MG-GPT: a multi-graph GPT-enhanced framework for dynamic fraud detection in cryptocurrency networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500475X.

Li:2025:DEI

[3074] Lanyining Li, Jiaxin Zhang, Hao Yuan, Jianghua Long, and Xing Zhang. DQN-enhanced interference mitigation for NGSO satellites in NTN based on spectrum sensing. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004761.

Jain:2025:SEH

[3075] Lilima Jain, Venkanna U., and Satyanarayana Vollala. SecTopo: Efficient hybrid model for detecting LLDP topology poisoning attack in programmable data plane. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004773.

Cai:2025:AAE

[3076] Saihua Cai, Gang Wang, Jinfu Chen, Shengran Wang, and Kun Wang. APT-ATT: an efficient APT attribution model based on heterogeneous threat intelligence representation and CTGAN. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004785.

Ghasemi:2025:DRA

[3077] Omid Abachian Ghasemi, Masoumeh Azghani, and Mehdi Chehel Amirani. The design of an RIS-assisted FDMA wireless sensor network for sum throughput maximization. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004797.

Li:2025:CMT

[3078] Yujie Li, Yiwei Liu, Peiyue Li, Yifan Jia, and Yanbin Wang. Continuous multi-task pre-training for malicious URL detection and webpage classification. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004803.

Zhang:2025:TPA

[3079] Xinwei Zhang, Chengzhe Lai, Guanjie Li, and Dong Zheng. Two-phase authentication for secure vehicular digital twin communications. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004815.

He:2025:MCF

[3080] Yongping He, Zihang Feng, Tijin Yan, Yufeng Zhan, and Yuanqing Xia. Meta-CAD: Few-shot anomaly detection for online social networks with meta-learning. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004827.

Guo:2025:GDM

[3081] Liang Guo, Chen-Khong Tham, Jie Jia, Jian Chen, and Xingwei Wang. Generative diffusion model-based QMIX for joint task offloading and resource allocation in VEC systems. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004839.

Haidar:2025:FTT

[3082] Ghulam Haidar, Murad ul Islam Khan, Sikander Ali, and Wali Khan. Faulttolerant topological indices of graph joins with applications in multi-swarm drone networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004840.

Wu:2025:BBE

[3083] Guangfu Wu, Su Xu, Daojing He, and Sammy Chan. Blockchain-based efficient and secure cloud cross-domain data sharing with dynamic revocation by multiple authorities. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004852.

Peng:2025:EES

[3084] Xiang Peng, Hua Xu, Zisen Qi, Dan Wang, Yue Zhang, and Yiqiong Pang. Energy-efficient strategy generation for smart jammer in nonzero-sum games: a deep reinforcement learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004876.

Wang:2025:VSM

[3085] Ke Wang, Yongjun Li, Xiang Wang, Yuanhao Liu, Kai Zhang, Fenglei Zhang, Zhiqiang Ma, and Zhe Zhao. A VNF sharing method based on node selection probability using reinforcement learning in air-ground network. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004888.

Chiaramello:2025:HCD

[3086] Emma Chiaramello, Carla Fabiana Chiasserini, Francesco Malandrino, Alessandro Nordio, Marta Parazzini, and Alvaro Valcarce. Human-centric decision-making in cell-less 6G networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500489X.

Liu:2025:NAE

[3087] Lei Liu, Yong Wang, Yangfan Liang, Junqi Chen, and Qian He. Innetwork aggregation enabled multiple sub-blocks parallel repair in erasurecoded storage system. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004906.

Li:2025:TEB

[3088] Linghao Li, Yan Zhu, Yun Li, Wei Qiao, Zelin Cui, Susu Cui, Bo Jiang, and Zhigang Lu. Towards effective black-box attacks on DoH tunnel detection systems. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004918.

Zhang:2025:ERA

[3089] Xiaoliang Zhang, Miao Wang, Mengxiong Wang, Liqiang Wang, and Hong Zhang. Enhanced RIS-assisted vehicular network with TDMA and Bayesian Compressive Sensing-based channel estimation. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500492X.

Fu:2025:PPW

[3090] Yanming Fu, Linfeng Peng, Ruikang Guo, Xian Zhang, Lin Li, and Wei Wei. Privacy-preserving worker recruitment scheme with MAB and reverse auction in crowdsensing to improve truth discovery performance. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004931.

Roohi:2025:CAI

[3091] Koorosh Roohi and Atena Roshan Fekr. A comparative analysis of indoor localization technologies. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625004943.

Mo:2025:OUS

[3092] Kaiwei Mo, Xianglong Li, Zongpeng Li, and Hong Xu. Optimizing UAV scheduling and trajectory planning: an online auction framework. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004955.

Wu:2025:ISS

[3093] Wenhao Wu, Xinsheng Ji, Jie Yang, Wei You, Xinyu Huang, and Cancan Tao. An intelligent scheduling strategy for security isolation of 5G network slice. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004967.

Fesl:2025:CML

[3094] J. Fesl and M. Naas. A comprehensive machine learning-based approach for virtual private network traffic detection, classification and hiding. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004979.

Zhang:2025:SEC

[3095] Milin Zhang, Mohammad Abdi, Venkat R. Dasari, and Francesco Restuccia. Semantic edge computing and semantic communications in 6G networks: a unifying survey and research challenges. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004980.

Fan:2025:JOD

[3096] Xiaokun Fan, Yali Chen, Min Liu, Yuchen Zhu, and Zhongcheng Li. Joint optimization of data sensing and computing in the air–ground collaborative inference framework: a multi-agent hybrid-action DRL approach. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005079.

Fernandez-Rodriguez:2025:LSB

[3097] Jose David Fernández-Rodríguez, Iván García-Aguilar, Rafael Marcos Luque-Baena, Ezequiel López-Rubio, Marcos Baena-Molina, and Juan Francisco Valenzuela-Valdés. Learning to shape beams: Using a neural network to control a beamforming antenna. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005110.

Gong:2025:SEA

[3098] Mengyi Gong, Ziling Wei, Shuhui Chen, Wanrong Yu, and Fei Wang. A survey of existing attacks on 5G SA. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005122.

Wang:2025:TCOb

[3099] Jian Wang, Qi Zhang, Guanzhi He, and Guosheng Zhao. Task combination optimization via dual-view heterogeneous graph contrastive learning for mobile crowdsensing. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005134.

Huang:2025:PBA

[3100] Wenwei Huang, Tianyu Kang, Li Guo, and Luo Deng. Permissioned blockchain architecture enabling boundedtime PBFT consensus over deterministic networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005146.

Shah:2025:EER

[3101] Syed Luqman Shah, Ziaul Haq Abbas, Ghulam Abbas, and Nurul Huda Mahmood. Energy-efficient and reliable data collection in receiverinitiated wake-up radio enabled IoT networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005158.

Kostas:2025:IGM

[3102] Kahraman Kostas, Mike Just, and Michael A. Lones. IoTGeM: Generalizable models for behaviour-based IoT attack detection. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005171.

Wu:2025:EID

[3103] Lin Wu, Yu-Lai Xie, Shi-Xun Zhao, Pan Zhou, Dan Feng, Avani Wildani, and Ya-Feng Wu. Efficient intrusion detection via heterogeneous graph attention networks and parallel provenance analysis. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005195.

Rasheed:2025:MAP

[3104] Iftikhar Rasheed and Hala Mostafa. Mobility-aware predictive split federated learning for 6G vehicular networks with ultra-low latency guarantees. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005201.

He:2025:CCO

[3105] Yuan He, Xie Jun, Yaqun Liu, and Xijian Luo. Collaborative computation offloading and trajectory planning in locally observable multi-UAV MEC networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005213.

Wang:2025:NBI

[3106] Tong Wang, Xiyao Liu, Zhixiang Wang, and Panfeng Huang. Networkbased integrated path planning for UAVs monitoring of dispersed targets. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005237.

Lv:2025:BEB

[3107] Quanfeng Lv, Yuan Chang, Tong Li, and Jingguo Ge. Betastack: Enhancing base station traffic prediction with network-specific Large Language Models. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005249.

Toka:2025:RLS

[3108] Laszlo Toka, Zoltan Illes, Endre Angelus Papp, Laszlo Hevizi, and Istvan Godor. Routing LEO satellite traffic in adverse weather. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005262.

Kagai:2025:PEL

[3109] Francis Kagai, Philip Branch, Jason But, Rebecca Allen, and Mark Rice. Performance evaluation of lowbitrate voice using spread spectrum techniques for satellite-based emergency communication. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005274.

Xiong:2025:ADA

[3110] Bing Xiong, Yu Chang, Yuhang Zhang, Jin Zhang, Baokang Zhao, and Keqin Li. AF-Detector: an accurate low-overhead method for detecting active flows in network traffic. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005298.

Zeng:2025:SRA

[3111] Yaoping Zeng, Shisen Chen, and Yimeng Ge. STAR-RIS-assisted UAVenabled MEC network: Minimizing long-term latency and system stability optimization. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005304.

Asif:2025:RIR

[3112] Muhammad Asif, Talha Ahmed Khan, and Wang-Cheol Song. R-IBN: a reinforcement learning-based intent-driven framework for end-to-end service orchestration and optimization. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005316.

Yang:2025:AAP

[3113] Yongqi Yang, Qianyi Huang, Yixue Liu, Xiaojie Huang, Liang Zhang, Jiaxin Tian, Li Wang, Chen Tian, Wanchun Dou, and Guihai Chen. APCC: Active precise congestion control for campus wireless networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005328.

Troia:2025:BNT

[3114] Sebastian Troia, Jean-Pierre Asdikian, Giacomo Sguotti, Enrico Gregorini, Mengyao Li, and Guido Maier. Inband network telemetry for softwaredefined wide area networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005341.

Chelghoum:2025:SEV

[3115] Massinissa Chelghoum, Gueltoum Bendiab, Mohamed Benmohammed, Mohamed-Aymen Labiod, Samia Bousalem, and Abdelhamid Mellouk. Securing 6g-enabled vehicleto-everything communications: a blockchain-enabled collaborative intrusion detection framework with reinforcement learning. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005353.

Uomo:2025:PPF

[3116] Domenico Uomo, Andrea Sgambelluri, Layal Ismail, and Francesco Paolucci. P4 programmability for fault-tolerant software defined flying ad-hoc network. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005390.

Leblebici:2025:GAI

[3117] Merih Leblebici and Ali Çalhan. Generative artificial intelligence for integrated sensing and communication in 6G. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005407.

Khan:2025:NTB

[3118] Waqar Khan, Jawad Ahmad, Nada Alasbali, Alanoud Al Mazroa, Mohammed S. Alshehri, and Muhammad Shahbaz Khan. A novel transformer-based explainable AI approach using SHAP for intrusion detection in vehicular ad hoc networks. Computer Networks (Amsterdam, Netherlands: 1999), 270 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005420.

Kilincer:2025:MSR

[3119] Ilhan Firat Kilincer. A modular system for real-time intrusion detection on local area networks. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005444.

Anonymous:2025:EBo

[3120] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005894.

Anonymous:2025:Oa

[3121] Anonymous. October 2025. Computer Networks (Amsterdam, Netherlands: 1999), 270(??):??, October 2025.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Pintor:2025:STN

[3122] Lucia Pintor, Luigi Atzori, and Antonio Iera. Sustainability in telecommunication networks and Key Value Indicators: a survey. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004335.

Munson:2025:NNE

[3123] Daisy Munson, Pierre Alain, and Guillaume Doyen. NEAT: a Nile-English Aligned Translation corpus based on a robust methodology for intent based networking. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625004864.

Dai:2025:INM

[3124] Xianlong Dai, Guang Cheng, Li Deng, Ziyang Yu, Zihan Chen, Xuman Zhang, Bingjie Duan, and Yuheng Li. ITD: a novel measure for detecting portable hotspot devices in modern cyberspace. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500516X.

Pincu:2025:TIA

[3125] Bar Pincu, Aviram Zilberman, Ilia Leibovich, Rami Puzis, Andikan Otung, Motoyoshi Sekiya, and Yuval Elovici. Threat impact analysis of man-in-the-middle attacks on delay-based geolocation on the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005183.

Wang:2025:EED

[3126] Qing Wang, Zhijun Wu, and YanRong Lu. EDVFL: an evaluable and decentralized privacy-preserving VFL for secure data sharing in ATM. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005225.

Lv:2025:RAP

[3127] Yali Lv, Xiaoxi Zhang, Yingsheng Peng, Jingpu Duan, Xiong Li, Bo Yi, and Qing Li. Resource allocation and pricing for SFC deployment in Space–Air–Ground-Integrated Networks: an innovative auction-based strategy. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005250.

Biyar:2025:ACH

[3128] Elham Dehghan Biyar, Mirko D’Angelo, Josué Castañeda Cisneros, Amadeu Nascimento, Marin Orlic, Ankita Likhyani, András Zahemszky, Ahmet Cihat Baktir, Dagnachew Azene Temesgene, and Dinand Roeland. Autonomous conflict handling in intent-based management. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005286.

Lei:2025:SLA

[3129] Lei Lei and Xudong Wang. Scaling laws of average end-to-end delay in multitier wireless mesh networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500533X.

Ngo:2025:RTP

[3130] Duc-Thinh Ngo, Ons Aouedi, Kandaraj Piamrat, Thomas Hassan, and Philippe Raipin. Rethinking traffic prediction in Mobile Network Digital Twins: a flexible inductive graphbased learning model for data-scarce scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005365.

Li:2025:EMA

[3131] Yi Li. An enhanced mutual authentication scheme with perfect forward secrecy for UAV networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005377.

Montes-Lopez:2025:DTB

[3132] Juan Manuel Montes-Lopez, Pablo Serrano, Marco Gramaglia, and Albert Banchs. DiWi: a transformerbased Digital twin for Wireless mobility. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005389.

Luo:2025:RFF

[3133] MengHao Luo, Xu Liu, XiaoQiang Di, and Hui Qi. Radio frequency fingerprint recognition method based on feature enhancement and Selective Channel–Spatial Kernel. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005419.

Hou:2025:MCE

[3134] Saifeng Hou, Yuxiang Hu, Pengshuai Cui, and Jiangxing Wu. MP-Cos: an efficient storage compression and optimization scheme of flow table for multiprotocol network scenarios. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005432.

Chen:2025:DBB

[3135] Hongwei Chen, Wei Wang, and Yujie Duan. A dual blockchainbased privacy-preserving authentication scheme for Vehicular Ad Hoc Networks. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005456.

Wondmagegn:2025:PAF

[3136] Ayalneh Bitew Wondmagegn, Dongwook Won, Quang Tuan Do, Demeke Shumeye Lakew, and Sungrae Cho. Performance analysis of FSObased communications in space–air– ground integrated networks: a comprehensive survey. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005468.

Liu:2025:DLN

[3137] Yixin Liu and Qiang Zhi. A duallayer network security model for smart factories based on the sentinel mechanism. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500547X.

Ramadan:2025:AEA

[3138] Ahmed Adel Ramadan, Muhammad El-Saba, and Mahmoud A. Shawky. Adaptive and efficient authentication for VANETs: a cross-layer framework for resilient communication. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005481.

Ning:2025:STC

[3139] Nianwen Ning, Wenbo Zhang, Wei Li, Hengji Li, Yanyu Zhang, Yi Zhou, and Fuqiang Liu. Spatial–temporal causal fusion graph neural networks for urban traffic prediction. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005493.

Lin:2025:TEM

[3140] Hai Lin, Jiayi Zhao, Xingchen Zhu, and Liyuan Wang. Trust evaluation mechanism for the Internet of Vehicles based on filtering algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500550X.

Wang:2025:LPH

[3141] Teng Wang, Shuai Wang, Jiangguo Liang, and Tengfei Yang. Locally private high-dimensional data synthesis based on optimal attribute pairs. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005511.

Sun:2025:PHP

[3142] Hesheng Sun, Zhuzhong Qian, Andong Zhu, Sheng Zhang, Sanglu Lu, Chu Xu, and Gangyi Luo. Provisioning high precision edge inference with runtime model reconfiguration. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005523.

AlOmar:2025:DTN

[3143] Ban AlOmar, Zouheir Trabelsi, and Saed Alrabaee. Detection of Tor network obfuscated traffic using Bidirectional Generative Adversarial Network. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005535.

Taimoor:2025:ROM

[3144] Shamim Taimoor, Muhammad Naeem, Sherali Zeadally, and Waleed Ejaz. Resource optimization for minimizing latency and cost in UAV-assisted mobile edge computing (MEC) networks. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005547.

Li:2025:RRD

[3145] Pengfei Li, Yujing Liu, Jinshu Su, and Bo Yu. ROVReco: an ROV deployment recommendation approach with GNN based on routing betweenness. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005559.

Akkal:2025:BBP

[3146] Maroua Akkal, Sarra Cherbal, Boubakeur Annane, and Hicham Lakhlef. BTMH: a blockchain-powered trust management system for IoMT in healthcare. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005560.

Charfi:2025:WFW

[3147] Emna Charfi, Ahlem Saddoud, and Lamia Chaari Fourati. From Wi-Fi 7 to Wi-Fi 8: a survey of technological evolution, emerging applications, challenges, and future aspects. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005572.

Lu:2025:TNE

[3148] Juncheng Lu, Zikun Zhou, Hua Wu, and Guang Cheng. TorVIA: a novel encrypted video identification method based on Tor transmission characteristics. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005584.

Li:2025:DDG

[3149] Jiayi Li, Xiaogang Wang, Haokun Chen, Zexin Wu, Ziqi Zhu, Jian Cao, and Rajkumar Buyya. DG-PAS: DQN-GRU guided distributed DNN pipeline training and adjacent scheduling in edge networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005596.

Tang:2025:FDM

[3150] Chenxi Tang, Jingwen Lu, Wen Yang, Wenjuan Xing, Chaowei Tang, Aobo Zou, and Jiayuan Guo. A frequencydomain multilayer perceptron with multiscale attention network for cellular traffic prediction. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005602.

Tariq:2025:SPB

[3151] Saman Tariq, Eva Rodriguez Luna, Xavier Masip Bruin, Rodrigo Diaz, Josep Martrat, and Panagiotis Trakadas. A survey on 5G private and B5G network threats and safeguarding AIbased security mechanisms through the layered analysis. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005614.

Xu:2025:AXV

[3152] Ke Xu, Haitao He, Yuxuan Zhao, Yufeng Jia, Peng Shi, and Bing Zhang. An adaptive XSS vulnerability detection method based on hierarchical multi-objective reward-enhanced Dueling Double Deep Q-Network. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005626.

Zhang:2025:TGA

[3153] Feng Zhang, Chenhao Luo, Winston K. G. Seah, Gang Xu, and Kailiang Zhao. Temporal graph attention and contrastive learning model for link prediction in dynamic networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005638.

Qian:2025:ROS

[3154] Pengcheng Qian, Liang Wang, Jiamin Fan, Zhenzheng Shi, and Mengge Li. Robust offloading strategy in VEC with computation uncertainty and imperfect CSI. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500564X.

Lamprou:2025:TFL

[3155] Vangelis Lamprou, George Doukas, Christos Ntanos, and Dimitris Askounis. On the trustworthiness of federated learning models for 5G network intrusion detection under heterogeneous data. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005651.

Hoque:2025:RCM

[3156] Md Muzammal Hoque, Ijaz Ahmad, Jani Suomalainen, Paolo Dini, and Mohammad Tahir. On resource consumption of machine learning in communications network security. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005675.

Baykara:2025:ZZT

[3157] Cem Ata Baykara, Ilgın Şafak, and Kubra Kalkan. ZETROS: a zero-trust IoT network security framework using distributed blacklisting, trust scoring and smart contracts. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005687.

Huang:2025:SSV

[3158] Zhongxin Huang, Zanbo Sun, Wei Tan, Houbing Herbert Song, Tian Wang, Shaobo Zhang, and Anfeng Liu. SPDTM: a Shapley-value based preference discovery and task matching to maximize satisfaction in mobile crowd sensing. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005699.

Kim:2025:IAA

[3159] Sungwook Kim. ISAC-aided adaptive spectrum sharing scheme based on new cooperative game solutions. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005705.

Hidouri:2025:DQI

[3160] Abdelhak Hidouri, Haifa Touati, Mohamed Hadded, Mohamed Amin Asri, Nasreddine Hajlaoui, Paul Muhlethaler, and Samia Bouzefrane. Deep Q-ICAN: a deep reinforcement learning-based approach for real-time CPA attack detection and mitigation in NDN architecture. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005717.

Chu:2025:EMT

[3161] Chu Chu, Sun Mao, Jinsong Wu, Gang Li, Zhenbing Li, and Guangjun Wen. Efficient missing tag identification for large-scale RFID systems via collision exploitation. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005729.

Qin:2025:EID

[3162] Zhenkai Qin, Lingying Li, Qining Luo, Xiaochuan Yu, and Qian Zhang. Enhancing intrusion detection performance using GCN-LOF: a hybrid graph-based anomaly detection approach. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005730.

Tang:2025:MSD

[3163] Jianzhi Tang, Luoyi Fu, Yu-E Sun, Xinbing Wang, and He Huang. Maintaining source–destination connectivity in uncertain networks under adversarial attack. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005742.

Li:2025:IMM

[3164] Bo Li and Hao Wang. Improving mobility management in LEO satellite networks utilizing soft actor-critic algorithms. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005754.

Darem:2025:IBR

[3165] Abdulbasit A. Darem, Asma A. Alhashmi, Abed Saif Ahmed Alghawli, Amal Alshardan, Mohammed Burhanur Rehman, and Mukhtar Ghaleb. Integrated BD-RIS and hybrid NOMA/ OMA optimization for secure, lowlatency vehicular networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005766.

Hu:2025:PAO

[3166] Xiaoxuan Hu, Xianyu Wei, Liang Shan, Jin Qi, Zhenjiang Dong, and Yanfei Sun. Power allocation optimization for hybrid IRS-assisted 6G V2V communication. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005778.

Yao:2025:OEH

[3167] Jie Yao, Li Tian, Ziyi Wei, and Guozi Sun. Overcoming emergency HTTP/3 DDoS attack detection: a domain adaptation solution with graph neural network. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500578X.

Tefera:2025:JBM

[3168] Mulugeta K. Tefera, Guilu Wu, and Feng Shu. Joint beamforming of MISO secure transmission with distributed intelligent reflecting surfaces. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005808.

Liu:2025:LLM

[3169] Fan Liu, Behrooz Farkiani, and Patrick Crowley. Large Language Models for computer networking operations and management: a survey on applications, key techniques, and opportunities. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S138912862500581X.

Alshami:2025:SSC

[3170] Faisal Alshami, Lin Yao, Xin Wang, and Guowei Wu. SCDFL: a spectral clustering-based framework for accelerating convergence in decentralized federated learning. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005821.

Huo:2025:SIW

[3171] Qitao Huo and Peng Zhou. Source identification for worm propagation: a graph neural network approach and evaluation on social network and internet datasets. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005833.

Xu:2025:NOD

[3172] Lixia Xu and Jie Yin. A novel optimization and deep learning enabled resource allocation algorithm using blockchainbased edge computing in IoT. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005857.

Cheng:2025:SLF

[3173] Wenhao Cheng, Xia Feng, Liangmin Wang, Zhan Xie, Liang Wang, and Siben Tian. SMUSAC: Lightweight federated learning framework for SUNETs with tolerance of data loss and node compromise. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005936.

Ledesma:2025:HIN

[3174] Oscar Ledesma and Paula Lamo. Hybrid IoT network for real-time monitoring of maritime containers. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625005948.

Sun:2025:PPD

[3175] Jin Sun, Mengna Kang, and Zhongpeng Sun. Privacy-preserving data sharing scheme based on federated learning in Internet of Vehicles. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500595X.

Akgul:2025:SAO

[3176] Özgür Umut Akgül, Antonio Varvara, Antonio de la Oliva, Panagiotis Charatsaris, Maria Diamanti, Pere Garau Burguera, Mårten Ericson, Stefan Wänstedt, Marcin Ziółkowski, Halina Tarasiuk, Hamed Hellaoui, Symeon Papavassiliou, Vasileios Tsekenis, Sokratis Barmpounakis, Panagiotis Demestichas, Bahare M. Khorsandi, and Hasanin Harkous. Sustainable 6G architecture: an organic evolution of 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005961.

Zhang:2025:STS

[3177] Yi Zhang, Zhenyu Xu, Xurui Ren, Hua Wu, and Guang Cheng. SCEP-TI: a side-channel attack on encrypted proxy video streams for video title identification. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625005973.

Li:2025:GNB

[3178] Shangsen Li, Lailong Luo, Changhao Qiu, Bangbang Ren, Yun Zhou, Deke Guo, and Richard T. B. Ma. GraphVeri: a NAR-based control plane verification framework for routing protocols. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625006024.

Zeng:2025:EEE

[3179] Fanyi Zeng, Dapeng Man, Yuhao Zhao, Yuchen Liu, Huanran Wang, and Wu Yang. EMTD: Efficient encrypted malware traffic detection based on adaptive meta-path guided graph propagation. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625006036.

Huy:2025:SSF

[3180] Trinh Gia Huy, Luong Nguyen Thanh Nhan, and Nguyen Tan Cam. secPEFL: Strengthening federated learning security for Portable Executable malware detection in distributed networks. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S138912862500605X.

Duan:2025:MCR

[3181] Yunge Duan, Zhenbo Liu, and Shuang Fu. Maximizing computation rate for NOMA-based WPT-MEC with user cooperation under nonlinear EH model. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625006061.

Khater:2025:OEC

[3182] Heba M. Khater, Farag Sallabi, Abdulmalik Alwarafy, Ezedin Barka, Mohamed Adel Serhani, Khaled Shuaib, and Mohamad Khayat. Optimized edge-cloud task offloading for WBANs: a hierarchical deep-reinforcement-learning approach. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625006073.

Zeng:2025:RAH

[3183] Yue Zeng, Pan Li, Shanshan Lin, Bin Tang, Xiaoliang Wang, Zhihao Qu, Baoliu Ye, Song Guo, and Junlong Zhou. Reliability-aware hybrid SFC backup and deployment in edge computing. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625006085.

Mezni:2025:FRP

[3184] Haithem Mezni, Mokhtar Sellami, Hela Elmannai, and Reem Alkanhel. Federated resource prediction in UAV networks for efficient composition of drone delivery services. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625006097.

Chen:2025:DLB

[3185] Peng Chen, Youtong Li, Hao Wu, and Jixian Zhang. A deep learningbased reverse auction mechanism for semantic communication in IoV crowdsensing services. Computer Networks (Amsterdam, Netherlands: 1999), 271 (??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625006103.

Silva:2025:RTA

[3186] Rui Silva, Filipe Antão, David Santos, André Perdigão, Tiago Barros, Fatma Marzouk, Alisson Chaves, Daniel Corujo, and Rui L. Aguiar. Real-time adaptive resource management for high-resolution computer vision over private 5G networks. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625006115.

Yang:2025:AOE

[3187] Pengfei Yang, Tianyang Zheng, Shuyu Zhang, Weidi Su, Bijie Yi, Wenkai Lv, and Quan Wang. Alternating optimization for energy consumption-oriented task offloading in SAGIN. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1389128625006139.

Jiang:2025:DPD

[3188] Haifeng Jiang, Yu Chen, Yingjie Zhu, Xin Xu, Yuejia Song, and Qiwei Chen. D-PhishNet: a dual-branch network for URL and HTML feature fusion in phishing webpage detection. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625006152.

Anonymous:2025:EBp

[3189] Anonymous. Editorial Board. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1389128625006528.

Anonymous:2025:Ob

[3190] Anonymous. October 2025. Computer Networks (Amsterdam, Netherlands: 1999), 271(??):??, October 2025. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).